A Bibliography of Publications about the Fortran Programming Language: Part 1: 1956–1980

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

19 March 2024 Version 2.86

Title word cross-reference

#6 [HJ74]. #8 [Hal75].

(a, a′) [SR73, SR73]. (a, b) [SMD71, SR73]. (a, bγ) [SR73]. (a, bγ − γ) [SR73]. (a, bσ) [SMD71]. (a, bσ − σ) [SMD71]. (a, γ) [SR73]. (a, γ − γ) [SR73]. + [AI79]. 0 [Fia73, MT78]. 1 [Fia73, MT78, RCL75]. 1/2 [SDH74]. 24 [Ein76, JR75]. 28 [JR76]. 37 [MT78]. 39 [Spä79a]. = [AI79]. _{N}C_{R} [Gen75c, Gen75a, Gen75b]. Ax = b [Bar79b]. χ^{2} [BR75b]. DO [Gen75a]. F (X) [Gaf80]. F_{17} [Gos80]. g [ES74b]. H^{p} [Sin73]. I_{n}(x + jy) [Sca71]. J_{n}(x) [Col80a]. J_{n}(x + jy) [Sca71]. k [Gar63]. K_{n}(x + iy) [Gar78]. L_{1} [Spä76, Abd80, RR73b]. ≤ 80 [CM66]. ⟶ [Pit79]. m [And73, LT73]. M × N [Sch72c]. N [Cam65, Ken65, LT73]. SU_{3} [AD73]. U [DB73]. we^{w} = x [FSC73]. w exp(w) = x [Ein74]. Y_{2n} [BH1 ]. Y_{i} [Gar63]. Y_{n}(x + iy) [Gar78].

*FTN [CW72, CW73b, CW76, CW77, CW78a, CW79]. *WATFIV [CW72, CW73b, CW76, CW77, CW78a, CW79].

-ary [Ken65]. -dimensions [Gar63]. -layer [And73]. -Splines [ES74b].

/6000 [Hon75a, Hon75b].

00 [Ent80a]. 01 [BPW72]. 051 [Int67a]. 052 [Int67a].

1 [Bro75, BDI72, DT74, ES78, Har69, Lea67, Lea75, Les72, LS76, Mac70b, Mac71, MW69, Per72a, Per72b, Rus78, Wei73]. 1- [Rin77]. 1/0 [Joh72]. 10 [Ano77a, Dig68, Dig70, Dig72a, Dig74, Dig75b, Edg79, Gra70b]. 10/ [Ste80]. 10/LSD [Les72]. 1011-tutorials [Mis78a]. 1022 [AE79]. 11 [CDGW76, Dig71a, Dig72c, Dig75d, Dig75f, Dig76b, Dig77c, Dig77d, Dig78b, Dig78c, Dig79b, Dig80c, Dig80d, Dig80e, Dig80f, Dig80g]. 11/ [San78]. 11/FORTRAN [Dig76c]. 11/RSTS/E [Dig75g, Dig77f]. 1107 [Abr65, Cha66]. 1108 [Hog70]. 112 [Bak77b]. 113 [BB77b]. 1130 [CF71, Hug69]. 1130/1800 [Int68a, Int69b, Int71e, Int73a, Pri69]. 1130/system [CF71]. 119 [Kri71]. 11R1 [Spe84]. 12 [Man71]. 12-statement [MHH71a, MHH71b, Man71]. 120 [RST78]. 130 [Gen78]. 131 [Lea78]. 132 [AK78]. 133 [Žil78]. 1401 [Hai65, Int67a, Int64b]. 1401-F0-051 [Int67a]. 1401-F0-052 [Int67a]. 1410 [Fed63]. 149 [Ste60a]. 15 [Dig71b, Dig71c]. 16/32K [Int65f, Int65d]. 1620 [Fas67, Gar65, Hor65, PD67, Poo62, SD67, Swi64b, Swi64c, SS66]. 170/6600 [BCE77]. 18 [Hil70]. 1900 [NY78, REC75]. 1966 [Ame66a, Ame78d]. 1969 [Ano69a]. 1971 [FGR72]. 1972 [Cha73]. 1975 [FJA80a, FJA80b]. 1978 [ACM78, Ame78e, Ano78d, FJA80a, FJA80b, U.S78]. 1979 [FJA80a, FJA80b, Lew79b, VP80a, VP80b]. 1979-1980 [Ano79]. 1V [Lec66a, McC65b, SM66b].

2 [Ano68c, Fri71a, MS74a, MV66, Sak79, Thr79]. 2.50 [Bar73a, Pat73b]. 2.O [Con71c]. 22 [Swa72]. 2A [KGY80]. 2d [MS75d].

3 [Bar72a, Fri71a, Spe80d]. 3-dimensional [Tur68]. 300 [Vic64]. 3100-3200-3300-3500 [Con68b]. 3100/3150/3170/3200/3300/3500 [Con73c]. 3100/3200/3300/3500 [Con68a]. 3170/3300/3500 [Con71b, Con73b]. 3200 [SH67]. 32K [Int65f, Int65d]. 3300/3500 [Con69a]. 34 [Fre74]. 3400/3600 [Con65]. 3400/3600/3800 [Con67b]. 343 [Nie72c]. 360 [CCHT67a, CCHT67b, Cod67, Int65c, Int65e, Int67b, Int68b, Int69c, Int70b, Int70c, Int71h, Int72j, Int73b, Int74g, Int75c, Int75f, Int76, Int77a, Int79, Kuk66, Kuk67, Lee72, Sun73]. 360/370 [Boy75, Fel75b, TC75]. 360/40 [ADT67]. 360/50 [Bri68b]. 360/65 [Ben69, Lou74]. 360/CDC [Shu75]. 3600 [Sta65, VR67]. 360F [Int68k, Int75f]. 360F-LM-619 [Int68k, Int75f]. 360N [Int68k, Int75f]. 360N-LM-480 [Int68k, Int75f]. 360S [Int68k, Int68b, Int70c, Int72k, Int75f]. 360S-F0-500 [Int72k]. 360S-F0-520 [Int72k]. 360S-FO-500 [Int68b, Int70c]. 360S-FO-520 [Int70c]. 360S-LM-501 [Int68k, Int75f]. 370 [FF75, Int72j, Int74g, Int76, Int77a, Int79, Lee72, Sun73, Uni79]. 370/155 [Kra74]. 370/360 [Kuo73, Kuo74, Sun73]. 3A [Con75a, Ste80]. 3A/B [Con75a]. 3D [Rog80]. 3j [CM66]. 3rd [IEE75].

4 [KC72, McC78a, Ste76a, Gar72]. 4-1-4-1 [Int59e]. 400 [Gen66b]. 4000 [Sta69]. 410 [Cha71a]. 418 [Ein72]. 419 [JT72]. 420 [Wil72d, Wil72e]. 421 [Kuk72a]. 422 [Whi72]. 423 [Mol72c]. 424 [Gen72]. 425 [HK72, Ste70]. 427 [Lin72]. 428 [Yoh72]. 43 [Yat71a, Yat71b]. 44 [Hol77]. 441 [Kno73]. 443 [Ein74, FSC73]. 448 [BS73b]. 449 [Fia73]. 450 [MM73a, MM73b]. 452 [LT73]. 453 [Pie73]. 454 [RK73]. 456 [Fen73]. 458 [RR73b]. 46 [Lil71]. 460 [SS73]. 461 [BH73b]. 462 [Don73a]. 463 [Lew73]. 468 [Pat73a]. 470 [Kub73]. 473 [Pie74]. 474 [Aki74]. 476 [Cli74a]. 479 [Pag74a]. 480 [Int68k, Int75f]. 485 [ES74b]. 487 [Pom74]. 488 [Bre74]. 4th [IEE78]. 4X [Hew80a].

5-9 [Ano73, Ano75c, Ano75d, Ano75e, Xer71a, Xer73]. 5.95 [Bar80b]. 5/6/7 [Ano70a, Ano70f, Ano70g]. 5/7 [Ano70b, Xer70a, Xer70b]. 50 [BB71, Tho72a]. 500 [Int68b, Int70c, Int72k]. 501 [Int68k, Int75f]. 51 [Hab72]. 52 [Spi72]. 520 [Int70c, Int72k]. 524 [Bre78a]. 550/560 [Ano75c, Ano75d]. 57 [Hab73]. 5702 [Int74f]. 5702-F01 [Int74f]. 5703 [Int74f]. 5703-F01 [Int74f]. 5704 [Int74f]. 5704-F01 [Int74f]. 5734 [Int74c, Int74d]. 5734-F03 [Int74c, Int74d]. 5734-LM3 [Int74c, Int74d]. 58 [Spa73]. 59 [Fre73].

6.00 [Ree73]. 60 [Gri78, ST73c]. 600 [Gen69, Sch68]. 600/6000 [Hon73d, Hon73c]. 6000 [Bur78, Lin71]. 6000/7000 [FL74]. 6000/7000/Cyber [FLM70, Fri73, FLM74, LM70]. 619 [Int68k, Int75f]. 62 [DB73]. 62/reference [Con62c, Con62d]. 63/reference [Con64d]. 635 [GS71]. 6400/6500/6600 [Con67c, Con69d, Con69b, WM9 ]. 65 [Håv71]. 65/6600 [Håv71]. 6600 [BH73a, Håv71]. 6700/B [Bur74]. 68 [HH79a, PH77]. 6j [CM66].

7 [Mor79]. 7.00 [Ree75]. 7000/B [Bur78]. 704 [IBM56, Int57b, Int57a, Int57c, Int57d, IBM58, Int58, Int59d]. 7040 [Bar66, HW67, Int63o, Int65f, Int65d]. 7040/7044 [Int63o, Int65f, Int65d]. 7044 [Int63o, Int65f, Int65d]. 7070-series [Int62b]. 709 [PG65, Saw62]. 709/7090 [Int60c, Int61c, Int61e, Int61f, Int62d, Moo60]. 7090 [BS64, Dah66, Gen66a, PB67, Smi63a]. 7090-4 [Jon64]. 7090/7094 [BC67, Int63p, Int63f, Int63e, Int63g, Int63h, Int64g, Int64c, Int64d, Int66a, Dah66, PB67]. 7094 [Abr65, Dah66, PB67]. 7094/1401 [Fox64, MI64]. 72 [Hyd66a, Hyd66b]. 72-J2-L205 [Hyd66b]. 72-J2-L208 [Hyd66a]. 73-17 [Lou73]. 74 [Fre74]. 74-34 [Fre74]. 76 [SIG76]. ’77 [Ste80, Ano77a, BM79b, BGG78, FW78, HH79b, Kat78b, MO80, Rei80, Sha77, VP80b, Wag80a, Zwa80, VP80a]. 77/FORTRAN [Gen77b]. 78-06-01 [Ame78f]. 7R1 [Ano78a].

8 [Ano80c, Dig64, LC75, Uni68a]. 80

[Mos78, Rad79, Rad88, Sch79b]. 82 [Gow75]. 83 [Mon75]. 84 [AI79, MZ75]. 85/86 [Sys73a, Sys73b]. 87 [MH75a]. 88 [Gen75c]. 89 [BR75a]. 8K [Int63o].

90 [Boh75]. 91 [BR75b]. 92 [BS75]. 9j [CM66]. 9R1 [Spe79, Spe80c].

= [LT75].

A. [Hil73, Hil79a, Kar77]. A.P.I [CS76]. A1 [BS73b, Bre78a, Bre79b]. AAEARITH [Bac72]. AAEC [Feh68, JS74]. AAEC-FMO [Feh68]. abbreviated [Cor61]. ABC [Mer79]. ability [Cla73a, New75]. Abkuerzungen [Amk73]. Absolute [AK78, Van73a]. absorbed [Won67]. absorption [Gus73]. abstract [ES75, SW75, SS75b]. Abstraction [Bur79, DM66c]. académique [Lou73]. Academy [Uni78]. Accelerated [KGY80]. access [IA78, Lit74]. Accessing [Ker80]. accompany [Col78b, Gro73c, HD78a, JSW77b, Moc71b, Mur80, MS75d, PD80b, Per72a, Per72b, SM72b, SS74, Wu73c]. according [Hun74]. accounting [Per72a, Per72b, Pil70, Per72a, Per72b]. accrescimento [SS68b]. accumulation [Mal70]. accuracy [Bri68a, OK72]. achsensymmetrischen [Rot71]. Acid [Nor66, Leo74]. acid-base [Leo74]. ACM [Aki74, BS73b, Bre74, BH73b, Cha71a, Cli74a, Don73a, ES74b, Ein72, Fen73, Fia73, FSC73, Gen72, Gri78, HK72, JT72, Jon79, KRS78, Kno73, Kub73, Kuk72a, Lew73, Lin72, LT73, MM73a, Mol72c, Pag74a, Pat73a, Pie73, Pie74, Pom74, RK73, RR73b, Ros78, SS73, Whi72, Wil72d, Wil72b, Yoh72]. ACM/SIGGRAPH [KRS78]. acquire [Smi79]. action [EB80, HBE80, ZD78]. activities [Lee77]. activity [O’D74]. Adaptable [Poo74]. adaptation [MW71a, MW71b]. Adaptive [KGY80, DP76a, Vic70a]. Addenda [Mer58a]. Addendum [Dat77a]. adding [CL70]. addition [Car66]. Additional [LS76, Lit74, Ver59]. Additions [Mas60, Ano67]. Additive [RST78, Krá72a]. Addressing [Gar63]. Adiabatic [RCL75]. adjustment [Bom67, Hor65]. administration [Int64f, Int69a, Int72i, MGL73, SM70]. Advanced [BE74, Bro74, Ent80a, HPR77, HPR78, Mis78a, êR76]. advising [Gil76]. advisor [Int8 ]. aeroelastic [PG67]. Aerotriangulation [Hor65]. AESYNTAX [Bar72b]. AFID [Hig79b]. AFIT [PH63]. After [Nee75]. Age [All66a, Cha66, VR67]. Age-Length [All66a]. agricultural [PMBK80]. agropcuaria [Oli71]. Aid [Gom79, Con79c, CS71e, CS77c, Cra75, JS74, Kaz78, PP77, Rad70, RS80, Wil76c]. aiding [KG76]. aids [Kau65, Uni77]. air [Kot72, ST73a, San74, SYR77, Van66]. air-cycle [Van66]. airborne [Har73]. aircraft [MI75a]. airfoil [Car77]. airplane [U.S61]. AIRPRO [SYR77]. al [Cou75, Ste72a]. Alfons [Han67]. Algebra [Bac78a, BF79, GIB65, Her72a, Law78, LHKK79a, LHKK79b, UK74, Ver65, Law77, dMdF73, Mat89]. algebraic [Car66, Dif72, Gra70b, Lee74a, Pow68, Pow70a, Sti62, Zim69, ZT76]. Algebras [BK72]. Algebrici [Amb65]. ALGOL [Din72, Jun69, KKU78, KTZ67, SS72, Baj77, AB69a, Ano69a, Blo71, Car78b, Cha70, Dig72a, DS62, Gri78, Hil69, Hil71, Kor77, KTZ68, KTZ71, Mat72b, PC71a, PC71b, Pul64, Smi78, Vei66, Vow74, Vow77, Vow78, Wil67, Baj72, Tec72, Fre74, Han67, HH79a, Per78, PH77, Sco77a, SMD71, SZ80, SF72].

ALGOL- [Jun69, KTZ67].

Algol-Programmierung [Tec72].

ALGOL68 [Hed77]. Algols

[FJA80b, Mee78d, Mee78a, Mee78b, Mee78c, Mee79, FJA80a]. Algorithm [Abd80, AB69a, Aki74, Bak77a, Ban78b, Ban78c, Ban78a, Ber76, BS73b, Bre74, BW75, BH73b, Cha71a, Cli74a, Day76, Dey76, Don73a, Dur80, ES74b, Ein72, Ein76, Fen73, Fia73, FHS78, FSC73, Gar78, Gen72, HK72, JR75, JR76, JT72, Kah80b, Kno73, Kub73, Kuk72a, LHKK79a, Lew73, Lin72, LT73, MM73a, MT78, McC71, McC78d, McL78, Mol72c, MC80a, MC80b, Osy76, Pag74a, Pat73a, Pie73, Pie74, Pom74, RK73, RR73b, RST78, Sal71a, SS73, She78e, She78b, Sim76b, Sim76a, Spä79a, Ste76b, Ste76c, SS79a, SS79b, Van73a, Whi72, Wil72d, Wil72b, Yoh72, Zak77, AGG61, BK77, CGH75, FS76, Hae77, Ham79a, Hor68, Kli73, Lep76, Les73, Mal70, Mue75, Rod76, SS78a, Sco78, Wil77b, Wil77c, Wil77d, AK78, Bak77b, BB77b, BB71, BR75a, BR75b, Boh75]. Algorithm [Bre78a, Bre79b, BS75, DB73, Ein74, Fre73, Fut78, Gen75c, Gen78, Gow75, Hab72, Hab73, Lea78, Lil71, MM73b, MZ75, MT78, MH75a, Mon75, Nie72c, RST78, Sim76b, Spa73, ST73c, Spi72, Ste79, Tho72a, Wil72e, Yat71a, Yat71b, Žil78]. Algorithmen [Spa75c, Spa75a]. Algorithmic [Mei68, Rab62]. Algorithms [Ano69a, AK78, Bak77b, BB77b, BJ74, BR75a, BR75b, Boh75, BS75, DB73, Fre73, Gav76, Gen75c, Gen78, Gow75, Hab72, Hab73, Lea78, Lew73, Mal72, MZ75, MH75a, Mon75, RST78, Sch66c, Sch66d, Sch66e, Sch66a, Sch66b, Spa73, ST73c, Spe80a, Spi72, Tho72a, Yat71a, Yat71b, Žil78, CJ77, Spa80]. Algoritm [AE79]. Algoritmicheskii [Kar76]. algorytmow [ATW77]. algunos [dMdF73]. alla [LMP77]. allocation [KW71, Sak64a, Sak64b, Sak65, Sak70]. alloys [BP76]. Almost [Kub73]. ALOMA [PT73]. ALP [Kan71]. alpha [SDH74]. Alphabetic [Yoh72]. also [Mur77a, Mur77b]. Alternating [SS73, Nav78]. Alternative [Kha77]. Alternatives [Fry71]. ALTRAN [Ano65, Bro71c, BH1 , BH72, Bro73b, FP73, HJ74, Hal74, Hal75, Fel75a, Hal71, RG76]. am [Rin77]. amendments [FRS77]. American [Ame78c, Ame78d, ANS78, Cad71, DH78, HD78a, U.S78, Per78]. AMPPL [FPB72, Fin72a]. AMPPL-II [FPB72, Fin72a]. AMTRAN [SWL68]. analise [Cad71]. analisi [Sic74]. Analisis [BB78]. analog [Tor69]. analyses [Bre79a, Fox67, MSNC61, Mel62, MSR66, Oer71]. analysing [KR69]. Analysis [AG80, Dah66, Dah78, Gin78a, Gre77, Gro73b, Hoa72, Hoa73, Hog70, Lou74, Mee74, Par75, Pen70, Rob67a, RT76, RT77, SFIK79, Spa73, SS66, Tip76, YS75, dL78, All67, Bai62, Bak68, BB72, BB77a, Bar79a, Bar72b, BC77, BC79, BY73, Bol76, BC67, Bro71b, Bry75, Car77, Cha79a, CK80, Cla73a, Clo72, Cri77, DS66, Dun75b, EF76, Edw76a, ESD68, Fin68, Fin72d, Fin72f, Fin72e, Fin77, GKB74, Gar63, Gen66a, GS70, Gro71, Har68a, HB63, Hem70, dPW80, Hil79b, IA80, Joh71, Kan79, Kan71, Khu68, Kra74, KRB78, Lar69, Lee69, Lou67, Lov68, Lyc80, Mac69, MI64, Mat72a, MC64, McC69b, McG76a, McG76b, Mit65, Mou70, NL75, Nor63, Nut76, Par70, PNK65a, PNK65b, PT67, PT69, PG66, PS78, Rau68, RPE79, Rey69, RR70]. analysis [Rin79, RW77, Kan68, She70a, SDZ80a, Sig80, Sik71, Spä76, Spa80, TC70, TB65, TD78, VP76, VP75, Wal68, WK77, Wil72a, Wis69, Wit79b, Wit79a, Wit79c, Wit79d, YHE69, vM75, vM78a, vM79, Kar77]. analytic [O’K64, Wat75]. Analytical [RG76, But66, Com80b, Kan79]. Analyze [Ant80, BP76, SG69, Whi68]. Analyzer [Fel76b, PJ75, Blu65, Jon64, Lyo74]. analyzing [Jul75, Sho76a]. Anatomy [Lee67a, Lee74b]. anaylsis [Ano70c, Ove72]. AND-OR [Nak77]. AND-OR-B [Nak77]. and/or [Car74b, SF72]. Anfanger [Con71a, Con73a]. Angeles [ACM78]. angewandten [Jun69]. angle [Maz77, WCT68]. angular [CM66, Ful74, RV78, SDH74]. Animal [Pat67, Whi71b]. animali [SS68b]. anisotropic [Zaa69]. ANL [Kuk66, Kuk67]. Anleitung [Dre70]. annular [KM77a, KM77b, Kat77]. anomalies [Plo75, Rap66c, Rap66d]. anomaly [KRB77]. ANS [Xer74a, Xer76a, Xer76b, ACM76, ANS76a, ANS76b, Ano76b, BS80a, Fel76a, FRS77, Hig79b, Kni76a, Kni76b, LHLM80a, LHLM80b, Lyo80, McI71, SIG76, ANS76c]. ANSI [Ame78d, Ano77c, Con73c, Ame66a, Ame66b, Ame78e, Ame78d, ABH^{+}71, Ano78d, Con71b, Con73b, Col78a, Col78b, Lam71a, Obr71, Sic74, SS76, Tan78a]. Answer [AB69b, Key73c, LJ71b, MHH71b]. answers [Haa69b]. Anthony [VP80a, VP80b]. Anwender [Her78]. Anwendung [Fri75b, Mac70a, Sto71]. any [Dav72b, Rus79, Ste72b]. AP.L. [Dav74]. aperiodic [Lov68]. APL [Cha70, CF71, Cor77, DT74, GS71, Lee74c, Mor71, Smi78]. APL/FORTRAN [Mor71]. aplicacion [Wei73]. aplicaciones [MD66a, MD73, Rid69]. aplicados [JSW70]. APLT [Ste60b]. Appendix [AS71, AW73b, HM62c]. Apple [Ano80a, Sil80]. Application [Int75a, SH67, BP76, BKW74, Boa69, Int78b, Int78c, Low76, RP74, SYR77, Weg66]. Applications [Day72b, HD78b, Kuo72, Lou73, May73a, McC67b, Mos64, Sas74a, See75, WB65, Ban75, Bar70, Car68, CW71, Day79, Gar71, HK75, Int64a, Ker72, LB68, LPJ79a, LPJ79b, Lud69, Mac70a, MD64, MD66b, MD68, Nol71, PS74, Sas74b, Sch69, SM70, Sch70, Sch71, Sho80, Vas72, Zim69, Hil73, Ung69, Wil75, Gow73, Bar73a].

Applicazioni [Rid67, Rid78]. Applied [Eng75, FJA80a, FJA80b, JSW67, JSW77a, JSW77b, Kar77, Lew80d, MS77b, SM72c, SM76b, Bri68a, Com80b, CK80, Clo72, Lew80c, Mat89]. Approach [CS73b, CS76, CS77b, DS77, Eng75, Har70, HM77, HH79b, Jay80, KC72, Lew80b, Mar80, OR75, Rul80, Swa72, Vic78a, Wol68b, AHP77, Ayc80, Bra79, Col78a, Col78b, CS68, CS71e, CS72, CS77c, DS72, Fis71, Gra79, Key73a, Key73c, Key73b, KS74, KS75b, Kro75, MH75b, MS70b, MS70c, MS75c, MS75d, Smi73a, SW79, Tea74, Vic70b, Vic78b, Zav73]. Approaches [Bar71b, BK72, And70]. approfondissement [Tho78]. Approved [AB66b, AB66e, Ame66c, Ame78c, Ame78d]. Approximation [McL78, RR73b, Dun79, Lil71, Rob79]. approximations [Lar63b]. April [Ame78c, Ame78d]. APSE [Nat70a]. arbitrarily [Dey76]. Arbitrary [Ono79a, Ono79b, Hol68, RS72, Sca71]. arc [MC70]. architectural [Rog80]. Architecture [LB80, PC78b]. architectures [BS78]. area [Gof74]. areas [Int78b, Int78c, Int80b]. argument [Ben78, Sca71]. argument-parameter [Ben78]. Arguments [Wan78, BH73a, FM76]. Arising [Mer78b, Wol78a, Mer78a, Mer78c]. Arithmetic [BY78, Bre75, Bre78a, Bre78b, Bre79b, BHY80, IEE75, IEE78, Mal72, NO75, Ono79a, Ono79b, RJAS78, She59, Ste77, WLO76, Yoh78, Yoh79a, Yoh79b, Yoh80, Bac72, Bre76a, Cla73a, Dat73, DD68, Kor79, Lin71, Per77, Smi71b, TG79b, Wal63, Wei65]. arithmetics [Cra79]. aritmetica [dMdF73]. arranging [Hil79c]. Array [Eri75, Flo78a, MR73, Ree79, Wet79, Wet80, Gar63]. arrays [Bai63, HV74, SS78a, SM73a, Tay76]. art [Kno72]. ary [Ken65]. ASA [Hei64, Hei66]. ASAP [MK73]. ASC [Mor79, Wed75]. ASC-7 [Mor79]. ASCII [Ano78a, Spe76a, Spe78b, Spe79, Spe80c, Spe84, Uni80c]. aspects [Ste74]. assemble [Ful73, Gro73d]. Assembler [Kuo73, PP77, Sun73, Ful73, Hug77, Joh72, Mos78, Har69, Rin77, Kuo74, LMP77]. ASSEMBLER-Programmen [Rin77]. Assembleur [Phe79]. assembly [Int62d, Int63g, Int64c, Mer60a, Moo60, Sti62, KF72]. assertion [SF75]. Assignment [BD80b, DZ78, Han72a, BS78, Han72b]. assignments [SDH74]. Assisted [DDM^{+}75, Hon73b, Mis78a, Pen68, Wid79, Bee75]. associated [Fis78, Gav76, Pol78]. association [Ben78, KPG63, Rei80]. Assumptions [Wag70]. astrology [Smi70d]. astronomical [PC67]. astronomy [Bal69]. Atlas [Yor64, Sch67]. atmosphere [FB73]. atmospheres [Gus73]. atmospheric [Fro63]. attitudes [Wil73]. attributes [Hil79c]. Aufgaben [Pau71a, Pau71b]. Aufl. [Bar74]. AUGMENT [BHY80, BJ77]. augmented [Pil70]. Augmenting [DZ78]. August [FGR72]. ausfuhrlicher [Sto71]. Ausgewaehlte [Spa75a]. Ausgewahlte [Spa75c, Spa79b]. Austin [Axf72]. Auswertung [Fri75b]. Authoring [Mis78a]. Authority [Int75a]. authors [Shn77]. Auto [Ent63, CS61]. auto-instructional [CS61]. Auto-primer [Ent63]. autocorrelation [Hem70]. Autoinstructional [Hui65, CS62]. Automata [Loe74]. automated [Car74b, JM76]. Automatic [AK80, Bac56, Bau56, Bee70a, Blu77, Cha74, DP76b, Håv71, HN58, IBM56, Int59e, Ked80, Ken70, Ken80, Pat73a, Sch72a, She59, BBB^{+}57, ES75, Gor64, Hae77, Her64, Int57b, Int57a, Int57c, Int57d, Int59a, Int59d, Int59f, Int63f, JK74, MSR66, MT72, PH77, SWL68, Ste70, vM75]. automatica [HDN74]. Automatically [Par75]. Automatique [IBM58]. Automatisierung [Tec72]. autoprogram [Pic66]. autotester [SJ62a, SJ62b, SJ63]. aux [Lou73]. Auxiliary [Her71]. av [Hus76]. available [CJ77, New72]. averages [She70a]. averaging [Hil79b, Kra74, Nut76, vM79]. Avoiding [Owe65a, BS64]. axial [KM73b, KM77a, KM77b, Kat77]. axial- [KM73b, KM77a, KM77b, Kat77]. axially [Per80, Tho65]. axis [McC71]. axisymmetric [Hol67, PNK65a, PNK65b, Ree72].

B [BK75, Bur78, Gar72, Nak77, BD71, Bur68b, Bur70b, Bur74, Con75a, Fre74, Han74a, NL71, Ste72a]. B1700 [Hug78]. B2DATL [AK77]. B4Tran [MH75b]. B5500 [Ben69, Bur67]. B72 [Swa72]. B72-22 [Swa72]. Back [MR70a]. Back-Calculation [MR70a]. Background [Kie66, CL80]. backward [Nut76]. balance [Leo74]. balances [Soy71]. Balancing [Day76]. Balfour [VP80a, VP80b]. bands [MW71a, MW71b]. bandwidth [Cla80]. Bank [Sch72b]. Barbara [SMD71]. Barraud [Ric73]. basal [Gof74]. Base [DW77, SM75, VP80a, VP80b, Con78a, CJM67, Fis78, Leo74, Sof80]. Based [Aki74, Axf72, BJ77, Cra76, Ina80a, Ina80b, Jam70, Ree79, DH78, Din69, FL76, Gut79a, Gut79b, HD78a, Jos78, Kal72a, Lee67b, Lee72, O’N74, PC78a, Phi71a, Phi71b, Plo77, PK67, PK69]. Bases [DPR70, Wol68b, Smi67a, Smi71a, Smi73a, Wol68a]. Basic [AB66a, AB66c, AB66e, Ame66c, All67, And64a, BB72, BB77a, BP74, Bla69, Bur69, DR70, For71, For74, FHS7x, Gue73b, Har70, Har66c, Har68b, HDBP68, Law77, Bro61, Dun80, Fos74, Gue73a, Hei72b, Int60b, Int65e, Int66b, Int66c, Int8 , IBM68, Int68h, Int68i, Int68a, Int69b, Int69c, Int71e, Int73a, Key73a, Key73c, Key73b, LS71b, Lee67b, Lee72, LJ71b, LJ71a, Mac73, Pra65, Pri69, Sla71, Sle75, Tip76, Xer70c, Dig72a, Dee74a, Dee74b, Dee74c, Dee74d, Kan77, Lee74c, Lyc80, Mur77c, Mur77a, Mur77b, PTM77, PC71a, PC71b, PC78b, SB78, Car79a, Car79b, Car79c, Coa80, EF76, Gen66b, Gor64, Gre77, HD75, Hec63, KC72, Law78, LHKK79a, LHKK79b, San73, Smi70b, Gar72]. BASIC/FORTRAN [Mur77a, Mur77b]. basico [BB78, Har74, SST72]. basics [Hol72]. basins [Oja70]. Basis [Nie72a, Nie72b]. Batch [Bee70b, RT77, Chi73, Coh74, Pin73]. batch-transmit [Coh74]. Batching [REC75]. Baukastenverfahren [Sch77]. BC [TB65]. BCD [Ost62]. Bd [Jun69]. Be [Bac78a, HM75, Hor75, Jon64, Per80, Wei66a, Wei66b, Ste72a]. Beach [Tou70]. beam [DW70]. beams [Sik71]. beginner [CS71e, CS77c]. Beginners [CS77b, DS77, Cou70, NS76, Tay77]. Beginning [Man71, MHH71a, MHH71b, MM80, McK80, Bre76b, SW74]. behalf [EO66]. Behavior [Boi74, BP76, Shn76]. behavioral [LB68, Vel67]. being [Dav72b]. Beispiel [Rin77]. believe [Hul73]. Benchmarking [Bas80]. benchmarks [Uni75a]. Benutzungsanleitung [Die72]. BEOS [Die72]. beräkning [Hus76]. Berechnung [Die72]. beregninger [Tju68]. Bertalanffy [Abr65, All66b]. Bessel [Col80a, FM76, Gar78, Gav76, Sca71]. Best [LaM72]. Bestimmung [RS69]. Betriebssystemen [Win79]. better [Ske79]. Bettis [Bri79]. between [BS64, Ber70a, Can77, LML69, Pal68, Sch79c, Sch80g, Tip76]. Beullasten [Die72]. BFORT [Ono79a, Ono79b]. BGFIX [CL80]. BGSUB [CL80]. BHYLD [PT73]. bibliothekarischer [Sto71]. bicubic [Gaf79]. Big [Smi71b]. bilimler [Yur76]. Binary [WB65, Ant77, Day76]. binomial [OG69]. biological [GKB74]. biologie [LS71a]. biologists [FB69]. biology [LS71b]. biorhythm [Smi70a, Smi72a]. Bit [Bee79a, Bee79b, Bee80b, Tob65, YP80, Ins76a, Ins76b, Ost64, Owe65a, ABH^{+}71]. Bits [DW71]. Bivariate [Aki74, Don73a, GKB74, ClW78, Wal68]. blade [MC70, Mei78]. blast [U.S61]. blending [Sch80f]. block [Kar73]. Blocks [Han72a, Fos74, Han72b, Var77]. blokowymi [ATW77]. BMDP [Fra79]. BNF [Mau72b]. body [Phi67]. boiling [Bak68, Sol64]. Book [Bar71a, Bar72a, Bar73a, Bar80b, Bar74, Bri67, Bru66, Cha73, Din72, Elt66, Flo70a, Flo70b, FJA80a, FJA80b, Fry71, Gar72, Gar74, Gow73, Hil73, Hil79a, Hui65, Jun68, Jun69, Kau78, Kre66b, Kre66a, May73b, Nak68, Pat73b, Ree73, Ree75, Ree76, Ung69, VP80a, VP80b, Van68b, Wil74, Wil75, Wol68b, AB69b, Bro71a, Bro73a, Haa69b, Hun74, Int63m, Int68e, MHH71b]. Boolean [Cam65, Les73, Mon78]. BOPTIC [DW70]. borehole [BC72a, Sco78]. both [NS76, Sca71]. Bottleneck [DZ78]. bottom [MG71]. Bouguer [KRB77]. boundaries [Wei75]. Boundary [Sal77b, Epp74, LP74, Sal78]. bounding [Sho76b]. Bounds [BCKT79, Ant72]. Brent [BHY80]. BRENTM [MC80a, MC80b, MC80c]. Bretschneider [Lal75]. Briandais [Bem61]. Brice [Ree75]. bridges [Sik71, BT76b]. BRIDGES-a [BT76b]. brief [Fel75a, Mat72b]. BRLESC [But66]. BRNANL [Fos74]. Bromwich [Pie73]. Bruno [Kli73]. BSOLVE [Bal73]. BSR [SIG76]. budget [Cra75, Wil77b, Wil77c, Wil77d]. budgeting [Lyt75]. Buffering [Fer60]. building [She78a]. buildings [FMC78, ST73a, San74]. bulletin [Con71d, Gol68a, Gol68b, Moo60, Sha71a, Sha71b]. Bürmann [Lou73]. burnable [Zaa69]. BURNAPAN [Zaa69]. Burroughs [Ben69]. bus [RSBR69]. Business [Car79b, Did78, Jay80, Lew80b, Mar80, May73a, Mur77c, Mur77a, Mur77b, Mur80, O’b75, See75, Wu73b, Wu73c, Wu77a, Wu77b, AB68, AB69b, Bar70, Cal69b, Car79a, Car79c, CS77a, DEN79, MG68, Nie68, Nie71, PS74, Rau68, Rob62, Sch69, SM70, Sch70, Sch71, Fai74]. Business-Oriented [Fai74].

C [Bar73a, CJ77, Din72, Fry71, Gar72, Hui65, Jun69, SS68b, CJ78, Ste72a]. C-256 [CJ78]. C.C.C [Ano70d, Ins64, Ins74]. C2 [JT72]. C5 [FSC73]. C6 [Pie74]. CA [ACM78, IEE78]. cable [SM73a]. cadre [Lou73]. CAI [Lin76, McA77a, McA77b, Mis78a, Pen68]. CAI/CMI [Lin76]. CAL [Uni69b, Uni73, Uni74a]. calcolo [Rid67, Rid78]. Calcomp [Sou71, Ste73, Wil80a, Wil80b]. Calcul [Lou73, Fer63]. calcular [dC73]. Calculate [Ste78a, Ste78b, WD79, Bar80a, Cam77, Col80b, GKB74, MG71, MS66, PV74, San74, Tho65, Win74]. Calculating [Man72a, Ste76b, Ste76c, Bea75, BC72a, DS67b, Die68, Fla71, Fox64, Gus73, HPB73, Joy77, Joy78, KM73b, KM77a, KM77b, Kat77, Kli73, LTB80, MW71a, MW71b, MW75, Mei78, MI75b, PB73b, PJT76b, PJT76a, RG68, Smi63b, SH78, ST73b, Tri73, Tro66a, Tro66b, Bur76, WCT68, Wil77b, Wil77c, Wil77d]. Calculation [DP74b, Ein72, Joh74, MH75a, MR70a, SS73, Spi72, Cas62, CM66, Epp74, FH71, Fri71a, Ful74, HM62a, HM62c, HM62b, HO64, Her64, HM64, HW75, KC60, Kol74, Lyn63, MC64, MU75, Moo76, Mot66, Pol78, PC67, SM73b, VV66, Wei75]. calculations [Der64, Las71, Phi67, Ste72a]. calculatorul [CJ78]. calculo [Gol76]. Calculus [AW73a, AW73b, Ful77]. Calibrating [Bas80]. Call [Ser71, Int72l]. CALL-OS [Int72l]. Call/360 [Ser71]. callable [Uni74b]. calls [Bid79]. Cambridge [Bar73a, Bar80b, Pat73b]. camera [Har73]. Campbell [BH1 ]. Can [Bac78a, Hor75]. Canada [LS71a, LS71b]. canonical [Lee69, RR70]. Canterbury [Eva72a, Eva72b]. capabilities [Car66]. capability [Blo68]. capacity [DP74b, Sch80a]. Capture [Whi71a, Whi71b]. capture-recapture [Whi71b]. card [BD80a, Int75b, Per80]. card-imaged [Per80]. cards [BCE77].

Carlo [Fer63]. Carlo/optimal [Feh68]. Carnahan [Ree75]. cartes [Dav74]. cartes-controle [Dav74]. cartography [Her69]. cascades [AK77, Der64]. Case [DM72c, Kar77, Mue66, Per72a, Per72b, Ree73, Sch69, Sch70, Sch71, Wei67]. caso [LMP77]. cataloging [Feu77]. Catalogued [Bee71c]. Catch [AS71, PB67, SH67]. category [Uni80a]. causes [Pat67]. CBL [Uni78]. CBL-FTN [Uni78]. CBT [Uni78]. CC [Ste60a]. CC-149 [Ste60a]. CCAE [Hur77, HF78]. CCP [Sch62]. CDC [Håv71, Ano72a, BH73a, BA72, Con69c, Cor79, Com80a, FLM70, Fri73, FLM74, FL74, Gre75, Hob67, KQS74, LM70, Lin71, MR70a, Rey69, RR70, Sei72, Shu75, Sta65, SH67, Uni75b, Uni77, Whi68, WM9 , Wra70, ZSW76, ZSW77]. CDC-3600 [Sta65]. CDC3300 [Mei68, Mei69]. Cell [BJ74]. Cenfor [Uni71]. Census [Uni71]. center [Pec77, Hyd66a]. centers [Tro64]. central [Wie75]. Centre [Bar72a]. centrifugal [Gal73]. Century [NCR69, NCR70, Nat73]. CEP [Man64, MM65]. CERN [Håv71, Lou73]. certain [Bec73, GS70]. certains [Tel80]. Certification [SSS78, Ste79]. cetiri [Bit75]. CFD [Ste75b]. CFT [Hig79a, Hig78]. chain [Kru67]. chaining [Arn65, Dra64, Har65c]. chains [Agh77]. chair [EF76]. Changing [Bee70c]. channel [KM73b, Las71, SDH74]. channel-spin- [SDH74]. channelized [Sig80]. Character [ABH^{+}71, Bee79c, Bee79d, Lew63, Poo62, Pyl62, Rey77, RH76, Smi63a, Lea67, Mor75, Wol78b]. Characteristics [Boy74b, Man64, Bec73, Har77, Her64]. Characters [WM77]. charged [Tro66b]. chart [Ste70]. Charts [McI71, BC77, BC79, Leu79b, Dil79, Fis79]. CHEBINQ [Win74]. Chebychev [Fut78, Sim76b]. Chebyshev [Sim76a, Win74]. Check [Spi65, Edg79]. checking [Ube76]. CHEMANAL [Col80b]. chemical [Col80b, PJT76b, PJT76a, SM73b, Tro66b, WS71]. chemically [Erd80]. chemischen [WS71]. Chemistry [Bee75, Dic68, DS62, DS63, DS76]. Chemists [Ise78, IJ79]. chen [Ano75b]. Cheng [jT79, Ano72d, pC79, lH80, Jam75]. Chernoff [TT80]. Chernoff-Type [TT80]. Chess [KC73, BB71, BW75, Gil70]. chi [Ano75b, pC79, lH80, jT79, Wie75]. chi-square [Wie75]. Chief [San70]. chih [Ano72d]. Chippewa [Con66a, Con66b]. CHLOE [Bra72a]. CHNGLC [Bar75]. Choice [ADG70]. Chromatic [JR75]. chromatograph [Bar75]. chromic [BP76]. ch’uan [Ano75b]. chungsim [JcK73]. ciencias [MD66a, MD73]. CINDY [SR73]. CIPW [VV66]. circles [Dil79]. circuits [U.S61]. circular [Cam77, MC70, TH62, WCT68]. circulation [U.S61]. Civil [McC75]. cladistic [Bar66]. Clarification [ANS69c, ANS71a, ANS69a, ANS71b, Ano69b, ANS69b, ANS71c]. Class [BH73b, Osy76, Ham74, Man63, RG76]. classes [Her74]. Classification [BB77b, Hil79c, Sal71b, Spa80]. classifications [Dem69]. Clear [Jaf78, Jaf72, Jaf79]. Clebsch [Tam66]. Clenshaw [Gen72]. climatic [Wil77b, Wil77c, Wil77d]. close [RG68]. close-packed [RG68]. CLUSTAN [Wis69]. Cluster [Spa73, Spa80, BC67, Par70, Spä76, TB65, Wis69]. Clustering [Pag74a, Sal71a]. Clusterwise [Spä79a]. CMI [Lin76]. CMS [Uni79, Int72a, Int72b, Int72c, Int72d, Int72e, Int72f, Int72g, Int74a, Int74b, Int75b, Int75e, Int75d, Uni80b]. coalescence [AJ69]. coastal [Lar69]. COBOL [Rin77, Ber64, BDI72, Can77, Car79b, Elk65, Har69, Kan71, KF72, Kha76, KG72, Kuo74, Lee74c, Mat72b, Mur77a, Mur77b, Sha65, Sof80, Taj65, Tha77, Vei66, Car79a, Car79c, DT74, Kuo73, LMP77, Obr70, Obr71, Slo68, Sun73, Uni78]. COBOL- [Rin77]. CODASYL [Con77a, Con80a, Sta74]. Code [Int74a, Lea78, NC76, Bak68, Bar77a, BS78, Cas62, ClW78, ES75, FH71, Gea65, Gro73d, HPB73, Hug77, Int71d, Int72h, Int72f, Kau69, Kro75, Lau80, LM69, McG67, Mot66, Mue75, Mur66, PB73b, Rus79, Sal70, Saw62, Sho76a, Sho76b, Sol64, Sta65, VV66, Zaa69, vM75]. Coded [Air77, Kno72, Kno75a, WM60, Kno75b]. Codes [LP73, DS75, Fis79, Fri71a, Jac78, Ske79]. CODEX [Wer72]. Coding [Bac56, HN58, IBM56, Int59e, She59, Yoh72, BBB^{+}57, Int57b, Int57a, Int57c, Int57d, Int59a, Int59d, Int59f, Int63f, Int75a, Jam73a, Kir79]. Coeff [KPG63]. coefficient [KPG63]. Coefficients [Pie74, AD73, CM66, Gus73, MC64, Pic66, RV78, Tam66, Vic70a, ZT76, Zoh72]. coherent [SDH74]. COKO [KC73]. Colin [Bar80b, Cha73, Gow73, Hil73, Hil79a, Pat73b, Wil75]. collaborative [Jul75]. collect [vM78b]. collection [Cli78a, Cli78b, Nor63]. collector [Ree72]. college [Zor68]. collision [TS73]. Colman [Hui65]. Coloring [Kau78]. com [Ano70d, CC70, Ins64, Ins70, Ins74]. combination [MU75]. Combinations [Gen75a, Gen75b, Gen75c, LT73, Wei66a, Wei66b, MB68a, MB68b, SMM65]. combining [Elk65]. comentate [HDN74]. Comes [Bri79]. COMFORT [Cla78]. commentary [Sch66e, Sch66b]. commentes [DG75, Tho72c, Tho78]. Comments [Bla60, Fel76a, Fel75b, Kuk66, Kuk67, Row76, HLS73, Owe65b]. commercial [Int8 , IBM68, Int68h, Int68i]. Committee [JV67a, JV67b]. commodity [Hol80]. Common [Ein76, Con75b, Con75c, Con79d, Har69]. Communicating [Jac73a, Jac73c, Jac73b]. communication [Phe76]. compact [Blu70]. companion [RR73c, RR73d]. Comparacão [dC73]. Comparative [Gre77, Raf79, Dav70, Dav72a, Kan71, Lee74c, Pin73]. comparing [MR78]. Comparison [Fit74, Pal68, Rod76, Woo77a, Woo77b, BS78, Car78b, Cha70, Ham74, Har69, Kee75, Mat72b, Nut78, Phi71a, Phi71b, Pot66, Sch79c, Sch80g, Slo68, ST73b, Tha77, Wri66]. comparisons [Gen77b]. compatibility [The68]. Compatible [Ano68c, Day78, Wil80b, DH78, Dav72b, HD78a, Mei74, Ste73, Wil80a, Hil79a, Bar80b]. compensation [Com80b]. Competence [Smi68]. competent [O’D74]. Compilation [Bee70b, RCM66, Die77, Gea65, Hai65, Rob68]. Compiled [GHG60, Han60, Mer60b]. Compiler [AU77, BE74, Bee71a, Con76d, FW78, Fel79, Hig79a, Int68b, Int74c, Lee67a, Lee74b, MM75, PB74b, REC75, She59, Bid79, Bla68a, Bla68b, Boa69, Bob70, Bur68b, Con66a, Cau78, Col75, Con77c, Cor61, Dig71a, Dig72b, Dig72c, Fit75, Hai65, Har78, Hea79, Int72d, Int74b, Int74d, Int80a, Kei69, KS68, Lar63a, Lin71, Low76, lAL72, MHM^{+}68, Mer58b, Moo75, Moo77, Sys73a, Sys73b, Ste80, San78, Sas69, SA73, Sch79b, SGM^{+}67, Sle75, Ste73, Ste72b, Sti62, Tay68, Wer72, BJ77]. Compilers [MS73a, Poo74, SFIK79, Bla71, Dav70, Dav72a, DH78, HD78a, KS70, Knu62, Kro75, Sco77a]. Compiling [Spe80a, Vic64]. complainants [Hal72b]. Complement [NO75, Mon78]. complemento [SS78b]. completa [Lec68]. Complete [Ano77a, BC70, Lec66a, Lec66b, Wat68]. Complex [JT72, Kuk72a, Kuk72b, Mil68, Mon75, Rei80, RK73, FM76, Mue66, Sca71, SM73b]. complexity [MR67, Sel77, Tan80c]. component [Dun75b]. components [Rap66a, Wah68]. composability [McC74b]. composicão [dC73]. Composite [FJA80a, FJA80b]. Composition [SH67, FB73, Hun74, Kot72]. compound [SMD71, SR73, SDH74]. compound-nuclear [SDH74]. compounds [Ste72a]. Comprehensive [Haa65, Haa69a, Haa69b, Jn69]. Compressed [MS73a]. compression [DP76a]. compressor [FH71]. compressors [Gal73]. computacion [Car78a, DG70, JSW70, dMdF73]. computador [Ano70d, CC70, Ins64, Ins70, Ins74]. computadoras [Far74]. Computation [AG80, AK77, BKW74, BKK^{+}80, BC72b, ES74b, Fro63, Pie74, SMD71, SR73, VS80, Coc80, DP76b, Gav76, Hol67, Hol68, JSW67, JSW77a, JSW77b, Kau69, Lil68, OLS66, PTM77, Rap66a, Rap66b, Sch66c, Sch66d, Sch66e, Sch66a, Sch66b, Wah68, ZD78, Zoh72]. Computational [Gin78a, Hun76, Mat89]. Computations [FMM77, FMM80, Mol72b, FP75, Mol71, Sim66, Wie75]. compute [EP67, Plo75, Plo77, Sca71]. Computer [AI78, Abr72, AS71, AB69a, AG80, AW73a, And64b, And66, Arm78, Axf72, BK72, Ben69, Bog74, Bog80, Boi74, Con68a, Com69, CPM72, CS62, Dah78, DDM^{+}75, DS76, Dic74a, Dic68, Edu70, EF76, Eld70, FJA80a, FJA80b, For70, For75, FMM77, Ful72, Geo80, GS70, GO75a, GO75b, Gue73b, HBJ76, Hog70, IEE75, IEE78, IEE79, IAAA57, Ise78, Jac73a, Jam73b, Ked80, KC72, KH75, Kuo72, Les72, Lew73, Lov75, MR70a, May72, Mer78b, MS75a, MS73d, MR70b, Pen70, Per75, Pil70, Rad75, Rad76a, Rad80, Raj77, Rob67a, Rus78, San70, SSS77, Sch73, Smi70b, Som71, Soy71, Sta75, SM64, SM68, SP70, TW71, Tou70, Ver65, Wag70, Wal72, Wal80b, Wei69, WB65, Whi71a, Wol78a, Wol68b, YS75, Agh77, AW73b, ADT67, And64a, Ano64, Ano72c, BP76]. computer [Bec72, BS78, BW78b, BB71, BC77, BC79, BC72a, Bis75, Blo68, BCS68, BR78, Bre79a, BC70, Bro71a, Bro73a, But66, Con62a, Con62b, Con64c, Con64b, Con64a, Con65, Con66a, Con66b, Con67a, Con67b, Con67c, Con68b, Con69d, Con69b, Con69a, Con70, Con71d, Con71b, Con71c, Con72a, Con73c, Con73d, Con73e, Con73b, Con75a, Con76c, Com80b, CB69, CCL69, CG73, Cha76, Cha71c, CS73a, Cla78, Cla68, Cle68, CJM67, CS77a, Cra75, DvC69, DS66, DS67a, Dav72b, DS62, DS63, Dun67, Dun69b, Dun69a, Dun74, Ent63, EKM74, FB79, FP75, Fis79, Fit75, FL76, Fos73, Fox64, Fro63, Ful73, Gra70b, GP73, Gel69, Gen66a, Gra79, Gre75, Gro73d, Gue73a, Gut79a, Gut79b, Hyd66a, Hyd66b, Hew76b, Hal65, HM75, Har68a, Har69, Har73, HG66, HPB73, Her74, Hob67, Høj69, Hor65, Ins76a, Ins76b, IA78]. computer [Irv60, IJ79, Jac73c, Jac73b, Jam73a, JM76, Jos78, KPG63, Key73a, Key73c, Key73b, Kha77, Kno72, Kru67, Kru68, Lou74, LJ71b, LJ71a, MI64, MH72, MH73, McC74b, MK70, McG76a, McG76b, McM66, MP65, Mer78a, Mer78c, Mer60a, Mis78a, Mur66, New73, Nyd68, O’D65, Obr71, OR77, OG69, Pan70, PG67, PB73b, Pen68, Per72a, Per72b, Phi71a, Phi71b, PG66, PJT76b, PJT76a, PC78b, Pol78, Ral71a, Rau68, Ren65, Rey69, RR70, RSBR69, Rub69b, Rub69a, Rub69c, Sco76b, SD66, SD67, SA74, Sch68, Sei72, She70a, She70b, SDZ80a, SDZ80b, Sik71, SYR77, Sle75, Slo68, Smi67b, Smi73a, SR76, Smi79, Sou71, Squ70, Sta65, Ste72b, SM73b, Swe67, SW79, TS73, Tay80, The68, Thr79, Tom71, Tri73, Tro64, Tug75, U.S61, Bur76, VV66]. computer [Vas72, Vic64, Wal80a, Wal63, War69, Wei75, Wei65, Whi71b, WB71, Whi68, Wid79, Wil69, Wit79b, Wit79a, Wit79c, Wit79d, Wit74, Wra70, ZSD80, Lin76, Bee75, FMM80, Hui65, Ree76]. Computer-Assisted [DDM^{+}75, Mis78a, Pen68, Wid79, Bee75]. Computer-Based [Axf72, FL76]. computer-generated [Her74]. Computer-Oriented [AG80, AW73a, AW73b]. COMPUTER-TUTOR [Lin76]. computer-use [Smi67b]. Computerized [Hei72a, Sch79c, Sch80g]. Computern [GG72]. Computerprogrammen [Bra75a, Bra75b]. Computers [BR74, DW71, Gin78a, Håv71, Ker72, LR77, MG68, MGL73, NO75, Obr70, Obr71, O’b75, OR75, Pac69, Wu73a, Ano70a, Ano70b, Ano70f, Ano73, Ano75c, Ano75d, Ano75e, BC67, Com80a, EB80, Ent80a, Ent80b, Feu77, FLM70, Fri73, FLM74, FL74, Han78, HW67, HBE80, Har66b, Int64a, Int78b, Int78c, Kno75a, Kno75b, LM70, MW69, Mye73, Paw65b, Pri69, SM66a, SM66b, Uni80c, Wil73, WM9 , Wri66, Xer70a, Xer70b, Xer71a, Xer73, vM75]. Computing [Bar72a, BH1 , Din73, DG67, Gaf77, Gaf80, Gar78, Jun68, MH78, Mon77b, Mon77a, Mon79, Ree75, Sci69, San73, Sch80c, Sch73, SP70, Tea72, AJ69, Car68, CW71, CW72, CW73b, CW73a, CW76, CW77, CW78a, CW79, DG68, DLS79, FB73, Gol65a, Gol65b, LK74, LB68, MV66, MC70, Nol71, Ott78, RR73a, SM73a, Sol78, Fai74, Hen67]. con [BB78, CDG73, CDG80a, DG70, JSW70, LP79, MD66a, MD73, Sic74]. concentration [Cam77]. concentric [Dil79]. Concepts [Dic74a, Els73, Bur68a, KF72, Sci69, SF75, Wid79]. concerning [Bec73, Kni76b, Lou67, PH77]. Concise [Ral71b]. concordance [Mer74]. condensing [FH71]. Condition [MR70a]. conditional [Sti72]. Conference [ACM78, ACM79, Eva72a, Eva72b, IEE79, IAAA57, Lew79b, Gri78, Ros78]. configuration [WD75]. confinement [BM74]. Conformal [HKK72]. CONFORT [Bri68b]. Congress [FGR72]. conjunction [McA77a, McA77b]. CONMIN [Van73b]. Considerations [Lee77, FK76]. constant [Cha71c, Pic66]. constants [ABB^{+}74, Nor66, O’D65]. constellations [Aub76]. constituencies [Van73a]. Constrained [RK73, Kra72b, MS79, Rob79, Van73b, Wor69]. construct [Bra77].

Constructing [GS79, MS73a, Wra70].

Construction

[BE74, Ant77, PC78b, War69]. consultant [NM74]. contained [IA80]. containers [U.S61]. Containing [GH72, Fic71]. Contemporary [BKK^{+}80]. contenant [Ray63]. Content [MR70a]. contention [IA78]. contents [Int80b]. context [Fri70]. context-free [Fri70]. Contingency [Hab72, MH75a, Sch72c]. continuation [Rap66d, RBp75, Wat75]. Continuous [Ano68c, Edw76a, Gus73, Kru68, Sik71]. continuous-time [Kru68]. continuum [SW75]. CONTOR [RS72]. Contour [Day63, Dic74b, LBG66, McM67, Nor66, SS78a, Tur68, McM67]. Contour-Map [Day63]. contours [RS72]. Control [BF79, Con71c, kC80, DPR70, Hon76, Kuk72a, Kuk72b, Mer78b, Mil73b, Mil75, Wol78a, BK77, BCE77, Bom67, Bri68a, CR74, Leu79a, Leu79b, IA78, Jos78, Lat79, MI75a, Mei75a, Mei76, Mer78a, Mer78c, Mod74, SA74, Sch80f, Sed77, TRW73a, TRW73b, TRW73c, TRW73d, Zal73, Gin78a, Gro73d, VR67]. controle [Dav74]. Controllable [LB80]. convenient [CF71]. conventions [BPW72]. conversation [Gro68a, Gro70]. Conversational [Dat67b, RCM66, Sch73, ADT67, Bar71b, Bri68b, Mar66, Spe9 ]. Conversationally [Ker80]. Conversion [BJ74, McG67, MR70b, Con79c, Cra68, Hyd66a, Hig79a, MM69, Shu75, Ste60b, Van68a]. Converter [Sal76]. converting [Gre75]. Convex [GS79]. Convolution [McC78d, Pic66]. Cookbook [Izz73, Den80]. coolant [Mei78]. cooled [Mei78]. Cooley [Bre67]. Cooley-Tukey [Bre67]. cooling [Van66]. Cooper [KC73]. Cooper-Koz [KC73]. coordinate [Ren65]. coordinates [LML69, MC70, MM69]. Coral [Web78]. CORALL [Veg74]. Core [Jon79, REC75, Bra79, BP78]. cores [PTM77]. COREST [BP76]. Coroutines [SP78]. Corporation [Hol77]. correct [CL80]. correcting [Sco78]. Correction [Pot66, Zoh80, Hei72a]. corrections [BC72a, Fer63, Plo77]. correctness [Mau77]. Correlated [HK72]. Correlation [Fas67, MH75a, Aub76, Die76, Die77, KRB78, Lee69, Mag71, RB76a, RBK76, Veg74]. Correlation-Regression [Fas67]. correlations [Ful74, MW75]. correspomdientes [dMdF73]. Correspondence [Art75, Cha72, Coo72, Dew72, Ehr72, Eng74, Fin72b, Fla72, Hal72b, HL70, HV74, Mee72, SS72, Tay68, Wel70a, DCHR76a, Hil79b]. Corrigenda [Yat71b]. Cosine [DA68, Lin72]. Cost [Axf72, Mid74, BA73, New72]. Cost-Oriented [Mid74]. costs [Cra75, Sin78]. count [OG69]. counter [Ken65]. counterflow [Saw62]. Counting [Van73a]. counts [Sit78]. coupled [Car74b, Las71]. Coupling [Mod74]. Cours [Ano78b, Cha67, Lou73, VG77]. Course [AW73a, BE74, Fre74, HPR78, Mon77a, Ste75a, SP70, AW73b, Baj72, Bau79, Blu70, Bur72, Bur71, Cal69a, CS71a, CS71b, CS71c, CS71d, CS75, Coo76b, Dig80a, Day72c, Ell80, ES78, GO75a, GO75b, Hew80b, HM75, Hon73b, HPR77, HcL78, Int63l, Int63m, Int63c, Int63i, IBM68, JOW72, Kha76, Kha77, Lee67b, Lee72, Lin76, Lot71, Mon77b, Mon79, New75, NBH70a, NBH70b, Plu65, PN68b, Rad76b, Ral71a, Rat72, Roh73, San73, SJ62a, SJ62b, SJ63, Smi67b, Smi68, SW79, TW71, Tom71, Uni78, Wat68, Wil76c]. courses [AHP77, Ful77, Sol78, Swi72]. covariance [Fin68, Fin72d, Fin72f, Fin72e, Fin77, Rey69]. coverage [Mei78]. CP [Xer76a]. CP-V [Xer76a]. craft [Hun74]. CRAY [Cra80, Hig78, Hig79a, Rus78]. CRAY-1 [Cra80, Rus78, Hig79a]. creating [Feu77]. Creation [DPR70]. criteria [Van73a]. critical [DCHR76a, DCHR76b, Mat72b, Tri73]. Critique [Cod67]. cross [Bla79, Cau78, Con77c, Fit75, Hea79, JM76, Kol74, Moo76, Mos78, PV74, SMD71, SR73, Wil65]. cross-compiler [Cau78, Con77c, Fit75]. cross-referencer [Bla79]. cross-system [JM76]. crossing [TS73]. CRT [Coh74]. cryogenic [Har66a]. crystal [FP75, Joh65a, Joh65b, Joh76, Nor63]. crystallographers [MK68]. crystallographic [BML62, BML64, Feh68, Ste72b]. Crystallography [Day63]. crystals [RG68]. CS [Mis78a]. CSMP [JSW77a, JSW77b]. CT [FH71]. Cubic [BH73b, Dur80, MP73, Sca70, Tay80, vM76]. culture [Ros71]. culverts [TH62]. Cumulative [Pom74, Zor68]. Current [Sli71, Kra74]. Curriculum [Mt.79, Fai74, HBJ76, SR76]. Curso [SST72, Far74]. Curtis [Gen72]. Curve [Abr65, Cli74a, Fut78, GM64, Sim76b, Sim76a, Tho72a, Cli74b, Cli74c, Kat68, MP72, MA78, McC71]. Curves [All66b, McL78, DM72a, Din69, Fox64, Maz77]. Cyber [Gin78a, Con76a, Con79a, Eld77, ZSW77, Con71d, Con72a, Con73d, Con73e, Con76c, Cor79, FLM70, Fri73, FLM74, FL74, LM70, ZSW76]. Cyber76 [SM75]. Cybernet [Con74]. cycle [Van66]. cyclindrical [Zaa69].

D [Bri67, Bru66, Elt66, Kar77, Kre66b, McC74c, Ree73, Ung69, VP80a, VP80b, Her78, Cle68, Fas67, RCL75]. D. [Bru66, Kre66b, Ung69]. D1 [Ein72, Gen72, Lin72, Pat73a, Pie73]. D2 [BH73b]. D3 [Ban78c, SS73, Ste79, SS79b]. d’absorption [Fer63]. Dale [Joh66b]. Dallas [IEE75]. Daniel [Bri67, Kar77, McC74c]. dans [Fer63, Lou73]. DAP [Fla77, Int78d, Lim78, RT77]. DAP-Fortran [Fla77, Lim78]. Darst [Her78]. Darstellung [KKU78]. Dashed [Ber76]. Data [AS71, BC72b, Bur79, Car74a, Car79b, CPR75, DW77, DPR70, EB80, Gin78a, Gro73b, HBE80, Has67, HN58, Int75a, Jet79, KG72, LG74, MS74b, Moo69, Rip77, Rul68b, Sch72b, SM75, VP80a, VP80b, VR67, ABH^{+}71, AB68, AB69b, BD71, Bea75, BC77, BC79, BY73, BC67, Bra72a, Bur68a, Bur69, CB69, Car79a, Car79c, CK80, Cla80, Col80b, Com80a, CL70, DP76a, DP74b, Don71, Edw76a, Ell78, Ent80a, Ent80b, ER79, Fis78, Fox67, Fro63, Gor64, Gro71, Gum77, Gut79a, Gut79b, Han75, Har68a, Hat78, Hei70, Hem70, dPW80, Hig79b, Hil79c, Hou62, Int57b, Int57c, Int58, Int59a, Int59d, Int59f, Owe62, Int63f, IA80, Jam66a, JK78, Jul75, KM73a, Kan79, Ker72, KW75, LGF75, Lou67, Lou74, Mac64, Mac69, MI75a, MI64, MP72, MA78, MT75]. data [McL73, Mit65, Moo60, Mur77c, Mur77a, Mur77b, Mur80, New73, NL75, Obr70, OG69, Per80, Pin80, RP74, RPE79, RG77, Rob62, RMM69, Rus79, Sal76, Sei75, SDZ80a, SDZ80b, Smi66, Smi71c, Smi77, Smi79, Sof80, Sou71, Spa80, Squ70, SS74, TC70, TD78, Tur68, Wal68, Wat75, Wer65, Wil65, Yar62, ZSD80, vM76, vM78a, Con62a, Con62b, Con64c, Con64b, Con66a, Con66b, Con68b, Con69d, Con71b, Con73c, Con75a, Gro73d]. Database [Par75, Sta74]. Datei [Rin77]. Datei- [Rin77]. Datenanalyse [Bra75a, Bra75b]. Datenbestande [Fri75b]. Datenverarbeitung [Amk73, Neb71]. Datenverarbeitungsanlagen [jSJ70]. dati [SS68b]. datorprogram [Hus76]. datos [SS78b]. DAVE [OF76]. Davidon [Lil71]. Davis [HD78a]. Day [Bar73a, Bar80b, Cha73, Pat73b, Gow73, Hil73, Hil79a, Wil75]. DC [Lew79b]. DCG [ES75]. DDC [Mod74]. debug [Ano70b, Ano75e, Con72b, Con76b, Con79a, Int72a, Int72g, Int75b, Xer75c]. Debugger [ASH73]. Debugging [AI80, Gra70a, Ho73, Hon71b, Kaz78, Mac68a, Mac68b, PP77, Ste80, Uni77]. debut [Int75e]. debutants [Tho72c]. DEC [BH73a, Com80a, CDGW76, Dig75a]. decays [Tro66a]. decimal [DD68]. Decision [IR78, MR70b, Gul71, Hou62, Lil68, Mac67, Mue75, Vei66]. decision-making [Hou62, Lil68]. Decks [LS75]. declarations [AGG61]. decomposition [Hea68b, Hor68]. deconvolution [Pin80]. Decorana [Hil79b]. DECsystem [Dig72a, Dig74, Dig75b, Edg79]. DECsystem-10 [Dig72a, Dig74, Dig75b, Edg79]. DECsystem10 [Dig77a]. deep [Lal75]. defined [Gut75, Gut76b, Rob67b]. definicao [Cad71]. Definite [Geo80, Rei72a]. Definition [KG72, Mcg80, KM64, Ovi77]. Definitions [Bur79]. deflections [SM73a]. deformation [Die68]. degenerate [MW71a, MW71b]. degrees [OLS66]. del [LMP77, Lec68, Rid67, Rid78, Sic74, SS78b]. deLaval [Tho65]. Delays [Ful72]. delta [Las71, She78a]. deltaic [BCS68]. Demand [Boy74b, Hol80]. dendrogram [Par70]. d’enseignement [Lou73]. densities [Bea75]. density [Kot72, Sal70]. Department [Int75a]. dependent [ABB^{+}74, Die68, PMBK80]. depletion [Sal70, Zaa69]. Derivation [Plo75]. derivative [War75]. Derivatives [Spe80a, Lil71, MI75a, Pow70b, Pow70c, RBp75, Spi65]. Derived [Bak77b]. dérivées [Lou73]. derives [Kot72]. Deriving [JR75]. desalination [FH71, Fri71a, Mot66]. descent [Lea67]. describing [Sch72b]. Description [Cor60, É67, JK78, Mer58b, Mer60a, Mot79, Win75, Ber64, Cor61, Fel75a, Int78a, Lap78, Lou67, Mat72b, She70a, Spe84]. descriptive [Die76, Moo69]. Design [AU77, ASH73, BSK67, DW77, Gil77b, GMPW79, JMG77, Kei69, Lar63a, Nic75a, Pra75, WDT76, Bob70, Boy75, Car77, Gal73, Gil77a, His75, Khu68, Kli70, Lit74, lAL72, Ree72, Rog80, RSBR69, Sal77a, Sch80f, TH62, TC75, Wal70, WG75, FJA80b, FJA80a]. designed [SJ62a, SJ62b, SJ63, Tro64]. Designing [Cra76]. designs [Bre79a, Bry75, Fri71a, Mot66, SA74, Sid72a, Sid72b]. desmearing [Maz77]. DET1 [Sta60]. detecting [Par78, Sco77b]. Detection [IR78, OF76, Hei72a, WK77]. detectors [U.S61]. Determinant [Gow75]. determinants [Mar71]. Determination [All66a, Lew73, Sin73, And73, Gof74, Har73]. determine [Gof74, Hob67, Sin78, Var77, Won67]. Determines [Sal77b, Sal78]. determining [BK77, MI75a, SMM65]. Deterministic [CC74]. detrended [Hil79b]. DEUG [Ano76a]. deutons [Ray63]. DEV1 [Kan79]. develop [Tur68]. Developing [FS76, Vas72]. Development [ACM79, Ano77a, CDGW76, Hei64, Hei66, Kal72a, RSBR69, Con77a, Con80a, Hin76, JS74, O’K64, Oja70, Pec77, Per78, Raf79, Sil61, Tan80b, Whi76]. développements [Lou73]. Deviates [Kno73]. Deviations [Spi72]. Device [War79]. Devices [Jon79, Ano70e, Coh66]. DHAMDI [Fic71]. Diagnostic [WDT76, Pan70, Sco77a]. Diagnostics [PB74b]. Diagonizable [Nik78]. diagram [Squ70]. diagramas [dMdF73]. Diagrams [AB69a, Wil80b, Bro71b, Fis79, War69, Wil80a, Wra70]. dialect [Gra70a, His75]. Dialects [Pyl63, RW77]. diameter [Sco78]. diatomic [SG69, VP75]. dictionary [Bro75]. Difference [GH72, Ada78, Cse75, LP74, Lil71]. Differences [Ano75a, Ber70a]. different [Fat78, FS76]. differentiable [Wer72]. Differential [BH73b, Kar77, Sch75, Sch76, Ste79, SS79a, SS79b, SS79c, Blu65, Bra72a, Cam77, Car74b, Cha74, Fic71, Pic66, RG76, SMD71, SR73, SSS78, SS75a, SS75b, Wil65]. Differentiation [Ked80, Wal63, Pat77]. diffraction [JV67a, JV67b, JV68, KC60, Smi63b, SH78, Tay80, vM76]. diffusion [Fer63, Ray63, Sal70]. diffusion-depletion [Sal70]. Digital [CW72, CW73b, CW73a, CW76, CW77, CW78a, CW79, Cla75, DS62, DS63, Fre76, LL65, LB68, Rob67a, TS76, Ver65, And73, Blu65, BM80, Car68, CW71, Cle68, Dun67, Dun69b, Dun69a, Fit75, GP73, Gre75, JSW67, JSW70, JSW77a, JSW77b, MW69, Paw65b, Tho68, Tor69, Vic64, Wal63, Wil69, Hol77, Ree75]. digitales [Far74]. Digitalrechner [RW69]. digitized [Plo77]. DIGRAF [WPK78, War79]. dilemma [Rei76]. dili [Yur76]. Dimensional [BF71, Žil78, Ant72, Cam77, Dey76, Erd80, Gut76a, Hem70, HM62a, HM62c, HM62b, HO64, HM64, JK78, KR69, MV66, OT80, RG77, RS72, Sal70, SD67, SS78a, SW75, Tur68, Wat73a, Wat73b]. dimensioning [Jef77]. dimensions [Gar63, Joy78]. dinamici [BT76a]. d’informatique [Ano78b, BV74]. Dipole [Kno73, And73, DS67b]. direct [Rei72a]. Directed [Fri69, Blo71, Has67]. Direction [SS73, MP65, Nav78]. directional [Fox67]. disc [Cla68]. disc-oriented [Cla68]. disciplined [DH78, Emb78, Fri75a, HD78a]. disciplines [CK80]. discontinuities [Fic71]. discovery [Tea74]. Discrete [Dur80, Fut78, LP73, Mon75, PC78a, RR73b, Sim76b, Sim76a, Bro80, Lov68, OG69]. discrètes [Ric73]. discriminant [DS66, PT67, PT69]. discrimination [TC70]. Discussion [DCHR76a, Gil60]. Disk [Int74f, Squ70, Int67a, Int70b, Int75c, WCT68]. Disparate [Par75]. displacement [PNK65a, PNK65b]. displacements [Pol78]. Display [BJ74, LG74, ADT67, EH68, Lou67, MCB^{+}62, PTM77, Phi71a, Phi71b, Wah68, Mac68a, Mac68b]. Displays [Les72]. Displaytran [ADT67]. DISPLOT [RG77]. dissociation [Nor66]. distance [Par70, Rey69]. Distances [Smi70c, Aub76]. distorted [vM77]. distortion [Dic74b]. distortions [Hol80]. distributed [Per80]. Distribution [All66a, BR75b, BS75, DB73, Don73a, Kno73, Par75, Pom74, AJ69, Bro80, Mac69, Mag71, McC78d, OG69, RG68, Sou71, Weg66, Wie75]. Distribution-Free [BS75]. Distributions [All66a, Bak77b, Dah78, Lea78, YS75, SDH74]. district [Cra75]. disturbance [Rap66a]. DIVERSE [MG71]. diversity [MG71]. DO [Gen75b, Gen75c, Mee74]. document [Int80b]. Documentation [Ano74a, Bre76b, JMG77, OF76, Ano78a, BPW72, Hon70b, Jam73a, Mic79a, Plo77, Pri77b, TRW73a, TRW73c]. DoD [U.S78]. DODES [Sch75]. Doing [Ver65]. Dokumentation [Sto76]. Dominant [Ste78a, Ste78b]. dominating [Coc80]. domination [CGH75]. donné [Lou73]. Dorn [Jun68, Kar77, Kre66b, Ree73]. dose [HW75, Won67, HW75]. Double [Sca71, Sch75, Fer63, FM76, Har68a, Jam66a, Rei80, SMD71]. double-differential [SMD71]. Dowding [Gar72]. down [CP80]. downward [RBp75]. Draft [ACM76, ANS76a, ANS76b, Kni76a, SIG76, ANS76c, Fel76a, FRS77, Kni76b]. drafting [Cal69c]. Draw [Wei66a, Wei66b, WM77]. Drawing [Ber76, MP72]. drawn [Fis79]. Drill [Org61a, McA77a, McA77b]. drill-and-practice [McA77a, McA77b]. Driven [WDT76, FH71, LGF75, Sho80]. drivers [Boa69]. droplet [Lyn63]. DTFORT [Gul71]. DTSS [Ano70e]. dual [ZD78]. Dubna [Kar76]. duct [KM77a, KM77b, Kat77]. due [AJ69, Hol67, Hol68, Won67]. DULPDX [Uni80c]. DULPDX/DULPLX [Uni80c]. DULPLX [Uni80c]. Dump [NY78, Rin77]. DUMP-Interpretationen [Rin77]. Dumps [Bee70d]. d’une [Lou73]. DWBA [PV74]. dyads [Lil68]. DYNA [KW71]. Dynamic [Bra74, CPR75, Par75, Ran65, RT77, Rog80, Arc76, Ell78, Huy77, KW71, Man63, Sak64a, Sak64b, Sak65, Sak70, Tro66a]. dynamically [LM76]. Dynamics [Cla75, CDH75, PT73]. DYNOSOR [Huy77]. DYSTAL [Sak64a, Sak64b, Sak65, Sak70, Sak79].

E1 [Dur80]. E2 [Aki74, Cli74a, ES74b, Fut78, RR73b, Sim76b]. E3 [Dur80]. E4 [MM73a, MM73b, RK73]. EA1 [Fit75]. each [SMM65]. EAI [Ele68]. earth [And73, Pol78, Rap66b, RZB77]. earthquake [McG76a]. easier [Owe79]. EAT [MU75]. Ebenen [Rot71]. echo [Rus79, Rus79]. ECODIV [GKB74]. econometrics [CK80]. economic [Sad72]. Economics [Weg64, Nie68, Nie71]. economists [MT72, Sla67, Sla72, Gar74, May73b, Van68b]. ecosystems [LGF75]. ed [Rid68]. EDB [Tju68]. edge [Cam77]. edge-loaded [Cam77]. edit [Dor79, Fed63, Bee76, Mor79]. Editing [Bee76, Bee78, Das74, Mur66, RP74, Bee77a]. edition [BK75, Dun67, MS75d, ZSW76, ZSW77]. editor [Jam73a, Kni76b, Lem75, Ube76, WM60, Bem61, Bus67, Elk65, Har65c, Owe65a, Owe65b]. Editorial [Nee75]. EDP [KP70a, KP70b]. EDPM [Int59e, Bac56]. EDSIM [PC78a]. EDSIM-Event [PC78a]. Education [BF79, HK75, Pec77, EH68]. educational [Bai72a, CDGW76, Wal70]. Edwards [HBE80]. effect [U.S61]. effective [KS72b]. effectiveness [McA77a, McA77b, Phi71a, Phi71b, Wit79b, Wit79a, Wit79c, Wit79d]. effects [Sal70, ZD78]. efficaces [Fer63]. efficiency [Low76]. Efficient [Kah80a, Kah80b, Par75, Ste79, SS79a, SS79b, SS75a, SS75b, SS79c, Gea65, HV74, SSS78]. effort [Int75a]. EGN1 [Sta60]. eigenfunctions [Cas62]. eigensystems [Mue66]. eigenvalue [Zak77]. Eigenvalues [Ste76b, Ste76c, Cas62, Nie72c]. Eigenvectors [Nie72c]. Einem [Neh74].

Einführung [Bar71a, Flo70b, Kle68, Kle69, Spi70, SR72, Kle77, SR74]. einheitlicher [KKU78]. Ekman [Sho80]. ekonomickych [Ham79b]. elaborazione [SS68b]. elastic [MSNC61, Mel62, MSR66, PNK65a, PNK65b, Rin79]. elasticity [Cha71c]. elastique [Ray63]. electric [And73]. electrical [LTB80, MS75a, Pin80]. electricity [Gro73a, Gro73c]. electromagnetic [Wat75]. Electronic [Int64a, Int78b, Int78c, Bur68a, Bur69, Har66b, MW69]. electronico [Ano70d, Ins64, Ins74]. electrostatic [Ree72]. elektroniczna [ATW77]. Elektronischer [Mue69]. Element [GF65, Hol67, Hol68, NM70, NM78]. élementaire [Tho78]. Elementary [Bog74, Bog80, Bur68a, Lyc80, Pet76, AW73b, FB69, Fri75a, FK76, Wid79]. elementi [PCR76]. Elements [BV74, Bez73, HH77b, HH78, KS72a, MH72, MH73, Pri69, Pri75, Uni75b, Uni68b, You76, BK77, Hei72b, KS72b, Smi72d, Sri69, Ste72a, Tro66b]. elemzese [Kor77]. Eliminating [Pag74b, She78c]. Elimination [She78e, She78d, She78b]. Ellips [Ber77]. ellipsoid [Joh65a, Joh65b, Joh76]. ellipsometer [Ber77, MC64, McC69b]. Elliptic [Ste79, SS79a, SS79b, SS79c, SSS78, SS75a, SS75b]. elliptical [Cam77]. emancipate [SJ62a, SJ62b, SJ63]. Embedding [BW64]. Emily [Ful74]. Empfehlungen [Fri75b]. emphasis [MW69, Nyd68, Sch68, Sei72]. Empirical [Hoa72, Hoa73, Knu70, Knu71, RT76, Cla80, Gut79a, Gut79b, SMM65]. employing [Nut76]. empresa [Oli71]. emulation [Bid79]. emulator [Fel75b]. Encapsulated [Bur79]. Encipherment [FH74]. enclosed [FMC78]. encoding [DD68]. ENDF [BD71]. ENDF/B [BD71]. ENEP [SW75]. energy [MS66, Soy71, Tro66b]. enfoque [Vic73, Vic77]. engine [Dic74b]. engineer [CG73, SJ62a, SJ62b, SJ63]. Engineering [AI78, FS78, Gro73b, McC67b, Ung69, AHP77, Ban75, CK80, Don71, Gar71, Gro71, Gro73a, Gro73c, KG76, Mac70a, MD64, MD66b, MD68, SM72a, SM72b, Tou70]. Engineers [McC75, CG68, Edw69, GC67, Gol66, Mei69, MT72, MS75a, MS68, MS73b, MS73c, MS73e, NL68, SSS77]. English [Hei74]. Enhanced [Gen78, WM77, Dar78, Wol78b]. Enhancement [Mar78b, Int78a]. enhancements [Lam71a]. ENORMSEP [YS75]. Enough [Boy74a, Boy76, Boy80]. enseignement [Ano68a]. entre [Ano75a]. Entrées [BDI72]. Entrées-sorties [BDI72]. entropy [Bar79a]. entry [BD80a]. Enumerating [AS71, LT73]. Environment [ACM79, DPR70, KRS78, Moh77, RT77, RCM66, BM74, Dig71c, PP77, Raf79, Wag80b]. Environments [RH76, Hin76]. EP37 [Ent80a]. EP37-01-00 [Ent80a]. EP37-10-00 [Ent80a]. EPS [Par77]. epsilon [Mik73]. EQRISK [McG76a]. Equalities [Lar63b]. Equation [Ein74, FSC73, Mol72c, Mol72a, Sch75, Bar77c, BW78a, Car74b, Lee74a]. Equations [Abd80, BH73b, MC80a, MC80b, MC80c, Sch76, Ste79, SS79a, SS79b, SS79c, Zoh80, Bar80c, Bra72a, Cha74, CR71, Dif72, Duf77b, Duf77a, Duf80, Fic71, Mar71, NM70, NM78, Nav78, Pic66, Pow68, Pow70a, Rei72a, RG76, SSS78, SS75a, SS75b, Vic70a, ZT76, Ric73, Kar77]. EQUIL [RCL75]. Equilibrium [PT68, Hol80, PJT76b, PJT76a, RCL75]. Equipment [Hol77]. equivalence [AGG61]. erganzungen [Sch74]. Erlauterungen [Sch74]. Error [Kuk72a, Kuk72b, OF76, BML64, Mal70, WK77]. Errors [IR78, Hei72a]. ES-1022 [AE79]. esercitazioni [Sic74]. Esercizi [BT76a, BG78, Rid68]. esposizione [Sic74]. essential [Har77]. Essentials [Bis75, Edu72c, Edu72f, Edu72e, Edu72b, Edu72a, Edu72d, GH73, Hir73, McG70, Edu70]. estensioni [Sic74]. Estimate [Gaf80, MH75a]. Estimates [AK78]. estimating [New72]. Estimation [Kle78, SH67, Bar66, Cha71c, Mac69, MI80, McC69a, Sei75]. Estimator [Ing71]. estructurado [LP79]. etc [Int60a]. ETL [Mis78a]. ETOT [BD71]. Euclidean [Blu7 , Blu78, Spa73]. Eugene [Int75a]. Eulerian [Die68]. Evalquote [CZ72]. evaluate [DS76, FM76]. Evaluating [Shn77]. Evaluation [HPLG79, Kuk66, Kuk67, Mon78, PB74b, Smi80, Wan78, And70, CJ77, Dem69, FP73, GKB74, Kal72a, LH65, Mar71, McG76a, O’K64, Par77, PG67, Rad76b, Sch80a]. evaporation [AJ69, Fri71a, Mot66]. Event [PC78a]. evolution [RW76, Tan78a]. Exact [Bak77b, FP73, Pom74, Pic66, Sch72c]. EXAFS [IA80]. examination [DCHR76a, DCHR76b]. examinations [Her74, Int68h]. example [Org61a, Que77]. examples [Wri77a, Wri77b, Wri77c, Wri77d]. exchanger [Saw62]. EXCHNG [Ste76b, Ste76c]. Excursion [Smi70d]. EXEC [LC75, Uni68a]. EXEC-8 [Uni68a]. Executable [KMC72]. executing [Lan80]. Execution [Boy74b, FH74, Ing71, RCM66, BS78, Mee74]. executive [Ano72c, Leu79a, Ins76a, Ins76b]. exemple [Dub77]. Exercices [Lam74, Lam77, RH78, Tho78, DG75, HH80]. Exercises [Tho72c, Hun74, Int64f, Org61a, Per72a, Per72b]. Exhibit [Par75, Int57e]. existing [Lju80, Rus79]. exospheric [Kot72]. expansion [Ell78, Zal73]. Expects [Ano77a]. Experience [Coo76a, Gin78a, Fri75a, Gle62]. Experiences [Bau79, Jac73a, Sab76, Jac73c, Jac73b]. Experiment [CS73a, PRO80]. Experimental [HH79a, And70, Bri68a, Pin80, TD78]. experimentation [Mis78a]. Experiments [RJAS78, Kru67, Shn76, Sid72a, Sid72b, Tan80c]. EXPLOR [Kno72, Kno75a, Kno75b]. Exploratory [Shn76, Hal65]. explorer [Smi70b]. Exposition [Ral71b]. expression [SMM65]. Expressions [BE69, Hal74, Mon78, Wal63, Wer72]. EXTENDED [Bee71b, Ano70a, Int72d, Int74c, Int74b, Int74d, Int71k, Owe79, Roc70, WLO76, Xer70a, Xer70b, Xer71b, Xer71c, Xer73, Xer75a, Xer75b, Zim69, Ano70f, Ano70g, Ano73, Ano75c, Ano75d, BN76, Con71e, Con71f, Con72b, Con73d, Con73e, Con75d, Con75e, Con75f, Con76a, Con76b, Con76c, Con77b, Con78b, Con79a, Con79b, Con79c, Con80b, Cra68, Leu79a, Leu79b, Far76, Gum77, Int78a, Kaz78, KA71, Mee74, Mei76, PT69, Wil76a, Wil76b, Wit74, Xer71a, Xer74b, Bee71d]. extended-precision [KA71]. extending [Mei75a, MT72, Ree68, Ree71]. Extensibility [Kno70, Gum77]. Extensible [Ker80, Coo76a]. Extension [BW64, CLS64, Fin72c, GH72, Jet79, LG74, Mau72b, Nag80b, Oll71, Cha77, Eld77, Ham69, Mei74, Pow74, Sak79]. Extensions [Ehr72, Fin72b, Mee72, Mil73a, Cha76, Coo76a, Dig76b, Dig76c, Dig77c, FPB72, Fin72a, Par77, Zwa75]. Extremal [Sin73]. Extremely [Kru69, Krá72a]. extremum [Coc80]. extrinsic [LTB80]. exzentrisch [Die72].

F [Flo70b, Sik71, Ste72a]. F. [Flo70a]. F0 [Int67a, Int72k]. F01 [Int74f]. F03 [Int74c, Int74d]. F1 [LHKK79a, NSB71]. F2 [Ste76c]. F2PM [Cor60]. F4 [Abd80, Kub73, Mol72c, She78b, She78c]. F40 [Dig75a]. F5 [MC80b]. F77 [Gen77a]. Faa [Kli73]. FABSTAT [Per80]. FACEDT [SDZ80a]. Faces [TT80, TT80]. facet [Bre79a]. Facilitate [CLS64, Mei75a, O’D74, Sal77a, Kan68]. Facilities [Wil80b, JS74, Sco77a, Tro64, Wil80a].

Facility [BW64, KG72, Nag80b, ADT67, Con78a, Con77a, Con80a, Rob67b]. Factor [Gos80, MR70a, Cla73a, MI64, Mat72a, TB65]. Factored [Hal74]. factorial [Cla73a, Sid72a, Sid72b]. Factorials [Smi70e]. Factorization [ES74b]. Factors [SH67, Cam77, MS66, Pol78, Smi70i, Wil76c]. facts [LS71b]. fallout [U.S61]. Fallstudien [Mac70a, Sto71]. FAP [Int62d, Int63g, Int64c, Ind60, Mer60a, Moo60, Sim66, Fox64]. farm [FMC78, dPW80, Pat67]. farms [Sin78]. FASAP [dPW80]. Fast [BM80, Mon75, RSD65, Spe80a, Ste77, Tay76, Bro80, Clo72, Dav70, Dav72a, Fis70, Sas69]. Faults [Hil71, SS72]. faunal [YHE69]. FDP [Ano70b, Ano75e, Xer75c]. Feasibility [Eld70, Ben69]. feature [Fed63]. Features [Par75, Ver59]. Featuring [MO80, KS70]. FEDIT [CS73a]. Feedback [Mer78b, Wol78a, Mer78a, Mer78c, Vas72]. Fehlern [Rin77]. feladatok [ZSF78]. FELIX [CJ78]. FESW [NM70, NM78]. FESW-a [NM78]. Fete [Ing71]. few [Sal70]. few-group [Sal70]. FFE [BML64]. field [And73, Der64, KR69, Rap66b, Sin78, Tho65]. fields [DS67b, RBp75]. figures [LS71b]. File [Ben77, dlB59, DW77, Har65b, BD71, IA78, JV67a, JV67b, Lat79, MK73]. File-handling [Har65b]. Files [Cor77, Hon71a, Lju80, MR78, RP74]. film [Mei78, Spe66a]. films [MC64]. filtering [And73, Kra74]. final [TS73]. Financial [Per72a, Per72b]. Find [Blu7 , Blu78, Wor69]. finding [Lil71, Sou71]. Finite [DA68, GH72, Pat73a, Ada78, Cse75, DP76b, LP74, NM70, NM78, PV74]. finite-difference [Ada78]. finite-element [NM70, NM78]. finite-range [PV74]. Finiteness [Wag70]. FINSYS [Thr79]. FINSYS-2 [Thr79]. fire [U.S61, MV66]. Fire-2 [MV66]. First [Sla71, SP70, Bur71, CS71a, CS71b, CS71c, CS71d, CS75, Leu79a, Leu79b, Fic71, FM76, HM75, Lot71, PMBK80, Pow70b, Pow70c, Ral71a, Sca71, Spi65, SW79]. Fishing [BA72]. Fit [Dah78, Fut78, Hab72, Sim76b, Sim76a, Tho72a, CJM67, Gut79a, Gut79b, Kat68]. Fitting [Aki74, All66b, Cli74a, MA78, Bra72a, ClW78, Cli74b, Cli74c, Din69, GM64, Jam66a, McC71, OG69]. five [Edu72a, Har69, Int57e]. Fixed [RH76]. fizikai [ZSF78]. FL [Tou70]. flash [Fri71a, Mot66]. Flex [Pow74]. Flexible [BS61, DW77, Fic73, HPB73, PB73b]. Flexural [O’D65]. FLIBL [Int60a]. flight [MI75a]. FLIP [Gre79]. Floating [Mal72, NC75, Rei79, VS80, Bid79, Mal70]. Floating-Point [Mal72, NC75, Rei79, VS80, Mal70]. FLOCHT [Fis79]. Florida [Blo68]. Flow [Rej72, Cha79a, Cla80, Leu79b, Der64, Dic74b, Fis79, Ham79a, Hol80, KM73b, KM77a, KM77b, Kat77, Mei78, vNS63, Ste70, Tho65]. flowchart [Cul80, McM66]. Flowcharting [Edw73]. flowcharts [HBJ76, Has78]. Flowgraphs [Bak77a]. flows [AK77, Erd80]. FLP [Bur73b]. FLS [BML62]. fluid [Pol78, Rin79]. fluid-pressurized [Pol78]. flujo [dMdF73]. fluorescence [Cri77]. fluorescent [Oer71]. FMO [Feh68]. FO [Int68b, Int70c]. FOIL [LO77]. foils [Mur77a]. fold [Whi68]. fonction [Lou73]. fondamental [Uni75b]. foods [U.S61]. For-train [Tho66]. FORALL [Ker80]. FORDATA [SM75]. FORDOC [BPW72]. foreground [Gen77b, Gen77a]. Forest [Gof74, SK69]. Forex [Rob68]. FORGO [Coc60]. FORLI. [Ste76a]. Form [AK80, BK77, Bon75, Bur72, Ken70, Ken80, Sal76]. FORMAC [Bah69, Ber70b, Lud69]. Formal [KM64, Car66, Nag80b]. formalin [LS71b]. Format [Car66, Ran65, Yat71a, Yat71b, Yat71c, Bai63, BD71, BS61, GM64, Gof74, Hea68b, Hig79b, New73, Sal76, Swi64a, Ver59, DM66c, Int62a, Int63a, Int63d, MR70a, Swi64b, Swi64c, SS66]. formatted [Car78b]. formol [LS71a]. FORmula [IBM54, Gaf77, Kli73]. Formulas [Pie73, Ada78, Cse75, Plo75]. formulation [Cla73a]. FORSIM [Car74b]. FORTAN [She78c]. FORTE [Lem75]. FORTED [Ube76]. FORTLEX [Fel76b]. FORTNEAT [Jam73a]. FORTRAN [Bar71a, Bar74, Bee77a, Bee77b, Bee78, Bee80a, BS64, BA72, Bri67, Bru66, Con70, Con73b, Cha72, Dig68, Dig70, Dig71a, Dig71b, Dig71c, Dig72c, Dig74, Dig75b, Dig75e, Dig75g, Dig77e, Dig77f, Dig78a, Dig78b, Dig78c, Dig80b, Dig80c, Dat77a, Deu73, Dew72, Die72, Din72, Dre70, Ehr72, Elt66, Fla72, FRS77, Fri75b, Gen66b, Gen66a, Hew79c, Hew80a, Int71c, Int72j, Int74c, Int74b, Int74d, Int74g, Int76, Int77a, Int78a, Int75a, Joh66b, Jun68, Jun69, Kar77, KKU78, Kre66a, KTZ67, McC74c, Mee72, Mos78, Nak68, Org61a, Owe79, Poo62, Ree73, Ree76, Rin77, Sta69, SS72, Sch74, SGM^{+}67, She59, SS76, Squ70, Ung69, Wil77d, Wu73c, Xer74a, Ame66a, Ame66b, Ame78e, AS71, All66a, All66b, And79a, And79b, Ano76b, Ano80c, Bar72a, Ben80, Tec72, Boy75, But66, CNW80, Col78a, Col78b, DM66b, Fel75b]. Fortran [Flo70a, FJA80b, Gar72, Gar74, Hew80b, HD75, Hen67, Int59e, Int60a, Ina80a, JMG77, Kau65, KC72, Kle68, Kle69, Kle77, Kre66b, Mac70a, ML70b, May73b, MSS78b, New73, Nie72a, Nie72b, Nor66, Ree75, RS69, Sch78b, Sch67, VP80a, VP80b, Van68a, VR67, WD75, Wil80a, Wol68b, ACM76, AB66a, AB66b, Ame78a, Ame78c, Ame78b, Ame78d, ANS78, AB66c, AB66d, AB66e, Ame66c, Ari76, Abd80, Abr72, Ada78, ATW77, AM79, Agh77, AB69a, ABH^{+}71, AW73b, All75, ADT67, Ano68b, Ano68c, Ano69a, Ano69b, Ano70a, Ano70b, Ano70e, Ano70f, Ano70g, Ano72c, Ano72d, Ano73, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano78c, Ano80a, Ant80, Ant77, AE79, Arc76, AI80, Arn65, Art75, ASH73, Aub76, Bus68, BBB^{+}57, BH64, Bac78b, Bac79, Bad77, Bah69, BBB64, BH73a]. FORTRAN [BK77, Bak68, BM79a, Bal69, BCKT79, Ban75, Ban78b, BPW72, Bar80a, Bar77a, Bar70, Bar77b, BP76, Bar79b, Bar71b, Bar73b, Bar73c, BSK67, BP74, Bea75, Bee70a, Bee70b, Bee70c, Bee70d, Bee71a, Bee71b, Bee71c, Bee71d, Bee76, Bee79a, Bee79c, Bee79b, Bee79d, BS78, BW78b, BV74, Bem61, BN76, Ber70a, Ber77, Ber64, Ber70b, BC77, BC79, BC72a, Bez75, Bid79, BS80a, Bis75, BLY70, Bla79, Bla60, Bla68a, Bla68b, Bla69, Blo71, Blo68, Blu70, BY73, Bol76, Bon75, BCS68, Boy74a, Boy76, Boy80, Bra78, BR78, Bre79a, Bre67, BT76b, Bri79, Bri68b, Bro61, Bro71a, Bro71b, Bro73a, Bro74, Bro75, BDI72, BS61, BD80b, Bur72, Bur68a, Bur69, Bur67, Bur68b, Bur70b, Bur73a, Bur74, Bur78, Bus67, Con62a, Con62c, Con62b, Con62d, Con64c, Con64b, Con64a, Con64d, Con65]. FORTRAN [Con66a, Con66b, Con67a, Con67b, Con67c, Con68b, Con68a, Con68c, Con69d, Con69b, Con69a, Con71b, Con71c, Con71e, Con71f, Con72a, Con73c, Con73e, Con73f, Con73g, Con75a, Con75b, Con75c, Con75e, Con75f, Con76a, Con76b, Con76c, Con78a, Con78b, Con79a, Con79b, Con79c, Con79d, Con80b, Con80c, Con77a, Con80a, Com80b, Cal72, Cal69b, Cal69c, Cal69d, Cal69e, Cal69a, Cam65, CS71a, CS71b, CS71d, CS75, Cam77, Can77, CB69, CCL69, Car66, CLS64, CG68, CG73, Car68, CW71, CW72, CW73b, CW73a, CW76, CW77, CW78a, CW79, Car78a, Car78b, Cas62, Cau78, CJ78, Cha79a, Cha71b, Cha76, Cha77, Cha70, CW75, CP80, CS73a, pC79, CK80, CW78b, CR74, Cla78, CCHT67a, Cla68, Cla80, Cle68, Cle70, Cli78a, Cli78b, CW63, Clo72, Coc60, Cod67, CZ72, Coh66, Coh74, Col75, Col76, Con71a, Con73a]. FORTRAN [Con76d, Con77c, Coo72, Coo76a, Cor60, Cor61, Cor77, CS68, CS71e, CS72, Cou75, CS77c, Cou70, Cra76, Cra68, CL70, CDG68, CDG73, CDG80a, CDG80b, Cse75, CL80, Dig64, Dig72a, Dig75a, Dig75c, Dig75d, Dig76b, Dig76a, Dig76c, Dig77a, Dig77c, Dig77b, Dig77d, Dig79b, Dig80a, Dig80d, Dig80e, Dig80f, Dig80g, Dat67a, Dat75, Dat77b, Dah78, Das74, Dat73, Dav74, DS67a, DH78, Dav72b, DCHR76b, Leu79a, Leu79b, DS62, Dea71, Dea77, Dee74a, Dee74b, Dee74c, Dee74d, Den71, Der64, Dey76, DS76, DS72, Dic74b, Dif72, Dig69, Din69, DO79, Doc72, Doc79, Don71, DG68, DG70, DM66c, Dra64, Dre72, Dre75b, DG75, Dun77, Dun80, Dun75b, Dun67, Dun69b, Dun69a, Dur80, DS75, É67, Edu70, Edu72c, Edu72f, Edu72e, Edu72b, Edu72a, Edu72d, ES74a, Edg79, EB80, Eld77, Elk65, EO66, Ent80a, Ent63, Era77]. FORTRAN [Erd80, EKM74, ER79, Fed70, Fed87, Flo78a, Fan65, FB79, Fel76a, Feu77, Fic73, FPB72, Fin72a, Fin72b, Fin77, Fis78, Fis71, Fis76, Fit74, Fit75, Fla71, Flo70b, Flo78b, FB73, For71, For74, For78, For70, For73, For75, For79, FHS7x, Fox75, Fox78b, FB69, FK76, FK77a, FK77b, FGH75, FGH80b, FH71, Fri71a, Fri75b, FLM70, Fri73, FLM74, FL74, Fri80, Fro63, Ful73, Ful77, Gra70b, Gen67, Gen80a, Gen80b, Gen73, Gen77b, Gen77a, Gen66c, Gen70a, Gen70b, Gen69, Gen80c, Gal75, Gal78, Gal73, GM64, Gar71, GH72, Gav76, Gea78, Gel69, Gil77b, Gil70, GH73, Gle62, Gof74, Gol68b, Gol65a, Gol76, Gom79, GM73, Gor64, Got73, Gra70a, GO75a, GO75b, Gre75, Gro68b, Gro71, Gro73a, Gro73c, Gro73d, Gul71, Gus73, Gut79a, Gut79b, Gut76a, Gut76b]. FORTRAN [Hyd66b, Hew71, Hew74, Hew76b, Hew76a, Hew79a, Hew79b, Hai65, HW72, Hal72b, HR73, Ham79b, Ham69, HRH78, Ham74, Han60, Han74a, lH80, Har68a, Har69, Har80, Har66a, Har78, Har65b, Has78, Has67, HV74, HLS73, HB63, Hea68a, Hea79, Hea68b, Hec63, Hei70, Hei72a, Hei64, Hel63, HPB73, HS69, Her72a, HM62a, HM62c, HM62b, HO64, Her64, HM64, Her78, Her69, Her70, Her72b, Her74, dPW80, HW75, Hig78, Hig75, Hig79b, Hil69, Hil71, Hil79b, Hil79c, Hil70, Hin76, Hir73, His75, Ho73, Hoa72, Hoa73, Hob67, HD78a, Hog70, Høj69, Høj70, HP74, Hol71, Hol80, HK75, Hol77, HH77a, HH80, Hol67, Hol68, Hon73d, Hon75a, Hon76, Hon70b, Hon72a, Hon73b, Hon75d, Hon75e, Hon77a, Hon77b, Hon79b, Hor68, Hou71, HN70, HN58, HPR77, Hug77]. FORTRAN [HH78, HcL78, Hun74, Hun76, Hur77, HF78, IBM56, Int57a, Int57c, Int57e, Int58, Int59a, Int59b, Int59d, Int59f, Int60b, Int60c, Int61g, Int61a, Int61b, Int61d, Int61e, Int61f, Int62a, Int62b, Owe62, Int62d, Int62c, Int63j, Int63a, Int63k, Int63b, Int63l, Int63m, Int63c, Int63n, Int63d, Int63o, Int63p, Int63f, Int63e, Int63g, Int63i, Int64a, Int64g, Int64c, Int64d, Int64f, Int65a, Int65b, Int65f, Int65d, Int65e, Int66b, Int66c, Int66d, Int66f, Int66g, Int66h, Int66i, Int66a, Int67c, Int68c, Int68d, Int68f, Int68g, Int8 , IBM68, Int68h, Int68i, Int68a, Int68k, Int68j, Int68b, Int69a, Int69b, Int69c, Int78b, Int78c, Int70a, Int70b, Int70c, Int71a, Int71e, Int71h, Int71f, Int71g, Int71b, Int71i, Int71j, Int71d, Int72l, Int72a, Int72b, Int72c, Int72d, Int72e, Int72i, Int72h, Int72f, Int72g, Int72k, Int72m, Int72n, Int73a, Int73b, Int74a, Int74b, Int74f, Int74e, Int75b, Int75e, Int75c, Int75f, Int75d]. FORTRAN [Int77b, Int79, Int80a, Int71k, Int78d, Ins64, Int80b, Ins76a, Ins76b, IA78, Ind60, Int75a, Irv60, Isa73, IJ79, Jac73c, Jac73b, Jaf72, Jam73a, JSW77a, JSW77b, JK78, JK74, Jef77, Joh80, JID80, Joh71, Joh65a, Joh65b, Joh72, JS74, Joh74, Joh76, Jon76, Joy77, Joy78, KL64, Kah66, Kah80b, KM73a, Kal72a, Kal72b, Kal71, Kal72c, KW71, Kan71, KF72, Kat68, KM73b, KM77a, KM77b, Kat77, Kaz78, KP70a, Kee75, Kei69, KS68, Ker72, Key73a, Key73c, Key73b, KQS74, Kie66, Kir79, Kir73, Kni76a, Kni76b, Kno72, Knu70, Knu71, KG72, Kol74, Kor77, Kot72, Kra74, KW75, KS74, KS75b, KMC72, KBC^{+}74, Kuk66, Kuk67, KA71, KTZ68, KTZ71, Kuo74, KRB78, Lea70, Lal75, Lap78, Lar73a, Lar73b, Lar63a, LTB80, LT75, Lau80, LK74, LBG66, Law77, Law78]. FORTRAN [Lea67, Lea80, Lea75, Lea64, Lec68, Led75, Lee69, Lee74c, LBM^{+}80, Lem75, Leo74, Lep76, Les72, Les73, Lew63, Lew79a, Lew80a, LP71, LM70, Lil68, Lin71, LPJ79b, LB70, Lit74, Lju80, LV77, Lov68, Lum77, LV73, LJ71b, LJ71a, LO77, Lyc80, Lyn63, Lyo74, LS75, LHLM80a, LHLM80b, U.S78, Mas60, Mis78b, Mas71, MS74a, Mac64, Mac67, Mac68a, Mac68b, Mac70b, Mac71, MI75a, MI80, MW69, MRS73, Mal77, Man64, MM65, MHH71a, MHH71b, Man72b, Man63, Mar77b, Mar66, Mar78a, MP73, MS69, Mar78b, MH72, MH73, Mat72a, Mat72b, MA78, MG71, Maz77, McA77b, MT75, McC79, McC70a, McC78b, McC68b, McC69a, MS64, McC61, McC62, McC63, McC64b, MD64, MD66a, MD68, MD73, McC78f, MK70, McG70, McG67, McI71, McK80, Mee74, MP79]. FORTRAN [Mei69, Mei71, Mei75a, Mei76, Mei77, Mei78, MP65, MSR66, MS77b, Mer78a, Mer78c, Mer74, Mer58a, Mer60a, Mer60b, Mes74, Met80, MS71, Mic79a, Mic79b, Mik73, Mil68, Mil73a, MI75b, MM75, MU75, Mis78a, Mit65, Moc69, Moc71b, Mon77b, Mon79, Moo71, Moo60, Moo77, MT80, Mor70, Mor79, MC80a, MM69, Mor75, Mor71, Mos64, Mt.79, Mue66, Mul80a, Mur77c, Mur77a, Mur77b, Mur80, Mur70, Mur66, Mur71, MS70b, MS70c, MS75c, Mye73, Nat70a, NCR69, Nih69, Nic78, Nag78, Nag80a, Nak77, NM78, Nav78, Nee75, NC75, New67, NS69, New72, New76, NY78, Nic74, Nic75b, Nic75c, Nic80a, Nic80b, Nic80c, Nie68, Nie71, Nir69, Nis78, Nol71, NSB71, Nor63, Nor0 , NM74, Nut78, Nut76, Nyd68, NL68, O’D65, O’D74, O’K64, OLS66]. FORTRAN [O’N74, OK72, Oja70, Oll71, Org61b, Org63, Org66b, Osi75, Osi77, Ost62, Ost64, Ove72, Ovi77, Owe65a, Owe65b, Owe79, PD80b, PB74a, Par77, PB73a, PTM77, Par70, Pat74, Paw65a, Paw65b, Pay64, PG67, PB73b, Pen68, PH63, PNK65a, PNK65b, PC71a, PC71b, Per72a, Per72b, Per80, PCR76, PS74, Pet80, Phe76, Phi71b, Pic66, PT67, PT69, PT73, PG66, Pin73, Pit79, PMBK80, Plo75, Plo77, Plu61, Plu63, Plu65, PN68c, PN68a, PN68b, PJT76b, PJT76a, PC78b, PC64, Pol78, PS78, Pow70b, Pow70c, Pra65, Pre70, Pri77a, Pri77b, PK69, Pul64, Pyl62, Pyl63, Que71, Que77, Rad70, Rab62, Rad80, Rad79, Rad88, RP74, Rap66a, Rap66b, Rap66c, Rap66d, Rat72, Rau78, Raw77, Ree68, Ree71, Ree72, Rei72b, Rei72a, Ren65, Rey69, RR70, Ric73, Rid79]. FORTRAN [Rin79, Rob62, Rob67b, RS72, Roc70, Rod76, REC75, RMM69, RSD65, RR73c, Rot71, Row76, Rub69b, RBK76, RB76c, Rul68b, Rus79, Ryd72, Ryd3 , Ser71, Sci64, Sys73a, Sys73b, SIG76, Ste80, Sab76, SG67, Oli71, Sak64a, Sak64b, Sak65, Sak70, Sak79, Sal71b, Sal76, Sal70, SD66, ST73a, San74, Sas69, Sca70, Sca71, SM72b, Sch77, SA73, SS70, Sch80f, Sch69, Sch79b, Sch72b, SB78, Sco78, Sea80, Sea79a, Kan68, Sei75, Sha76, Sha65, Sha71a, Sha71b, She78a, SR73, She78e, She78b, SDZ80a, SDZ80b, Sho76a, Sho76b, Sho80, Shu69, Shu75, Sil61, Sil71, Sim66, Sin78, Ske79, SP78, Sle75, Smi78, Smi63a, Smi63b, Smi67a, Smi71a, SJ72, Smi73a, SM75, SH78, Smi79, Smi80, Sol64, Som71, Sou67, Sou68, Soy71, ST73b, Spa75b, Spa75c, Spa79b, Spe69a]. FORTRAN [Spe69b, Spe70a, Spe77a, Spe77b, Spe77c, Spe78a, Spe67, Spe69c, Spe9 , Spe70b, Spe73a, Spe73b, Spe74a, Spe74b, Spe76a, Spe76b, Spe78b, Spe79, Spe80c, Spe84, SR72, Spi80, Sta65, Sta60, SMM65, Ste76a, Ste70, Ste72a, Ste60a, Ste60b, SG78, Ste73, Ste74, SS74, Ste60c, SSS78, SD73, Ste72b, Ste78a, Ste78b, Sti62, Sti72, Sto71, Sto76, SD72, Str78, SM73b, Swa72, SS79a, SS75a, SS75b, Swe67, Swi72, Swi64a, TRW73a, TRW73b, TRW73c, TRW73d, Taj65, Tan78a, Tan78b, Tan80a, TH64, Tay68, Tay76, Tel80, TH62, Tha77, The68, Tho71, Tho72b, TC70, Tho68, Thr79, Tip76, Tju68, Tok68, Tor69, TR77, TB80, Tro66a, Tro66b, TB65, TD78, jT79, Tur73, TT80, TC75, Tym68, Tym70, Uni69a, Uni69c, Uni68b, U.S61, Uni80a, Uni75a, Uni68a]. FORTRAN [Uni74b, Uni70, Upc72, VP76, Van66, Van73b, Var77, Vas72, Veg71, Veg74, Vei66, VS80, Ver59, Vic64, VAB62, Vin75, Vow74, Vow77, Vow78, VP75, Wag75, Wag80b, Wal80a, Wal68, Wal70, WPK78, WG75, Wat76, WM60, Weg66, Wei75, Wei67, Wei66a, Wei66b, WK77, Wes69, Wet80, Whi76, Whi68, Wid79, Wie75, Wil69, Wil72a, Wil76c, Wil77a, Wil77b, Wil77c, Wil67, Win74, Wis69, WS73, Wit74, Wol78b, Won67, Woo77a, Woo77b, Wor76a, Wri66, Wu73b, Wu77a, Wu77b, Xer70a, Xer70b, Xer70c, Xer71b, Xer71c, Xer71a, Xer73, Xer74b, Xer75a, Xer75b, Xer75c, Xer76a, Xer76b, Yar62, YS75, Yor64, YP80, Yur76, Zaa69, Zal73, ZSF78, Zim69, ZSD80, ZSW76, ZSW77, ZSW79, Zoh72, Zor68, vM79, ANS69c, ANS71a, ANS76a, Ame78f, ANS69a, ANS71b, ANS76b, AJ69, Abr65]. Fortran [Ack64, AR72, Air77, AD73, AM72, AK80, All67, AK77, And64a, And64b, And66, And70, And73, Ano64, Ano67, Ano68a, Ano70c, Ano70d, Ano72a, Ano72b, Ano72e, Ano74b, Ano74c, Ano74a, Ano76a, Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78d, Ano79, Ano80b, Ano80d, AB68, AB69b, AI79, Ars64, Ayc80, Aye63, Bac72, Bac56, Bac80, Bai62, Bai63, Bai72a, Baj72, Baj77, Bak75, BB72, BB77a, BB78, BM79b, Ban78c, Ban78a, BF72, BM74, Bar61, Bar77c, BW78a, Bar79a, Bar75, Bar72b, Bar66, BD71, Bec73, Bec72, Bee80b, Bee75, BG78, Ben77, Ben78, Ben69, Bez73, Bit75, BCE77, Bla67, Bla71, Blu7 , Blu78, Blu65, Boa69, Bob70, Bod77, BJ77, BY78, Bog74, Bog80, BKK^{+}80, Boi78, Boi75, Bom67, BC67, BT76a, BP78]. Fortran [Bor67, Bor69, BK75, Bra72a, Bra77, BGG78, Bra79, Bra72b, Bra72c, Bra74, Bre75, Bre76a, Bre78a, Bre78b, Bre79b, Bur71, Bur70a, Bur73b, Bur79, BE69, BML62, BML64, Con69c, Con71d, Con72b, Con73d, Con74, Con75d, Con77b, Cor79, Com69, CC70, Cad71, Cad79, CS71c, Car77, CF60, dC73, Car69, Car74b, CJ77, Car79a, Car79b, Car79c, CCN^{+}79, CM66, Cha66, Cha67, CR69, CR73, Cha79b, ClW78, Chi73, CM79, CPR75, CCHT67b, Cla73a, Cle66, CF71, Coa80, Coc80, CC74, CJM67, Col80a, CS61, CS62, Col80b, Com80a, CA78, Con79e, Coo76b, CS73b, CS76, CS77b, Cou76, Cra75, Cra80, CDG70, Cri77, CR71, Dig72b, Dig75f, Dig79a, Dat67b, Dah66, DDM^{+}75, DW70, DS77, Dar78, Dat66, DvC69, DS66, Dav70, Dav72a, DW71, Day72a, Day72b, Day72c, DCHR76a, Day78, Day79, DPR70]. Fortran [De 72, Dem69, Den80, DS67b, DP73, DP74a, DP77, Did78, Die74a, Die74b, Die76, Dil79, DM66a, DM67, DM72b, Doc76, DH79, DG67, DM72c, Dor79, DR70, Dre67, DB68, DB69, DB70, Dre75a, Dub77, Duf77b, Duf77a, Duf80, Dun75a, Dun79, Dun74, DT74, Ele68, EF76, Edw69, Edw73, Edw76a, Edw76b, Ein76, Ell78, Ell80, Emb78, Eng74, Eng75, Epp74, EP67, ESD68, Far66, Far74, Far76, Fas67, Fat78, Fed63, FMC78, Feh68, Fel76b, FW78, Fel79, Fer60, Fer63, Fic71, FP75, Fin68, Fin72d, Fin72f, Fin72e, Fin72c, Fis79, Fis70, Fla77, FM76, Fle70, Fle72, FS78, FS80, Fos74, Fox64, Fox67, Fra77, Fra79, Fre76, Fre74, Fri70, FGH80a, Ful74, Fut78, Gaf77, Gaf79, Gaf80, Gaj66, GKB74, Gar63, Gar65, Mer77, GHG60, Gen75a, Gen75b, Gen75c]. Fortran [Gil77a, GMPW79, Gil60, GC67, Gin78a, Gin78b, Gol66, Gol68a, Gol65b, Goo64, Got72, Gre77, Gre79, Gro68a, Gro70, Gro69, Gro73b, Gue73a, Gum77, Gut75, Hyd66a, Haa65, Haa69a, Haa69b, Hal65, Hal69, HV66, HL70, HRH76, Han72a, Han72b, HDN74, Han78, Han67, Han74b, Han75, HW67, Har63, Har64a, Har64b, Har65a, Har65c, Har71, Har73, Har66b, Har77, Har66c, Har68b, Har74, Hat78, HDBP68, HG66, Hei63, Hei66, Hei72b, Hem70, Her71, Hig79a, Hil73, Hil79a, Hol70, Hol72, Hon70a, Hon75b, Hon72b, Hon71a, Hon71b, Hon73a, Hon73c, Hon74, Hon75c, Hon79a, Hor75, Hor65, Hou62, HPR78, Hug78, Hul73, HD78b, HH77b, HH79b, Hut80, Huy77, Int57b, Int57d, IBM58, Int59c, Int60a, Int61c, Int63h, Int67a, Int64b, Int64e, Int65c, Int66e, Int67b, Int68e, Tab66, Int66j]. Fortran [Ins70, Ins74, Ina80b, IA80, Ing71, Ame77, Ise78, Izz73, Jet74, Jac75, Jac73a, Jaf78, Jaf79, Jah80, Jak73, Jam66a, JSW67, JSW70, JOW72, Jam78, Jam66b, Jam70, JcK73, Jam75, Jay80, Jet79, Jn69, JV67a, JV67b, JV68, JCMS76, JCMS77a, JCMS77b, Jon64, Jul75, KPG63, Kah80a, Kan79, Kan77, Kar73, Kar76, Kar68, Kat78a, Kat78b, Kau69, Kau78, KC60, KP70b, Ken70, KS70, Ken74, KS75a, Ken80, Ken65, Ker70, Ker4 , Ker75a, Ker75b, Ker79, Ker80, Khu68, KGY80, Kli70, Kno70, Kno75a, Kno75b, Kra72b, KS72a, KS72b, Kri71, KR69, Kro75, Kru67, Kru68, Jac78, KBC^{+}73, KM73c, Kuo73, KRB77, Lam71a, Lam71b, Lam71c, LG73, Lam74, Lam77, LG78, Lan80, LT76, LP73, Lan72a, Lan72b, LMP77, LS71a, LS71b, Lar67a]. Fortran [Lar67b, Lar67c, Lar69, Lar63b, Las71, Lat79, LML69, LHKK79a, LHKK79b, LB77, Lec66a, Lec66b, LC78, LW66, Lee67b, Lee72, Lee74a, LB68, LH65, Lel74, Ler72, Lev71, Lew80b, Lim78, Lip77, Lip78, LP78, LPJ79a, LP79, LM76, LG74, LGF75, dMdF73, Lot71, Lou67, Lou74, Low76, Lue66, lAL72, LR77, Lyo80, Lyt75, Mas62, Mac70a, Mac73, Mac69, Mac74, MV66, ML70a, MG68, MGL73, Man72a, Man71, Man69, Man74, MG70, MI64, Mar77a, Mar71, MP72, MR78, MS77a, MS78, Mar80, MW71a, MW71b, MR70a, May73a, Maz78, McA77a, MW75, McC67a, McC68a, McC70b, McC74a, McC78a, MM58, McC64a, MCB^{+}62, McC65a, McC65b, MD66b, McC67b, McC67c, McC72a, McC72b, McC73, McC74d, McC74e, McC78e, MC64, McC69b, McC75, McD80, McG76a, McG76b, MM80]. Fortran [McL73, McM66, McM67, MC70, Mee78d, Mee78a, Mee78b, Mee78c, Mee79, Mei68, Mei74, MH75b, MO80, MSNC61, Mel62, Mer79, Mer78b, Mer58b, Mes73, MK68, Mil73b, Mil75, MS66, MS79, MT72, Moc70, Moc71a, Mod74, Moh77, Mol71, Mol72b, Mon77a, Moo75, Moo76, MM78, MC80b, MC80c, Mot66, Mou70, Mue75, MS70a, Mul68a, Mul68b, Mul80b, MS68, MS73b, MS73c, MS73e, MS75b, MS75d, Nat70b, Nat72, NCR70, Nat73, NL71, NO72, NO75, NM70, Neh74, NS76, NC76, Nie72a, Nie72b, Nie75, Nik78, NL75, NBH70a, NBH70b, Obr70, Obr71, Oer71, OG69, Ono79a, Ono79b, Org66a, Org72, OM74, OF76, Ott78, PD80a, Pal68, Par78, PC78a, Pat77, Pat67, PG65, Pau71a, Pau71b, PP77, Pay70, PV74, Per77, PB67, Pet76, PH71, Phe79, Phi67, Phi71a, PT68]. Fortran [PD67, Pin80, Plu64, Pol65a, Pol65b, PC67, Pot66, Pow68, Pow69, Pow70a, Pow74, Pre79, Pri69, Pri75, PK67, PH77, Rad75, Rad76a, Rad76b, Raf79, Raj77, Ral71a, Ral71b, RW76, êR76, RV78, RG68, Rau68, Ray63, Ree79, Rei80, Rei76, Rej72, RPE79, Rey77, RR73a, Rid67, Rid68, Rid69, Rid78, RG77, RW77, Rit68, Rob69, RCL75, Rob79, Rob68, RT76, RT77, Roh73, Ros73, RH76, RR73d, Ros71, RSBR69, Rub69a, Rub69c, RST78, RBp75, RB76a, RB76b, RZB77, Rul66a, Rul66b, Rul68a, RFP73, Rul80, Sco76a, Sco76b, Sci65, Sch66c, Sch66d, Sch66e, Sch66a, Sch66b, Sil80, Rom75, Sal77a, Sal77b, Sal78, SM76a, SD67, SA74, San78, Sas74a, Sas74b, Saw62, SK80, SC79, vNS63, SM72a, SK69, Sch78a, Sch80b, Sch80d, Sch80c, Sch68, Sch79a, SM70]. Fortran [Sch70, Sch71, Sch62, Sch72c, Sch79c, Sch80g, Sco77a, Sco77b, Sea79b, Sed77, See75, Sei72, Sel77, Sel72, Sep75, SST72, Sha77, SMD71, SDH74, She70a, She78d, Sic74, Sid72a, Sid72b, Sie74, Sig80, Sik71, SYR77, Sim76b, Sim76a, Sin73, SM73a, Sla67, Sla71, Sla72, Slo68, SJ62a, SJ62b, SJ63, Smi66, Smi70b, Smi77, Sof80, SS68b, Sol69, Sou71, Spa75a, Spe77d, Spe80b, Spe66b, Spe66a, Spe80d, Spi70, SR74, SE74, Spi65, Sri69, Sta74, SH67, SZ80, SM66a, SM66b, Ste75a, SP70, SS78b, Ste79, Ste75b, Ste76b, Ste76c, Stu68, Stu70a, Stu70b, Stu71, SM72c, SM76b, SG69, SF72, Sun73, SS79b, SS79c, Swi64b, Swi64c, SS66, Tam66, TI72, TS73, TS76, Tay80, Tea72, Tea74, Tho65, Tho66, Tho72c, Tho78, Tob65, TW71, Tom71]. Fortran [Tri79, Tri73, Tro64, Tug75, Tur68, Tur69a, Tur69b, Uni69b, Uni73, Uni74a, Uni75b, Uni72, Uni71, Uni78, Uni77, Uni79, Uni80b, Uni80c, Ube76, Bur76, Van68a, VV66, Vel67, Vic70a, Vic70b, Vic73, Vic77, Vic78a, Vic78b, VHP69, VL72, VG77, VR67, Wag80a, Wah68, Wal72, Wal75, Wal80b, WMM72, Wal63, War69, War79, Wat68, Wat75, Wed75, Wei65, Wei69, Wei73, Wel70a, Wel70b, Wer65, Wet79, Whi71b, Whi71a, WD75, Wil76a, Wil76b, WCT68, Wil80b, Wil65, Win75, Wit79b, Wit79a, Wit79c, Wit79d, WM77, Wol78a, Wol68a, Wol73, WM9 , Wor76b, Wor69, Wra70, Wri77a, Wri77b, Wri77c, Wri77d, WLO76, ANS69b, ANS71c, ANS76c, Yat71a, Yat71b, Yat71c, Yoh78, Yoh79b, You76, YHE69, ZD78, Zav73, Zin79, ZT76, ZN79a, ZN79b, Zoh80, Zwa80, vM75, vM76]. Fortran [vM77, vM78b, vM78a, Bar80b, HBE80, Joh66b, Joh66a, Hui65, Cha73, Wil74, Wil75, FJA80a, Gow73, Pat73b, Van68b, Bar73a]. FORTRAN- [Rin77, Tec72]. FORTRAN-80 [Sch79b]. Fortran-Based [Ina80a, O’N74, Ina80b, Ree79].

Fortran-Basis [Nie72a, Nie72b].

FORTRAN-coded

[Kno72, Kno75b, Kno75a]. Fortran-Compiled [GHG60]. Fortran-Dubna [Kar76]. FORTRAN-IV [Ano68c, CW73a, SDH74]. FORTRAN-Like [BCKT79, KMC72, Ste75b]. FORTRAN-Programmen [Jun69, KTZ67]. FORTRAN-programs [BD80b]. FORTRAN-Standards [Fri75b]. Fortran-System [Neh74]. Fortran-to-Pcode [CCN^{+}79]. Fortran-to-Universal [CNW80]. Fortran-Training [Pau71a, Pau71b]. Fortran-Triplex-Pre-Compiler [BJ77]. FORTRAN. [Ame66c]. FORTRAN/ [Bid79, Con64c, Hew76b, Hew79b, PN68c]. Fortran/ANSI [Ano78d]. Fortran/Format [MR70a, Swi64b, Swi64c, SS66]. FORTRAN/MASTER [Con69a]. FORTRAN/RT [Dig76b, Dig77c]. FORTRAN/RT-11 [Dig76b, Dig77c]. FORTRANe [Cal78, DG78, Lam78]. FORTRANe-IV [Cal78]. FORTRANIE [ATW77]. Fortranner [Fel76a]. forward [Nut76]. Four [FPB72, Fin72a, Sas74a, Ant72, Edu72b]. four-dimensional [Ant72]. Fourier [Bre67, Clo72, DA68, Ein72, Fis70, Har68a, IA80, Jam66a, Lin72, Lov68, Mon75, PS78, Rej72]. fractionation [ZD78]. fracture [Pol78]. frames [EF76]. framework [FHS78]. Franconian [Rub69b, Rub69a, Rub69c]. Free [Bon75, BS75, Bai63, Fri70, Hig79b, New73]. free-format [Hig79b]. French [Ric73]. Frequency [Dah66, Dah78, Lea78, Par75, YS75, DM72a, Din69, Gen66a, Mac69, Mag71, RG68, Sei75]. Freudenthal [AB69a]. FRG [PT68]. FROTRAN [Rob67b]. FRTDAP [McA77a, McA77b]. Fry [AS71]. FTN [Ano75a, Uni78]. fu [Ano75b]. fuel [Hol67, Hol68]. full [Mei78, Pow69]. full-coverage-film-cooled [Mei78]. Function [Kuk66, Kuk67, Kuk72a, Kuk72b, MM73a, MM73b, MP73, Bar77b, BML64, Cha71c, Col80a, Fla71, Hal71, Kot72, Lil71, MS79, Par70, Pow70b, Pow70c, Sei75, Sou71, Spi65, Van73b, Wor69]. Functional [Bac78a, BH73b, DO79]. Functions [FM76, Lag74, LS76, Rei79, Sch72b, Sin73, Spe80a, Wan78, Žil78, Ano72c, BH73a, Fan65, Gar78, Gav76, Gut75, Gut76b, Hil71, Hol80, Ins76a, Ins76b, MM58, McC78d, MU75, O’D65, OK72, RG68, Sca71, SS72, Tay80, Uni69c, Bur76]. fundamental [Nic78, Sch69]. Fundamentals [AM79, AG80, HH77a, KP70a, KP70b, MH78, Nic74, Nic75b, Nic75c, Nic80a, Nic80b, Nic80c, Sam69, Ste75a, Uni70, Wal75, Obr70, Spe70b, Spe74b]. funnies [Jac75]. Funstat [Ros73]. Furi [Bar77b]. FUTIL [Cor77]. future [Lee77].

G [Din72, Flo70b, FJA80a, FJA80b, Joh66a, Jun69, SS68b, Dea71, Dea77, Int66i, Int70c, Int72k, Int72n, Int73b, Int75a, Ste73]. G. [Bar71a, SS68b, Whi71b, Whi71a]. G1 [Int71i, Int71j, Int72b, Int72c, Int72h]. G5 [Bre74, HK72, Kno73]. G6 [LT73]. gage [Var77]. Gagne [Kal72a]. gains [Bar80a]. Game [Smi70k, Hou62]. gaming [Smi73e]. GAMM [WD79]. Gamma [Kuk72a, Kuk72b, CL80, Ful74, Cha67]. gamma-ray [CL80]. GAPID [Bre79a]. gas [Bar75, Joh74, Wer65]. GASP [MSS78b, PK67, PK69]. GASP-II [PK69]. Gaussian [Bre74, Pie73, SA74, She78e, She78b, She78d, She78c, Win74]. GC [Gra70b]. GC-10 [Gra70b]. GCARS [Tur69a, Tur69b]. GE [EP67, Gen66b, Gen69, Har68a, Sch68, Ste70, CL80]. GE-400 [Gen66b]. GE-425 [Ste70]. GE-600 [Gen69, Sch68]. GE225 [Cla68]. Gebrauch [Dre70]. General [Bro80, Fel76b, Har78, Int57c, Int61a, Int63n, Nie72c, PC78a, See75, Agh77, BK77, Bla79, But66, Cal69d, Dat73, DS67b, EH68, Ent80a, Ent80b, Fel75b, Fle70, HN70, Int63k, Int63b, MI80, MA78, MM69, MS70b, MS70c, MS75c, MS75d, Sak79, Sou71, Wei73]. general-purpose [EH68]. Generalisation [Par75]. Generalised [Zak77, Hat78]. generalizability [Bre79a]. Generalized [Ban78b, Ban78c, Ban78a, Bor69, Rey69, Zoh72]. generate [Cse75, Dic74b]. Generated [Lew73, Her74, Soy71]. generates [Hun74]. Generating [DD68, JR76, MS73a, TT80, Lan80, Pan70]. Generation [Gen75a, Gen75b, Gen75c, Hon76, HK72, Les72, Sti72, Bro80, Leu79a, Leu79b, ES75, Fri69, Han78, Hug77, JM76, Kan71, KG76, Kro75, RB76b, RB76c, RZB77]. Generator [DB73, Kru69, Sch80e, Sch79a, Bre74, Edg79, Fel75b, Krá72a, Pay70, Ste70, War75]. Generators [NO75, Ano70c, Cle66, Gro69, Kir79, MB68a, MB68b, NO72, Ove72, U.S61]. Gentle [CA78, Con79e]. geochemical [Hei70]. geographic [LML69, Plo77]. geologic [ESD68, Har73, Hem70, MI64, Oja70]. Geological [Smi66, EP67, Lou74]. geology [Kru67]. Geometric [BF71, Bor69, Lep76]. geometrical [BM80]. geometry [Hun76, Whi68]. geotechnical [PTM77]. German [KTZ67, Ant72, Kas74, WS71]. GHX [RG68]. Given [JR76, Spe80a, MP72, Sid72a, Sid72b]. global [DS75]. Go [Int72h, Col75, Int71d, Int72f, Int74a, Sil61]. Golden [Hen67]. Goodness [Dah78, Gut79a, Gut79b]. Gordan [Tam66]. Gould [Wil80a, Wil80b]. government [Pec77]. GPAK [Hun76]. GPSS [MSS78b, Nie72a, Nie72b, Sch80d, Sch80c, Sch77, Sch78b, Sch78a, Sch80b]. GPSS-Fortran [MSS78b, Sch77, Sch78a, Sch80b, Sch78b].

Grade [New75, TS73]. gradient [Ber70b]. graduate [Fai74]. Grammar [MR73, Mac73, Rau78]. Graph [JR75, Rej72, Squ70]. Graphic [WM77, Gra70b, SW75]. graphical [ADT67, Ree68, Ree71, Rin79]. Graphics [Jon79, KRS78, Les72, Rul68b, Dig76c, Hol77, Hun76, Kno72, KW75, RP74, TS73, War79, Wol78b]. Graphs [JR76, Coc80, Lau80, Tan80b]. gravimeter [KRB77]. gravimetrii [SZ80]. gravitational [Rap66b]. gravity [Bea75, BC72a, Plo75, Plo77, Rap66a, Rap66b, Rap66d, RZB77]. Greenberg [Jun68]. Grenzschichten [Rot71]. grey [MT75]. grid [Plo77]. gridded [Har68a]. Grit [Lew80d, Lew80c]. ground [Joy77, Joy78]. Group [AB69a, Mac67, MV66, Sal70, She70a, Wol68a, Sta74]. groups [BKW74]. Growth [Abr65, All66b, Cha66, VR67, SK69]. Growth-Curves [All66b]. GSPC [WPK78]. GT [Hol77]. GT-44 [Hol77]. guidance [Zor68]. Guide [Bee79e, Bri67, Bru66, Elt66, Fly73, McC62, McC72a, McC72b, McC74e, Pac69, Ryd3 , San70, Sch75, ZN79a, ZN79b, Ari76, AD73, AK77, Ano68b, Con68c, Con72b, Con73d, Con76a, Con76b, Con76c, Con77b, Con79a, CW78b, Dig75c, Dig75g, Dig76c, Dig77f, Dig78a, Dig78c, Dig79a, Dig80b, Dig80c, Dig80d, Dig80f, Dey76, FMC78, Fin77, Fis76, Flo78b, Fra79, Gen80c, Got73, Hew79a, Haa69b, HBE80, Hei72b, dPW80, Hig79b, Hon75a, Hon72b, Hon72a, Hon79b, Hug77, Hun76, Int66f, Int66g, Int66h, Int66i, Int67c, Int68g, Int8 , Int70b, Int70c, Int71d, Int72h, Int72k, Int72n, Int73b, Int74d, Int75e, Int75c, Int75d, Int75a, Jac73c, Key73c, Lew79a, Lew80a, LJ71b, Mac68a, Mac68b, MHH71b, McC61, McC64b, McC65a, McC65b, McC67c, McC74c, MS79, Mt.79, New73, PJT76b, PC64, Pol65a, Pol65b, Pri77a]. guide [Pri77b, Rei72b, Shn77, SF72, The68, Tho71, Uni69b, Uni73, Uni74a, Uni79, Uni80b, Bee80c]. guide/release [Dig80b, Dig80d]. guiding [CS77a]. gyros [Wil72a].

H [Bar74, Bee71b, Bee71d, CCHT67a, CCHT67b, Cod67, Din72, Hil70, Hui65, Int68b, Jun68, Jun69, Kar77, Bee71a, Dea71, Dea77, Int70c, Int72d, Int72k, Int72n, Int73b, Int74c, Int74b, Int74d, Int78a, Int75a, MS71, Ste72a, Tym70, Whi72, Xer70a, Xer71b]. H. [Jun69]. H1 [Fia73]. H1640 [Hon70b]. Haar [Rej72]. Hammer [Les73]. Hamming [Kar77]. hand [Kra72b]. Handbook [AM72, Bre79a, Com80b, Sta75, Dig72a, Day72c]. handled [Ber70b]. Handler [DW77]. Handling [Rey77, Har65b, Wol68a]. handwriting [Smi71e]. Handy [Gen80c]. Hans [Din72]. Hardware [Amk73, Dic74a, NC75, Dav74, Gel69, Kei69, lAL72]. harmonic [Har68a, Tay80]. Harris [Joh66b]. harvesting [New67, NS69]. Having [Ful72, Ste72b]. Headers [Bee70a]. Heap [Kah80a, Kah80b]. Heat [Gro73b, Bar80a, Bec72, Gro71, Hol67, Hol68, Pin80, Saw62, vNS63, Bec72]. heating [FMC78]. heavy [Nor66]. Height [Han72a, Han72b, Kot72, PC67, vM79]. held [Lew79b, Weg64]. Herbert [VP80a, VP80b]. here [Smi70c]. Hermitian [Mue66]. Hessenberg [Ste76b, Ste76c]. hexagonal [RG68]. Hidden [Wil72d, Wil72b, Wil72c, Wil72e]. Hidden-Line [Wil72d, Wil72b, Wil72c, Wil72e]. hierarchic [Mod74]. Hierarchical [BB77b, Kal72a]. High [Gea65, Mt.79, BLY70, FPB72, Fin72a, Hei74, Mag71, O’D74, PC67, Sch79c, Sch80g, SDH74]. high-level [BLY70, FPB72, Fin72a]. high-lying [SDH74]. Higher [Sch72a]. highway [Sik71]. Hilfe [Nie72a, Nie72b]. Hindemith [Hun74]. histograms [vM78b]. History [ACM78, Bac78b, Bac79, Hei64, Hei66, Knu62, McC78c, Sam69, Gri78, Ros78]. Hit [Tho72a]. Hochschulrechenzentrums [jH78]. Hoffmann [HD78a]. Höhere [KKU78, Sie74]. Hollerith [Int60a]. holocaust [Smi70h, Smi73c]. homme [Gro68a, Gro70]. homme-machine [Gro68a, Gro70]. Homogeneity [Dah78]. homogeneous [Hol67, Hol68]. Homogenisation [Sch80f]. Homogenisation/blending [Sch80f]. Honeywell [GS71, Mar78a]. horizontal [Bom67]. host [Sof80]. hour [New72]. hourly [PC67, RPE79]. Household [Sad72]. Householder [Mue66]. housing [BM74]. Houston [IEE79]. HP [Hew71, Hew76b]. HQR3 [Ste76b, Ste76c]. hsieh [Ano72d]. HSPF [Joh80, JID80]. HSSR [SDZ80b]. hsu [pC79, lH80, jT79]. Hu [Yoh72]. Hu-Tucker [Yoh72]. hub [KM73b, KM77a, KM77b, Kat77]. hub-shroud [KM73b, KM77a, KM77b, Kat77]. Hull [GS79]. Human [vO78, Leo74]. Humans [DP74a, DP73, DP77, PD76, PD80a, PD80b]. Hybrid [Høj69, Høj70]. hydraulic [TH62]. hydraulics [Sol64]. Hydro [Ano72a, Lel74]. HYDRODAT [BC79, BC77]. hydrodynamic [GS70]. hydrogeochemical [SDZ80a, ZSD80]. Hydrologic [Hyd66b, Pat74]. hydrological [BC77, BC79, Joh80, JID80]. hyperbolic [ES75]. Hypergeometric [Fre73]. hypernuclei [Gaj66]. hyperthermal [Erd80].

I.B.M. [BDI72, Kra74, MS71]. i.e [ATW77, Bit75, Dre67, Dre75b, Uni78]. I.T.S. [Wit74]. I/O [Bai72b]. IAS [Dig78a]. IAS/RSX [Dig78a]. iazyk [Kar76]. iazyke [SM76a]. IBJOB [Ber64]. IBM [Boy75, Fel75b, Rin77, Abr65, AS71, All66a, All66b, ADT67, And64a, Ano64, Bac54, Bac56, Bar66, Bar73c, Ben69, Ber70a, Bla69, BC67, Bri68b, Bro71a, Bro73a, CC70, Cla78, CCHT67a, CCHT67b, Cle68, Cod67, CJM67, CS61, Com80a, Dah66, DS66, DS67a, Dea71, Dea77, Den71, Ent63, Fas67, FF75, Fox64, Fro63, Gar65, Gen66a, Goo64, Gre75, HW67, Hec63, HG66, Hil70, Hor65, Hou62, HN58, Hug69, Int57b, Int57a, Int57c, Int57d, Int58, Int59a, Int59d, Int59e, Int59f, Int60b, Int60c, Int61b, Int61c, Int62b, Owe62, Int62d, Int62c, Int63d, Int63o, Int63p, Int63f, Int63e, Int63g, Int63h, Int67a, Int64b, Int64g, Int64c, Int64d, Int64e, Int65c, Int65f, Int65d, Int65e, Int66b, Int66c, Int66d, Int66e, Int66f, Int66g, Int66h, Int66i, Int66a, Int67b, Int68c, Int68d, Int68e, Int8 , Int68h, Int68i, Int68a, Int68k, Int68j]. IBM [Int68b, Int69b, Int69c, Int70b, Int70c, Int71e, Int71c, Int71h, Int71f, Int71g, Int71b, Int71i, Int71j, Int71d, Int72h, Int72f, Int72g, Int72k, Int72m, Int72n, Int72j, Int73a, Int73b, Int74c, Int74b, Int74d, Int74f, Int74g, Int75e, Int75c, Int75f, Int75d, Int76, Int77a, Int78a, Int79, Ins70, Int75a, Irv60, Izz73, Jam70, JcK73, KPG63, KW71, Kan71, KF72, KS68, KS70, KR69, Kuo73, Kuo74, Lea67, Lee67b, Lee72, Lou74, Man69, Man74, MI64, Mar66, Moo60, Mor70, New73, Nir69, PG65, PB67, PH71, PG66, PD67, Plu61, Plu63, Plu65, PN68c, PN68a, PN68b, Pri69, Ren65, Rey68, Rob68, SD66, SD67, Saw62, Sch62, She59, Shu75, Sil61, Squ70, Sun73, Swi64a, Swi64b, Swi64c, SS66, TC70, TB65, TC75, Wei66a, Wei66b]. IBM-1620 [Hor65]. IBM-709 [Saw62]. IBM-7090 [Gen66a]. IBM/ [KW71]. IBM/360 [Rin77]. IBM/370 [Rin77]. IBM/ASSEMBLY [KF72]. IBM360 [Bac72]. IBMAP [Ber64]. IBP [Gof74]. IBSYS [Ber64, Int66a]. IC [Sta69]. IC-4000 [Sta69]. icin [Yur76]. ICL [REC75]. ideas [Owe65b]. identification [Gaj66, JV67a, JV67b, JV68, TI72]. identificazione [BT76a]. Identifiers [LV73, Par78, Sco77b]. identify [Fos74, Tro64]. Identifying [LaM72]. IEEE [Bid79]. IEEE/KCS [Bid79]. IFIP [FGR72]. IFOR [Rad70]. If’s [DW71]. IFTRAN [Bez75, Eld77]. II [Hun74, Joh66b, Van68a, Wes69, Abr65, Ack64, AK77, Arn65, Aye63, Bac78b, Bac79, BS64, But66, CC70, Car69, Cla73b, Cle68, CJM67, Dat67a, Dah66, DS66, DS67a, Dey76, Dra64, Fan65, Fas67, FPB72, Fin72a, Gar65, Gol66, Goo64, Har64a, Har65a, Har65c, HM62c, HM62b, Int58, Int60a, Int63b, Int63f, Int63e, Int63g, Int63h, Int64g, Int64c, Int64e, Int71c, Int71h, Int72e, Int74c, Int74b, Ins70, JV67a, Joh76, Jon64, KPG63, Kau65, KR69, Lec66b, Lec68, MW71b, MA78, Mer60b, Mit65, Nor66, O’D65, Ost64, PG65, PG66, PD67, Poo62, PK67, PK69, Rab62, Rey68, Sci64, Sci65, SD66, SD67, Sch78b, Shu69, Swe67, Swi64c, TB65, Uni75b, Wal68, Wer65, Wis69]. II-D [Cle68, Fas67]. III [Bac78b, Bac79, CF60, Gen77b, Gen77a, KC73, PNK65b, Rad88, SMD71]. Iktisat [Yur76]. ILLIAC [MM75, Mil73b, Mil75, Ste75b]. ILLOD [NL71, Nak77]. ILLOD- [Nak77, NL71]. illustrating [Sch69]. illustrations [Int63l, Joh65a, Joh65b, Joh76]. illustrative [Har69]. im [Sto76]. imaged [Per80]. images [BM80]. IMB [Int68f]. Impact [Cob75, LR77]. implement [Bid79]. Implementation [AB69a, Ban78b, Ban78c, Ban78a, Blo68, Com78, Fel79, Håv71, Jon79, Kah80a, Kah80b, KRS78, KGY80, NO72, NO75, Nik78, NSB71, Pra75, Ryd72, WDT76, Fri70, His75, LC75, Lit74, Lum77, Mis78a, O’N74, PMBK80, Pot66, Sle75, Sli71, WPK78, Wri66]. Implementations [BP78]. implemented [Pri69]. implicants [DD68]. Implicit [SS73, Nav78]. Improved [Com78, HKK72, SK80, HM62b, Sal71a]. Improvement [PH63, Dem69, Low76]. In-Core [REC75]. Inc. [Ame78d]. incidence [Zaa69]. Includes [Ano69a, Sal70]. Including [CS76, FS78, HM64, Jaf79, SM66a, SM66b]. inclusions [Cam77]. incluye [LP79]. Incluyendo [FS80]. incomplete [Ske79]. incore [Bob70]. incorporating [Ell78, FH71]. Incorporation [Coh66]. increasing [BT76b]. Incremental [RCM66, Sch62]. independence [Sch72c]. Independent [Bee79a, Bee79c, Bee79b, Bee79d, Bee80b, KG72, BN76, CW78b, Coc80, Hew80b, Kir79, MA78, Sle75, War79, WD75, Bee77b]. index [Ott78, Sho80]. indexes [Cla73b]. indices [dC73, GKB74, Kli73, MG71]. individualized [GO75a, GO75b]. individuals [Hil79c]. induced [Kol74, Moo76]. induction [Bol76]. inductive [Zav73]. Industrial [Ano72c, Ins76a, Ins76b, IA78, CK80]. Industrie [WS71]. industry [WS71]. inequalities [Les73, Win74]. inequality [MS79]. Inference [Boh75]. infiltration [San74]. infinite [DP76b]. infinity [Rob79]. inflow [WG75]. Informatik [jH78]. Informatik-rechner [jH78]. Information [AI78, Bee71d, Int75a, Tou70, Bur67, Bur68b, Bur70b, Gol68a, Gol68b, Int57c, Int61a, Int63k, Int63b, Int63n, Sha71a, Sha71b, Sid72a, Sid72b, FGR72]. informational [Gor64]. Informatique [Dav74, Ano76a, Ano79, CR69, CR73, Cha79b]. infrared [MW71a, MW71b]. ing [NBH70a, NBH70b]. Ingenieria [FS80, MD66a, MD73]. Ingenieurstudenten [And79a, And79b]. inhalation [Won67]. Initial [ANS69c, ANS69a, ANS69b, Ano69b, CR74]. initial-value [CR74]. Initiation [Ano76a, Ano79, Dub77, ES74a, Gro68a, Gro70, Wel70b, CR69, CR73, Cha79b, Phe79]. inlet [Dic74b]. Innforing [Tju68]. INP3F [New73]. Input [Eld70, Fer60, Int60a, TR77, Wan78, Yar62, Ano72c, Bai63, Cle66, Coh66, Fla71, Hyd66a, Has67, Ins76a, Ins76b, Mye73, Rus79, SD66, Sha77, Tay76, Uni68a].

Input-Output [Fer60, Ano72c, Has67, Sha77, Uni68a]. Input/Output [TR77, Coh66, Hyd66a, Ins76a, Ins76b, Tay76]. insights [Gin78b]. instalado [Ano70d, Ins64, Ins74]. Installation [Bee79e, Bee80c, HR73, Dig80b, Dig80d, Int72a, Int72b, Int72f, Int72g, Int74b]. instant [Con75d]. Instantaneous [PD67]. Institute [Ame78c, Ame78d, Axf72, Cad71, McC64a]. Instruction [CS73a, DDM^{+}75, Mar80, Plu64, Spi80, Ano72b, CS68, CS72, FMC78, Hed77, Int63l, Int63m, Int63c, Int63i, Int68f, IBM68, Int79, Mis78a, Pen68, Plu65, PN68b, Tho71, Tho72b, Wid79, Wit79b, Wit79a, Wit79c, Wit79d]. Instructional [Ben69, Dun69a, BP74, CS61, Dun69b, FL76, Hal65, HDBP68, KP70a, KP70b, Lov75, RS80, Tro64]. Instructions [Swi64a, Rob67b]. Instructor [Col78b, Gro73c, HD78a, Mar77b, McC74c, Nic75c, Nic80c, PD80b, Spe69a, Spe77c, BK75, Key73c, LJ71b]. instructors [Shn77]. Instrumenting [LS75]. Instruments [Mor79, Wed75, Wil72a]. Integer [Ono79a, Ono79b, RH76, Sca71]. Integral [Lin72, Pie73]. Integrals [Ein72, DP76b, FP73, McC78d]. integrand [DP76b]. integrate [Fic71]. Integrated [SW79, Bra76, Lar63a, Mac70b]. Integration [Ant72, Håv71, Pat73a, And73, Hae77, Win74]. Intel [Sch79b]. Intellectual [Wag70]. intensities [KC60]. intensity [MT75, Pol78]. Interaction [NC76, DO79, Las71, Wil76c]. interactions [Tro66a, vM78b]. Interactive [ASH73, Boi74, GP73, Gom79, Hol77, Les72, Rad70, RP74, Rul68b, San73, SW74, WDT76, Bas80, Bri68a, Con76c, Con79a, Cla78, FL76, Gil76, Gra70a, Gre79, Int72a, Int72g, Int75b, Int75e, Kaz78, LaP72, Maz78, Moo77, Ree68, Ree71, SW75, WS73]. interchange [MS66]. INTERCOM [Con76c]. Interdata [Gro73d]. Interest [Smi70f]. Interface [BHY80, Sta74, Boy75, Cle70, MS77a, MS78, Sof80, TRW73b, TRW73d, TC75, Uni68a]. interfaces [Kal72b]. Interfacing [SD72]. interferometry [Var77]. interindustry [CB69]. intermediate [LO77]. Internal [DPR70, Mal77, MM65, DS67b, Erd80]. International [IEE79]. interpolating [Sca70]. Interpolation [Aki74, Dur80, Gaf77, Lag74, ClW78, Gaf79]. interpret [vM76]. interpretation [RPE79]. Interpretationen [Rin77]. interpreter [Blo71, Hea68b, Mar66, WS73]. Interpreteur [BM73]. Interpreting [CZ72]. ‘Interpretive’ [TR77]. Interval [BS75, Gin78a, Jet79, Pat73a, RJAS78, RR73b, Yoh78, Yoh79a, Yoh80]. Intervall [Ben80]. Intervall-Fortran-Präcompiler [Ben80]. Intervallarithmetik [Neh74]. Introduccion [Car78a, Cou75, DM72b, FS80, dMdF73, Wei73]. Introducing [Smi78, BC70]. Introduction [Ano78c, Ano80b, Bee75, BCE77, Bla67, Bla68a, Bla68b, Bla71, BC72b, Car68, CW71, Car69, Car74a, Car79a, Car79b, Car79c, Chi73, CM79, kC80, CA78, Con79e, DW71, Dic74a, Dic68, DS72, DM66a, DM67, Din73, Dun67, Dun69b, Dun69a, DLS79, FS78, FB69, Gen70b, Gra79, Gre77, HRH76, HRH78, Har70, Hol70, Hol71, Hui65, HD78b, Ise78, IJ79, Jaf78, Jam73b, Kal71, Kal72c, Kan77, Kar77, KC72, Key73a, Key73c, Key73b, KH75, Kre66a, LK74, LP71, Lim78, Lue66, MW69, MS69, Mau72a, McC67a, McD80, Mes73, MS73d, MS75b, Nak68, Nir69, O’b75, Per75, Pet74, Plu64, PC78b, Pra65, Que71, RR73a, Rul66b, Sch73, Tur73, Vin75, Wal72, Weg64, Bar70, BS73a, BW78b, Cle68, CS62, Dun80, Edw69, ER79, Fis71, Int77b, Int78d, Jaf72, Jaf79, Jam78, MG68]. introduction [MGL73, Mur80, MS70b, MS70c, MS75c, MS75d, Nyd68, Obr71, Oll71, Rau68, SS68a, San73, SB78, Shu69, SM66a, SM66b, Tho72b, WB71, Wri77a, Wri77b, Wri77c, Wri77d, Wu73a, DM66b]. Introductory [Int59c, Jac73a, Kir73, MS74b, Pen70, AHP77, Bau79, Coo76b, Dun75b, ES78, Her74, Jac73c, Jac73b, Kha76, Kha77, New75, Nut78, OR77, Pin73, Shn77, Sol78]. Introduzione [LMP77]. invalid [BH73a]. Invariant [Ste78a, Ste78b, McC71]. Inventarisierung [Sto76]. inventory [Hyd66b]. Inversion [DA68, Mur71, Fit74, Kar73, Mar71]. invert [Pow69]. investigation [Bec73, Mag71, Spa75b]. investigations [Ham79a]. involved [Wil76c]. involving [CK80, McC78d, Ste72a, Wie75]. ionization [Kol74]. IPI [Fis71]. irasa [Kor77]. IRIS [Int75a]. IRIS/TCDMS [Int75a]. irradiated [U.S61]. irradiation [LH65]. irredundant [Coc80]. irregularly [Jam66a]. ISO [Ano77c]. Isolation [Wag70]. isometric [Wra70]. isotopic [Smi79]. Issues [VP80b, VP80a]. Iteration [GF65, Nik78, Mat89]. Iterative [Kau69, KGY80]. ITPACK [KGY80]. ITRAN [Mil73a]. IV [And79a, And79b, Bri67, CCHT67a, CCHT67b, Cod67, Dig80a, Die72, DM66b, Ehr72, Elt66, Hen67, Hil70, Int68b, Joh66b, Kar77, Mee72, Mil73b, Mil75, MM75, Nie72a, Nie72b, Ree73, Ree75, Spe80d, SG78, Whi68, Wra70, Con79e, Ari76, Abr65, AS71, AM72, All66a, All66b, And64b, And66, And73, Ano64, Ano68c, Ano70f, Ano70g, Ano72b, Ano72e, Ano73, Ano75b, Ano75c, Ano75d, Ant77, Ars64, Art75, Ban75, Bar80a, BM74, Bar71a, Bar66, BP74, BD71, Bee75, BS64, Ben69, Ber70a, Ber64, BA72, BC72a, Bez73, Bis75, Bit75, Bla67, Bla68a, Bla68b, Bla69, Bla71, BY73, Bog74, Bog80, BC67, BCS68, BK75, Bre79a, BDI72, Bru66, Bur68a, Bur69, Cad71, Cad79, Cal69a, Cal78, CS71a, CS71b, CS71c, CS71d, CS75, CB69, CCL69]. IV [Car68, CW71, CW72, CW73b, CW73a, CW76, CW77, CW78a, CW79, dC73, Car69, Cha66, Chi73, CK80, CPR75, Cla68, Clo72, Col78a, Col78b, CS68, CS72, CS76, Cou76, CDG68, CDG70, CDG73, CDG80a, Dig68, Dig70, Dig71a, Dig71b, Dig71c, Dig72c, Dig74, Dig75b, Dig75a, Dig75c, Dig75e, Dig75g, Dig77e, Dig77f, Dig78a, Dig78b, Dig78c, Dig79a, Dig80b, Dig80c, Dat75, Dat77a, Dat77b, Dat67b, Dah66, Dah78, DW70, DvC69, DW71, Leu79a, Leu79b, Dea71, Dea77, Dem69, Den71, DS72, DM66a, DM67, DM72b, Doc72, Doc76, Doc79, DM72c, DR70, Dre67, DB68, DB69, Dre70, DB70, Dre72, Dre75a, Dre75b, Dun67, Dun69b, Dun69a, Ele68, Edw69, Edw73, Ent80a, EP67, ESD68, EKM74, Far66, Far74, FMC78, FPB72, Fin72a, FP75, Fin68, Fin72b, Fin72d, Fin72f, Fin72e, Fin72c, Fin77, Fla71]. IV [For71, For74, For75, Fox67, FGH75, FGH80a, FGH80b, Gra70b, Gen66b, Gen73, Gen77b, Gen69, GKB74, Gar71, Gen66a, Gol66, Gol65a, Gol76, Got72, Got73, GO75a, GO75b, Gue73a, Hew74, Hew79c, Hew80b, Hal69, Ham69, HRH76, HRH78, Han74b, Han75, HW67, Har68a, Har64a, Har65a, Har71, Har73, HV74, HDBP68, HD75, HG66, Hei70, Hem70, HPB73, Her69, Hir73, Hob67, Hog70, Hol70, Hol71, Hol67, Hol68, Hon73b, Hon74, Hor68, Int63o, Int63p, Int67a, Int64d, Int65c, Int65f, Int65d, Int65e, Int66b, Int66c, Int66d, Int66e, Int66f, Int66g, Int66h, Int66i, Int66a, Int67b, Int68d, Int68e, Int68f, Int8 , IBM68, Int68h, Int68i, Int68a, Int68k, Int68j, Int69b, Int69c, Int70b, Int70c, Int71e, Int71c, Int71h, Int71f, Int71g, Int71b, Int71i, Int71j, Int72b, Int72c, Int72d, Int72e, Int72h, Int72k, Int72m, Int72n, Int72j, Int73a]. IV [Int73b, Int74c, Int74b, Int74d, Int74f, Int74g, Int75c, Int75f, Int75d, Int76, Int77a, Int79, Int75a, Izz73, Jac73a, Jac73c, Jac73b, Jaf72, Jaf78, Jaf79, Jah80, Jam66a, Jam70, JcK73, Jam75, JV67b, JV68, Joh74, JCMS76, JCMS77a, JCMS77b, JMG77, Kah66, Kan77, Kar73, KS68, KS70, Ken74, KS75a, Ker72, Key73a, Key73c, Key73b, KQS74, Kle68, Kle69, Kle77, Kli70, Kru67, Kru68, Jac78, Lam71b, Lam71c, LG73, Lam74, Lam77, LG78, Lam78, Lan72a, Lan72b, Lar63a, LBG66, Lea64, Lec66b, Lec68, LW66, Lee67b, Lee69, Lee72, LB68, LH65, LB70, LGF75, dMdF73, Lou74, Lue66, LJ71b, LJ71a, Mas62, MW69, MV66, MGL73, Man72b, MG70, Mat72a, MA78, MG71, May73a, MW75, MT75, McC70a, McC70b, McC74a, McC78b, McC69a, McC65a, McC67c]. IV [McC72a, McC73, McC78e, MK70, McG67, McL73, Mei68, Mei69, Mes74, MS71, MS66, Mon77b, Mon79, MM69, Mos78, Mou70, Mt.79, Mul68a, Mul68b, Mul80b, Mur71, MS68, MS70b, MS70c, MS73b, MS73c, MS73e, MS75c, MS75b, MS75d, MSS78b, NM70, NM78, Nav78, New73, NL75, Nol71, Nyd68, O’D74, OLS66, Obr71, Oja70, OG69, Org66a, Org72, OM74, Osi75, Osi77, Owe79, Par70, Pat74, PB73b, Pen68, Per77, PB67, Per72a, Per72b, PH71, Phe76, Phi71a, Phi71b, PMBK80, PJT76b, PJT76a, PC78b, PC64, Pol65a, Pol65b, Pre70, Pri69, Pri75, Raj77, Ral71b, Rey69, RR70, RG77, RMM69, Ros73, Rot71, Rul66a, Rul68a, RFP73, Sta69, Sys73a, Sys73b, SG67, Oli71, Sal70, Sal77b, Sal78, ST73a, San74, Sas74b, SK69, Sch68, Sch74, See75]. IV [Kan68, Sel72, SST72, SGM^{+}67, Sha76, She78a, SR73, SDH74, Shu75, Sid72a, Sid72b, Sik71, Sil71, SS76, SM73a, Smi66, SH78, Sou67, Sou68, Spa75b, Spe69a, Spe69b, Spe66a, Squ70, SH67, Ste72a, SM66a, Ste74, SP70, Ste75b, SD73, SM72c, SM76b, SM73b, Swa72, TRW73a, TRW73b, TRW73c, TRW73d, Tay76, Tay80, Tho71, Tok68, TW71, Tom71, Tri79, Tri73, Tug75, Tur68, Tur69a, Tur69b, Tur73, Tym68, Tym70, Uni69c, Uni68b, Van68a, VV66, Vas72, Vic70b, Vic73, Vic77, Vic78a, Vic78b, VG77, VR67, Vow74, Vow77, Vow78, Wah68, WMM72, War69, Wat75, Wei75, Wil72a, Wil80a, Wil80b, Wil77b, Wil77c, Wil77d, Wit79b, Wit79a, Wit79c, Wit79d, WM77, Wol78b, WM9 , Wu73b, Wu77a, Wu77b, Xer70a, Xer70b, Xer70c, Xer71b, Xer71a, Xer74a, Xer74b, Xer75a]. IV [Xer75b, YS75, ZT76, Zor68, Zwa80, Wu73c]. IV-F [Sik71]. IV-H [Xer70a, Xer71b]. IV-PLUS [Dig75c, Dig78b, Dig78c, Dig79a]. IV-Programm [Die72]. IV-Rechenprogramm [Rot71]. IV. [Dig72b, Gen66a]. IV/CDC3300 [Mei68]. IVF [Bai62]. IVPS [Ske79]. IVTRAN [PJ75].

J [FJA80a, FJA80b, Gar72, Gar74, Jun68, Tam66]. J2 [Hyd66a, Hyd66b]. J6 [Lew73, Wil72d, Wil72e]. jack [McL73]. James [Ree75, Hen67]. January [Ano77a, Con77a]. jazik [Bit75]. jazyku [Ham79b]. JCL [êR76]. jezyku [ATW77, Bad77, BF72]. ji [Jam75]. Joan [May73b, Van68b]. job [BD80a, BCE77]. Jobstream [HPLG79]. John [Ree75, Joh66a]. Joint [IAAA57, Int75a, JV67a, JV67b]. Jolly [Whi71b, Whi71a]. JOSS [Gra70a]. JOSTRAN [Gra70a]. journal [Con77a, Con80a]. Jovial [SF72]. ju [Ano72d]. July [Eva72a, Eva72b, TB65, Weg64, Ari76]. June [ACM78, Fin72f, Lew79b, Ste72b]. junior [Zor68].

K-7 [Sta65]. K. [Bar74]. kagakuno [Nis78]. KANS [MM69]. Karzanov [Ham79a]. Kay [Fri70]. KCS [Bid79]. kemiai [ZSF78]. Kent [Eva72a, Eva72b]. Ketnyelvu [Kor77]. Keys [All66a, dlB59, Pan70]. keyword [New73]. keyword-labelled [New73]. kind [Sca71]. kindergarten [Zor68]. kinds [FM76]. kinematic [Phi67, Tro66a]. kinetic [MA78]. Kingston [Kau65]. klassov [AE79]. knapsack [MT78]. Knuth [Hoa72]. Knuth’s [Hoa73]. Kokanee [AS71]. Kolmogorov [Pom74]. kommentariiami [DG78]. kompressiblen [Rot71]. kompyuto [JcK73]. Koppelanordnungen [RS69]. Kopplung [jH78]. Koz [KC73]. Kriegspiel [BW75]. KRONOS [Con71c, Con73f, Con73g, Con74]. Kunzi [Din72, Jun69]. Kurs [Cal78]. Kursmaterialien [Tec72]. Kurtosis [MZ75]. KWIKR8 [ESD68].

L [CJ77, Hui65, Pic66, Rob79, BF79]. L-A-S [BF79]. L. [Joh66b]. l/s [Pic66]. L205 [Hyd66b]. L208 [Hyd66a]. labelled [New73]. Laboratories [Con66a]. laboratorio [SS68b]. Laboratory [Jon79, SDZ80b, ZSD80]. Labs [JMG77]. LALR [Rau78]. Lament [Fel76a]. Lamprecht [Bar71a]. Land [Wil74, MM69]. Langage [Lam71b, Lam71c, Ano76a, BM73, CR69, CR73, Cha79b, Gro68a, Gro70, LG73, LG78, Lap78, Lev71, Tho72c, Uni68b, Wel70b]. langages [DT74]. Language [ANS78, Ano68c, Ano77b, Ano77c, Ano78d, Ben69, BF79, BW64, BF71, Cor79, GHG60, Gum77, Han60, HM80, Hui65, Int71b, Kuo73, MS74a, Nag80b, OFP78, Par75, Sch72a, Ste75b, Sun73, UK74, Wal72, Woo77a, Woo77b, Ame78a, Ame78c, Ame78b, Ame78d, Ano68b, Ano70f, Ano72b, Ano72e, Ano73, Ano75c, Ano80a, Bar61, Bec72, BLY70, Con68c, Con74, CS62, Cor61, Dig71b, Dig75d, Dig77d, Dig78b, Dig79b, Dig80e, Ele68, Edu70, For70, FP73, Ful73, Gra70b, Gen70a, Gea78, Gil60, Gor64, Gre79, Gri78, Hei74, Int63o, Int63p, Int67a, Int64d, Int65f, Int66b, Int66c, Int66d, Int66e, Int66a, Int68a, Int68j, Int69b, Int69c, Int71e, Int71f, Int71g, Int72j, Int73a, Int74g, Int76, Int77a, Kha77, Kno75a, Kno75b, Kuo74, Lea70, LK74, Lea64, Lee77, Lue66, LO77, Lyn63, Mas60, Man64, MH75b, Mis78a, MT72]. language [MK73, O’D65, O’N74, Org61b, Per77, Pot66, Pow74, PK67, PK69, Rab62, Ros78, Sak64a, Sak64b, Sak65, Sak70, San73, Sch67, Sch79c, Sch80g, Sof80, Spa75b, Spe9 , Ste74, Sto80, Tob65, Uni69a, Uni72, Uni80a, Wil69, Wri66, Xer70c, Xer75a, Xer76a, Gal78, Org61a]. Language-Oriented [UK74]. Languages [ACM78, Els73, Loe74, Mcg80, Nau75, Nic75a, PC78a, PRO80, Pet74, Pra75, Ros66, Ros72, Sam69, San70, Sch73, Tuc77, BS73a, Dig72a, GB76, Gri78, Har69, Kan71, Kee75, Lee74c, Mat72b, Mis78a, Moo69, Nut78, Obr70, Obr71, Rod76, Ros78, Sli71, Slo68, SR76, Sol78]. Laplace [DA68, Mur71]. LARC [Axf72]. Large [Geo80, KGY80, REC75, Wan78, Boy75, Cla73b, DO79, Kar73, Kno75a, Kno75b, TC75, Wag80b, ZT76]. Latent [ST73c, ST73b]. latitude [MM69]. lattice [Pin80]. lattices [DS67b]. Lawrence [SDZ80b]. layer [And73, Epp74]. learner [Bur70a, Hur77, HF78]. learning [Can77, CS71e, CS77c, Hon73b, Int72l, Kal72a, O’D74, Sch79c, Sch80g, Tip76, Wil76c]. Least [AK78, ZN79a, ZN79b, Bar79a, Bom67, BML62, Tay80]. least-squares [Bar79a, BML62]. Lebanon [Mt.79]. Lecture [Fox75, Fox78b, Hal65]. Lectures [Sri69, DEN79, Roh73]. legal [BB71]. Legendre [Gav76, Pie74]. Lehmer [Kir79]. Lehrbuch [Mac70a]. Leistungsfähigkeit [RS69]. Leitfaden [Rin77, Jun69]. Length [All66a, dlB59, Dah66, Swi64b, Swi64c, Art75, Han75, Krá72a, Mac69, Swi64a, Tay76, Var77]. Length-Frequency [Dah66, Mac69]. length-weight [Swi64a]. Lenguaje [FGH80b, Cou75, Mer77, Wei73]. Leontief [CCL69]. Letter [Bem61, Bus67, Elk65, Har65c, Owe65a, Owe65b, Kni76b]. Level [Hon77a, Sch72a, Tym70, Ano78a, Bar77a, Bar61, BLY70, Cla73b, FPB72, Fin72a, Fri75a, Hei74, Hon75a, Hon75b, Hon77b, Hon79b, Low76, Sid72a, Sid72b, Spe79, Spe80c, Spe84]. levels [SDH74, SMM65]. Lexical [Fel76b]. Li [CL80, Ste72a]. Liberated [Bac78a].

Libraries [FHS7x, Hil70, Ker80, Cla73b, FF75, Hon75e, Hon77a, Int71c, Int71h]. Library [AI78, Ano80c, CGG^{+}91, CCHT67a, Cod67, GMPW79, Int60a, Int74c, Int80b, Kuk66, Kuk67, Lov75, Sou67, Sou68, WLO76, Xer71c, Xer73, Ano67, Ano70a, BH73a, BD71, Con75b, Con75c, Con79d, CJ77, CCHT67b, Dat73, Feu77, FS76, FHS78, Fox78a, Gil77a, Gil77b, Hog72, Int63g, Int63h, Int64c, Int64d, Int65d, Int66a, Int68k, Int72e, Int72m, Int74b, Int75f, Int80a, Jos78, Kau65, KA71, LT75, Lou67, Nag78, OK72, Spe73a, TD78, Uni69c]. Lie [AB69a, BK72]. life [Int64a, Int65a, Int78b, Int70a, McL73]. light [Gro73a, Gro73c, Nor66]. Like [BCKT79, KMC72, Ste75b, Kro75]. likelihood [BY73, MI80]. limbajul [CJ78]. Limited [Bar72a, Ost64]. Line [Eld70, Gen69, Sad72, TT80, Wil72d, Wil72e, Wil72b, DP76a, MB68a, MB68b, Sel72, Squ70, Sto76, Whi76, Wil72c, Dat77a]. lineal [Oli71]. Linear [Abd80, AK78, Bur73b, DZ78, Fia73, Hab72, Kas74, KGY80, Kub73, LP73, Law78, LHKK79a, LHKK79b, Mol72c, Mol72a, PG65, RR73b, Smi72b, UK74, Wol73, Zoh80, And73, BK77, Bar79a, BY73, CGH75, CJM67, CR71, Dav72b, Duf77b, Duf77a, Duf80, Edw76a, JCMS76, Law77, Les73, Lyt75, Mar71, MU75, Nut76, PNK65a, PNK65b, Pic66, Pow68, Rei72a, Rob79, SA74, Spä79a, SD72, Bur76, Weg66, Zim69, ZT76, BF79]. lineare [Kas74]. lineinykh [SZ80]. Lines [Ber76]. linguagem [Cad71]. Linguaggi [Amb65]. linguaggio [Sic74]. Linkage [Com80a]. linking [BA73, Lyt75, WM60]. LISP [CZ72, Ina80a, Ina80b, Mat72b, McC78c, NSB71, Pit79, Fat78, Mau72a]. List [AI78, Bee71c, BW64, GHG60, Han60, Moo71, LaP72, Rit68, SS68a, Ske79, Ste76a, LC75]. List-Processing [GHG60, Moo71]. listing [HM62c, Lou67, Whi71b]. Listings [Ano74b, Bee80d, Ano74c]. literacy [CS77a].

Livermore [SDZ80b]. livestock [BM74]. Ljubjana [FGR72]. LLLSRT [SDZ80b]. llS [Col76]. LLSS01 [ZN79a, ZN79b]. LM [Int68k, Int75f]. LM3 [Int74c, Int74d]. Load [Sil61, Ack64, Col75]. load-and-go [Col75]. load-time [Ack64]. loaded [Cam77]. Loader [Int60a, Boa69, WM60]. loading [BA73, PNK65a, PNK65b]. Local [Aki74, DS75, McC71, Sal70]. location [Jac78]. locations [MM69]. Loesungen [Pau71a, Pau71b]. Log [Hab72, BY73]. Log-Linear [Hab72, BY73]. logging [New72, New76]. Logic [Int68b, Lew80d, Gil60, Lar67c, Lew80c, McC68a, Wei65]. logica [Cad79]. Logical [IR78]. LOGPLAN [New76]. logs [Sco78]. London [Ree73, Weg64, Day72c]. longitude [MM69]. look [Smi70g]. Loop [DFO79, Hoa72, Hoa73, Bra77, Cul80, EKM74, MR67, Sol64]. Loops [BCKT79, DH79, Gen75a, Gen75b, Gen75c, Bak68]. l’Ordinateur [IBM58]. Lösung [Rin77]. Low [Bar61, Jon79, PC67]. Low-level [Bar61]. Low-Performance [Jon79]. lower [Sei72]. LPI [Rit68]. LPSUB [Dav72b]. LPTOR [JCMS76]. LR [Pag74b]. LRLTRAN [MHM^{+}68, Zwa75]. LSD-1 [Les72]. Lucy [Gar74, May73b, Van68b]. Ludeman [Ber70b]. lying [SDH74].

M [Joh66a, Whi71b, Whi71a, NBH70a, NBH70b]. M1 [Cha71a]. M6 [Hal72a]. M6800 [Cau78]. M77 [Fri80]. MA28 [Duf77b, Duf77a, Duf80]. MAC [Ano68b, Con68c]. Machine [BSK67, Bee79a, Bee79c, Bee79b, Bee79d, Bee80b, GM73, Kir79, REC75, ABB^{+}74, BN76, Bur68a, CW78b, CF71, Fat78, FS76, Gro68a, Gro70, Gro73d, Hug78, Kei69, New72, Sle75, WD75, Bee77b]. Machine-Independent [Bee79a, Bee79c, Bee79b, Bee79d, Bee80b, Kir79, BN76, Bee77b]. machinery [Sin78]. machines [Lin71, Mon78, New67, NS69].

MACLISP [Ste77]. Macro [Hal72a, Nag80b, Ell78, Rob67b, Rob67b]. MACRO-FROTRAN [Rob67b]. macro-instructions [Rob67b]. Macro-Oriented [Nag80b]. macrogenerator [HN70]. macroprocessor [Mac71]. macros [Par77, Joh72]. MACSYMA [Lew79b, Lan80]. MAD [Bry75]. made [May72]. Maechanik [Jun69]. MAGEN [Fis78]. magnetic [And73, KC60, vM78a]. magnetism [Gro73a, Gro73c]. Main [Høj70, Sha65, Taj65]. mainframes [McC79]. maintaining [Feu77]. Maintenance [Kah80a, Kah80b, Ho73, Mor79]. Major [GM73, Lea67]. majority [Van73a]. making [Hou62, Lil68, Mac67, Owe79, SDH74]. male [Cla73a]. Management [DW77, MS74b, O’b75, SM75, AM79, CK80, HK75, Hou62, Obr70, Obr71, Sof80, VP80b, VP80a]. manager [Con73d, Con76a]. Managerial [Per72a, Per72b]. Manchester [Bar72a]. Mandy [SMD71]. Manipulating [Rei79, CCL69, RP74]. Manipulation [Bee79a, Bee79c, Bee79b, Bee79d, Bee80b, BBB64, FP73, Hal71, Han75, Ins76a, Ins76b, Kee75, Lew63, Mor75, Ost64, Owe65a, Poo62, Pyl62, Smi63a, Tob65]. manipulations [BM80]. manipulative [Car66]. manipulator [Zim69]. Mann [DB73]. Manova [RST78]. MANTIS [ASH73]. Manual [Ano80c, Bro71c, Bro73b, CGG^{+}91, Cra80, Flo78a, Hec63, IBM56, Int59c, Int59e, Int68b, LF78, MS74b, NL71, AR72, Ano70a, Ano70b, Ano70f, Ano70g, Ano73, Ano75c, Ano75d, Ano75e, Ano80a, Ano80d, Bus68, BP74, Bla68b, BK75, Bur67, Bur68b, Bur70b, Bur73b, Bur73a, Bur74, Bur78, Con62a, Con62c, Con62b, Con62d, Con64c, Con64a, Con64d, Con65, Con66b, Con67b, Con67c, Con68b, Con68a, Con69d, Con69c, Con69b, Con69a, Con70, Con71b, Con71c, Con71e, Con71f, Con72a, Con73c, Con73e, Con73b, Con73f, Con73g, Con74, Con75a, Con75d, Con75e, Con75f, Con78a, Con78b, Con79b, Con79c, Con79d, Con80b, Con80c, Cal69b, Cal69c, Cal69d, Cal69e, Car74b, CS73a, Coc60, Col78b, Dig64, Dig68, Dig70, Dig71b, Dig72b, Dig72c, Dig74, Dig75b, Dig75a, Dig75d, Dig75e, Dig75f, Dig76b, Dig76a, Dig77a, Dig77c, Dig77b, Dig77d, Dig77e, Dig78b, Dig79b, Dig80e, Dat67a]. manual [Dat75, Dat77a, Dat77b, Dat67b, DM66b, Dun74, Ele68, Ent80a, FB79, FLM70, Fri71b, Fri73, FLM74, FL74, Fri80, Gra70b, Gen67, Gen66b, Gen77b, Gen77a, Gen66c, Gen69, Gea78, Gre79, Gro73c, Hew71, Hew74, Hew76b, Hew76a, Hew79b, Hew79c, Hew80a, Han74a, HDBP68, HD78a, Hon70a, Hon70b, Hon75d, Hon79a, Int57c, Int58, Int59a, Int59b, Int59d, Int59f, Int60b, Int60c, Int61a, Int61b, Int61c, Int61d, Int61e, Int61f, Int62d, Int62c, Int63k, Int63b, Int63n, Int64e, Int74f, Int78a, Int66j, Int75a, JSW77b, Joh80, JID80, JCMS76, JCMS77a, JCMS77b, JMG77, KP70a, KP70b, Key73c, Lea70, Lar67a, Lar67b, LM70, LJ71b, Mas71, MI80, Mar77b, MS69, MH73, McC70b, McC74c, Mei68, Mer74, Mer58a, Mic79b, Moc71b, MS73e, MS75d, Nat70a, Nat70b, Nat72, NCR69, Nic78, Nag78, Nag80a, Nak77, Nic75c, Nic80c]. manual [PD80b, Pat74, Per72b, Raw77, Rid79, Ser71, Sta69, Sci64, Sci65, Sys73a, Sys73b, Sak64a, Sak64b, Sak65, Sak70, SA74, SM72b, Sch67, Sho76a, Sho76b, Sik71, Sof80, Spe69a, Spe77c, Spe66b, Spe66a, Squ70, Ste76a, TRW73b, TRW73d, Tan78b, Tan80a, Tok68, TB65, Tur69a, Tur69b, Uni69a, Uni69c, Uni71, Uni80c, Van73b, Wes69, Wu73c, Wu77b, Xer70a, Xer70b, Xer70c, Xer71b, Xer71c, Xer71a, Xer73, Xer74a, Xer74b, Xer75a, Xer75b, Xer75c, Xer76a, Xer76b, Yor64, BLF80, Bar72a]. Manuel [Ars64, HV66, IBM58, Rou75, Uni75b]. Manufacturers [GM73]. Map [Day63, Edw76b, LML69, OLS66]. Maple [CGG^{+}91]. MAPLIB [Sch72b]. Mapping [HKK72, MT75, Ren65, Wri77a, Wri77b, Wri77c, Wri77d]. MAPPROJ2 [Edw76b]. maps [Dic74b, McM67, Moc69, Moc70, Moc71a, Moc71b, Nor66, Tur68]. March [AB66b, AB66e, Ame66c, Ano77a]. Marching [Ban78b, Ban78c, Ban78a]. marine [HW67]. Mario [Joh66a]. Mark [Ano80c, Gen77b, Gen77a]. market [Hol80]. Markov [Agh77, Kru67, Kru68]. Martin [BK75]. Marwick [VP80a, VP80b]. MASH [Sad72]. Masinnye [FMM80]. Masked [Wat73a, Wat73b]. mass [Fro63, Smi79]. MASTER [Con69a, Con70, Con71b, Con73c, Con73b]. MASTER/MSOS [Con70, Con71b, Con73c, Con73b]. masters [Spe78a]. Match [KPG63, DS76]. Match-Coeff [KPG63]. matching [BD80b, Lau80]. Matematicas [DG70]. matematiceskich [FMM80]. Material [Sch72b, HM64, Int72b, Int72f, Int72g, Int74b, Soy71]. materials [Tro64]. Math [Kuk66, Kuk67]. Mathematical [Air77, BCS68, Din72, FMM77, FMM80, IBM54, LP73, Wat76, Wil74, Con75b, Con75c, Con79d, CJ77, Dig72a, HW67, Int65d, Int71c, Int71h, Int72m, KTZ68, KTZ71, LT75, SWL68]. Mathematics [DG67, DG68, Int80b, Jam73b, Jun68, LJ71b, LJ71a, Nau75, CG73, CK80, CDH75, Ham74, Sci69]. Mathematik [Jun69]. Mathematische [RW69]. mathematischen [Jun69, KTZ67]. Matrices [Nik78, DO79, Mar71, Mue66, Rey69]. Matrix [Gow75, Kub73, Mol71, Mol72b, Nie72c, ST73c, Ste76b, Ste76c, Ste78a, Ste78b, BS64, DM66c, Fit74, Kar73, Lee74a, McM67, MU75, PNK65a, PNK65b, Pow69, Rei72a]. matter [Oer71]. Maturation [PD67]. MAX [Gil70]. maximal [Ham79a]. maximax [DS75]. maximum [Bar79a, BY73, Lau80, MI80]. McCalla [Nak68]. McCluskey [Cam65]. McCormick [Joh66a]. McCracken [Bru66, Kre66b, Ree73, Ung69, Bri67, Elt66, Kar77]. MCP [Hug77]. me [Nor0 ]. Mean [Spi72, Rap66c]. measurable [Joh74]. Measure [WD79]. measured [SMM65]. Measurement [KBC^{+}73, Rip77, U.S61]. Measurements [KBC^{+}74, MC64, McC69b]. Measures [MZ75, Fin77, Veg74]. mechanical [SW75]. Mechanics [Gro73b, Gro71]. mechanism [Khu68]. mechanisms [Leo74, O’K64]. Mechanization [GIB65]. mechanized [NS69]. media [Phi71a, Phi71b]. mediante [DM72b]. medical [Boa69]. Medium [Bur73b, Bur73a, vM75]. medium-scale [vM75]. medizinischer [Fri75b]. Meek [FJA80a, FJA80b]. Meeting [Ano77a]. megoldasara [ZSF78]. Mehtods [Pen70]. member [Van73a]. members [O’D65]. memento [Ano78b]. memory [Gel69, Hug78, Huy77, Spe66a]. men [Ano72d]. MESS [Hol80]. messages [Int74c]. mesures [Fer63]. metal [Mei78]. Metalanguage [Bur65]. metals [Pin80]. meteorizacão [dC73]. Meteorological [Cra76]. meter [Kra74, Low76]. Method [DZ78, ES74b, HKK72, Pag74a, RK73, Yoh72, dL78, Bar77c, BW78a, BY73, CL70, Die68, Din69, Gre75, HV74, Hat78, Her64, LP74, Lil71, Lil68, Lit74, Mik73, Mue66, PNK65a, PNK65b, Ree68, Ree71, Ree72, TI72, Vas72, Vic70a]. méthode [Fer63]. Methoden [Bra75a, Bra75b, Jun69, KTZ67, RW69, Neh74]. methodischer [Rin77]. Methodist [IEE75]. Methodology [OR77, GKB74, McC74b]. Methods [AI80, Bee75, Dic68, Din72, DM72c, FMM77, FMM80, Håv71, Joh66b, Joh66a, Kar77, KGY80, Kre66b, Kuo72, McC67a, Nak68, Ree75, Ban75, BD80b, CW73a, Der64, Fle72, Ham74, Har63, Har64b, JSW67, JSW77a, JSW77b, KTZ68, KTZ71, MS64, MD64, MD66b, MD68, Pra65, Ree73, SSS77, vNS63, Veg71, Wis69, Wit79b, Wit79a, Wit79c, Wit79d, Kre66a]. Methuselah [Van68a]. Metodika [SZ80]. Metodos [JSW70, MD66a, MD73]. metody [FMM80]. metric [SF75]. Metrology [Smi72c]. Mg [Ste72a]. MHD [RCL75]. MI [Sta60, Ste60a, Ste60b]. Miami [Tou70]. Michael [BK75]. Michigan [Uni0 ]. Micro [IEE79, TG79b, JM76]. micro/mini [JM76]. microcoding [LM76]. microcomputer [Hea79]. microprocessor [Bra76, Cau78]. microprocessors [Sch79b]. micros [McC79]. microscopic [vM78b]. Microsimulation [Sad72]. mid [KM73b]. mid-channel [KM73b]. midchannel [KM77a, KM77b, Kat77]. MIFLL [WM60]. Migrating [AS71]. Military [U.S78]. MILTIQUAL [BY73]. Milton [Fry71]. MIMD [LB80]. mineralogica [dC73]. minerals [Sch80f]. Mini [IEE79, JM76, Jos78, Kno75a, Kno75b, MH75b, Nag80a, Sle75, Kno75a, Kno75b]. mini-computer [Jos78, Sle75]. MINI-EXPLOR [Kno75a, Kno75b]. mini-language [MH75b]. Minicomputer [GB76, Jon79, BN76, Gro73d, Hin76, RW77, San78, Upc72]. MINIFOR [Upc72]. Minimal [Pag74a, Whi72]. minimax [Dun79, DS75, Gro73d]. Minimization [MM73a, MM73b, Fle72, Lar67c, MS79, Pow70b, Pow70c, Van73b]. Minimum [Yoh72, Lil71, Wor69]. minis [McC79]. MiNnesota [FLM70, Fri73, FL74, LM70, FLM74, Fri80]. Miscellanea [Bry75, LaM72]. misspelt [Par78, Sco77b]. Mixed [MS74a, NC76, KM73b, KM77a, KM77b, Kat77]. mixed-flow [KM73b, KM77a, KM77b, Kat77]. mixtures [Mac69]. ML [Par77]. ML/I [Par77]. MMLE3 [MI80]. MNF [DH78, FLM70, Fri71b, Fri73, FLM74, FL74, HD78a, LM70]. mnogomernoi [AE79]. Mod [Int72e, Int74c, Int74b, Int71c, Int71h]. mode [BC67, MI75b, Par70]. Model [Ben69, Int74f, Rad79, Rad88, RST78, Squ70, Whi71a, Bas80, BCS68, Con72b, Con73e, HG66, Leo74, Ler72, MSNC61, Mel62, MSR66, New76, Pat67, PMBK80, Rey68, SYR77, Whi71b, Gro73d, Pat74]. modele [Ray63]. modeled [Per77]. Modeling [LL65, Rog80]. modelirovaniia [AE79]. Modelle [WS71]. Modellerstellung [Sch77]. modelling [Dey76]. models [Con71d, Con72a, Con72b, Con73d, Con73e, Con76c, CCL69, FB73, Kru68, OG69, WS71]. Modem [Swa72]. Modern [Fel79, Mul68a, Mul68b, Mul80b, SA74, Vic78a, HS69, Vic70b, Vic78b]. moderno [Vic73, Vic77]. Modifications [CF71, Int60a, AJ69, Ver59]. modified [Bid79, Den71, Kal72a, Lil71, Sca71, TI72]. modify [Thr79]. Modular [Cha76, Ken74, BK75, CP80, KS75a, Lum77, Pat77]. Modularized [LB77]. modulation [Ber77]. module [RCL75]. modules [Bra79]. molecules [SG67, SG69, VP75]. momenta [CM66]. momentum [RV78]. MONECS [BW78b]. MONECS/FORTRAN [BW78b]. money [Smi70f]. Monica [IEE78]. Monitor [Jon64, Squ70, Int60c, Mer60a]. monitoring [Tan80b]. Monte [Feh68, Fer63]. Monte-Carlo [Feh68, Fer63]. Monte-Carlo/optimal-shift [Feh68]. Morphological [Ste74]. morphometric [Wal68]. Mortality [PD67]. Mortem [KL64, NY78, Cor60]. MOSES [SG67]. Mössbauer [Kan68]. most [Har69]. motion [O’K64]. Motivation [Ryd72]. motor [Bol76]. Motorola [Cau78]. MOTUS [SS68b]. MOUSE4 [Com78]. movement [ST73a]. Moving [Gen78]. MP [Bre78a, Bre79b, Mac71]. MP/1 [Mac71]. MPLIB [LT75]. MPS [SD72]. MPS/ [SD72]. MS [Con75a]. MSOS [Con70, Con71b, Con73c, Con73b, Con75a].

Msufor [Bai72a]. Mt. [Mt.79]. MTS [Car68, CW71, CW72, CW73b, CW76, CW77, CW78a, CW79, Uni0 ]. mud [Sco78]. Müller [Bar74, Bar77c, BW78a]. Multi [Boh75, Gut76a, Ano64, EKM74, Fri71a, Hol80, MV66, Mot66, ST73a, Sch68, Spe66b, Spe9 , Van73a]. Multi-dimensional [Gut76a]. multi-group [MV66]. multi-loop [EKM74]. multi-market [Hol80]. multi-member [Van73a]. multi-processing [Ano64]. multi-processor [Spe66b, Spe9 ]. multi-programming [Sch68]. multi-region [MV66]. multi-stage [Fri71a, Mot66]. multi-storey [ST73a]. Multi-Variable [Boh75]. Multichannel [Rob67a, vM79]. multicrop [Sin78]. Multics [Mar78a, Col80b, Hon77b, Hon79b]. Multidimensional [Hab73]. Multifit [MA78]. multilingual [Fos73]. Multimodal [Žil78]. multiphase [JV67a, JV67b, JV68]. Multiple [Bre75, Bre78a, Bre78b, Bre79b, BHY80, VS80, Yoh79a, Bac72, Bre76a, Die76, ESD68, PT67, PT69, TC70]. Multiple-Precision [Bre75, Bre78a, Bre78b, Bre79b, Bre76a]. Multiplicators [ADG70]. Multiply [BS73b]. Multiply-Restricted [BS73b]. multipoint [LP74]. Multiprozessor [Sch80e]. Multiprozessor-Systeme [Sch80e]. Multisystem [KRS78]. Multivariance [Fin68, Fin72d, Fin72f, Fin72e, Fin77]. Multivariate [Dem69, MZ75, DS66, Fin68, Fin72d, Fin72f, Fin72e, Fin77, Hil79c, Jon64, Nut76, SDZ80a, WS71]. musical [Hun74]. mutual [DS76]. Myers [FJA80a, FJA80b].

n [Pic66, LC75, Ste72a]. nach [Sch77]. Nag [Ano80c]. Name [ABB^{+}74, Bee70c]. NAMER [Sid72a, Sid72b]. nao [Ano75b]. National [ANS78, Bar72a, Cad71, DH78, HD78a, U.S78]. Natural [Nau75, FB69, PJT76b, PJT76a, U.S61].

NBS [Ano74b, Ano74c, Ano74a, HP74]. NCR [Nat73]. near [Pol78]. NEAT [PC71a, PC71b]. need [SJ62a, SJ62b, SJ63]. Negative [Gar65, OG69]. nel [Rid67, Rid78]. nelineinykh [SZ80]. nella [Rid67, Rid78]. Neoclassical [OR75]. nerka [AS71]. Nested [Gen75a, Gen75b, Gen75c]. nesting [Han67, Kau69]. net [Bar80a]. Network [Kri71, DP74b, Sho76b, Tho68]. Networks [LL65, NL71, Osy76, Lar67c, Nak77, Sho76a]. Neumann [Bac78a]. neutron [KC60, McG67, Sta65, Zaa69]. neutrons [Fer63]. Newcomb [GIB65]. Newsletter [Ano77a]. NEWSUMT [MS79]. Newton [Fle72, TI72]. ni [Mur70]. Niggli [VV66]. nilpotent [BKW74]. nine [Tam66]. nine-j [Tam66]. NLPROG [Wor69]. NMR [vM77]. No [Ano77c, Bar72a, Bar73a, Bar80b, Law79, Ree73, Ree75, Ano70d, Gal75, Gol68a, Gol68b, Ins64, Ins74, PV74, PT68]. no-recoil [PV74]. NOAA [RPE79]. nominal [BY73]. Non [BB77b, CC74, Gow73, Hil73, LS76, Wil75, Wor76a, Bar73a, Day79, O’D65, PG67, Pow68, SS78a, Wie75, BC72b, Day72b, TR77, Wor76b]. non-central [Wie75]. Non-Deterministic [CC74]. Non-Hierarchical [BB77b]. Non-‘Interpretive’ [TR77]. non-linear [Pow68]. Non-Numeric [BC72b]. Non-Numerical [Gow73, Hil73, Wil75, Bar73a, Day79, Day72b]. Non-Rational [LS76]. non-rolling [PG67]. non-smoothing [SS78a]. Non-technical [Wor76a, Wor76b]. non-uniform [O’D65]. Nonlinear [MC80a, MC80b, MC80c, Mot79, Bar77c, BW78a, Hol80, Joy77, Joy78, LP74, McC69a, Nav78, Pow70a, RG76]. nonparametrical [Veg71]. nonquantitative [BC67]. Nonstandard [Yoh79b, Coh66, Cra79]. NOR [NL71]. NOR-B [NL71]. Norm [Blu7 , Blu78]. Normal [Don73a, Fry71, HK72, MI75b, Rap66b, Sco78]. normality [Mar78a, Pre70]. normalnykh [AE79]. norms [VV66]. Note [BB71, Boy75, CPR75, Lil71, Ver65, Bur72, Kor79]. Notes [Bee76, Ben78, Cou76, Mer60b, Swi72, Dig80b, Dig80d, Fox75, Fox78b]. Notions [Boi75]. NOVA [Dat77a]. NOVA-LINE [Dat77a]. November [IEE75, IEE79, U.S78]. nozzle [Tho65]. NSF [McC64a]. NSPIV [She78e, She78b, She78d, She78c]. nuclear [Ful74, MSNC61, Mel62, MSR66, Phi67, SMD71, SR73, SDH74, U.S61]. nucleon [SDH74]. Number [Ano77a, BS73b, Int68b, Kru69, KMC72, NO75, Sch79a, Sch62, Ano70c, Bre74, CGH75, Edg79, Fel75b, Gro69, Int68k, Int75f, Kir79, Krá72a, MB68a, MB68b, Mil68, NO72, Ove72, Pay70, U.S61]. Number-writing [Sch62]. Numbers [ADG70, Rei79, WB65, Int70c, Int72k, Int74c, Int74d, Int74f, Smi70j, Tri79]. Numeric [BC72b]. numerica [Sic74]. Numerical [ACM79, AG80, Ban75, Blu77, BKK^{+}80, Der64, Dic68, DM72c, DA68, Gin78a, Gow73, Ham79a, Har63, Har64b, HKK72, HH79a, Hil73, Joh66b, Joh66a, Kar77, Kre66b, Kre66a, KTZ68, KTZ71, Kuo72, McC67a, MS64, MD64, MD66b, MD68, MC80a, MC80b, MC80c, Nak68, Pat73a, Pen70, Ree72, Ree75, Sch76, UK74, Wil75, And73, BB72, BB77a, Bar73a, Bri68a, CW73a, Day79, FS76, Hol80, Jet74, JSW67, JSW77a, JSW77b, Pra65, Ree68, Ree71, Rin79, SSS77, Day72b, Din72, Ree73]. numerico [BB78, Rid67, Rid78]. numericos [JSW70, MD66a, MD73]. Numerische [Jun69, KTZ67]. numerology [Smi70g]. NUSC [Con77c]. nyelv [LV77]. nyelvu [Kor77]. nyumon [Nih69].

O [Ree75, Ano70d, Ins64, Ins74, Bai72b, Ste72a]. Ober [Ant72]. Object [LM69, SK80, Dig71a, Dig72b, Dig72c, Gea65]. objective [Pow70b, Pow70c]. Objects [LT73, Spa80]. obliczeniowa [ATW77]. obratnykh [SZ80]. observations [Kra74, KRB77, Swe67]. observed [Bra72a, OG69]. ocean [Fro63, MI75b]. oceanographic [KW75, Swe67]. October [IEE78, Kie66]. ODES [Ske79]. OERs [ZD78]. off [Gal73, Sto76, WG75, Whi76]. off-design [Gal73, WG75]. off-line [Whi76]. off-line-Verfahren [Sto76]. office [MM69]. Oklahoma [Hed77]. OLB [Ste76a]. Old [Woo77a, Woo77b]. Olika [Hus76]. OLYMPUS [CR74]. On-Line [Eld70, Sad72, Sel72]. on-resonance [SDH74]. Oncorhynchus [AS71]. One [Boh75, MB68a, MB68b, NO75, Sch76, She70a, Tho72a, Žil78, Edu72c, HM62a, HM62c, HM62b, HM64, Kan79, MV66, Mon78, Rin79, Sal70, SW75, WCT68, Sli71]. One-Dimensional [Žil78, HM62a, HM62c, HM62b, HM64, MV66, Sal70, SW75]. One-Hit [Tho72a]. One-line [MB68a, MB68b]. One-Sided [Boh75]. one-way [Kan79]. Operating [Gle62, Int64b, Int68b, Con66b, CF71, Dig71c, Int67a, Int65f, Int65d, Int66f, Int66g, Int66h, Int66i, Int66a, Int70b, Int70c, Int71h, Int72k, Int72m, Int75c, KW71, Lar63a, Lee67b, New72, Raf79, WH73]. operation [Chi73, CK80, Int72n, Int73b]. Operations [KMC72, Sch72a, Spa75a, Spa75c, Spa79b, Ano70g, Ano75d, Int59b, Int63e, Int78a, New76, Sci65, Sin78, TG79b, Xer70c, Xer71b, Xer71a, Xer74a, Xer75b, Xer76b]. Operations-research-Software [Spa79b]. operativa [Rid67, Rid78]. Operator [Int59e]. Operators [GH72, GIB65]. OPSCAL [NL75]. optical [MSNC61, Mel62, MSR66]. Optimal [Gaf77, Gaf80, NL71, Ant77, BK77, Lyt75, Mue75, Nak77, NL75, Wei75]. optimal-shift [Feh68]. Optimierung [Jun69, KTZ67]. Optimisation [DFO79, Mid74]. optimising [SA73]. Optimization [BE69, Con77c, Dea71, Dea77, Din72, GMPW79, Hoa72, Hoa73, KM73c, Mer78b, Osy76, RK73, SK80, Sea79a, Sea79b, Wai74, Wol78a, Žil78, Con76d, Gil77a, Gil77b, Int78a, KTZ68, KTZ71, LT75, LF75, LM69, Mer78a, Mer78c, Met80, Mik73]. optimizations [Bar77a]. optimize [Van66]. Optimized [BN76]. optimizing [Sid72a, Sid72b]. Optimum [SS73]. optique [Ray63]. Order [Ano77c, Fry71, BK77, Col80a, Fic71, LP74, New73, Sca71]. ordered [Hil79c]. Ordering [Ste76b, Ste76c]. ordinal [BY73]. Ordinary [Kar77, KBC^{+}73, KBC^{+}74, Sch75, Bra72a, Car74b, Fic71, Pic66, Sca71]. Ordinateur [Uni75b, Phe76]. ordinateurs [Lev71]. Oregon [Int75a]. organ [Won67]. organigrammes [Boi75]. organisms [MG71]. organization [Huy77, Lum77, Yar62]. Organized [BJ74]. orientation [DS76, Smi67b]. Oriented [AG80, AW73a, Mid74, Nag80b, UK74, AW73b, ADT67, Cla68, Feu77, Int72l, LJ71b, LJ71a, Fai74]. Orion [TH64, Tay68]. ORSEF [Fri71a, Mot66]. ORSEF-2 [Fri71a]. ORSEF-3 [Fri71a]. Orthogonal [Gow75, Kas74]. Orthogonale [Kas74]. Orthographic [Wei66a, Wei66b]. orthotroper [Die72]. ORVAC [FH71]. ORVAC-CT [FH71]. Orwell [Pat67]. OS- [Rin77]. OS/3 [Spe80d]. OS/360 [Hil70, Kuk66, Kuk67]. OS/VS [Rin77]. OS/VS-Programmen [Rin77]. Osiemdziesiat [ATW77]. other [CS73a, CK80, Sig80, Thr79, Wil72a]. Ouchless [Bai72b]. outline [LP78]. outlined [CW63]. Outlines [Lip78]. Output [Bee70a, Fer60, Int60a, TR77, Ano72c, Bar75, Coh66, Hyd66a, Has67, Ins76a, Ins76b, Mer58b, Sha77, Tay76, Uni68a]. Overdetermined [Abd80]. Overhead [MP79, Mur77a]. overlaid [Cla73b]. overlay [Leu79a, Leu79b]. overview [ES78, Fra79, Har77]. oxidation [BP76]. oxide [BP76].

P [Din72, Jun69]. P1 [War69]. paced [Coo76b, Dig80a, DS72, Lin76, Mul80a]. Package [Bre75, Bre78a, Bre78b, Bre79b, BHY80, Gin78a, SM75, Wei66a, Wei66b, WLO76, Yoh78, Yoh79a, Ano70b, Ano70c, Ano75e, BLY70, Bre76a, BC70, Car74b, CR74, Dig75f, DS67a, Fis76, Flo78b, IA80, Lan80, Lou67, Lou74, MCB^{+}62, Ove72, Ros73, Rul68b, Sch79c, Sch80g, Sig80, Ste76a, Xer75c, Zor68, Kle78]. Packages [Yoh79b, BPW72, Boy75, Ent80a, Ent80b, O’D74, Sta60, TC75]. packed [RG68]. Packing [Coo72]. Page [Bee70a, Boy74b]. Page-On-Demand [Boy74b]. pages [Bar72a, Bar73a, Bar80b, Ree73, Ree75]. Paging [Mol72b, Mol71, SS68a]. Palm [Smi73b]. paper [Sol78]. para [dC73, Far74, Oli71, Wei73]. paralinear [BP76]. Parallel [Arm78, BCKT79, DFO79, Han72a, NC75, Sch72a, Han72b, Kro75]. parallèle [DT74]. Parallelism [Cob75, KBC^{+}73, KBC^{+}74, PJ75]. Paralyzer [PJ75]. PARAM [Jef77]. parameter [Ben78, MI80, Sei75]. Parameters [Kas74, SS73, Dic74b, Joh74, LTB80, Mal77, Win74]. parametric [McC71, WH73]. parity [SDH74]. PARMS [Thr79]. Parser [MS73a, Fri70, Rau78]. Parser-Generating [MS73a]. Parsers [Pag74b]. Part [Dey76, Edu72c, Lar73b, Lar63a, Nor66, Edu72f, Edu72e, Edu72b, Edu72a, Edu72d, KRB78, Lam71a, HS69]. Partial [Cha71a, Sch76, She78e, She78d, Spe80a, Ste79, SS79a, SS79b, SS79c, Car74b, She78b, She78c, Sid72a, Sid72b, SSS78, SS75a, SS75b, War75]. particle [Tro66a]. particles [Tro66b]. partir [Phe79]. Partition [JR76]. partitioned [Com80a]. partitioning [Kar73]. Partitions [BS73b]. PASCAL [KKU78, Bau79, Lju80, Nut78, Raw77, Moh77, SFIK79]. past [Lyt75]. past-optimal [Lyt75]. Path [DZ78, Zal73].

PATRICIA [Cla73b]. PATRICIA-II [Cla73b]. pattern [CF60, Mas62]. Patterns [BF71, JV67a, JV67b, JV68, Smi63b, SH78, Tay80, vM77]. Paul [Hun74]. PCFORT [CCN^{+}79]. PCODE [CNW80, CCN^{+}79]. PDP [CDGW76, Dig64, Dig68, Dig70, Dig71a, Dig71b, Dig71c, Dig72c, Dig75d, Dig77d, Dig79b, Les72, San78]. PDP-10 [Dig68, Dig70, Les72]. PDP-10/LSD-1 [Les72]. PDP-11 [CDGW76, Dig71a, Dig72c, Dig75d, Dig77d, Dig79b, San78]. PDP-11/ [San78]. PDP-15 [Dig71b, Dig71c]. PDP-8 [Dig64]. PDP10 [BH73a]. PDP9L [Bar72b]. PDR3057 [Pri77b]. Pearson [DM72a]. pedagogic [FK76]. pedagogical [SR76]. pela [dC73]. pelo [Cad71]. People [Did78]. Percentage [BR75b]. perception [Tan80c]. perform [Bre67, Gut79a, Gut79b, vM79]. Performance [CCHT67a, CCHT67b, HPLG79, Hil70, Jon79, Mor73, RT77, Gal73, Sch80a, Tho65, WG75]. Periodic [Wan78]. perspective [OT80]. Perspectives [Rip77]. PERT [Sho76a]. PERTNET [Sho76a]. Peter [FJA80a, FJA80b]. PFORT [Ryd74]. PFortran [WD75]. phase [Wei75, Wer65]. phenol [Nor66]. philosophy [MK73]. photogrammetry [But66]. photopeaks [CL80]. Phrase [Hou71]. physical [DS62, DS63, Int65b, Int78c, KW75, Tro64]. physicists [MT72]. Physics [Gro73b, Bor67, CK80, Gro71, Gro73a, Gro73c, Zim69]. Physictran [MT72]. Picard [Mat89]. PIDGIN [Ste72b]. pie [Dil79, Dil79]. p’ien [Ano72d]. pin [Zaa69]. pine [McL73]. Pivoting [She78e, She78d, She78b, She78c]. PL [FJA80a, FJA80b, Rin77, Ano69a, Ber70b, Bro75, BDI72, Car78b, Cha70, Cle70, DT74, ES78, Har69, Hil70, Int67c, Int68g, Kan71, Kuo74, LaP72, Lea67, Lea75, Lud69, MW69, Mat72b, Mee78d, Mee78a, Mee78b, Mee78c, Mee79, NS76, Per72a, Per72b, Rod76, Ros78, Sea80, Sun73, Tha77, Wei73, Kuo73]. PL-I [Mee79, Tha77]. PL/1 [FJA80a, FJA80b, Rin77, Bro75, BDI72, DT74, ES78, Har69, Lea67, Lea75, MW69, Per72a, Per72b, Wei73]. PL/1- [Rin77]. PL/I [Ano69a, Ber70b, Car78b, Cha70, Cle70, Hil70, Int67c, Int68g, Kan71, Kuo74, LaP72, Lud69, Mat72b, Mee78d, Mee78a, Mee78b, Mee78c, NS76, Rod76, Ros78, Sea80, Sun73, Kuo73]. PL1 [Som71]. Plaid [CS77b, DS77]. Plain [Moo71, The68]. plan [Smi67b]. planar [Cli74b, Cli74c, Per80]. planar-structural [Per80]. planar-valued [Cli74b, Cli74c]. Plane [GS79, Wei66a, Wei66b, MC70, O’K64, Rin79]. Planetary [GIB65]. Plankton [Cla75]. planning [New76]. plans [SW80]. plant [FH71, Fri71a, Mot66]. plate [Cam77, Rin79]. play [BB71]. player [Gil70]. Plot [Ste73, VAB62, Gof74, Joh65a, Joh65b, Joh76, RG77, RS72, Kan68, Uni74b, Wat73a, Wat73b]. Plots [Gen78, Lew73]. plotter [Ano70e, Sch62, Squ70]. Plotting [MP73, Wil80b, Wil72d, Wil72e, Wil72b, Dil79, Fox64, Hat78, Hol77, JK78, Kra74, LBG66, Nic78, Nor66, OLS66, RMM69, Squ70, Wil80a, Wil72c, Wri72]. Plus [Ken74, BK75, KS68, KS70, KS75a, Wer72, Dig75c, Dig78b, Dig78c, Sto80, Dig79a]. Pocket [AM72, Pac69, Hew79a, Hon75a]. Poets [CA78, Con79e]. POIF [Ste60a]. Point [Mal72, NC75, Rei79, VS80, Bid79, Mal70, McC71]. pointer [Gum77]. Points [BR75b, GS79, MP72, McC71]. POIS [Ste60a]. poisons [Zaa69]. Poisson [OG69]. POL [CW63]. Polarization [Bee80d, Ber77]. polarization-modulation [Ber77]. Policy [Ano69a]. POLISH [Dor79]. pollen [Squ70]. pollutant [Ott78]. polyatomic [SG67]. Polychoric [MH75a]. polycrystals [DS76]. Polygon [Sal77b, Sal78]. Polygonal [Sal77b, Sal78]. Polymodal [YS75]. Polynomial [JR75, JT72, Dif72, Vic70a]. Polynomials [BH1 ]. POP [MS74a]. POP-2 [MS74a].

POPSS [WH73]. Population [Cla75, Smi70h, Smi73c, SS66, PT73]. populations [Whi71b]. por [dC73]. PORT [Fox78a]. Portability [Air77, Bro77, Sab76]. Portability-Some [Sab76]. Portable [Bee80a, Blu7 , Blu78, FW78, Fel79, FHS7x, Hal69, Kru69, Poo74, RH76, Sch79a, WLO76, Yoh78, Yoh80, Coo76a, FHS78, Fox78a, WD75]. portions [Bai63]. Portland [Int75a]. position [Sol78]. Positive [Geo80, Fan65, Rei72a]. positive-definite [Rei72a]. possibilities [MT72]. Post [KL64, NY78, Cor60, LH65, WM60]. post-editor [WM60]. post-irradiation [LH65]. Post-Mortem [KL64, Cor60]. Posters [Smi73d]. potential [Cas62, Las71, RBp75]. pour [Aub76, Fer63, Ray63, Ric73]. powder [JV67a, JV67b, JV68, Smi63b, SH78, Tay80, vM76, vM77]. Powell [Wil74]. Power [BA72, Spi72, EP67, KRB78, Sal70]. powered [U.S61]. pp [Ano77a]. Pracniques [Fed63, Poo62]. Präcompiler [Ben80]. practica [Rom75]. Practical [CTC72, Cor73, Eng75, KC72, Mon77a, Rul80, San70, Key73a, Key73c, Key73b, Mon77b, Mon79, San73, TW71, Tom71]. practice [LF75, McA77a, McA77b, Per72a, Per72b, SW79]. Prager [Kre66a]. Praktika [DG78]. praktischen [Dre70]. pratique [DG75, HV66, Lap78, VHP69, VL72]. Praxis [Mac70a]. Pre [BJ77, Flo78b, Jet74, Oll71, BJ77]. Pre-Compiler [BJ77]. pre-processor [Flo78b, Oll71]. pre-processors [Jet74]. precipitation [Mit65, RPE79]. Precision [Bre75, Bre78a, Bre78b, Bre79b, BHY80, Jet79, Ono79a, Ono79b, Sch75, VS80, WLO76, Yoh79a, Bac72, Bre76a, FM76, KA71, Rei80, Sca71]. Precompiler [WLO76, Cra79, O’N74]. precompilers [Mal77]. predict [BM74, DS76, FMC78]. predicting [Gal73, WG75, Wil76c]. prediction [Bar79a, Lar69, Rap66c]. predictions [New75]. predictive [Nut76]. predstavliaiushchei [AE79]. preferential [Van73a]. Preliminary [Ber64, Int59e, Plo77, Con64c, Dun75b, Dun67, Fri71b, Pri77b]. Preparation [CLS64, Her70, BC77, BC79, Cra75]. prepare [Soy71]. preparing [CB69]. Preprocessor [Bur79, Com78, Gin78a, Ker4 , Ker75a, Ker75b, Ker79, Wil77a, AI79, Bod77, Ell78, Gal75, Jef77, Ker70, Lea80, Lum77, Mar78b, Mue75, Str78, Wag80b]. Preprocessors [MP79, Mei75b]. presentation [Squ70]. Presentations [Joh66b, Joh66a]. presents [Ent80b]. Press [Bar73a, Bar80b, Pat73b]. pressure [Fla71, Fro63, Rin79, RCL75]. pressurized [Pol78]. PRETTY [Bee80a]. Prettyprinter [Bee80a]. Price [Bar72a, Bar73a, Bar80b, Ree73, Ree75]. Prime [Pri77b, DD68]. Primer [Gre79, Ent63, Int57c, Int57d, Owe62, Int64f, Int78b, Int78c, Man69, Man72b, Man74, Nie68, Nie71, Org61a, Org61b, Org63, Org66a, Org66b, Sel72]. Primes [Smi70i]. Primitives [Bee79a, Bee79c, Bee79b, Bee79d, Bee80b]. Princeton [Ube76]. principal [Wah68]. Principios [SS78b]. Principles [AU77, De 72, Jam78, Lew80d, SS74, Wag80a, Lew80c]. printed [Kra72b, Par70]. Printer [TT80, Squ70]. Printing [Hab73, Fic73]. prior [Sid72a, Sid72b]. prisms [Plo75]. private [DP74b]. Probabilities [BR75a, Fre73, Fry71, Sti72]. Probability [Wri77a, Wri77b, Wri77c, Wri77d, McC78d]. probes [PB74a]. Problem [AK78, CS73b, Fen73, FK77a, Lew80c, MG70, Mou70, Spe77d, Ber70b, But66, Dav72b, DLS79, DS75, Elk65, FK77b, HJ74, Hal75, Int63m, Int68e, Man63, MT78, Pec77, Ree72, Rob79, WMM72, Weg66, Zak77, Eng75, Lew80d]. Problem-Solving [CS73b, Lew80c, Eng75, Lew80d]. problemas [LP79, dMdF73]. problemes [LPJ79b, Tel80]. problemorientierte [SG78]. Problems [AW73a, AW73b, DZ78, Fia73, Mer78b, SW80, Wal80a, Wal80b, Wol78a, ZN79a, ZN79b, Ano72e, Ban75, BD80b, Dun67, Dun69b, Dun69a, ES75, GS70, Hol80, LP74, Lep76, LP78, Mer78a, Mer78c, Org61a, Per72a, Per72b, PH77, Ree68, Ree71, RR73c, RR73d, SW75, Tea72, Zim69]. Proc [Eva72a, Eva72b]. Procedure [KL64, Lyt75, Mal77, McC71, Rus79, Sit78, VP76]. Procedures [Aki74, Bee71c, Int64b, SS73, Ano72c, Dem69, Int67a, Ins76a, Ins76b, IA78, Smi71d]. Proceedings [ACM79, Axf72, IEE78, Lew79b, FGR72, Tou70, Weg64]. procesamiento [SS78b]. process [Ano72c, Bar75, BD71, Ful73, Ins76a, Ins76b, Mei68, Mod74, SDZ80a, Smi79, Soy71, ZSD80]. Processing [AS71, Ben77, BW64, Car74a, Car79b, DFO79, Eld70, Fre76, GHG60, Han72a, Han60, HN58, Int75a, Moo71, NC75, Ree79, RH76, Rul68b, TS76, Ano64, AB68, AB69b, BLY70, Bur68a, Bur69, Bur67, Bur68b, Bur70b, Car79a, Car79c, Dar78, EB80, EH68, Eri75, FGR72, Gor64, Han72b, HBE80, Hei70, Hou62, Int57b, Int57c, Int58, Int59a, Int59d, Int59f, Owe62, Int63f, KM73a, Ker72, LaP72, MP65, Moo60, MK73, Mur77c, Mur77a, Mur77b, Mur80, Obr70, Pin73, Rit68, Rob62, Sch80f, SDZ80b, Smi66, Ste76a, SS74, Swe67, Wet79, Wet80, YP80]. Processor [BCKT79, Flo78a, Hal72a, Bai72a, Ber64, Cla68, Eri75, Flo78b, Gre79, Int71i, Int71j, Int71d, Int72b, Int72c, Int72f, Mac70b, Oll71, SS68a, Spe66b, Spe9 , Wal70, LC75]. Processor-N [LC75]. processors [Han78, Jet74]. produce [McM67]. product [Int72h, Int74c, Int74d, Int75a]. Production [Wil80b, Cha71c, Moo76, Wil80a]. Productions [Pag74b]. productivity [SYR77]. products [Bus67, Int75d]. produtos [dC73]. professional [SJ62a, SJ62b, SJ63]. Profile [SFIK79, Epp74, Tay80]. profiles [Tip76]. Progammation [IBM58]. Program [Ant80, Bak75, Bee70c, Bee76, Bee80a, Blu7 , Blu78, Cha66, Dah78, Day63, Ful74, GMPW79, Hoa72, Hoa73, Hog70, Int68b, Int70c, Int72k, Kli70, KC73, LaM72, Lov75, LF75, Mid74, Mor73, Mot79, NL71, PB67, Plu64, Rej72, Rip77, RT77, Sal77b, SFIK79, Sin73, TT80, VR67, Whi71a, WD79, Wil80b, Wil72d, Wil72e, Wil72b, WM77, YS75, AJ69, Agh77, AK77, Ant77, BK77, Bar80a, BM74, BP76, Bar77c, BW78a, Bar79a, Bar79b, Bar75, Bar66, BD71, Bea75, Bec72, BB71, BC72a, BY73, Bol76, Bom67, BC67, BCS68, Bor69, Bre79a, Bry75, BW75, BML62, BML64, Con71d, Con79c, Cam65, CB69, CCL69, Car77, Cha71c, CS73a, Cla73a, Coc80, CJM67, CS61, Col80b, Cra75, Cri77, Cse75, DW70, DS66, Leu79b, Dem69, Der64]. program [DS67b, Dic74b, Dif72, Dor79, Dun79, Edw76a, Epp74, Erd80, EP67, ESD68, EKM74, Fan65, FMC78, Feh68, Fin68, Fin72d, Fin72f, Fin72e, Fin77, Fis70, Fla71, FB73, Fos74, Fox64, Fox67, Fro63, Ful73, GKB74, Gal73, Gil77a, Gil77b, Gil76, Gof74, Goo64, Gro73d, Gus73, Gut79a, Gut79b, Hyd66a, Hyd66b, HW67, Har68a, Har80, Har73, Hat78, HG66, Hei70, Hem70, HM62a, HM62c, HM62b, HO64, Her64, HM64, dPW80, HW75, Hil79b, Hil79c, Hol80, Hun74, Int62d, Int63g, Int67a, Int64c, Int68k, Int72h, Int74c, Int74d, Int74f, Int75f, Int75d, Int78a, IA80, Jam66a, Jam73a, Joh80, JID80, Joh65a, Joh65b, Joh76, Joy77, KPG63, Kan79, Kar73, Kat68, KM73b, KM77a, KM77b, Kat77, KC60, Kot72, KR69, Kru68, KRB77, KRB78, Lal75, LTB80, Lar67c, LT75, Las71]. program [Lat79, Lee74a, LH65, Les73, LB70, LGF75, Lov68, Low76, Lyn63, Mac67, MI75a, MI80, MV66, MI64, Man63, Mar78a, MW71a, MW71b, MA78, Mau77, MG71, Maz77, MW75, MT75, McC69a, MM58, MC64, McC69b, MK70, McG76b, McM66, McM67, MC70, Mei78, MSNC61, Mel62, MSR66, Mer60a, Mer60b, Met80, Mik73, MU75, Mit65, MS79, Moo60, Mor79, MM69, Nak77, NM70, NM78, Nav78, New73, NS69, New72, New76, NL75, Nor66, NBH70a, NBH70b, Nut76, OLS66, Oja70, OT80, OG69, Ott78, Pan70, Par70, PV74, PG67, PNK65a, PNK65b, Per80, Phi67, PT67, PT68, PT69, PG66, Pin80, Plo77, PJT76b, PJT76a, Pol78, PS78, PC67, Pre70, RP74, Rap66a, Rap66b, Rap66c, Rap66d, RG68, Ree72, Ren65, Rey69, RR70, RCL75, Rob79, RMM69]. program [RSBR69, Rub69b, Rub69a, Rub69c, RB76a, RBK76, RB76b, RB76c, RZB77, Sal78, SD66, SD67, ST73a, San74, SK69, Sch72c, Sei75, She78a, SR73, SDH74, SDZ80a, SDZ80b, Shu75, Sid72a, Sid72b, Sik71, Sil61, SYR77, Sin78, SM73a, Smi63b, SH78, Sou71, Squ70, SMM65, Ste72a, Sti62, SG69, SW75, SM73b, Tan80c, TI72, TS73, Tay80, TH62, Tho65, Thr79, Tor69, Tri79, Tri73, Tro64, Tro66a, Tug75, VP76, Bur76, Van66, Van73b, Var77, Veg71, Veg74, Vic70a, VP75, Wag80b, Wah68, War69, WG75, Wat73a, Wat73b, Wat75, Weg66, Wei75, Wer65, Whi71b, Whi68, Wie75, Wil72a, WCT68, Wil80a, Wil65, Win74, Won67, Wra70, ZD78, Zoh72, vM76, vM77, vM78b, vM78a, vM79, Bee80d]. program-package [IA80]. Programacão [PC71a, CC70, Cad79, Ins64, Ins70, Ins74, Ano70d]. Programacion [FS80, For73, For79, McC63, McC73, McC74d, McC78e, McC78f, SS78b, Car78a, DM72b, DG70, Far74, LP79, MD66a, MD73, Oli71]. programado [SJ72]. programador [Lec68]. programare [Pet80]. Programarea [CJ78, Pet80]. programas [dC73, dMdF73]. Programe [HDN74]. programlama [Yur76]. Programm [Die72]. programma [SS68b]. Programmacão [PC71b]. programmacion [Gol76]. Programmatheque [Die76, É67]. Programmation [Die74b, Era77, LPJ79a, LPJ79b, Phe79, Ano79, BM73, BV74, Cha67, Die74a, DT74, HV66, HH80, HH77b, HH78, Lam74, Lam77, Lue66, RH78, Tho72c, VHP69, VL72, VG77]. programmazione [LMP77, PCR76]. Programmbeschreibung [Die72]. Programmdokumentation [Sto71]. Programme [Fer63, Fri75b, Ray63, Gaj66, Lou73, New67, Ric73, She70a, Ano68a]. Programmed [CS71e, CS77c, Fre74, Int63i, Mar80, Spi80, Baj72, Ben69, CS68, CS72, Dun69b, Int63l, Int63m, Int63c, Int68f, IBM68, Int79, Phi71a, Phi71b, Plu65, PN68b, Tho71, Tho72b]. Programmen [Jun69, KTZ67, Rin77, Rin77]. Programmentwicklung [SF76]. Programmer [Hig79b, IBM56, Int57d, Int59c, Owe62, Int72n, Int73b, LF78, Mau72a, Dig70, Dig71b, Dig72b, Dig72c, Dig74, Dig75b, Dig75a, Dig77a, Fra79, Got73, Hew71, Int57c, Int66f, Int66g, Int66h, Int66i, Int70b, Int70c, Int72h, Int72k, Int74d, Int75c, Int75a, Lec66a, Lec66b, Lew79a, Lit74, Mis78b, Mer58a, Pri77a, Pri77b, Rob67b, Shn76, Shu69, SJ62a, SJ62b, SJ63, Spe73a, Spe73b, Spe74a, Spe74b, Spe76a, Spe76b, Spe78b, Spe79, Spe80d, SF72, Uni78]. programmer-defined [Rob67b]. programmering [Tju68]. Programmers [BLF80, Kuo73, Lew80d, Sun73, Fis76, Int75a, Kuo74, Led75, Lew80c, Raw77, Spe66b, Spe66a, Spe69c, Tan80c, Tur69a]. programmer’smanual [Dig71a]. Programmes [Aub76, Hol67, MK68, Oer71]. Programmeur [IBM58, Ars64]. programmi [BG78, BT76a]. Programmieranleitung [Bar74]. Programmieren [Flo70b, GG72, Spi70, SR72, SR74, And79a, And79b].

Programmierkurs [Nie75]. Programmiersprache [Bar71a, Kle68, Kle69, Kle77, RS69, SG78]. Programmiersprachen [KKU78, Hig72, Win79]. Programmieruebungen [SE74]. Programmierung [Bra72b, Mue69, Sie74, Tec72, Bra72c, jSJ70, SS70]. Programmierungsanleitung [MS70a]. Programmierwerkzeuge [KP80]. Programming [ACM78, ACM79, ANS78, Ano78d, Arm78, Bac78a, BM79a, BM79b, Bau56, Bog74, Bog80, BGG78, Bro61, Bur72, Bur73b, Com69, Car79b, CW78b, CA78, Con79e, Cor73, CS76, DPR70, Den71, Dic68, Doc76, Don73b, DG67, Els73, Fia73, FS78, Got72, Gue73b, Gut75, Gut76b, Har70, HD75, Hen67, Her71, Her72b, HM80, Hog72, Hug69, HM77, HPR78, HD78b, HH79b, Int64e, Ise78, Jam70, Jn69, Joh66b, Ken74, KC72, KH75, Kre66a, LP73, Lan72a, Lan72b, Led75, LC78, LW66, Lip77, Lip78, Loe74, LR77, ML70b, ML70a, Man71, May73a, McC67a, McC72a, McC72b, McC74e, McC75, Mcg80, MO80, MR73, Mer79, Mot79, MS75b, Nau75, Nic75a, OR75, OFP78, Pet74, Pra75, Rad80, Ral71b, Ree75, Ree76, RR73b, Roh73, Ros66, Ros72, Sam69, San70]. Programming [Sas74a, SC79, Sha76, Sit78, Sol78, Spe69b, Spe80b, Sta75, SM64, SM68, SD73, Stu70b, Stu71, Tuc77, VP80b, Vei66, Ver65, Wag80a, Wal72, Wal75, Web78, Weg64, WDT76, Wil69, Zwa80, Ame78a, Ame78c, Ame78b, Ame78d, AHP77, AW73b, All75, And64a, And64b, And66, Ano68b, Ano72b, Ano72e, Ano80d, Arc76, Ayc80, Aye63, Bar70, Bar72a, BS73a, BW78b, BN76, Bez75, Bis75, Bla67, Bla68a, Bla68b, Bla69, Bla71, Blu70, BK75, Bur69, Con68c, Con71d, Con74, Cal69a, CS71a, CS71b, CS71c, CS71d, CS75, CG68, CW73a, Car69, Car79a, Car79c, CW75, CP80, CK80, Cle68, Coh74, Col78a, Col78b, CS61, Coo76b, CTC72, Cul80, Dig64, Dig68, DEN79, Dav72b, De 72, DS72, DM66a, DM67, Doc72, Doc79, DG68, Dun80, Dun75b, Dun74, DS75, Edw69, Ell80]. programming [Emb78, Ent63, ES78, ER79, FB79, Fis71, Fle70, Flo75, For71, For74, For78, For75, FB69, Fri75a, FK76, FK77a, FK77b, Ful77, Gar65, Gil60, GC67, Gol65a, Gol65b, GB76, Gor64, GO75a, GO75b, Gri78, Gro68b, Gue73a, Haa65, Haa69a, Haa69b, Hal65, Ham74, Har69, Har63, Har64a, Har65a, Har71, Har66b, Har66c, Har68b, Has78, HDBP68, Hei74, Hei72b, Her72a, Her69, His75, HH77a, HPR77, HcL78, Int61b, Int61c, Int61d, Int61e, Int61f, Int62b, Int62d, Int63o, Int63p, Int63f, Int63e, Int63g, Int63h, Int64a, Int64g, Int64c, Int64d, Int65e, Int71a, Int77b, Int78d, Ind60, IJ79, JOW72, Jam78, Jam66b, JCMS76, JCMS77a, JCMS77b, Jon76, Kal72a, Kan77, Kan71, KF72, Kar68, KS75a, Key73a, Key73c, Key73b, Kha76, KQS74, KS72b, KS74, KS75b, Lea70, Lar63a, Lar63b, LB77]. programming [Lee67b, Lee72, Lee74c, Lee77, LBM^{+}80, Lep76, LP71, LP78, LJ71b, LJ71a, Lyc80, Lyt75, MHH71a, MHH71b, Man63, MS77a, MS78, MS69, MH72, MH73, Mat72b, May72, McC68a, McC61, McC62, McC64b, MD64, McC65a, McC65b, MD66b, McC67c, MD68, Mei75a, MS77b, Mil73a, MT72, Moc69, Moc70, Moc71a, Moc71b, Moo75, Mos64, Mul68a, Mul68b, Mul80b, MS75a, MS68, MS70b, MS70c, MS73b, MS73c, MS73e, MS75c, MS75d, Nat70b, Nat72, New75, Nic74, Nic75b, Nic75c, Nic80a, Nic80b, Nic80c, Nir69, Nut78, Nyd68, Obr70, Obr71, Org61a, Org61b, PB73a, Paw65b, Pay64, Pin73, PC78b, Pra65, Pri69, Pri75, Rad75, Rad76a, Raj77, êR76, Rau68, Rei72b, RR73c, RR73d, Ros78, Rul66a, Rul66b, Rul68a, RFP73, Sal77a, SSS77, Sas74b, Sch68, SB78, Sch79c]. programming [Sch80g, SW74, She70b, Shn77, Sil71, SS76, Smi72b, Smi73a, Spe69a, Spe77a, Spe77c, Spe78a, Sri69, SM66a, SM66b, Ste60c, Sto80, SD72, Stu68, Stu70a, SM72c, SM76b, Tan78b, Tay77, TG79b, The68, Tok68, TW71, Tom71, TB80, Uni80a, Upc72, Vel67, Wag75, WMM72, Wat68, Weg66, Wei69, WB71, Wid79, Wil76c, Wit79b, Wit79a, Wit79c, Wit79d, Wol73, Wu73b, Wu73a, Wu73c, Wu77a, Wu77b, Zav73, Bee77b, Flo70a, Jun68, Kre66b, Nak68, DM66b, Bri67, Elt66, Wil74, Wol68b, McC74c, VP80a]. programming/Fortran [Rad76a]. Programmirovanie [SM76a]. programmirovaniia [Cal78, DG78]. programmirovaniiu [Lam78]. Programmproblemen [Rin77]. Programmsystem [Deu73]. Programmsystems [Sto76]. programmy [SZ80]. programok [Kor77, ZSF78]. Programovanie [Ham79b]. Programowanie [Bad77, BF72]. programozasi [Kor77, LV77]. Programs [AS71, AK80, Bac78a, Bee70d, Bee76, Bee78, Boy74b, Dea71, FH74, Gar74, Hig79a, Hig78, Int67a, IR78, Kah80a, Kah80b, Ked80, Knu70, Knu71, KMC72, KBC^{+}73, KBC^{+}74, MR70a, May73b, Mer78b, MR70b, OF76, Rob67a, RT76, SK80, Sha65, SS66, Van68b, Wat76, Wei66a, Wei66b, Wil74, Wol78a, dL78, Ari76, AD73, And73, Ano64, Ano70e, Ano74b, Ano74c, Ano74a, Bec73, BC77, BC79, Bor69, Bra74, Bre67, Bro71a, Bro73a, Bro74, But66, Com80b, Cam77, CM66, Cha79a, CS73a, CR74, Cla80, Com80a, CL80, Das74, DvC69, DS62, DS63, Dea77, DS76, Dor79, FP75, Fis78, Fit74, Fos74, Gen66a, GS70, Gra70a, HJ74, Har80, Har66a, Hei72a, Ho73, Hob67, HP74, Hol68, Huy77, Irv60, Jam73a, Joh71, JV67a, JV67b, JV68, Joh72, JS74, Jon64]. programs [Joy78, Jul75, Ken70, Ken80, Kra74, KW75, Kru67, KTZ68, KTZ71, Lan80, Lar69, LML69, Lee69, LM76, Lud69, Mac69, McC64a, McC74b, McG76a, Mee74, Mer78a, Mer78c, MR67, MI75b, MS66, Moc69, Moc70, Moc71a, Moc71b, Mur66, Nor63, O’K64, PB74a, PT73, Plo75, RV78, RPE79, Rin79, Rob69, Rob67b, RS80, Sch80f, Kan68, Sel77, SMD71, Sla67, Sla72, Smi79, Smi80, SW80, Soy71, Spi65, Ste72b, Sti72, SD72, Swe67, Swi64a, Taj65, Tam66, TC70, Thr79, TB65, Tur68, Vas72, Wal68, Wat68, Whi76, Wis69, Wor69, Wri77a, Wri77b, Wri77c, Wri77d, YHE69, ZSD80, Bee77a, BD80b, Din72]. programski [Bit75]. Progress [ANS69c, ANS69a, ANS69b, Ano69b, DEN79, PMBK80]. progressive [CJM67]. Project [Int75a, Cha76, Clo72, Ho73, Hon76, Jam73a, Lin76, Man63, Ube76, Wil77a, Zal73]. projection [Edw76b, SK69]. projector [Mur77a]. Projektorganisation [SF76]. prolate [Bur76]. Prolog [BM73, Rou75]. promotion [Ano78b]. prompter [Int71i, Int71j, Int72b, Int74e]. propagation [HM62a, HM62c, HM62b, HO64, Her64, HM64, MI75b]. Propagator [Bee80d]. Properties [Mal72, Sch72b, HPB73, PTM77, PB73b, SM73b]. property [Mac73]. Proposal [ADG70, Bee79b, Bee79d, Bee80b, CLS64, ABH^{+}71, Bra77, Bee79a, Bee79c]. Proposals [Smi77]. Proposed [ANS76a, ANS76b, KRS78, Mei76, Woo77a, Woo77b, ANS76c, ACM76, Bid79, Bri68a, DCHR76a, DCHR76b, Fel76a, FRS77, Kni76a, Kni76b, SIG76, Tan78a, WPK78]. proposta [Cad71]. proton [Kol74, Moo76]. Proverbs [LC78, Led75]. provide [CF71, McC71]. Provided [GM73, Raf79]. Provision [Ano64]. provisional [Spe9 ]. Prozesskonzept [Win79]. przykladow [ATW77]. Pseudo [ADG70, Bre74]. Pseudo-Random [ADG70, Bre74]. pseudorandom [Ano70c, Edg79, Krá72a, Ove72, Pay70, Tri79]. PSI [Ott78]. Psychology [Lew80d, Lew80c]. psychometric [Ler72]. psychometrics [CK80]. public [Tro64]. publication [Rob69]. Publications [Joh66b, Joh66a]. publishers [Shn77]. PUFFT [RSD65]. pulpwood [New67]. pulse [vM79]. pulse-height [vM79]. Purdue [RSD65]. puroguraeming [JcK73]. puroguramingu [Mur70, Nih69]. Purpose [Fel76b, PC78a, Bla79, EH68, HN70, MA78, Sak79]. PWR [EKM74].

Q [BC67, Par70]. Q-mode [BC67, Par70]. QDRTC [Kah66]. QPTOR [JCMS77a, JCMS77b]. QRT1 [Sta60]. Quadratic [LP73, CJM67, Fle70, JCMS77a, JCMS77b, SA74]. Quadrature [Blu77, Gen72, Kar77, Pie73, CJ77, EE77]. Quadric [Wei66a, Wei66b]. quadrupole [vM77]. quadrupole-distorted [vM77]. qualitative [BY73, NL75]. Quality [kC80, GKB74, RMM69]. quantities [Joh74, Tro66a]. quantity [SMM65]. quasi [FL76, Fle72]. quasi-interactive [FL76]. quasi-Newton [Fle72]. quatre [Dre67, Dre75b]. Queue [MSS78a]. quick [McC79, FL76]. Quine [Cam65]. Quine-McCluskey [Cam65]. QWIKTRAN [McC79].

R [Nak68, Wol68b]. R2D2 [Erd80]. Racah [AD73, Tam66]. radial [KM77a, KM77b, Kat77, WG75]. radial- [KM77a, KM77b, Kat77]. radial-inflow [WG75]. radiant [MS66]. radiated [Rin79]. radiation [HW75, U.S61, ZD78]. radiations [Ful74]. radioactive [U.S61]. radioisotope [U.S61]. radiolysis [Wer65]. radionuclides [CL80, PMBK80, Won67]. radiopharmaceuticals [HW75]. RAGTIME [PMBK80]. raindrop [AJ69]. Raising [SH67]. Raman [MW71a, MW71b]. Ramona [Bak68]. Ramp [Lag74]. RANDHIND [Hun74]. Random [ADG70, HK72, Kno73, Kru69, NO75, Sch79a, Smi70j, Bre74, Bro80, Fel75b, Fri69, Gro69, Hon71a, Kir79, MB68a, MB68b, NO72, New73, Sou71, U.S61]. randomness [Tri79]. ‘RANDU’ [Fel75b]. range [DP76b, PV74, Tro66b]. range-energy [Tro66b]. ranges [FS76, Mee74]. rangs [Aub76]. rank [Mag71, Pow69, Fry71]. RANTEST [Tri79]. Raphson [TI72]. Rapid [Roc70, Wie75]. rapides [Fer63]. RASO [Har80]. Raster [BJ74]. RATFOR [Com78, Ker4 , Ker75a, Ker75b, Ker79, Ker70]. Rating [Ant80]. Rational [Ker4 , Ker75a, Ker75b, Ker79, LS76, Flo78b, Hal71, Hal74, Ker70, Kor79, Mur71, Per77, TG79b]. rationale [ES78]. ratios [ZD78]. Raulefs [FJA80a, FJA80b]. RAVE [HO64]. ray [CL80, Maz77, Moo76, vM76, Cri77, Day63, Oer71, Smi63b, SH78, Ste72b, TC70]. RBAD [Gof74]. RBEOER [ZD78]. RBEs [ZD78]. RC [HW72]. RCA [Sas69]. RDUMP [Bee70d]. Re [Bem61]. reacting [Erd80]. reactions [Phi67, SMD71, SR73, SDH74]. reactor [MV66]. reactors [Bak68, MV66, U.S61]. read [Bai63, Hea68b]. readability [Ho73, McC74b]. reader [BD80a]. Reading [Art75, LT76, New73, HV74, Smi73b]. Real [Nie72c, Nik78, Ono79a, Ono79b, Ste76b, Ste76c, Ste78a, Ste78b, Cas62, Fan65, Hin76, Joh74, MU75, Mod74, Upc72, Vic70a, Whi76, ZT76]. real-time [Hin76, Mod74, Upc72, Whi76]. Realisierung [Neh74]. Realization [Ree79, Gav76]. REBOUND [Sho76b]. Recapture [Whi71a, Whi71b]. receiver [Sig80]. Rechenanlagen [Mue69]. Rechenprogramm [Rot71]. rechner [jH78]. Rechneranlagen [Sch78a]. Rechnermodelle [Sch78a]. reciprocal [Hat78, Hil79b]. recoding [Hou62]. Recognition [Kra72b, Sch72a, CF60, Mas62]. recoil [PV74]. reconnaissance [SDZ80a, ZSD80]. record [Con73d, Con76a, IEE79]. records [Art75, Zor68]. recovery [Spa75b]. recruitment [PT68]. rectangular [LML69, Pow69]. Recursive [AI80, Aye63, Kal72b, Lea67]. recursively [Gut75, Gut76b]. REDUCE [Ina80a, Ina80b]. reduces [Gel69]. Reducing [Bre73]. Reduction [Han72a, Cam65, Han72b, Hat78, KRB77, Mac64, Spa80, Wer65, Wil65]. Redundancy [Yoh72]. réelle [Lou73]. Reellen [Neh74]. referee [BW75]. Reference [BLF80, CGG^{+}91, Day72b, Flo78a, Gen66b, Gen69, Gow73, Hil73, IBM56, Int58, Int59f, Int60c, Int61d, Int61e, Int61f, Int62d, Int62c, Int64e, LF78, NL71, Nak77, Rou75, Wil75, AR72, Ano68b, Ano70b, Ano70f, Ano70g, Ano73, Ano75c, Ano75d, Ano75e, Ano80a, Bar73a, Bur67, Bur68b, Bur70b, Bur73b, Bur73a, Bur74, Bur78, Con62a, Con62c, Con62b, Con62d, Con64c, Con64a, Con64d, Con65, Con66b, Con67b, Con67c, Con68b, Con68a, Con68c, Con69d, Con69c, Con69b, Con69a, Con70, Con71b, Con71c, Con71e, Con71f, Con72a, Con73c, Con73e, Con73b, Con73f, Con73g, Con74, Con75a, Con75e, Con75f, Con78a, Con78b, Con79b, Con79c, Con79d, Con80b, Con80c, Dig70, Dig71b, Dig74, Dig75b, Dig75a, Dig75d, Dig75f, Dig76a, Dig77a, Dig77b, Dig77d, Dig78b, Dig79b, Dig80e, Dat67a, Dat67b, Day79, Ele68, FLM70, Fri71b, Fri73, FLM74, FL74, Fri80, Gen67, Gen80a, Gen77b]. reference [Gen77a, Gen66c, Got73, Hew71, Hew74, Hew76b, Hew76a, Hew79b, Hew79c, Hew80a, Hon75d, Hon79a, Int59a, Int59b, Int59d, Int60b, Int61b, Int61c, Int63g, Int63h, Int64c, Int64d, Int66a, Int72a, Int72b, Int72f, Int72g, Int74b, Int74f, Int75b, Int75a, Lea70, Lar67a, Lar67b, Lec66a, Lec66b, Lew80a, LM70, Mis78b, Mer58a, MM78, NCR69, Nic78, Rid79, Ser71, Sta69, Sci64, Sys73a, Sys73b, Sof80, Spe66b, Spe66a, Spe69c, Spe73a, Spe73b, Spe74a, Spe74b, Spe76a, Spe76b, Spe78b, Spe79, Spe80d, Tan80a, Uni69a, Uni69c, Uni80c, Xer70a, Xer70b, Xer70c, Xer71b, Xer71a, Xer74a, Xer75a, Xer75b, Xer75c, Xer76a, Xer76b, Uni75b]. referencer [Bla79]. references [BT76b]. referencia [Lec68]. Refernce [Cra80]. refinement [Tay80]. reflected [Rin79]. reflection [MC64]. refraction [Hat78]. Regarding [Ein76]. regdata [Sea80, Sea80]. Region [Sal77b, MV66, Sal78]. Regional [Int75a, CB69]. regionalisation [DO79]. Register [Bee70d]. Regression [AK78, Fas67, Kas74, PG65, ESD68, Fin72d, Fin72f, Fin72e, Fin77, Kru68, Spä79a, VP76]. Regressions [LaM72]. regularly [Har68a]. regulatory [Leo74]. Reinwald [Mue75]. Reinwald-Soland [Mue75]. Relating [DA68]. relation [Tro66b]. Relational [Rip77]. Relationships [Swi64b, Swi64c, Bar66, Lal75, Tip76]. Relative [BA72, Gof74, Mal70]. relativi [SS68b]. relativistic [Phi67]. relaxation [RCL75]. Release [Bee78, Hil70, Dig80b, Dig80d, Fin72f, Pri77b, Spe84]. reliability [BT76b, Sho76b]. Reliable [FJA80a, FJA80b, FJA80a]. Remark [Bre79b, Ein74, Ein76, Fut78, Nie72c]. Remarks [SS72]. Remez [Rod76]. remote [BD80a, Har73]. remotely [Hem70]. remotely-sensed [Hem70]. Removes [Par75]. René [Bem61]. renumbering [Hyd66b, Har80, Mur66]. reorganizer [Sep75]. repeatable [Her74]. repeated [Fin77]. replacing [Fra77]. Report [ANS71a, ANS71b, Jun68, Kre66a, Nak68, Rab62, ANS71c, Cle66, DEN79, Fed70, Fed87, Kno72, PMBK80, SSS78, TS73]. Reports [Her70, U.S61]. Representation [BK72, Rip77]. Representations [AB69a, SS78a, Wol68a]. representing [Han74b]. required [BS78]. Requirements [FH74, FMC78, Gel69, Sin78]. requiring [Pow70b, Pow70c, Squ70]. Research [BF79, Gue73b, Spa75c, CK80, Clo72, Flo75, Ho73, Man63, Spa79b, Ube76, Zal73]. Research-Algorithmen [Spa75c]. Researd [Spa75a]. Researd-algorithmen [Spa75a]. resheniia [SZ80]. resident [Int67a].

RESIS2D [Dey76]. Resistivity [Dey76, Pin80, Sco78]. resolus [LPJ79b]. résolution [Ric73]. resonance [SDH74]. resource [FB69]. respect [Cra75]. Response [TG79a, Joy77, Joy78, Rin79, SD67, Sei75]. Response/360 [Lea70]. Restricted [BS73b]. Restructuring [Ho73]. Resulting [KMC72]. Results [BH73a]. Retrieval [Bre73]. retrieve [LGF75]. Reveal [Mal72]. reversals [BS64]. Review [Bar71a, Bar72a, Bar73a, Bar80b, Bar74, Bri67, Bru66, Din72, Elt66, Flo70a, Flo70b, Gar72, Gar74, Gow73, Hui65, Jun69, Kre66b, May73b, Pat73b, Ree73, Ree75, Ree76, Ung69, Van68b, Wil74, Wil75, Wol68b, Cha73, Fry71]. Reviews [Hen67, Hil73, Hil79a, Kar77, FJA80a, FJA80b, VP80a, VP80b]. Revised [Ano69a, Ano77a, KM77a, KM77b, Kat77]. Revision [Ame78d, Con64d, Dat73]. revolution [Pat67, PNK65a, PNK65b]. RFC [Kri71]. Rheingans [Flo70b]. Rho [BR75a]. RHS [Squ70]. rhythm [Rub69b, Rub69a, Rub69c]. Riccati [Lee74a, Ric73]. ricerca [Rid67, Rid78]. Richard [Kar77]. Riemann [Spi65]. ring [BLY70]. risk [McG76a, McG76b]. River [Cla75, ZSD80]. Robot [Arm78]. RObust [Kle78]. rock [Bea75]. rocket [PG67]. rocks [Lou67]. Rogers [Gar72]. ROKDOC [Lou74, Lou67]. rolling [PG67]. Romanian [Ste74]. Romberg [Ant72, Ant72]. Romberg-Integration [Ant72, Ant72]. Roots [ST73c, ST73b]. Rose [Wil80a, Wil80b]. Rosenbrock [MM73a, MM73b]. ROSENET [Wil80a, Wil80b]. ROSEPACK [Kle78]. rotating [AK77, Tri73]. rotation [Mat72a]. Rotational [Ful72, SG69, VP75]. rotations [Wat73a, Wat73b]. roughness [Hob67]. round [Cha71b, Cha72, Dew72, Fla72]. Rounded [Lag74, Lin71]. routes [RSBR69]. Routine [Int60a, Bai63, Das74, JCMS76, JCMS77b, New73, Sep75, SD72, vM78a].

Routines [Dur80, LaM72, MM58, WB65, Bac72, BPW72, Bar73b, Bid79, Bra74, Con75b, Con75c, Con79d, Cli78a, Cli78b, Gen80b, Gen73, Høj69, Høj70, Hol77, Joh74, Lou67, Mor75, Ste73, Uni80c]. Routing [Fen73]. Roy [Fry71]. rozwiazanych [ATW77]. RPG [Har69, Kan71]. RSDUMP [Bee70d]. RSTS [Dig75g, Dig77f, Dig75e, Dig77e, Dig80b, Dig80c]. RSTS/E [Dig75e, Dig77e, Dig80b, Dig80c]. RSX [Dig78a]. RT [Dig75f, Dig75g, Dig76c, Dig77f, Dig80c]. RT-11 [Dig75f, Dig75g, Dig76b, Dig76c, Dig77c, Dig77f, Dig80c]. RT-11/FORTRAN [Dig76c]. RT-11/RSTS/E [Dig75g, Dig77f]. RTE [Hew79c, Hew80b, Hew80a]. Rudeanu [Les73]. Rudiments [Mei71]. rules [Hun74]. RUN [Uni73, Uni74a, NC76, Con66a, Hon71b, Jon64, Ano75a]. run-time [Hon71b]. RUNT [Han74a]. RUNW [Han74a]. Russian [BK77, FMM80, Gav76]. Rutinas [Oli71]. RW [Vic64]. RW-300 [Vic64].

S

[Jun68, Kar77, Kre66b, Ree73, Wil74, DG78, Mee74, Pic66, BF79, Cha67, Col76, CDG80b, DH78, DLS79, HD78a, HH77a, HH80, HH77b, HH78, Int58, Int59d, Int59f, Sea80, TB80]. S14 [Kuk72a, Kuk72b]. S15 [Don73a]. SABOR [PNK65a, PNK65b]. SAC [Hor72, LS76]. SAC-1 [LS76, Hor72]. Sadwichschalen [Die72]. SAGES [KG76]. SAHYB [Ano68c]. SAHYB-2 [Ano68c]. salary [Cra75]. Salmon [AS71]. Salvadori [Joh66a]. Sample [BS75, Fry71, Kas74, SH67, Pre70]. sampled [Edw76a]. Samples [Pom74]. Sampling [DB73, Mag71]. samt [Hus76]. Santa [IEE78]. SAP [WM60]. SASP [BBB64]. SATRAP [Las71]. saturable [MA78]. SAU [Dat73]. Savannah [ZSD80]. SAVE [Her64]. saver [Ack64]. sawfly [McL73].

Scalar [Cli74b, Cli74c, Bus67, vM78b]. Scalar- [Cli74b, Cli74c]. scale [MT75, vM75]. SCALE1 [Lew73]. SCALE2 [Lew73]. SCALE3 [Lew73]. Scales [Ant80, Lew73]. scaling [NL75, vM79]. Scan [BJ74]. scanner [Har73]. Scatter [Bre73, Gen78, LV73]. Scattergram [Fas67]. scattering [Maz77, MSNC61, Mel62, MSR66]. Schaum [Lip78, LP78]. schedule [Sin78]. Schedules [PD67]. schematami [ATW77]. schematic [Moc69, Moc70, Moc71a, Moc71b]. schemes [LM76]. Schnelldrucker [Gut79c]. Schnelldrucker-Setzprogramm [Gut79c]. School [Smi67b, Weg64, CDH75, CS77a, Cra75, O’D74, RSBR69, Sch79c, Sch80g, Tro64, Mt.79]. schools [Int69a, Int72i]. Schranken [Ant72]. Science [AI78, CPM72, Fre74, HBJ76, Jam73b, MS73d, Per75, Ree76, SP70, Wal72, Abr72, Baj72, BC70, Cha76, Don71, For70, For75, Gra79, HM75, Jam73a, LB68, MD64, MD66b, MD68, Obr71, OR77, Ral71a, SR76, SW79, FJA80b, FJA80a]. Sciences [BR74, MH78, Tou70, Bur68a, Int64a, Int65a, Int65b, Int78b, Int78c, Int70a, Vel67]. scientific [Cal69e, Dig75f, Dat73, FB69, Rob69]. scientifically [Feu77]. scientifique [CR69, CR73, Cha79b, Lev71]. scientist [CG73, SJ62a, SJ62b, SJ63]. Scientists [Wag70, Gol66, Mei69, MS68, MS73b, MS73c, MS73e, NL68]. scope [Ste60b]. Scores [Hoa72, Hoa73]. SCP [Ari76]. SCRIPTOR [Gut79c]. SDS [Sci64, Sci65]. se [Ano75b]. sea [Lar69]. searches [MSR66]. Searching [dlB59]. seas [Lar69]. Second [ANS71a, ANS71b, Han78, Hon76, HPR78, ANS71c, BK77, BK75, FM76, HPR77]. secondary [CB69, CDH75, Tro64]. section [SR73, Wil65]. sections [Fer63, Kol74, MC70, Moo76, PV74, SMD71]. Sector [Sad72]. sediment [Hei70, PTM77, SDZ80a, ZSD80]. sedimentary [Lou67, Oja70]. sedimentation [BCS68, HW67]. sediments [PG66]. SEEK [MSR66]. seems [Sha77]. seismic [Hat78, Joy77, Joy78, McG76a, McG76b]. seismograms [RB76b, RB76c]. SELECT [LaM72]. Selecting [SR76]. Selection [LM76, CJ77]. Self [Far66, Plu64, Ste60c, Wol68b, BP74, Bur70a, Coo76b, Dig80a, DS72, Dun69b, Dun74, Hal65, HDBP68, Hur77, HF78, Int72l, IA80, KP70a, KP70b, Lin76, Mul80a, NBH70a, NBH70b, Plu61, Plu63, SA74, SJ62a, SJ62b, SJ63, Smi73a, SF75]. self-contained [IA80]. Self-Instruction [Plu64]. self-instructional [BP74, Dun69b, Hal65, HDBP68, KP70a, KP70b]. self-learner [Bur70a]. self-metric [SF75]. self-paced [Coo76b, Dig80a, DS72, Lin76, Mul80a]. Self-programming [Ste60c]. self-study [Dun74, Int72l, NBH70a, NBH70b, SA74]. Self-Taught [Far66]. self-teaching [Plu63]. Self-Training [Wol68b, SJ62a, SJ62b, SJ63, Smi73a]. Semantic [Mau72b]. semester [ES78]. Semi [Yoh78, Zal73, Her64]. semi-automatic [Her64]. Semi-Portable [Yoh78]. Semi-structured [Zal73]. semicircles [Dun79]. semiclosed [Lar69]. sensed [Hem70]. Sensible [Ken74, BK75, KS75a]. sensing [Har73]. sentences [Fri69]. Separable [Ste79, SS79a, SS79b, SS79c]. September [Axf72, Ano69a, Dat73]. sequence [ES78, Han67, MP72, Mee74]. Sequential [Ben77, Hor68, Lar67c, Raw77]. Serial [Hai65]. série [Jak73]. Series [Gen66b, Hon71a, Hon71b, Hon73c, Hon77b, Hon79b, Pie74, Rob67a, Sch80c, Spe84, Bus68, Bar80c, Bur78, Con66a, Con66b, Con71c, Con76c, Clo72, Fla71, FLM70, Fri73, FLM74, FL74, Ful73, Har68a, HB63, Hon73d, Hon75a, Hon75b, Int62b, Int66j, Jam66a, KW75, KRB78, LM70, Mac64, O’D74, RB76a, RBK76, Sci64, Sci65, Sch68, Sei72, Sim66, Smi70b, Uni80c, Wil72a, Dat73, Gin78a]. Serious [Lar73a, Lar73b]. Service [Gof74, Gra70b, Gen77b, Gen77a, Int71c, Int71h, Int72m, Rad76b, Spe9 ]. Set [Bee79a, Bee79c, Bee79b, Bee79d, Bee80b, GS79, BPW72, Duf77b, Duf77a, Duf80, Fic71, Huy77, Wol68a, Wor69]. Sets [Lea78, Zoh80, Coc80, Com80a, CR71, Sou71]. Setzprogramm [Gut79c]. several [Sou71, SMM65, Wri66]. SF/ [HH77b, HH78]. SF/k [HH77a]. SFOR [O’N74]. SFTRAN [Bee79e, Fly73]. SFTRAN3 [Bee80c, Bee80d, BLF80, LF78]. shaded [Dil79]. shadow [Bai62, CF60, Mas62]. shafts [Tri73]. shallow [NM70, NM78, Nav78]. shallow-water [NM78, Nav78]. shaped [Dey76]. Shapiro [Mar78a]. Shapiro-Wilk [Mar78a]. Shared [Sch73, Hea68b]. sharing [Blo68, BR78, Con73f, Con73g, Con74, Gra70b, Gen67, Gen66c, Hon73d, Hon70b, Mar66, Pin73, RCM66, Uni69a, Wes69, Wit74, ZSW76, ZSW77, ZSW79, CS73a, Zin79]. shells [PNK65a, PNK65b]. shelters [U.S61]. shi [Jam75]. Shift [Fry71, Feh68]. Shih [Ano72d, Ano75b]. shock [Gen66a]. Short [Ste75a, Krá72a, Lee67b, Lee72]. Should [Hor75, HM75]. shroud [KM73b, KM77a, KM77b, Kat77]. side [Per78, Rin79]. Sided [Boh75]. siedem [ATW77]. SIEGEL [Veg71]. SIFT [Con71d]. SIGGRAPH [KRS78]. Sigma [Ano70a, Ano70b, Ano75e, Ano70f, Ano70g, Ano73, Ano75c, Ano75d, BR78, Xer70a, Xer70b, Xer71a, Xer73]. SIGMUS [Sch80e]. Signal [Fre76, Rej72, TS76, vM79]. SIGNUM [ACM79]. SIGPLAN [ACM78, Gri78, Ros78]. SIGSAM [HJ74]. silicates [Oer71]. silicon [LTB80]. Silini [SS68b]. SIM [MSS78a]. SIM-Queue [MSS78a]. similar [BLY70]. SIMPDX [Uni80c]. SIMPDX/SIMPLX [Uni80c].

Simple [AK78, CZ72, DDM^{+}75, LS75, CL70, Han74b, May72, Rus79, vM76]. SIMPLETRAN [Spe70a]. simplificada [McC78f]. Simplified [Eng75, KF72, Man71, McC74e, RR73c, RR73d, Sil71, SS76, Zav73, Ayc80, Lep76, MHH71a, MHH71b, McC74c, Mye73]. simplifying [Cle66]. SIMPLX [Uni80c]. SIMSCRIPT [Shu69, Wei67]. SIMSCRIPT-FORTRAN [Wei67]. SIMUDELT [She78a]. Simula [Pal68]. simulate [Joh72, New67, NS69, ST73a, vM77]. Simulated [Smi73e]. Simulating [Gen75a, Gen75b, Gen75c, Oja70]. Simulation [Ano68c, Cla75, Hin76, LL65, Nie72a, Nie72b, PC78a, PRO80, PK69, Sch78a, Wei65, Ber77, Blu65, BCS68, EKM74, GP73, HW67, Hei74, Joh80, JID80, Kau65, Kru68, MK70, PK67, She78a, Sig80, SYR77, Tor69, Upc72, Vas72, Whi76]. Simulationsprogramme [RS69]. Simulationssprachen [MSS78b]. Simulator [Ful72, Sch80e, Tho68, WH73]. Simulator-Generator [Sch80e]. simulators [SW74]. Simultaneous [Nik78, Dif72, Les73, Mat72a]. Simultaneously [KMC72]. single [Bre79a, FM76, FS76, Hol80, MT78, MC70, RS72, SDH74, Var77]. single-facet [Bre79a]. single-valued [RS72]. singular [DP76b]. sinusoidal [Fla71]. Sistema [Ano70d, CC70, Ins64, Ins70, Ins74]. sistemi [BG78, BT76a, Sic74]. site [McG76a]. SITGO [Ste80]. SITGO-10 [Ste80]. SITGO-10/ [Ste80]. sitizyusiti [Nis78]. Six [Cli74a, BS78, Edu72d]. Size [BS75, AJ69, Dav72b, PG66]. Skewness [MZ75]. skills [Dun75b, Tip76]. Slater [FP73, Gar74, May73b, Van68b]. slender [PG67]. SLIP [FPB72, Fin72a, SS68a]. slopes [McC71]. sluttningar [Hus76]. Small [Pom74, GP73, Mal70, Maz77, MP65]. small-angle [Maz77]. smaller [Bah69].

Smallwood [Hui65]. smes [AE79]. Smirnov [Pom74]. Smith [Wol68b]. Smooth [Aki74, ClW78, RS72, Bar73b]. Smoothing [Dur80, Bar73b, SS78a]. SNOBOL [Kee75]. soboi [AE79]. Social [BR74, Gue73b, Bur68a]. Socio [Sad72]. Socio-economic [Sad72]. sociologists [Bur69]. Software [ACM79, Air77, Bro77, kC80, Cob75, Dic74a, Eva72a, Eva72b, FJA80a, FJA80b, Fos73, Gin78a, HH79a, KP76, Spa79b, Tan80b, Tou70, Yoh80, AHP77, Bra76, CDGW76, Leu79a, Leu79b, Ent80a, Ent80b, Feu77, Flo78b, Hin76, Hon73d, Hon76, Hon73c, Hon77a, Jet74, JM76, KG76, PB74a, Shu75, SF75, TRW73a, TRW73b, TRW73c, TRW73d, Uni80a, Amk73, SF76]. soils [Oer71]. SOL [KM64]. Soland [Mue75]. solid [WCT68]. Solomon [BK75]. solpositioner [Hus76]. Solution [Abd80, Ano72e, Ein74, Fia73, FSC73, Geo80, Kar77, MC80a, MC80b, MC80c, Sch76, Ste79, SS79a, SS79b, SS79c, Zoh80, Bor69, BD80b, Car74b, Cha74, Dif72, DS75, Hal75, Hol80, Les73, Mar71, MT78, Mou70, Ree68, Ree71, Ree72, Rei72a, Rob79, SSS78, SS75a, SS75b, Tea72, Vic70a, ZT76]. Solutions [Bla68b, BH73b, DM66b, JSW77b, McC70b, Mer78b, Moc71b, MS73e, MS75d, Per72b, SM72b, Tel80, Wal80b, Wol78a, Wu73c, But66, CR71, Mer78a, Mer78c, RG76, Wal80a, Wu77b]. solve [Lal75, Lep76, Pec77]. Solver [MG70, Mol72c, Mol72a, Sch75]. Solving [CS73b, DZ78, KGY80, Spe77d, ZN79a, ZN79b, Bar77c, BW78a, Bar79b, Dav72b, DLS79, Elk65, FK77a, FK77b, Lee74a, Lew80c, MT72, NM70, NM78, Nav78, Pic66, Pow68, Pow70a, SW75, WMM72, Eng75, Lew80d]. Some [Bal69, BP78, FK76, Gin78a, GM73, Håv71, HLS73, PH77, RJAS78, Row76, Sab76, Swe67, Fit74, LS71b, Mac69]. Sondage [Die77]. sorties [BDI72]. Sorting [Cha71a, Roc70, Smi71c]. sosyal [Yur76]. SOTRAN [Ham69]. sound [MI75b, Rin79]. Source [FF75, Bar77a, Hyd66b, Har80, Joh71, Low76, Mas60, Mue75, Mur66, Spa75b]. sources [CB69]. sous [Ric73]. sous-programme [Ric73]. Southern [IEE75]. Sozialwissenschaften [Neb71]. SP [Mac70b]. SP/1 [Mac70b]. Space [BF71, Sch76, BS78]. spaced [Jam66a]. SPAN [Lil68, Mac67]. Spanning [Pag74a, Whi72]. SPARKS [Mar78b, Str78]. Sparse [Geo80, KGY80, She78e, She78d, CR71, DO79, Duf77b, Duf77a, Duf80, Rei72a, She78b, She78c, ZT76]. spatial [ER79, MT75]. Speak [Nor0 ]. Spearman [BR75a]. Special [Day72b, Gow73, Hil73, Osy76, Wil75, Bar73a, Day79, Weg66]. species [Gof74]. specific [Pin80]. Specification [Ano67, Bur65, ABB^{+}74, Dat73, Har78, Spe9 ]. Specifications [IBM54, Int64b, Ran65, Sta74, Int67a]. specified [Bai63]. spectra [CL80, Kan68, VP75, Sas69]. spectral [Bar79a, Kra74, KRB78, Nut76, Sig80, Cha77]. spectrometric [Smi79]. spectrophotometry [Cha77]. spectroscopy [TC70]. spectrum [EP67, Gen66a, vM75]. speed [Gea65, Mag71]. Speedcoding [Bac54]. Speeding [Hig78]. speeds [Low76, Tri73]. Speedup [KMC72]. spelling [Hei72a]. sperimentali [SS68b]. Spezielle [Tec72]. spherical [Gav76, Hol67, Hol68, MW71b]. Spiess [Flo70b]. spin [SDH74]. spiral [KS74, KS75b]. SPITBOL [Tha77]. Spitsym [MK70]. Spline [BH73b, Dur80, MP73, Gaf79, Kat68, Sca70]. Splines [Cli74a, ES74b, Cli74b, Cli74c, PS78]. SPLINS [Bar73b]. split [MK70]. Sponsored [Eva72a, Eva72b]. square [Wie75]. Squares [ZN79a, ZN79b, Bar79a, Bom67, BML62, Tay80]. SQUIRREL [Sal70]. SRRIT [Ste78a, Ste78b]. SRU [NO72]. ST [ZT76]. stability [GS70, MI75a, U.S61]. stage [Fri71a, Mot66]. stand [SK69]. STANDARD [Ber70a, AB66a, Ame78c, ANS78, AB66c, AB66d, Ano80d, BM79b, Bee79a, Bee79c, Bee79b, Bee79d, Bee80b, CS73b, Dat66, DH78, For78, HD78a, Ame77, Jn69, Lam71a, U.S78, Nat70b, Nat72, Nik78, REC75, Sta69, VP80a, VP80b, Woo77b, AB66b, Ame66c, Ano72c, BM79a, Bid79, CR74, Dav72b, DCHR76a, DCHR76b, Fel76a, Fri80, Gof74, Haa69a, Haa69b, Ins76b, IA78, Jaf79, KS75b, MS77b, PG66, RW76, Sal76, Sch74, Spe69a, Spe69b, Tan78a, Uni80a, WPK78, Ame78a, Ame78b, Ame78d, Ano76b, Ein76, FRS77, KRS78, Woo77a, Bar72a]. standard-size [PG66]. Standardisation [HM80]. standardisierte [Fri75b]. Standardization [Hei64, Hei66, ABB^{+}74, Bra79, Eng74, Jam73a]. Standards [ANS69c, ANS71a, ANS69a, ANS71b, Cad71, ANS69b, ANS71c, Ano69b, Bre76b, Lee77, LBM^{+}80, Ott78, Fri75b]. standpoint [Mat72b]. Star [Har77]. State [Hed77, AJ69, Day70, FB73, MU75, PG67, Pec77, Sol64, Wil72a]. stated [Hun74]. Statement [Ken74, BK75, FF75, KS68, KS70, KS75a, MHH71a, MHH71b, Mee74, Sit78, Sti72, Man71]. Statements [Han72a, BS78, Han72b, Hea68b, Lat79, MP65, Sal71b, Sal76]. static [BA73, SM73a, WK77]. Station [CS73a]. stationary [AK77, Clo72]. Statistic [DB73, Pom74]. statistica [Rid67, Rid78]. Statistical [AK78, Bak77b, BB77b, BR75a, BR75b, Boh75, BS75, DB73, Fre73, Gen75c, Gen78, Gow75, Hab72, Hab73, Lea78, MZ75, MH75a, Mon75, RST78, Smi71d, Spa73, ST73c, Spi72, Tho72a, Yat71a, Yat71b, Žil78, All67, Edg79, GKB74, Hog72, Joh71, Jon64, Lou67, Lyc80, Mac69, Rey69, Sch79c, Sch80g, SDZ80a, TD78]. Statistics [CCHT67a, Cra76, Gen78, Hil70, Int80b, Kle78, KH75, LV73, CK80, CCHT67b, FB79, Kir73, Lyc80, Sch79c, Sch80g]. statistique [Die76]. Statistischen [Bra75a, Bra75b]. STATLIB [TD78]. steady [AJ69, FB73, PG67, Sol64, Wil72a]. steady-state [AJ69, FB73, Wil72a]. steam [HPB73, PB73b]. stellar [Gus73]. step [O’D65]. steps [Sla71]. Stepwise [SH67, VP76]. STGPAK [Fis76]. Stichprobenparameter [Kas74]. stiff [Bar80c]. stil [Pet80]. Still [Arn65]. stima [BT76a]. Stirling [BPW72]. Stochastic [Whi71a, DS75, Sho76a, Whi71b]. Stock [Hog70]. STOFI [MS74b]. Stomach [MR70a]. Storage [Bre73, CPR75, Ful72, LV73, Mor73, Bra74, KW71, Moc69, Moc70, Moc71a, Moc71b, Sak64a, Sak64b, Sak65, Sak70]. store [LGF75]. storey [ST73a]. storm [Sho80]. strålningsenergi [Hus76]. Strachey [HN70]. strategy [Kha77, Vic70a]. stratified [And73]. stratigraphic [KRB78, RB76a, RBK76]. stratospheric [U.S61]. stream [Joh74, KM77a, KM77b, Kat77, MG71, SDZ80a, ZSD80]. stream-sediment [SDZ80a, ZSD80]. streamlines [KM73b, KM77a, KM77b, Kat77]. Streams [AS71]. Strecher [Bar74]. STREDO [Zor68]. strength [Die68, HM64]. stress [Cam77, Hol67, Hol68, Pol78]. stress-intensity [Pol78]. stresses [Pol78]. String [Lam71a, Mor75, Rey77, Fis76, Han75, Kau69, Kee75, Mac70b]. strings [Han74b, Lea67]. strip [Hor65]. Stromungen [Rot71]. Stroud [Kar77]. Struct [Bak75]. structural [DS76, Per80, SM73a]. Structure [GMPW79, MM65, Nic75a, Rul80, FP75, Gil77a, Gil77b, Joh65a, Joh65b, Joh76, Kal72a, MT80, Nor63, Sel77, SG69, Tor69, Tel80]. Structured [All75, AI80, Bod77, CM79, Col76, CS76, CDG80b, Doc79, Ell80, Gal75, Hor75, HM77, HH79b, Jay80, Jon76, Ken74, Lea78, Lew80b, LR77, MO80, MM78, RW76, Row76, TB80, Wag75, Wil76c, Wil77a, Zwa80, Ayc80, Bez75, Bon75, BK75, CW75, Col78a, Col78b, Cul80, Dar78, DH78, DLS79, Eld77, Flo75, FK76, FK77a, FK77b, Gra79, HBJ76, Has78, Hig75, His75, HD78a, HH77a, Hul73, Int77b, KS75a, Kha76, Mal77, Mei74, MH75b, Mei75a, MS77b, Mil73a, O’N74, Ovi77, Par77, Rei76, Sal77a, Zal73]. structurée [Era77, HH80, HH77b, HH78, RH78]. Structures [Bak75, BC72b, Mil73b, Mil75, OFP78, Cha70, Dey76, Don71, Ell78, Hou71, Mei75a, Mei76, Rin79, Smi77]. Structuring [Bak77a, Gom79, Sed77]. strutturata [LMP77]. Stuart [Flo70a]. Student [Car79b, Car79c, Int64f, Sch66a, Wil73, Ano64, Day72c, Int65a, Int65b, Int67c, Int68g, Int70a, Lin76, Mur77b, Nat73, RS80, Sch66d, Tok68, Zor68]. Studenten [Neb71]. Students [Fre74, Mar80, Baj72, Bre76b, Cla73a, DEN79, Don71, Gen80c, Hut80, MS69, O’D74, Sch79c, Sch80g, SW74, Shn77]. Studie [Hig72]. Studies [Axf72, DM72c, Sch69, Sch70, Sch71, TS73, Wer65, Whi71b, Kar77, Ree73]. Study [HBE80, Hoa72, Hoa73, Knu70, Knu71, Cla73a, Cla80, Dun74, Hew80b, Hal65, Hei72b, Hon72b, Hon72a, Int72l, Lee74c, McA77a, McA77b, NBH70a, NBH70b, Per77, Pin73, Raf79, SA74, Wei67]. Style [Bac78a, LC78, Rul80, DH78, HD78a, KS72b, Tho78, Tip76]. sua [Ano70d, CC70, dC73, Ins64, Ins70, Ins74]. sub [New73, She70a]. sub-group [She70a]. sub-routine [New73]. subcategory [Uni80a]. subcommittee [ABH^{+}71]. subject [SA74]. subprogram [Cor60, Mat72a]. Subprograms [Cli74a, Ste79, SS79a, SS79b, SS79c, Ada78, Bal69, BA73, Hun76, Int68k, Int71c, Int71h, Int72m, Int75f, Kal72b, Kau65, Kri71, Law77, SSS78, SS75a, SS75b, Law78, LHKK79a, LHKK79b]. Subroutine [Abd80, Gaf80, Ker80, MP73, MC80a, MC80b, MC80c, Roc70, She78e, She78d, Ste78a, Ste78b, ZN79a, ZN79b, Bal73, Blo68, Boy75, Bra72a, Col80a, Dav72b, Dil79, Fic71, Fle70, Fox78a, Hon75e, Int65d, JK78, Kah66, KM73a, MP72, MR78, Pow68, Pow69, Pow70a, Pow70b, Pow70c, RG77, Rob68, RS72, Rul68b, Sca70, Sch62, She78b, She78c, Tro66b, TC75, Wes69, ZT76]. Subroutines [Gaf77, Hel63, Sha65, Ste76b, Ste76c, Zoh80, BBB64, Bar61, BS61, Bus67, CR71, Dig75f, DO79, Duf77b, Duf77a, Duf80, Edw76b, Fic73, FM76, Fle72, Gaf79, HB63, Hon77a, Huy77, Int65d, Int80b, LBG66, Mac64, Mar71, Mue66, Mye73, Nic78, RP74, Rei72a, SA74, Sca71, Sou67, Sou68, ST73b, Sta60, Taj65, TD78, Uni74b, VAB62, Wol68a]. subscripted [HV74]. subscripts [Gar65]. Subset [Hal69, LaM72, Hor68, Spe70a]. Subspaces [Ste78a, Ste78b]. substitution [Cha71c]. subsurface [Whi68]. Subsystem [LS76]. subtended [WCT68]. success [Wil76c]. successives [Lou73]. such [PC78a]. Suggested [Ehr72, Fin72b, Fin72c, Mee72]. suite [Hun76]. summaries [McL73]. Summarization [PB67, LB70]. Summary [Hei64, Hei66, Leu79a, Fed70, Fed87, Gri78, Lou67, Ros78, SYR77, Spe80c]. Summated [Ant80]. Summations [Gut76a]. Summer [McC64a]. Sums [Spi72, Mal70]. SUNY [Han74a]. SUNY/B [Han74a]. SUPER [Tym70, Wer72]. SUPER-CODEX [Wer72]. superset [Tym70]. supersonic [Der64, Joh74]. Supervisor [Wer72]. Supplement [U.S78, Coc60, Dun69a, Ful77, Int8 , Int71c, Int71h, Int71i, Int71j, Sch66c, Sch66d, Sch66e, Sch66a, Sch66b, SS74]. Supplementary [MH78]. supplementing [Irv60]. supplied [Uni69c]. supply [Hol80]. Support [Bur79, DDM^{+}75, Bra76, Cli78a, Cli78b, Dat73, Hon76, Int65e, LaP72, Mil73a]. supporting [JM76]. Surface [Aki74, Her72b, ClW78, Goo64, HG66, Hob67, Hol67, Hol68, Jam66a, KM73b, KM77a, KM77b, Kat77, Las71, Lee69, Pol78, RS72, SD66, SD67, Wri72, Wri77a, Wri77b, Wri77c, Wri77d]. Surfaces [Wei66a, Wei66b, CJM67, EP67, OLS66, OT80]. Survey [BP78, LB70, UK74, Bea75, Dav70, Dav72a, Hal71, dPW80, RW77]. Surveys [SH67, Bom67]. survival [LH65, ZD78]. susceptibility [vM78a]. Swaine [McL73]. swell [Lar69]. SWIFT [Cal72]. switching [U.S61]. Symbol [BBB64, CF60, Mas62, Day70, FP73, Tam66]. symbol-state [Day70]. SYMBOLANG [Ber70b, FPB72, Fin72a]. Symbolic [Mat89, Pat77]. symbols [Kra72b]. Symmetric [ST73c, Kar73, MW71a, Rei72a, SS68a, Tho65]. Symposium [IEE75, IEE78, Tou70, Weg64]. SYMTRAN [Car66]. Syntactic [Bai62, Hea68b, Ovi77]. Syntax [Bur65, Bar72b, Blo71, Bro71b, Can77, Lea64, Ube76]. syntax-directed [Blo71]. Syntaxe [Tho78]. synthesis [BK77, RMM69]. Synthesizer [PJ75]. Synthetic [HPLG79, RB76b, RB76c]. System [Ano68c, Bac54, Bac56, BSK67, BF79, Boi74, Boy74b, Bro73a, CF60, Cra76, DW77, Eld70, Her70, Her71, Hon70b, HN58, IBM54, IBM56, Int59e, Int66b, Int66c, Int66d, Int66e, Int66f, Int66g, Int66h, Int66i, Int68d, Int68e, Int68f, Int68k, Int68j, Int68b, Int71c, Int71g, Int71b, Int71i, Int71j, Int71d, Int72k, Int72m, Int72n, Int74f, Int78a, Ina80a, Ina80b, Jam70, Ker80, Lea67, Lee67b, Les72, lAL72, Mas62, MS74a, MW69, MS74b, Mar66, MR73, MS73a, NY78, Ono79a, Ono79b, OF76, PN68a, PN68b, Ree79, Rus78, RCM66, Sys73a, Sys73b, Sad72, Sch80a, SFIK79, Sof80, Squ70, Wat76, Wed75, Weg64, WDT76, Ano72c, Arc76, BBB^{+}57, Bah69, Bai62, BK77, Bar75, Bar72b, Bar73c, BC77, BC79, BCE77, Blo68, Bra76, Bri68a, Bri68b, Con64c, Con64b, Con64a, Con66b, Cla78, CF71, Cle70, CS61, Dig71a]. system [Dig72b, Dig72c, Dig75a, DP76a, Dar78, DM72a, Leu79a, Leu79b, ES75, EH68, Feu77, FP75, Fis79, FL76, Fox64, FH71, Gen80b, Gen73, Gil76, Hal71, Hea79, Hig79b, Hol77, Hou62, Hug78, Int57b, Int57a, Int57c, Int57d, Int58, Int59a, Int59d, Int59f, Int61b, Int61c, Int61d, Int61e, Int61f, Int63f, Int67a, Int64e, Int65f, Int65d, Int66a, Int70b, Int70c, Int71h, Int71f, Int73b, Int75c, Ins76a, Ins76b, IA78, Irv60, JK74, JM76, Jon64, Jos78, KW71, KG76, LaP72, Lar63a, Mac68a, Mac68b, MI64, McA77a, McA77b, McL73, Mer60a, Moo60, Moo77, Mor79, Mye73, OK72, Pat77, Plu65, PN68c, Rit68, SG67, Sil61, Spe66b, Spe66a, Spe9 , Ste72b, TRW73a, TRW73b, TRW73c, TRW73d, TH64, TB65, Tur69a, Tur69b, Ube76, Vic64, WH73, WM60, Whi76, Wid79, BF79, BP78, CCHT67a, CCHT67b, Cod67, Fis78]. System [FF75, Hon76, HN58, Int65c, Int65e, Int67b, Int68b, Int69c, Int70b, Int70c, Int71h, Int72j, Int73b, Int74g, Int75c, Int75f, Int76, Int77a, Int79, Jon79, Lee72, Mar78a, Neh74, She59, Uni0 ]. System/3 [Int74f]. System/360 [Int66b, Int66c, Int66d, Int66e, Int66f, Int66g, Int66h, Int66i, Int68d, Int68e, Int68f, Int68k, Int68j, Int71c, Int71g, Int71b, Int71i, Int71j, Int71d, Int72k, Int72m, Int72n, Lea67, Lee67b, MW69, Mar66, PN68a, PN68b, CCHT67a, CCHT67b, Cod67, Int65c, Int65e, Int67b, Int68b, Int69c, Int70b, Int70c, Int71h, Int72j, Int73b, Int74g, Int75c, Int75f, Int76, Int77a, Int79, Lee72]. System/370 [Int68d, Int68f, Int71g, Int71b, Int78a, FF75, Int72j, Int74g, Int76, Int77a, Int79, Lee72]. Systemablaeufen [Nie72a, Nie72b]. systemat [Her78]. Systeme [Rin77, Sch80e]. Systems [AI78, Abd80, Bur73b, Bur73a, Don73b, Ful72, Geo80, Int75a, KGY80, Kub73, MC80c, Mor73, Ros66, Ros72, TI72, UK74, Ano68b, Bas80, Bri68a, Bur68b, Bur70b, Con65, Con66a, Con66b, Con67a, Con67b, Con67c, Con68b, Con68a, Con68c, Con69d, Con69b, Con69a, Con70, Con71d, Con71b, Con71c, Con72a, Con73c, Con73d, Con73e, Con73b, Con75a, Con76c, Car74b, Coh66, Edw76a, Gor64, Hew76b, Hal65, HLS73, Int62b, Owe62, Int62d, Int63o, Int63p, Int63f, Int63e, Int63g, Int63h, Int64g, Int64c, Int64d, Int66a, Int76, Int77a, JM76, Pec77, PMBK80, PC78b, Pow68, Pow70a, Raf79, Sch80f, Sch68, Sig80, Sli71, SM73b, Van66, Vas72, ZT76]. szamitogepes [Kor77].

T [Hen67]. T. [Nak68]. T3 [Lar67a, Lar67b]. Table [IR78, Kah80a, Kah80b, WDT76, CB69, Fic73, Gul71, Hil79c, Lil68, Rey68, SK69, Sch72c, Sea80, Thr79]. TABLE-2 [Thr79]. Tables [Bak77b, Hab72, Hab73, MH75a, MR70b, Day70, Mue75, Vei66]. Tabulating [Lea78]. tabulation [LB70]. Tail [BR75a]. Talmi [Zoh72]. Tandem [Tan80a, MC70]. tangential [And73]. tanker [SYR77]. TAP [Wes69]. tape [Int67a, TB65]. tapes [Gof74, McA77a, McA77b, RPE79]. Task [Sta74]. Taught [Far66]. Tausworthe [Pay70]. Taylor [Bar80c, Lou73]. TCDMS [Int75a]. TDC [McG67]. Teacher [Sch66b, Haa69b, Jac73c, MHH71b, MH73, Sch66e]. Teaching [Bez75, Emb78, FK77b, NS76, Tay77, And70, CDH75, DEN79, Fri75a, FK76, Hal65, Ham74, Kha77, Kno72, Mau77, McA77a, McA77b, MH75b, MK73, OR77, Phi71a, Phi71b, Pin73, Plu61, Plu63, Rad76b, Sch79c, Sch80g, Sil61, Spe78a, Wid79, Wit79b, Wit79a, Wit79c, Wit79d]. technical [Ano70a, Wor76a, Xer71c, Xer73, Xer74b, JMG77, U.S61, Wor76b]. technicians [Edw69]. technika [ATW77]. Technique [AS71, Lap78, Cle66, DM66c, Han74b, Mac67, SS68a, ZT76]. Techniques [Bre73, Day72a, Day72b, Gow73, Hil73, HPR78, KM73c, LV73, Sea79b, She70b, Wil75, Bar73a, Con76d, Con77c, Day79, Dea71, Dea77, HPR77, Int71a, KS72b, LT75, Mac69, O’K64, Sea79a, Cha73, Pat73b]. technischen [Mac70a]. Technology [Fre74, Baj72]. Tedious [GIB65, Ver65]. Tekniske [Tju68]. tekunikku [Mur70]. tele [BR78]. tele-type [BR78]. Telecourse [MS69]. television [Day72c, Roh73]. temperature [Hol67, Hol68, Kot72]. temperatures [Mei78]. Ten [BK75, KS68, KS70, Ken74, KS75a, Roh73, Sti72]. Tension [Cli74a, Cli74b, Cli74c]. tensoriels [Ray63]. Teoria [BG78, LP79, Rid68, Rid69, Rom75]. TEP [Joh65a, Joh65b, Joh76]. TEP-II [Joh76]. termes [Ray63]. Terminal [Uni0 , Coh74, Int71i, Int71j, Int71d, Int72l, Int75e, Int75d, Lat79, Rad70, SW75]. terminal-oriented [Int72l]. terminals [BD80a, BR78]. terms [MA78]. ternary [DD68]. terrain [BC72a, Plo77]. tervezese [Kor77]. test [Ano74b, Ano74c, Ano74a, HP74, Jul75, Mar78a, OK72, RCL75, Sch72c]. Testbed [HH79a]. Testing [McC74b, Ano70c, Fan65, Gut79a, Gut79b, Ost62, Ove72, PB74a, Smi80, Tri79, U.S61]. Tests [Tho72a, Pre70]. tetrachoric [MW75]. TETREST [MW75]. Texas [Axf72, IEE75, IEE79, Mor79, Wed75]. text [Ano72b, BP74, Dun69b, Hal65, HDBP68, Int64f, Int65a, Int65b, Int67c, Int68g, Int68i, Int70a, Int72l, Int79, Tab66, Kal72a, Lem75, MM78, Nat73, Phi71a, Phi71b, Sch66d, Sch66a, Tho72b]. textbooks [Shn77]. texto [SJ72, SS78b]. Their [KMC72, LR77, Gav76, Håv71, Han78, MB68a, MB68b, Kan68, Ste72a]. Them [DA68]. Themen [Tec72]. theoretical [CK80]. Theorie [VHP69, VL72, LPJ79a, LPJ79b]. Theory [GIB65, Gil60, LP78, LF75, SA74, Sho80, Smi70a, Smi70k, Smi72a, SW79, ZD78]. there [Smi70c]. thermal [BM74, BD71, Joh65a, Joh65b, Joh76]. thermal-ellipsoid [Joh65a, Joh65b, Joh76]. thermalization [Sta65]. thermochemical [Ste72a]. thermodynamic [SM73b]. thermoelastic [Hol67, Hol68]. thermometry [Har66a]. THERMOS [Sta65]. thesis [Dar78, Gel69, Kei69]. Thiele [Lou73]. thin [MC64, PNK65a, PNK65b, Spe66a]. thin-film [Spe66a]. thinning [NS69]. Third [Tou70, Han78, Kan71]. Three [Bre67, FRS77, MR70a, PT73, SD67, Sta60, Edu72e, JK78, McC71, OT80, RG77, RS72, Spi65, Wat73a, Wat73b]. Three-dimensional [SD67, JK78, OT80, RG77, RS72, Wat73a, Wat73b]. three-point [McC71]. TICK [Moo77]. tide [PC67, RZB77, Sho80]. TIDY [Mur66]. Tien [Ano75b]. Time [BCKT79, Bre73, CS73a, FH74, Gen67, Gen66c, Hon73d, Hon70b, Ing71, NC76, Rob67a, RCM66, Sch76, Sch73, Sim66, Uni69a, dL78, Ack64, BS78, Blo68, BR78, Con73f, Con73g, Con74, Clo72, Dig71a, Dig72b, Dig72c, DS67a, Fox64, Gra70b, HB63, Hea68b, Hin76, Hon71b, KW75, Kru68, KRB78, Mac64, Mar66, MU75, Mod74, Pin73, PMBK80, PC67, RB76a, RBK76, Upc72, Wes69, Whi76, Wil72a, Wit74, ZSW76, ZSW77, ZSW79, Zin79]. time-dependent [PMBK80]. Time-series [Sim66]. Time-Shared [Sch73, Hea68b]. Time-sharing [Gen67, Gen66c, Hon73d, Hon70b, RCM66, Uni69a, BR78, Con73f, Con73g, Con74, Gra70b, Mar66, Pin73, Wes69, Wit74, ZSW76, ZSW77, ZSW79, Zin79]. time-trend [DS67a, Fox64]. Time-Varying [Sch76]. times [LH65]. Timeshare [Ano70e, Chi73]. timings [Sit78]. title [Law79]. TMS [Hea79]. Toeplitz [Zoh80]. Tolerance [BS75]. tool [BT76b, Obr70, Obr71]. toolkit [Khu68]. Tools [Cha79a, Fel79, KP76, Lea67, Pre79, Sit78, SR76, Smi80]. Top [CP80].

Top-down [CP80]. topics [Mis78a]. topography [Hob67, Plo77]. tops [MW71a, MW71b]. Torso [vO78]. total [BM74, SMD71, SR73]. totally [FMC78]. touch [Sha77]. tracers [U.S61]. TRACK [Mac68a, Mac68b]. Traffic [Hon76, Leu79a, Leu79b, DP74b, TRW73a, TRW73b, TRW73c, TRW73d]. train [Tho66]. trainer [Tho66]. Training [Wol68b, Con67a, Pec77, SJ62a, SJ62b, SJ63, Smi73a, Pau71a, Pau71b]. Trains [Her72b]. trajectories [Lyn63]. Trandes [Car77]. transactions [CB69]. Transcendental [Ein74, FSC73]. transcribing [Rub69b, Rub69a, Rub69c]. Transfer [BB77b, Sei75, Bec72, Fla71, Hol67, Hol68, PV74]. Transfer-function-parameter [Sei75]. Transferability [Uni75a, Can77]. Transferable [Roc70]. Transform [Bre67, DA68, Mon75, Rej72, Clo72, Fis70]. transformation [LML69]. Transforms [DA68, LS75, Mur71, Pic66]. transgression [Kru68]. transient [AJ69, And73, Bak68]. transients [EKM74, Saw62]. transition [Bro71b, MU75]. translate [Mue75]. translating [Bar73c, IBM54]. Translation [AK80, Fut78, Sim76b, Sim76a, Str78, Ken70, Ken80, Lea67, PH77, SWL68, Som71]. translations [DS62, Mor71, SMD71]. Translator [CCN^{+}79, CNW80, MM65, Pit79, RSD65, She59, Wil67, Bon75, Con71d, Gul71, Hig75, Hil69, Lea67, Lea75, Pul64, Spe74a]. transmit [Coh74]. transonic [AK77, Car77]. Transport [Pat74, DW70, Fro63, Sho80, SM73b]. transportability [CW78b]. transput [Car78b]. trapezoidal [Pic66]. trapiantabili [SS68b]. Trawl [PB67]. Treating [BF71]. Tree [Han72a, Han72b, Pag74a, Whi72, Ant77, CGH75, Day76]. TREETRAN [FPB72, Fin72a]. trend [DS67a, ESD68, Fox64, Fox67, Goo64, Har68a, HG66, Lee69, OLS66, SD66]. trend-surface [Goo64, HG66, Lee69, SD66]. triangle [Hae77]. Triangulating [vO78]. Tridiagonal [Kub73]. TRIGMAN [Jef72]. trigonometry [Her72a]. Triplex [BJ77, BY78]. TRS [Rad79, Rad88]. TRS-80 [Rad79, Rad88]. Truck [Rey68]. Truck-weight [Rey68]. TRY [TB65]. TSO [Int71i, Int71j, Int71d, Int72a, Int72b, Int72g, Int74e, Int75b, Int75e, Isa73]. tubes [Fla71]. Tucker [Yoh72]. Tukey [Bre67]. tumori [SS68b]. turbine [FH71]. turbines [WG75]. turbomachine [KM73b, KM77a, KM77b, Kat77]. turbomachinery [MC70]. turning [MM58]. Tutor [Spi80, Lin76]. Tutorial [CZ72, EH68, Mul80a]. tutorials [Mis78a]. Twenty [Int57e]. Twenty-five [Int57e]. twin [SDH74]. TWINSPAN [Hil79c]. TWISK [NM78, Nav78]. Two [Bak77b, BF71, Cla73b, Fri71a, Fry71, Kal72b, PG65, PRO80, Rei72a, RST78, Sho80, Tan80c, Tho72a, Wal68, And70, Bar73b, Cam77, Dav70, Dav72a, Dey76, Dif72, Edu72f, ES78, Erd80, Ham74, Hem70, HO64, Hil79c, Hun74, Joy78, KR69, MR78, MA78, Phi67, SS78a, Sid72a, Sid72b, SMM65]. two-body [Phi67]. Two-Dimensional [BF71, Cam77, Erd80, Hem70, HO64, SS78a]. Two-Hit [Tho72a]. Two-level [Cla73b, Sid72a, Sid72b]. two-semester [ES78]. two-voiced [Hun74]. Two-Way [Bak77b, RST78, Hil79c]. Type [Jet79, TT80, BR78, CCL69, CL70, SMD71, SR73]. types [ABH^{+}71, Gum77]. typesetting [JK74]. Typographic [Her70, Her71]. typography [Her69]. Tzschach [Din72, Jun69].

U.S. [Sad72]. ubersichtlicher [KKU78]. UFORT [CNW80]. uloh [Ham79b]. UMASS [Mas71]. Unbalanced [RST78, Bry75]. unconsolidated [PG66]. unconstrained [Pow70b, Pow70c]. Undergraduate [AI78, Ham74, HBJ76].

Understanding [Boi78, BS80b, McC68b, Wat76]. Underwater [Gen66a]. unfort [Emb78]. uniform [O’D65, Tri79]. unique [Ent80b]. UNIRUN [Deu73]. Unit [Axf72, Pag74b, Dat73, Phi71a, Phi71b]. Units [Ful72, Soy71]. Univac [Jak73, Spe74b, Abr65, Ber70a, Ful73, Mor75, Sou71, Spe70b, Cha66, Hog70, War69]. univariate [Fin68, Fin72d, Fin72f, Fin72e, Fin77, Wal68]. Universal [CNW80]. University [Axf72, Bar73a, Bar80b, Eva72a, Eva72b, IEE75, Pat73b, Pec77, Blo68, Hed77, RSD65, SGM^{+}67]. UNIWAFT [Rei72b]. UNIX [Ube76, Raf79, Win75]. unknown [JV67a, JV67b, JV68]. unknowns [Dif72]. unofficial [Ame78f]. unrestricted [SD66]. Unrolling [DH79]. unsteady [AK77]. unsymmetric [Duf77b, Duf77a, Duf80]. Untereinander [jH78]. Unwanted [Par75]. upon [Kal72a]. Upper [BR75a, Ste76b, Ste76c, FB73]. Uprazhneniia [Lam78, DG78]. upward [Rap66d, RBp75]. upwelling [Sho80]. Urban [Leu79a, Leu79b, TRW73a, TRW73b, TRW73c, TRW73d, Hon76]. USA [ACM78, Dig69, Lew79b, AB66b, Ame66c, Ber70a, Dat66, IEE78, Ame77, Nik78, Spe69a, Spe69b, Tou70]. Usage [Law78, LHKK79a, LHKK79b, Bur68a, Lat79, Law77, Tho72c, Wei69]. USAS [Cal69b, Cal69c, Cal69d, Cal69e]. USASI [Con69a, Con70]. Use [Bau56, BF79, Han60, HPLG79, HBJ76, LR77, Ver65, Arn65, Bar61, Ber64, BLY70, BR78, Day70, Dra64, Her74, Kno72, LGF75, MB68a, MB68b, Maz78, MCB^{+}62, Oer71, Owe79, PB74a, Sco76b, Sha65, Sid72a, Sid72b, Smi67b, Squ70, Taj65, Uni78, Wit74, ZT76]. used [Bal69, Dav72b, Eld77, Gra70b, LaP72, McA77a, McA77b, Mer60a, Per80, PT73, Wit79b, Wit79a, Wit79c, Wit79d]. useful [Wil76c]. User [AK77, Boi74, Bro71c, Bro73b, Dat75, Dey76, Fin77, Fis76, Flo78b, Fly73, Hun76, JCMS76, JCMS77a, JCMS77b, MI80, Mor73, Ryd3 , Sch75, TB65, Ari76, AD73, Bac72, Bla60, Boy75, Con72b, Con73g, Con76b, Con77b, Cal69b, Cal69c, Cal69d, Cal69e, Car74b, CF71, Dig75c, Dig75g, Dig76c, Dig77f, Dig78a, Dig78c, Dig79a, Dig80c, Dig80f, Dat77a, Dat77b, Ent80a, FMC78, Gre79, Hon70a, Hon70b, Hon79b, Int71i, Int71j, Int71d, Int75e, Int75d, Mas71, Mic79b, Nat70a, New73, Pat74, PJT76b, Rad70, Sho76a, Sho76b, Sik71, Sli71, Sof80, Ste76a, TC75, Uni79, Uni80b, Van73b, ZN79a, ZN79b]. Users [Joh80, JID80, Lew79b, Con73d, Con76a, Con76c, Con79a, dPW80, Int67c, Int68g, Int8 , IBM68, Int68h, Int68i, MS79, Mor75, Tur69b]. uses [Mue75]. Using [BD80a, dlB59, Cli74a, CA78, Con79e, Fel79, HPR78, Joh66b, Lat79, LHLM80a, LHLM80b, Lyo80, Moh77, NC76, PRO80, RJAS78, Wal80b, Art75, Bar66, Bla68a, Bla68b, Bla71, BC67, Bro71b, CS73a, Cli74b, Cli74c, Clo72, CJM67, CDGW76, DS66, Dun67, Dun69b, Dun69a, DLS79, ES78, Fox64, FK76, Fri71a, FP73, Gav76, Har68a, Har63, Har64b, Has78, HV74, Hat78, HH77a, Hor65, HPR77, Jam66a, JV67a, JV67b, Joh74, KPG63, KF72, KS68, Las71, LK74, Lea67, Lep76, Les72, Lil71, Lou74, MI64, MT75, Mik73, Mou70, Mur71, Ost64, PTM77, PC78a, PNK65a, PNK65b, Pic66, PG66, Ren65, RG76, Sch79c, Sch80g, Smi66, Sti72, SW75, TI72, TS73, Tho72b, Tho68, Tug75, Van73a, Var77, Wal80a, ZD78, Zal73]. ustoichivogo [SZ80]. usuarios [Wei73]. UT [Uni77]. UT-CDC [Uni77]. Utah [Bee80d]. UTCS [Hon76]. Utilisation [Rou75, É67, Isa73]. Utilities [Cor77, FHS7x, Dig75e, Dig77e, Ent80a]. Utility [Bee70c, Bee78, CR74, Lud69]. utilizacao [CC70, Ins64, Ins70, Ins74, Ano70d]. utilizacion [Oli71]. Utilizing [Jon79, O’D65]. Utopia [AI79].

v [Ham79b, Ano67, Ano70c, Bec72, Ber70a, CGG^{+}91, Fin72f, HM64, MT72, Ove72, Ree68, Ree71, Sch67, Sou71, Spe66b, Spe69c, Spe9 , Spe73a, Spe73b, Spe76b, Uni68a, VP76, Xer76a]. V001C [Fis76]. V02A [Hun76]. Validation [OF76, Fed70, Fed87, SF75]. Value [AK78, ABB^{+}74, CR74, LP74]. valued [Cli74b, Cli74c, RS72]. values [PC67, RZB77, Spi65, VV66]. vane [Mei78]. vapor [FH71]. vari [Sic74]. Variable [Boh75, dlB59, Jet79, Sch76, Yat71a, Yat71b, Yat71c, Art75, Han75, Jef77, LP74, Lou73, Tay76]. variable-length [Han75, Tay76]. Variables [Fia73, HK72, Lea78, PG65, Bai63, BT76b, Bro80, Cam65, Gum77, Har73, SMM65]. variance [Bry75, Fin68, Fin72d, Fin72f, Fin72e, Fin77, Gar63, Kan79, She70a]. variates [RR70]. variation [Hol67, Hol68]. varierande [Hus76]. varimax [Mat72a]. Varycord [Bom67]. Varying [Sch76]. VAX [Dig78b, Dig78c, Dig80d, Dig80e, Dig80f, Dig80g, Raf79]. VAX-11 [Dig78b, Dig78c, Dig80d, Dig80e, Dig80f, Dig80g]. VAX/UNIX [Raf79]. VAX/VMS [Raf79]. ve [Yur76]. Vector [AK80, Blu7 , Blu78, McL78, Sch72a, Zwa75, Fox67, Han78, Ken70, Ken80, MI64, MS78, MS77a]. Vectorization [Hig79a]. Vectors [ST73c, ST73b]. vegetable [Oer71]. vehicles [PG67, U.S61]. velocities [KM73b, KM77a, KM77b, Kat77]. ventilation [FMC78]. Verfahren [Sto76]. Vergleich [MSS78b]. vergleichende [Hig72]. verification [Smi80]. Verifier [HR73, Ryd74, Ryd72, Ryd3 ]. Versatile [Whi71a, Cra79]. Version [Hil70, Sch78b, Yoh79a, Ari76, Ano74b, Ano74c, Bla69, Con71f, Con72a, Con72b, Con73e, Con73f, Con73g, Con75a, Con75d, Con75e, Con75f, Con76a, Con76b, Con77b, Con78a, Con78b, Con79a, Con79b, Con79c, Con79d, Con80b, Con80c, Dig80g, Leu79a, Leu79b, Edw76b, Fin72f, Fis76, Fri71b, Fri80, Hun76, Int66a, JV67b, JV68, JMG77, Kno75a, Kno75b, LO77, MS71, Nat70a, Pat74, PT68, PT69, PJT76b, PJT76a, Ste80, SH78, Squ70, Sta65, Ste60a, Ste72b, TB65]. versions [Ano74a, Con72b, Con73c, Mur77a, Mur77b]. versus [BD80a]. vertical [And73]. Very [Ein76, Hei74]. VI [HS69, Deu73, HL70, Hea68a, Wel70a]. via [Bid79, DD68, ES74b, Emb78]. video [Hon73b, McA77a, McA77b]. video-assisted [Hon73b]. video-tapes [McA77a, McA77b]. vierdimensionale [Ant72]. view [Sli71]. Views [Wei66a, Wei66b, OT80]. Violation [Ein76]. Virtual [Mor73, dC73, Hug78]. Visible [Wri72]. Vision [Arm78, Wei66a, Wei66b]. Visual [Smi72d, Spe78a, MCB^{+}62, SW74, WS73]. VLF [Wat75]. VM [Hug78, Int72b, Int72c, Int72d, Int72e, Int72f, Int72g, Int74a, Int74b, Int75b, Int75e, Int75d, Uni79, Uni80b]. VM/ [Int72b, Int72c, Int72d, Int72e, Int72f, Int72g, Int74a, Int74b, Int75b, Int75e, Int75d, Uni80b]. VM/370 [Uni79]. VMS [Raf79]. Voice [Eld70]. voiced [Hun74]. VOIFLO [Sol64]. Volume [Ano77a, MH78, Leu79b]. votes [Van73a]. vs [Gor64, Pec77, Pin73, Tho72a, Int80a]. VS-Programmen [Rin77]. vyborki [AE79]. vycislenij [FMM80].

W [Kar77, Kre66b, Ree73, ATW77, Bad77, BF72, Mar78a, Rey68]. W-3 [Rey68]. W. [Flo70b, Jun68, Kre66a]. Warteschlangensysteme [Ger80]. Washington [Lew79b]. WASP [HPB73, PB73b]. Wasserman [VP80a, VP80b]. WATBUG [Wil77b, Wil77c, Wil77d]. WATEQ [PJT76b, PJT76a]. WATEQF [PJT76b, PJT76a]. water [Bak68, GKB74, HPB73, Lal75, NM70, NM78, Nav78, Nor66, PB73b, PC67, RMM69, Sal70, Sho80, Wil77b, Wil77c, Wil77d]. water-quality [RMM69]. Waterloo [SGM^{+}67]. waters [Lar69, PJT76b, PJT76a]. Waterways [CS73a]. Watf [Con79e]. Watfiv [Ree75, BS73a, BP74, BS80b, BK75, CS71a, CS71b, CS71d, CS75, CW72, CW73b, CW73a, CW76, CW77, CW78a, CW79, CP80, Col76, CDG73, CDG80a, CDG80b, DH78, DLS79, For75, Gea78, His75, HD78a, HH77a, HH80, HH77b, HH78, Jaf79, KS70, KS75a, KS75b, MS77b, Mes74, Moo75, MM78, MT80, PD76, PB73a, Ste73, Tel80, Tho71, Tho72b, TB80, Bla71, CS71c, CA78, CDG70, Stu71, Wal75, Ree76]. WATFIV-S [Col76, CDG80b, DH78, DLS79, HD78a, HH77a, HH80, HH77b, HH78, TB80]. WATFIV/WM [His75]. WATFOR [Ree76, Ree75, BS73a, BP74, Bla68b, BK75, CW73a, Col76, CDG68, CDG73, CDG80a, DH78, For75, HD78a, Jaf79, KS68, KS70, KS75a, KS75b, MS77b, SGM^{+}67, Tho71, Tho72b, Uni69c, Bla68a, Bla71, CDG70, Ken74, Stu71, Wal75]. WATFOR-llS [Col76]. WATFOR.WATLIB [Uni69c]. WATFOR/WATFIV [KS75b, Bla71, Stu71, Wal75]. Wativ [Ken74]. Wave [Her72b, Gen66a, HM62a, HM62c, HM62b, HO64, Her64, HM64, Lal75, HM62a, HM62c, HM62b, HM64]. waveforms [Lov68]. wavelength [Var77]. Way [Bak77b, RST78, Hil79c, Kan79, She70a]. weather [LGF75]. weathering [Col80b]. Weber [VP80a, VP80b]. Weight [AB69a, Swi64b, Swi64c, Rey68, Swi64a]. Weighted [PG65, Weg66]. wells [Cas62]. Western [IAAA57]. Westinghouse [Bri79]. Westinghouse-Bettis [Bri79]. WFLASH [EKM74]. Which [Par75, Sal77b, Bak75, Gel69, Mue75, Sal70, Sal78]. Whitney [DB73]. whose [Rei72a]. Wigner [AD73, CM66]. Wijngaarden [Mac73]. Wiley [Ree73, Ree75, Sch80c]. Wilf [EE77]. Wilf-quadrature [EE77]. Wilk [Mar78a]. Wilkes [Ree75]. will [Bid79]. William [Kar77]. wind [Lal75, Sho80]. wind-driven [Sho80]. wind-wave [Lal75]. windows [Bar80a]. Wirtschafts [Neb71]. Wirtschafts- [Neb71]. Wisconsin [Pat74]. Within [Sal77b, Bai62, BM74, Bar61, CF60, Har65b, Mas62, MCB^{+}62, RP74, Sal78]. without [Bod77, DEN79]. WM [His75]. word [Krá72a]. word-length [Krá72a]. words [Ost62]. Workbook [Car79b, Gue73a, Car79c, Mur77b, Spe77b]. Works [Win75]. Workship [Jet74]. Workshop [FJA80b, FJA80a]. World [Eva72a, Eva72b, HS69]. Would [Hul73]. wraz [ATW77]. write [Hea68b]. writer [Bar77b]. writing [CS73a, Gra70a, HV74, Knu62, Sch62, Wat68]. Written [Lea67, Dea71, Dea77, McC64a, Pat77, Rod76, Squ70]. WVONB [PT73].

X [Cri77, Day63, Maz77, Moo76, Oer71, Smi63b, SH78, Ste72b, TC70, vM76]. X-ray [Cri77, Day63, Oer71, Smi63b, SH78, Ste72b, TC70, Maz77, Moo76, vM76]. X3 [Ano77c]. X3-9 [Ano77c]. X3.9 [Ame78e, Ame78d, U.S78, Ame66a, Ame78d, Ano78d, SIG76, Ame66b]. X3.9-1966 [Ame66a, Ame78d, Ame66b]. X3.9-1978 [Ame78d, U.S78, Ano78d]. X3J3 [Ame78f, ABH^{+}71, Ano77a, SIG76]. X3J3/ [Ame78f]. X3J3/76 [SIG76]. XDS [Ano70a, Ano70b, Owe79, Xer70a, Xer70b]. Xerox [Ano75e, Ano70f, Ano70g, Ano73, Ano75c, Ano75d, BR78]. XFOR [Mos78]. XFOR-80 [Mos78]. XLFIT [Wei75].

years [Int57e]. yen [Ano75b, jT79]. yield [PT68]. yoru [Mur70]. yu [Ano75b, jT79]. Yugoslavia [FGR72].

z [Pic66, FHS78, Kah80b, Pag74a, Yoh72]. ztransforms [Pic66]. zadach [SZ80]. zadaniach [ATW77]. ze [ATW77]. Zehnder [Din72, Jun69]. zero [Gar65]. Zeros [JT72, Spi65]. zeta [Spi65]. Zinbun [Nis78]. Zugangsstatistik [Sto76]. zum [Dre70]. Zur [Rin77, Tec72, Die72, Fri75b, RS69]. Zuwachsverzeichnis [Sto76].

References

ANSI:1966:ANS

[AB66a] American National Standards Institute and Business Equipment Manufacturers Association (U.S.). American National Standard Basic FORTRAN. The Institute, New York, NY, USA, 1966. 32 pp.

ANSI:1966:USF

[AB66b] American National Standards Institute and Business Equipment Manufacturers Association (U.S.). USA standard FORTRAN Approved March 7, 1966. United States of America Standards Institute, New York, NY, USA, 1966. 36 pp.

ASA:1966:ASB

[AB66c] American Standards Association and Business Equipment Manufacturers Association (U.S.). American Standard Basic FORTRAN. ASA X3.10-1966. American Standards Association, Inc., New York, NY, USA, 1966. 32 pp.

ASA:1966:ASF

[AB66d] American Standards Association and Business Equipment Manufacturers Association (U.S.). American Standard FORTRAN. ASA X3.9-1966 American Standards Association. American Standards Association, Inc., New York, NY, USA, 1966. 38 pp.

ASA:1966:BFA

[AB66e] American Standards Association and Business Equipment Manufacturers Association (U.S.). Basic FORTRAN Approved March 7, 1966. American Standards Association, Inc., New York, NY, USA, 1966. 32 pp.

Anton:1968:FBD

[AB68] Hector R. Anton and Wayne S.

Boutell. Fortran and business data processing. McGraw-Hill accounting series. McGraw-Hill, New York, NY, USA, 1968. ix + 348 pp. LCCN HF5548.5.F2 A6.

Agrawala:1969:WDL

[AB69a] Vishnu K. Agrawala and Johan G.

Belinfante. Weight diagrams for Lie group representations: a computer implementation of Freudenthal’s algorithm in ALGOL and FORTRAN. BIT (Nordisk tidskrift for informationsbehandling), 9(4):301–314, December 1969. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL

http://www.springerlink.com/
openurl.asp?genre=article&issn=
0006-3835&volume=9&issue=4&
spage=301.

Anton:1969:ABF

[AB69b] Hector R. Anton and Wayne S.

Boutell. Answer book for Fortran and business data processing. Mc-Graw-Hill, New York, NY, USA, 1969. iii + 130 pp.

Aird:1974:NSV

[ABB^{+}74] Thomas J. Aird, Edward L. Battiste, Nancy E. Bosten, Henry L. Darilek, and Walton C. Gregory. Name standardization and value specification for machine dependent constants. ACM SIGNUM Newsletter, 9(4):11–13, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Abdelmalek:1980:AFS

[Abd80] Nabih N. Abdelmalek. Algorithm 551: A FORTRAN subroutine for the L_{1} solution of overdetermined systems of linear equations [F4]. ACM Transactions on Mathematical Software, 6(2):228–230, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ahl:1971:CBD

[ABH^{+}71] Kenneth Ahl, John Barrington, John Hillier, Evelyn Mack, and Walter Whipple. Character and BIT data types for FORTRAN: a proposal to ANSI subcommittee X3J3. ACM SIGPLAN Notices, 6 (10):22–40, November 1971. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:/

                                        /dl.acm.org/citation.cfm?id=
                                        1317450.

Abramson:1965:BGC

[Abr65] Norman J. Abramson. Von Bertalanffy growth curve II, IBM 7094, UNIVAC 1107, Fortran IV. Transactions of the American Fisheries Society, 94(2):195–196, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abrahams:1972:CSF

[Abr72] John R. Abrahams. Computer science with FORTRAN. Computer studies series. Griffin House, Toronto, Ontario, Canada, 1972. ISBN 0-88760-026-3. 147 pp.

Ackermann:1964:FIL

[Ack64] A. Frank Ackermann. A Fortran II load-time saver. Comm. ACM, 7(5):310, May 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ACM:1976:DPA

[ACM76] ACM. Draft proposed ANS FOR-TRAN. ACM SIGPLAN Notices, 11(3):1–223, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1978:ASH

[ACM78] ACM, editor. ACM SIGPLAN History of Programming Languages Conference. Los Angeles, CA, USA, 1–3 June, 1978, volume 13(8) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, August 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1979:PSC

[ACM79] ACM, editor. Proceedings of the SIGNUM Conference on the Programming Environment for Development of Numerical Software. ACM Press, New York, NY 10036, USA, 1979.

Akiyama:1973:UGF

[AD73] Yoshimi Akiyama and J. P. Draayer. A user’s guide to Fortran programs for Wigner and Racah coefficients of SU_{3}. Computer Physics Communications, 5 (6):405–406, June 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465573900775.

Adams:1978:FSF

[Ada78] John C. Adams. FORTRAN subprograms for finite-difference formulas. Journal of Computational Physics, 26(1):113–116, January 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999178901031.

Ahrens:1970:PRN

[ADG70] J. H. Ahrens, Ulrich Dieter, and A. Grube. Pseudo-random numbers. A new proposal for the choice of multiplicators. Computing, 6 (1–2):121–138, March 1970. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Ammerman:1967:DGD

[ADT67] Anne B. Ammerman, Larry Diesen, and Hermon W. Thombs. Displaytran: a graphical display oriented conversational FOR-TRAN facility for an IBM 360/ 40 computer. Technical memorandum (Naval Weapons Laboratory) K-39/67, The Laboratory, Dahlgreen, VA, USA, 1967. 105 pp.

Aprausheva:1979:AMM

[AE79] N. N. Aprausheva and E. B. Ershov. Algoritm modelirovaniia mnogomernoi vyborki, predstavliaiushchei soboi smes K normalnykh klassov: FORTRAN, ES-1022. Programmy i algoritmy vyp. 90, TSentr. ekonomikomatematicheskii in-t AN SSSR, Moskva, USSR, 1979. 14 pp.

Alefeld:1980:FNC

[AG80] G. Alefeld and R. D. Grigorieff, editors. Fundamentals of Numerical Computation (Computer-Oriented Numerical Analysis), volume 2 of Computing. Supplementum. Springer, Wien / New York, 1980. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel.

Arden:1961:AED

[AGG61] Bruce W. Arden, Bernard A.

Galler, and Robert M. Graham. An algorithm for equivalence declarations. Comm. ACM, 4(7):310–314, July 1961. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aghakhani:1977:GFC

[Agh77] Mohammad Kazem Aghakhani. A general FORTRAN computer program for Markov chains. Thesis (m.s.), Department of Industrial Engineering, Wichita State University, Wichita, KS, USA, December 1977. vi + 68 pp.

Alford:1977:SEA

[AHP77] M. Alford, P. Hsia, and F. Petry. A software engineering approach to introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1): 157–161, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

ACM:1978:LLU

[AI78] ACM and IEEE. A library list on undergraduate computer science, computer engineering, and information systems. Technical report, ACM Press, New York, NY 10036, USA, 1978.

Arisawa:1979:FPU

[AI79] Makoto Arisawa and Minoru Iuchi. Fortran + preprocessor = Utopia 84. ACM SIGPLAN Notices, 14 (1):12–15, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arisawa:1980:DMR

[AI80] Makato Arisawa and Minoru Iuchi. Debugging methods in recursive structured FORTRAN. Software —Practice and Experience, 10(1): 29–43, January 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aird:1977:PMS

[Air77] Thomas J. Aird. Portability of mathematical software coded in Fortran. ACM Transactions on Mathematical Software, 3(2):113– 127, June 1977. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Abraham:1969:FPC

[AJ69] Farid F. Abraham and Stanley K. Jordan. Fortran program for computing the transient and steady-state modifications of a raindrop size distribution due to evaporation and coalescence. Journal of Computational Physics, 4(1):144–145, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999169900461.

Alzner:1977:CUT

[AK77] Edgar Alzner and Paul P. Kalben.

Computation of unsteady transonic flows through rotating and stationary cascades: II — user’s guide to Fortran program B2DATL. NASA contractor report NASA CR-2901, National Aeronautics and Space Administration, Scientific and Technical Information Office, Washington, DC, USA, 1977. various pp. For sale by the National Technical Information Service.

Armstrong:1978:SAA

[AK78] Ronald D. Armstrong and Mabel Tam Kung. Statistical algorithms: Algorithm AS 132:

Least absolute value estimates for a simple linear regression problem. Applied Statistics, 27(3): 363–366, September 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

                                        URL http://lib.stat.cmu.edu/
                                        apstat/132.

Allen:1980:ATF

[AK80] John R. Allen and Ken Kennedy. Automatic translation of Fortran programs to vector form. Rice Technical Report 476-029-4 (rereleased as Rice COMP TR84-9, July 1984), Rice University, Houston, TX, USA, October 1980.

Akima:1974:AAB

[Aki74] Hiroshi Akima. ACM Algorithm 474: Bivariate interpolation and smooth surface fitting based on local procedures [E2]. Comm. ACM, 17(1):26–31, January 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [?].

Allen:1966:DAD

[All66a] K. Radway Allen. Determination of age distribution from age-length keys and length distributions, IBM 7090, 7094, Fortran IV. Transactions of the American Fisheries Society, 95(2):230–231, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1966:FBG

[All66b] K. Radway Allen. Fitting of von Bertalanffy growth-curves, IBM 709, 7094, Fortran IV. Transactions of the American Fisheries Society, 95(2):231–232, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alluisi:1967:BFS

[All67] Earl A. Alluisi. Basic Fortran for statistical analysis. Dorsey Press, Homewood, IL, USA, 1967. vii + 126 pp. LCCN QA76.5 .A38.

Allensworth:1975:SPF

[All75] James A. Allensworth. Structured programming in FORTRAN. Technical report, Sandia Laboratories, Albuquerque, NM, USA, 1975. 55 pp.

Alexander:1972:FIP

[AM72] Daniel E. (Daniel Edward) Alexander and Andrew C. Messer. Fortran IV Pocket Handbook. Mc-Graw-Hill, New York, NY, USA, June 1972. ISBN 0-07-001015-3. 91 pp. LCCN QA76.73.F25 A43. US$4.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070010153.

Ageloff:1979:FFM

[AM79] Roy Ageloff and Richard Mojena.

Fundamentals of FORTRAN for management. Wadsworth, Pacific Grove, CA, USA, 1979. ISBN 0-534-00710-4. xvi + 443 pp. LCCN QA 76.73 F25 A34.

Ambrosio:1965:LA

[Amb65] Silvano Ambrosio. Linguaggi Algebrici. Number 4 in Serie di recerca operativa. P. Boringhieri, Torino, Italy, 1965. 170 pp. LCCN QA76.73.A24 A42.

ANSI:1966:AF

[Ame66a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9-1966, 1966. ?? pp.

URL http://www.fh-jena.de/~
kleine/history/languages/ansi-
x3dot9-1966-Fortran66.pdf. Ap-

proved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [ANS69c, ANS71a, ANS69b, ANS71c], and history [Hei64, Hei66].

ANSI:1966:AFX

[Ame66b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9-1966, 1966. Approved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [ANS69c, ANS71a].

ASA:1966:BFU

[Ame66c] American Standards Association. Basic FORTRAN USA standard FORTRAN. Approved March 7, 1966. USAS X 3.9-25ag71 1966; X 3.10-1966. American Standards Association, Inc., New York, NY, USA, 1966. 32 + 36 pp.

Institute:1977:USF

[Ame77] American National Standards Institute. USA Standard Fortran. Technical report, American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1977.

ANSI:1978:ANSa

[Ame78a] American National Standards Institute. American National Standard programming language FOR-TRAN. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. ca. 400 pp.

ANSI:1978:ANSc

[Ame78b] American National Standards Institute. American National Standard programming language, FOR-TRAN. American National Standard; ANSI X3.9-1978 CSA standard; Z243.18-1980 American National Standards Institute. American National Standard; ANSI X3.9-1978. Canadian Standard Association. CSA standard; Z243.18-1980. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1978. 438 pp. Available on the World-Wide Web

at http://observer.gsfc.nasa.
gov/iteams/doc/ansi_f77.ps,
http://observer.gsfc.nasa.gov/
iteams/doc/f77.doc, http://
observer.gsfc.nasa.gov/iteams/
doc/f77_cov.pdf, and http://
observer.gsfc.nasa.gov/iteams/
doc/f77_doc.pdf.

ANSI:1978:ANSb

[Ame78c] American National Standards Institute. American National Standard programming language FOR-TRAN: approved April 3, 1978, American National Standards Institute, Inc. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. ca. 250 pp.

ANSI:1978:ANSd

[Ame78d] American National Standards Institute. American National Standard programming language FORTRAN: approved April 3, 1978, American National Standards Institute, Inc.: ANSI X3.9-1978. Revision of ANSI X3.9-1966. World-Wide Web document., 1978. URL http://www.

                                        fortran.com/fortran/F77_std/

rjcnf0001.html. This is an online hypertext of the Fortran 77 Standard.

ANSI:1978:AFX

[Ame78e] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9–1978, 1978. URL http://

                                        www.fh-jena.de/~kleine/history/
                                        languages/ansi-x3dot9-1978-

Fortran77.pdf. Approved April 3, 1978 (also known as Fortran 77). See [ANS76a, ANS76c].

ANSI:1978:FUX

[Ame78f] American National Standards Institute.Committee on Computers X3 Information Processing. Fortran 77: unofficial, X3J3/ 90.5, (78-06-01). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. 371 pp.

Amkreutz:1973:ADH

[Amk73] Carl Amkreutz. Abkuerzungen der Datenverarbeitung, Hardware — Software. J. J. Amkreutz, Bergisch Gladbach, Germany, 1973.

Anderson:1964:BCP

[And64a] Decima M. Anderson. Basic computer programming; IBM 1620 Fortran. ACC administration series. Appleton-Century-Crofts, New York, NY, USA, 1964. x + 245 pp. LCCN QA76.8.I16 A5 1964.

Anderson:1964:CPF

[And64b] Decima M. Anderson. Computer programming; Fortran IV. Appleton-Century-Crofts, New York, NY, USA, 1964. ix + 435 pp. LCCN QA76.5 .A49.

Anderson:1966:CPF

[And66] Decima M. Anderson. Computer programming; Fortran IV. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1966. ISBN 0-13-164822-5. ix + 435 pp.

Anderson:1970:EET

[And70] Mary Perry Anderson. An experimental evaluation of two approaches to teaching Fortran. Thesis (m.a.), San Diego State College, San Diego, CA, USA, 1970. 96 pp.

Anderson:1973:FIP

[And73] Walter L. Anderson. Fortran IV programs for the determination of the transient tangential electric field and vertical magnetic field about a vertical magnetic dipole for an m-layer stratified earth by numerical integration and digital linear filtering. Technical report, U.S. Geological Survey, Denver, CO, USA, 1973. 82 pp.

Andresen:1979:PFI

[And79a] K. Andresen. Programmieren mit Fortran IV fuer Ingenieurstudenten. Technical report, Technische Universität Braunschweig (??), Braunschweig, Germany, 1979.

Andresen:1979:PMF

[And79b] K. Andresen. Programmieren mit Fortran IV fuer Ingenieurstudenten. Technical report, Technische Universität Braunschweig (??), Braunschweig, Germany, 1979.

Anonymous:1964:PMF

[Ano64] Anonymous. Provision for multiprocessing of Fortran IV student programs on the IBM 7094 computer. Thesis (m.s.), University of Georgia, Athens, GA, USA, 1964. 37 pp.

Anonymous:1965:A

[Ano65] Anonymous. ALTRAN. Wikipedia article, 1965.

Anonymous:1967:SAF

[Ano67] Anonymous. Specification for additions to the Fortran V library. Technical report, Systems Programming, UNIVAC Division, Sperry Rand Corporation, New York, NY, USA, June 1967. 22 pp.

Anonymous:1968:FEP

[Ano68a] Anonymous. Fortran: enseignement programmé. Technical report, IBM France, Paris, France, 1968. 7 fasc. pp.

Anonymous:1968:FPL

[Ano68b] Anonymous. FORTRAN programming language for MAC systems reference guide. CEIR Multi-Access Computer Service, Washington, DC, USA, 1968. various pp.

Anonymous:1968:SCS

[Ano68c] Anonymous. SAHYB-2: a continuous system simulation language compatible with FORTRAN-IV. IEEE Transactions on Computers, C-17(2):187–188, February 1968. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1687311.

Anonymous:1969:APR

[Ano69a] Anonymous. Algorithms policy, revised September, 1969 (includes ALGOL, FORTRAN, and PL/I). Comm. ACM, 12(9):513, September 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1969:CFS

[Ano69b] Anonymous. Clarification of FORTRAN standards — initial progress. Comm. ACM, 12 (5):289–294, May 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1970:EFL

[Ano70a] Anonymous. Extended FORTRAN library: technical manual for XDS Sigma 5/6/7 computers. Xerox Data Systems, El Segundo, CA, USA, 1970. iv + 201 pp.

Anonymous:1970:FDP

[Ano70b] Anonymous. FORTRAN debug package (FDP) reference manual for XDS Sigma 5/7 computers. Xerox Data Systems, El Segundo, CA, USA, 1970. iv + 65 pp.

Anonymous:1970:FVP

[Ano70c] Anonymous. A Fortran V package for testing and anaylsis of pseudorandom number generators. Technical report CP-700011, Computer Science/Operations Research Center, Institute of Technology, Southern Methodist University, Dallas, TX, USA, 1970. 32 pp.

Anonymous:1970:SPF

[Ano70d] Anonymous. Sistema de programaçao Fortran 2: e sua utilizaçao com o computador electronico instalado no C.C.C. Instituto Gulbenkian de Ciencia, Centro de Calculo Cientifico, Lisboa, Portugal, 1970. 150 pp.

Anonymous:1970:TDP

[Ano70e] Anonymous. Timeshare devices plotter programs in DTSS FOR-TRAN. Dartmouth College, Hanover, NH, USA, 1970. 53 pp.

Anonymous:1970:XEFa

[Ano70f] Anonymous. Xerox extended FOR-TRAN IV, Sigma 5/6/7 computers: language reference manual. Xerox Data Systems, El Segundo, CA, USA, 1970. v + 151 pp.

Anonymous:1970:XEFb

[Ano70g] Anonymous. Xerox extended FOR-TRAN IV; Sigma 5/6/7 operations reference manual. Xerox Data Systems, El Segundo, CA, USA, revised edition, 1970. iv + 108 pp.

Anonymous:1972:FHC

[Ano72a] Anonymous. Fortran Hydro CDC 6500. Water Resources and Hydromechanics Laboratory. Technical Report 36, School of Civil Engineering, Purdue University, West Lafayette, IN, USA, 1972. various pp.

Anonymous:1972:FIP

[Ano72b] Anonymous. Fortran IV programming language; text for instruction. Technical report, Chinese University of Hong Kong, Hong Kong, 1972. 119 pp.

Anonymous:1972:ICS

[Ano72c] Anonymous. Industrial computer system FORTRAN procedures for executive functions and process input-output; standard. Technical Report ISA-S61.1, Instrument Society of America, Pittsburgh, PA, USA, 1972. 10 pp.

Anonymous:1972:SJM

[Ano72d] Anonymous. Shih i ju men FOR-TRAN chih cheng shih p’ien hsieh. Hsu shih chi chin, Taipei, Taiwan, 1972. 162 pp.

Anonymous:1972:SPF

[Ano72e] Anonymous. Solution to problems in Fortran IV programming language. Technical report, Chinese University of Hong Kong, Hong Kong, 1972. 75 pp.

Anonymous:1973:XEF

[Ano73] Anonymous. Xerox extended FOR-TRAN IV, Sigma 5-9 computers: language reference manual. –. Xerox Corp., El Segundo, CA, USA, second edition, 1973. 174 pp.

Anonymous:1974:NFTc

[Ano74a] Anonymous. NBS Fortran test programs: Documentation for versions 1 and 3. Government Printing, Washington, DC, USA, 1974. various pp.

Anonymous:1974:NFTa

[Ano74b] Anonymous. NBS Fortran test programs: Listings for version 1. Government Printing, Washington, DC, USA, 1974. VII + 212 pp.

Anonymous:1974:NFTb

[Ano74c] Anonymous. NBS Fortran test programs: listings for version 3. Government Printing, Washington, DC, USA, 1974. IX + 215 pp.

Anonymous:1975:DEF

[Ano75a] Anonymous. Differences entre le FORTRAN 6600 RUN et le FOR-TRAN 7600 FTN. Compagnie internationale de services en informatique, Saclay, France, 1975. 18 pp.

Anonymous:1975:TNY

[Ano75b] Anonymous. Tien nao yu yen fu ch’uan FORTRAN IV chen shih se chi. Hsu shih Foundation, Taipei, Taiwan, 1975. 177 pp.

Anonymous:1975:XEFa

[Ano75c] Anonymous. Xerox extended FOR-TRAN IV, Xerox 550/560 and Sigma 5-9 computers language reference manual. Xerox Corporation, El Segundo, CA, USA, 1975. ???? pp.

Anonymous:1975:XEFb

[Ano75d] Anonymous. Xerox extended FOR-TRAN IV, Xerox 550/560 and Sigma 5-9 computers operations reference manual. Xerox Corporation, El Segundo, CA, USA, 1975. ???? pp.

Anonymous:1975:XFD

[Ano75e] Anonymous. Xerox FORTRAN debug package (FDP), Sigma 5-9 computers: reference manual. Xerox Corporation, El Segundo, CA, USA, 1975. ???? pp.

Anonymous:1976:IAL

[Ano76a] Anonymous. Initiation au langage Fortran: DEUG A 2 informatique. Université Claude Bernard, Lyon, France, 1976. 35 pp.

Anonymous:1976:NAF

[Ano76b] Anonymous. New ANS Fortran Standard. The Computer Journal, 19(1):7, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/
content/19/1/7.full.pdf+html.

Anonymous:1977:FDN

[Ano77a] Anonymous. Fortran Development Newsletter: Volume 3, Number 1, January 1977, pp. 1–10:

X3J3 expects to complete Fortran 77 revised at March meeting. ACM SIGPLAN Notices, 12 (4):21–30, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1977:LF

[Ano77b] Anonymous. Language Fortran. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, June 1977. ISBN 999779-773-6. ???? pp. LCCN ???? US$22.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
9997797736.

Anonymous:1977:LFA

[Ano77c] Anonymous. Language Fortran (ANSI X3-9 Order No. ISO 1539). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, June 1977. ISBN 999779-773-6. ???? pp. LCCN ???? US$22.00.

Anonymous:1978:AFD

[Ano78a] Anonymous. ASCII Fortran documentation for level 7R1. ????, ????, 1978. 107 pp.

Anonymous:1978:CDM

[Ano78b] Anonymous. Cours d’informatique: memento Fortran: promotion 1977. Ecole Polytechnique, Paris, France, 1978. various pp.

Anonymous:1978:IF

[Ano78c] Anonymous. Introduction to FOR-TRAN, 1978.

Anonymous:1978:PLF

[Ano78d] Anonymous. Programming Language Fortran/ANSI X3.9-1978. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, June 1978. ISBN 999848-763-3. ???? pp. LCCN ???? US$75.00; US$40.00. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=9998487633.

Anonymous:1979:IPF

[Ano79] Anonymous. Initiation a la programmation Fortran: informatique (1979-1980). Université Claude Bernard, Lyon, France, 1979. 83 pp.

Anonymous:1980:AFL

[Ano80a] Anonymous. Apple FORTRAN:

language reference manual. Apple Computer, Cupertino, CA, USA, 1980. xii + 230 pp.

Anonymous:1980:IF

[Ano80b] Anonymous. Introduction to Fortran. United States Government Printing Office, Washington, DC, USA, June 1980. ISBN 0-8470-0306-4, 99973-850-5-5. ???? pp. LCCN ???? US$8.12.

URL http://www.amazon.com/
exec/obidos/ISBN=9997385055/
wholesaleproductA/;   http:
//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
9997385055.

Anonymous:1980:NFL

[Ano80c] Anonymous. Nag Fortran Library Manual Mark 8. Numerical Algorithms Group, February 1980.

ISBN 0-317-52215-9. ???? pp. LCCN ???? US$35.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0317522159.

Anonymous:1980:SFP

[Ano80d] Anonymous. Standard Fortran programming manual. The National Computing Center, Manchester, UK, 1980. 152 pp.

ANSI:ftn69

[ANS69a] ANSI Subcommittee X3J3. Clarification of Fortran standards — initial progress. Comm. ACM, 12:289–294, 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [?].

X3J3:1969:CFS

[ANS69b] ANSI Subcommittee X3J3. Clarification of Fortran standards — initial progress. Comm. ACM, 12(5):289–294, May 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66a].

ANSI:1969:CFS

[ANS69c] ANSI Subcommittee X3J3. Clarification of Fortran standards — initial progress. Comm. ACM, 12(5):289–294, May 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66a, Ame66b].

ANSI:1971:CFS

[ANS71a] ANSI Subcommittee X3J3. Clarification of Fortran standards — second report. Comm. ACM, 14 (10):628–642, October 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66a, Ame66b].

ANSI:ftn71

[ANS71b] ANSI Subcommittee X3J3. Clarification of Fortran standards — second report. Comm. ACM, 14:628– 642, 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [?].

X3J3:1971:CFS

[ANS71c] ANSI Subcommittee X3J3. Clarification of Fortran standards — second report. Comm. ACM, 14 (10):628–642, October 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66a].

ANSI:1976:DPA

[ANS76a] ANSI Subcommittee X3J3. Draft proposed ANS Fortran. ACM SIGPLAN Notices, 11(3):various, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also final standard [Ame78e].

ANSI:ftn76

[ANS76b] ANSI Subcommittee X3J3. Draft proposed ANS Fortran. ACM SIG-PLAN Notices, 11(3), 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also final standard [?].

X3J3:1976:DPA

[ANS76c] ANSI Subcommittee X3J3. Draft proposed ANS Fortran. ACM SIG-PLAN Notices, 11(3), 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also final standard [Ame78e].

ANSIA:1978:ANS

[ANS78] ANSI. American National Standard Programming Language FORTRAN. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978.

Antes:1972:OVR

[Ant72] H. Antes. Ober die vierdimensionale Romberg-Integration mit Schranken. (German) [On the four-dimensional Rombergintegration with bounds]. Computing, 9(1):45–52, 1972. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Antonelli:1977:FIP

[Ant77] Rocco A. Antonelli. A FORTRAN IV program for optimal binary tree construction. Thesis (m.s.), Florida Institute of Technology, Melbourne, FL, USA, 1977. 181 pp.

Antonak:1980:FPA

[Ant80] Richard F. Antonak. A FOR-TRAN program to analyze summated rating scales. The American Statistician, 34(1):59, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.
org/stable/2683004.

Adams:1972:FRM

[AR72] J. C. Adams and P. A. Rotar. Fortran reference manual. Technical report, National Center for Atmospheric Research, Boulder, CO, USA, 1972. 14 + 12 pp.

Archibald:1976:TDF

[Arc76] J. A. Archibald, Jr. Towards a dynamic (FORTRAN) programming system. ACM SIGPLAN Notices, 11(7):17–24, July 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ASDC:1976:SSC

[Ari76] Arizona.State Dept.of Corrections.Division of Research and Program Planning. Evaluation and Statistical Classification Programs. SCP: statistical classification programs: a FORTRAN IV program, version 1, July 1976, user’s guide. Technical report, The Division, ????, 1976. 86 pp.

Armstrong:1978:PPC

[Arm78] J. L. Armstrong. Programming a parallel computer for robot vision. The Computer Journal, 21 (3):215–218, August 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Arnold:1965:SAU

[Arn65] P. H. Arnold. Still another use for FORTRAN II chaining. Comm. ACM, 8(4):234, April 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arsac:1964:MDP

[Ars64] Jacques Arsac. Manuel du programmeur — Fortran IV. Observatoire, Paris, France, 1964. 86 pp.

Arthur:1975:CRV

[Art75] D. F. Arthur. Correspondence: Reading variable length records using FORTRAN IV. The Computer Journal, 18(3):283, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/283.

tif. See [HV74, ?].

Acara:1971:TEK

[AS71] A. H. Acara and H. D. Smith. A technique for enumerating Kokanee salmon (Oncorhynchus nerka) fry migrating through streams, with an appendix for processing catch data by IBM 360 Fortran IV computer programs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):573–585, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/
10.1139/f71-081.

Ashby:1973:DID

[ASH73] Gordon Ashby, Loren Salmonson, and Robert Heilman. Design of an interactive debugger for FOR-TRAN:MANTIS. Software—Practice and Experience, 3(1):65–74, January/March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Adamski:1977:IOS

[ATW77] Andrzej Adamski, Andrzej Turnau, and Jan Werewka. ??87 [i.e. Osiemdziesiat siedem] przykladow w FORTRANIE: elektroniczna technika obliczeniowa w zadaniach rozwiazanych w jezyku FORTRAN wraz ze schematami blokowymi algorytmow. Skrypty uczelniane — Akademia Gorniczo-Hutnicza im. S. Staszica; Nr.

559. Akademia Gorniczo-Hutnicza, Krakow, Poland, wyd. 1 edition, 1977. 351 pp.

Aho:1977:PCD

[AU77] Alfred V. Aho and Jeffrey D. Ullman. Principles of Compiler Design. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-00022-9 (hardcover), 0-201-10073-8. x + 604 pp. LCCN QA76.6 .A285 1977. See also the much expanded subsequent book [?].

Aubry:1976:PFP

[Aub76] Anne Marie Aubry. Programmes FORTRAN pour distances de rangs, constellations et correlation. Initiations, documentations techniques; no 30. O.R.S.T.O.M., Paris, France, 1976. ISBN 2-7099-0420-9. 56 pp. LCCN QA278.2.A9.

Allen:1973:PCC

[AW73a] R. C. Allen, Jr. and G. Milton Wing. Problems for a Computer-Oriented Calculus Course. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1973. ISBN 0-13-716423-8. xiii + 206 pp. LCCN QA303.A43.

Allen:1973:PCO

[AW73b] Richard C. Allen, Jr. and G. Milton Wing. Problems for a computer-oriented calculus course, with an appendix on elementary FORTRAN programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1973. ISBN 0-13-716423-8. xiii + 206 pp. LCCN QA303.A43.

Axford:1972:PLC

[Axf72] H. William Axford, editor. Proceedings of the LARC Computer-Based Unit Cost Studies Institute (University of Texas, Austin, September 16–17, 1971). The LARC Association, P. O. Box 27235, Tempe, Arizona 85282, USA, 1972.

Aycock:1980:SAS

[Ayc80] Doris S. Aycock. A simplified approach to structured Fortran programming. Thesis (sp. a.), Eastern Michigan University, Ypsilanti, MI, USA, 1980. vii + 458 pp.

Ayers:1963:RPF

[Aye63] James A. Ayers. Recursive programming in Fortran II. Comm. ACM, 6(11):667–668, November 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berude:1972:RFP

[BA72] Catherine L. Berude and Norman J. Abramson. Relative fishing power, CDC 6600, FORTRAN IV. Transactions of the American Fisheries Society, 101(1):133, January 1972. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beck:1973:CSL

[BA73] Norman Beck and Gordon Ashby.

On cost of static linking and loading of subprograms. ACM SIG-METRICS Performance Evaluation Review, 2(3):17–20, September 1973. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Backus:1954:ISS

[Bac54] J. W. Backus. The IBM 701 Speedcoding system. Journal of the ACM, 1(1):4–6, January 1954. CODEN JACOAH. ISSN 0004-5411. URL http:

//community.computerhistory.
org/scc/projects/FORTRAN/paper/
p4-backus.pdf.

Backus:1956:FAC

[Bac56] J. W. Backus. The Fortran Automatic Coding System for the IBM 704 EDPM. ????, ????, 1956. ???? pp. LCCN ????

Backstrom:1972:AMP

[Bac72] R. P. Backstrom. AAEARITH, multiple precision arithmetic routines for the IBM360 Fortran user. AAEC/TM633. Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1972. ISBN 0-642-99512-5. 6 + A2 pp.

Backus:1978:CPL

[Bac78a] John Backus. Can programming be liberated from the von Neumann style? A functional style and its algebra of programs. Comm. ACM, 21(8): 613–641, August 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.stanford.edu/
class/cs242/readings/backus.

pdf. Reproduced in “Selected Reprints on Dataflow and Reduction Architectures” ed. S. S. Thakkar, IEEE, 1987, pp. 215-243.

Backus:1978:HFI

[Bac78b] John Backus. The history of FORTRAN I, II, and III. ACM SIGPLAN Notices, 13(8): 165–180, August 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//community.computerhistory.
org/scc/projects/FORTRAN/paper/
p165-backus.pdf.

Backus:1979:HFI

[Bac79] John Backus. The history of FOR-TRAN I, II, and III. Annals of the History of Computing, 1(1):21– 37, July/September 1979. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/
books/an1979/pdf/a1021.pdf;
 http://www.computer.org/annals/
an1979/a1021abs.htm.

Backus:1980:F

[Bac80] John W. Backus. Fortran, 1980.

Badach:1977:PWJ

[Bad77] Anatol Badach. Programowanie w jezyku FORTRAN 1900. Skrypty uczelniane — Uniwersytet Gdanski ??, UG, Gdansk, Poland, 1977. 224 + [1] pp.

Bahr:1969:SFF

[Bah69] K. Bahr. A smaller FORTRAN– FORMAC system. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(12):59–60, July 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bailey:1962:SAW

[Bai62] Michael J. Bailey. Syntactic analysis within Fortran — the shadow IVF system. Technical report 15, M.I.T. Cooperative Computing Laboratory, Cambridge, MA, USA, 1962. 75 pp.

Bailey:1963:FFI

[Bai63] Michael J. Bailey. A free format input routine to read Fortran variables and specified portions of arrays. Technical Report 30, M.I.T. Cooperative Computing Laboratory, Cambridge, MA, USA, 1963. 41 pp.

Bailey:1972:MEF

[Bai72a] James Lawrence Bailey. Msufor, an educational Fortran processor. Thesis, Dept. of Computer Science, Mississippi State University, Mississippi State, MS, USA, 1972. v + 29 pp.

Bailey:1972:O

[Bai72b] T. E. Bailey. Ouchless I/O. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(3):8–18, October 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bajpai:1972:FAP

[Baj72] Avinash Chandra Bajpai. Fortran and Algol: a programmed course for students of science and technology. A Series of programmes on mathematics for scientists and technologists. John Wiley and Sons, New York, London, Sydney, 1972. ISBN 0-471-04371-0 (paperback). 21 + 133 + 87 pp.

Bajpai:1977:FA

[Baj77] Avi C. Bajpai. Fortran y Algol. Editorial Limusa, Mexico, DF, Mexico, 1977. 265 pp.

Bakstad:1968:RIF

[Bak68] Pal Bakstad. “Ramona I” a FOR-TRAN code for transient analysis of boiling water reactors and boiling loops. Kjeller report 135, Institutt for atomenergi, Kjeller, Norway, 1968. various pp.

Baker:1975:PSP

[Bak75] Brenda S. Baker. Struct — a program which structures Fortran. Technical Memorandum TM-75-1271-12, Bell Laboratories, Murray Hill, NJ, USA, December 22, 1975. See also [Bak77a].

Baker:1977:ASF

[Bak77a] Brenda S. Baker. An algorithm for structuring flowgraphs. Journal of the ACM, 24(1):98– 120, January 1977. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baker:1977:SAA

[Bak77b] R. J. Baker. Statistical algorithms:

Algorithm AS 112: Exact distributions derived from two-way tables. Applied Statistics, 26(2):199–206, June 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/112. See

remarks [?, ?].

Ball:1969:SFS

[Bal69] John A. Ball. Some FOR-TRAN subprograms used in astronomy. Technical note 1969-42, M.I.T. Lincoln Laboratory, Lexington, MA, 1969. iii + 19 pp.

Ball:1973:BS

[Bal73] William E. Ball. The BSOLVE subroutine. In James L. Kuester and Joe H. Mize, editors, Optimization Techniques with FOR-TRAN, pages 240–250. McGraw-Hill, New York, NY, USA, 1973.

Banham:1975:NMF

[Ban75] James W. Banham. Numerical methods and FORTRAN IV applications in engineering problems. Drexel University Press, Philadelphia, PA, USA, 1975. 168 pp.

Bank:1978:FIG

[Ban78a] R. E. Bank. A Fortran implementation of the generalized marching algorithm. ACM Transactions on Mathematical Software, 4(2):165– 176, June 1978.

Bank:1978:AAF

[Ban78b] Randolph E. Bank. Algorithm 527: A FORTRAN implementation of the generalized marching algorithm. ACM Transactions on Mathematical Software, 4(2):165– 176, June 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bank:1978:AFI

[Ban78c] Randolph E. Bank. Algorithm 527: A Fortran implementation of the generalized marching algorithm [D3]. ACM Transactions on Mathematical Software, 4(2):165– 176, June 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barnett:1961:LLL

[Bar61] M. P. Barnett. Low-level language subroutines for use within Fortran. Comm. ACM, 4(11): 492–495, November 1961. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bartcher:1966:FIP

[Bar66] Ronald L. Bartcher. Fortran IV program for estimation of cladistic relationships using the IBM 7040. Computer contribution 6, University of Kansas, Lawrence, KS, USA, 1966. 54 pp.

Barlow:1970:FIP

[Bar70] H. A. Barlow. FORTRAN: introduction to programming for business applications. Barlow, Beaumont, TX, USA, 1970. iv + 153 pp.

Baron:1971:BRG

[Bar71a] W. Baron. Book review: G.

Lamprecht, Einführung in die Programmiersprache FORTRAN IV. Computing, 8(1–2):202, 1971. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barron:1971:ACF

[Bar71b] D. W. Barron. Approaches to conversational FORTRAN. The Computer Journal, 14(2): 123–127, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_02/140123.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/123.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
02/tiff/124.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/
125.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_02/tiff/126.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_02/tiff/127.tif.

Barron:1972:BRBa

[Bar72a] D. W. Barron. Book review: Standard Fortran programming manual, The National Computing Centre Limited, Manchester. No. of pages: 152. Price: £3. Software —Practice and Experience, 2(1): 98–99, January 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barry:1972:AFS

[Bar72b] J. M. Barry. AESYNTAX: a Fortran syntax analysis system for the PDP9L. AAEC E251, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1972. ISBN 0-642-99522-2. 6 + [3] pp.

Barron:1973:BRBe

[Bar73a] D. W. Barron. Book review: FOR-TRAN techniques with special reference to non-numerical applications, A. C. Day, Cambridge University Press, 1972. No. of pages:

96. Price: £2.50. Software—Practice and Experience, 3(4):404, October/December 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barry:1973:SST

[Bar73b] J. M. Barry. SPLINS and SMOOTH, two FORTRAN smoothing routines. AAEC E253, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1973. ISBN 0-642-99532-X. 11 + [3] pp.

Bartlett:1973:FTS

[Bar73c] Michael L. Bartlett. A FORTRAN translating system for the IBM 1620. Thesis (m.s.), Southern Illinois University, Dept. of Engineering and Technology, Carbondale, IL, USA, 1973. v + 105 pp.

Barth:1974:BRK

[Bar74] W. Barth. Book review: K.

H. Müller und I. Strecher, FOR-TRAN, Programmieranleitung, 2 Aufl. Computing, 12(4):386, 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barron:1975:CFP

[Bar75] J. L. Barron. CHNGLC: a Fortran program to process output from a gas chromatograph system. Technical report, Bedford Institute of Oceanography, Dartmouth, NS, Canada, 1975. ii + 17 pp.

Barber:1977:FOA

[Bar77a] Willie D. Barber. FORTRAN optimizations at the source code level. Thesis (m.s.), North Texas State University, Denton, TX, USA, August 1977. iv + 70 pp.

Barnard:1977:FFF

[Bar77b] A. J. Barnard. Furi — a FOR-TRAN function writer. Computer Physics Communications, 13 (4):271–280, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465577900054.

Barrodale:1977:FPS

[Bar77c] Ian Barrodale. A Fortran program for solving a nonlinear equation by Muller’s method. Internal report 108, University of Victoria, Dept. of Mathematics, Victoria, BC, Canada, 1977. 9. pp.

Barrodale:1979:FPL

[Bar79a] Ian Barrodale. A Fortran program for least-squares linear prediction and maximum entropy spectral analysis. Internal report 182, University of Victoria, Dept. of Mathematics, Victoria, BC, Canada, 1979. 18. pp.

Barrodale:1979:FPS

[Bar79b] Ian Barrodale. A FORTRAN program for solving Ax = b. Internal report 120, University of Victoria, Dept. of Mathematics, Victoria, BC, Canada, 1979. 17. pp.

Barakat:1980:FIP

[Bar80a] S. A. Barakat. A FORTRAN IV program to calculate net heat gains through windows. DBR computer program 47, National Research Council of Canada, Division of Building Research, Ottawa, Ontario, Canada, 1980. 12 + [6] + A-18 pp.

Barron:1980:BRBa

[Bar80b] D. W. Barron. Book review: Compatible FORTRAN, A. Colin Day. Cambridge University Press, 1979. No. of pages: 107. Price: £5.95. Software—Practice and Experience, 10(10):850, October 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barton:1980:TSS

[Bar80c] David Barton. On Taylor series and stiff equations. ACM Transactions on Mathematical Software, 6 (3):280–294, 1980. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bashioum:1980:BIS

[Bas80] Douglas L. Bashioum. Benchmarking interactive systems: Calibrating the model. ACM SIGMET-RICS Performance Evaluation Review, 9(2):35–41, Summer 1980. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bauer:1956:UAP

[Bau56] Walter F. Bauer. Use of automatic programming. Computers and Automation, 5(11):6–11, November 1956. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549. The online edition of the Oxford English Dictionary cites this as the earliest mention of the name FORTRAN, with the quote: “John Backus’ group at IBM has prepared FOR-TRAN (FORmula TRANslation) for the IBM-704 computer. FOR-TRAN will translate into computer language a program written very close [sic] the language of the mathematician or scientist.” The OED citation references issue “Nov. 9/2”; that refers to page 9, column 2, of the article. There are no further references to Fortran in this journal up to at least 1962.

Bauer:1979:EPI

[Bau79] Michael A. Bauer. Experiences with PASCAL in an introductory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1): 158–161, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Berry:1971:NAH

[BB71] J. L. Berry and A. G. Bell. Note on Algorithm 50: How to program a computer to play legal chess. The Computer Journal, 14(1):106–107, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

                                        oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_01/tiff/
                                        106.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_14/Issue_01/tiff/107.

tif. See [?].

Balfour:1972:BNA

[BB72] Alexander Balfour and Walter Thomson Beveridge. Basic numerical analysis with Fortran. Heinemann Educational Books, London, UK, 1972. ISBN 0-435-77481-6. 239 pp.

Balfour:1977:BNA

[BB77a] Alexander Balfour and Walter Thomson Beveridge. Basic numerical analysis with Fortran. Heinemann Educational Books, London, UK, second edition, 1977. ISBN 0-435-77484-0. viii + 287 pp.

Banfield:1977:SAA

[BB77b] C. F. Banfield and L. C. Bassill. Statistical algorithms: Algorithm AS 113: a transfer for nonhierarchical classification. Applied Statistics, 26(2):206–210, June 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/113.

Balfour:1978:ANB

[BB78] Alexander Balfour and Walter Thompson Beveridge. Analisis numerico basico con Fortran. Compania Editorial Continental, Mexico, DF, Mexico, 1978. ISBN 968-26-0027-8. 270 pp.

Backus:1957:FAC

[BBB^{+}57] J. W. Backus, R. J. Beeber, S. Best, R. Goldberg, L. M. Haibt, H. L. Herrick, R. A. Nelson, D. Sayre, P. B. Sheridan, H. Stern, I. Ziller, R. A. Hughes, and R. Nutt. The FORTRAN automatic coding system. In Proceedings of the Western Joint Computer Conference, February 26–28, 1957, Los Angeles, CA, USA, pages 188–198. Institute of Radio Engineers, 1 East 79th Street, New York 21, NY, USA, 1957. ISSN 0449-1173. LCCN TK7885.A1 J6. URL http:

//community.computerhistory.
org/scc/projects/FORTRAN/paper/
BackusEtAl-FortranAutomaticCodingSystem-

1957.pdf. The online edition of the Oxford English Dictionary cites this as the second earliest mention of the name FORTRAN, with the extract “The programmer attended a one-day course on FORTRAN and … then programmed the job in four hours using 47 FORTRAN statements.”.

Bailey:1964:SMF

[BBB64] M. J. Bailey, M. P. Barnett, and P. B. Burleson. Symbol manipulation in FORTRAN: SASP I subroutines. Comm. ACM, 7 (6):339–346, June 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bonham-Carter:1967:FIP

[BC67] Graeme Bonham-Carter. Fortran IV program for Q-mode cluster analysis of nonquantitative data using IBM 7090/7094 computers. Computer contribution 17, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 28 pp.

Brillinger:1970:CPI

[BC70] P. C. Brillinger and D. D. Cowan. A complete package for introducing computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 2(3):118–126, November 1970. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beyer:1972:FIC

[BC72a] L. A. Beyer and C. E. Corbato. A FORTRAN IV computer program for calculating borehole gravity terrain corrections. Computer contribution 18, U.S. Geological Survey, Geological Division, Menlo Park, CA, USA, 1972. ???? pp.

Brillinger:1972:IDS

[BC72b] Peter C. Brillinger and Doron J.

Cohen. Introduction to Data Structures and Non-Numeric Computation. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1972. ISBN 0-13-479899-6. 629 pp. LCCN QA76.6 .B74.

Beven:1977:HSF

[BC77] K. J. Beven and J. L. Callen. Hydrodat, a system of FORTRAN computer programs for the preparation and analysis of hydrological data from charts. Working paper 175, University of Leeds, School of Geography, Leeds, UK, 1977. 32 + A1–A4 pp.

Beven:1979:HSF

[BC79] K. J. Beven and J. L. Callen. HY-DRODAT: a system of FORTRAN computer programs for the preparation and analysis of hydrological data from charts. Technical bulletin 23, British Geomorphological Research Group, Norwich, UK, 1979. ISBN 0-86094-023-3. 72 pp.

Blair:1977:IFJ

[BCE77] J. M. Blair, M. B. Carver, and L. E. Evans. Introduction to Fortran and job control cards on the 170/6600 system. Technical report, Chalk River Nuclear Laboratories, Chalk River, Ontario, Canada, 1977. 53 pp.

Banerjee:1979:TPP

[BCKT79] Utrpal Banerjee, Shyh-Ching Chen, David J. Kuck, and Ross A.

Towle. Time and parallel processor bounds for FORTRAN-like loops. IEEE Transactions on Computers, C-28(9):660–670, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bonham-Carter:1968:MMF

[BCS68] Graeme Bonham-Carter and Alexander James Sutherland. Mathematical model and FORTRAN IV program for computer simulation of deltaic sedimentation. Computer contribution 24, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1968. 56 pp.

Beard:1971:EFI

[BD71] Charles L. Beard and Robert A. Dannels. ETOT a Fortran IV program to process data from the ENDF/B file to thermal library format. Technical report, Westinghouse Electric Corporation, Nuclear Energy Systems, Pittsburgh, PA, USA, 1971. 69 pp.

Bailey:1980:UTV

[BD80a] M. Gene Bailey and Lloyd Davis. Using terminals versus card reader in remote job entry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 12(1):181–183, February 1980. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 11th SIGCSE Symposium on Computer Science Education.

Burkard:1980:AMP

[BD80b] Rainer E. Burkard and Ulrich Derigs. Assignment and matching problems: solution methods with FORTRAN-programs. Lecture notes in economics and mathematical systems ; 184. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-10267-1. 148 pp. LCCN QA402.5.B86.

Brun:1972:EFI

[BDI72] Gerard Brun, Pierre Dubois, and Michel Isambert. Entrées-sorties: FORTRAN IV, PL/1, COBOL (I.B.M. 360). Compagnie internationale de services en informatique, Saclay, France, 1972. 267 pp.

Busam:1969:OEF

[BE69] V. A. Busam and D. E. Englund. Optimization of expressions in Fortran. Comm. ACM, 12 (12):666–674, December 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bauer:1974:CCA

[BE74] F. L. Bauer and J. Eickel, editors. Compiler Construction — An Advanced Course. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. xiv + 621 pp. LCCN QA76.6 .C6281.

Beattie:1975:FPC

[Bea75] R. D. Beattie. A FORTRAN program for calculating rock densities from gravity survey data. Technical communication 56, CSIRO Division of Mineral Physics, North Ryde, NSW, Australia, 1975. ISBN 0-643-01080-7. 9 + [15] pp.

Beckman:1972:HHT

[Bec72] William A. Beckman. Heat: heat transfer computer program in Fortran V language. EES report 37, Engineering Experiment Station, Madison, WI, USA, 1972. 65 pp.

Beck:1973:ICC

[Bec73] Tommy Ray Beck. An investigation concerning certain characteristics of Fortran programs. Thesis (m.s.), Dept. of Computer Science. University of Houston, Houston, TX, USA, 1973. 71. pp.

Beebe:1970:APH

[Bee70a] Nelson H. F. Beebe. Automatic page headers in FORTRAN output. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 1 pp.

Beebe:1970:BCF

[Bee70b] Nelson H. F. Beebe. Batch compilation of FORTRAN. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 2 pp.

Beebe:1970:FNC

[Bee70c] Nelson H. F. Beebe. FOR-TRAN name changing utility program. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, November 30, 1970. 1 pp.

Beebe:1970:RRR

[Bee70d] Nelson H. F. Beebe. RDUMP and RSDUMP: Register dumps for FORTRAN programs. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 1 pp.

Beebe:1971:FHC

[Bee71a] Nelson H. F. Beebe. FORTRAN H compiler. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 2, 1971. 1 pp.

Beebe:1971:FHE

[Bee71b] Nelson H. F. Beebe. FORTRAN H (EXTENDED). Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, June 15, 1971. 5 pp.

Beebe:1971:LFC

[Bee71c] Nelson H. F. Beebe. List of FORTRAN catalogued procedures. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, February 1, 1971. 1 pp.

Beebe:1971:MIF

[Bee71d] Nelson H. F. Beebe. More information on FORTRAN H EX-TENDED. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 11, 1971. 1 pp.

Beech:1975:FIC

[Bee75] Graham Beech. Fortran IV in Chemistry: An Introduction to Computer-Assisted Methods. John Wiley and Sons, New York, London, Sydney, December 1975. ISBN 0-471-06165-4. x + 303 pp. LCCN QD39.3 E46B43 1975. US$62.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471061654.

Beebe:1976:NEP

[Bee76] Nelson H. F. Beebe. Notes on EDIT — a program for editing FORTRAN programs. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1977:EFP

[Bee77a] Nelson H. F. Beebe. Editing FORTRAN Programs. Technical Report QTP.770927.02, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 27 1977. 6 pp.

Beebe:1977:MIF

[Bee77b] Nelson H. F. Beebe. Machineindependent FORTRAN Programming. Technical Report QTP.771205.03, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, December 5 1977. 10 pp.

Beebe:1978:UEF

[Bee78] Nelson H. F. Beebe. A Utility for Editing FORTRAN Programs [New Release]. Technical Report QTP.780825.09, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 25 1978. 18 pp.

Beebe:1979:PSSa

[Bee79a] Nelson H. F. Beebe. Proposal for a Standard Set of Primitives for Machine-Independent Bit Manipulation in FORTRAN. In Proceedings of the NRCC Conference on Software Standards in Chemistry, pages 106–115. National Resource for Computation in Chemistry, Lawrence Berkeley Laboratory, Berkeley, CA, USA, July 25– 27 1979.

Beebe:1979:PSSc

[Bee79b] Nelson H. F. Beebe. Proposal for a standard set of primitives for machine-independent bit manipulation in FORTRAN. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, December 13, 1979. 10 pp.

Beebe:1979:PSSb

[Bee79c] Nelson H. F. Beebe. Proposal for a Standard Set of Primitives for Machine-Independent Character Manipulation in FORTRAN. In Proceedings of the NRCC Conference on Software Standards in Chemistry, pages 116–128. National Resource for Computation in Chemistry, Lawrence Berkeley Laboratory, Berkeley, CA, USA, July 25–27 1979.

Beebe:1979:PSSd

[Bee79d] Nelson H. F. Beebe. Proposal for a standard set of primitives for machine-independent character manipulation in FORTRAN. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, December 13, 1979. 13 pp.

Beebe:1979:SIG

[Bee79e] Nelson H. F. Beebe. SFTRAN 3 Installation Guide. College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, July 1979.

Beebe:1980:PPF

[Bee80a] Nelson H. F. Beebe. PRETTY— A Portable FORTRAN Program Prettyprinter. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, May 7 1980. 33 pp.

Beebe:1980:PSS

[Bee80b] Nelson H. F. Beebe. Proposal for a standard set of primitives for machine-independent bit manipulation in Fortran. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, 1980. This document is included in the <PLOT79> distribution for online access.

Beebe:1980:SIG

[Bee80c] Nelson H. F. Beebe. SFTRAN3 Installation Guide. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, March 4 1980. 69 pp.

Beebe:1980:SLU

[Bee80d] Nelson H. F. Beebe. SFTRAN3 Listings of the Utah Polarization Propagator Program. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, May 13 1980. 158 pp.

Bemer:1961:LER

[Bem61] R. W. Bemer. Letter to the Editor:

Re: René De La Briandais’ letter on FORTRAN. Comm. ACM, 4 (3):A12–A13, March 1961. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [?].

Benson:1969:FIM

[Ben69] Jimmie Dan Benson. On the feasibility of an instructional model of the Burroughs B5500 computer programmed in the Fortran IV language for the IBM 360/65 computer. Thesis (m.s. in computer science), Texas A and M University., College Station, TX, USA, 1969. 138 pp.

Benediktsson:1977:SFP

[Ben77] O. Benediktsson. Sequential file processing in Fortran. Software— Practice and Experience, 7(5):655– 659, September/October 1977.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Benediktsson:1978:NAP

[Ben78] Oddur Benediktsson. Notes on argument-parameter association in Fortran. ACM SIGPLAN Notices, 13(1):16–20, January 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bendzulla:1980:I

[Ben80] C. Bendzulla. Intervall-Fortran-Präcompiler. Rechentech. Datenverarb., 11:10–12, 1980. CODEN RTDVAQ. ISSN 0300-3450.

Berns:1964:PDU

[Ber64] Gerald M. Berns. Preliminary description; use of IBJOB processor under IBSYS (FORTRAN IV, COBOL, IBMAP). Technical report, Computer Science Center, University of Maryland, College Park, MD, USA, 1964. various pp.

Bergman:1970:DBI

[Ber70a] Floyd R. Bergman. Differences between IBM 7094 FORTRAN IV, UNIVAC 1108 FORTRAN V, and USA STANDARD FOR-TRAN. Technical report, Systems Programming Branch, Computer Division, Analysis and Computation Directorate, White Sands Missile Range, NM, USA, 1970. various pp.

Bernstein:1970:LPF

[Ber70b] Herbert J. Bernstein. Ludeman’s PL/I FORMAC gradient problem as handled by FORTRAN SYMBOLANG. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(??):19–34, January 1970. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Berry:1976:ADD

[Ber76] G. Berry. Algorithm 92: The drawing of dashed lines. The Computer Journal, 19(4):361–363, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/
361.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/362.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
19/Issue_04/tiff/363.tif.

Bermudez:1977:EFS

[Ber77] Victor M. Bermudez. Ellips — a FORTRAN simulation of a polarization-modulation ellipsometer. Computer Physics Communications, 13(3):207–224, September/October 1977. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465577900157.

Bezner:1973:EFI

[Bez73] Hart C. Bezner. The Elements of Fortran IV. Goodyear Pub. Co., Pacific Palisades, CA, USA, January 1973. ISBN 0-87620-270-9. 249 pp. LCCN ???? US$13.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0876202709.

Bezanson:1975:TSP

[Bez75] William R. Bezanson. Teaching structured programming in FOR-TRAN with IFTRAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):196–199, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Bracchi:1971:LTG

[BF71] Giampio Bracchi and Domenico Ferrari. Language for treating geometric patterns in a twodimensional space. Comm. ACM, 14(1):26–32, January 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bankowski:1972:PWJ

[BF72] Jacek Bankowski and Konrad Fialkowski. Programowanie w jezyku Fortran. Seria: Biblioteka informatyki. Panstwowe Wydawn. Naukowe, Warszawa, Poland, wyd. 2 edition, 1972. 131 + [1] pp.

Bingulac:1979:LAS

[BF79] S. P. Bingulac and M. A. Farias. L-A-S (Linear Algebra and System) language and its use in control system education and research. Computing, 23(1):1–23, 1979. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Beghelli:1978:TSE

[BG78] Sergio Beghelli and Roberto Guidorzi. Teoria dei sistemi: esercizi e programmi Fortran. Patron, Bologna, Italy, 2 edition, 1978. 318 pp.

Brainerd:1978:FP

[BGG78] Walter S. Brainerd, Charles H.

Goldberg, and Jonathan L. Gross. Fortran 77 Programming. Harper & Row, New York, NY, USA, August 1978. ISBN 0-06-042394-3. xii + 360 pp. LCCN QA76.73 .F25 B73 1978. US$18.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0060423943.

Backus:1964:F

[BH64] J. W. Backus and W. P. Heising.

FORTRAN. IEEE Transactions on Electronic Computers, EC-13: 382–385, August 1964. CODEN IEECA8. ISSN 0367-7508.

Brown:1972:A

[BH72] W. S. Brown and A. D. Hall. AL-TRAN. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (24):4–7, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Bailey:1973:RCD

[BH73a] C. B. Bailey and T. H. Herrera.

Results of CDC 6600 and DEC PDP10 FORTRAN library functions for invalid arguments. Technical Report SLA–73-859, Sandia Labs, Albuquerque, NM, USA, December 1, 1973. 112 pp.

Burkowski:1973:AAC

[BH73b] Forbes J. Burkowski and W. D. Hoskins. ACM Algorithm 461: Cubic spline solutions to a class of functional differential equations [D2]. Comm. ACM, 16 (10):635–637, October 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:1971:CCP

[BH1 ] W. S. Brown and A. D. Hall. Computing Campbell’s Y_{2n} polynomials with ALTRAN. Computing Science Technical Report 6, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1971 (??). ?? pp.

Brent:1980:AIB

[BHY80] Richard P. Brent, Judith A. Hooper, and J. Michael Yohe. An AUGMENT interface for Brent’s multiple precision arithmetic package. ACM Transactions on Mathematical Software, 6(2):146–149, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Bre78a, Bre79b].

Biddulph:1979:MFC

[Bid79] Thomas P. Biddulph. A modified FORTRAN/ 77 compiler that will implement the proposed IEEE/ KCS floating point standard via calls to emulation routines. Master of science, plan ii., Dept. of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1979. ?? pp.

Biswas:1975:ECP

[Bis75] Nripendra Nath Biswas. Essentials of computer programming in FOR-TRAN IV. Radiant Books, Bangalore, India, 1975. xi + 125 pp.

Bitrakov:1975:FII

[Bit75] Dimitar Bitrakov. Fortran IV [i.e. cetiri]: programski jazik. Prosvetno delo, Univerzitetska pecatnica Kiril i Metodij, Skopje, Yugoslavia, 2., dop. izd. edition, 1975. 279 pp.

Barrett:1974:SCA

[BJ74] R. C. Barrett and B. W. Jordan, Jr. Scan conversion algorithms for a cell organized raster display. Comm. ACM, 17(3):157–163, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boehmer:1977:FBA

[BJ77] K. Boehmer and R. T. Jackson. A Fortran-Triplex-Pre-Compiler based on the Augment Pre-Compiler. MRC Technical Summary 1732, University of Wisconsin — Madison, Madison, WI, USA, 1977.

Beck:1972:CAR

[BK72] Robert E. Beck and Bernard Kolman. Computer approaches to the representation of Lie algebras. Journal of the ACM, 19(4):577– 589, October 1972. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bowdon:1975:TSF

[BK75] Edward K. Bowdon and Michael Kennedy. Ten statement Fortran plus Fortran IV: sensible, modular, and structured programming with WATFOR and WATFIV, second edition, [by] Michael Kennedy, Martin B. Solomon: instructor’s manual. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1975. ISBN 0-13-903427-7. iv + 359 pp.

Baitman:1977:AFP

[BK77] M. M. Baĭtman and H. I. Kilov. An algorithm and a FORTRAN program for determining the elements of optimal synthesis for a second order system of general form that is linear in the control. (Russian). Differencial′nye Uravnenija, 13(2): 358–361, 1977.

Bohlender:1980:FCN

[BKK^{+}80] G. Bohlender, E. Kaucher, R. Klatte, U. Kulisch, W. L. Miranker, Ch. Ullrich, and J. Wolff Von Gudenberg. Fortran for contemporary numerical computation. Research Report RC 8348, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1980. Published In: Computing 26, 277–314, 1981.

Bayes:1974:CNG

[BKW74] A. J. Bayes, J. Kautsky, and J. W. Wamsley. Computation in nilpotent groups (application). In Proc. Second Internat. Conf. Theory of Groups (Canberra, Australia, 1973), volume 372 of Lecture Notes in Math., pages 82– 89. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974.

Blatt:1960:CFU

[Bla60] John M. Blatt. Comments from a FORTRAN user. Comm. ACM, 3 (9):501–505, September 1960. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blatt:1967:IFI

[Bla67] John Markus Blatt. Introduction to Fortran IV programming. Goodyear Pub. Co., Pacific Palisades, CA, USA, 1967. 343 pp.

Blatt:1968:IFIa

[Bla68a] John Markus Blatt. Introduction to FORTRAN IV programming, using the Watfor compiler. Goodyear computer series. Goodyear Pub. Co., Pacific Palisades, CA, USA, 1968. xi + 313 pp.

Blatt:1968:IFIb

[Bla68b] John Markus Blatt. Introduction to FORTRAN IV programming, using the WATFOR compiler. Solutions manual. Goodyear Pub. Co., Pacific Palisades, CA, USA, 1968. 65 pp.

Blatt:1969:BFI

[Bla69] John Markus Blatt. Basic FOR-TRAN IV programming [version IBM 360]. Computer Systems (Aust.), Sydney, NSW, Australia, 1969. viii + 183 pp.

Blatt:1971:IFI

[Bla71] John Markus Blatt. Introduction to Fortran IV programming; using the Watfor/Watfiv compilers. Goodyear Pub. Co., Pacific Palisades, CA, USA, 1971. ISBN 0-87620-440-X. ix + 325 pp. LCCN QA76.5 .B54 1971 Sci-Eng.

Blanks:1979:GPC

[Bla79] Leonard Thomas Blanks. A general purpose cross-referencer for FORTRAN. Thesis (m.s. in s.s.), Louisiana State University, Baton Rouge, LA, USA, 1979. v + 228 pp.

Beebe:1980:SPR

[BLF80] Nelson H. F. Beebe, C. L. Lawson, and J. A. Flynn. SFTRAN3 Programmers Reference Manual. Technical report, College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, March 3 1980. 44 pp.

Blough:1968:IFS

[Blo68] David Kennedy Blough. Implementation of FORTRAN subroutine capability for the University of Florida time sharing computer system. Thesis (m.s. in engin.), University of Florida, Gainesville, FL, USA, 1968. iv + 79. pp.

Bloom:1971:SFI

[Blo71] Howard M. Bloom. A syntaxdirected FORTRAN interpreter for ALGOL 60. Technical report, H. Diamond Laboratories, Washington, DC, USA, 1971. 159 pp.

Reproduced by the National Technical Information Service, Springfield, VA.

Blumenberg:1965:FSD

[Blu65] Wayne Edward Blumenberg. Fortran simulation of a digital differential analyzer. Thesis (m.s.), Iowa State University, Ames, IA, USA, 1965. 87 pp.

Blum:1970:CCF

[Blu70] Harold A. Blum. A compact course in FORTRAN programming. Audio Tutorial Associates, inc., Dallas, TX, USA, 1970. vii + 90 pp.

Blue:1977:ANQ

[Blu77] James L. Blue. Automatic numerical quadrature. The Bell System Technical Journal, 56(9): 1651–1678, November 1977. CO-DEN BSTJAN. ISSN 0005-

8580. URL http://bstj.bell-
labs.com/BSTJ/images/Vol56/
bstj56-9-1651.pdf;   http://
www.alcatel-lucent.com/bstj/
vol56-1977/articles/bstj56-
9-1651.pdf.

Blue:1978:PFP

[Blu78] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. ACM Transactions on Mathematical Software, 4(1):15–23, March 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Blue:1977:PFP

[Blu7 ] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. Computing Science Technical Report 45, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Blake:1970:RPP

[BLY70] L. F. Blake, R. E. Lawson, and I. M. Yuille. A ring processing package for use with FOR-TRAN or a similar high-level language. The Computer Journal, 13 (1):40–47, February 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_01/130040.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/40.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
41.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/42.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
43.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/44.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
45.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/46.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/
47.tif.

Battani:1973:IDL

[BM73] G. Battani and H. Meloni. Interpreteur du langage de programmation Prolog. Technical report, Groupe d’Intelligence Artificielle, Université d’Aix-Marseille II, Marseille, France, 1973.

Barlott:1974:FIP

[BM74] P. J. Barlott and J. B. McQuitty.

Fortran IV program to predict the thermal environment within total confinement livestock housing. Research bulletin 74-1, Dept. of Agricultural Engineering, University of Alberta, Edmonton, Alberta, Canada, 1974. v + 177 pp.

Balfour:1979:PSFa

[BM79a] Alexander Balfour and David Howie Marwick. Programming in standard FORTRAN 77. Heinemann Educational Books, London, UK, 1979. ISBN 0-435-77485-9, 0-435-77486-7 (paperback). ix + 388 pp.

Balfour:1979:PSFb

[BM79b] Alexander Balfour and David Howie Marwick. Programming in Standard Fortran 77. North-Holland, Amsterdam, The Netherlands, August 1979. ISBN 0-444-19465-7. ix + 388 pp. LCCN QA76.73.F25 B34. US$39.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0444194657.

Braccini:1980:FGM

[BM80] C. Braccini and G. Marino. Fast geometrical manipulations of digital images. Computer Graphics and Image Processing, 13, 2:127– 141, 1980. CODEN CGIPBG. ISSN 0734-189X.

Busing:1962:FFC

[BML62] William R. Busing, K. O. Martin, and H. A. Levy. OR FLS, a Fortran crystallographic least-squares program. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1962. iii + 75 pp.

Busing:1964:FFC

[BML64] William R. Busing, K. O. Martin, and H. A. Levy. OR FFE, a Fortran crystallographic function and error program. Technical report, Oak RIdge National Laboratory, Oak Ridge, TN, USA, 1964. iii + 83 pp.

Bennett-Novak:1976:OMI

[BN76] Gordon Bennett-Novak. Optimized machine-independent extended FORTRAN for minicomputer programming. ACM SIG-PLAN Notices, 11(4):40–44, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boam:1969:FCD

[Boa69] William David Boam. Fortran compiler, drivers and loader for medical application. Thesis (m.a. 1969), Dept. of Computer Sciences, University of Utah, Salt Lake City, UT, USA, 1969. x + 106 pp.

Bober:1970:DIF

[Bob70] Vincent Ernest Bober. The design of an incore Fortran compiler. Thesis (m.s.), Newark College of Engineering, Newark, NJ, USA, 1970. 47 pp.

Boddy:1977:SFP

[Bod77] David E. Boddy. Structured Fortran: with or without a preprocessor. ACM SIGPLAN Notices, 12(4):34–39, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boguslavsky:1974:ECP

[Bog74] Boris W. Boguslavsky. Elementary Computer Programming in Fortran IV. Reston Publishing Co., Inc., Reston, VA, USA, August 1974. ISBN 0-87909-251-3. x + 325 pp. LCCN QA76.73.F25 B63. US$8.75; US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0879092513.

Boguslavsky:1980:ECP

[Bog80] Boris W. Boguslavsky. Elementary Computer Programming in Fortran IV. Reston Publishing Co., Inc., Reston, VA, USA, second edition, August 1980. ISBN 0-8359-1648-0. xi + 482 pp. LCCN QA76.73.F25 B63 1980. US$21.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0835916480.

Bohrer:1975:SAA

[Boh75] Robert Bohrer. Statistical algorithms: Algorithm AS 90: Onesided multi-variable inference. Applied Statistics, 24(3):380–384, September 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/90.

Boies:1974:UBI

[Boi74] B. J. Boies. User behavior on an interactive computer system. IBM Systems Journal, 1(13):2–18, 1974. CODEN IBMSA7. ISSN 0018-8670.

Boisson:1975:NOF

[Boi75] C. Boisson. Notions sur les organigrammes et sur le Fortran. INSA, Villeurbanne, France, 1975. 152 pp.

Boillot:1978:UF

[Boi78] Michel H. Boillot. Understanding Fortran. West Publishing Company, St. Paul, MN, USA, January 1978. ISBN 0-8299-0205-8. xi + 490 pp. LCCN QA76.73.F25B64. US$11.50.

Bollenbacher:1976:FPI

[Bol76] Gary Bollenbacher. FORTRAN program for induction motor analysis. NASA technical note; NASA TN D-8184 United States. National Aeronautics and Space Administration. NASA technical note; NASA TN D-8184. U.S. National Aeronautics snd Space Administration, Washington, DC. USA, 1976. iii + 81 pp. For sale by the National Technical Information Service.

Bomford:1967:VFP

[Bom67] Anthony Gerald Bomford. Varycord; a Fortran program for the least squares adjustment of horizontal control surveys. Technical report 6, Division of National Mapping, Dept. of National Development, Canberra, Australia, 1967. 86 pp.

Bond:1975:FFS

[Bon75] R. Bond. Free form structured FORTRAN translator. ACM SIG-PLAN Notices, 10(10):12–15, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bork:1967:FP

[Bor67] Alfred M. Bork. Fortran for physics. Addison-Wesley series in physics. Addison-Wesley, Reading, MA, USA, 1967. viii + 85 pp. LCCN QC20.2 .B6.

Bornzin:1969:FPS

[Bor69] Grant Owen Bornzin. A Fortran program for solution of generalized geometric programs. Thesis (m.s.e.), Arizona State University, Tempe, AZ, USA, 1969. 116 pp.

Boyle:1974:EF

[Boy74a] Thomas A. Boyle. Enough FOR-TRAN. Technical Publications, West Lafayette, IN, USA, 1974. v + 74 pp.

Boyse:1974:ECP

[Boy74b] John W. Boyse. Execution characteristics of programs in a pageon-demand system. Comm. ACM, 17(4):192–196, April 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyle:1975:NDU

[Boy75] James M. Boyle. Note on “The design of a user interface to large Fortran subroutine packages under IBM 360/370 OS”. ACM SIGNUM Newsletter, 10(2–3):40, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [TC75].

Boyle:1976:EF

[Boy76] Thomas A. Boyle. Enough FOR-TRAN. Technical Publications, West Lafayette, IN, USA, second edition, 1976. 89 pp.

Boyle:1980:EF

[Boy80] Thomas A. Boyle. Enough FOR-TRAN. Technical Directions, West Lafayette, IN, USA, fourth edition, 1980. ii + 66 pp.

Bauer:1974:BFI

[BP74] Charles R. Bauer and Anthony P. Peluso. Basic FORTRAN IV with WATFOR and WATFIV: selfinstructional manual and text. Addison-Wesley, Reading, MA, USA, June 1974. ISBN 0-201-00411-9. viii + 296 pp. LCCN QA76.73.F25 B38. US$23.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201004119.

Barrett:1976:CFC

[BP76] Charles A. Barrett and Alden F. Presler. COREST: a FORTRAN computer program to analyze paralinear oxidation behavior and its application to chromic oxide alloys. NASA technical note NASA TN D-8132, National Aeronautics and Space Administration, Washington, DC, USA, 1976. 79 pp. For sale by the National Technical Information Service.

Bono:1978:SSF

[BP78] P. Bono and M. Polisher. A survey of some Fortran implementations of the Core System. Computer Graphics, 12(4):67–73, December 1978. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Banks:1972:SFS

[BPW72] D. Banks, I. C. Percival, and J. McB. Wilson. Stirling FORDOC

01. a set of documentation conventions for FORTRAN packages and routines. Computer Physics Communications, 3(3):180–196, April/ May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465572900677.

Brier:1974:CSS

[BR74] Alan Brier and Ian C. Robinson. Computers and the Social Sciences. Columbia University Press, New York, NY, USA, 1974. ISBN 0-231-03914-X, 0-231-03915-8 (paperback). 284 pp. LCCN H61 .B6451 1974b WITHDRAWN.

Best:1975:SAAa

[BR75a] D. J. Best and D. E. Roberts.

Statistical algorithms: Algorithm AS 89: The upper tail probabilities of Spearman’s rho. Applied Statistics, 24(3):377–379, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/89.

Best:1975:SAAb

[BR75b] D. J. Best and D. E. Roberts. Statistical algorithms: Algorithm AS 91: The percentage points of the χ^{2} distribution. Applied Statistics, 24(3):385–388, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.
                                        cmu.edu/apstat/91.

Brant:1978:FTT

[BR78] Jeff Brant and Elizabeth L. Ross. FORTRAN: tele-type time-sharing computer terminals for use on Xerox Sigma 6 computer. Technical report, Clarion State College, Clarion, PA, USA, 1978. 20 pp.

Bradley:1972:CFS

[Bra72a] James W. Bradley. CHLOE: a Fortran subroutine for fitting ordinary differential equations to observed data. Memorandum report 2184, Marine Research Laboratories, ????, 1972. 45 pp. Distributed by U.S. National Technical Information Service, Springfield, VA, no. AD-743 878.

Brauch:1972:PF

[Bra72b] W. Brauch. Programmierung mit Fortran. Teubner, Stuttgart, Leipzig, 1972. ISBN 3-519-00013-X.

Brauch:1972:PMF

[Bra72c] W. Brauch. Programmierung mit Fortran. Teubner, Stuttgart, Leipzig, 1972. ISBN 3-519-00013-X.

Bray:1974:DSR

[Bra74] David W. Bray. Dynamic storage routines for Fortran programs. ACM SIGPLAN Notices, 9(9):2– 13, September 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brandt:1975:DMS

[Bra75a] S. Brandt. Datenanalyse — mit Statistischen Methoden und Computerprogrammen. Bibliographisches Institut, Mannheim, Germany, 1975. ISBN 3-411-01488-1.

Brandt:1975:DSM

[Bra75b] S. Brandt. Datenanalyse — mit statistischen Methoden und Computerprogrammen. Bibliographisches Institut, Mannheim, Germany, 1975. ISBN 3-411-01488-1.

Braun:1976:IMS

[Bra76] Christine L. Braun. An integrated microprocessor support software system. ACM SIGPLAN Notices, 11(4):57–65, April 1, 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brainerd:1977:PFL

[Bra77] Walt Brainerd. A proposal for a Fortran loop construct. ACM SIG-PLAN Notices, 12(12):60–67, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brainerd:1978:F

[Bra78] W. Brainerd. FORTRAN 77.

Comm. ACM, 21(10):??, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brainerd:1979:CMA

[Bra79] Walt Brainerd. A “core + modules” approach to Fortran standardization. In ACM [ACM79], pages 32–33.

Brenner:1967:TFP

[Bre67] N. M. Brenner. Three FORTRAN programs that perform the Cooley-Tukey Fourier transform. Technical note 1967-2, Lincoln Laboratory; available from the Clearinghouse for Federal Scientific and Technical Information, Lexington, MA, USA, 1967. iii + 30 pp.

Brent:1973:RRT

[Bre73] Richard P. Brent. Reducing the retrieval time of scatter storage techniques. Comm. ACM, 16(2): 105–109, February 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Modification of open addressing with double hashing to reduce the average number of probes for a successful search.

Brent:1974:AAG

[Bre74] Richard P. Brent. ACM Algorithm 488: a Gaussian pseudo-random number generator [G5]. Comm. ACM, 17(12):704–706, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brent:1975:FMP

[Bre75] R. P. Brent. A Fortran multipleprecision arithmetic package.

Technical report, Department of Computer Science, Australian National University, Canberra, Australia, 1975.

Brent:1976:FMA

[Bre76a] Richard P. Brent. A Fortran multiple-precision arithmetic package. Technical report, Department of Computer Science, Carnegie-Mellon University, Pittsburgh, PA, USA, 1976. 29 pp.

Brewer:1976:DSB

[Bre76b] Richard K. Brewer. Documentation standards for beginning students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1): 69–73, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Brent:1978:AMF

[Bre78a] Richard P. Brent. Algorithm 524: MP, A Fortran multipleprecision arithmetic package [A1]. ACM Transactions on Mathematical Software, 4(1):71–81, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Bre79b, BHY80, ?].

Brent:1978:FMP

[Bre78b] Richard P. Brent. A Fortran multiple-precision arithmetic package. ACM Transactions on Mathematical Software, 4(1):57– 70, March 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brennan:1979:HGF

[Bre79a] Robert L. Brennan. Handbook for GAPID: a FORTRAN IV computer program for generalizability analyses with single-facet designs. ACT Technical bulletin 34, Navy Personnel Research and Development Center, San Diego, CA, USA, 1979. 136 pp.

Brent:1979:RMF

[Bre79b] R. P. Brent. Remark on “Algorithm 524: MP, A Fortran multiple-precision arithmetic package [A1]”. ACM Transactions on Mathematical Software, 5(4): 518–519, December 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Bre78a, BHY80].

Brillinger:1967:BRB

[Bri67] D. R. Brillinger. Book review: A Guide to FORTRAN IV Programming, by Daniel D. McCracken. Journal of the Royal Statistical Society. Series A (General), 130(2): 258–259, ???? 1967. URL http://

                                        www.jstor.org/stable/2343418.

Bright:1968:PNA

[Bri68a] H. S. Bright. A proposed numerical accuracy control system interactive systems for experimental applied. Mathematics, pages 315– 330, 1968.

Bristol:1968:CCF

[Bri68b] William Arthur Bristol. CON-FORT conversational FORTRAN for the IBM system 360/50. Thesis (m.s. in engin.), University of Florida, Gainesville, FL, USA, 1968. vii + 111 pp.

Bright:1979:FCW

[Bri79] Herb Bright. FORTRAN comes to Westinghouse-Bettis, 1957. Annals of the History of Computing, 1(1): 72–74, July/September 1979. CO-DEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1979/pdf/a1072.
pdf;   http://www.computer.
org/annals/an1979/a1072abs.
htm.

Brown:1961:PBF

[Bro61] Brown University.Division of Applied Mathematics.Computation Center. Programming in basic FORTRAN. Brown University Computing Laboratory, Providence, RI, USA, 1961. 35 pp.

Brown:1971:BFP

[Bro71a] Charles E. Brown. A book of FOR-TRAN programs for the IBM 1130 computer. Technical report, Plymouth State College, Plymouth, NH, USA, 1971. 209 pp.

Brown:1971:FSA

[Bro71b] Richard S. Brown. FORTRAN syntax analysis using transition diagrams. Thesis (m.s.), Washington University, Sever Institute of Technology. Dept. of Applied Mathematics and Computer Science, St. Louis, MO, USA, 1971. vi + 56 pp.

Brown:1971:AUM

[Bro71c] W. S. Brown. ALTRAN user’s manual. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1971. ?? pp.

Brown:1973:BFP

[Bro73a] Charles E. Brown. A book of FOR-TRAN programs for the IBM system 3 computer. Technical report, Plymouth State College, Plymouth, NH, USA, 1973. 105 pp.

Brown:1973:AUM

[Bro73b] W. S. Brown. ALTRAN user’s manual. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973. ?? pp.

Brown:1974:AFP

[Bro74] Charles E. Brown. Advanced FORTRAN programs. Technical report, Plymouth State College, Plymouth, NH, USA, 1974. 46 pp.

Brown:1975:FPD

[Bro75] Gary Deward Brown. FOR-TRAN to PL/1 dictionary, PL/1 to FORTRAN dictionary. John Wiley and Sons, New York, London, Sydney, January 1975. ISBN 0-471-10796-4. xi + 204 pp. LCCN QA76.73.F25 B761. US$23.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0471107964.

Brown:1977:SP

[Bro77] P. J. Brown. Software Portability. Cambridge University Press, Cambridge, UK, 1977. ISBN 0-521-21485-8. xiv + 328 pp. LCCN QA76.6 .S635.

Brown:1980:GFG

[Bro80] T. Brown. General fast generation of random variables for discrete distribution. ACM Simuletter, 11, 4:73–75, 1980. CODEN SIMUD5. ISSN 0163-6103.

Brusberg:1966:BRD

[Bru66] H. Brusberg. Book review: D.

D. McCracken, A Guide to FOR-TRAN IV. Computing, 1(4): 371, 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Bryce:1975:MMA

[Bry75] G. Rex Bryce. Miscellanea:

MAD: an analysis of variance program for unbalanced designs. Applied Statistics, 24(3):350–352, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bryan:1961:FFS

[BS61] Norman William Bryan and John H. Suhrbier. Flexible format subroutines for 1620 FORTRAN. Publication 155; Technical report T62-3, Massachusetts Institute of Technology, School of Engineering Dept. of Civil Engineering, Civil Engineering Systems Laboratory, Cambridge, MA, USA, 1961. 10 + [2] pp.

Benson:1964:AMR

[BS64] Bill Benson and Dave Stevens. On avoiding matrix reversals between 7090 FORTRAN II and 7090 FOR-TRAN IV. Comm. ACM, 7(4):242, April 1964. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Basso:1973:IWW

[BS73a] David T. Basso and Ronald D. Schwartz. An introduction to the WATFOR and WATFIV programming languages. Charles E. Merrill Publishing Co., Columbus, OH, USA, 1973. ISBN 0-675-08944-1. vi + 341 pp. LCCN QA76.73.F25B37.

Beyer:1973:AAN

[BS73b] Terry Beyer and D. F. Swinehart. ACM Algorithm 448: Number of multiply-restricted partitions [A1]. Comm. ACM, 16(6):379, June 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brooker:1975:SAA

[BS75] P. Brooker and M. J. P. Selby. Statistical algorithms: Algorithm AS 92: The sample size for a distribution-free tolerance interval. Applied Statistics, 24(3):388–390, September 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/92.

Belgard:1978:CCS

[BS78] Richard A. Belgard and Victor B. Schneider. A comparison of the code space and execution time required for FOR-TRAN assignment statements on six computer architectures. ACM SIGMICRO Newsletter, 9(4):56– 64, December 1978. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/
10.1145/1014198.804308.

Bielecki:1980:AF

[BS80a] Jan Bielecki and Marek A. Suchenek. ANS FORTRAN. Publishing House of Warsaw Technical University, Warsaw, Poland, 1980. Textbook in Polish.

Boillot:1980:UW

[BS80b] Michel H. Boillot and Carol Rochlen Shingles. Understanding WAT-FIV. West Publishing Company, St. Paul, MN, USA, June 1980. ISBN 0-8299-0232-5. xiii + 567 pp. LCCN QA76.73.F25 B65 Sci-Eng. US$23.95; US$11.95.

Bashkow:1967:SDF

[BSK67] T. R. Bashkow, A. Sasson, and A. Kronfeld. System design of a FORTRAN machine. IEEE Transactions on Electronic Computers, EC-16(4):485–499, August 1967. CODEN IEECA8. ISSN 0367-7508.

Bonivento:1976:EPF

[BT76a] Claudio Bonivento and Alberto Tonielli. Esercizi e programmi Fortran per l’identificazione e la stima dei sistemi dinamici. Quaderni di controllo dei processi; 2. Patron, Bologna, Italy, 1976. 122 pp.

Bridge:1976:BTI

[BT76b] R. F. Bridge and E. W. Thompson. BRIDGES-a tool for increasing the reliability of references to FORTRAN variables. ACM SIG-PLAN Notices, 11(9):2–9, September 1976. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burkhardt:1965:MSS

[Bur65] Walter H. Burkhardt. Metalanguage and syntax specification. Comm. ACM, 8(5):304–305, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burroughs:1967:BBI

[Bur67] Burroughs Corporation.Equipment and Systems Marketing Division. Burroughs B5500, information processing systems FORTRAN reference manual. Burroughs Corp., Detroit, MI, USA, 1967. ix + 67 pp.

Burnside:1968:ECE

[Bur68a] R. Wayne Burnside. Elementary concepts of electronic data processing: FORTRAN IV and machine usage in the social sciences. Technical report, Indiana University, Bloomington, IN, USA, 1968. ii + 104 pp.

Burroughs:1968:BBI

[Bur68b] Burroughs Corporation. Burroughs B 5700 information processing systems: FORTRAN compiler reference manual. Burroughs Corporation, Detroit, MI, USA, 1968. various pp.

Burnside:1969:BED

[Bur69] R. Wayne Burnside. Basic electronic data processing and FOR-TRAN IV programming for sociologists. Technical report, University of Toronto, Toronto, Ontario, Canada, 1969. iv + 129 pp.

Burroughs:1970:BFS

[Bur70a] Burroughs Corporation. Burroughs Fortran self-learner. Burroughs Corporation, Detroit, MI, USA, 1970. 147 pp.

Burroughs:1970:BBI

[Bur70b] Burroughs Corporation.Systems Documentation. Burroughs B 6500 information processing systems: FORTRAN reference manual. Burroughs Corp., Detroit, MI, USA, 1970. various pp.

Burr:1971:FCF

[Bur71] Edmund John Burr. A first course in Fortran. Dept. of University Extension, University of New England, Armidale, NSW, Australia, 1971. ISBN 0-85834-010-0. 83 pp. LCCN QA76.73.F25B87.

Burkhardt:1972:PFC

[Bur72] Diana Burkhardt. Programming in FORTRAN: a course in note form. University of Birmingham, Computer Centre, Birmingham, England, 1972. ISBN 0-7044-0005-7. [3] + 63 pp. LCCN QA76.73.F25B85.

Burroughs:1973:MSF

[Bur73a] Burroughs Corporation. Medium Systems FORTRAN reference manual. The Corporation, Detroit, MI, USA, 1973. 231 pp.

Burroughs:1973:BMS

[Bur73b] Burroughs Corporation.Documentation Dept. Burroughs Medium Systems Fortran Linear Programming (FLP) reference manual. The Corporation, Detroit, MI, USA, 1973. 98 pp.

Burroughs:1974:BBB

[Bur74] Burroughs Corporation. Burroughs B 6700/B 7700 FORTRAN reference manual. The Corporation, Detroit?, 1974. various pp.

VanBuren:1976:FCP

[Bur76] Van A. L. (Arnie Lee) Buren. A Fortran computer program for calculating the linear prolate functions. NRL report 7994, U.S. Dept. of Defense, Dept. of the Navy, Office of Naval Research, Naval Research Laboratory; Springfield, VA, USA, Washington, DC. USA, 1976. iii + 41 pp. For sale by the National Technical Information Service.

Burroughs:1978:BBB

[Bur78] Burroughs Corporation. Burroughs B 7000/B 6000 series: FORTRAN reference manual. Burroughs Corp., Detroit, MI, USA, 1978. various pp.

Burton:1979:FPS

[Bur79] W. Burton. Fortran preprocessor to support encapsulated data abstraction definitions. The Computer Journal, 22(4):307–312, November 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Businger:1967:LEF

[Bus67] Peter A. Businger. Letter to the Editor: FORTRAN subroutines for scalar products. Comm.

ACM, 10(5):260, May 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

BITI:1968:FMS

[Bus68] Business Information Technology, Inc. FORTRAN manual: series 480. Business Information Technology, Inc., Natick, MA, USA, 1968. ???? pp.

Butler:1966:BFI

[But66] Lila M. Butler. BRLESC Fortran II computer programs for analytical solutions to the general problem of photogrammetry. BRL memorandum report 1769, Ballistic Research Laboratories, Aberdeen Proving Ground, MD, USA, 1966. 153 pp.

Bellehumeur:1974:EDP

[BV74] Marcel Bellehumeur and Jean-Pierre Valard. Elements d’informatique et programmation FORTRAN. Lidec, Montréal, PQ, Canada, 1974. ISBN 0-7762-6080-4. 238 pp.

Bobrow:1964:LPE

[BW64] Daniel G. Bobrow and Joseph Weizenbaum. List processing and extension of language facility by embedding. IEEE Transactions on Electronic Computers, EC-13: 395–400, August 1964. CODEN IEECA8. ISSN 0367-7508.

Buckholtz:1975:APR

[BW75] T. J. Buckholtz and C. S.

Wetherell. Algorithm 89: a program to referee Kriegspiel and chess. The Computer Journal, 18(2):177–183, May 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.
oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/
177.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/178.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
18/Issue_02/tiff/179.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_
02/tiff/180.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/
181.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/182.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
18/Issue_02/tiff/183.tif.

Barrodale:1978:FPS

[BW78a] I. Barrodale and K. B. Wilson. A Fortran program for solving a nonlinear equation by Muller’s method. Journal of Computational and Applied Mathematics, 4(2):159–166, June 1978. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0771050X78900414.

Bellamy:1978:ICP

[BW78b] Clifford John Bellamy and Leonard George Whitehouse. An introduction to computer programming in FOR-TRAN: MONECS/FORTRAN. Sorrett Publishing, Malvern, Victoria, Australia, second edition, 1978. ISBN 0-909752-41-9. 366 pp.

Bock:1973:MLA

[BY73] R. Darrell Bock and George Yates.

MILTIQUAL: log-linear analysis of nominal or ordinal qualitative data by the method of maximum likelihood: a FORTRAN IV program. National Educational Resources, Inc., Chicago, IL, USA, version II edition, 1973. ISBN 0-89498-008-4. 35 pp.

Boehmer:1978:TAF

[BY78] K. Boehmer and J. M. Yohe.

Triplex arithmetic for Fortran. MRC Technical Summary 1901, University of Wisconsin — Madison, Madison, WI, USA, 1978. 38 pp.

Conway:1978:PPG

[CA78] Richard Walter Conway and James Elson Archer. Programming for Poets: a Gentle Introduction Using Fortran With Watfiv. His Programming for poets series. Conway, Richard Walter, 1931- Programming for poets series. Winthrop Publishers, Cambridge, MA, USA, June 1978. ISBN 0-87626-722-3. xiv + 332 pp. LCCN QA 76.73 F25 C657. US$12.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0876267223. Contributions by Charles Van Loan and Kathryn Conway.

Cadete:1971:LFI

[Cad71] Maria Odete Rodrigues Cadete.

A linguagem Fortran IV; analise da definiçao proposta pelo American National Standards Institute. Technical Report ????, Instituto Gulbenkian de Ciencia, Centro de Calculo Cientifico, Oeiras, Portugal, 1971. vii + 176 pp.

Cadete:1979:FIP

[Cad79] Antonio M. F. Cadete. Fortran IV — programacão logica. Technical Report ????, Centro de calculo cientifico, Oeiras, Portugal, 1979. 38 pp.

Calderbank:1969:CPF

[Cal69a] Valerie Joyce Calderbank. A course on programming in FOR-TRAN IV. Science paperbacks. Chapman and Hall, Ltd., London, UK, 1969. ISBN 0-412-09250-6, 0-412-20640-4. viii + 88 pp. LCCN QA76.5.C34 1969; TK 7888.3 C123c. Distributed in the U.S.A. by Barnes and Noble.

Calcomp:1969:UFB

[Cal69b] California Computer Products, Inc. USAS FORTRAN business: user’s manual. CalComp, Anaheim, CA, USA, revised edition, 1969. 22 pp.

Calcomp:1969:UFD

[Cal69c] California Computer Products, Inc. USAS FORTRAN drafting: user’s manual. CalComp, Anaheim, CA, USA, revised edition, 1969. 13 pp.

Calcomp:1969:UFG

[Cal69d] California Computer Products, Inc. USAS FORTRAN general: user’s manual. CalComp, Anaheim, CA, USA, revised edition, 1969. 18 pp.

Calcomp:1969:UFS

[Cal69e] California Computer Products, Inc. USAS FORTRAN scientific: user’s manual. CalComp, Anaheim, CA, USA, 1969. 25 pp.

CSU:1972:SF

[Cal72] California State University, Sacramento, Computer Center. SWIFT FORTRAN. Technical report, Computer Center, California State University, Sacramento, Sacramento, CA, USA, 1972. various pp.

Calderbank:1978:KPF

[Cal78] Valerie Joyce Calderbank. Kurs programmirovaniia na FORTRANe-IV. Biblioteka po avtomatike; vyp. 592 Biblioteka po avtomatike; vyp.

592. Energiia, Moskva, USSR, izd. 2., dop. edition, 1978. 86 pp.

Campbell:1965:FPQ

[Cam65] Robert William Campbell. A FORTRAN program for a Quine-McCluskey reduction of N Boolean variables. Thesis (m.s.), George Washington University, Washington, DC, USA, 1965. vi + 81 pp.

Campbell:1977:FPC

[Cam77] David L. Campbell. FORTRAN programs to calculate differential stress concentration factors around two-dimensional circular and elliptical inclusions in an edge-loaded plate. Reports-Open file series 77-459, U.S. Geological Survey, Denver, CO, USA, 1977. 16 pp.

Cannon:1977:TLS

[Can77] Francis Robert Cannon. The transferability of learning of syntax between COBOL and FOR-TRAN. Thesis, Temple University, Philadelphia, PA, USA, 1977. vii + 134. pp.

Carey:1966:SAF

[Car66] Mary Clo Carey. SYMTRAN: the addition of formal algebraic manipulative capabilities to FOR-TRAN with format. Thesis (m.s.), University of Southwestern Louisiana, Lafayette, LA, USA, 1966. 1962 pp.

Carnahan:1968:IDC

[Car68] Brice Carnahan. Introduction to digital computing and FORTRAN IV with MTS applications. Technical report, University of Michigan, Ann Arbor, MI, USA, 1968. various pp.

Carver:1969:IFI

[Car69] D. K. Carver. Introduction to Fortran II and Fortran IV programming. John Wiley and Sons, New York, London, Sydney, 1969. ISBN 0-471-13860-6. xv + 224 pp. LCCN QA76.5.C38.

Carver:1974:IDP

[Car74a] D. K. Carver. Introduction to Data Processing. John Wiley and Sons, New York, London, Sydney, 1974. ISBN 0-471-13854-1. xxi + 518 pp. LCCN QA76 .C362.

Carver:1974:FFP

[Car74b] M. B. Carver. FORSIM: a Fortran package for the automated solution of coupled partial and/or ordinary differential equation systems user’s manual. Technical report, Chalk River Nuclear Laboratories, Chalk River, Ontario, Canada, 1974. 91 pp.

Carlson:1977:TFP

[Car77] Leland A. Carlson. Trandes: a Fortran program for transonic airfoil analysis or design. NASA contractor report NASA CR-2821, National Aeronautics and Space Administration, Washington, DC. USA, 1977. 109 pp. For sale by the National Technical Information Service.

Carril:1978:ICP

[Car78a] Horacio J. Carril. Introduccion a la computacion y programacion FORTRAN. Coleccion Ciencia y tecnologia. Ediciones Marymar, Buenos Aires, Argentina, 1978. 252 pp.

Carroll:1978:CFT

[Car78b] Johnny Glen Carroll. A comparison of the formatted transput of ALGOL 68, PL/I, and FOR-TRAN. Thesis (m.s.), Oklahoma State University, Stillwater, OK, USA, 1978. vii + 197 pp.

Carver:1979:IBDa

[Car79a] D. K. Carver. Introduction to business data processing: with Basic, Fortran, and Cobol programming. John Wiley and Sons, New York, London, Sydney, second edition, 1979. ISBN 0-471-03091-0. x + 366 pp. LCCN QA 76 C3619 1979.

Carver:1979:IBDb

[Car79b] D. K. Carver. Introduction to Business Data Processing: With Basic, Fortran, and COBOL Programming: Student Workbook. John Wiley and Sons, New York, London, Sydney, second edition, April 1979. ISBN 0-471-03998-5. ???? pp. LCCN ???? US$12.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0471039985.

Carver:1979:SWI

[Car79c] D. K. Carver. Student workbook [to] Introduction to business data processing: with Basic, Fortran, and Cobol programming. John Wiley and Sons, New York, London, Sydney, second edition, 1979. 147 pp.

Caswell:1962:FCC

[Cas62] Randall S. Caswell. A FORTRAN code for calculation of eigenvalues and eigenfunctions in real potential wells. Technical note 159, National Bureau of Standards, Washington, DC. USA, 1962. ii + 29 pp. For sale by the Supt. of Docs. U.S. Govt. Print. Off.

Cauthen:1978:FCM

[Cau78] Sibyl Sumner Cauthen. A FOR-TRAN cross-compiler for the Motorola M6800 microprocessor. Thesis (m.s.), University of South Carolina, Columbia, SC, USA, 1978. 122 pp.

Carasso:1969:FICa

[CB69] Mary Carasso and Jona Bargur. A FORTRAN IV computer program for preparing a regional interindustry transactions table from secondary data sources. Open file report 69-18, California University, [Berkeley Computer Center], Berkeley, CA, USA, 1969. 97–123. pp.

Cadete:1970:SPF

[CC70] Maria Odete Rodrigues Cadete and Antonio M. F. Cadete. Sistema de programacão Fortran II; e sua utilizacao com um computador IBM 1620. Cursos e seminarios, Instituto Gulbenkian de Ciencia, Centro de Calculo Cientifico, Lisboa, Portugal, 1970. 160 pp.

Cohen:1974:NDF

[CC74] J. Cohen and Eileen Carton. Nondeterministic Fortran. The Computer Journal, 17(1):44–51, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Clark:1967:PSFa

[CCHT67a] N. A. Clark, W. J. Cody, K. E.

Hillstrom, and E. A. Thieleker. Performance statistics of the FOR-TRAN IV (H) library for the IBM System/360. Technical Report ANL-7231, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1967. Reprinted in SHARE Secretary Distribution, SDD 169, C4473, pp. 12 46.

Clark:1967:PSFb

[CCHT67b] N. A. Clark, W. J. Cody, K. E. Hillstrom, and E. A. Thieleker. Performance statistics of the Fortran IV (H) library for the IBM System/360. ACM SIGNUM Newsletter, 2(3):??, December 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Carasso:1969:FICb

[CCL69] Mary Carasso, Meir Carasso, and Everard Mervyn Lofting. A FOR-TRAN IV computer program for manipulating Leontief type models. Open file report 69-19, Computer Center, University of California, Berkeley, Berkeley, CA, USA, 1969. 124–159 pp.

Castaneda:1979:PTP

[CCN^{+}79] Fernando Castaneda, Frederick Chow, Peter Nye, Dan Sleator, and Gio Wiederhold. PCFORT — a Fortran-to-Pcode translator. Technical Report STAN-CS-79-714, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. vi +

85 pp. URL http://i.stanford.
edu/pub/cstr/reports/cs/tr/
79/714/CS-TR-79-714.pdf.

Cress:1968:FIW

[CDG68] Paul Cress, Paul Dirksen, and James Wesley Graham. FOR-TRAN IV with WATFOR. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1968. xii + 380 pp.

Cress:1970:FIW

[CDG70] Paul Cress, Paul Dirksen, and James Wesley Graham. Fortran IV With Watfor and Watfiv. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, revised edition, April 1970. ISBN 0-13-329433-1. xv + 447 pp. LCCN QA76.5.C73 1970. US$32.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

0133294331. First ed. published under title: FORTRAN IV with WATFOR.

Cress:1973:FIC

[CDG73] Paul Cress, Paul Dirksen, and James Wesley Graham. FOR-TRAN IV con WATFOR y WAT-FIV. Editorial Prentice/Hall Internacional, Bogota, Colombia, traduccion y adaptacion; Ing. Manuel F. Diaz edition, 1973. ISBN 0-13-329698-9. xvi + 407 pp.

Cress:1980:FIC

[CDG80a] Paul Cress, Paul Dirksen, and James Wesley Graham. FOR-TRAN IV con WATFOR y WAT-FIV. Prentice-Hall International, Englewood Cliffs, NJ 07632, USA, 1980. ISBN 84-237-0476-9. 407 pp.

Cress:1980:SFW

[CDG80b] Paul Cress, Paul Dirksen, and James Wesley Graham. Structured FORTRAN with WATFIV-S. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1980. ISBN 0-13-854752-1 (paperback). xi + 403 pp.

Cowan:1976:DES

[CDGW76] D. D. Cowan, P. H. Dirksen, J. W. Graham, and J. W. Welch. Development of educational software using the DEC PDP-11. ACM SIGPLAN Notices, 11(4):109–112, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooney:1975:DTS

[CDH75] Thomas J. Cooney, Edward J. Davis, and Kenneth B. Henderson. Dynamics of teaching secondary school mathematics. Houghton Mifflin, Boston, MA, USA, 1975. ISBN 0-395-14028-5. x + 448 pp. LCCN QA 11 C783d.

Carter:1960:SPR

[CF60] Edward J. D. Carter and Robert Peel Futrelle. Symbol pattern recognition within Fortran: the SHADOW III system. Programming note 30, Massachusetts Institute of Technology, Solid State and Molecular Theory Group, Cambridge, MA, USA, 1960. 87 pp.

Clementi:1971:MAS

[CF71] J. F. Clementi and A. P. Fletcher. Modifications to the APL 1130/ system to provide more convenient operating on a Fortran user’s machine. ACM SIGPLAN Notices, 6(6):16–18, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlile:1968:FPE

[CG68] Robert E. Carlile and Billy E. Gillett. FORTRAN programming for engineers. University of Missouri–Rolla, Rolla, MO, USA, 1968. 94 pp.

Carlile:1973:FCM

[CG73] Robert E. Carlile and Billy E.

Gillett. FORTRAN and computer mathematics for the engineer and scientist. Petroleum Pub. Co., Tulsa, OK, USA, 1973. ISBN 0-87814-016-6. xv + 520 pp. LCCN QA76.73.F25C37.

Char:1991:MLVb

[CGG^{+}91] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple Library V Reference Manual. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97592-6, 3-540-97592-6. xxv + 698 pp. LCCN QA155.7.E4 M353 1991.

Cockayne:1975:LAD

[CGH75] E. Cockayne, S. Goodman, and S. Hedetniemi. A linear algorithm for the domination number of a tree. Information Processing Letters, 4(2):41–44, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chadwick:1966:AGP

[Cha66] Harold K. Chadwick. Age and growth program, Univac 1107, Fortran IV. Transactions of the American Fisheries Society, 95(2):229– 230, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Challe:1967:CPF

[Cha67] A. Challe. Cours de programmation Fortran: Gamma 30 S. Technical Report ????, Bureau des Longitudes, Paris, France, 1967. 90 pp.

Charmonman:1970:CSA

[Cha70] S. Charmonman. A comparison of the structures of APL, FORTRAN, ALGOL and PL/I. ACM SIGPLAN Notices, 5(3):22– 24, March 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:1971:AAP

[Cha71a] J. M. Chambers. ACM Algorithm 410: Partial sorting [M1]. Comm. ACM, 14(5):357–358, May 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chambers:1971:ARF

[Cha71b] J. M. Chambers. Another round of FORTRAN. The Computer Journal, 14(3):312– 314, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        14/Issue_03/140312.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_03/tiff/312.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/313.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_03/tiff/

314.tif. See correspondence [Dew72, Fla72, Cha72].

Charatsis:1971:CPE

[Cha71c] E. G. Charatsis. A computer program for estimation of the constant elasticity of substitution production function. Applied Statistics, 20(3):286–296, 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chambers:1972:CAR

[Cha72] J. M. Chambers. Correspondence:

On “Another round of FOR-TRAN”. The Computer Journal, 15(4):381, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
15/Issue_04/150381.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/381.tif. See

[Cha71b, Dew72, Fla72].

Chance:1973:BRB

[Cha73] E. M. Chance. Book review: FOR-TRAN Techniques, by A. Colin Day, 1972. The Computer Journal, 16(3):265, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
16/Issue_03/tiff/265.tif.

Chang:1974:ASD

[Cha74] Y. F. Chang. Automatic solution of differential equations. In D. L. Colton and R. P. Gilbert, editors, Constructive and Computational Methods for Differential and Integral Equations, volume 430 of Lecture Notes in Mathematics, pages 61–94. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974.

Chang:1976:MEF

[Cha76] Julie Chou Chang. Modular extensions to FORTRAN: a project in computer science. Thesis (m.s.), University of Missouri, Columbia, Columbia, MO, USA, 1976. 32 pp.

Chapman:1977:SEF

[Cha77] Robert Claude Chapman. SPEC-TRAL: an extension of FORTRAN for spectrophotometry. Thesis (m.s.), University of California, Los Angeles, Los Angeles, CA, USA, 1977. ix + 165 pp.

Chady:1979:TFA

[Cha79a] Stephen G. Chady. Tools for the flow analysis of FORTRAN programs. Thesis (meng), University of Louisville, Department of Applied Mathematics and Computer Science, Louisville, KY, USA, 1979. vii + 145 pp.

Charet:1979:FIA

[Cha79b] Gerard Charet. Fortran: initiation au langage de l’informatique scientifique. Sedes-Informatique. Societé d’éditions d’enseignement supérieur, Paris, France, 3e edition, 1979. xiii + 271 pp.

Chirlian:1973:IFI

[Chi73] Paul M. Chirlian. Introduction to Fortran IV; with timeshare and batch operation. Academic Press, New York, NY, USA, 1973. ISBN 0-12-172850-1. xii + 291 pp. LCCN QA76.73.F25 C47.

Carver:1977:EAQ

[CJ77] M. B. Carver and V. J. Jones.

An evaluation of available quadrature algorithms and selection for the A E C L Fortran mathematical library. Technical report, Chalk River Nuclear Laboratories, Chalk River, Ontario, Canada, 1977. 137 pp.

Cazacu:1978:PLF

[CJ78] Constantin Cazacu and T. Jucan. Programarea in limbajul FORTRAN: calculatorul FELIX C-256. Junimea, Iasi, ??, 1978. viii + 213 pp.

Cole:1967:FIP

[CJM67] Alfred John Cole, C. Jordan, and Daniel Francis Merriam. Fortran II program for progressive linear fit of surfaces on a quadratic base using an IBM 1620 computer. Computer contribution 15, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 54 pp.

Chopra:1980:FIP

[CK80] M. G. Chopra and Ram Kumar. FORTRAN IV programming: for engineering, theoretical physics, applied mathematics, statistics, operation research, industrial management, econometrics, psychometrics and other disciplines involving data analysis. Vikas, Ghaziabad, U.P., India, 1980. ISBN 0-7069-1040-0. 248 pp.

Crary:1970:SMA

[CL70] F. D. Crary and T. D. Ladner. A simple method of adding a new data type to FORTRAN. MRC Technical Summary 1065, Mathematics Research Center, University of Wisconsin, Madison, WI, USA, 1970. 120 pp.

Cutshall:1980:BBF

[CL80] N. H. Cutshall and I. L. Larsen. BGSUB and BGFIX, FOR-TRAN programs to correct Ge(Li) gamma-ray spectra for photopeaks from radionuclides in background. Environmental Sciences Division publication no. 1415 ORNL/TM 7051, Dept. of Energy, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1980. xi + 20 pp. For sale by the National Technical Information Service.

Clark:1968:DFI

[Cla68] Howard Evans Clark. A discoriented FORTRAN IV processor for the GE225 computer. Thesis (m.s.), Lehigh University, Bethlehem, PA, USA, 1968. iii + 33 pp.

Clark:1973:FSA

[Cla73a] Deborah Livingston Clark. A factorial study of arithmetic ability of male students and a formulation of a Fortran program for factor analysis. Thesis (m.a.), Oklahoma State University, Stillwater, OK, USA, 1973. vii + 32 pp.

Clark:1973:PTO

[Cla73b] J. L. Clark. PATRICIA-II. twolevel overlaid indexes for large libraries. International Journal of Computer and Information Sciences, 2(4):269–292, December 1973. CODEN IJCIAH. ISSN 0091-7036.

Claudson:1975:DSR

[Cla75] R. Mark Claudson. Digital simulation of river plankton population dynamics. Comm. ACM, 18 (9):517–523, September 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clancy:1978:CIF

[Cla78] B. E. Clancy. COMFORT, an interactive FORTRAN system for the IBM 360 computer. AAEC/E;

454. Research Establishment, Australian Atomic Energy Commission, Lucas Heights, NSW, Australia, 1978. ISBN 0-642-59659-X. 20 pp.

Clay:1980:ESD

[Cla80] Laurence E. Clay. An empirical study of data flow bandwidth in FORTRAN programs. Thesis (m.a.), University of Texas at Austin, Austin, TX, USA, 1980. vi + 56 pp.

Cleary:1966:FTS

[Cle66] J. G. Cleary. A Fortran technique for simplifying input to report generators. Comm. ACM, 9(6):441–442, June 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cleveland:1968:FID

[Cle68] S. Thomas Cleveland. FORTRAN II-D on the IBM 1620; introduction to digital computer programming. Dickenson Pub. Co., Encino, CA, USA, 1968. ix + 128 pp.

Cleveland:1970:PFI

[Cle70] George Lindsay Cleveland. PL/I FORTRAN interface system. Thesis (m.s.), University of North Carolina at Chapel Hill, Chapel Hill, NC, USA, 1970. various pp.

Cline:1974:AAS

[Cli74a] A. K. Cline. ACM Algorithm 476: Six subprograms for curve fitting using splines under tension [E2]. Comm. ACM, 17(4):220–223, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cline:1974:SPC

[Cli74b] A. K. Cline. Scalar- and planar-valued curve fitting using splines under tension. Comm. ACM, 17:218–220, 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cline:1974:SPV

[Cli74c] A. K. Cline. Scalar- and planarvalued curve fitting using splines under tension. Comm. ACM, 17 (??):218–220, ?? 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cliff:1978:CFSa

[Cli78a] Steven B. Cliff. A collection of FORTRAN support routines. Technical Report K/CSD/TM-20, Oak Ridge Gaseous Diffusion Plant; Available from the National Technical Information Service, U.S. Dept. of Commerce, Oak Ridge, TN, USA, 1978. 136 pp.

Cliff:1978:CFSb

[Cli78b] Steven B. Cliff. A collection of FORTRAN support routines. Technical Report K/CSD/TM-20, Union Carbide Corp., Nuclear Division, ????, January 1978. 114 pp.

Cloete:1972:FFT

[Clo72] Johannes H. Cloete. The fast Fourier transform applied to stationary time series analysis using FORTRAN IV: research project. Thesis (m.s. in electrical engineering), University of California, Berkeley, Berkeley, CA, USA, June 1972. 41 pp.

Carleton:1964:FEF

[CLS64] J. T. Carleton, P. E. Lego, and R. M. Suarez. A FORTRAN extension to facilitate proposal preparation. IEEE Transactions on Electronic Computers, EC-13:456–462, August 1964. CODEN IEECA8. ISSN 0367-7508.

Chen:1978:FCB

[ClW78] Suan Chen and Chiu lung Wang. A Fortran code of bivariate interpolation and smooth surface fitting. INER 0236, Institute of Nuclear Energy Research, Atomic Energy Council, Lungtan, Taiwan, 1978. iii + 51 pp.

Caswell:1966:FPC

[CM66] Randall Smith Caswell and Leonard C. Maximon. Fortran programs for the calculation of Wigner 3j, 6j, and 9j coefficients for angular momenta ≤ 80. Technical note 409, United States Government Printing Office, Washington, DC, USA, 1966. 65 pp.

Chirlian:1979:ISF

[CM79] Paul M. Chirlian and Merl K. Miller. An Introduction to Structured Fortran. Matrix Publishers, Portland, OR, USA, March 1979. ISBN 0-916460-07-X. xii + 458 pp. LCCN QA76.73.F25 C48. US$21.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
091646007X.

Chow:1980:UFU

[CNW80] Frederick Chow, Peter Nye, and Gio Wiederhold. UFORT: a fortran-to-universal PCODE translator. Technical Report CSL-TR-79-168, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1980. iv + 88

pp. URL http://i.stanford.
edu/pub/cstr/reports/csl/tr/
79/168/CSL-TR-79-168.pdf.

Coan:1980:BF

[Coa80] James S. Coan. Basic Fortran. Hayden computer programming series. Hayden Book Co., Rochelle Park, NJ, USA, October 1980. ISBN 0-8104-5168-9. 235 pp. LCCN QA76.73.F25 .C62. US$17.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0810451689.

Cobb:1975:IPS

[Cob75] Gary W. Cobb. The impact of parallelism on software. In IEEE SCA ’75 [IEE75], pages 220– 222. LCCN QA76.6.S919 1975.

URL http://www.acsel-lab.
com/arithmetic/arith3/papers/
ARITH3_Cobb.pdf. IEEE order

number CH1017-3C.

Cochran:1960:FMF

[Coc60] Charles Nelson Cochran. FOR-TRAN manual, with FORGO supplement. Technical report, West Virginia University, Morgantown, WV, USA, 1960. various pp.

Cockayne:1980:FPC

[Coc80] E. J. Cockayne. A Fortran program for computation of extremum dominating, independent and irredundant sets in graphs. Internal report 209, University of Victoria, Dept. of Mathematics, Victoria, BC, Canada, 1980. 2 + [6]. pp.

Cody:1967:CFI

[Cod67] W. J. Cody. Critique of the FOR-TRAN IV (H) library for the IBM System/360. SHARE Secretary Distribution, SSD 169(C4473):4– 11, 1967.

Cohn:1966:INI

[Coh66] Charles Erwin Cohn. Incorporation of nonstandard input/output devices into FORTRAN systems. Comm. ACM, 9(5):343–344, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohn:1974:FPB

[Coh74] Charles E. Cohn. FORTRAN programming for a batch-transmit CRT terminal. Software—Practice and Experience, 4(4):402–404, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cole:1975:LFC

[Col75] John Patrick Cole. A load-and-go FORTRAN compiler. Thesis (m.s.), Illinois Institute of Technology, Chicago, IL, USA, 1975. iv + 52 pp.

Collens:1976:SFW

[Col76] P. A. Collens. Structured FOR-TRAN with WATFIV-S and WATFOR-llS. Charles Babbage Research Centre, St. Pierre, Manitoba, Canada, 1976. 517 pp.

Cole:1978:AFI

[Col78a] J. W. Perry Cole. ANSI Fortran IV: a structured programming approach. Wm. C. Brown Publishers, Dubuque, IA, USA, March 1978. ISBN 0-697-08125-7. xv + 420 pp. LCCN QA76.73.F25C64. US$32.48. URL http://www.

                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=0697081257.

Cole:1978:IMA

[Col78b] J. W. Perry Cole. Instructor’s manual to accompany ANSI Fortran IV: a structured programming approach. Wm. C. Brown Publishers, Dubuque, IA, USA, 1978. iv + 101 pp.

Coleman:1980:FSB

[Col80a] J. P. Coleman. A Fortran subroutine for the Bessel function J_{n}(x) of order 0 to 10. Computer Physics Communications, 21 (1):109–118, December 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Colman:1980:CMF

[Col80b] S. M. Colman. CHEMANAL: a MULTICS Fortran program to calculate chemical weathering data. Open-file report 80-844, U.S. Geological Survey, Denver, CO, USA, 1980. 22 pp.

CUC:1969:CUF

[Com69] Computer Usage Company. Computer Usage; 360 Fortran Programming. McGraw-Hill, New York, NY, USA, January 1969. ISBN 0-07-012381-0. ???? pp. LCCN ???? US$20.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070123810.

Comer:1978:MII

[Com78] D. Comer. MOUSE4: An improved implementation of the RATFOR preprocessor. Software —Practice and Experience, 8(1): 35–40, January/February 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Comfort:1980:LFP

[Com80a] J. R. Comfort. Linkage of Fortran programs with partitioned data sets on IBM, CDC, and DEC computers. Computer Physics Communications, 19(1): 43–49, January/March 1980. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/001046558090065X.

CI:1980:HAC

[Com80b] Compensation Institute. Handbook of analytical computer programs (FORTRAN) applied in compensation. Compensation Institute, Pacific Palisades, CA, USA, 1980. v + 167 pp.

CDC:1962:FRMa

[Con62a] Control Data Corporation. 160 FORTRAN reference manual; Control Data 160 computer. Control Data Corporation, Minneapolis, MN, USA, 1962. 91 + 5 pp.

CDC:1962:FRMc

[Con62b] Control Data Corporation. 160 FORTRAN reference manual; Control Data 160 computer. Control Data Corporation, Minneapolis, MN, USA, 1962. 91 + 5 pp.

CDC:1962:FRMb

[Con62c] Control Data Corporation. FOR-TRAN 62/reference manual. Control Data Corporation, Minneapolis, MN, USA, 1962. vii + 101 pp.

CDC:1962:FRMd

[Con62d] Control Data Corporation. FOR-TRAN 62/reference manual. Control Data Corporation, Minneapolis, MN, USA, 1962. vii + 101 pp.

CDC:1964:CSF

[Con64a] Control Data Corporation. 3200 computer system FORTRAN reference manual. Control Data Corporation, Palo Alto, CA, USA, 1964. 129 pp.

CDC:1964:CDCb

[Con64b] Control Data Corporation. 3400; Control Data 3400 computer system FORTRAN. Control Data Corporation, Minneapolis, MN, USA, 1964. various pp.

CDC:1964:CDCa

[Con64c] Control Data Corporation. 3600:

Control Data 3600 computer system FORTRAN/ reference manual: preliminary. Control Data Corp., Palo Alto, CA, USA, 1964. 186 pp.

CDC:1964:FRM

[Con64d] Control Data Corporation. FOR-TRAN 63/reference manual [Revision A]. Control Data Corporation, Minneapolis, MN, USA, 1964. various pp.

CDC:1965:CSF

[Con65] Control Data Corporation. 3400/ 3600: computer systems FOR-TRAN reference manual. The Corporation, Palo Alto, CA, USA, 1965. vii + [192] pp.

CDC:1966:CDSa

[Con66a] Control Data Corporation. Control Data 6000 series computer systems: Chippewa Laboratories FORTRAN compiler run. Control Data Corp., Minneapolis, MN, USA, prelim. edition, 1966. various pp.

CDC:1966:CDSb

[Con66b] Control Data Corporation. Control Data 6000 series computer systems; Chippewa operating system FORTRAN reference manual. Control Data Corporation, Palo Alto, CA, USA, 1966. various pp.

CDC:1967:CSFa

[Con67a] Control Data Corporation. 3100, 3200, 3300, 3500 computer systems FORTRAN training. Control Data Corporation, Minneapolis, MN, USA, 1967. v + [285] pp.

CDC:1967:CSFb

[Con67b] Control Data Corporation. 3400/ 3600/3800 computer systems: FORTRAN reference manual. Control Data Corporation, St. Paul, MN, USA, revised edition, 1967. various pp.

CDC:1967:CSFc

[Con67c] Control Data Corporation. 6400/ 6500/6600 computer systems; FORTRAN reference manual. Control Data Corp., Documentation Dept., Palo Alto, CA, USA, revised May 1967 edition, 1967. various pp.

CDC:1968:CSF

[Con68a] Control Data Corporation. Computer systems FORTRAN; reference manual: 3100/3200/3300/ 3500. Control Data Corporation, St. Paul, MN, USA, new edition, 1968. ix + [234] pp.

CDC:1968:CDC

[Con68b] Control Data Corporation. Control Data 3100-3200-3300-3500 computer systems FORTRAN reference manual. Control Data Corporation, Minneapolis, MN, USA, 1968. various pp.

CDC:1968:FPL

[Con68c] Control Data Corporation. FOR-TRAN programming language for MAC systems: reference guide. Control Data Corporation, St. Paul, MN, USA, preliminary edition, 1968. various pp.

CDC:1969:CSU

[Con69a] Control Data Corporation. 3300/ 3500 computer systems USASI FORTRAN/MASTER reference manual. Control Data Corp., St. Paul, MN, USA, 1969. ix + [171] pp.

CDC:1969:CSF

[Con69b] Control Data Corporation. 6400/ 6500/6600 computer systems; FORTRAN reference manual. Control Data Corporation, Palo Alto, CA, USA, 1969. various pp.

CDC:1969:CFR

[Con69c] Control Data Corporation. CDC Fortran reference manual. CDC publication 60174900A, State University of New York, Computing Center Press, Buffalo, NY, USA, 1969. various pp.

CDC:1969:CDC

[Con69d] Control Data Corporation. Control Data 6400/6500/6600 computer systems: FORTRAN reference manual. Control Data Corporation, Palo Alto, CA, USA, 1969. viii + [181] pp.

CDC:1970:UFM

[Con70] Control Data Corporation. USASI FORTRAN MASTER/MSOS reference manual: 3100, 3200, 3300, 3500 computer systems. Control Data Corporation, St. Paul, MN, USA, 1970. various pp.

Constam:1971:FA

[Con71a] M. Constam. FORTRAN für Anfanger, volume 48 of Lecture notes in operations research and mathematical systems. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 3-540-05471-5 (print), 3-642-96076-6. 145 pp. LCCN QA75.5-76.95. URL https:

//link.springer.com/book/10.
1007/978-3-642-96076-5).

CDC:1971:CDCb

[Con71b] Control Data Corporation. Control Data 3170/3300/3500 computer systems: ANSI FORTRAN, MASTER/MSOS reference manual. Control Data Corporation, St. Paul, MN, USA, 1971. various pp.

CDC:1971:CDS

[Con71c] Control Data Corporation. Control Data 6000 series computer systems: KRONOS 2.O FOR-TRAN reference manual. Control Data Corporation, Minneapolis, MN, USA, revised edition, 1971. various pp.

CDC:1971:CDCa

[Con71d] Control Data Corporation. Control Data Cyber 70 computer systems models 72, 73, 74, 76, 7600 computer systems, 6000 computer systems: SIFT (Fortran translator program) programming systems bulletin. Software Documentation, CDC, Sunnyvale, CA, USA, 1971. 20 pp.

CDC:1971:FER

[Con71e] Control Data Corporation. FOR-TRAN extended reference manual. CDC, Sunnyvale, CA, USA, 1971. ca. 200 pp.

CDC:1971:FEV

[Con71f] Control Data Corporation. FOR-TRAN extended version 4 reference manual. Control Data Corporation, Sunnyvale, CA, USA, 1971. various pp.

CDC:1972:CDC

[Con72a] Control Data Corporation. Control Data Cyber 70 computer systems, models 72, 73, 74, 6000 computer systems; FORTRAN reference manual, models 72, 73, 74, version 2.3, 6000 version 2.3. The Corp., Sunnyvale, CA, USA, 1972. various pp.

CDC:1972:FED

[Con72b] Control Data Corporation. Fortran extended debug user’s guide: models 72, 73, 74 version 4, model 76 version 2, 7600 versions 1 and 2, 6000 versions 3 and 4. The Corp., Sunnyvale, CA, USA, 1972. ???? pp.

Constam:1973:FA

[Con73a] M. Constam. FORTRAN für Anfanger. Lecture notes in economics and mathematical systems,

48. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2., verb. Aufl. edition, 1973. vi + 148 pp.

CDC:1973:CSA

[Con73b] Control Data Corporation. 3170/ 3300/3500 computer systems: ANSI FORTRAN MASTER/ MSOS reference manual. Control Data Corp., St. Paul, MN, USA, revised edition, 1973. 197 pp.

CDC:1973:CDCa

[Con73c] Control Data Corporation. Control Data 3100/3150/3170/3200/ 3300/3500 computer systems: MASTER/MSOS ANSI FOR-TRAN versions 1 and 2 reference manual. Control Data Corp., Arden Hills, MN, USA, 1973. various pp.

CDC:1973:CDCb

[Con73d] Control Data Corporation. Control Data Cyber 70 computer systems models 72, 73, 74 6000 computer systems: record manager guide for users of Fortran extended. Software Documentation, CDC, Sunnyvale, CA, USA, 1973. various pp.

CDC:1973:CDCc

[Con73e] Control Data Corporation. Control Data Cyber 70 computer systems models 72, 73, 74, 76, 7600 computer systems, 6000 computer systems: FORTRAN extended reference manual, models 72, 73, 74 version 4, model 76 version 2, 7600 version 2, 6000 version 4. Software Documentation, CDC, Sunnyvale, CA, USA, 1973. various pp.

CDC:1973:KVTa

[Con73f] Control Data Corporation. KRO-NOS version 2 time-sharing FOR-TRAN reference manual. The Corp., Arden Hills, MN, USA, revised edition, 1973. viii + [125] pp.

CDC:1973:KVTb

[Con73g] Control Data Corporation. KRO-NOS version 2 time-sharing FOR-TRAN user’s reference manual. The Corp., St. Paul, MN, USA, 1973. ???? pp.

CDC:1974:CKT

[Con74] Control Data Corporation. Cybernet, Kronos 2.1 time-sharing Fortran programming language: reference manual. Control Data Corporation, Minneapolis, MN, USA, 1974. various pp.

CDC:1975:CDC

[Con75a] Control Data Corporation. Control Data 1700 computer systems: 1700 MSOS 4 MS FORTRAN version 3A/B reference manual. Control Data Corp., La Jolla, CA, USA, 1975. various pp.

CDC:1975:FCLa

[Con75b] Control Data Corporation. FOR-TRAN common library mathematical routines. The Corporation, Sunnyvale, CA, USA, 1975. vi + 155 pp.

CDC:1975:FCLb

[Con75c] Control Data Corporation. FOR-TRAN common library mathematical routines. CDC, Sunnyvale, CA, USA, 1975. ???? pp.

CDC:1975:FEVa

[Con75d] Control Data Corporation. Fortran extended version 4 instant manual. Control Data Corporation, Minneapolis, MN, USA, revised edition, 1975. 43 pp.

CDC:1975:FEVb

[Con75e] Control Data Corporation. FOR-TRAN extended version 4 reference manual. CDC, Sunnyvale, CA, USA, 1975. ???? pp.

CDC:1975:FEVc

[Con75f] Control Data Corporation. FOR-TRAN extended version 4 reference manual. The Corp., Sunnyvale, CA, USA, 1975. ???? pp.

CDC:1976:CRM

[Con76a] Control Data Corporation. CY-BER record manager version 1 guide for users of FORTRAN extended version 4. CDC, Minneapolis, MN, USA, 1976. ???? pp.

CDC:1976:FEV

[Con76b] Control Data Corporation. FOR-TRAN extended version 4 debug user’s guide. CDC, Sunnyvale, CA, USA, 1976. ???? pp.

CDC:1976:IIG

[Con76c] Control Data Corporation. IN-TERCOM interactive guide for users of FORTRAN extended: Control Data Cyber 170 series, Cyber 70 models 72, 73, 74, 6000 series computer systems. Control Data Corporation, Publications and Graphic Division, Sunnyvale, CA, USA, 1976. ???? pp.

Converse:1976:COT

[Con76d] Robert A. Converse. Compiler optimization techniques for a FORTRAN compiler. Thesis (m.s.), University of Rhode Island, Kingston, RI, USA, 1976. v + 72 pp.

CDSL:1977:CFD

[Con77a] Conference on Data Systems Languages.Fortran Data Base and Manipulation Language Committee. CODASYL FORTRAN data base facility: journal of development, January 1, 1977. Technical Report ????, Department of Supply and Services, Ottawa, Ontario, Canada, 1977. 135 pp.

CDC:1977:FEV

[Con77b] Control Data Corporation. Fortran extended version 4 user’s guide. The Corporation, Sunnyvale, CA, USA, 1977. ???? pp.

Converse:1977:OTN

[Con77c] Robert A. Converse. Optimization techniques for the NUSC FOR-TRAN cross-compiler. Technical Report ????, Naval Underwater Systems Center, Newport, RI, USA, 1977. 26 pp.

CDC:1978:FDB

[Con78a] Control Data Corporation. FOR-TRAN data base facility version 1: reference manual. CDC, Sunnyvale, CA, USA, 1978. ???? pp.

CDC:1978:FEV

[Con78b] Control Data Corporation. FOR-TRAN extended version 4 reference manual. The Corp., Sunnyvale, CA, USA, revised edition, 1978. ???? pp.

CDC:1979:CID

[Con79a] Control Data Corporation. CY-BER interactive debug version 1 guide for users of FORTRAN extended version 4. CDC, Sunnyvale, CA, USA, 1979. ???? pp.

CDC:1979:FEVa

[Con79b] Control Data Corporation. FOR-TRAN extended version 4 reference manual. CDC, Sunnyvale, CA, USA, 1979. ???? pp.

CDC:1979:FEVb

[Con79c] Control Data Corporation. FOR-TRAN extended version 4 to FOR-TRAN version 5 conversion aid program version 1 reference manual. CDC, Sunnyvale, CA, USA, 1979. ???? pp.

CDC:1979:FVC

[Con79d] Control Data Corporation. FOR-TRAN version 5 common library mathematical routines reference manual. Control Data Corporation, Sunnyvale, CA, USA, revision B edition, 1979. various pp.

Conway:1979:PPG

[Con79e] Richard Conway. Programming for Poets: a Gentle Introduction Using Fortran (Watf Iv). Scott, Foresman and Company, Glenview, IL, USA, June 1979. ISBN 0-316-15421-0. ???? pp. LCCN ???? US$19.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0316154210.

CDSL:1980:CFD

[Con80a] Conference on Data Systems Languages.Fortran Data Base and Manipulation Language Committee. CODASYL FORTRAN data base facility: journal of development, 1980. Technical Report ????, Secretariat of the Canadian Government EDP Standards Committee on behalf of the CODASYL Organisation, Hull, Québec, 1980. 104 pp.

CDC:1980:FEV

[Con80b] Control Data Corporation. FOR-TRAN extended version 4 reference manual. CDC, Sunnyvale, CA, USA, 1980. ???? pp.

CDC:1980:FVR

[Con80c] Control Data Corporation. FOR-TRAN version 5 reference manual. Control Data Corporation, Sunnyvale, CA, USA, revised edition, 1980. 287 pp.

Cook:1972:CPF

[Coo72] D. B. Cook. Correspondence:

‘Packing’ in FORTRAN. The Computer Journal, 15(4):361, November 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/
361.tif.

Cook:1976:EEP

[Coo76a] A. J. Cook. Experience with extensible, portable FORTRAN extensions. ACM SIGPLAN Notices, 11(9):10–17, September 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1976:SPI

[Coo76b] Curtis R. Cook. A self-paced introductory Fortran programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):78– 79, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Corbato:1960:DFP

[Cor60] F. J. Corbató. Description of a FORTRAN post-mortem subprogram, (F2PM). Memorandum CC-151, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 5 pp.

Corbato:1961:ADF

[Cor61] F. J. Corbató. An abbreviated description of the FORTRAN compiler language. Memorandum CC-164-4, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1961. 8 pp.

Corlett:1973:PP

[Cor73] Peter Norman Corlett. Practical Programming. Cambridge University Press, Cambridge, UK, February 1973. ISBN 0-521-08198-X. ???? pp. LCCN ???? US$27.95.

Corkill:1977:FFU

[Cor77] Daniel D. Corkill. FUTIL: FOR-TRAN utilities for APL files. Technical report, APL Group, University Computing Center, University of Massachusetts, Amherst, MA 01003, USA, July 1977.

CDC:1979:CCF

[Cor79] Control Data Corporation. CDC Cyber 200 Fortran language 1.4. Document 60457040, Control Data Corporation, Minneapolis, MN, USA, 1979.

Court:1970:FB

[Cou70] R. A. Court. FORTRAN for beginners. Computer studies series. Holmes McDougall, Edinburgh, Scotland, 1970. ISBN 0-7157-0824-4 (paperback). 128 pp.

Couger:1975:IAL

[Cou75] J. Daniel Couger. Introduccion al lenguaje FORTRAN. Serie de Ensenanza programada “El Ateneo”. Libreria “El Ateneo” Editorial, Buenos Aires, Argentina, 1975. xiii + 161 pp.

Courchesne:1976:NFI

[Cou76] Andre W. Courchesne. Notes on Fortran IV. Technical report, Suffolk University, Boston, MA, USA, 1976. 28 pp.

Chattergy:1980:TMP

[CP80] R. Chattergy and Udo W. Pooch. Top-down, modular programming in FORTRAN with WATFIV. Winthrop Publishers, Cambridge, MA, USA, 1980. ISBN 0-87626-879-3. xvi + 217 pp. LCCN QA76.73.F25 .C43.

Cardenas:1972:CS

[CPM72] Alfonso F. Cardenas, Leon Presser, and Miguel A. Marin. Computer Science. John Wiley and Sons, New York, London, Sydney, March 1972. ISBN 0-471-13468-6. xii + 522 pp. LCCN QA76.5 .C365. US$133.50.

Chung-Phillips:1975:NDD

[CPR75] Alice Chung-Phillips and R. W. Rosen. Note on dynamic data storage in Fortran IV. The Computer Journal, 18(4):342–343, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Charet:1969:FIA

[CR69] Gerard Charet and Alain Riche. Fortran; initiation au langage de l’informatique scientifique [par]. Societé d’édition d’enseignement supérieur, Paris, France, 1969. xv + 239 pp.

Curtis:1971:FSS

[CR71] A. R. Curtis and John Ker Reid. Fortran subroutines for the solutions of sparse sets of linear equations. Report / United Kingdom Atomic Energy Authority Research Group; AERE-R 6844 Report (Atomic Energy Research Establishment (Harwell, Oxfordshire)); AERE-R 6844. Theoretical Physics Division, Atomic Energy Research Establishment, Harwell, Berkshire, UK, 1971. ISBN 0-7058-0460-7. 38 pp.

Charet:1973:FIA

[CR73] Gerard Charet and Alain Riche.

Fortran; initiation au langage de l’informatique scientifique [par]. Societé d’enseignement supérieur, Paris, France, 1973. xiii + 271 pp.

Christiansen:1974:OSC

[CR74] J. P. Christiansen and K. V.

Roberts. OLYMPUS a standard control and utility package for initial-value FORTRAN programs. Computer Physics Communications, 7(5):245–270, May 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465574900265.

Cramton:1968:FFE

[Cra68] Dale A. Cramton. 3600 FOR-TRAN to 6500 FORTRAN, extended conversion. Technical Bulletin 10, Computer Laboratory, Michigan State University, East Lansing, MI, USA, 1968. 9 pp.

Cranor:1975:FCP

[Cra75] Ralph Sherman Cranor. A Fortran computer program to aid in a school district budget preparation with respect to salary costs. Thesis (m.a.), Western State College of Colorado, Gunnison, CO, USA, August 1975. vi + 55 pp.

Craddock:1976:DFB

[Cra76] J. M. Craddock. Designing a FOR-TRAN based system for meteorological statistics. Software—Practice and Experience, 6(3):301–319, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Crary:1979:VPN

[Cra79] F. D. Crary. A versatile precompiler for nonstandard arithmetics. ACM Transactions on Mathematical Software, 5(2):204– 217, 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cray:1980:CFR

[Cra80] Cray Research, Inc., Minneapolis, MN. CRAY-1 Fortran Refernce Manual, 1980.

Criss:1977:FPX

[Cri77] J. W. Criss. A Fortran program for X-ray fluorescence analysis. Technical report, Naval Research Laboratory, Washington, DC, USA, 1977. various pp.

Colman:1961:API

[CS61] Harry L. Colman and Clarence Smallwood. An auto-instructional program for the IBM Fortran programming system. Technical report, Western Data Processing Center, University of California, Los Angeles, CA, USA, 1961. various pp.

Colman:1962:CLA

[CS62] Harry L. Colman and Clarence Smallwood. Computer language; an autoinstructional introduction to Fortran. McGraw-Hill, New York, NY, USA, 1962. 196 pp.

Couger:1968:FIP

[CS68] J. Daniel Couger and Loren E.

Shannon. FORTRAN IV; a programmed instruction approach. Irwin, Homewood, IL, USA, 1968. x + 244 pp.

Campbell:1971:FCPa

[CS71a] Graham M. Campbell and Wilson E. Singletary. A first course in programming; FORTRAN IV with WATFIV. Technical report, Pennsylvania State University, Dept. of Computer Science, University Park, PA, USA, 1971. 200 pp.

Campbell:1971:FCPb

[CS71b] Graham M. Campbell and Wilson E. Singletary. A first course in programming: FORTRAN IV with WATFIV. Auerbach, Princeton, NJ, USA, 1971. ISBN 0-87769-067-7 (paperback). 137 pp.

Campbell:1971:FIWa

[CS71c] Graham M. Campbell and Wilson E. Singletary. Fortran IV with Watfiv; a first course in programming. Auerbach, Princeton, NJ, USA, 1971. ISBN 0-87769-067-7. 144 pp. LCCN QA76.5 .C35 1971.

Campbell:1971:FIWb

[CS71d] Graham M. Campbell and Wilson E. Singletary. FORTRAN IV with WATFIV: a first course in programming. Petrocelli Books, New York, NY, USA, 1971. 144 pp.

Couger:1971:PLA

[CS71e] J. Daniel Couger and Loren E. Shannon. Programmed learning aid for FORTRAN; a beginner’s approach. Irwin PLAID series. Learning Systems Company, Homewood, IL, USA, 1971. xi + 160 pp.

Couger:1972:FIP

[CS72] J. Daniel Couger and Loren E. Shannon. FORTRAN IV; a programmed instruction approach. Irwin-Dorsey information processing series. R. D. Irwin, Homewood, IL, USA, revised edition, 1972. x + 278 pp.

Cheek:1973:IMU

[CS73a] James B. Cheek and Paul K. Senter. Instruction manual for using Waterways Experiment Station Time Sharing computer program “FEDIT” for writing other FORTRAN programs. Miscellaneous paper K-73-1, Waterways Experiment Station, Vicksburg, MS, USA, 1973. vii + 9 + [3] + 1 pp.

Cooper:1973:SFP

[CS73b] Laura G. Cooper and Marilyn Z. Smith. Standard Fortran: a Problem-Solving Approach. Houghton Mifflin, Boston, MA, USA, June 1973. ISBN 0-395-14028-5. vii + 251 pp. LCCN ???? US$24.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0395140285.

Campbell:1975:FIW

[CS75] Graham M. Campbell and Wilson E. Singletary. FORTRAN IV with WATFIV: a first course in programming. Petrocelli/Charter, New York, NY, USA, revised edition, 1975. ISBN 0-88405-306-7. 160 pp. LCCN QA76.6.C34 1975.

Couger:1976:FIP

[CS76] J. Daniel Couger and Loren E.

Shannon. Fortran IV: A.P.I. Approach, Including Structured Programming. Irwin series in information and decision sciences. R. D. Irwin, Homewood, IL, USA, third edition, March 1976. ISBN 0-256-01632-1. xii + 302 pp. LCCN QA 76.73 F25 C83. US$12.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0256016321.

Costello:1977:GBS

[CS77a] Donald F. Costello and Richard J.

Schonberger. On guiding the business school toward computer literacy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):180–183, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Couger:1977:PFB

[CS77b] Daniel Couger and Loren E. Shanon. Plaid for Fortran a Beginners Approach. R. D. Irwin, Homewood, IL, USA, June 1977. ISBN 0-256-01986-X. ???? pp. LCCN ???? US$11.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        025601986X.

Couger:1977:PLA

[CS77c] J. Daniel Couger and Loren E. Shannon. Programmed learning aid for FORTRAN; a beginner’s approach. Irwin PLAID series. Learning Systems Company, Homewood, IL, USA, 1977. ISBN 0-256-01986-X. xi + 180 pp.

Csendes:1975:FPG

[Cse75] Z. J. Csendes. A FORTRAN program to generate finite difference formulas. International Journal for Numerical Methods in Engineering, 9(3):581–599, 1975. CO-DEN IJNMBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Corlett:1972:PP

[CTC72] Peter Norman Corlett, J. D. Tinsley, and R. A. Court. Practical programming. School Mathematics Project Handbooks. University Press, Cambridge [Eng.], second edition, 1972. ISBN 0-521-08198-X. x + 264 pp. LCCN QA76.6 .C66 1972.

Culik:1980:WFL

[Cul80] Karel Culik. What is a flowchart loop and about structured programming. ACM SIGPLAN Notices, 15(1):45–57, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cline:1963:POF

[CW63] J. L. Cline and D. L. Whitlow. POL outlined FORTRAN. POL Outlined, Orlando, FL, USA, 1963. ???? pp.

Carnahan:1971:IDC

[CW71] Brice Carnahan and James O.

Wilkes. Introduction to digital computing and FORTRAN IV with MTS applications. Technical report, University of Michigan, Ann Arbor, MI, USA, 1971. 344 pp.

Carnahan:1972:DCF

[CW72] Brice Carnahan and James O.

Wilkes. Digital computing, FOR-TRAN IV, WATFIV, and MTS (with *FTN and *WATFIV). Technical report, Chemical Engineering Dept., University of Michigan, Ann Arbor, MI, USA, 1972. vi + 538 pp.

Carnahan:1973:DCN

[CW73a] Brice Carnahan and James O.

Wilkes. Digital computing and numerical methods: with FORTRAN-IV, WATFOR and WATFIV programming. John Wiley and Sons, New York, London, Sydney, January 1973. ISBN 0-471-13500-3. xi + 477 pp. LCCN TA345 .C38 1973. US$45.95. URL http:

//www.cbooks.com/sqlnut/SP/

                                        search/gtsumt?source=&isbn=
                                        0471135003.

Carnahan:1973:DCF

[CW73b] Brice Carnahan and James O. Wilkes. Digital computing, FOR-TRAN IV, WATFIV, and MTS (with *FTN and *WATFIV). Technical report, Chemical Engineering Dept., University of Michigan, Ann Arbor, MI, USA, 1973. 505 pp.

Charmonman:1975:SPF

[CW75] S. Charmonman and J. L. Wagener. On structured programming in FORTRAN. ACM SIGNUM Newsletter, 10(1):21–23, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Carnahan:1976:DCF

[CW76] Brice Carnahan and James O. Wilkes. Digital computing, FOR-TRAN IV, WATFIV, and MTS (with *FTN and *WATFIV). Technical report, Chemical Engineering Dept., University of Michigan, Ann Arbor, MI, USA, 1976. vi + 538 pp.

Carnahan:1977:DCF

[CW77] Brice Carnahan and James O. Wilkes. Digital computing, FOR-TRAN IV, WATFIV, and MTS (with *FTN and *WATFIV). Technical report, Chemical Engineering Dept., University of Michigan, Ann Arbor, MI, USA, 1977. ???? pp.

Carnahan:1978:DCF

[CW78a] Brice Carnahan and James O.

Wilkes. Digital computing, FOR-TRAN IV, WATFIV, and MTS (with *FTN and *WATFIV). Technical report, Chemical Engineering Dept., University of Michigan, Ann Arbor, MI, USA, 1978. ???? pp.

Chrisman:1978:PTG

[CW78b] Nicholas Chrisman and Denis Naldrett White. Programming for transportability: a guide to machine independent FORTRAN. Technical report, Laboratory for Computer Graphics and Spatial Analysis, Graduate School of Design, Harvard University, Cambridge, MA, USA, 1978. i + 12 pp.

Carnahan:1979:DCF

[CW79] Brice Carnahan and James O.

Wilkes. Digital computing, FOR-TRAN IV, WATFIV, and MTS (with *FTN and *WATFIV). Technical report, Chemical Engineering Dept., University of Michigan, Ann Arbor, MI, USA, 1979. ???? pp.

Cohen:1972:ESF

[CZ72] Jacques Cohen and Carl Zuckerman. Evalquote in simple FOR-TRAN: a tutorial on interpreting LISP. BIT (Nordisk tidskrift for informationsbehandling), 12(3):299–317, 1972. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Dubner:1968:NIL

[DA68] H. Dubner and J. Abate. Numerical inversion of Laplace transforms by relating them to the finite Fourier cosine transform. Journal of the ACM, 15(1):115– 123, January 1968. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dahlberg:1966:LFA

[Dah66] Michael L. Dahlberg. Lengthfrequency analysis, IBM 7090/ 7094, Fortran II or Fortran IV. Transactions of the American Fisheries Society, 95(3):331, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dahlberg:1978:CPA

[Dah78] Michael L. Dahlberg. Computer program for analysis of the homogeneity and goodness of fit of frequency distributions, FOR-TRAN IV. Fishery Bulletin, 76 (1):??, 1978. CODEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic). URL https:

//spo.nmfs.noaa.gov/content/
computer-program-analysis-homogeneity-
and-goodness-fit-frequency-
distributions-fortran-iv;
https://spo.nmfs.noaa.gov/sites/
default/files/pdf-content/fish-
bull/dahlberg.pdf.

Darby:1978:EPS

[Dar78] Mary Virginia Darby. An enhanced processing system for structured Fortran: a thesis. Thesis (m.s.), University of Alabama in Huntsville, Huntsville, AL, USA, 1978. vii + 97 l pp.

Dasenbrock:1974:ERF

[Das74] Robert R. Dasenbrock. An editing routine for FORTRAN programs. Technical Report AD 779 841, Naval Research Library, Washington, DC, USA, 1974. 48 pp.

Data:1966:USF

[Dat66] Control Data. USA Standard Fortran. Technical Report USAS X3.9-1966, Control Data Corp., New York, NY, USA, March 1966.

DGC:1967:FIR

[Dat67a] Data General Corporation. FOR-TRAN II: reference manual. Data General Corporation, Southboro, MA, USA, 1967. vii + 106 pp.

DNC:1967:CFI

[Dat67b] Data Network Corporation. Conversational Fortran IV reference manual. Data Network, ????, 1967. vi + 109 pp.

Datacraft:1973:SSA

[Dat73] Datacraft Corporation. Series 6000 scientific arithmetic unit (SAU): FORTRAN support library, general specification: revision A, September 1973. Technical Report AA61641-00A, Datacraft Corp., Fort Lauderdale, FL, USA, 1973. 5 + 86 [i.e. 89] + 5 pp.

DGC:1975:UMF

[Dat75] Data General Corporation. User’s manual, FORTRAN IV. Data General Corporation, Southboro, MA, USA, eighth edition, 1975. ca. 250 pp.

DGC:1977:ANF

[Dat77a] Data General Corporation. Addendum to NOVA-LINE FORTRAN IV: user’s manual. Data General Corporation, Southboro, MA, USA, 1977. 94 pp.

DGC:1977:FIU

[Dat77b] Data General Corporation. FOR-TRAN IV: user’s manual. Data General Corporation, Southboro, MA, USA, 1977. various pp.

Davis:1970:CST

[Dav70] Mitchell Davis. A comparative survey of two fast Fortran compilers. Thesis (m.s.), University of North Carolina at Chapel Hill, Chapel Hill, NC, USA, 1970. 64 pp.

Davis:1972:CST

[Dav72a] Mitchell Davis, Jr. A comparative survey of two fast Fortran compilers. Thesis (m.s.), University of North Carolina at Chapel Hill, Chapel Hill, NC, USA, 1972. iv + 64. pp.

Davisson:1972:LFS

[Dav72b] Charlotte M. Davisson. LPSUB, a FORTRAN subroutine for solving any standard linear programming problem of a size compatible with the computer being used. NRL memorandum report 2383; NRL computer bulletin 27, Naval Research Laboratory, Washington, DC, USA, 1972. 64 pp.

David:1974:IHF

[Dav74] Daniel Jean David. Informatique: hardware, FORTRAN, AP.L., cartes-controle. Collection scientifique contemporaine. Editions techniques et scientifiques françaises: Agence parisienne de distribution, Paris, France, 1974. 331 pp.

Dayhoff:1963:CMP

[Day63] M. O. Dayhoff. A contour-map program for X-ray crystallography. Comm. ACM, 6(10):620–622, October 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Day:1970:USS

[Day70] A. C. Day. The use of symbol-state tables. The Computer Journal, 13 (4):332–339, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_04/130332.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/332.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
04/tiff/333.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/
334.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_04/tiff/335.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_04/tiff/336.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_

                                        04/tiff/337.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_13/Issue_04/tiff/
                                        338.tif;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_13/Issue_04/tiff/339.

tif. See note [?].

Day:1972:FT

[Day72a] A. C. Day. Fortran Techniques. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08549-7. viii + 96 pp. LCCN QA76.73.F25 D39.

Day:1972:FTS

[Day72b] A. Colin Day. Fortran Techniques With Special Reference to Non-Numerical Applications. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08549-7, 0-521-09719-3 (paperback). viii + 96 pp. LCCN QA76.73.F25 D39. US$16.95; US$19.95. URL http://www.

                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=0521097193.

Day:1972:LFC

[Day72c] A. Colin Day. A London Fortran course: students’ handbook for the television course. Athlone Press of the University of London, London, UK, 1972. ISBN 0-485-12018-6. v + 85 pp.

Day:1976:ABB

[Day76] A. C. Day. Algorithm 91: Balancing a binary tree. The Computer Journal, 19(4):360–361, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/
360.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/361.
tif.

Day:1978:CF

[Day78] A. Colin Day. Compatible Fortran. Cambridge University Press, Cambridge, UK, June 1978. ISBN 0-521-22027-0. vii + 107 pp. LCCN QA76.73 F25 D38 1978. US$22.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0521220270.

Day:1979:FTS

[Day79] A. Colin Day. Fortran techniques with special reference to non-numerical applications. Cambridge University Press, Cambridge, UK, 1979. viii + 96 pp.

Dreyfus:1968:FI

[DB68] Michel Dreyfus and H. Boucher.

Fortran IV. Centre interarmées de recherche operationnelle. Dunod, Paris, France, 2e edition, 1968. xii + 178 pp.

Dreyfus:1969:FI

[DB69] M. Dreyfus and H. Boucher. Fortran IV. Centre interarmées de recherche operationnelle; 1. Dunod, Paris, France, 3e edition, 1969. xii + 184 pp.

Dreyfus:1970:FI

[DB70] M. Dreyfus and H. Boucher. Fortran IV. Centre interarmées de recherche operationnelle. 1. Dunod, Paris, France, 4eme edition, 1970. xii + 184 pp.

Dinneen:1973:SAA

[DB73] L. C. Dinneen and B. C. Blakesley. Statistical algorithms: Algorithm AS 62: a generator for the sampling distribution of the Mann– Whitney U statistic. Applied Statistics, 22(2):269–273, June 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/62.

Carvalho:1973:CPM

[dC73] Gaspar Soares de Carvalho. Comparacão de produtos de meteorizacão pela sua composicão mineralogica virtual e por indices de meteorizacão e programas em Fortran IV para os calcular. Trabalhos do Instituto de Investigacão Cientifica de Moçambique 27/28, Instituto de Investigacão Cientifica de Moçambique, Lourenço Marques, Mozambique, 1973. 25–45 pp.

Day:1976:DCP

[DCHR76a] A. C. Day, P. A. Clarke, D. Hill, and J. K. Reid. Discussion and correspondence: the proposed new standard for Fortran: a critical examination. The Computer Journal, 19(3):268–271, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
19/3/268.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/

268.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
19/Issue_03/tiff/270.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/271.tif.

Day:1976:PNS

[DCHR76b] A. C. Day, P. A. Clarke, D. Hill, and J. K. Reid. The proposed new standard for FORTRAN: a critical examination. The Computer Journal, 19(3):268–271, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/
268.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
19/Issue_03/tiff/270.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/271.tif.

Dietmeyer:1968:GPI

[DD68] D. L. Dietmeyer and J. R. Duley. Generating prime implicants via ternary encoding and decimal arithmetic. Comm. ACM, 11(7):520–523, July 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DalBono:1975:SFS

[DDM^{+}75] Ivan Dal Bono, Mauro Diligenti, Concetta Mosca, Antonio Ricci, and Antonio Villani. Simple Fortran support for computer-assisted instruction. Information Processing Letters, 3(3):88–90, January 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeTar:1972:PFP

[De 72] De Los Fletcher De Tar. Principles of Fortran programming. W. A. Benjamin, Menlo Park, CA, USA, 1972. ix + 136 pp.

Dean:1971:OTF

[Dea71] John L. Dean. Optimization techniques for FORTRAN IV (G and H) programs written for the IBM 360 under OS. Technical report, Goddard Space Flight Center, Greenbelt, MD, USA, 1971. 42 pp.

Dean:1977:OTF

[Dea77] John L. Dean. Optimization techniques for FORTRAN IV (G and H) programs written for the IBM 360 under OS. NASA technical memorandum X-70477, U.S. Goddard Space Flight Center, Greenbelt, MD, USA, 1977. iii + 42 pp. Reproduced by National Technical Information Service.

Deever:1974:BFa

[Dee74a] David Deever. From BASIC to FORTRAN, 1, 1974.

Deever:1974:BFb

[Dee74b] David Deever. From BASIC to FORTRAN, 2, 1974.

Deever:1974:BFc

[Dee74c] David Deever. From BASIC to FORTRAN, 3, 1974.

Deever:1974:BFd

[Dee74d] David Deever. From BASIC to FORTRAN, 4, 1974.

Demirmen:1969:MPF

[Dem69] Ferruh Demirmen. Multivariate procedures and Fortran IV program for evaluation and improvement of classifications. Computer contribution 31, State Geological Survey, Lawrence, KS, USA, 1969. 51 pp.

Dennis:1971:PFI

[Den71] David M. Dennis. Programming in FORTRAN IV (IBM 1130 modified). Technical report, Western New Mexico University, Silver City, NM, USA, 1971. 12 pp.

Daly:1979:PRT

[DEN79] Cecilia Daly, David W. Embley, and George Nagy. A progress report on teaching programming to business students without lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1): 247–250, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Dence:1980:FC

[Den80] Thomas P. Dence. The Fortran cookbook. Tab Books, Blue Ridge Summit, PA, USA, 1980. ISBN 0-8306-9914-7, 0-8306-1187-8 (paperback). 334 pp. LCCN QA76.73.F25 D45. US$14.95 (hardcover), US$8.95 (paperback).

Der:1964:NMF

[Der64] J. Der. Numerical methods and FORTRAN program for flow field calculations in supersonic cascades. Arl; 64-146, Aerospace Research Laboratories, Wright-Patterson Air Force Base, OH, USA, 1964. vi + 221 pp.

Deutsch:1973:PUF

[Deu73] Edwin Deutsch. Das Programmsystem UNIRUN in FORTRAN VI. Forschungsbericht 76, Institut fur Höhere Studien und Wissenschaftliche Forschung, Wien, Austria, 1973. 58 pp.

Dewar:1972:CAR

[Dew72] R. B. K. Dewar. Correspondence: On “Another round of FORTRAN”. The Computer Journal, 15(1):7, February 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
15/Issue_01/150007.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/7.tif. See

[Cha71b, Fla72, Cha72].

Dey:1976:RMA

[Dey76] Abhijit Dey. Resistivity modelling for arbitrarily shaped two dimensional structures part II: User’s guide to the FORTRAN algorithm RESIS2D. Technical Report LBL 5283, Energy and Environment Division, Lawrence Berkeley Laboratory University of California, Berkeley, CA, USA, 1976. 56 pp.

DiManzo:1979:LOP

[DFO79] M. Di Manzo, A. L. Frisiani, and G. Olimpo. Loop optimisation for parallel processing. The Computer Journal, 22(3):234–239, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Dorn:1967:MCF

[DG67] William S. Dorn and Herbert J.

Greenberg. Mathematics and Computing: With Fortran Programming. John Wiley and Sons, New York, London, Sydney, annotated instructor’s edition, January 1967. ISBN 0-471-21915-0. xvi + 595 (or xix + 861??) pp. LCCN ???? US$24.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471219150.

Dorn:1968:MCF

[DG68] William S. Dorn and Herbert J.

Greenberg. Mathematics and computing: with FORTRAN programming. John Wiley and Sons, New York, London, Sydney, corr. print. edition, 1968. xvi + 595 pp.

Dorn:1970:MCC

[DG70] William S. Dorn and Herbert J.

Greenberg. Matematicas y computacion, con programacion FOR-TRAN. Editorial Limusa, Mexico, DF, Mexico, 1970. 583 pp.

Dreyfus:1975:PDF

[DG75] Michel Dreyfus and Claude Gangloff. La pratique du FORTRAN: exercices commentes. Dunod technique Serie violette, Informatique. Dunod, Paris, France, 1975. ISBN 2-04-009946-8. x + 198 pp.

Dreyfus:1978:PPF

[DG78] Michel Dreyfus and Claude Gangloff. Praktika programmirovaniia na FORTRANe: uprazhneniia s kommentariiami. Mir, Moscow, Russia, 1978. 224 pp.

Davis:1978:FSD

[DH78] Gordon Bitter Davis and Thomas Russell Hoffmann. FORTRAN: a structured, disciplined style: based on 1977 American National Standard FORTRAN and compatible with WATFOR, WATFIV, WATFIV-S, and MNF FOR-TRAN compilers. McGraw-Hill, New York, NY, USA, July 1978. ISBN 0-07-015901-7. vii + 368 pp. LCCN QA76.73.F25 D385. US$25.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0070159017.

Dongarra:1979:ULF

[DH79] J. J. Dongarra and A. R. Hinds. Unrolling loops in Fortran. Software—Practice and Experience, 9 (3):219–226, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dickson:1968:CCI

[Dic68] Thomas R. Dickson. The Computer and Chemistry: An Introduction to Programming and Numerical Methods. W. H. Freeman, New York, NY, USA, July 1968. ISBN 0-7167-0141-3. ???? pp. LCCN ????

Dickey:1974:ICC

[Dic74a] Larry W. Dickey. Introduction to Computer Concepts: Hardware and Software. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1974. ISBN 0-13-480004-4. x + 342 pp. LCCN QA76.5 .D461.

Dicus:1974:FPG

[Dic74b] John H. Dicus. FORTRAN program to generate engine inlet flow contour maps and distortion parameters. NASA technical memorandum NASA TM X-2967, National Aeronautics and Space Administration, Washington, DC, USA, 1974. 61 pp. For sale by the National Technical Information Service.

Didday:1978:FBP

[Did78] Richard L. Didday. Fortran for Business People. West Publishing Company, St. Paul, MN, USA, June 1978. ISBN 0-8299-0101-9. ix + 342 pp. LCCN HF 5548.5 F2 F78. US$29.95; US$10.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0829901019.

Dienes:1968:EMC

[Die68] J. K. Dienes. An Eulerian method for calculating strength dependent deformation. Technical Report ????, Gulf General Atomic, Incorporated, ????, 1968. 172 pp.

Diestelmann:1972:BFI

[Die72] M. L. Diestelmann. BEOS — ein FORTRAN IV-Programm zur Berechnung der Beullasten exzentrisch orthotroper Sadwichschalen; Programmbeschreibung und Benutzungsanleitung. Deutsche Luftund Raumfahrt. Mitteilung 72-08, Deutsche Forschungs-und Versuchsanstalt fur Luft-und Raumfahrt, Porz-Wahn, 1972. 146 pp.

Diegel:1974:FP

[Die74a] Adolf Diegel. Fortran, programmation. Éditions Science moderne, Chicoutimi, Québec, 1974. 119 pp.

Diegel:1974:PF

[Die74b] Adolf Diegel. Programmation Fortran. Éditions Science moderne, Chicoutimi, Québec, 1974. 146 pp.

Diegel:1976:PFS

[Die76] Adolf Diegel. Programmatheque Fortran: statistique descriptive, correlation multiple. Editions IOMA, Ste-Foy, Québec, 1976. ISBN 0-88562-006-2. 263 pp.

Diegel:1977:SCC

[Die77] Adolf Diegel. Sondage: compilation et correlation. Programmatheque FORTRAN. Presses de l’Université du Québec, Montréal, PQ, Canada, 1977. ISBN 0-7770-0211-6. xii + 162 pp.

Diffely:1972:FPS

[Dif72] Robert Elwin Diffely. A FOR-TRAN program for the simultaneous solution of two algebraic polynomial equations in two unknowns.

Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1972.

[6] + 59 pp.

DEC:1964:PFP

[Dig64] Digital Equipment Corporation.

PDP-8 FORTRAN programming manual. Digital Equipment Corporation, Maynard, MA, USA, 1964. vi + 54 pp.

DEC:1968:PFI

[Dig68] Digital Equipment Corporation.

PDP-10 FORTRAN IV programming manual. Digital Equipment Corporation, Maynard, MA, USA, 1968. viii + 77 pp.

Digitek:1969:DUF

[Dig69] Digitek Corporation. Digitek USA FORTRAN. Digitek Corporation, Marina del Rey, CA, USA, 1969. 31 pp.

DEC:1970:PFI

[Dig70] Digital Equipment Corporation.

PDP-10 FORTRAN IV programmer’s reference manual. Digital Equipment Corporation, Maynard, MA, USA, 1970. various pp.

DEC:1971:PFIa

[Dig71a] Digital Equipment Corporation.

PDP-11 FORTRAN IV compiler and object time system: programmer’smanual. Digital Equipment Corp., Maynard, MA, USA, 1971. various pp.

DEC:1971:PFIb

[Dig71b] Digital Equipment Corporation.

PDP-15 FORTRAN IV language programmer’s reference manual.

Digital Equipment Corporation, Maynard, MA, USA, 1971. various pp.

DEC:1971:PFIc

[Dig71c] Digital Equipment Corporation. PDP-15 FORTRAN IV operating environment. Digital Equipment Corporation, Maynard, MA, USA, 1971. various pp.

DEC:1972:DML

[Dig72a] Digital Equipment Corporation. DECsystem-10 mathematical languages handbook: FORTRAN, BASIC, ALGOL. Digital Equipment Corporation, Maynard, MA, USA, second edition, 1972. 388 pp.

DEC:1972:FIC

[Dig72b] Digital Equipment Corporation. Fortran IV. compiler and object time system programmer’s manual. DEC, Maynard, MA, USA, 1972. various pp.

DEC:1972:PFI

[Dig72c] Digital Equipment Corporation. PDP-11 FORTRAN IV compiler and object time system programmer’s manual. Digital Equipment Corporation, Maynard, MA, USA, 1972. x + [241] pp.

DEC:1974:DFI

[Dig74] Digital Equipment Corporation. DECsystem-10 FORTRAN IV programmer’s reference manual. Digital Equipment Corporation, Maynard, MA, USA, 1974. various pp.

DEC:1975:DSF

[Dig75a] Digital Equipment Corporation.

DEC system 10: FORTRAN IV (F40) programmer’s reference manual. The Corporation, Maynard, MA, USA, 1975. 129 pp.

DEC:1975:DFI

[Dig75b] Digital Equipment Corporation.

DECsystem-10 FORTRAN IV programmer’s reference manual. The Corporation, Maynard, MA, USA, 1975. various pp.

DEC:1975:FIU

[Dig75c] Digital Equipment Corporation.

FORTRAN IV-PLUS user’s guide. Digital Equipment Corporation, Maynard, MA, USA, 1975. 138. pp.

DEC:1975:PFL

[Dig75d] Digital Equipment Corporation.

PDP-11 FORTRAN language reference manual. The Corporation, Maynard, MA, USA, 1975. various pp.

DEC:1975:RFI

[Dig75e] Digital Equipment Corporation.

RSTS/E FORTRAN IV utilities manual. The Corporation, Maynard, MA, USA, 1975. various pp.

DEC:1975:RFS

[Dig75f] Digital Equipment Corporation.

RT-11 Fortran scientific subroutines package reference manual. Digital Equipment Corporation, Maynard, MA, USA, 1975. 218 pp.

DEC:1975:RRF

[Dig75g] Digital Equipment Corporation. RT-11/RSTS/E FORTRAN IV user’s guide. The Corporation, Maynard, MA, USA, 1975. various pp.

DEC:1976:FRM

[Dig76a] Digital Equipment Corporation. FORTRAN reference manual. Digital Equipment Corp., Maynard, MA, USA, 1976. 226 pp.

DEC:1976:FRE

[Dig76b] Digital Equipment Corporation. FORTRAN/RT-11: extensions manual. The Corp., Maynard, MA, USA, 1976. v + 74 pp.

DEC:1976:RFG

[Dig76c] Digital Equipment Corporation. RT-11/FORTRAN: graphics extensions user’s guide. The Corp., Maynard, MA, USA, 1976. vi + 36 pp.

DEC:1977:DFP

[Dig77a] Digital Equipment Corporation. DECsystem10 FORTRAN programmer’s reference manual. The Corporation, Maynard, MA, USA, 1977. [ca. 200] pp.

DEC:1977:FRM

[Dig77b] Digital Equipment Corporation. FORTRAN reference manual. The Corporation, Maynard, MA, USA, 1977. 288 pp.

DEC:1977:FRE

[Dig77c] Digital Equipment Corporation. FORTRAN/RT-11: extensions manual. The Corp., Maynard, MA, USA, 1977. vii + 131 pp.

DEC:1977:PFL

[Dig77d] Digital Equipment Corporation.

PDP-11 FORTRAN language reference manual. The Corporation, Maynard, MA, USA, 1977. various pp.

DEC:1977:RFI

[Dig77e] Digital Equipment Corporation.

RSTS/E FORTRAN IV utilities manual. The Corporation, Maynard, MA, USA, 1977. various pp.

DEC:1977:RRF

[Dig77f] Digital Equipment Corporation.

RT-11/RSTS/E FORTRAN IV user’s guide. The Corporation, Maynard, MA, USA, revised edition, 1977. various pp.

DEC:1978:IRF

[Dig78a] Digital Equipment Corporation.

IAS/RSX FORTRAN IV user’s guide. The Corporation, Maynard, MA, USA, revised edition, 1978. 118 pp.

DEC:1978:VFIa

[Dig78b] Digital Equipment Corporation.

VAX-11 FORTRAN IV-PLUS: language reference manual. Digital Equipment, Maynard, MA, USA, 1978. various pp.

DEC:1978:VFIb

[Dig78c] Digital Equipment Corporation.

VAX-11 FORTRAN IV-PLUS: user’s guide. Digital Equipment, Maynard, MA, USA, 1978. various pp.

DEC:1979:FIU

[Dig79a] Digital Equipment Corporation. Fortran IV-plus user’s guide. The Corporation, Maynard, MA, USA, 1979. various pp.

DEC:1979:PFL

[Dig79b] Digital Equipment Corporation. PDP-11 FORTRAN language reference manual. The Corporation, Maynard, MA, USA, 1979. various pp.

DEC:1980:FIS

[Dig80a] Digital Equipment Corporation. FORTRAN IV A self-paced course. Digital Equipment Corporation, Maynard, MA, USA, 1980. x + 244 pp.

DEC:1980:RFI

[Dig80b] Digital Equipment Corporation. RSTS/E FORTRAN IV installation guide/release notes. The Corporation, Maynard, MA, USA, revised edition, 1980. 28 pp.

DEC:1980:RRF

[Dig80c] Digital Equipment Corporation. RT-11, RSTS/E FORTRAN IV user’s guide. Digital Equipment Corporation, Maynard, MA, USA, revised edition, 1980. 134 pp.

DEC:1980:VFI

[Dig80d] Digital Equipment Corporation. VAX-11 FORTRAN installation guide/release notes. Digital Equipment Corporation, Maynard, MA, USA, 1980. various pp.

DEC:1980:VFL

[Dig80e] Digital Equipment Corporation.

VAX-11 FORTRAN: language reference manual. Digital Equipment Corporation, Maynard, MA, USA, revised edition, 1980. xii + 254 pp.

DEC:1980:VFU

[Dig80f] Digital Equipment Corporation.

VAX-11 FORTRAN: user’s guide. Digital Equipment Corp., Maynard, MA, USA, revised edition, 1980. various pp.

DEC:1980:VFV

[Dig80g] Digital Equipment Corporation.

VAX-11 FORTRAN, version 2.1. Digital Equipment Corporation, Maynard, MA, USA, 1980. various pp.

Dillard:1979:PFS

[Dil79]  Carol A. Dillard.  PIE, a

Fortran subroutine for plotting circles, concentric circles, pie charts and shaded pie charts. ORNL/CSD/TM 82, Dept. of Energy, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1979. v + 17 pp. For sale by the National Technical Information Service.

Dineen:1969:FBM

[Din69] Robert Anthony Dineen. A FOR-TRAN based method for fitting frequency curves. Thesis (m.s.e.), Arizona State University, Tempe, AZ, USA, 1969. 109 pp.

Dinkel:1972:BRB

[Din72] John Dinkel. Book review: Numerical Methods of Mathematical Optimization with ALGOL and FOR-TRAN Programs (Hans P. Kunzi, H. G. Tzschach and C. A. Zehnder). SIAM Review, 14(2):359– 361, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dinter:1973:IC

[Din73] Heinz Dinter. Introduction to Computing. MacMillan Publishing Company, New York, NY, USA, February 1973. ISBN 0-02-329690-9. ???? pp. LCCN ????

deFreitas:1978:MTA

[dL78] Simplicio Lopes de Freitas and Pierre J. Lavell. A method for the time analysis of programs. IBM Systems Journal, 17(1):26– 38, 1978. CODEN IBMSA7. ISSN 0018-8670.

Briandais:1959:FSU

[dlB59] Renee de la Briandais. File searching using variable length keys. In Proc. WJCC, AFIPS, volume 15. AFIPS Press, Montvale, NJ, USA, 1959.

Dyck:1979:ICS

[DLS79] V. A. Dyck, J. D. Lawson, and J. A. Smith. Introduction to computing: structured problem solving using WATFIV-S. Reston Publishing Co., Inc., Reston, VA, USA, 1979. ISBN 0-8359-3158-7. xx + 604 pp. LCCN QA76.6 .D93 Sci-Eng.

Dimitry:1966:IFI

[DM66a] Donald L. Dimitry and Thomas H. Mott. Introduction to Fortran IV programming. Holt, Rinehart, and Winston, New York, NY, USA, 1966. xiv + 334 pp.

Dimitry:1966:SMI

[DM66b] Donald L. Dimitry and Thomas H.

Mott. Solutions manual for Introduction to Fortran IV programming. Holt, Rinehart, and Winston, New York, NY, USA, 1966. 48 pp.

DouglasAircraft:1966:FFM

[DM66c] Douglas Aircraft Company.Aircraft Division and McDonnell Douglas Corporation. FORMAT — FOR-TRAN matrix abstraction technique. Technical Report AFFDL-TR-66-207, Air Force Flight Dynamics Laboratory, Air Force Systems Command, Wright-Patterson Air Force Base, OH, USA, 1966. various pp.

Dimitry:1967:IFI

[DM67] Donald L. Dimitry and Thomas H.

Mott. Introduction to Fortran IV programming. Holt, Rinehart, and Winston, New York, NY, USA, 1967. xiv + 334 pp.

Dawson:1972:PSF

[DM72a] John A. Dawson and Paul M.

Mather. The Pearson system of frequency curves. Computer applications in the natural and social sciences 13, Dept. of Geography, University of Nottingham, Nottingham, UK, 1972. 42 pp.

Dimitry:1972:IPM

[DM72b] Donald L. Dimitry and Thomas H.

Mott. Introduccion a la programacion mediante Fortran IV. Interamericana, Mexico, DF, Mexico, 1972. xx + 355 pp.

Dorn:1972:NMF

[DM72c] William S. Dorn and Daniel D. McCracken. Numerical Methods With Fortran IV Case Studies. John Wiley and Sons, New York, London, Sydney, May 1972. ISBN 0-471-21918-5. xii + 447 pp. LCCN QA297 .D68. US$45.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471219185.

LopezdeMedinadeFader:1973:ICD

[dMdF73] Isabel Lopez de Medina de Fader. Introduccion a la computacion; diagramas de flujo y programas correspomdientes a algunos problemas de aritmetica y algebra, Fortran IV. Serie Cuadernos: Seccion matematicas y estadistica 46, Universidad Nacional de Cuyo Facultad de Ciencias Economicas, Mendoza, Argentina, 1973. 40 pp.

Dixon:1979:FSF

[DO79] John S. Dixon and Stan Openshaw. FORTRAN subroutines for the functional regionalisation of large sparse interaction matrices. Discussion paper 24, Centre for Urban and Regional Development Studies, University of Newcastle upon Tyne, Newcastle upon Tyne (The University), 1979. 10 pp.

Dock:1972:FIP

[Doc72] V. Thomas Dock. FORTRAN IV programming. Reston Publishing Co., Inc., Reston, VA, USA, 1972. ISBN 0-87909-271-8. xii + 287 pp. LCCN QA76.73.F25 D63.

Dock:1976:FIP

[Doc76] V. Thomas. Dock. Fortran IV Programming. Reston Publishing Co., Inc., Reston, VA, USA, second edition, January 1976. ISBN 0-87909-279-3. ix + 230 pp. LCCN QA76.73.F25 D631 1975. US$19.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0879092793.

Dock:1979:SFI

[Doc79] V. Thomas Dock. Structured FOR-TRAN IV programming. West series in data processing and information systems. West Publishing Company, St. Paul, MN, USA, 1979. ISBN 0-8299-0249-X. xii + 344 pp. LCCN QA76.73.F25D64.

Donaghey:1971:FDS

[Don71] Charles Edmund Donaghey. FOR-TRAN data structures for science and engineering students. ????, Dept. of Industrial and Systems Engineering, Cullen College of Engineering, University of Houston, Houston, TX, USA, 1971. 106 pp.

Donnelly:1973:AAB

[Don73a] Thomas G. Donnelly. ACM Algorithm 462: Bivariate normal distribution [S15]. Comm. ACM, 16(10):638, October 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Donovan:1973:SP

[Don73b] J. J. Donovan. Systems Programming. McGraw-Hill, New York, NY, USA, 1973. ISBN 0-07-017603-5.

Dorrenbacher:1979:PFP

[Dor79] John Dorrenbacher. POLISH: a Fortran program to edit Fortran programs. Technical report, University of Colorado, Dept. of Computer Science, Boulder, CO, USA, August 1979. 38 pp.

Didday:1973:FH

[DP73] Richard L. Didday and Rex L. Page. Fortran for humans. West Publishing Company, St. Paul, MN, USA, 1973. 255 pp.

Didday:1974:FH

[DP74a] Richard L. Didday and Rex L. Page. Fortran for Humans. West Publishing Company, St. Paul, MN, USA, 1974. ISBN 0-8299-0020-9. xv + 430 pp. LCCN QA76.73.F25 D52.

Din:1974:CTC

[DP74b] M. E. F. El Din and L. Peiram. Calculation of the traffic capacity of a private data network. Ericcson technics no. 2, 1974.

Dance:1976:ALD

[DP76a] D. L. Dance and U. W. Pooch. An adaptive on line data compression system. The Computer Journal, 19(3):216–224, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.

oxfordjournals.org/content/
19/3/216.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/
216.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/217.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
19/Issue_03/tiff/218.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/219.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/
220.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/221.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
19/Issue_03/tiff/222.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_
03/tiff/223.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/
224.tif.

DeDoncker:1976:ACI

[DP76b] E. De Doncker and R. Piessens.

Automatic computation of integrals with singular integrand, over a finite or an infinite range. Computing, 17(3):265–279, 1976. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Didday:1977:FH

[DP77] Richard L. Didday and Rex L.

Page. Fortran for humans. West Publishing Company, St. Paul, MN, USA, second edition, 1977. ISBN 0-8299-0125-6. xi + 468 pp.

DeSalvio:1970:CCI

[DPR70] A. J. DeSalvio, J. G. Purdy, and J. Rau. Creation and control of internal data bases under a Fortran programming environment. Comm. ACM, 13 (4):211–215, April 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

HessedePolanco:1980:UGF

[dPW80] Edith Hesse de Polanco and Peter Walker. A users guide to FASAP: a FORTRAN program for the analysis of farm survey data. CIM-MYT economics program working paper 80/3, International Maize and Wheat Improvement Center, Mexico, DF, Mexico, 1980. 50 pp.

Dowding:1970:BFI

[DR70] B. J. Dowding and C. B. Rogers. Basic Fortran IV. Blackie, Glasgow, Scotland, 1970. ISBN 0-216-88874-3. v + 90 pp.

Drath:1964:AUF

[Dra64] D. G. Drath. Another use of FOR-TRAN II chaining. Comm. ACM, 7(12):726, December 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dreyfus:1967:FIQ

[Dre67] Michel Dreyfus. Fortran IV [i.e. quatre]. Centre interarmées de recherche operationnelle. A 1. Dunod, Paris, France, 1967. 174 pp.

Dreyfus:1970:APG

[Dre70] Michel Dreyfus. Anleitung zum praktischen Gebrauch von FOR-TRAN IV. Verfahren der Datenverarbeitung. R. Oldenbourg, München, Germany, 1970. 222 pp.

Dreyfus:1972:FI

[Dre72] Michel Dreyfus. FORTRAN IV.

Centre Interarmées de Recherche Operationnelle; 1. Dunod, Paris, France, fifth edition, 1972. xii + 228 pp.

Dreyfus:1975:FI

[Dre75a] Michel Dreyfus. Fortran IV. Centre interarmées de recherche operationnelle. Dunod, Paris, France, 5e edition, 1975. ISBN 2-04-007726-X. xii + 228 pp.

Dreyfus:1975:FII

[Dre75b] Michel Dreyfus. FORTRAN IV [i.e. quatre]. Dunod Informatique Phase formation. Dunod, Paris, France, fifth edition, 1975. ISBN 2-04-007726-X. xii + 228 pp.

DeMaine:1962:DCP

[DS62] Paul A. D. De Maine and Robert D. Seawright. Digital computer programs for physical chemistry ALGOL 60 translations of the FORTRAN programs. MacMillan Publishing Company, New York, NY, USA, 1962. 2 v pp.

DeMaine:1963:DCP

[DS63] Paul A. D. De Maine and Robert D. Seawright. Digital computer programs for physical chemistry. Macmillan, Collier-Macmillan, New York, NY, USA, 1963. xxiii + 423 pp.

Davis:1966:FIP

[DS66] John C. Davis and Robert J. Sampson. Fortran II program for multivariate discriminant analysis using an IBM 1620 computer. Computer contribution 4, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1966. 8 pp.

Davis:1967:FIT

[DS67a] John C. Davis and Robert J. Sampson. FORTRAN II timetrend package for the IBM 1620 computer. Computer contribution 19, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 28 pp.

Dickmann:1967:FPC

[DS67b] D. B. Dickmann and G. E. Schacher. A Fortran program for calculating internal fields in general dipole lattices. Journal of Computational Physics, 2 (2):87–89, November 1967. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0021999167900277.

Dickson:1972:IFI

[DS72] Gary W. Dickson and Harry R. Smith. Introduction to FORTRAN IV programming; a self-paced approach. Rinehart Press, San Francisco, CA, USA, 1972. ix + 291 pp.

Dyson:1975:FCG

[DS75] Robert Dyson and Gillian Swaithes. FORTRAN codes for a global maximax and a local minimax solution to a stochastic programming problem. Warwick research in industrial and business studies 62, Centre for Industrial Economic and Business Research, University of Warwick, Coventry, UK, 1975. 55 pp.

Dickens:1976:CPS

[DS76] Brian Dickens and LeRoy W.

Schroeder. Computer programs for structural chemistry: MATCH 1 and MATCH 2, FORTRAN programs to predict and evaluate mutual orientation of polycrystals. NBS Technical note 893, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC. USA, 1976. 66 pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Daniel:1977:PFB

[DS77] Daniel and Loren E. Couger Shanon. Plaid for Fortran a Beginners Approach. R. D. Irwin, Homewood, IL, USA, June 1977. ISBN 0-256-01986-X. ???? pp. LCCN ????

Dupont:1974:LPP

[DT74] Philippe Dupont and Yves Tallineau.

Les langages de programmation en parallèle: Fortran, Cobol, PL/1, APL. Masson, Masson, France, 1974. ISBN 2-225-38368-5. vii + 149 pp.

Dubois:1977:IAF

[Dub77] Philippe Dubois. Initiation au Fortran par l’exemple. Publications de l’Institut français du petrole Cours de l’Ecole nationale supérieure du petrole et des moteurs. Editions Technip, Paris, France, 1977. ISBN 2-7108-0328-3. 111 pp.

Duff:1977:NSF

[Duf77a] I. S. Duff. MA28 — a set of Fortran subroutines for sparse unsymmetric linear equations. Technical Report R8730, HMSO, AERE Harwell, Harwell, Berkshire, UK, 1977.

Duff:1977:MSF

[Duf77b] Iain S. Duff. MA28 – a set of Fortran subroutines for sparse unsymmetric linear equations. Technical Report AERE R8730, Her Majesty’s Stationary Office, London, UK, 1977.

Duff:1980:MSF

[Duf80] Iain S. Duff. MA28, a set of Fortran subroutines for sparse unsymmetric linear equations. AERE-R 8730 Report (Atomic Energy Research Establishment (Harwell, Oxfordshire)); AERE-R 8730. Computer Science and Systems Division, AERE Harwell, Harwell, Berkshire, UK, 1980 revision edition, 1980. ISBN 0-7058-0593-X. 104 pp.

Dunn:1967:IDC

[Dun67] Walter L. Dunn. Introduction to digital computer problems using FORTRAN IV; preliminary edition. McGraw-Hill General Engineering series. McGraw-Hill, New York, NY, USA, 1967. various pp.

Dunn:1969:ISI

[Dun69a] Walter L. Dunn. Instructional supplement for Introduction to digital computer problems using FOR-TRAN IV. McGraw-Hill, New York, NY, USA, 1969. 33 pp.

Dunn:1969:IDC

[Dun69b] Walter L. Dunn. Introduction to digital computer problems using FORTRAN IV; a self-instructional programmed text. McGraw-Hill, New York, NY, USA, 1969. xii + 162 + 97 pp.

Dunn:1974:FCP

[Dun74] Walter L. Dunn. Fortran computer programming; self-study manual. Asuw lecture notes, University of Washington, Seattle, WA, USA, 1974. ca. 175 pp.

Duncan:1975:F

[Dun75a] A. K. Duncan. Fortran. International Computers (Australia) Pty Ltd, North Sydney, NSW, Australia, student edition, 1975. ISBN 0-9598529-0-5. vi + 106 pp.

Dunlap:1975:PAC

[Dun75b] Jackson Michael Dunlap. A preliminary analysis of component skills in introductory FORTRAN programming. Thesis (Ph.D.), University of Oregon, Eugene, OR, USA, 1975. vii + 122 pp.

Duncan:1977:F

[Dun77] A. K. Duncan. FORTRAN. International Computers, North Sydney, NSW, Australia, student edition, 1977. ISBN 0-9598529-1-3 (spiral binding). vi + 106 pp.

Dunham:1979:FPM

[Dun79] Charles Burton Dunham. Fortran program for minimax approximation by semicircles. Report 50, Dept. of Computer Science, University of Western Ontario, London, UK, 1979. ISBN 0-7714-0119-1. 3 + [11] pp.

Duncan:1980:BIF

[Dun80] A. K. Duncan. A basic introduction to FORTRAN programming. Techsearch, Adelaide, South Australia, Australia, 1980. ISBN 0-909386-23-4. iv + 68 pp.

Duris:1980:AFR

[Dur80] Charles S. Duris. Algorithm 547: FORTRAN routines for discrete cubic spline interpolation and smoothing [E1], [E3]. ACM Transactions on Mathematical Software, 6(1):92–103, March 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Davies:1969:FIC

[DvC69] A. F. Davies and D. A. van Cawenberghe. Fortran IV computer programs. Technical memoranda TEC 715, Canada, Dept. of Transport, Meteorological Branch, Toronto, Ontario, Canada, 1969. 15 pp.

Dangerfield:1970:BBT

[DW70] G. R. Dangerfield and R. L. Walsh. BOPTIC — a beam transport program in Fortran IV. Technical Report AAEC/TM 544, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1970. 14 + [24] pp.

Dawson:1971:BII

[DW71] Clive B. Dawson and Thomas C.

Wool. From Bits to If’s: An Introduction to Computers and Fortran IV. HarperCollins College Publishers, New York, NY, USA, January 1971. ISBN 0-06-041594-0. xii + 157 + [2] pp. LCCN ???? US$5.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0060415940.

Davis:1977:FHD

[DW77] H. Davis and L. Winslow. File handler design for a flexible data base management system. Technical Report WSU-CS-77-02, Washington State University, Pullman, WA, USA, 1977.

Derigs:1978:APM

[DZ78] U. Derigs and U. Zimmermann.

Augmenting path method for solving linear bottleneck assignment problems. Computing, 19(4):285– 295, 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

EP:1967:DUP

[É67] École polytechnique, Montréal, Québec.Laboratoire de calcul électronique. Description et utilisation de la programmatheque FORTRAN. Technical report, École polytechnique, Laboratoire de calcul électronique, Montréal, PQ, Canada, 1967. 125 pp.

Edwards:1980:DPC

[EB80] Perry Edwards and Bruce Broadwell. Data processing: computers in action with FORTRAN. Wadsworth, Pacific Grove, CA, USA, 1980. ISBN 0-534-00805-4. 472 pp. LCCN QA76.E29.

Edgell:1979:SCD

[Edg79] Stephen E. Edgell. A statistical check of the DECsystem-10 FORTRAN pseudorandom number generator. Behavior Research Methods and Instrumentation, 11(5):529–530, September 1979. CODEN BR-MIAC. ISSN 0005-7878. URL

                                        http://www.springerlink.com/
                                        content/ah71321478333g10/.

ESI:1970:CLE

[Edu70] Edutronics Systems International. Computer language: essentials of FORTRAN, 1970.

ESI:1972:EFPe

[Edu72a] Edutronics Systems International. Essentials of FORTRAN, part five, 1972.

ESI:1972:EFPd

[Edu72b] Edutronics Systems International. Essentials of FORTRAN, part four, 1972.

ESI:1972:EFPa

[Edu72c] Edutronics Systems International. Essentials of FORTRAN, part one, 1972.

ESI:1972:EFPf

[Edu72d] Edutronics Systems International. Essentials of FORTRAN, part six,

ESI:1972:EFPc

[Edu72e] Edutronics Systems International.

Essentials of FORTRAN, part three, 1972.

ESI:1972:EFPb

[Edu72f] Edutronics Systems International.

Essentials of FORTRAN, part two, 1972.

Edwards:1969:IFI

[Edw69] Anthony John Edwards. An introduction to Fortran IV programming for engineers and technicians. Draughtsmen’s and Allied Technicians’ Association, Richmond, VA, USA, 1969. ISBN 0-900328-02-9. 50 pp. LCCN QA76.5.E37.

Edwards:1973:FFI

[Edw73] Perry Edwards. Flowcharting and Fortran IV. McGraw-Hill, New York, NY, USA, January 1973. ISBN 0-07-019042-9. 147 pp. LCCN QA76.6 .E4. US$15.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070190429.

Edwards:1976:FPA

[Edw76a] John W. Edwards. A Fortran program for the analysis of linear continuous and sampled — data systems. NASA TM 56038, NASA Dryden Flight Research Center, Edwards, CA, USA, 1976. vi + 229 pp. National Technical Information Service. N; 76-18823.

Edwards:1976:MMP

[Edw76b] R. G. Edwards. MAPPROJ2– Fortran map projection subroutines — version 2. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1976. vii + 44 pp.

Engels:1977:WQ

[EE77] H. Engels and U. Eckhardt. Wilfquadrature. Computing, 18(3): 271–279, 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckelman:1976:CAF

[EF76] Carl A. Eckelman and David A. Fergus. Computer analysis (Fortran and Basic) of chair frames. Research bulletin 937, Purdue University Agricultural Experiment Station, West Lafayette, IN, USA, 1976. 24 pp.

Engvold:1968:EGP

[EH68] K. J. Engvold and J. L. Hughes. Education: a general-purpose display processing and tutorial system. Comm. ACM, 11(10): 697–702, October 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ehrman:1972:CSE

[Ehr72] J. R. Ehrman. Correspondence: On “Suggested Extensions to FORTRAN IV”. The Computer Journal, 15(4): 332, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
15/Issue_04/150332.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/332.tif. See

[Fin72b].

Einarsson:1972:AAC

[Ein72] Bo Einarsson. ACM Algorithm 418: Calculation of Fourier integrals [D1]. Comm. ACM, 15 (1):47–48, January 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?, ?].

Einarsson:1974:STE

[Ein74] Bo Einarsson. Remark on “Algorithm 443: Solution of the transcendental equation w exp(w) = x”. Comm. ACM, 17(4):225, April 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [FSC73].

Einarsson:1976:RAR

[Ein76] B. Einarsson. Remark on algorithm 24, regarding a very common violation of the Fortran Standard. Computing, 16(3):291–292, 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See [JR75].

Esposito:1974:WFI

[EKM74] V. J. Esposito, K. Kesavan, and B. Maul. WFLASH, a FORTRAN — IV computer program for simulation of transients in a multiloop PWR. Technical Report WCAP-8261 Revision 1, Westinghouse Electric Corporation, Nuclear Energy Systems, Pittsburgh, PA, USA, 1974. various pp.

Elder:1970:FVI

[Eld70] H. A. Elder. On the feasibility of voice input to an on-line computer processing system. Comm. ACM, 13(6):339–346, June 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eldring:1977:ISE

[Eld77] S. Eldring. IFTRAN, a structured extension of FORTRAN as used on a CYBER 74. Technical note 25, Environment Canada, Fisheries and Marine Service, Ottawa, Ontario, Canada, 1977. 4 + [5] pp.

EA:1968:EFI

[Ele68] Electronic Associates. EAI 640 Fortran IV language reference manual. Publication; no. 00 827 0022-0. EAI, West Long Branch, NJ, USA, 1968. various pp.

Elkin:1965:LEC

[Elk65] Sanford Elkin. Letter to the Editor: On combining FORTRAN and COBOL in problem solving. Comm. ACM, 8(10):594, October 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ellis:1978:FPI

[Ell78] David Brande Ellis. A Fortran preprocessor incorporating macro expansion and dynamic data structures. Thesis (m.s.), South Dakota School of Mines and Technology, Rapid City, SD, USA, 1978. ???? pp.

Ellis:1980:SFF

[Ell80] T. M. R. Ellis. Structured Fortran: a Fortran 77 programming course. Technical report, University of Sheffield Computing Services, Sheffield, UK, 1980. ISBN 0-9506910-1-1. 121 pp.

Elson:1973:CPL

[Els73] Mark Elson. Concepts of Programming Languages. Science Research Associates, Chicago, IL, USA, 1973. xii + 333 pp. LCCN QA76.7 .E471.

Elton:1966:BRBa

[Elt66] H. D. Elton. Book review: A Guide to FORTRAN IV Programming, by D. D. McCracken. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2): 218–219, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2986741.

Embley:1978:TDF

[Emb78] David W. Embley. Teaching disciplined Fortran programming via unfort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10 (2):34–38, June 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engel:1974:CFS

[Eng74] Frank Engel, Jr. Correspondence on Fortran standardization. ACM SIGNUM Newsletter, 9(4):9–10, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Engelsohn:1975:PFA

[Eng75] Harold S. Engelsohn. Practical Fortran: An Applied and Simplified, Problem-Solving Approach. MacMillan Publishing Company, New York, NY, USA, May 1975. ISBN 0-02-333730-3. xii + 434 pp. LCCN ???? US$9.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0023337303.

Entwisle:1963:APC

[Ent63] Doris R. Entwisle. Auto-primer in computer programming for the IBM 1620 in FORTRAN. A Blaisdell book in the pure and applied sciences. Blaisdell Pub. Co., Waltham, MA, USA, 1963. x + 345 pp.

Entropy:1980:AFI

[Ent80a] Entropy Limited. Advanced FOR-TRAN IV utilities for data general computers: user’s manual: for software packages EP37-01-00 [through] EP37-10-00. Entropy Limited, Lincoln, MA, USA, 1980. 231 pp.

Entropy:1980:ELP

[Ent80b] Entropy Limited. Entropy Limited presents unique software packages for data general computers. Entropy Limited, Lincoln, MA, USA, 1980. 17 pp.

Engel:1966:BF

[EO66] Frank A. Engel III and Lewis A. Ondis II. On behalf of FOR-TRAN. Comm. ACM, 9(4):257, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Esler:1967:FIP

[EP67] J. E. Esler and Floyd W. Preston.

Fortran IV program for the GE 625 to compute the power spectrum of geological surfaces. Computer contribution 16, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 23 pp.

Eppler:1974:FPP

[Epp74] Richard Eppler. A Fortran program for profile and boundary layer calculation. Technical Report FTD-HC-23-897-74, Translation Division, Foreign Technology Division, Wright-Patterson AFB, OH, USA, 1974. 25 pp.

Ramden:1976:JAF

[êR76] H. êA. Ramdén. JCL and advanced Fortran programming. Methods in geomathematics ; 2. American Elsevier and Elsevier, New York, NY, USA and Amsterdam, The Netherlands, 1976. ISBN 0-444-41415-0. viii + 169 pp. LCCN QA76.73.F25 R35.

Eyton:1979:IFP

[ER79] J. Ronald Eyton and Curtis C.

Roseman. An introduction to FORTRAN and the programming of spatial data. Occasional publication 13, Geography Graduate Student Association, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1979. iv + 139 pp.

Erard:1977:PSF

[Era77] P. J. Erard. Programmation structurée en FORTRAN. Centre de calcul, Université de Neuchatel 2, Centre de calcul, Université de Neuchatel, Neuchatel, Switzerland, 1977. 26 pp.

Erdos:1980:RFP

[Erd80] John Erdos. R2D2 — a FORTRAN program for twodimensional chemically reacting hyperthermal, internal flows. Technical report / air force flight dynamics laboratory (u.s.); affdltr-79-3162 technical report (air force flight dynamics laboratory (u.s.)); affdl-tr-79-3162, Air Force Filught Dynamics Laboratory, Wright-Patterson AFB, OH, USA, 1980. ???? pp.

Erickson:1975:APA

[Eri75] David B. Erickson. Array processing on an array processor. ACM SIGPLAN Notices, 10(3):17–24, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edeline:1974:IF

[ES74a] M. Edeline and Jacques Subsol. Initiation a FORTRAN. Compagnie internationale de services en informatique, Saclay, France, 1974. 152 pp.

Eidson:1974:AAC

[ES74b] Harold D. Eidson and Larry L. Schumaker. ACM Algorithm 485: Computation of g-splines via a factorization method [E2]. Comm.

ACM, 17(9):526–530, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Engquist:1975:DSA

[ES75] Björn Engquist and Tom Smedsaas. DCG: a system for automatic code generation for hyperbolic problems (abstract). ACM SIGNUM Newsletter, 10(4):20– 22, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Epley:1978:TSC

[ES78] Donald Epley and Ted Sjoerdsma.

A two-semester course sequence in introductory programming using PL/1 — a rationale and overview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(3):113–119, August 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 9th SIGCSE symposium on Computer science education.

Esler:1968:KFI

[ESD68] J. E. Esler, Paul F. Smith, and John C. Davis. KWIKR8, a Fortran IV program for multiple regression and geologic trend analysis. Computer contribution 28, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1968. 31 pp.

Evans:1972:SPA

[Eva72a] D. J. Evans, editor. Software 72, Proc. of a Conference Sponsored By Software World at the University of Kent at Canterbury on 24–26 July 1972. Transcripta Books, London, UK, 1972. ISBN 0-903012-07-3.

Evans:1972:SPC

[Eva72b] D. J. Evans, editor. Software 72, Proc. of a Conference Sponsored by Software World at the University of Kent at Canterbury on 24–26 July 1972. Transcripta Books, London, UK, 1972. ISBN 0-903012-07-3.

Fairley:1974:GCB

[Fai74] Richard E. Fairley. A graduate curriculum in Business-Oriented Computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6 (1):37–39, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Fang:1965:FIP

[Fan65] David Hsien-Chen Fang. FOR-TRAN II program for testing of positive real functions. Thesis (m.s.), George Washington University, Washington, DC, USA, 1965. viii + 64 pp.

Farina:1966:FIS

[Far66] Mario V. Farina. Fortran IV Self-Taught. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, May 1966. ISBN 0-13-329722-5. xi + 426 pp. LCCN ???? US$17.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0133297225.

Farina:1974:FIC

[Far74] Italo Hilario Farina. Fortran IV:

curso de programacion para computadoras digitales. Manuales. Editorial Universitaria de Buenos Aires, Buenos Aires, Argentina, second edition, 1974. xv + 212 pp.

Farino:1976:EF

[Far76] Carol S. Farino. An extended Fortran. Thesis (m.s.), Virginia Commonwealth University, Dept. of Mathematical Sciences, Richmond, VA, USA, 1976. 100 pp.

Fast:1967:CRS

[Fas67] Arlo W. Fast. Correlationregression with scattergram, IBM 1620, Fortran II-D. Transactions of the American Fisheries Society, 96(2):230, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fateman:1978:LMD

[Fat78] Richard J. Fateman. Is a Lisp machine different from a Fortran machine? SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 12 (3):8–11, August 1978. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Frayer:1969:IES

[FB69] Warren E. Frayer and Joseph E.

Barnard. Introduction to elementary scientific FORTRAN programming for natural resource biologists. Technical report, Dept. of Forest and Wood Sciences, College of Forestry and Natural Resources, Colorado State University, Fort Collins, CO, USA, 1969. 53 pp.

Forbes:1973:FPC

[FB73] J. M. Forbes and A. S. Bramson. A FORTRAN program for computing steady-state composition models of the upper atmosphere. Air force surveys in geophysics 276, Air Force Cambridge Research Laboratories, Bedford, MA, USA, 1973. 39 pp.

Fegan:1979:FCP

[FB79] Richard C. Fegan and Susan L. Brosche. FORTRAN computer programming for statistics: a manual. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1979. ISBN 0-13-329326-2 (paperback). vii + 79 pp.

Fedako:1963:PFE

[Fed63] John E. Fedako. Pracniques: 1410 Fortran edit feature. Comm. ACM, 6(6):310–311, June 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

FCCTS:1970:FCV

[Fed70] Federal Cobol Compiler Testing Service (U.S.). FORTRAN compiler validation summary report. Validation number FCVS66-VSR205. ADA040385, Federal Cobol Compiler Testing Service (U.S.), Washington, DC, USA, 1970. 20 pp. Available from National Technical Information Service.

FCCTS:1980:FCV

[Fed87] Federal Cobol Compiler Testing Service (U.S.). FORTRAN compiler validation summary report. Technical Report FCVS66-VSR210. ADA040392, Federal Cobol Compiler Testing Service (U.S.), Washington, DC, USA, 1980 (or 1987??). 16 pp. Available from National Technical Information Service.

Fehlmann:1968:AFC

[Feh68] M. Fehlmann. AAEC-FMO Fortran crystallographic Monte-Carlo/optimal-shift program. AAEC/TM 434, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1968. various pp.

Feldman:1975:BDA

[Fel75a] Stuart I. Feldman. A brief description of Altran. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):12–20, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fellows:1975:CGF

[Fel75b] David Michael Fellows. Comments on “A general Fortran emulator for IBM 360/370 random number generator ‘RANDU’”. Technical report 6, School of Computer Science, University of New Brunswick, Fredericton, NB, Canada, 1975. b + [4] pp.

Feldman:1976:FLC

[Fel76a] S. I. Feldman. A Fortranner’s lament: Comments on the draft proposed ANS FORTRAN standard. ACM SIGPLAN Notices, 11 (12):25–34, December 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1976:FGP

[Fel76b] Stuart I. Feldman. FORTLEX — a general purpose lexical analyzer for Fortran. Technical Report 51, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1976.

Feldman:1979:IPF

[Fel79] Stuart I. Feldman. Implementation of a portable Fortran 77 compiler using modern tools. ACM SIGPLAN Notices, 14(8):98–106, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the ACM SIG-PLAN ’79 Symposium on Compiler Construction.

Fencl:1973:AAR

[Fen73] Zdeněk Fencl. ACM Algorithm 456: Routing problem. Comm. ACM, 16(9):572–574, September 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?].

Ferguson:1960:IOB

[Fer60] David E. Ferguson. Input-output buffering and Fortran. Journal of the ACM, 7(1):1–9, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fernandez:1963:PFP

[Fer63] B. Fernandez. Programme Fortran pour le calcul des corrections d’absorption et de double diffusion dans les mesures de sections efficaces pour les neutrons rapides par la méthode de Monte-Carlo. Commissariat a l’énergie atomique. Rapport 2374, Centre d’études nucléaires de Saclay, Gif sur Yvette, France, 1963. 10 pp.

Feuchter:1977:FSC

[Feu77] Christopher A. Feuchter. FOR-TRAN software for creating and maintaining a library cataloging system on scientifically oriented computers. United States Air Force Systems Command Directorate of Aerospace Studies. DAS TR-77-2, U.S. Air Force. Systems Command, Kirtland AFB, NM, USA, 1977. ???? pp.

Flores:1975:SSL

[FF75] I. Flores and M. Feuerman. Source statement libraries and IBM System/370. Computer Languages, 1 (2):139–150, June 1975. CODEN COLADA. ISSN 0096-0551.

Friedmann:1975:FI

[FGH75] Jehosua Friedmann, Philip J.

Greenberg, and Alan Hoffberg. FORTRAN IV. Wiley selfteaching guides. John Wiley and Sons, New York, London, Sydney, 1975. ISBN 0-471-28082-8. xii + 452 pp. LCCN QA76.73.F25 F75.

Friedmann:1980:FI

[FGH80a] Jehosua Friedmann, Philip J.

Greenberg, and Alan Hoffberg. Fortran IV. Wiley self-teaching guides. John Wiley and Sons, New York, London, Sydney, second edition, December 1980. ISBN 0-471-07771-2. xii + 499 pp. LCCN QA76.73.F25 .F75 1980. US$14.95; US$8.95 (est.). URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471077712.

Friedmann:1980:LFI

[FGH80b] Jehosva Friedmann, Philip J. Greenberg, and Alan Hoffberg. Lenguaje FORTRAN IV. Serie instruccion programada Limusa. Editorial Limusa, Mexico, DF, Mexico, 1980. ISBN 968-18-0978-5. xvi + 515 pp.

Freiman:1972:IPP

[FGR72] C. V. Freiman, J. E. Griffith, and J. L. Rosenfeld, editors. Information processing 71: proceedings of IFIP Congress 71, Ljubjana, Yugoslavia, August 23–28, 1971, volume 1 of IFIP congress series. North-Holland, Amsterdam, The Netherlands, 1972. ISBN 0-7204-2063-6. LCCN QA75.5 .I532

1971. URL https://dblp.org/
db/conf/ifip/ifip71-1.html.

Friedrich:1971:OFC

[FH71] R. O. Friedrich and J. A. Hafford. ORVAC-CT: a FORTRAN code for the calculation of a desalination plant incorporating a vapor compressor system driven by a condensing turbine. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1971. vii + 212 pp.

Friedman:1974:ETR

[FH74] Theodore D. Friedman and Lance J. Hoffman. Execution time requirements for encipherment programs. Comm. ACM, 17 (8):445–449, August 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [?].

Fox:1978:AFP

[FHS78] P. A. Fox, A. D. Hall, and N. L.

Schryer. Algorithm 528, framework for a portable library (Z). ACM Transactions on Mathematical Software, 4:177–188, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fox:197x:BUP

[FHS7x] P. A. Fox, A. D. Hall, and N. L. Schryer. Basic utilities for portable FORTRAN libraries. Computing Science Technical Report 37, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? 197x. 53 pp. URL https://

archive.computerhistory.org/
resources/access/text/2022/
08/102805039-05-01-acc.pdf.

Fiala:1973:AAS

[Fia73] Frantisek Fiala. ACM Algorithm 449: Solution of linear programming problems in 0-1 variables [H1]. Comm. ACM, 16(7):445–447, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fick:1971:DFS

[Fic71] Goran Fick. DHAMDI: a Fortran subroutine to integrate a set of first order, ordinary differential equations containing discontinuities. FOA 2 rapport C 2504-E5. N74-13290, Forsvarets National Technical Information Service, Stockholm, Sweden, 1971. 22 pp.

Fick:1973:FTP

[Fic73] Goran Fick. Flexible table printing subroutines in FORTRAN. Technical Report N75-12657, Forsvarets Forskningsanstalt, Avdelning 2, Stockholm, Sweden, 1973. 64 pp. Reproduced by National Technical Information Service.

Finn:1968:MUM

[Fin68] Jeremy D. Finn. Multivariance; univariate and multivariate analysis of variance and covariance: a Fortran IV program. Technical report, Department of Educational Psychology, State University of New York at Buffalo, Buffalo, NY, USA, 1968. i + 109 pp.

Findler:1972:FHL

[Fin72a] N. V. Findler. Four high-level extensions of FORTRAN IV: SLIP, AMPPL-II, TREETRAN, SYM-BOLANG. Spartan Books, New York, NY, USA, 1972. ISBN 0-87671-168-9. ???? pp. LCCN ????

Finn:1972:CSE

[Fin72b] A. J. Finn. Correspondence: Suggested extensions to FOR-TRAN IV. The Computer Journal, 15(1):12, February 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        15/Issue_01/150012.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/12.tif. See

correspondence [Mee72, Hal72b, Ehr72].

Finn:1972:SEF

[Fin72c] A. J. Finn. Suggested extension to Fortran IV. The Computer Journal, 15(1):12, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
15/1/12.full.pdf+html.

Finn:1972:MUMa

[Fin72d] Jeremy D. Finn. Multivariance; univariate and multivariate analysis of variance, covariance, and regression: a Fortran IV program. Technical report, Dept. of Educational Psychology, State University of New York at Buffalo, Buffalo, NY, USA, 1972. 134 pp.

Finn:1972:MUMc

[Fin72e] Jeremy D. Finn. Multivariance; univariate and multivariate analysis of variance, covariance, and regression: a Fortran IV program. National Educational Resources, Inc., Chicago, IL, USA, 1972. 134 pp.

Finn:1972:MUMb

[Fin72f] Jeremy D. Finn. Multivariance:

univariate and multivariate analysis of variance, covariance, and regression: a Fortran IV program, version V, release 2, June 1974. International Educational Services, Chicago, IL, USA, 1972. 133 pp.

Finn:1977:MUM

[Fin77] Jeremy D. Finn. Multivariance: univariate and multivariate analysis of variance, covariance, regression and repeated measures: a FORTRAN IV program: [User’s guide]. National Educational Resources, Inc., Chicago, IL, USA, 1977. 166 + 61 pp.

Fisher:1970:FPF

[Fis70] James R. Fisher. Fortran program for fast Fourier transform. NRL Report 7041, Naval Research Laboratory, Washington, Springfield, VA, USA, 1970. 22 pp. Reproduced by National Technical Information Service.

Fisher:1971:IFP

[Fis71] D. D. Fisher. An introduction to FORTRAN programming: an IPI approach. Technical report, Oklahoma State University, Stillwater, OK, USA, 1971. vii + 267 pp.

Fisher:1976:UGS

[Fis76] W. Burns Fisher. Users’ guide to STGPAK: a string package for FORTRAN programmers: version V001C. Computer-related development memo 5a, Production Automation Project, College of Engineering and Applied Science, University of Rochester, Rochester, NY, USA, 1976. 14 pp. Reproduced by National Technical Information Service, PB-272 431.

Fishbone:1978:SDB

[Fis78] Leslie G. Fishbone. A System 2000 data base for MAGEN data and its associated FORTRAN programs.

Informal report BNL 25680., U.S. Dept. of Energy, Washington, DC, USA, December 29, 1978. 48 pp.

Fischer:1979:FSF

[Fis79] Steven K. Fischer. FLOCHT, a system of Fortran codes for computer drawn flow charts and diagrams. ORNL/CSD/TM 61, Dept. of Energy, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1979. ix + 87 pp. For sale by the National Technical Information Service.

Fitzgerald:1974:CSF

[Fit74] Kenneth E. Fitzgerald. Comparison of some FORTRAN programs for matrix inversion. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 78B:15–33, 1974. CO-DEN JNBBAU. ISSN 0022-4340.

Fitzgerald:1975:FCE

[Fit75] John Anthony Fitzgerald. A FORTRAN cross-compiler for the EA1 640 digital computer. Thesis (m.sc.), University of New Brunswick, Fredericton, NB, Canada, 1975. vi + 44 pp.

Ford:1980:BRB

[FJA80a] Brian Ford, David Jacobs, and A. M. Addyman. Book reviews: Workshop on Reliable Software: Applied Computer Science, by Peter Raulefs, 1979; Reliable Software Through Composite Design, by G. J. Myers, 1975; Fortran, PL/1 and the Algols, by Brian Meek, 1978. The Computer Journal, 23(2):146, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
23/Issue_02/tiff/146.tif.

Ford:1980:BRW

[FJA80b] Brian Ford, David Jacobs, and A. M. Addyman. Book reviews: Workshop on Reliable Software: Applied Computer Science, by Peter Raulefs, 1979; Reliable Software Through Composite Design, by G. J. Myers, 1975; Fortran, PL/1 and the Algols, by Brian Meek, 1978. The Computer Journal, 23(2):146, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
23/Issue_02/tiff/146.tif.

Friedman:1976:SPC

[FK76] Frank L. Friedman and Elliot B. Koffman. Some pedagogic considerations in teaching elementary programming using structured FORTRAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):1–10, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Friedman:1977:PSS

[FK77a] Frank L. Friedman and Elliot B. Koffman. Problem solving and structured programming in FOR-TRAN. Addison-Wesley series in computer science and information processing. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-01967-1. xvi + 404 + [68] pp. LCCN QA76.73.F25 .F74.

Friedman:1977:TPS

[FK77b] Frank L. Friedman and Elliot B.

Koffman. Teaching problem solving and structured programming in FORTRAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9 (1):63–68, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Frisch:1974:MMFb

[FL74] Michael J. Frisch and Lawrence Liddiard. MNF (MiNnesota FOR-TRAN) reference manual for CDC 6000/7000 Cyber series computers. Technical report, University of Minnesota Computer Center, Minneapolis, MN, USA, 1974. 544 pp.

Fletcher:1976:QQI

[FL76] Sharon Fletcher and Thomas Luce. QUICK: a quasi-interactive computer-based instructional system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):275–279, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Flack:1971:FIP

[Fla71] P. E. Flack. A FORTRAN IV program for calculating the transfer function of a series of tubes, for a sinusoidal pressure input. Structures and materials note 370, Aeronautical Research Laboratories, Melbourne, Victoria, Australia, 1971. ISBN 0-642-97709-7. 12 + [14] pp.

Flavell:1972:CAR

[Fla72] A. J. Flavell. Correspondence: On “Another round of FOR-TRAN”. The Computer Journal, 15(1):92, February 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        15/Issue_01/150092.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        15/Issue_01/tiff/92.tif. See

[Cha71b, Dew72, Hal72b, Cha72].

Flanders:1977:D

[Fla77] P. M. Flanders. DAP-Fortran. Technical report, Heinrich-Heine-Universität Düsseldorf (??), Düsseldorf, Germany, 1977.

Fletcher:1970:FSG

[Fle70] R. Fletcher. A Fortran subroutine for general quadratic programming. AERE-R 6370, United Kingdom Atomic Energy Authority, Harwell, Berkshire, UK, 1970. 14 pp.

Fletcher:1972:FSM

[Fle72] R. Fletcher. Fortran subroutines for minimization by quasi-Newton methods. Technical report, United Kingdom Atomic Energy Authority, Harwell, Berkshire, UK, 1972. 29 pp.

Frisch:1970:MMF

[FLM70] Michael J. Frisch, Lawrence Liddiard, and E. James Mundstock. MNF (MiNnesota FOR-TRAN) reference manual for CDC 6000/7000/Cyber series computers. Control Data Corporation, Minneapolis, MN, USA, 1970. ix + 523 pp.

Frisch:1974:MMFa

[FLM74] Michael J. Frisch, Lawrence Liddiard, and E. James Mundstock. MNF (Minnesota FOR-TRAN) reference manual for CDC 6000/7000/Cyber series computers. Technical report, University Computer Center, University of Minnesota, Minneapolis, MN, USA, 1974. ix + 337 pp.

Florian:1970:BRF

[Flo70a] H. Florian. Book review: F. Stuart, Fortran Programming. Computing, 5(1):95, 1970. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Florian:1970:BRW

[Flo70b] H. Florian. Book review: W.

E. Spiess und F. G. Rheingans, Einführung in das Programmieren in FORTRAN. Computing, 6(3– 4):376, 1970. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Flon:1975:RSP

[Flo75] L. Flon. On research in structured programming. ACM SIG-PLAN Notices, 10(10):16–17, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

FPS:1978:APF

[Flo78a] Floating Point Systems, Inc. Array Processor FORTRAN Reference Manual. Beaverton (??), OR, USA, document number 860-7408-000 edition, 1978.

Floyd:1978:UGR

[Flo78b] Benzell Floyd. User’s guide for the rational FORTRAN pre-processor software package. Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1978. iv + 165 pp.

Flynn:1973:SUG

[Fly73] John A. Flynn. SFTRAN user guide. Section 914, Internal Computing Memorandum 337, Jet Propulsion Laboratory, Pasadena, CA, USA, July 1973. Also catalogued as JPL Document No.1846-7.

Flax:1976:FSE

[FM76] Lawrence Flax and Janet P. Mason. Fortran subroutines to evaluate, in single or double precision, Bessel functions of the first and second kinds for complex arguments. NRL report 7997, U.S. Dept. of Defense, Dept. of the Navy, Office of Naval Research, Naval Research Laboratory, Springfield, VA, Washington, DC. USA, 1976. iii + 26 pp. For sale by the National Technical Information Service.

Feddes:1978:FIP

[FMC78] J. J. R. Feddes, J. B. McQuitty, and W. D. Campbell. Fortran IV program to predict heating and ventilation requirements of totally enclosed farm buildings: a user instruction guide. Research bulletin 78-1, Dept. of Agricultural Engineering, The University of Alberta, Edmonton, Alberta, Canada, 1978. ii + 31 pp.

Forsythe:1977:CMM

[FMM77] George E. (George Elmer) Forsythe, Michael A. Malcolm, and Cleve B. Moler. Computer Methods for Mathematical Computations. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1977. ISBN 0-13-165332-6. xi + 259 pp. LCCN QA297 .F5681. US$16.95. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous

ftp from math.liu.se in pub/
references.

Forsythe:1980:MMM

[FMM80] George E. Forsythe, Michael A.

Malcolm, and Cleve B. Moler. Mašinnye metody matematičeskich vyčislenij. (Russian) [Computer Methods for Mathematical Computations]. Izdatel’stvo Mir, Moscow, USSR, 1980. 280 pp. Translated from the English by Kh. D. Ikramov.

Forsythe:1970:CSF

[For70] Alexandra I. Forsythe. Computer science: FORTRAN language. John Wiley and Sons, New York, London, Sydney, 1970. ISBN 0-471-26679-5. ix + 181 pp. LCCN QA76.5.C61265.

Ford:1971:BFI

[For71] Donald H. Ford. Basic FORTRAN IV programming. Irwin-Dorsey series in information processing. R. D. Irwin, Homewood, IL, USA, 1971. xi + 254 pp.

Forsythe:1973:PF

[For73] Alexandra I. Forsythe. Programacion FORTRAN. Tecnicas de computacion. Editorial Limusa, Mexico, DF, Mexico, 1973. ISBN 968-18-0814-2. 194 pp.

Ford:1974:BFI

[For74] Donald H. Ford. Basic FORTRAN IV programming. Irwin-Dorsey information processing series. R. D. Irwin, Homewood, IL, USA, revised edition, 1974. ISBN 0-256-01580-5. xi + 278 pp. LCCN QA76.73.F25 .F66 1974. Third ed. (1978) published under title: Standard FORTRAN programming.

Forsythe:1975:CSP

[For75] Alexandra I. Forsythe. Computer science: programming in FOR-TRAN IV with WATFOR WAT-FIV. John Wiley and Sons, New York, London, Sydney, 1975. ISBN 0-471-26685-X. vi + 210 pp.

Ford:1978:SFP

[For78] Donald H. Ford. Standard FOR-TRAN programming. The Irwin series in information and decision sciences. R. D. Irwin, Homewood, IL, USA, third edition, 1978. ISBN 0-256-01998-3. xv + 330 pp. LCCN QA76.73.F25,F664 1978. Earlier ed. published under title: Basic FORTRAN IV programming. Includes index.

Forsythe:1979:PF

[For79] Alexandra I. Forsythe. Programacion FORTRAN. Tecnicas de computacion. Editorial Limusa, Mexico, DF, Mexico, 1979. ISBN 968-18-0815-0. 194 pp.

Foster:1973:SMC

[Fos73] W. E. Foster. Software for a multilingual computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(5): 15–19, January 1973. CODEN HPJOAX. ISSN 0018-1153.

Fosdick:1974:BFP

[Fos74] Lloyd Dudley Fosdick. BRNANL, a Fortran program to identify basic blocks in Fortran programs. Technical report, University of Colorado, Boulder, CO, USA, 1974. 29 pp.

Fox:1964:FFP

[Fox64] William T. Fox. Fortran and Fap program for calculating and plotting time-trend curves using an IBM 7090 or 7094/1401 computer system. Special distribution publication 12, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1964. 24 pp.

Fox:1967:FIP

[Fox67] William T. Fox. Fortran IV program for vector trend analyses of directional data. Computer contribution 11, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 36 pp.

Foxworth:1970:LNF

[Fox75] Mike Foxworth. Lecture notes on FORTRAN. Technical report, College of Business Administration, Univ. of South Carolina, Columbia, SC, USA, 1970 (or 1975??). 136 pp.

Fox:1978:PPS

[Fox78a] Phyllis A. Fox. PORT: a portable subroutine library. ACM SIGNUM Newsletter, 13(1):14–15, March 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Foxworth:1978:LNF

[Fox78b] Mike Foxworth. Lecture notes on FORTRAN. Technical report, College of Business Administration, Univ. of South Carolina, Columbia, SC, USA, 1978. 186 pp.

FroeseFischer:1973:EES

[FP73] C. Froese Fischer and D. W. B. Prentice. Exact evaluation of Slater integrals using ALTRAN, a symbol manipulation language. Computer Physics Communications, 6(4):157–164, October 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465573900970.

Finger:1975:SFI

[FP75] Larry W. Finger and E. Prince.

A system of Fortran IV computer programs for crystal structure computations. NBS Technical note 854, U.S. Department of Commerce, National Bureau of Standards, Washington, DC. USA, 1975. iv + 128 pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Findler:1972:FHE

[FPB72] N. V. Findler, John L. Pfaltz, and Herbert J. Bernstein. Four highlevel extensions of FORTRAN IV: SLIP, AMPPL-II, TREETRAN, SYMBOLANG. Spartan Books, New York, NY, USA, 1972. ISBN 0-87671-168-9. xiii + 392 pp. LCCN QA76.73.F25 F55.

Fraley:1977:RF

[Fra77] R. A. Fraley. On replacing Fortran. ACM SIGPLAN Notices, 12 (9):130–132, September 1977. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frane:1979:BFP

[Fra79] James W. Frane. The BMDP Fortran programmer’s guide: an overview. BMDP technical report 66, BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 6 pp.

Freeman:1973:SAA

[Fre73] P. R. Freeman. Statistical algorithms: Algorithm AS 59: Hypergeometric probabilities. Applied Statistics, 22(1):130–133, March 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.
                                        cmu.edu/apstat/59.

Freeman:1974:BFA

[Fre74]  P. Freeman.  B 74-34 For-

tran and Algol: a programmed course for students of science and technology. IEEE Transactions on Computers, C-23(5):559–560, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672583.

Fred:1976:DSP

[Fre76] Taylor Fred. Digital Signal Processing in Fortran. Lexington Books, Lexington, MA, USA, June 1976. ISBN 0-669-00330-1. ???? pp. LCCN ????

Friedman:1969:DRG

[Fri69] Joyce Friedman. Directed random generation of sentences. Comm. ACM, 12(1):40–46, January 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Friedman:1970:FIK

[Fri70] Joyce Friedman. Fortran implementation of the Kay context-free parser. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (16):19–45, October 1970. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Friedrich:1971:OOT

[Fri71a] R. O. Friedrich. ORSEF-2 and ORSEF-3: Two FORTRAN codes for the calculation of desalination plant designs using multistage flash evaporation. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1971. vi + 358 pp.

Frisch:1971:MRM

[Fri71b] Michael Frisch. MNF reference manual: preliminary version. Technical report, University Computer Center, University of Minnesota, Minneapolis, MN, USA, 1971. ca. 250 pp.

Frisch:1973:MMF

[Fri73] Michael J. Frisch. MNF (MiNnesota FORTRAN) reference manual for CDC 6000/7000/Cyber series computers. Technical report, University Computer Center, University of Minnesota, Minneapolis, MN, USA, 1973. 337 pp.

Friedman:1975:ETD

[Fri75a] Frank L. Friedman. An experience in teaching disciplined programming at an elementary level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(3):38– 43, September 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Friedrich:1975:FES

[Fri75b] H. J. Friedrich. FORTRAN-Standards: Empfehlungen fur die standardisierte Anwendung von FORTRAN fur Programme zur Auswertung medizinischer Datenbestande. Beiheft zum Statistical software newsletter; Heft B1, Universität Giessen, Giessen, Germany, 1975. 58 pp.

Frisch:1980:MRM

[Fri80] Michael J. Frisch. M77 reference manual: Minnesota FORTRAN, 1977 standard version. Technical report, University Computer Center, University of Minnesota, Minneapolis, MN, USA, 1980. v + [338] pp.

Froese:1963:CMT

[Fro63] C. Froese. Computation of mass transport in the ocean from atmospheric pressure data: FORTRAN I program for IBM 1620 computer. Manuscript report series (oceanographic and limnological) / Fisheries Research Board of Canada 163, Pacific Oceanographic Group, Nanaimo, BC, Canada, 1963. 16 pp.

Ford:1977:TPA

[FRS77] B. Ford, J. K. Reid, and B. T. Smith. Three proposed amendments to the draft proposed ANS FORTRAN Standard. ACM SIGNUM Newsletter, 12(1):18–20, March 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fryer:1971:BRB

[Fry71] J. G. Fryer. Book review: Rank Order Probabilities: Two Sample Normal Shift Alternatives, by Roy C. Milton, 1970. The Computer Journal, 14(1):107, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/107.
tif.

Ford:1976:DSN

[FS76] B. Ford and D. K. Sayers. Developing a single numerical algorithm library for different machine ranges. ACM Transactions on Mathematical Software, 2:115–131, 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fletcher:1978:IEI

[FS78] Leroy S. Fletcher and Terry E.

Shoup. Introduction to Engineering: Including Fortran Programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1978. ISBN 0-13-501858-7. xii + 290 pp. LCCN TA145.F53. US$40.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0135018587.

Fletcher:1980:III

[FS80] L. S. Fletcher and T. E. Shoup.

Introduccion a LA Ingenieria: Incluyendo Programacion Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1980. ISBN 0-13-501874-9. ???? pp. LCCN ???? US$12.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0135018749.

Fritsch:1973:AAS

[FSC73] Fred N. Fritsch, R. E. Shafer, and W. P. Crowley. ACM Algorithm 443: Solution of the transcendental equation we^{w} = x [C5]. Comm. ACM, 16(2): 123–124, February 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Ein74].

Fuller:1972:SCS

[Ful72] Samuel H. Fuller. A simulator for computer systems with storage units having rotational delays. Technical report, Stanford University, Digital System Laboratory, Stanford, CA, USA, August 1972.

Fuller:1973:FPA

[Ful73] Robert Ray Fuller. A FORTRAN program to assemble process assembler language on the UNI-VAC 1100 series computer. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1973. vii + 120 pp.

Fultz:1974:EFP

[Ful74] Rick Charles Fultz. Emily: a Fortran program for the calculation of angular correlations of nuclear gamma radiations. Thesis (m.s.), University of Minnesota, Minneapolis, MN, USA, June 1974. i + 31 pp.

Fuller:1977:FPS

[Ful77] William Richard Fuller. FOR-TRAN programming: a supplement for calculus courses. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-90283-X. ca. 200 p. in various pagings pp. LCCN QA303 .F963.

Futrell:1978:RTA

[Fut78] R. Futrell. Remark on “Fortran translation of Algorithm 409: Discrete Chebychev curve fit [E2]”. ACM Transactions on Mathematical Software, 4(1):95, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Sim76b].

Feldman:1978:PFC

[FW78] Stuart I. Feldman and Peter J.

Weinberger. A portable Fortran 77 compiler. Report 1248, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1978. ?? pp.

Gaffney:1977:FSC

[Gaf77] P. W. Gaffney. Fortran subroutines for computing the optimal interpolation formula. Technical Report A.E.R.E. 8781, A.E.R.E. Harwell, Oxford, Oxfordshire, UK, 1977.

Gaffney:1979:FSB

[Gaf79] Pat W. Gaffney. Fortran subroutines for bicubic spline interpolation. ORNL/CSD/TM 67, Dept. of Energy, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1979. v + 31 pp. For sale by the National Technical Information Service.

Gaffney:1980:FSC

[Gaf80] P. W. Gaffney. A Fortran subroutine for computing the optimal estimate of F (X). Technical Report ORNL/CSD-59, Computer Sciences Division At Oak Ridge Nat. Lab., Oak Ridge, TN, USA, 1980.

Gajewski:1966:FPI

[Gaj66] W. Gajewski. A Fortran programme for the identification of hypernuclei. Number 29 in Bulletin — Université libre de Bruxelles, Institut de physique, Faculté des sciences appliquées, Service de physique nucléaire, Physique des particules élémentaires; no. 29 Brussels (Belgium). Université libre. Services de physique nucléaire et de metrologie nucléaire. Bulletin. Presses académiques européennes, Bruxelles, Belgium, 1966. 107 pp.

Galvas:1973:FPP

[Gal73] Michael R. Galvas. FORTRAN program for predicting off-design performance of centrifugal compressors. NASA technical note NASA TN D-7487, National Aeronautics and Space Administration, Washington, DC, USA, 1973. 57 pp. For sale by the National Technical Information Service.

Gales:1975:SFN

[Gal75] L. E. Gales. Structured FOR-TRAN with no preprocessor. ACM SIGPLAN Notices, 10(10):17–24, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Galler:1978:FL

[Gal78] Bernard A. Galler. The FOR-TRAN Language. ACM SIG-PLAN Notices, 13(8):163–164, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garber:1963:AAD

[Gar63] M. J. Garber. Addressing an array Y_{i} in k-dimensions by Fortran for analysis of variance. Comm. ACM, 6(3):100–101, March 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garber:1965:NZS

[Gar65] M. J. Garber. Negative and zero subscripts in Fortran II programming for the IBM 1620. Comm. ACM, 8(8):515–516, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garg:1971:FIE

[Gar71] Satya Prakash Garg. FORTRAN IV and engineering applications. Nem Chand and Bros., New Delhi, India (??), 1971. 551 pp.

Garside:1972:BRBb

[Gar72] M. J. Garside. Book review: Basic Fortran 4, by B. J. Dowding; C. B. Rogers. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):143–144, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.
org/stable/2987328.

Garside:1974:BRB

[Gar74] M. J. Garside. Book review: More Fortran Programs for Economists, by Lucy J. Slater. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):140–141, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.
org/stable/2988018.

Garg:1978:ACB

[Gar78] V. K. Garg. Algorithm 103: Computing the Bessel functions Y_{n}(x + iy) and K_{n}(x + iy). The Computer Journal, 21(3):272–276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/
272.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/273.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
21/Issue_03/tiff/274.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/275.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/
276.tif.

Gavris:1976:ACS

[Gav76] I. B. Gavris. Algorithms for the computation of spherical Bessel functions and associated Legendre functions, and their realization using FORTRAN. (Russian). Vestnik Beloruss. Gos. Univ. Ser I, 3:8–16, 1976.

Gorman:1976:MPL

[GB76] Walter Gorman and Michael Broussard. Minicomputer programming languages. ACM SIG-PLAN Notices, 11(4):4–15, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gillett:1967:FPE

[GC67] Billy E. Gillett and Robert E.

Carlile. Fortran programming for engineers. Technical report, University of Missouri–Rolla, Rolla, MO, USA, 1967. 221 pp.

Gear:1965:HSC

[Gea65] C. W. Gear. High speed compilation of efficient object code. Comm. ACM, 8(8):483–488, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gear:1978:FWL

[Gea78] Charles William (Charles William)

Gear. FORTRAN and WATFIV language manual. Introduction to computers, structured programming, and applications. Science Research Associates, Chicago, IL, USA, 1978. ISBN 0-574-21192-6. 108 pp. LCCN QA76.73 .F25G4.

Gellman:1969:FHC

[Gel69] Leonard Jack Gellman. A FOR-TRAN hardware computer which reduces memory requirements: a thesis. Thesis (m.s.e.), University of Alabama in Huntsville, Huntsville, AL, USA, 1969. ix + 168 l pp.

Genau:1966:USW

[Gen66a] Mary Alice Genau. Underwater shock wave frequency spectrum analysis: IV. FORTRAN IV programs for the IBM-7090 computer. Technical report, Explosions Research Dept., Naval Research Laboratory, Whiteoak, MD, USA, 1966. iii + 15 + [17] pp.

GECIS:1966:GSB

[Gen66b] General Electric Company.Information Systems. GE-400 Series Basic FORTRAN IV Reference manual. General Electric Company, Phoenix, AZ, USA, revised edition, 1966. vii + I-59 pp.

GECISD:1966:TFR

[Gen66c] General Electric Company.Information Systems Division. Time-sharing FORTRAN: reference manual. General Electric, Phoenix, AZ, USA, 1966. viii + 126 pp.

GEC:1967:TFR

[Gen67] General Electric Company. Timesharing FORTRAN: reference manual. General Electric, Bethesda, MD, USA, 1967. viii + 126 pp.

GECISED:1969:GLF

[Gen69] General Electric Company.Information Systems Equipment Division. GE-600 Line FORTRAN IV Reference manual. General Electric Company, Phoenix, AZ, USA, revised edition, 1969. viii + 198 pp.

GECISD:1970:FL

[Gen70a] General Electric Company.Information Service Dept. FORTRAN language. General Electric, Bethesda, MD, USA, 1970. v + 68 pp.

GECISD:1970:IF

[Gen70b] General Electric Company.Information Service Dept. Introduction to FORTRAN. General Electric Company, Bethesda, MD, USA, 1970. iv + 40 pp.

Gentleman:1972:AAC

[Gen72] W. Morven Gentleman. ACM Algorithm 424: Clenshaw–Curtis quadrature [D1]. Comm. ACM, 15(5):353–355, May 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?, ?].

GECISBD:1973:FIS

[Gen73] General Electric Company.Information Services Business Division. FOR-TRAN IV system routines. The Division, Bethesda, MD, USA, revision C edition, 1973. v + 132 pp.

Gentleman:1975:GAC

[Gen75a] J. F. Gentleman. Generation of all NC_{R} combinations by simulating nested Fortran DO loops. Applied Statistics, 24(3):374–376, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gentleman:1975:GAN

[Gen75b] J. F. Gentleman. Generation of all NC_{R} combinations by simulating nested Fortran DO loops. Applied Statistics, 24(3):374–376, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gentleman:1975:SAA

[Gen75c] Jane F. Gentleman. Statistical algorithms: Algorithm AS 88: Generation of all _{N}C_{R} combinations by simulating nested Fortran DO loops. Applied Statistics, 24(3):374–376, September 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/88.

GECISBD:1977:FFM

[Gen77a] General Electric Company.Information Services Business Division. FOR-TRAN 77 (F77): MARK III foreground service reference manual. General Electric, Information Services Business Division, Rockville, MD, USA, revised Mar. 1977 edition, 1977. viii + 253 pp.

GECISBD:1977:FFI

[Gen77b] General Electric Company.Information Services Business Division. FOR-TRAN 77/FORTRAN IV comparisons: MARK III foreground service reference manual. General Electric, Information Services Business Division, Rockville, MD, USA, revised Mar. 1977 edition, 1977. iv + 65 pp.

Gentleman:1978:SAA

[Gen78] Jane F. Gentleman. Statistical algorithms: Algorithm AS 130: Moving statistics for enhanced scatter plots. Applied Statistics, 27(3):354–358, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/130.

GEC:1980:FR

[Gen80a] General Electric Company. FOR-TRAN 77: reference. General Electric, Rockville, MD, USA, revised edition, 1980. 276 pp.

GEC:1980:FSR

[Gen80b] General Electric Company. FOR-TRAN 77: system routines. General Electric, Rockville, MD, USA, 1980. 185 pp.

GEISC:1980:HGF

[Gen80c] General Electric Information Services Company. Handy guide to FORTRAN 77 for students. General Electric Information Services Company, Rockville, MD, USA, 1980. iv + 198 pp.

George:1980:CSL

[Geo80] Alan. George. Computer Solution of Large Sparse Positive Definite Systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1980. ISBN 0-13-165274-5. ???? pp. LCCN ????

Gernoth:1980:W

[Ger80] Brigitte Gernoth. Warteschlangensysteme. R. Oldenbourg, München, Germany, 1980. ISBN 3-486-24541-4.

Galler:1965:IE

[GF65] Bernard A. Galler and Michael J.

Fischer. The iteration element. Comm. ACM, 8(6):349, June 1965.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gritsch:1972:PC

[GG72] R. Gritsch and H. Gritsch. Das Programmieren von Computern. Carl Hanser, München, Germany, 1972. ISBN 3-446-11519-6.

Gary:1972:EFC

[GH72] John Gary and Richard Halgason. An extension of FORTRAN containing finite difference operators. Software—Practice and Experience, 2(4):321–336, October/December 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gleason:1973:EF

[GH73] Gary M. Gleason and Lister Wayne Horn. Essentials of FORTRAN. Rinehart Press, San Francisco, CA, USA, 1973. ISBN 0-03-091400-0. vii + 177 pp. LCCN QA76.73.F25G54.

Gelernter:1960:FCL

[GHG60] Herbert Gelernter, J. R. Hansen, and C. L. Gerberich. A Fortrancompiled list-processing language. Journal of the ACM, 7(2):87–101, April 1960. CODEN JACOAH. ISSN 0004-5411.

Gerad:1965:MTA

[GIB65] J. M. Gerad, I. G. Izsak, and M. P. Barnett. Mechanization of tedious algebra: The Newcomb operators of planetary theory. Comm. ACM, 8(1):27–32, January 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gillett:1966:DLL

[Gil60] Billy E. Gillett. A discussion on the language, logic and theory of Fortran programming. Technical report, University of Missouri– Rolla, Rolla, MO, USA, 1966 (or 1960??). 42 pp.

Gillogly:1970:MFC

[Gil70] James J. Gillogly. MAX: a FOR-TRAN chess player. Technical Report P-4428, Rand Corporation, Santa Monica, CA, USA, 1970. 8 pp.

Gillett:1976:IPA

[Gil76] Will Gillett. An interactive program advising system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(1):335–341, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Gill:1977:DSFa

[Gil77a] Philip E. Gill. The design and structure of a Fortran program library for optimization. Technical report SOL 77-7 (NTIS AD/A-044 905), Dept. of Operations Research, Stanford University; Available through National Technical Information Service, Stanford, CA, USA, 1977. ii + 64 pp.

Gill:1977:DSFb

[Gil77b] Philip E. Gill. The design and structure of a FORTRAN program library for optimization. NPL report NAC 82, National Physical Laboratory, Teddington, Middlesex England, 1977. ii + 64 pp.

Ginsberg:1978:SNC

[Gin78a] M. Ginsberg. Some numerical computational experience with a software interval analysis package and Fortran preprocessor on Control Data Cyber 70 Series computers. Technical Report CS 7808, Department of Computer Science, Southern Methodist University, Dallas, TX, USA, 1978. See Also: Cohn, D. A.; Potter, J. B.; Ginsberg, M., Implementation and Evaluation of Interval Arithmetic Software.

Ginsburg:1978:FI

[Gin78b] Marvin J. Ginsburg. Fortran insights. Data Tactics, Gaithersburg, MD, USA, 1978. various pp.

Galat:1974:EFI

[GKB74] D. L. Galat, T. J. Keefe, and E. P. Bergersen. ECODIV, a Fortran IV program to calculate biological indices and bivariate statistical analysis methodology for evaluation of water quality. Technical publication — Thorne Ecological Institute 11, Thorne Ecological Institute, Boulder, CO, USA, 1974. iii + 70 pp.

Glennie:1962:OEF

[Gle62] A. E. Glennie. Operating experience with FORTRAN. The Computer Journal, 5(2):132– 139, July 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
05/Issue_02/050132.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/132.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/133.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
134.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/135.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
05/Issue_02/tiff/136.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_
02/tiff/137.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/
138.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/139.
tif.

Garber:1964:CFF

[GM64] M. J. Garber and Conrad Miziumski. Curve fitting with format FORTRAN. Comm. ACM, 7(12):721, December 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Good:1973:FPS

[GM73] J. Good and B. A. M. Moon.

FORTRAN — as provided by some major machine manufacturers in 1970. Software—Practice and Experience, 3(1):9–14, January/March 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gill:1979:DSF

[GMPW79] Philip E. Gill, Walter Murray, Susan M. Picken, and Margaret H. Wright. The design and structure of a Fortran program library for optimization. ACM Transactions on Mathematical Software, 5 (3):259–283, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Grame:1975:CPIa

[GO75a] Carl A. Grame and Daniel J. O’Donnell. Computer programming: an individualized course in FORTRAN IV. Cambridge Book, New York, NY, USA, 1975. [6] 182 pp.

Grame:1975:CPIb

[GO75b] Carl A. Grame and Daniel J. O’Donnell. Computer programming; an individualized course in FORTRAN IV, 1975.

Goff:1974:RRB

[Gof74] F. Glenn Goff. RBAD, relative basal area determination: a FOR-TRAN program to determine relative basal area by species and plot from IBP standard format forest service plot tapes. International Biological Programme. Eastern Deciduous Forest Biome EDFB-IBP 74-6, Oak Ridge National Laboratory, Environmental Sciences Division, Oak Ridge, TN, USA, 1974. iii + 31 pp.

Golden:1965:FIP

[Gol65a] James T. Golden. FORTRAN IV: programming and computing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1965. 270 pp.

Golden:1965:FPC

[Gol65b] James Thomas Golden. Fortran 4: programming and computing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1965. 270 pp.

Golde:1966:FII

[Gol66] Hellmut Golde. Fortran II and IV for engineers and scientists. MacMillan Publishing Company, New York, NY, USA, 1966. xvi + 224 pp.

Golde:1968:FIBa

[Gol68a] Hellmut Golde. Fortran information bulletin no. 1. ACM SIG-PLAN Notices, 3(2):1–5, February 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Golde:1968:FIBb

[Gol68b] Hellmut Golde. FORTRAN information bulletin no. 3. ACM SIG-PLAN Notices, 3(7–8):1–19, July– August 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Golden:1976:FIP

[Gol76] James T. Golden. FORTRAN IV:

programmacion y calculo. Urmo, S.A. de Ediciones, ??, Spain, 1976. 271 pp.

Gomez:1979:IFS

[Gom79] J. E. Gomez. An interactive FOR-TRAN structuring aid. In Proceedings of the 4th International Conference on Software Engineering, pages 241–244. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1979.

Good:1964:FIT

[Goo64] Donald I. Good. Fortran II trend-surface program for the IBM 1620. Special distribution publication 14, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1964. 53 pp.

Gorn:1964:FVB

[Gor64] S. Gorn. FORTRAN vs. Basic FORTRAN: a programming language for informational processing on automatic data processing systems. Comm. ACM, 7 (10):591–624, October 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gostin:1980:F

[Gos80] Gary B. Gostin. A factor of F_{17}. Mathematics of Computation, 35 (151):975–976, July 1980. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gottfried:1972:PFI

[Got72] Byron S. Gottfried. Programming with Fortran IV. QPI series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1972. ISBN 0-13-730119-7. 272 pp.

Gottfried:1973:FIP

[Got73] Byron S. Gottfried. FORTRAN IV: programmer’s reference guide. QPI reference guides in computer languages. Quantum, New York, NY, USA, 1973. 15 pp.

Gower:1973:BRB

[Gow73] J. C. Gower. Book review: Fortran Techniques, with Special Reference to Non-Numerical Applications, by A. Colin Day. Journal of the Royal Statistical Society. Series A (General), 136(2):266–267, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345122.

Gower:1975:SAA

[Gow75] J. C. Gower. Statistical algorithms: Algorithm AS 82: The determinant of an orthogonal matrix. Applied Statistics, 24(1):150– 153, March 1975. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/82.

Gay:1973:IDS

[GP73] B. Gay and S. G. Payne. Interactive digital simulation on a small computer. The Computer Journal, 16(2):118–121, May 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
16/Issue_02/160118.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/118.tif;
http://www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_
02/tiff/119.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/
120.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_02/tiff/121.
tif.

Graham:1970:JIJ

[Gra70a] W. R. Graham. JOSTRAN: an interactive JOSS dialect for writing and debugging FORTRAN programs. Research memorandum RM-6248-PR, Rand Corp., Santa Monica, CA, USA, 1970. vii + 8 pp.

GCCCSD:1970:FIG

[Gra70b] Graphic Controls Corporation.Computer Systems Division. FORTRAN IV for the GC-10, a manual for Fortran IV, the algebraic language used with the graphic controls computer time-sharing service. Graphic Controls Corporation, New York, NY, USA, 1970. various pp.

Graham:1979:ICS

[Gra79] Neill Graham. Introduction to computer science: a structured approach. West Publishing Company, St. Paul, MN, USA, 1979. ISBN 0-8299-0187-6. xiii + 511 pp. LCCN QA76.6.G68. US$15.95.

Green:1975:MCC

[Gre75] Fred Murray Green. A method for converting CDC FORTRAN to IBM FORTRAN by digital computer. Thesis (m.s.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1975. vi + 134 pp.

Green:1977:IBF

[Gre77] Leroy Green. Introduction to Basic and Fortran: a Comparative Analysis. Monograph publishing on demand: sponsor series. University Microfilms International, Ann Arbor, MI, USA, 1977. ISBN 0-8357-0289-8. 160 pp. LCCN QA76.73.B3G73. US$10.75. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0835702898.

Greenburg:1979:PUM

[Gre79] Harvey J. Greenburg. Primer and user manual for a Fortran language interactive processor (FLIP). Technical memorandum — Energy Information Administration TM/OA/79-13, Dept. of Energy, Energy Information Administration, Office of Applied Analysis, Washington, DC, USA, March 1979. 49 pp. For sale by the National Technical Information Service.

Gries:1978:ASH

[Gri78] David Gries. ACM SIGPLAN history of programming languages conference ALGOL 60 language summary. ACM SIGPLAN Notices, 13(8):1, August 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Groboillot:1968:IAL

[Gro68a] Jean Louis Groboillot. Initiation au langage Fortran: conversation homme-machine. Monographies des hautes etudes commerciales.

Dunod, Paris, France, 1968. x + 129 pp.

Gross:1968:FP

[Gro68b] P. D. Gross. FORTRAN programming. Douglas United Nuclear, Richland, WA, USA, 1968. iii + 160 pp.

Grosenbaugh:1969:MFR

[Gro69] L. R. Grosenbaugh. More on Fortran random number generators. Comm. ACM, 12(11):639, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Groboillot:1970:IAL

[Gro70] Jean Louis Groboillot. Initiation au langage Fortran, conversation homme-machine. Monographies des hautes etudes commerciales, 1. Dunod, Paris, France, second edition, 1970. x + 131 pp.

Grossberg:1971:FEP

[Gro71] Alan B. Grossberg. FORTRAN for engineering physics: mechanics, data analysis, and heat. Mc-Graw-Hill, New York, NY, USA, 1971. ISBN 0-07-024971-7. ix + 211 pp. LCCN QC20.2 .G76.

Grossberg:1973:FEPa

[Gro73a] Alan B. Grossberg. FORTRAN for engineering physics: electricity, magnetism, and light. Mc-Graw-Hill, New York, NY, USA, 1973. ISBN 0-07-024972-5. viii + 246 pp. LCCN QC37 .G76.

Grossberg:1973:FEPb

[Gro73b] Alan B. Grossberg. Fortran for Engineering Physics: Mechanics, Data Analysis, and Heat. McGraw-Hill, New York, NY, USA, January 1973. ISBN 0-07-024971-7. ???? pp. LCCN ???? US$6.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070249717.

Grossberg:1973:IMA

[Gro73c] Alan B. Grossberg. Instructor’s manual to accompany FORTRAN for engineering physics, electricity, magnetism and light. McGraw-Hill, New York, NY, USA, 1973. ISBN 0-07-024973-3. 83 pp.

Grote:1973:MFC

[Gro73d] Heinz H. Grote. MINIMAX: a FORTRAN computer program to assemble the machine code for the Interdata Model 70 minicomputer on the Control Data 3800. NOAA technical report ERL 269-APCL 28, U.S. Dept. of Commerce, National Oceanic and Atmospheric Administration, Environmental Research Laboratories, Boulder, CO, USA, 1973. iv + 124 pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Gersting:1970:CPA

[GS70] John Marshall Gersting and S. Silverston. Computer programs for the analysis of certain problems in hydrodynamic stability. Technical report, Engineering Research Center, College of Engineering Sciences Arizona State University, Tempe, AZ, USA, 1970. vii + 217 + 16 [i.e. 17] pp.

Glick:1971:AH

[GS71] Norman Glick and Richard Schrader. APL on the Honeywell 635. ACM SIGPLAN Notices, 6(10):23–30, November 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL https://dl.acm.
                                        org/citation.cfm?id=1317455.

Green:1979:CCH

[GS79] P. J. Green and B. W. Silverman. Constructing the convex hull of a set of points in the plane. The Computer Journal, 22(3):262–266, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Guernsey:1973:WBC

[Gue73a] Elwood Guernsey. Workbook in basic computer programming (Fortran IV). Charles C. Thomas, Springfield, IL, USA, 1973. ISBN 0-398-02592-4. 185 pp.

Guernsey:1973:SRB

[Gue73b] Elwood W. Guernsey. Social Research and Basic Computer Programming. Charles C Thomas Pub Ltd, Springfield, Ill., June 1973. ISBN 0-398-02591-6. xxix + 305 pp. LCCN H61 .G821.

Gulledge:1971:DDT

[Gul71] Barbara Jean Vervenne Gulledge. DTFORT: a decision table FOR-TRAN translator. Thesis (m.s.), University of Southwestern Louisiana, Lafayette, LA, USA, 1971. 49 pp.

Gumb:1977:LEE

[Gum77] Raymond D. Gumb. Language extensibility in extended Fortran: data types and pointer variables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):53–54, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gustafsson:1973:FPC

[Gus73] Bengt Gustafsson. A FORTRAN program for calculating “continuous” absorption coefficients of stellar atmospheres. Uppsala Astronomiska Observatoriums Annaler, Bd. 5, no. 6, 1973. Landstingets Verkstader, Uppsala, Sweden, 1973. 31 pp.

Guttmann:1975:PRD

[Gut75] A. J. Guttmann. Programming recursively defined functions in Fortran. Research report 155, University of Newcastle, Dept. of Mathematics, Newcastle, NSW, Australia, 1975. ISBN 0-7259-0175-6. 14 pp.

Guttmann:1976:MDS

[Gut76a] A. J. Guttmann. Multi-dimensional summations in FORTRAN. Software—Practice and Experience, 6 (2):221–224, April/June 1976. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guttmann:1976:PRD

[Gut76b] A. J. Guttmann. Programming recursively defined functions in FORTRAN. International Journal of Computer and Information Sciences, 5(2):111–122, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Guthrie:1979:FBCa

[Gut79a] Sue D. Guthrie. A FORTRAN based computer program to perform goodness to fit testing on empirical data. Technical report, University of Southern Mississippi, Hattiesburg, MS, USA, 1979. 105 pp.

Guthrie:1979:FBCb

[Gut79b] Sue D. Guthrie. A FORTRAN based computer program to perform goodness to fit testing on empirical data. Thesis (m.s.), University of Southern Mississippi, Hattiesburg, MS, USA, 1979. vii + 280 pp.

Gutzmann:1979:SS

[Gut79c] T. Gutzmann. Das Schnelldrucker-Setzprogramm SCRIPTOR. Angewandte Informatik, 6:256–258, 1979. ISSN 0013-5704.

Haag:1965:CFP

[Haa65] James N. Haag. Comprehensive Fortran programming. Hayden Book Co., Rochelle Park, NJ, USA, 1965. x + 246 pp.

Haag:1969:CSFa

[Haa69a] James N. Haag. Comprehensive standard Fortran programming. Hayden computer programming series. Hayden Book Co., Rochelle Park, NJ, USA, 1969. vi + 312 pp.

Haag:1969:CSFb

[Haa69b] James N. Haag. Comprehensive standard Fortran programming: teacher’s guide and answers book. Hayden Book Co., Rochelle Park, NJ, USA, 1969. 32 pp.

Haberman:1972:SAA

[Hab72] S. J. Haberman. Statistical algorithms: Algorithm AS 51: Loglinear fit for contingency tables. Applied Statistics, 21(2):218–225, June 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/51.

Haberman:1973:SAA

[Hab73] S. J. Haberman. Statistical algorithms: Algorithm AS 57: Printing multidimensional tables. Applied Statistics, 22(1):118–126, March 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/57.

Haegemans:1977:AAI

[Hae77] A. Haegemans. An algorithm for the automatic integration over a triangle. Computing, 19(2):179– 187, 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Haines:1965:SCF

[Hai65] L. H. Haines. Serial compilation and the 1401 FORTRAN compiler. IBM Systems Journal, 4(1):73–80, 1965. CODEN IBMSA7. ISSN 0018-8670. URL https://en.

wikipedia.org/wiki/Fortran.

Hall:1965:ESS

[Hal65] Dale Junior Hall. An exploratory study of self-instructional text and lecture systems of teaching Fortran computer programming. Thesis (ed. d.), Indiana University, Bloomington, IN, USA, 1965. 143 pp.

Hall:1969:PFI

[Hal69] A. D. Hall. A portable Fortran IV subset. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1969.

Hall:1971:ASR

[Hal71] Andrew D. Hall, Jr. The Altran system for rational function manipulation — a survey. Comm. ACM, 14(8):517–521, August 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:1972:MMP

[Hal72a] Andrew D. Hall. The M6 macro processor. Computing Science Technical Report 2, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 12, 1972. ii + 13 pp.

Hall:1972:CFC

[Hal72b] D. T. Hall. Correspondence: On FORTRAN complainants. The Computer Journal, 15(3): 273, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
15/Issue_03/150273.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/273.tif. See

[?, Fin72b, Fla72, ?].

Hall:1974:FRE

[Hal74] Andrew D. Hall, Jr. Factored rational expressions in AL-TRAN. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):35–45, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hall:1975:ASP

[Hal75] A. D. Hall. ALTRAN solution to problem #8. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(4):39–40, November 1975. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hamm:1969:SEF

[Ham69] Robert Bryan Hamm. SOTRAN:

an extension of FORTRAN IV. Thesis (m.s.), George Washington University, Washington, DC, USA, 1969. iv + 73 [i.e. 165] pp.

Hamrock:1974:CTM

[Ham74] Josephine Stephanie Hamrock. A comparison of two methods of teaching FORTRAN programming in an undergraduate mathematics class. Thesis (Ph.D.), Purdue University, Lafayette, IN, USA, 1974. ix + 154 pp.

Hamacher:1979:NIM

[Ham79a] H. Hamacher. Numerical investigations on the maximal flow algorithm of Karzanov. Computing, 22(1):17–29, 1979. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Hamala:1979:PEU

[Ham79b] Milan Hamala. Programovanie ekonomickych uloh v jazyku FOR-TRAN. Vydavatelstvo technickej a ekonomickej literatury, Bratislava, Czechoslovakia, 1979. 447 pp.

Hansen:1960:UFC

[Han60] J. R. Hansen. The use of the FORTRAN compiled list processing language. Research report RC-282, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, June 1960.

Hansen:1967:AAF

[Han67] Erik Hansen. Alfons (Algol Fortran nesting sequence). Technical Report Risø-M-564, Atomenergikommissionens Bibliotek, Risø, Denmark, 1967. 10 pp.

Han:1972:THRa

[Han72a] Joseph Ching-Chi Han. Tree height reduction for parallel processing of blocks of Fortran assignment statements. Technical Report UIUCDCS-R-72-493, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, February 1972. 78 pp.

Han:1972:THRb

[Han72b] Joseph Ching-Chi Han. Tree height reduction for parallel processing of blocks of Fortran assignment statements. Thesis (m.s.), Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1972. iv + 73 pp.

Hansen:1974:SBR

[Han74a] Sally Hansen. The SUNY/B RUNT and RUNW FORTRAN manual. Technical report, State University of New York at Buffalo, University Computing Services Press, Buffalo, NY, USA, 1974. ???? pp.

Hanson:1974:STR

[Han74b] David R. Hanson. A simple technique for representing strings in Fortran IV. Comm. ACM, 17(11): 646–647, November 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1975:MVS

[Han75] David R. (David Roy) Hanson.

The manipulation of variablelength string data in Fortran IV. Technical report, Dept. of Computer Science, University of Arizona, Tucson, AZ, USA, 1975. 11 pp.

Hansborough:1978:STG

[Han78] Kay Anderson Hansborough. Second and third generation vector computers and their Fortran processors. Thesis (m. sc.), University of New Mexico, Albuquerque, NM, USA, 1978. v + 143 pp.

Harris:1963:NMU

[Har63] L. Dale Harris. Numerical methods using Fortran programming. Technical report, University of Utah, Electrical Engineering Dept., Salt Lake City, UT, USA, 1963. 237 + 15 + 18. pp.

Harris:1964:FPI

[Har64a] L. Dale Harris. Fortran programming (II and IV). Charles E. Merrill Publishing Co., Columbus, OH, USA, 1964. x + 146 pp.

Harris:1964:NMU

[Har64b] L. Dale Harris. Numerical methods using Fortran. Charles E. Merrill Publishing Co., Columbus, OH, USA, 1964. xi + 244 pp.

Harris:1965:FPI

[Har65a] L. Dale Harris. Fortran programming (II and IV). Charles E. Merrill Publishing Co., Columbus, OH, USA, 1965. x + 146 pp.

Harrison:1965:FHW

[Har65b] Malcolm C. Harrison. Filehandling within FORTRAN. Comm. ACM, 8(8):514–515, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harrison:1965:LEF

[Har65c] Malcolm C. Harrison. Letter to the Editor: Fortran II chaining. Comm. ACM, 8(5):262, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harris:1966:FPC

[Har66a] Richard E. Harris. FORTRAN programs for cryogenic thermometry. Technical report, University of Illinois (Urbana-Champaign campus). Dept. of Physics, Urbana, IL, USA, 1966. 24 pp.

Hartkemeier:1966:FPE

[Har66b] Harry P. (Harry Pelle) Hartkemeier. Fortran programming of electronic computers. Charles E. Merrill Publishing Co., Columbus, OH, USA, 1966. xi + 200 pp.

Harvill:1966:BFP

[Har66c] John B. Harvill. Basic Fortran programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1966. 264 pp.

Harbaugh:1968:FIP

[Har68a] John Warvelle Harbaugh. FOR-TRAN IV program for harmonic trend analysis using double Fourier series and regularly gridded data for the GE 625 computer. Computer contribution 29, University of Kansas, Lawrence, KS, USA, 1968. 30 pp.

Harvill:1968:BFP

[Har68b] John B. Harvill. Basic Fortran programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, revised edition, 1968. ISBN 0-13-061457-2. 268 pp.

Harkins:1969:ICF

[Har69] Paul Houston Harkins. An illustrative comparison of five of the most common computer programming languages: ASSEMBLER, COBOL, FORTRAN, PL/1 and RPG. Thesis (m.b.a.), Drexel University, Philadelphia, PA, USA, 1969. 145 pp.

Hare:1970:IPB

[Har70] Van Court Hare, Jr. Introduction to Programming: a Basic Approach. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, June 1970. ISBN 0-15-543600-7. xii + 436 pp. LCCN QA76.5 .H3574.

Hart:1971:FIP

[Har71] Maurice I. Hart. Fortran IV programming. International Correspondence Schools, Scranton, PA, USA, 1971. v pp.

Hart:1973:FIC

[Har73] Stephen S. Hart. A Fortran IV computer program for determination of airborne camera and scanner variables in geologic remote sensing. Technical Report USGS-GD-73-030, U. S. Geological Survey, Denver, CO, USA, 1973. 1 card pp.

Harvill:1974:FB

[Har74] John B. Harvill. Fortran basico. Coleccion informativa. Ediciones Deusto, Bilbao, Spain, 1974. 234 pp.

Hartnett:1977:SFO

[Har77] Edward T. Hartnett. Star Fortran: an overview of essential characteristics. ACM SIGPLAN Notices, 12(4):57–66, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:1978:GSH

[Har78] Harris Corporation.Computer Systems Division. General specification Harris FORTRAN compiler. Harris Corporation, Needham, MA, USA, 1978. 174 pp.

Harrach:1980:RPR

[Har80] George H. Harrach. RASO: program for renumbering FORTRAN source programs. Open-file report 80-1274, U.S. Dept. of the Interior, Geological Survey, Denver, CO, USA, 1980. i + 60 pp.

Hassitt:1967:DDI

[Has67] A. Hassitt. Data directed inputoutput in FORTRAN. Comm. ACM, 10(1):35–39, January 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Haskell:1978:FPU

[Has78] Richard E. Haskell. FOR-TRAN programming using structured flowcharts. Science Research Associates, Chicago, IL, USA, 1978. ISBN 0-574-21135-7. xii + 280 pp. LCCN QA76.73.F25,H381.

Hatherly:1978:FPR

[Hat78] P. J. Hatherly. A Fortran program for the reduction and plotting of seismic refraction data using the generalised reciprocal method. Geological Survey report GS1976/236, Geological Survey of New South Wales, Dept. of Mines, Sydney, NSW, Australia, 1978. 9 + [20] + [1] pp.

Haavie:1971:SMA

[Håv71] T. Håvie. Some methods for automatic integration and their implementation on the CERN CDC 65/ 6600 computers. Report CERN 71-26, CERN, Geneva, Switzerland, December 1, 1971. URL https:

//cds.cern.ch/record/190218/
files/CERN-71-26.pdf.

Healy:1963:FST

[HB63] M. J. R. Healy and B. P. Bogert. FORTRAN subroutines for time series analysis. Comm. ACM, 6 (1):32–34, January 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harrison:1980:SGE

[HBE80] William L. Harrison, Bruce Broadwell, and Perry Edwards. Study guide for Edwards and Broadwell’s Data processing: computers in action with FORTRAN. Wadsworth, Pacific Grove, CA, USA, 1980. ISBN 0-534-00879-8 (paperback). vi + 244 pp.

Haskell:1976:USF

[HBJ76] R. E. Haskell, D. E. Boddy, and G. A. Jackson. Use of structured flowcharts in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):67–74, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Hung:1978:CPF

[HcL78] Hing Sum Hung and Shiu chang Loh. A course on programming in FORTRAN. Chinese University Press, Hong Kong, 1978. ISBN 962-201-168-3. viii + 444 pp. LCCN QA76.73.F25H87.

Healy:1975:BFI

[HD75] Jeremiah J. Healy and Dalward J.

DeBruzzi. Basic Fortran IV Programming. Addison-Wesley, Reading, MA, USA, revised edition, May 1975. ISBN 0-201-02827-1. x + 294 pp. LCCN QA76.73.F25 H41 1975. US$26.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201028271.

Hoffmann:1978:IMA

[HD78a] Thomas Russell Hoffmann and Gordon Bitter Davis. Instructor’s manual to accompany Davis and Hoffmann: FORTRAN, a structured, disciplined style; based on 1977 American National Standard Fortran and compatible with WAT-FOR, WATFIV,WATFIV-S and MNF Fortran compilers. McGraw-Hill, New York, NY, USA, 1978. 380 pp.

Hull:1978:IPA

[HD78b] T. E. Hull and D. D. F. (David D. F.) Day. An Introduction to Programming and Applications With Fortran. Addison-Wesley, Reading, MA, USA, August 1978. ISBN 0-201-03066-7. xvii + 254 + 12 pp. LCCN QA76.6 .H84. US$16.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201030667.

Healy:1968:BFI

[HDBP68] Jeremiah J. Healy, Dalward J. De-Bruzzi, Charles R. Bauer, and Anthony P. Peluso. Basic Fortran IV programming: self-instructional manual and text. Addison-Wesley, Reading, MA, USA, 1968. 264 pp.

Hanganut:1974:PFC

[HDN74] Marius Hanganut, Ioan Dancea, and Olimpiu Negru. Programe Fortran comentate, in automatica. Number 35 in Colectia Automatica, informatica. Editura tehnica, Bucuresti, Romania, 1974. 312 pp.

Healy:1968:TFV

[Hea68a] M. J. R. Healy. Towards FOR-TRAN VI? The Computer Journal, 11(2):169–172, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_11/Issue_02/110169.sgm.
                                        abs.html;   http://www3.oup.
                                        co.uk/computer_journal/hdb/
                                        Volume_11/Issue_02/tiff/169.
                                        tif;   http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        11/Issue_02/tiff/170.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_11/Issue_
                                        02/tiff/171.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_11/Issue_02/tiff/
                                        172.tif.

Hearne:1968:SDR

[Hea68b] Oliver Kirtley Hearne. Syntactic decomposition of read, write, and format statements for a timeshared FORTRAN interpreter. Thesis (m.s. in engin.), University of Florida, Gainesville, FL, USA, 1968. vi + 113 pp.

Heard:1979:FCC

[Hea79] Steven Roger Heard. A FOR-TRAN cross compiler for a TMS 9900 microcomputer system. Thesis (m.s.), Oklahoma State University, Stillwater, OK, USA, 1979. vii + 85 + pp.

Heck:1963:IBF

[Hec63] G. F. Heck. IBM 1620 Basic FOR-TRAN Manual. Bethlehem Steel Co., Research Department, Bethlehem, PA, USA, 1963. various pp.

Hedrick:1977:AIO

[Hed77] G. E. Hedrick. ALGOL68 instruction at Oklahoma State University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):16–20, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Heising:1963:F

[Hei63] W. P. Heising. Fortran. Comm.

ACM, 6(3):85–86, March 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heising:1964:HSF

[Hei64] W. P. Heising. History and summary of FORTRAN standardization development for the ASA. Comm. ACM, 7(10):590, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also final standard [Ame66a].

Heising:1966:HSF

[Hei66] W. P. Heising. History and summary of Fortran standardization development for the ASA. Comm. ACM, 7(??):590–625, ?? 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also final standard [Ame66a].

Heiner:1970:FIP

[Hei70] Lawrence E. Heiner. FORTRAN IV program for processing geochemical sediment data. M.I.R.L. report 23, Mineral Industry Research Laboratory, University of Alaska, College, Alaska, 1970. 10 + [9] pp.

Heinselman:1972:CDC

[Hei72a] Russell Craig Heinselman. Computerized detection and correction of spelling errors in FORTRAN programs. Thesis (m.s.), University of Minnesota, Minneapolis, MN, USA, 1972. 76 pp.

Heitzman:1972:SGF

[Hei72b] William R. Heitzman. A study guide for Fortran programming: the basic elements. Technical report, California State University, Fullerton, CA, USA, 1972. 61 pp.

Heidorn:1974:EVH

[Hei74] G. E. Heidorn. English as a very high level language for simulation programming. ACM SIGPLAN Notices, 9(4):91–100, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hellwig:1963:SF

[Hel63] Jessica Hellwig. Subroutines in FORTRAN. Memorandum (Massachusetts Institute of Technology. Computation Center) CC-200-2, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, February 1963. 17 pp.

Hempkins:1970:FIP

[Hem70] W. B. Hempkins. A Fortran IV program for two-dimensional autocorrelation analysis of geologic and remotely-sensed data. Report 21, Dept. of Geological Sciences, Northwestern University, Evanston, IL, USA, 1970. iv + 54 pp.

Henney:1967:RFI

[Hen67] Dagmar Henney. Reviews: Fortran IV. Programming and Computing, by James T. Golden. American Mathematical Monthly, 74(7): 886, August/September 1967. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Herrmann:1964:SFP

[Her64] Walter Herrmann. SAVE 1: FOR-TRAN program for the semiautomatic calculation of wave propagation by the method of characteristics. M.I.T. Aeroelastic and Structures Research Laboratory. Technical report ARRL TR 1018, M.I.T. Aeroelastic and Structures Research Laboratory, Cambridge, MA, USA, 1964. 34 pp.

Hershey:1969:FIP

[Her69] Allen V. Hershey. FORTRAN IV programming for cartography and typography. Technical report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, 1969. iv + 19 + 11 + [66] pp.

Hershey:1970:PRF

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:1971:APF

[Her71] Allen V. Hershey. Auxiliary programming for the Fortran typographic system. Technical Report TR-2645, U.S. Naval Weapons Laboratory, Dahlgren, VA, USA, October 1971.

Heres:1972:ATF

[Her72a] Celestino Heres. Algebra and trigonometry with FORTRAN programming. Thesis (ed. d.), Lawrence University, Appleton, WI, USA, 1972. ???? pp. OCLC catalog has Laurence University, but no such institution seems to exist.

Hershey:1972:FPS

[Her72b] Allen V. Hershey. FORTRAN programming for surface wave trains. Technical Report TR-2714, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1972.

Hershkowitz:1974:UCR

[Her74] Sandra J. Weiner Hershkowitz.

The use of computer-generated repeatable examinations in introductory FORTRAN classes. Thesis (m.a.), St. Bonaventure University, Saint Bonaventure, NY ,USA, 1974. [5] 42 [4] pp.

Herschel:1978:FSD

[Her78] Rudolf Herschel. FORTRAN:

systemat. Darst. fur d. Anwender. Reihe Datenverarbeitung. R. Oldenbourg, München, Germany, 1978. ISBN 3-486-22431-X. 135 pp.

HP:1971:HFP

[Hew71] Hewlett–Packard Company. HP FORTRAN programmer’s reference manual. Hewlett Packard, Cupertino, CA, USA, third edition, 1971. 108 pp.

HP:1974:FIR

[Hew74] Hewlett–Packard Company. FOR-TRAN IV reference manual. Hewlett–Packard, Cupertino, CA, USA, 1974. 160 pp.

HP:1976:FRMb

[Hew76a] Hewlett–Packard Company. FOR-TRAN reference manual. Hewlett Packard, Santa Clara, CA, USA, 1976. various pp.

HP:1976:FRMa

[Hew76b] Hewlett–Packard Company. FOR-TRAN/ 3000 reference manual: HP 3000 computer systems. Hewlett Packard, Santa Clara, CA, USA, 1976. various pp.

HP:1979:FPG

[Hew79a] Hewlett–Packard Company. FOR-TRAN pocket guide. Hewlett– Packard, Santa Clara, CA, USA, third edition, 1979. 26 pp.

HP:1979:FRM

[Hew79b] Hewlett–Packard Company. FOR-TRAN/ 3000 reference manual. Hewlett–Packard, Santa Clara, CA, USA, 1979. ???? pp.

HP:1979:RFI

[Hew79c] Hewlett–Packard Company. RTE FORTRAN IV reference manual. The Company, Cupertino, CA, USA, sixth edition, 1979. ???? pp.

HP:1980:RFR

[Hew80a] Hewlett–Packard Company. RTE FORTRAN 4X reference manual. Hewlett–Packard, Cupertino, CA, USA, 1980. various pp.

HP:1980:RFI

[Hew80b] Hewlett–Packard Company. RTE Fortran IV independent study course, 1980.

Hurst:1978:CFS

[HF78] Barbara Joan Hurst and Jeremy William Firth. CCAE FORTRAN self learner. CCAE, Canberra, Australia, third edition, 1978. ISBN 0-85889-093-3. 126 + [4] pp.

Heiner:1966:FIT

[HG66] Lawrence E. Heiner and Stephen P. Geller. Fortran IV trend-surface program for the IBM 360 model 40 computer. M.I.R.L. Report 9, Mineral Industry Research Laboratory, University of Alaska, College, AK, USA, 1966. 69 pp.

Holt:1977:FSP

[HH77a] R. C. (Richard C.) Holt and J. N. P. Hume. Fundamentals of structured programming using FORTRAN with SF/k and WATFIV-S. Reston Publishing Co., Inc., Reston, VA, USA, 1977. ISBN 0-87909-303-X, 0-87909-302-1 (paperback). xiii + 349 pp. LCCN QA76.73.F25 .H64.

Hume:1977:EPS

[HH77b] James N. P. Hume and R. C.

(Richard C.) Holt. Elements de programmation structurée en Fortran avec SF/ k et WATFIV-S. École polytechnique de Montréal, Montréal, PQ, Canada, 1977. xii + 350 pp.

Hume:1978:EPS

[HH78] James N. P. Hume and R. C.

(Richard C.) Holt. Elements de programmation structurée en FORTRAN avec SF/ k et WATFIV-S. École polytechnique de Montréal, Montréal, PQ, Canada, ed. corr edition, 1978. xii + 350 pp.

Hennell:1979:ETN

[HH79a] M. A. Hennell and D. Hedley. Experimental testbed for numerical software — 2. Algol 68. The Computer Journal, 22(1):53–56, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Hume:1979:PFS

[HH79b] J. N. P. Hume and R. C. (Richard C.) Holt. Programming Fortran 77: a Structured Approach. Reston Publishing Co., Inc., Reston, VA, USA, June 1979. ISBN 0-8359-5671-7. 340 pp. LCCN QA76.73.F25 H86. US$32.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0835956717.

Holt:1980:FWP

[HH80] R. C. (Richard C.) Holt and J. N. P. (James Nairn Patterson) Hume. FORTRAN WATFIV-S: programmation structurée et exercices. Collection Systemes d’information. G. Morin, Chicoutimi, Québec, 1980. xvii + 448 pp.

Higman:1972:PVS

[Hig72] B. Higman. Programmiersprachen, eine vergleichende Studie. Carl Hanser, München, Germany, 1972. ISBN 3-446-10547-6.

Higgins:1975:SFT

[Hig75] Donald S. Higgins. A structured FORTRAN translator. ACM SIGPLAN Notices, 10(2):42–48, February 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Higbie:1978:SFC

[Hig78] L. Higbie. Speeding up FORTRAN (CFT) programs on the CRAY–1. Pub. 2240207, Cray Research, Inc., Minneapolis, MN, 1978.

Higbee:1979:VCF

[Hig79a] L. Higbee. Vectorization and conversion of Fortran programs for the Cray-1 CFT compiler. Publication 2240207, Cray Research, Inc., Minneapolis, MN, June 1979.

Higgins:1979:PGA

[Hig79b] J. P. (James Patrick) Higgins. Programmer’s guide to the AFID system: an ANS FORTRAN 66 system for free-format data. Division of Entomology report; no. 5. Australia, Canberra, Australia, 1979. ISBN 0-643-02298-8. 40 pp.

Hill:1969:FAT

[Hil69] Edward Burlingame Hill. A FOR-TRAN to ALGOL translator. Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1969. 271 pp.

Hillstrom:1970:PSF

[Hil70] K. E. Hillstrom. Performance statistics for the FORTRAN IV (H) and PL/I (version 5) libraries in the IBM OS/360 Release 18. Report ANL-7666, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1970.

Hill:1971:FFA

[Hil71]  I. D. Hill.  Faults in func-

tions, in ALGOL and FORTRAN. The Computer Journal, 14(3): 315–316, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_03/140315.sgm.abs.

html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/315.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/316.tif.  See notes

[SS72].

Hill:1973:BRF

[Hil73] I. D. Hill. Book reviews: Fortran Techniques, With Special Reference to Non-Numerical Applications, by A. Colin Day. Applied Statistics, 22(3):404, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1979:BRC

[Hil79a] I. D. Hill. Book reviews: Compatible Fortran, by A. Colin Day. Applied Statistics, 28(3):304, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1979:DFP

[Hil79b] M. O. Hill. Decorana: a FOR-TRAN program for detrended correspondence analysis and reciprocal averaging. Technical report, Cornell University, Ithaca, NY, USA, 1979.

Hill:1979:TFP

[Hil79c] M. O. Hill. TWINSPAN: a FOR-TRAN program for arranging multivariate data in an ordered twoway table by classification of the individuals and attributes. Cornell ecology programs series, Cornell University, Section of Ecology and Systematics, Ithaca, NY, USA, 1979. iv + 90 pp.

Hinchman:1976:SME

[Hin76] Richard Keith Hinchman. Simulation of minicomputer environments for real-time FORTRAN software development. Thesis (m.s.

— chemical engineering), University of Arizona, Tucson, AZ, USA, 1976. ix + 148 pp.

Hirsch:1973:EFI

[Hir73] Seymour C. Hirsch. Essentials of FORTRAN IV. Reston Publishing Co., Inc., Reston, VA, USA, 1973. ISBN 0-87909-255-6. xii + 291 pp. LCCN QA76.73.F25 H57.

Hisgen:1975:DIW

[His75] Andrew L. Hisgen. The design and implementation of WATFIV/WM: a FORTRAN dialect for structured programming. Thesis (honors), College of William and Mary, Williamsburg, VA, USA, 1975. ii + 66 + [11] pp.

Hall:1974:APS

[HJ74] A. D. Hall and S. C. Johnson.

ALTRAN programs for SIGSAM problem #6. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(2):12–36, May 1974. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Hurst:1972:AAG

[HK72] Rex L. Hurst and Robert E. Knop.

ACM Algorithm 425: Generation of random correlated normal variables [G5]. Comm. ACM, 15(5):355–357, May 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?].

Hollman:1975:FAM

[HK75] Kenneth W. Hollman and C. Bruce Kavan. FORTRAN applications in management education. Technical report, Bureau of Business and Economic Research, University of Mississippi, Oxford, MS, USA, 1975. vii + 228 pp.

Hayes:1972:IMN

[HKK72] John K. Hayes, David K. Kahaner, and Richard G. Kellner. An improved method for numerical conformal mapping. Mathematics of Computation, 26(118):327–334, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hammersley:1970:CTF

[HL70] P. Hammersley and J. Larmouth. Correspondence: Towards Fortran VI? The Computer Journal, 13(2):220, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_02/tiff/220.tif. See

[?, Wel70a].

Hazel:1973:SCF

[HLS73] P. Hazel, J. Larmouth, and A. Stoneley. Some comments on FORTRAN systems. Software— Practice and Experience, 3(2):185– 187, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Herrmann:1962:WFP

[HM62a] Walter Herrmann and Evelyn Mack. WAVE I: FORTRAN program for calculation of onedimensional wave propagation. ASRL report 1004, Aeroelastic and Structures Research Laboratory, Department of Aeronautics and Astronautics, Massachusetts Institute of Technology, Cambridge, MA, USA, 1962. iv + 29 pp.

Herrmann:1962:WII

[HM62b] Walter Herrmann and Evelyn Mack. WAVE II: an improved FORTRAN program for calculation of one-dimensional wave propagation. ASRL Report 1005, Aeroelastic and Structures Research Laboratory, Department of Aeronautics and Astronautics, Massachusetts Institute of Technology, Cambridge, MA, USA, 1962. vii + 59 pp.

Herrmann:1962:WIF

[HM62c] Walter Herrmann and Evelyn Mack. WAVE II: FORTRAN program for calculation of onedimensional wave propagation, appendix: program listing. ASRL Report 1005, Aeroelastic and Structures Research Laboratory, Department of Aeronautics and Astronautics, Massachusetts Institute of Technology, Cambridge, MA, USA, 1962. 38 pp.

Herrmann:1964:WVF

[HM64] Walter Herrmann and Evelyn Mack. WAVE V: a FORTRAN program for calculation of onedimensional wave propagation including material strength. M.I.T. Aeroelastic and Structures Research Laboratory. Technical report ASRL TR 1020, M.I.T. Aeroelastic and Structures Research Laboratory, Cambridge, MA, USA, 1964. 95 pp.

Hanson:1975:FCC

[HM75] Allan Hanson and Kurt Maly. A first course in computer science: What it should be and why. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):95–101, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Hughes:1977:SAP

[HM77] Joan K. Hughes and Jay I. Michtom. A Structured Approach to Programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1977. ISBN 0-13-854356-9. viii + 264 pp. LCCN QA76.6 .H83.

Hill:1980:PLS

[HM80] I. D. Hill and B. L. Meek. Programming Language Standardisation. John Wiley and Sons, New York, London, Sydney, 1980. ISBN 0-85312-188-5. 261 pp. LCCN QA76.7 .P75.

Hughes:1958:FAC

[HN58] R. A. Hughes and R. Nutt. FOR-TRAN Automatic Coding System for the IBM 704 Data Processing System. ????, ????, 1958. ???? pp. LCCN ????

Hsu:1970:SGP

[HN70] Ju-Tung Hsu and Albert Newhouse. Strachey’s general purpose macrogenerator in FOR-TRAN. Technical report, University of Houston, Houston, TX, USA, 1970. 43 pp.

Herrmann:1964:RFP

[HO64] Walter Herrmann and Margot O’Brien. RAVE I: a FOR-TRAN program for the calculation of two-dimensional wave propagation. M.I.T. Aeroelastic and Structures Research Laboratory. Technical report ASRL TR 1021, M.I.T. Aeroelastic and Structures Research Laboratory, Cambridge, MA, USA, 1964. 74 pp.

Ho:1973:RFP

[Ho73] Ka Keung Ho. Restructuring FORTRAN programs for readability, debugging and maintenance: research project. Thesis (m.s. in electrical engineering), University of California, Berkeley, Berkeley, CA, USA, December 1973. 33 pp.

Hoaglin:1972:ALO

[Hoa72] D. C. Hoaglin. An analysis of the loop optimization scores in Knuth’s ‘empirical study of FOR-TRAN programs’. Software— Practice and Experience, 3(2):161– 169, April/June 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hoaglin:1973:ALO

[Hoa73] David C. Hoaglin. An analysis of the loop optimization scores in Knuth’s ‘empirical study of FOR-TRAN programs’. Software— Practice and Experience, 3(2):161– 169, April/June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hobson:1967:FIP

[Hob67] R. D. (Richard David) Hobson. FORTRAN IV programs to determine surface roughness in topography for the CDC 3400 computer. Computer contribution 14, University of Kansas, Lawrence, KS, USA, 1967. 28 pp.

Hogman:1970:CPS

[Hog70] Walter J. Hogman. A computer program for stock analysis in FOR-TRAN IV, Univac 1108. Transactions of the American Fisheries Society, 99(2):426–427, April 1970. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hogge:1972:PSL

[Hog72] James H. Hogge. Programming the statistical library. Auerbach Publishers, Philadelphia, 1972. ISBN 0-87769-136-3. viii + 208 pp. LCCN QA276.4 .H63.

Hojberg:1969:HCF

[Høj69] K. Sø Højberg. Hybrid computer FORTRAN routines. Technical Report Risø-M-957, Atomenergikommissionens Bibliotek, Risø, Roskilde, Denmark, 1969. 14 pp.

Hojberg:1970:MFH

[Høj70] K. Sø Højberg. Main FORTRAN hybrid routines. Risø-M-1234 Risø-M; 1234. Danish Energy Commission, Research Establishment, Electronics Dept., Risø, Denmark, 1970. ISBN 87-550-0030-4. 9 pp. LCCN QC770.D42 no. 1234.

Holy:1967:FIP

[Hol67] Z. J. Holy. FORTRAN IV programmes for computation of temperature and thermoelastic stress in a homogeneous spherical fuel element due to axisymmetric heat transfer variation over the surface. Technical Report AAEC/TM 366, Research Establishment, Australian Atomic Energy Commission, Lucas Heights, NSW, Australia, 1967. 18 + [21] pp.

Holy:1968:FIP

[Hol68] Z. J. Holy. FORTRAN IV programs for computation of temperature and thermoelastic stress in a homogeneous spherical fuel element due to arbitrary heat transfer variation over the surface. Technical Report AAEC/TM470, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1968. 26 pp.

Holden:1970:IFI

[Hol70] Herbert L. Holden. Introduction to Fortran IV. MacMillan Publishing Company, New York, NY, USA, October 1970. ISBN 0-02-355990-X. viii + 134 pp. LCCN ???? US$16.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
002355990X.

Holden:1971:IFI

[Hol71] Herbert L. Holden. Introduction to FORTRAN IV. MacMillan Publishing Company, New York, NY, USA, 1971. viii + 134 pp.

Hollocks:1972:FB

[Hol72] Brian Hollocks. Fortran basics. IPC Electrical-Electronic Press, London, UK, 1972. 70 pp.

Holsclaw:1977:IFG

[Hol77] Lynden Edward Holsclaw. Interactive FORTRAN graphics plotting routines for the Digital Equipment Corporation GT-44 system. Thesis (m.s. — electrical engineering), University of Arizona, Tucson, AZ, USA, 1977. viii + 83 pp.

Holland:1980:MFP

[Hol80] F. D. Holland. MESS: FORTRAN program for numerical solution of single commodity multi-market equilibrium problems with nonlinear supply and demand functions and flow distortions. Station bulletin 296, Dept. of Agricultural Economics, Agricultural Experiment Station Purdue University, West Lafayette, IN, USA, 1980. 79 pp.

Honeywell:1970:FUM

[Hon70a] Honeywell, Inc. Fortran, user’s manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1970. 189 pp.

HoneywellIS:1970:TDH

[Hon70b] Honeywell Information Systems, Inc. Time-sharing documentation: H1640 System FORTRAN user’s manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1970. 222 pp.

HoneywellIS:1971:SFRa

[Hon71a] Honeywell Information Systems, Inc. Series 1640 Fortran random files. Honeywell Information Systems, Inc., Waltham, MA, USA, 1971. ca. 89 pp.

HoneywellIS:1971:SFRb

[Hon71b] Honeywell Information Systems, Inc. Series 1640 Fortran runtime debugging. Honeywell Information Systems, Inc., Waltham, MA, USA, 1971. 22 pp.

HoneywellIS:1972:FSG

[Hon72a] Honeywell Information Systems, Inc. FORTRAN: study guide. Honeywell Information Systems, Inc., Waltham, MA, USA, 1972. vi + 136 pp.

HoneywellIIS:1972:FSG

[Hon72b] Honeywell Institute of Information Sciences. Fortran, a study guide. Honeywell Information Systems, Inc., Waltham, MA, USA, second edition, 1972. various pp.

HoneywellIS:1973:F

[Hon73a] Honeywell Information Systems, Inc. Fortran. Honeywell Information Systems, Inc., Waltham, MA, USA, 1973. 313. pp.

HoneywellIS:1973:FIV

[Hon73b] Honeywell Information Systems, Inc. FORTRAN IV: video-assisted learning course, 1973.

HoneywellIS:1973:SFS

[Hon73c] Honeywell Information Systems, Inc. Series 600/6000 Fortran software. Honeywell Information Systems, Inc., Waltham, MA, USA, 1973. various pp.

Honeywell:1973:TFS

[Hon73d] Honeywell Information Systems, Inc. Time-sharing FORTRAN, series 600/6000 software. Honeywell Information Systems, Inc., Waltham, MA, USA, 1973. xii + 220 pp.

HoneywellIS:1974:FI

[Hon74] Honeywell Information Systems, Inc. Fortran IV. Honeywell Information Systems, Inc., Waltham, MA, USA, 1974. ???? pp.

Honeywell:1975:FPG

[Hon75a] Honeywell, Inc. FORTRAN pocket guide: series 60 (level 66)/6000. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. 36 pp.

Honeywell:1975:FS

[Hon75b] Honeywell, Inc. Fortran: series 60(level 66)/6000. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. 266 pp.

HoneywellIS:1975:F

[Hon75c] Honeywell Information Systems, Inc. Fortran. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. 301. pp.

HoneywellIS:1975:FRM

[Hon75d] Honeywell Information Systems, Inc. FORTRAN reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. various pp.

HoneywellIS:1975:FSL

[Hon75e] Honeywell Information Systems, Inc. FORTRAN subroutine libraries. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. various pp.

Honeywell:1976:SGC

[Hon76] Honeywell, Inc.Traffic Management Center. Second Generation Control FORTRAN software: Urban Traffic Control System (UTCS), software support project. Federal Highway Administration; Available through the National Technical Information Service, Washington, DC. USA, 1976.

HoneywellIS:1977:LSF

[Hon77a] Honeywell Information Systems, Inc. Level 66 software: FOR-TRAN subroutines libraries. Honeywell Information Systems, Inc., Waltham, MA, USA, 1977. various pp.

HoneywellIS:1977:SLM

[Hon77b] Honeywell Information Systems, Inc. Series 60 (level 68) Multics FORTRAN. Honeywell Information Systems, Inc., Waltham, MA, USA, revised edition, 1977. vii + [200] pp.

HoneywellIS:1979:FRM

[Hon79a] Honeywell Information Systems, Inc. Fortran reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1979. various pp.

HoneywellIS:1979:SLM

[Hon79b] Honeywell Information Systems, Inc. Series 60 (level 68) Multics FORTRAN users’ guide. Honeywell Information Systems, Inc., Waltham, MA, USA, 1979. vii + [100] pp.

Horsfall:1965:ASA

[Hor65] C. T. Horsfall. Aerotriangulation strip adjustment using Fortran and the IBM-1620 computer. Technical report, U. S. Dept. of Commerce; Coast and Geodetic Survey, Washington, DC, USA, 1965. 14 + 2.5 pp.

Horton:1968:ASD

[Hor68] David L. Horton. An algorithm for the sequential decomposition of a subset of FORTRAN IV. Thesis (m.s.), Department of Computer Science, Mississippi State University, Mississippi State, MS, USA, 1968. 45 pp.

Horowitz:1972:S

[Hor72] E. Horowitz. SAC-1. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):22–24, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Horowitz:1975:FCI

[Hor75] E. Horowitz. Fortran can it be structured — should it be? Computer, 8(6):30–37, June 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:1962:IMD

[Hou62] Roger Wolcott Hough. The IBM management decision-making game: a recoding into Fortran for the IBM 1620 data processing system. Thesis (m.s.), University of Oregon, Dept. of Marketing and Industrial Management, Eugene, OR, USA, 1962. 34 + 25 pp.

Housden:1971:PSF

[Hou71] R. J. W. Housden. Phrase structures in FORTRAN. The Computer Journal, 14(3):224– 228, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_03/140224.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/224.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_
03/tiff/225.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/
226.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_03/tiff/227.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_03/tiff/228.tif.

Holberton:1974:NFT

[HP74] Frances E. Holberton and Elizabeth G. Parker. NBS FORTRAN test programs. NBS special publication 399, National Bureau of Standards, Washington, DC. USA, 1974. various pp. For sale by the Supt. of Docs. U.S. Govt. Print. Off.

Hendricks:1973:WFF

[HPB73] Robert C. Hendricks, Ildiko C. Peller, and Anne K. Baron. WASP: a flexible FORTRAN IV computer code for calculating water and steam properties. NASA technical note NASA TN D-7391, National Aeronautics and Space Administration; available from National Technical Information Service, Washington, DC, USA, 1973. iii + 116 pp.

Haring:1979:USJ

[HPLG79] G. Haring, R. Posch, C. Leonhardt, and G. Gell. Use of a synthetic jobstream in performance evaluation. The Computer Journal, 22(3):209–219, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Hughes:1977:APT

[HPR77] Charles Edward Hughes, Charles P. Pfleeger, and Lawrence L. Rose. Advanced programming techniques: a second course in programming using FORTRAN. John Wiley and Sons, New York, London, Sydney, preprint edition, 1977. 3 + 810 + 15 pp.

Hughes:1978:APT

[HPR78] Charles Edward Hughes, Charles P. Pfleeger, and Lawrence L. Rose. Advanced Programming Techniques — a Second Course in Programming Using Fortran. John Wiley and Sons, New York, London, Sydney, April 1978. ISBN 0-471-02611-5. xiii + 287 pp. LCCN QA76.73 F25 H81 1978. US$44.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471026115.

Hall:1973:IFV

[HR73] A. D. Hall, Jr. and B. G. Ryder.

Installation of the FORTRAN verifier. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973. ?? pp.

Hammond:1976:IFI

[HRH76] Robert H. Hammond, William B.

Rogers, and Byard Houck, Jr. Introduction to Fortran IV. Mc-Graw-Hill, New York, NY, USA, 1976. ISBN 0-07-025895-3. viii + 133 pp. LCCN QA76.73.F25 H36. US$34.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070258953.

Hammond:1978:IFI

[HRH78] Robert H. Hammond, William B.

Rogers, and Byard Houck. Introduction to FORTRAN IV. Mc-Graw-Hill, New York, NY, USA, second edition, 1978. ISBN 0-07-025897-X. x + 198 pp.

Hendry:1969:TFV

[HS69] D. F. Hendry and P. A. Samet.

Towards FORTRAN VI? Part 2. FORTRAN in the modern world. The Computer Journal, 12(3): 218–220, August 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        12/Issue_03/120218.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        12/Issue_03/tiff/218.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_12/Issue_
                                        03/tiff/219.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_12/Issue_03/tiff/
                                        220.tif.

Hughes:1969:PI

[Hug69] Joan Kirkby Hughes. Programming the IBM 1130. John Wiley and Sons, New York, London, Sydney, 1969. ISBN 0-471-42040-9. ix + 512 pp. LCCN QA76.8 .I125.

Hughes:1977:FAG

[Hug77] Christopher J. Hughes. The FOR-TRAN assembler: a guide to code generation for MCP. Technical report 131M, State University of New York at Buffalo, Dept. of Computer Science, Buffalo, NY, USA, 1977. 25 pp.

Hughes:1978:FVV

[Hug78] Christopher J. Hughes. Fortran VM: a virtual memory system for a FORTRAN machine on the B1700. Thesis (m.s.), State University of New York at Buffalo, Dept. of Computer Science, Buffalo, NY, USA, 1978. 30 pp.

Huitson:1965:BRBa

[Hui65] A. Huitson. Book review: Computer Language — An Autoinstructional Introduction to Fortran, by H. L. Colman; C. Smallwood. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):86–87, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.
org/stable/2987246.

Hull:1973:WYB

[Hul73] T. E. Hull. Would you believe structured Fortran? ACM SIGNUM Newsletter, 8(4):13–16, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hunt:1974:RFP

[Hun74] Michael F. Hunt. RANDHIND:

a FORTRAN program that generates two-voiced exercises according to the rules stated in Paul Hindemith’s The craft of musical composition (book II). Thesis, Washington University, St. Louis, MO, USA, 1974. 75 pp.

Hunt:1976:UGG

[Hun76] William A. Hunt. Users’ guide to GPAK: a suite of FORTRAN subprograms for computational geometry and graphics: version V02A. Computer-related development memo 9, Production Automation Project, College of Engineering and Applied Science, University of Rochester, Rochester, NY, USA, 1976. 41 pp. Reproduced by National Technical Information Service.

Hurst:1977:CFS

[Hur77] Barbara Joan Hurst. CCAE FOR-TRAN self learner. Canberra College of Advanced Education, Canberra, Australia, second edition, 1977. ISBN 0-85889-058-5. 116 pp.

Huszar:1976:ODB

[Hus76] Anders Huszar. Olika datorprogram for beräkning av solpositioner samt strålningsenergi for varierande sluttningar. Rapporter och uppsatser (Research notes) 74, Skogshögskolan, Institutionen for skogsforyngring, Stockholm, Sweden, 1976. 81. pp.

Hutty:1980:FS

[Hut80] Roger Hutty. Fortran for students. Macmillan computer science series. MacMillan Publishing Company, New York, NY, USA, 1980. ISBN 0-333-25331-0. x + 99 pp.

Huybrechts:1977:DSS

[Huy77] M. Huybrechts. DYNOSOR: a set of subroutines for dynamic memory organization in Fortran programs. ACM SIGPLAN Notices, 12(4):67–74, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hallopeau:1966:MPP

[HV66] B. Hallopeau and J. Vignes. Manuel pratique de programmation Fortran. Publications de l’Institut français du petrole. Societé des Editions Technip, Paris, France, 1966. 132 pp.

Hathaway:1974:CEM

[HV74] P. J. Hathaway and D. Van Vliet. Correspondence: An efficient method for reading and writing subscripted arrays using FORTRAN IV. The Computer Journal, 17(2):190–191, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.
oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/
190.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/191.

tif. See correspondence [Art75, ?].

Harbaugh:1967:FIP

[HW67] John Warvelle Harbaugh and Warren J. Wahlstedt. Fortran IV program for mathematical simulation of marine sedimentation with IBM 7040 or 7094 computers. Computer contribution 9, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 40 pp.

Hald:1972:RF

[HW72] Jens Hald and Alan Wessel. RC 4000 FORTRAN. Regnecentralen, Copenhagen, Denmark, 1972. ISBN 87-7557-000-9. 58 + [35] pp. LCCN QA76.73.F25H34.

Hetherington:1975:DFP

[HW75] E. L. R. Hetherington and Nigel Robert Wood. DOSE — a FORTRAN program for the calculation of radiation dose from radiopharmaceuticals. AAEC/E328. Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1975. ISBN 0-642-99651-2. 4 + [13] pp.

HECUS:1966:FIO

[Hyd66a] Hydrologic Engineering Center (U.S.). Fortran input/output conversion: Hydrologic Engineering Center computer program 72-J2-L208. Technical report, U.S. Army Engineer District, Sacramento, CA, USA, 1966. 35 pp.

HECUS:1966:FSI

[Hyd66b] Hydrologic Engineering Center (U.S.). FORTRAN source inventory and renumbering: Hydrologic engineering center computer program 72-J2-L205. Technical report, Dept. of the Army, Sacramento District, Corps of Engineers, Sacramento, CA, USA, 1966. 68 pp.

ISA:1978:ICS

[IA78] Instrument Society of America and American National Standards Institute. Industrial computer system FORTRAN procedures for file access and the control of file contention: standard. Instrument Society of America, Pittsburgh, PA, USA, 1978. iv + 7 pp.

Indrea:1980:FAE

[IA80] E. Indrea and N. Aldea. Fourier analysis of EXAFS data, a self-contained Fortran programpackage. Computer Physics Communications, 21(1):91–96, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0010465580900788.

IRE:1957:WJC

[IAAA57] IRE, AIEE, ACM, and AFIPS, editors. Western Joint Computer Conference. ????, New York, NY, USA, 1957.

IBM:1954:SIM

[IBM54] IBM. Specifications for the IBM Mathematical FORmula TRANSlating system. Preliminary report, IBM Corp., Programming Research Group, Applied Sciences Division, New York, NY, USA, November 10, 1954.

IBM:1956:PRM

[IBM56] IBM. Programmer’s reference manual: FORTRAN automatic coding system for the IBM

704. Technical report C28-6000, IBM Corporation, New York, NY, USA, October 15, 1956. 51 pp. URL http://www.

fh-jena.de/~kleine/history/
languages/FortranAutomaticCodingSystemForTheIBM704.
pdf.

IBM:1958:MDP

[IBM58] IBM. Manuel du Programmeur: Fortran: Progammation Automatique de l’Ordinateur 704 IBM. Institut de Calcul Scientifique, IBM France, 5, Place Vendome, Paris 1^{er}, France, February 1958. 56 pp. URL http://

archive.computerhistory.org/
resources/text/Fortran/102663111.
05.01.acc.pdf.  This man-

ual describes a French version of Fortran in which the library functions appear to match those in English, but the keywords are changed: aller a,

                                        continuer, imposer k en n, si,
                                        voyant, cle, arr^et, faire, lire
                                        [bande ou tambour], rebobiner,
                                        imprimer, esp arr, inscr bande,
                                        perforer, modèle . . ..

IBM:1968:IBFb

[IBM68] IBM Education Development. IBM basic FORTRAN IV for commercial users: programmed instruction course. IBM Corporation, Education Development, Endicott, NY, USA, 1968. 3 pp.

IEEE:1975:SCA

[IEE75] 3rd Symposium on Computer Arithmetic, November 19–20, 1975, Southern Methodist University, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1975. LCCN QA76.6.S919 1975. IEEE order number CH1017-3C.

IEEE:1978:PSC

[IEE78] Proceedings of the 4th Symposium on Computer Arithmetic, Santa Monica, CA, USA, 25–27 October 1978. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978. ISSN 1063-6889. IEEE catalog no. 78CH1412-6C.

IEEE:1979:CRI

[IEE79] Conference record: 1979 International Micro & Mini Computer Conference, Houston, Texas, November 14–16, 1979. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1979.

LCCN QA76.5 .I578 1979. IEEE Catalog no. 79CH474-6 MINI.

Isenhour:1979:ICP

[IJ79] Thomas L. Isenhour and Peter C.

Jurs. Introduction to computer programming for chemists: FOR-TRAN. Allyn and Bacon, Needham Heights, MA, USA, second edition, 1979. ISBN 0-205-05897-3. x + 356 pp. LCCN QD39.3.C6I83 1979. US$15.95.

Inada:1980:FBL

[Ina80a] Nobuyuki Inada. Fortran-based LISP system for REDUCE. Technical report, Information Science Laboratory, The Institute of Physical and Chemical Research, 1980.

Inada:1980:FLS

[Ina80b] Nobuyuki Inada. Fortran-based LISP system for REDUCE. Technical report, Information Science Laboratory, The Institute of Physical and Chemical Research, ????, 1980.

Indiana:1960:FFP

[Ind60] Indiana State Highway Dept.Planning Division. [FAP and FORTRAN programming]. State Highway Dept. of Indiana, Planning Division, Indianapolis, IN, USA, 1960.

[72] + 19 + 12 pp.

Ingalls:1971:FAF

[Ing71] D. Ingalls. Fete: A Fortran execution time estimator. Technical report, Stanford University, Stanford, CA, USA, February 1971. 10 pp.

IGCCCC:1964:SPF

[Ins64] Instituto Gulbenkian de Ciencia.Centro de Calculo Cientifico. Sistema de programacão FOR-TRAN 2 e sua utilizacao com o computador electronico instalado no C.C.C. Cursos e seminarios., Instituto Gulbenkian de Ciencia. Centro de Calculo Cientifico, Lisboa, Portugal, 1964. 150 pp.

IGCCCC:1970:SPF

[Ins70] Instituto Gulbenkian de Ciencia.Centro de Calculo Cientifico. Sistema de programacão Fortran II e sua utilizacao com um computador IBM 1620. Cursos e seminarios, Instituto Gulbenkian de Ciencia. Centro de Calculo Cientifico, Lisboa, Portugal, 1970. 160 pp.

IGCCCC:1974:SPF

[Ins74] Instituto Gulbenkian de Ciencia.Centro de Calculo Cientifico. Sistema de programacão Fortran 2: e sua utilizacao com o computador electronico instalado no C.C.C. Cursos e seminarios / Instituto Gulbenkian de Ciencia, Centro de Calculo Cientifico Cursos e seminarios (Instituto Gulbenkian de Ciencia. Centro de Calculo Cientifico). O Centro, Lisboa, Portugal, 1974. 150 pp.

ISA:1976:ICSa

[Ins76a] Instrument Society of America. Industrial computer system FOR-TRAN procedures for executive functions, process input/output, and bit manipulation. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised: February 1976 edition, 1976. ISBN 0-87664-393-4. 11 pp.

ISA:1976:ICSb

[Ins76b] Instrument Society of America.

Industrial computer system FOR-TRAN procedures for executive functions, process input/output, and bit manipulation: standard. Instrument Society of America, Pittsburgh, PA, USA, 1976. 11 pp.

IBM:1957:FACb

[Int57a] International Business Machines Corporation. FORTRAN: an automatic coding system for the IBM 704. IBM Corporation, New York, NY, USA, 1957. 53 pp.

IBM:1957:FACa

[Int57b] International Business Machines Corporation. Fortran automatic coding system for the IBM 704 data processing system. IBM Corporation, New York, NY, USA, 1957. 3 pts. pp.

IBM:1957:GIM

[Int57c] International Business Machines Corporation. General information manual; programmer’s primer for FORTRAN automatic coding system for the IBM 704 data processing system. IBM Corporation, New York, NY, USA, 1957. 64 pp.

IBM:1957:PPF

[Int57d] International Business Machines Corporation. Programmer’s primer for Fortran automatic coding system for the IBM 704.

IBM Corporation, New York, NY, USA, 1957. 64 pp.

                                        URL http://www.science.uva.
                                        nl/faculteit/museum/704fortranprimer.
                                        gif.

IBM:1957:TYF

[Int57e] International Business Machines Corporation. Twenty-five years of FORTRAN exhibit, 1957.

IBM:1958:RMF

[Int58] International Business Machines Corporation. Reference manual [S]: FORTRAN II for the IBM 704 data processing system. IBM Corporation, New York, NY, USA, 1958. 63 pp.

IBM:1959:FAC

[Int59a] International Business Machines Corporation. 709 FORTRAN automatic coding system for the IBM 709 data processing system; reference manual. IBM Corporation, New York, NY, USA, 1959. 82 pp.

IBM:1959:FOR

[Int59b] International Business Machines Corporation. 709 FORTRAN operations; reference manual. IBM Corporation, White Plains, NY, USA, 1959. 46 pp.

IBM:1959:FIP

[Int59c] International Business Machines Corporation. Fortran Introductory Programmer’s Manual. IBM Corporation, New York, NY, USA, 1959. 37 + 31 + 28 + i + 6 pp. Section 1 (20 March 1959); Section 2 (10 April 1957); Section 3 (7 June 1957); Addenda to Fortran Programmer’s Reference Manual.

IBM:1959:IRM

[Int59d] International Business Machines Corporation. IBM reference manual [S]: FORTRAN, automatic coding system for the IBM 704 data processing system. IBM Corporation, New York, NY, USA, 1959. 56 pp.

IBM:1959:POM

[Int59e] International Business Machines Corporation. Preliminary Operator’s Manual: Automatic Coding System for IBM 704 EDPM (Fortran 4-1-4-1). IBM Corporation, New York, NY, USA, 1959. 37 pp. Includes copy of [BBB^{+}57], plus two letters by J. Backus.

IBM:1959:RMF

[Int59f] International Business Machines Corporation. Reference manual [S]: 709 FORTRAN, automatic coding system for the IBM 709 data processing system. IBM Corporation, New York, NY, USA, 1959. 82 pp.

IBM:1960:FIM

[Int60a] International Business Machines Corporation. Fortran II Modifications. FLIBL (Fortran Library Loader), Input Output Hollerith Routine, etc. IBM Corporation, New York, NY, USA, April 11, 1960. Modification to 704 Fortran II, with a letter by Bekish explaining contents of 111pages of “listings”.

IBM:1960:IRM

[Int60b] International Business Machines Corporation. IBM reference manual: 7070 basic FORTRAN. IBM Corporation, New York, NY, USA, 1960. 68 pp.

IBM:1960:RMI

[Int60c] International Business Machines Corporation. Reference manual IBM 709/7090 FORTRAN monitor. IBM Corporation, White Plains, NY, USA, 1960. 22 pp.

IBM:1961:GIM

[Int61a] International Business Machines Corporation. General information manual, FORTRAN. IBM Corporation, White Plains, NY, USA, revised edition, 1961. 103 pp.

IBM:1961:IRMa

[Int61b] International Business Machines Corporation. IBM reference manual 704 FORTRAN programming system. IBM Corporation, Data Processing Division, White Plains, NY, USA, revised edition, 1961. 99 pp.

IBM:1961:IRMb

[Int61c] International Business Machines Corporation. IBM reference manual 709/7090 Fortran programming system. IBM Corporation, New York, NY, USA, revised edition, 1961. 104 pp.

IBM:1961:RMFa

[Int61d] International Business Machines Corporation. Reference manual, 704 FORTRAN programming system. IBM Corporation, White Plains, NY, USA, major revision edition, 1961. 99 pp.

IBM:1961:RMFb

[Int61e] International Business Machines Corporation. Reference manual, 709/7090 FORTRAN programming system. Technical report, IBM Corporation, Data Processing Division, White Plains, NY, USA, 1961. 104 pp.

IBM:1961:RMFc

[Int61f] International Business Machines Corporation. Reference manual, 709/7090 FORTRAN programming system. IBM Corporation, Data Processing Division, White Plains, NY, USA, revised edition, 1961. 67 pp.

IBM:1961:F

[Int61g] International Business Machines Corporation.Data Processing and Division. FORTRAN. IBM Corporation, New York, NY, USA, corrected printing edition, 1961. 103 pp.

IBM:1962:FF

[Int62a] International Business Machines Corporation. 1620 FORTRAN (with FORMAT). IBM Corporation, White Plains, NY, USA, 1962. 95 pp.

IBM:1962:IPS

[Int62b] International Business Machines Corporation. IBM 7070-series programming systems: FORTRAN.

IBM Corporation, Poughkeepsie, NY, USA, 1962. 100 pp.

IBM:1962:RMIb

[Int62c] International Business Machines Corporation. Reference manual, IBM 1620 FORTRAN. IBM Product Publications Dept., San Jose, CA, USA, 1962. 94 pp.

IBM:1962:RMIa

[Int62d] International Business Machines Corporation.Data Processing Division. Reference manual; IBM 709/ 7090 programming systems: FOR-TRAN assembly program (FAP). IBM Corporation, White Plains, NY, USA, major revision, Sept. 1962 edition, 1962. 77 pp.

IBM:1963:FF

[Int63a] International Business Machines Corporation. 1620 FORTRAN (with FORMAT). IBM Corporation, White Plains, NY, USA, 1963. 107 pp.

IBM:1963:FIG

[Int63b] International Business Machines Corporation. FORTRAN II: general information manual. IBM Corporation, Poughkeepsie, NY, USA, 1963. 103 pp.

IBM:1963:FPIb

[Int63c] International Business Machines Corporation. FORTRAN: programmed instruction course. IBM Corporation, Endicott, NY, USA, 1963. various pp.

IBM:1963:IFF

[Int63d] International Business Machines Corporation. IBM 1620 FOR-TRAN (with Format). IBM Corporation, San Jose, CA, USA, 1963. 106 pp.

IBM:1963:IPSd

[Int63e] International Business Machines Corporation. IBM 7090/7094 programming systems: FORTRAN II operations. IBM Corp., Programming Systems Publications, Poughkeepsie, NY, USA, 1963. various pp.

IBM:1963:IPSc

[Int63f] International Business Machines Corporation. IBM 7090/7094 programming systems, FORTRAN II programming … an automatic coding system for the IBM 7090/ 7094 data processing system. IBM Corporation, Programming Systems Publications, Poughkeepsie, NY, USA, 1963. 51 pp.

IBM:1963:ISRa

[Int63g] International Business Machines Corporation. IBM systems reference library: IBM 7090/7094 programming systems, FORTRAN II assembly program (FAP). IBM Corporation, New York, NY, USA, 1963. 75 pp.

IBM:1963:ISRb

[Int63h] International Business Machines Corporation. IBM systems reference library: IBM 7090/7094 programming systems; Fortran II programming. IBM Corporation, Poughkeepsie, NY, USA, 1963. 51 pp. URL http://www.

                                        fh-jena.de/~kleine/history/
                                        languages/C28-6054-4_7090_FORTRANII.
                                        pdf.

IBM:1963:PIC

[Int63i] International Business Machines Corporation. Programmed instruction course — FORTRAN. IBM Corporation, Education Development, Endicott, NY, USA, 1963. various pp.

IBM:1963:F

[Int63j] International Business Machines Corporation.Data Processing and Division. FORTRAN. IBM Corporation, New York, NY, USA, minor revision, June 1963 edition, 1963. 103 pp.

IBM:1963:FGI

[Int63k] International Business Machines Corporation.Data Processing and Division. FORTRAN; general information manual. IBM Corporation, White Plains, NY, USA, 1963. 103 pp.

IBM:1963:FIP

[Int63l] International Business Machines Corporation.Data Processing and Division. FORTRAN illustrations: programmed instruction course. IBM Corporation, Data Processing Division, White Plains, NY, USA, 1963. various pp.

IBM:1963:FPIa

[Int63m] International Business Machines Corporation.Data Processing and Division. FORTRAN programmed instruction course: problem book.

IBM Corporation, White Plains, NY, USA, 1963. 46 pp.

IBM:1963:GIM

[Int63n] International Business Machines Corporation.Data Processing and Division. General information manual — FORTRAN. IBM Corporation, Poughkeepsie, NY, USA, revised edition, 1963. vii

+ 103 pp. URL http://www.
fh-jena.de/~kleine/history/
languages/F28-8074-3_FORTRANII_
GenInf.pdf.

IBM:1963:IPSa

[Int63o] International Business Machines Corporation.Data Processing and Division. IBM 7040/7044 programming systems (8K) FOR-TRAN IV language. IBM Corporation, White Plains, NY, USA, 1963. 27 pp.

IBM:1963:IPSb

[Int63p] International Business Machines Corporation.Data Processing and Division. IBM 7090/7094 programming systems FORTRAN IV language. IBM Corporation, New York, NY, USA, 1963. 37 pp.

IBM:1964:ECL

[Int64a] International Business Machines Corporation. Electronic computers in the life sciences; FOR-TRAN programming and applications. IBM Corporation, Los Angeles, CA, USA, 1964. various pp.

IBM:1964:FSO

[Int64b] International Business Machines Corporation. Fortran Specifications and Operating Procedures, IBM 1401. IBM Corporation, White Plains, NY, USA, 1964.

61 pp. URL http://bitsavers.
org/pdf/ibm/1401/C24-1455-2_
Fortran_Specifications_and_
Operating_Procedures_Apr65.
pdf.

IBM:1964:ISRa

[Int64c] International Business Machines Corporation. IBM systems reference library: IBM 7090/7094 programming systems, FORTRAN II assembly program (FAP). IBM Corporation, Poughkeepsie, NY, USA, minor rev. edition, 1964. 75 pp.

IBM:1964:ISRb

[Int64d] International Business Machines Corporation. IBM systems reference library: IBM 7090/ 7094 programming systems, FOR-TRAN IV language. IBM Corporation, Poughkeepsie, NY, USA, minor rev. edition, 1964.

37 pp.  URL http://www.
fh-jena.de/~kleine/history/
languages/C28-6274-1_7090_FORTRANIV.
pdf.

IBM:1964:RMI

[Int64e] International Business Machines Corporation. Reference manual: IBM 1620 Fortran II Programming system. IBM Corporation, White Plains, NY, USA, 1964. 44 pp.

IBM:1964:STF

[Int64f] International Business Machines Corporation. Student text; a FOR-TRAN primer with business administration exercises. IBM Corporation, White Plains, NY, USA, 1964. 125 pp.

IBM:1964:IPS

[Int64g] International Business Machines Corporation.Data Processing and Division. IBM 7090/7094 programming systems FORTRAN II programming. IBM Corporation, White Plains, NY, USA, major revision, April 1964 edition, 1964. 51 pp.

IBM:1965:FLS

[Int65a] International Business Machines Corporation. FORTRAN in the life sciences: student text. IBM Corporation, White Plains, NY, USA, 1965. 112 pp.

IBM:1965:FPS

[Int65b] International Business Machines Corporation. FORTRAN in the physical sciences: student text. IBM Corporation, White Plains, NY, USA, 1965. 87 pp.

IBM:1965:FII

[Int65c] International Business Machines Corporation. Fortran IV for IBM System/360. Education Development, Endicott, NY, USA, 1965. 8 pp.

IBM:1965:IOSb

[Int65d] International Business Machines Corporation. IBM 7040/7044 operating system (16/32K) subroutine library (FORTRAN IV mathematical subroutines). IBM Corporation, New York, NY, USA, 1965. 32 pp.

IBM:1965:ISB

[Int65e] International Business Machines Corporation. IBM System/360 basic programming support: FOR-TRAN IV. IBM Corporation, Poughkeepsie, NY, USA, 1965. 81 pp.

IBM:1965:IOSa

[Int65f] International Business Machines Corporation.Data Processing and Division. IBM 7040/7044 operating system (16/32K) FORTRAN IV language. IBM Corporation, White Plains, NY, USA, minor rev. edition, 1965. 43 pp.

IBM:1966:ISR

[Int66a] International Business Machines Corporation. IBM systems reference library: IBM 7090/7094 IB-SYS operating system, version 13, FORTRAN IV language. IBM Corporation, New York, NY, USA, minor rev. edition, 1966. various pp.

IBM:1966:ISBa

[Int66b] International Business Machines Corporation. IBM System/360; basic FORTRAN IV language. IBM Corporation, New York, NY, USA, ???? edition, 1966. 93 pp.

IBM:1966:ISBb

[Int66c] International Business Machines Corporation. IBM System/360: basic FORTRAN IV language. IBM Corporation, New York, NY, USA, second edition, 1966. 84 pp.

IBM:1966:ISFa

[Int66d] International Business Machines Corporation. IBM System/360: FORTRAN IV language. IBM Corporation, New York, NY, USA, seventh edition, 1966. 125 pp.

IBM:1966:ISFb

[Int66e] International Business Machines Corporation. IBM System/360; Fortran IV language. IBM Corporation, White Plains, NY, USA, 1966. 121 pp.

IBM:1966:ISOa

[Int66f] International Business Machines Corporation. IBM System/360 operating system, FORTRAN IV (E) programmer’s guide. IBM Corporation, White Plains, NY, USA, 1966. 117 pp.

IBM:1966:ISOb

[Int66g] International Business Machines Corporation. IBM System/360 operating system: FORTRAN IV (E) programmer’s guide. IBM Corporation, New York, NY, USA, third edition, 1966. 123 pp.

IBM:1966:ISOc

[Int66h] International Business Machines Corporation. IBM System/360 operating system, FORTRAN IV (E) programmer’s guide. IBM Corporation, White Plains, NY, USA, 1966. 117 pp.

IBM:1966:ISOd

[Int66i] International Business Machines Corporation. IBM System/360 operating system: FORTRAN IV (G) programmer’s guide. IBM Corporation, New York, NY, USA, second edition, 1966. 116 pp.

ICL:1966:FSM

[Int66j] International Computers, Ltd. Fortran 1900 series manual. I.C.L., London, UK, 1966. 27 pp.

IBM:1964:FIL

[Int67a] International Business Machines Corporation. Fortran IV language specifications, program specifications, and operating procedures IBM 1401, 1440, and 1460: Programs: 1401-F0-051 (disk resident system), 1401-F0-052 (tape resident system). IBM Corporation, Data Processing Division, White Plains, NY, USA, fourth edition, 1964 (or 1967??). 100 pp.

IBM:1967:FII

[Int67b] International Business Machines Corporation.Data Processing and Division. Fortran IV for IBM System/360. IBM Corporation, New York, NY, USA, minor rev. edition, 1967. various pp.

IBM:1967:GPF

[Int67c] International Business Machines Corporation.Data Processing and Division. A guide to PL/I for FORTRAN users; student text. IBM Corporation, White Plains, NY, USA, 1967. 36 pp.

IBM:1968:IBFe

[Int68a] International Business Machines Corporation. IBM 1130/1800 basic FORTRAN IV language. IBM Corporation, White Plains, NY, USA, 1968. 57 pp.

IBM:1968:ISO

[Int68b] International Business Machines Corporation. IBM System/360 Operating System: FORTRAN IV (H) Compiler Program Logic Manual: Program Number 360S-FO-500. IBM Programming Publications, New York, NY, USA, fourth edition, November 1968. 257

pp. URL http://www.bitsavers.
org/pdf/ibm/360/fortran/Y28-
6642-3_FortH_PLM_Nov68.pdf.

IBM:1968:FI

[Int68c] International Business Machines Corporation.Data Processing and Division. FORTRAN for the IBM 1130. IBM Corporation, Education Development, Endicott, NY, USA, minor revision edition, 1968. 7 v.in 1 pp.

IBM:1968:FIIa

[Int68d] International Business Machines Corporation.Data Processing and Division. FORTRAN IV for IBM System/360 and System/370. IBM Corporation, Poughkeepsie, NY, USA, 1968. ???? pp.

IBM:1968:FIIb

[Int68e] International Business Machines Corporation.Data Processing and Division. Fortran IV for IBM System/360: problem book. Education Development, Endicott, NY, USA, minor rev. edition, 1968. 53 pp.

IBM:1968:FIIc

[Int68f] International Business Machines Corporation.Data Processing and Division. FORTRAN IV for IMB System/360 and System/370; programmed instruction. IBM Corporation, DPD Education Development, Publications Services, Poughkeepsie, NY, USA, 1968. various pp.

IBM:1968:GPF

[Int68g] International Business Machines Corporation.Data Processing and Division. A guide to PL/I for FORTRAN users: student text. IBM Corporation, New York, NY, USA, revised May 1968 edition, 1968. 36 pp.

IBM:1968:IBFc

[Int68h] International Business Machines Corporation.Data Processing and Division. IBM 1130 basic FOR-TRAN IV for commercial users: examinations. Education Development, Endicott, NY, USA, 1968. 36 pp.

IBM:1968:IBFd

[Int68i] International Business Machines Corporation.Data Processing and Division. IBM 1130 basic FOR-TRAN IV for commercial users: text. Education Development, Endicott, NY, USA, 1968. 163 pp.

IBM:1968:ISFb

[Int68j] International Business Machines Corporation.Data Processing and Division. IBM System/360: FORTRAN IV language. IBM Programming Publications, New York, NY, USA, eighth edition, 1968. 126 pp.

IBM:1968:ISFa

[Int68k] International Business Machines Corporation.Data Processing and Division. IBM System/360. FORTRAN IV library subprograms; program number 360S-LM-501, 360F-LM-619, 360N-LM-480. IBM Corporation, White Plains, NY, USA, fifth edition, 1968. various pp.

IBM:1969:FSB

[Int69a] International Business Machines Corporation.Data Processing and Division. FORTRAN for schools of business administration. IBM Corporation, White Plains, NY, USA, 1969. 174 pp.

IBM:1969:IBF

[Int69b] International Business Machines Corporation.Data Processing and Division. IBM 1130/1800 basic FORTRAN IV language. IBM Corporation, White Plains, NY, USA, fifth edition, 1969. 57 pp.

IBM:1969:ISB

[Int69c] International Business Machines Corporation.Data Processing and Division. IBM System/360 basic FORTRAN IV language. IBM Corporation, White Plains, NY, USA, third edition, 1969. various pp.

IBM:1970:FLS

[Int70a] International Business Machines Corporation. FORTRAN in the life sciences: student text. IBM Corporation, New York, NY, USA, 1970. 112 pp.

IBM:1970:ISD

[Int70b] International Business Machines Corporation.Data Processing Division. IBM System/360 disk operating system: FORTRAN IV programmer’s guide. IBM Corporation, White Plains, NY, USA, third edition, 1970. 139 pp.

IBM:1970:ISO

[Int70c] International Business Machines Corporation.Data Processing Division. IBM System/360 operating system: FORTRAN IV (G and H) programmer’s guide. (Program numbers 360S-FO-500 [and] 360S-FO-520). IBM Corporation, White Plains, NY, USA, third edition, 1970. 207 pp.

IBM:1971:FPT

[Int71a] International Business Machines Corporation. 1130 FORTRAN programming techniques. IBM Corporation, White Plains, NY, USA, 1971. 17 pp.

IBM:1971:ISSc

[Int71b] International Business Machines Corporation. IBM System/360 and System/370 FORTRAN IV Language. IBM Corporation, New York, NY, USA, 9th edition, 1971. various pp.

IBM:1971:ISF

[Int71c] International Business Machines Corporation. IBM System/360 OS FORTRAN IV mathematical and service subprograms supplement for Mod I and Mod II libraries. IBM Corporation, New York, NY, USA, 1971. 33 pp.

IBM:1971:ISTc

[Int71d] International Business Machines Corporation. IBM System/360 OS (TSO) code and go FORTRAN processor: terminal user’s guide. IBM Corporation, New York, NY, USA, second edition, 1971. 182 pp.

IBM:1971:IBF

[Int71e] International Business Machines Corporation.Data Processing and Division. IBM 1130/1800 basic FORTRAN IV language. IBM Corporation, White Plains, NY, USA, eighth edition, 1971. 57 pp.

IBM:1971:ISSa

[Int71f] International Business Machines Corporation.Data Processing and Division. IBM system 360 and system 370: FORTRAN IV language. IBM Corporation, White Plains, NY, USA, nineth edition, 1971. various pp.

IBM:1971:ISSb

[Int71g] International Business Machines Corporation.Data Processing and Division. IBM System/360 and System/370: FORTRAN IV language. IBM Corporation, White Plains, NY, USA, nineth edition, 1971. various pp.

IBM:1971:ISO

[Int71h] International Business Machines Corporation.Data Processing and Division. IBM System/360 operating system FORTRAN IV mathematical and service subprograms, supplement for Mod I and Mod II libraries. IBM Corporation, White Plains, NY, USA, 1971. 33 pp.

IBM:1971:ISTa

[Int71i] International Business Machines Corporation.Data Processing and Division. IBM System/360 OS (TSO) terminal user’s supplement for FORTRAN IV (G1) processor and TSO FORTRAN prompter. IBM Corporation, White Plains, NY, USA, second edition, 1971. 126 pp.

IBM:1971:ISTb

[Int71j] International Business Machines Corporation.Data Processing and Division. IBM System/360 OS (TSO) terminal user’s supplement for FORTRAN IV (G1) processor and TSO FORTRAN prompter. IBM Corporation, White Plains, NY, USA, second edition, 1971. 126 pp.

ICL:1971:EF

[Int71k] International Computers, Ltd. Extended FORTRAN. International Computers Ltd., London, UK, ICL student edition, 1971. xiii + 92 pp.

IBM:1972:FID

[Int72a] International Business Machines Corporation. FORTRAN interactive debug for CMS and TSO installation reference. IBM Corporation, New York, NY, USA, second edition, 1972. ???? pp.

IBM:1972:FIGa

[Int72b] International Business Machines Corporation. FORTRAN IV (G1) processor and TSO FORTRAN prompter for OS and VM/ 370 (CMS): installation reference material. IBM Corporation, New York, NY, USA, second edition, 1972. 53 pp.

IBM:1972:FIGb

[Int72c] International Business Machines Corporation. FORTRAN IV (G1) processor for OS and VM/ 370 (CMS). IBM Corporation, New York, NY, USA, 1972. 2 pp.

IBM:1972:FIH

[Int72d] International Business Machines Corporation. FORTRAN IV (H Extended) compiler for OS and VM/ 370 (CMS). IBM Corporation, New York, NY, USA, 1972. 2 pp.

IBM:1972:FIL

[Int72e] International Business Machines Corporation. FORTRAN IV library (Mod II) for OS and VM/ 370 (CMS). IBM Corporation, New York, NY, USA, 1972. 2 pp.

IBM:1972:ICGb

[Int72f] International Business Machines Corporation. IBM code and go FORTRAN processor for OS and VM/ 370 (CMS): installation reference material. IBM Corporation, New York, NY, USA, second edition, 1972. 72 pp.

IBM:1972:IFI

[Int72g] International Business Machines Corporation. IBM FORTRAN interactive debug for OS (TSO) and VM/ 370 (CMS): installation reference material. IBM Corporation, New York, NY, USA, second edition, 1972. 80 pp.

IBM:1972:ICGa

[Int72h] International Business Machines Corporation. IBM OS code and Go FORTRAN and FORTRAN IV (G1) programmer’s guide program product. IBM Corporation, White Plains, NY, USA, second edition, 1972. 188 pp.

IBM:1972:FSB

[Int72i] International Business Machines Corporation.Data Processing and Division. FORTRAN for schools of business administration. IBM Corporation, White Plains, NY, USA, second edition, 1972. 174 pp.

IBM:1972:ISS

[Int72j] International Business Machines Corporation.Data Processing and Division. IBM System/360 and System/370 FORTRAN IV language. IBM Corporation, White Plains, NY, USA, tenth edition, 1972. various pp.

IBM:1972:ISOa

[Int72k] International Business Machines Corporation.Data Processing and Division. IBM System/360 operating system FORTRAN IV (G and H) programmer’s guide. Program numbers 360S-F0-500, 360S-F0-520. IBM Corporation, White Plains, NY, USA, fourth edition, 1972. various pp.

IBM:1972:CLF

[Int72l] International Business Machines Corporation.Data Processing Division. CALL-OS: learning FOR-TRAN; terminal-oriented selfstudy text. IBM Corporation, White Plains, NY, USA, 1972. various pp.

IBM:1972:ISOb

[Int72m] International Business Machines Corporation.Data Processing Division. IBM System/360 operating system FORTRAN IV library; mathematical and service subprograms. IBM Corporation, White Plains, NY, USA, second edition, 1972. various pp.

IBM:1972:ISOc

[Int72n] International Business Machines Corporation.Data Processing Division. IBM System/360 operation system: FORTRAN IV (G and H) Programmer’s guide. IBM Corporation, White Plains, NY, USA, 4th edition, 1972. various pp.

IBM:1973:IBF

[Int73a] International Business Machines Corporation.Data Processing and Division. IBM 1130/1800 basic FORTRAN IV language. IBM Corporation, White Plains, NY, USA, nineth edition, 1973. iii + 57 pp.

IBM:1973:ISO

[Int73b] International Business Machines Corporation.Data Processing Division. IBM System/360 operation system: FORTRAN IV (G and H) Programmer’s guide. IBM Corporation, New York, NY, USA, fifth edition, 1973. 257 pp.

IBM:1974:CGF

[Int74a] International Business Machines Corporation. Code and go FOR-TRAN for OS and VM/ 370 (CMS). IBM Corporation, New York, NY, USA, 1974. 2 pp.

IBM:1974:IFIb

[Int74b] International Business Machines Corporation. IBM FORTRAN IV (H Extended) compiler and FOR-TRAN library (Mod II) for OS and VM/ 370 (CMS): installation reference material. IBM Corporation, New York, NY, USA, third edition, 1974. 85 pp.

IBM:1974:IFIa

[Int74c] International Business Machines Corporation. IBM OS FORTRAN IV (H Extended) Compiler and Library (Mod II ) messages: program numbers 5734-F03, 5734-LM3: program product. IBM Corporation, White Plains, NY, USA, second edition, 1974. 146 pp.

IBM:1974:IFIc

[Int74d] International Business Machines Corporation. IBM OS FORTRAN IV (H Extended) compiler programmer’s guide: program numbers 5734-F03, 5734-LM3: program product. IBM Corporation, White Plains, NY, USA, third edition, 1974. 206 pp.

IBM:1974:TFP

[Int74e] International Business Machines Corporation. OS (TSO) FOR-TRAN prompter. IBM Corporation, New York, NY, USA, 1974. 1 pp.

IBM:1974:ISF

[Int74f] International Business Machines Corporation.Data Processing and Division. IBM System/3. FOR-TRAN IV reference manual; program numbers: 5702-F01 Model 10 Disk System; 5703-F01 Model 6; 5704-F01 Model 15. IBM Corporation, San Jose, CA, USA, third edition, 1974. 218 pp.

IBM:1974:ISS

[Int74g] International Business Machines Corporation.Data Processing and Division. IBM System/360 and System/370 FORTRAN IV language. IBM Corporation, Data Processing Division, White Plains, NY, USA, eleventh edition, 1974.

169 pp.  URL http://www.
fh-jena.de/~kleine/history/
languages/GC28-6515-10-FORTRAN-
IV-Language.pdf.

Inter-RegionalInformationSystem:1975:APM

[Int75a] Inter-Regional Information System, Eugene, OR, USA. Application programmer’s manual, FORTRAN: a reference guide for TCDMS application programmers coding in IBM FORTRAN IV (G and H). a product of the IRIS/ TCDMS Project, a joint effort of the Data Processing Authority, Portland, Oregon, and Regional Information Systems Department, Eugene, Oregon, 1975. vii + 88 pp.

IBM:1975:FID

[Int75b] International Business Machines Corporation. FORTRAN interactive debug for OS (TSO) and VM/ 370 (CMS) reference card. IBM Corporation, New York, NY, USA, 1975. 1 pp.

IBM:1975:ISD

[Int75c] International Business Machines Corporation. IBM System/360 disk operating system: FORTRAN IV programmer’s guide. IBM Corporation, White Plains, NY, USA, fourth edition, 1975. 139 pp.

IBM:1975:IVC

[Int75d] International Business Machines Corporation. IBM VM/ 370 (CMS) terminal user’s guide for FORTRAN IV program products. IBM Corporation, New York, NY, USA, second edition, 1975. 208 pp.

IBM:1975:IFI

[Int75e] International Business Machines Corporation.Data Processing and Division. IBM FORTRAN interactive debut for OS (TSO) and VM/ 370 (CMS) terminal user’s guide. IBM Corporation, White Plains, NY, USA, third edition, 1975. 135 pp.

IBM:1975:ISF

[Int75f] International Business Machines Corporation.Data Processing and Division. IBM System/360. FORTRAN IV library subprograms; program number 360S-LM-501, 360F-LM-619, 360N-LM-480. IBM Corporation, San Jose, CA, USA, 6th edition, 1975. various pp.

IBM:1976:ISS

[Int76] International Business Machines Corporation. IBM System/360 and System/370 FORTRAN IV language: systems. IBM Corporation, Data Processing Division, White Plains, NY, USA, 11th, rev. Apr. 30, 1976 / by IBM Technical newsletter no. GN26-0805 edition, 1976. ???? pp.

IBM:1977:ISS

[Int77a] International Business Machines Corporation. IBM System/360 and System/370 FORTRAN IV language: systems. IBM Corporation, Programming Publishing Division, San Jose, CA, USA, 11th edition, 1977. ???? pp.

IBM:1977:ISP

[Int77b] International Business Machines Corporation. An introduction to structured programming in FOR-TRAN. IBM Corporation, New York, NY, USA, 1977. iv + 40 pp.

IBM:1978:ISF

[Int78a] International Business Machines Corporation. IBM System/370 FORTRAN H extended: optimization enhancement: program description and operations manual. IBM Corporation, New York, NY, USA, 1978. iii + 49 pp.

IBM:1970:ECL

[Int78b] International Business Machines Corporation.University Dept. Electronic computers in the life sciences; a FORTRAN primer and areas of application. Technical report, IBM University Dept., Los Angeles, CA, USA, 1970 (or 1978??). various pp.

IBM:1970:ECP

[Int78c] International Business Machines Corporation.University Dept. Electronic computers in the physical sciences; a FORTRAN primer and areas of application. Technical report, IBM University Dept., Los Angeles, CA, USA, 1970 (or 1978??). various pp.

ICL:1978:DIF

[Int78d] International Computers, Ltd. DAP: introduction to FORTRAN programming. International Computers, Ltd., London, UK, 1978. various pp.

IBM:1979:FII

[Int79] International Business Machines Corporation.Data Processing and Division. FORTRAN IV for IBM System/360 and System/ 370: programmed instruction, text. IBM Corporation, DPD Education-Publishing/Media Support, Poughkeepsie, NY, USA, minor rev. edition, 1979. 248 pp.

IBM:1980:VFC

[Int80a] International Business Machines Corporation. VS FORTRAN compiler and library. IBM Corporation, New York, NY, USA, 1980. 3 pp.

IMSL:1980:LCD

[Int80b] International Mathematical Statistical Libraries, Inc. Library contents document: FORTRAN subroutines in the areas of Mathematics and Statistics. IMSL, Inc., Houston, TX, USA, eighth edition, 1980. ???? pp.

IBM:1968:IBFa

[Int8 ] International Business Machines Corporation.Data Processing and Division. IBM 1130 basic FOR-TRAN IV for commercial users: advisor guide supplement. IBM Corporation, Education Development, Endicott, NY, USA, 1968 (??). 15 pp.

Ibramsha:1978:DLE

[IR78] M. Ibramsha and V. Rajaraman.

Detection of logical errors in decision table programs. Comm. ACM, 21(12):1016–1025, December 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irvine:1960:SIC

[Irv60] Crawford Allen Irvine. A system of IBM 650 computer programs supplementing the 650 FORTRAN system. Thesis (m.s.), Oklahoma State University, Stillwater, OK, USA, 1960. v + 47 pp.

Isambert:1973:TUF

[Isa73] M. Isambert. TSO: utilisation de FORTRAN. Compagnie internationale de services en informatique, Saclay, France, 1973. 26 pp.

Isenhour:1978:ICP

[Ise78] Thomas L. Isenhour. Introduction to Computer Programming for Chemists: Fortran. Allyn and Bacon, Needham Heights, MA, USA, July 1978. ISBN 0-205-05897-3. ???? pp. LCCN ???? US$24.61. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0205058973.

Izzo:1973:IFI

[Izz73] Paul John Izzo. An IBM 1130 Fortran IV cookbook. Thesis (m.a.), Glassboro State College, Glassboro, NJ, USA, 1973. v + 67 pp.

Jacobs:1973:CCIa

[Jac73a] Zeney P. Jacobs. Communicating With the Computer: Introductory Experiences, Fortran IV. Allyn and Bacon, Needham Heights, MA, USA, June 1973. ISBN 0-205-03819-0. vi + 345 pp. LCCN ???? US$19.63. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0205038190.

Jacobs:1973:CCIc

[Jac73b] Zeney P. Jacobs. Communicating with the computer; introductory experiences: FORTRAN IV. Holbrook Press, Boston, MA, USA, 1973. vi + 345 pp.

Jacobs:1973:CCIb

[Jac73c] Zeney P. Jacobs. Communicating with the computer: introductory experiences: FORTRAN IV: teacher’s guide. Allyn and Bacon, Needham Heights, MA, USA, 1973. 97 pp.

Jackson-Lorriman:1975:FF

[Jac75] Jackson-Lorriman Films. Fortran funnies, 1975.

KruseJacobsen:1978:FIC

[Jac78] Soren Kruse Jacobsen. Fortran IV codes for location. IMSOR, Lyngby, Denmark, 1978. 148 pp.

Jaffe:1972:CIF

[Jaf72] Richard M. Jaffe. A clear introduction to FORTRAN IV. Dickenson Pub. Co., Encino, CA, USA, 1972. ISBN 0-8221-0089-4. xv + 202 pp. LCCN QA76.73.F25J33.

Jaffe:1978:CIF

[Jaf78] Richard M. Jaffe. Clear Introduction to Fortran IV. Wadsworth, Pacific Grove, CA, USA, second edition, August 1978. ISBN 0-87872-175-4. ???? pp. LCCN ???? US$23.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0878721754.

Jaffe:1979:CIF

[Jaf79] Richard M. Jaffe. A clear introduction to Fortran IV: including standard FORTRAN, WATFOR, and WATFIV. Duxbury Press, North Scituate, MA, USA, second edition, 1979. ISBN 0-87872-175-4. xv + 289 pp.

Jahosua:1980:FI

[Jah80] Friedmann Jahosua. Fortran IV.

John Wiley and Sons, New York, London, Sydney, December 1980. ISBN 0-471-07771-2. ???? pp. LCCN ????

Jakolew:1973:FUS

[Jak73] Y. Jakolew. Fortran 5 (Univac série 1100). Université Paris-Sud, Orsay, France, 1973. 120 pp.

James:1966:FIP

[Jam66a] William R. James. Fortran IV program using double Fourier series for surface fitting of irregularly spaced data. Computer contribution 5, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1966. 19 pp.

Jamison:1966:FP

[Jam66b] Robert V. Jamison. Fortran programming. McGraw-Hill, New York, NY, USA, 1966. ix + 214 pp.

Jamison:1970:FIP

[Jam70] Robert V. Jamison. Fortran IV Programming: Based on the IBM System 1130. McGraw-Hill, New York, NY, USA, June 1970. ISBN 0-07-032270-8. ix + 369 pp. LCCN ???? US$31.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070322708.

James:1973:FPE

[Jam73a] Richard J. James. FORTNEAT: a program editor for standardization of coding and documentation in FORTRAN programs; a project in computer science. Thesis (m.s.), University of Missouri–Columbia, Columbia, MO, USA, 1973. 69 pp.

Jamison:1973:ICS

[Jam73b] R. V. Jamison. Introduction to Computer Science Mathematics. McGraw-Hill, New York, NY, USA, second edition, 1973. ISBN 0-07-032276-7. viii + 273 pp. LCCN QA39.2 .J35.

Jamison:1975:FIC

[Jam75] Robert V. Jamison. Fortran IV cheng shi she ji. Li Ming, Taipei, Taiwan, 1975. 566 o. pp.

James:1978:PPI

[Jam78] Edward B. James. Principles of programming: an introduction with Fortran. Pitman Publishing Ltd., London, UK, 1978. ISBN 0-273-01221-5. viii + 93 pp. LCCN QA76.6.J35.

Jay:1980:BFS

[Jay80] Lewis Robert Jay. Business Fortran: a Structured Approach. Wadsworth, Pacific Grove, CA, USA, December 1980. ISBN 0-534-00778-3. ???? pp. LCCN ???? US$30.75.

Jamison:1973:FIK

[JcK73] Robert V. Jamison and Tok chin Kim. Fortran IV kompyuto puroguraeming: IBM 1130 kompyuto chungsim. Chonpa Kwahaksa, Soul Tukpyolsi, Korea, 1973. 6 + 379 pp.

Jonch-Clausen:1976:UML

[JCMS76] Torkil Jonch-Clausen and Hubert Jean Morel-Seytpux. User’s manual for LPTOR, a Fortran IV linear programming routine. Technical Report PB-291 762 CER75-76TJ-HJM36, Colorado State University, Dept. of Civil Engineering, Fort Collins, CO, USA, 1976. 39 pp. Available through the National Technical Information Service.

Jonch-Clausen:1977:UMQa

[JCMS77a] Torkil Jonch-Clausen and Hubert J. Morel-Seytoux. User’s manual for QPTOR, a Fortran IV quadratic programming manual. Colorado State Un., Dept. of Civil Engineering, Ft. Collins, CO, USA, 1977. 68 pp.

Jonch-Clausen:1977:UMQb

[JCMS77b] Torkil Jonch-Clausen and Hubert Jean Morel-Seytoux. User’s manual for QPTOR, a Fortran IV quadratic programming routine. Technical Report PB-291 725 CER76-77TJ-HJM48, Colorado State University, Dept. of Civil Engineering, Fort Collins, CO, USA, 1977. iii + 72 pp. Available through the National Technical Information Service.

Jeffreys:1972:T

[Jef72]  W. H. Jeffreys.  TRIGMAN.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24): 20–21, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Jefferson:1977:PFP

[Jef77] Thomas H. Jefferson. PARAM: a FORTRAN preprocessor for variable dimensioning. Technical report, Applied Mathematics Division, Sandia Laboratories, Livermore, CA, USA, 1977. 20 pp.

JPL:1974:WFP

[Jet74] Jet Propulsion Laboratory. Workship on Fortran pre-processors for numerical software. Technical report, Jet Propulsion Laboratory, Pasadena, CA, USA, 1974. 52 pp.

Jeter:1979:VPI

[Jet79] J. P. Jeter. A variable precision interval data type extension to Fortran. Project report, Faculty and Graduate School of The University of Southwestern Louisiana, Lafayette, LA, USA, 1979.

Hofmann:1978:KIU

[jH78] Uwe jens Hofmann. Kopplung der Informatik-rechner Untereinander und mit dem Rechner des Hochschulrechenzentrums. Master’s thesis, Technische Universität Braunschweig (??), Braunschweig, Germany, 1978.

Johanson:1980:UMHb

[JID80] Robert C. Johanson, John C.

Imhoff, and Harley H. Davis. Users manual for hydrological simulation program — FORTRAN (HSPF). Research reporting series. 9, miscellaneous reports; epa-600/9-80-015, Environmental Research Laboratory, Office of Research and Development, U.S. Environmental Protection Agency, Athens, ??, USA, 1980. v + 678 pp. Available to the public through the National Technical Information Service.

Janiczek:1974:FAT

[JK74] P. M. Janiczek and George H. Kaplan. FORTRAN automatic typesetting system. United States Naval Observatory Circular 149, U.S. Naval Observatory, Washington, DC. USA, 1974. 71 pp.

James:1978:DFS

[JK78] Michael R. James and Wendell W. Kelley. Description of a FOR-TRAN subroutine for plotting three-dimensional data. Technical memorandum N78-32751 NASA TM-7877, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, August 1978. 27 pp. Available through the National Technical Information Service.

Johnson:1976:AGC

[JM76] Gearold R. Johnson and Robert A. Mueller. The automated generation of cross-system software for supporting micro/mini computer systems. ACM SIGPLAN Notices, 11(4):45–56, April 1, 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jurich:1977:TDL

[JMG77] Dale Jurich, Dennis Mickunas, and Raymond Groek. Technical Design Labs Fortran IV; documentation manual for version 2.0. Technical report, Small System Services, Inc., Urbana, IL, USA, 1977. 66 pp.

Jn:1969:CSF

[Jn69] Haag Jn. Comprehensive Standard Fortran Programming. Howard W. Sams, Indianapolis, IN 46268, USA, June 1969. ISBN 0-8104-5812-8. ???? pp. LCCN ???? US$10.30. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0810458128.

Johnson:1965:TFTa

[Joh65a] Carroll K. Johnson. OR TEP: a FORTRAN thermal-ellipsoid plot program for crystal structure illustrations. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1965. iv + 165 pp.

Johnson:1965:TFTb

[Joh65b] Carroll K. Johnson. OR TEP: a FORTRAN thermal-ellipsoid plot program for crystal structure illustrations. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1965. iv + 127 pp.

Johnston:1966:RPPb

[Joh66a] J. B. Johnston. Recent publications and presentations: Numerical Methods in FORTRAN, by John M. McCormick and Mario G. Salvadori. American Mathematical Monthly, 73(1):104–105, January 1966. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Johnston:1966:RPPa

[Joh66b] J. B. Johnston. Recent publications and presentations: Numerical Methods Using FORTRAN, by L. Dale Harris; FORTRAN Programming (II and IV), by L. Dale Harris. American Mathematical Monthly, 73(1):104, January 1966. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Johnsey:1971:SAF

[Joh71] Alan R. Johnsey. A statistical analysis of FORTRAN source programs. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1971. 242 pp.

Johnson:1972:MSF

[Joh72] Susan Johnson. MACROS to simulate FORTRAN 1/0 in assembler programs. AAEC/E248. Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1972. ISBN 0-642-99514-1. 7 pp.

Johnson:1974:CSS

[Joh74] Robert C. (Robert Clinton) Johnson. Calculation of supersonic stream parameters of a real gas from measurable quantities using FORTRAN IV routines. NASA technical note TN D-7653, National Aeronautics and Space Administration, Washington, DC, USA, 1974. 87 pp. For sale by the National Technical Information Service.

Johnson:1976:TFT

[Joh76] Carroll K. Johnson. OR TEP-II: a FORTRAN thermal-ellipsoid plot program for crystal structure illustrations. Technical report, U.S.

Oak Ridge National Laboratory, Chemistry Division, Oak Ridge, TN, USA, 1976. 128 pp. For sale by U.S. Dept. of Commerce, National Technical Information Service.

Johanson:1980:UMHa

[Joh80] Robert C. Johanson. Users manual for hydrological simulation program — FORTRAN (HSPF). Hydrocomp, Inc., Mountain View, CA, USA, 1980. 686 pp.

Jones:1964:MSA

[Jon64] Kenneth J. Jones. The multivariate statistical analyzer; a system of Fortran II programs to be run under 7090-4 FORTRAN Monitor System. ????, ????, 1964. iii + 180 pp.

Jones:1976:SPF

[Jon76] Daniel E. Jones. Structured programming and FORTRAN. Thesis (m.s.), Mississippi State University. Department of Computer Science, Mississippi State, MS, USA, 1976. viii + 66 pp.

Jones:1979:ACS

[Jon79] B. F. Jones. An ACM Core System implementation on a laboratory minicomputer utilizing lowperformance graphics devices. In IEEE IMMCC’79 [IEE79], pages 321–324. LCCN QA76.5 .I578 1979. IEEE Catalog no. 79CH474-6 MINI.

Josephs:1978:MCB

[Jos78] William H. Josephs. A minicomputer based library control system. ACM SIGMETRICS Performance Evaluation Review, 7(3– 4):126–132, November 1978. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

James:1972:FPC

[JOW72] Edward B. James, Fergus O’Brien, and Peter Whitehead. A Fortran programming course. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1972. ISBN 0-13-329748-9. v + 176 pp. LCCN QA76.73.F25J35 1972.

Joyner:1977:FPC

[Joy77] William B. Joyner. A FORTRAN program for calculating nonlinear seismic ground response. Open-file series 77-671, U.S. Geological Survey, Reston, VA, USA, 1977. 50 pp.

Joyner:1978:FPC

[Joy78] William B. Joyner. FORTRAN programs for calculating nonlinear seismic ground response in two dimensions. Open-file series 78-287, U.S. Geological Survey, Reston, VA, USA, 1978. 69 pp.

James:1975:AAD

[JR75] K. R. James and W. Riha. Algorithm 24. algorithm for deriving the chromatic polynomial of a graph. Computing, 14(1–2):195– 203, 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See remark [Ein76].

James:1976:AAG

[JR76] K. R. James and W. Riha. Algorithm 28. algorithm for generating graphs of a given partition. Computing, 16(1–2):153– 161, 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Johnson:1974:AFA

[JS74] S. G. Johnson and P. L. Sanger.

AAEC facilities to aid in the development of FORTRAN programs. AAEC/E318. Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1974. ISBN 0-642-99637-7. 65 pp.

Schneider:1970:PD

[jSJ70] H. j. Schneider and D. Jurksch. Programmierung von Datenverarbeitungsanlagen. Number 1225/1225a in Einführung in das Programmieren von Digitalrechnern. Walter de Gruyter, New York, NY, USA, second edition, 1970. ???? pp.

James:1967:ANM

[JSW67] M. L. (Merlin L.) James, G. M.

(Gerald M.) Smith, and J. C. Wolford. Applied numerical methods for digital computation with Fortran. International Textbook Co., Scranton, PA, USA, 1967. ix + 514 pp.

James:1970:MNA

[JSW70] M. L. (Merlin L.) James, Gerald M. Smith, and J. C. Wolford. Metodos numericos aplicados a la computacion digital con Fortran. Representaciones y Servicios de Ingenieria, Mexico, DF, Mexico, 1970. vii + 575 pp.

James:1977:ANM

[JSW77a] M. L. (Merlin L.) James, G. M. (Gerald M.) Smith, and J. C. Wolford. Applied numerical methods for digital computation with FOR-TRAN and CSMP. Harper & Row, New York, NY, USA, second edition, 1977. ISBN 0-7002-2499-8. x + 687 pp. LCCN QA 297 J235a 1977. Published in 1967 under title: Applied numerical methods for digital computation with FORTRAN. Includes bibliographical references and index.

James:1977:SMA

[JSW77b] M. L. (Merlin L.) James, G. M. (Gerald M.) Smith, and J. C. Wolford. Solutions manual to accompany Applied numerical methods for digital computation with FOR-TRAN and CSMP. Thomas Y. Crowell, New York, NY, USA, second edition, 1977. various pp.

Jenkins:1972:AAZ

[JT72] Michael A. Jenkins and J. F. Traub. ACM Algorithm 419: Zeros of a complex polynomial [C2]. Comm. ACM, 15(2):97–99, February 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?].

Tung:1979:CHS

[jT79] Chi jun Tung. Cheng hsu she chi yu yen: FORTRAN. Tien tzu chi shu tsung shu Tien tzu chi shu tsung shu (Tsinan, China). Shantung ko hsueh chi shu chu pan she: Shan-tung sheng hsin hua shu tien fa hsing, Chi-nan, ti 1 pan edition, 1979. 2 + 216 pp.

Julian:1975:FPA

[Jul75] Elmo C. Julian. Fortran programs for analyzing collaborative test data. Environmental monitoring series ????, National Environmental Research Center, Cincinnati, OH, USA, April 4, 1975. ???? pp.

Junginger:1968:BRWa

[Jun68] W. Junginger. Book report: W.

S. Dorn and H. J. Greenberg, Mathematics and Computing with FORTRAN Programming. Computing, 3(2):158, 1968. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Junginger:1969:BRH

[Jun69] W. Junginger. Book review: H.

P. Künzi. H. G. Tzschach und C. A. Zehnder, Numerische Methoden der mathematischen Optimierung mit ALGOL- und FORTRAN-Programmen (Leitfäden der angewandten Mathematik und Maechanik, Bd. 8). Computing, 4(1):87, 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Johnson:1967:FIPa

[JV67a] G. G. Johnson and Vladimir Vand.

Fortran II programs for the identification of multiphase unknown powder diffraction patterns using the Joint Committee’s powder diffraction file. American Society for Testing and Materials, Philadelphia, PA, USA, 1967. vi + 35 pp.

Johnson:1967:FIPb

[JV67b] G. G. Johnson and Vladimir Vand. Fortran IV programs (version 2) for the identification of multiphase unknown powder diffraction patterns using the Joint Committee’s powder diffraction file. American Society for Testing and Materials, Philadelphia, PA, USA, 1967. vi + 47 pp.

Johnson:1968:FIP

[JV68] G. G. Johnson and Vladimir Vand. Fortran IV programs (version 7) for the identification of multiphase unknown powder diffraction patterns. American Society for Testing and Materials, Philadelphia, PA, USA, 1968. vii + 55 pp.

Kuki:1971:FEP

[KA71] H. Kuki and J. Ascoly. FOR-TRAN extended-precision library. IBM Systems Journal, 10(1):39– 61, 1971. CODEN IBMSA7. ISSN 0018-8670.

Kahan:1966:FIS

[Kah66] W. Kahan. The FORTRAN IV subroutine QDRTC. Technical report, Computer library, McLennan Laboratories, University of Toronto, Toronto, ON, Canada, 1966.

Kahamer:1980:FIH

[Kah80a] D. A. Kahamer. Fortran implementation of heap programs for efficient table maintenance. ACM Transactions on Mathematical Software, 6(3):444–449, September 1980.

Kahaner:1980:AFI

[Kah80b] D. K. Kahaner. Algorithm 561:

FORTRAN implementation of heap programs for efficient table maintenance [Z]. ACM Transactions on Mathematical Software, 6 (3):444–449, September 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kallin:1971:IF

[Kal71] Sten Kallin. Introduction to FOR-TRAN. Studentlitteratur and Auerbach, Lund, Sweden and Princeton, NJ, USA, 1971. ISBN 0-87769-130-4. 168 pp.

Kaled:1972:DEF

[Kal72a] Albert Richard Kaled. Development and evaluation of a FOR-TRAN programming text based upon a modified Gagne hierarchical structure of learning. Thesis (ed. d.), Indiana University, Bloomington, IN, USA, 1972. 150 pp.

Kalish:1972:TIR

[Kal72b] D. (Daniel) Kalish. Two interfaces for recursive subprograms in FOR-TRAN. Technical note 1972-43, Massachusetts Institute of Technology, Lincoln Laboratory, Lexington, MA, USA, 1972. iii + 35 pp.

Kallin:1972:IF

[Kal72c] Sten Kallin. Introduction to FOR-TRAN. Studentlitteratur and Auerbach, Lund, Sweden and Princeton, NJ, USA, 1972. ISBN 0-87769-130-4 (Auerbach). 168 pp. LCCN QA76.73.F25 K35.

SeiInKang:1968:FIP

[Kan68] Sei In Kang. FORTRAN IV programs to plot Mössbauer spectra and to facilitate their analysis. Thesis (m.s.), Department of Physics, Mississippi State University, Mississippi State, MS, USA, 1968. 96 pp.

Kantrowitz:1971:CAI

[Kan71] Marvin Joel Kantrowitz. A comparative analysis of IBM third generation programming languages: ALP, FORTRAN, COBOL, RPG and PL/I. Thesis (m.b.a.), St. John’s University, New York, NY, USA, 1971. v + 65 pp.

Kantaris:1977:IBF

[Kan77] E. Kantaris. Introduction to BA-SIC and Fortran IV programming. Camborne School of Mines, Redruth (Trevenson, Redruth, Cornwall TR15 3SE), 1977. iv + 116 pp.

Kane:1979:DFP

[Kan79] Jean S. Kane. DEV1, a Fortran program for one-way analysis of variance of analytical data. Openfile series 79-282, U.S. Geological Survey, Reston, VA, USA, 1979. 18 pp.

Karunaratne:1968:FP

[Kar68] Samarajeewa Karunaratne. Fortran programming. Frewin, Colombo, Ceylon, 1968. 110 + [4] pp.

Karki:1973:FIP

[Kar73] Pentti Karki. A Fortran IV program for the inversion of a large symmetric matrix by block partitioning. Technical report, The Ohio State University, Dept. of Geodetic Science, Columbus, OH, USA, 1973. 39 pp. Distributed by U.S. National Technical Information Service.

Karpov:1976:AIF

[Kar76] V. Ia. (Vladimir Iakovlevich) Karpov. Algoritmicheskii iazyk Fortran: Fortran-Dubna. Bibliotechka programmista. Nauka, Moscow, Russia, 1976. 192 pp.

Karon:1977:RIA

[Kar77] John Karon. Reviews: Introduction to Applied Numerical Analysis, by Richard W. Hamming; Numerical Methods with FORTRAN IV Case Studies, by William S. Dorn and Daniel D. McCracken; Numerical Quadrature and Solution of Ordinary Differential Equations, by A. H. Stroud. American Mathematical Monthly, 84(4): 304–307, April 1977. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kasack:1974:OLR

[Kas74] U. Kasack. Orthogonale lineare Regression: Stichprobenparameter. (German) [Orthogonal linear regression: Sample parameters]. Computing, 13(1):17– 20, 1974. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic).

Katsanis:1968:FPS

[Kat68] Theodore Katsanis. FORTRAN program for spline fit curve. NASA technical memorandum X-1707, National Aeronautics and Space Administration; Clearinghouse for Federal Scientific and Technical Information, Washington, DC, USA, 1968. 9 pp.

Katsanis:1977:RFPc

[Kat77] Theodore Katsanis. Revised FOR-TRAN program for calculating velocities and streamlines on the hub-shroud midchannel stream surface of an axial-, radial-, or mixed-flow turbomachine or annular duct. NASA technical note TN D-8430, D-8431, National Aeronautics and Space Administration, Washington, DC, USA, 1977. ???? pp. For sale by the National Technical Information Service.

Katjan:1978:F

[Kat78a] Harry Katzan, Jr. Fortran 77. Computer Science Series. Van Nostrand Reinhold, New York, NY, USA, 1978. ISBN 0-442-24278-6. xvi + 207 pp. LCCN QA76.73 .F25 K373 1978.

Katzan:1978:F

[Kat78b] Harry Katzan, Jr. Fortran 77. Computer Science Series. Van Nostrand Reinhold, New York, NY, USA, 1978. ISBN 0-442-24278-6. xvi + 207 pp. LCCN QA76.73 .F25 K373 1978.

Kauffman:1965:KFI

[Kau65] Richard Henry Kauffman. Kingston Fortran II library subprograms as simulation aids. Thesis (m.s.), Air Force Institute of Technology, Wright-Patterson Air Force Base, OH, USA, 1965. vii + 97 pp. GRE/MATH/65-6. AD628335.

Kaufman:1960:ICS

[Kau69] Bruria Kaufman. Iterative computation of string nesting (Fortran code). Transformations and discourse analysis projects 20, University of Pennsylvania, Philadelphia, PA, USA, 1960 (or 1969??). 58 pp.

Kaufman:1978:FCB

[Kau78] Roger Emanuel Kaufman. A Fortran Coloring Book. MIT Press, Cambridge, MA, USA, March 1978. ISBN 0-262-61026-4. 285 pp. LCCN QA76.73 F25 K38

1978. URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0262610264.

Kazek:1978:FEI

[Kaz78] Chester S. Kazek. FORTRAN extended interactive debugging aid. Informal report LA 7467-MS, [Dept. of Energy], Los Alamos Scientific Laboratory, Los Alamos, NM, USA, September 1978. 11 pp. For sale by the National Technical Information Service.

Kuck:1973:MPO

[KBC^{+}73] D. Kuck, P. Budnick, S. Chen, E. Davis, J. Han, P. Kraska, D. Lawrie, Y. Muraoka, R. Strehendt, and R. Towle. Measurement of parallelism in ordinary Fortran programs. In Proceedings of the Sagamore Conference on Parallel Processing, pages 23– 36. ????, ????, 1973.

Kuck:1974:MPO

[KBC^{+}74] D. Kuck, P. Budnik, S. C. Chen, E. Davis.Jr., J. Han, P. Kraska, D. Lawrie, Y. Muraoka, R. Strebendt, and R. Towle. Measurements of parallelism in ordinary FORTRAN programs. Computer, 7(1):37–46, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kay:1960:FPC

[KC60] M. I. Kay and Don T. Cromer. Fortran program for the calculation of neutron diffraction magnetic intensities. PRNC 78, University of Puerto Rico, Puerto Rico Nuclear Center, Mayaguez, Puerto Rico, 1960. 6 + [6] pp.

Keys:1972:ICP

[KC72] William J. Keys and Thomas J. Cashman. Introduction to Computer Programming — Basic Fortran 4: a Practical Approach. Anaheim Pub. Co., Fullerton, CA, USA, June 1972. ISBN 0-88236-151-1. ???? pp. LCCN ???? US$27.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0882361511.

Kozdrowicki:1973:CIC

[KC73] Edward W. Kozdrowicki and Dennis W. Cooper. COKO III: the Cooper-Koz chess program. Comm. ACM, 16(7):411–427, July 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cho:1980:ISQ

[kC80] Chin kuei Cho. An Introduction to Software Quality Control. John Wiley and Sons, New York, London, Sydney, 1980. ISBN 0-471-04704-X. xxii + 445 pp. LCCN QA76.6 .C45.

Kedem:1980:ADC

[Ked80] G. Kedem. Automatic differentiation of computer programs. ACM Transactions on Mathematical Software, 6(2):150–165, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Keeler:1975:CSF

[Kee75] Gary Norman Keeler. A comparison of SNOBOL and FOR-TRAN as string manipulation languages. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1975. 65 pp.

Keith:1969:DHC

[Kei69] Samuel Palmer Keith. Design of a hardware compiler for a FOR-TRAN machine: a thesis. Thesis (m.s.e.), University of Alabama, Tuscaloosa, AL, USA, 1969. x + 109 pp.

Kent:1965:FAC

[Ken65] William Kent. A Fortran n-ary counter. Comm. ACM, 8(6):378, June 1965. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Kennedy:1970:ATF

[Ken70] Ken Kennedy. Automatic translation of Fortran programs to vector form. Technical Report Rice COMP TR476-029-4, Rice University, Houston, TX, USA, 1970.

Kennedy:1974:TSF

[Ken74] Michael Kennedy. Ten Statement Fortran Plus Fortran IV: Sensible, Modular, and Structured Programming With Watfor and Wativ. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, December 1974. ISBN 0-13-903385-8. ???? pp. LCCN ???? US$35.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0-13-903385-8.

Kennedy:1980:ATF

[Ken80] Ken Kennedy. Automatic translation of Fortran programs to vector form. Technical report, Rice University, Dept. of Mathematical Sciences, Houston, TX, USA, 1980. 94 pp.

Kernighan:1970:RPR

[Ker70] Brian W. Kernighan. Ratfor: a preprocessor for a rational Fortran. UNX 12.2.6, University of California, Berkeley, Computing Services, Berkeley, CA, USA, 1970. 11 pp. Reissued 1980, with series statement.

Keros:1972:CFI

[Ker72] John W. Keros. Computers, FOR-TRAN IV, and data processing applications. Allyn and Bacon, Needham Heights, MA, USA, 1972. vi + 280 pp.

Kernighan:1975:RAP

[Ker75a] B. W. Kernighan. RATFOR — a preprocessor for a rational Fortran. Software—Practice and Experience, 5(4):395–406, October 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1975:RPR

[Ker75b] Brian W. Kernighan. RATFOR — a preprocessor for a rational Fortran. Software—Practice and Experience, 5(4):395–406, October/December 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1979:RPR

[Ker79] Brian W. Kernighan. RATFOR — a preprocessor for a rational Fortran. Report 1193, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Kerr:1980:FEF

[Ker80] John D. Kerr. FORALL: An extensible Fortran system for conversationally accessing subroutine libraries. Software—Practice and Experience, 10(11):889– 896, November 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1974:RPR

[Ker4 ] Brian W. Kernighan. RATFOR — a preprocessor for a rational Fortran. Computing Science Technical Report 55, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974 (??). ?? pp.

Keys:1973:ICPa

[Key73a] William J. Keys. Introduction to computer programming: basic FORTRAN IV: a practical approach. Anaheim Pub. Co., Fullerton, CA, USA, 1973. ISBN 0-88236-151-1. 383 pp.

Keys:1973:ICPc

[Key73b] William J. Keys. Introduction to computer programming: basic FORTRAN IV; a practical approach. Anaheim Pub. Co., Fullerton, CA, USA, 1973. ISBN 0-88236-151-1. various pp.

Keys:1973:ICPb

[Key73c] William J. Keys. Introduction to computer programming: basic FORTRAN IV: a practical approach: instructor’s guide and answer manual. Anaheim Pub. Co., Fullerton, CA, USA, 1973. v + 93 pp.

Kapperman:1972:SPC

[KF72] Richard Kapperman and Robert Fedrick. Simplified programming concepts using IBM/ASSEMBLY, COBOL, and FORTRAN. Allyn and Bacon, Needham Heights, MA, USA, 1972. 258 pp.

Koch:1972:IDD

[KG72] H. S. Koch and M. H. Gotterer. An independent data definition facility for COBOL and FORTRAN. ACM SIGFIDET, Dean(ed), 1972.

Korncoff:1976:SSA

[KG76] A. R. Korncoff and C. H. Goodspeed. SAGES: system aiding the generation of engineering software. ACM SIGPLAN Notices, 11(12): 56–73, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kincaid:1980:IFI

[KGY80] David R. Kincaid, Roger G.

Grimes, and David M. Young. IT-PACK 2A: A Fortran implementation of adaptive accelerated iterative methods for solving large sparse linear systems. Report CNA-164, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1980.

Kossack:1975:ISC

[KH75] Carl Fredrick Kossack and Claudia I. Henschke. Introduction to Statistics and Computer Programming. Holden-Day, San Francisco, CA, USA, pilot edition, October 1975. ISBN 0-8162-4754-4. 651 pp. LCCN QA276.K683.

Khailany:1976:ICC

[Kha76] Asad Khailany. An introductory COBOL course with structured programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8 (1):11–16, February 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the SIGCSE–SIGCUE joint symposium on Computer science education.

Khailany:1977:ATS

[Kha77] Asad Khailany. Alternative teaching strategy for an introductory computer language course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):93–95, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Khuanghlawn:1968:FFE

[Khu68] Wheatley Khuanghlawn. Fortran toolkit for mechanism design and analysis. Report (m. eng.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1968. v + 80 pp.

Kiewit:1966:BFO

[Kie66] Kiewit Computation Center, Hanover, NH, USA. Background FORTRAN (October, 1966), 1966. various pp.

Kirch:1973:ISF

[Kir73] Allan M. Kirch. Introductory statistics with FORTRAN. International series in decision processes. Holt, Rinehart, and Winston, New York, NY, USA, 1973. ISBN 0-03-086392-9. xv + 458 pp. LCCN QA276.4 .K571.

Kirby:1979:MFC

[Kir79] William H. Kirby. Machineindependent FORTRAN coding of Lehmer random number generators. Open-file series 80- 004, U.S. Geological Survey, Reston, VA, USA, 1979. [i] + 11 pp.

Kaucher:1978:HPA

[KKU78] Edgar W. Kaucher, Rudi Klatte, and Christian Ullrich. Höhere Programmiersprachen ALGOL, FOR-TRAN, PASCAL in einheitlicher und ubersichtlicher Darstellung. Reihe Informatik; Bd. 24. Bibliographisches Institut, Mannheim, Germany, 1978. ISBN 3-411-01544-6. 258 pp.

Kahan:1964:FPM

[KL64] W. Kahan and J. J. Leppik. A FORTRAN post-mortem procedure. Comm. ACM, 7(1):15, January 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klein:1968:EPF

[Kle68] Gunter Klein. Einfuhrung in die Programmiersprache Fortran IV. Allgemeine Elektricitats-Gesellschaft AEG-Telefunken, Berlin, Germnay, 1968. 59 pp.

Klein:1969:EPF

[Kle69] Gunter Klein. Einfuhrung in die Programmiersprache Fortran IV. Allgemeine Elektricitats-Gesellschaft AEG-Telefunken, Berlin, Germany, second edition, 1969. 57 pp.

Klein:1977:EPF

[Kle77] G. Klein. Einführung in die Programmiersprache Fortran IV. Elitera, Berlin, Germany, 1977. ISBN 3-87087-101-6. ???? pp.

Klema:1978:RRS

[Kle78] Virginia Klema. ROSEPACK: RObust Statistics Estimation PACKage. ACM SIGNUM Newsletter, 13(2):18–19, June 1978. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kliphardt:1970:PDF

[Kli70] Raymond A. Kliphardt. Program design in Fortran IV. Allyn and Bacon, Needham Heights, MA, USA, 1970. xvi + 224 pp.

Klimko:1973:ACI

[Kli73] E. M. Klimko. An algorithm for calculating indices in Faa di Bruno’s formula. BIT (Nordisk tidskrift for informationsbehandling), 13:38–49, 1973. CODEN BITTEL. ISSN 0006-3835 (print), 1572-9125 (electronic).

Knuth:1964:FDS

[KM64] D. E. Knuth and J. L. Mc-Nely. Formal definition of SOL. IEEE Transactions on Electronic Computers, EC-13:409–414, August 1964. CODEN IEECA8. ISSN 0367-7508.

Kaiser:1973:FSD

[KM73a] Kenneth W. Kaiser and Marietta A. Marchitelli. A FOR-TRAN subroutine for data processing. Report E-2779, M.I.T. Charles Stark Draper Laboratory, Cambridge, MA, USA, 1973. vi + 26 pp.

Katsanis:1973:FPC

[KM73b] Theodore Katsanis and William D. McNally. FORTRAN program for calculating velocities and streamlines on the hub-shroud midchannel flow surface of an axial- or mixed-flow turbomachine. NASA technical note NASA TN D-734, National Aeronautics and Space Administration, Washington, DC, USA, 1973. ???? pp. For sale by the National Technical Information Service.

Kuester:1973:OTF

[KM73c] James L. Kuester and Joe H.

Mize. Optimization Techniques With Fortran. McGraw-Hill, New York, NY, USA, 1973. ISBN 0-07-035606-8. 500 pp. LCCN QA402.5 .K84. US$73.55. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070356068.

Katsanis:1977:RFPa

[KM77a] Theodore Katsanis and William D.

McNally. Revised FORTRAN program for calculating velocities and streamlines on the hub-shroud midchannel stream surface of an axial-, radial-, or mixed-flow turbomachine or annular duct. NASA technical note NASA TN D-843, National Aeronautics and Space Administration, Washington, DC, USA, 1977. ???? pp. For sale by the National Technical Information Service.

Katsanis:1977:RFPb

[KM77b] Theodore Katsanis and William D.

McNally. Revised FORTRAN program for calculating velocities and streamlines on the hub-shroud midchannel stream surface of an axial-, radial-, or mixed-flow turbomachine or annular duct. NASA technical note NASA TN D-8430, D-8431, National Aeronautics and Space Administration, Washington, DC, USA, 1977. ???? pp. For sale by the National Technical Information Service.

Kuck:1972:NOS

[KMC72] D. J. Kuck, Y. Muraoka, and S.- C. Chen. On the number of operations simultaneously executable in FORTRAN-like programs and their resulting speedup. IEEE Transactions on Computers, C-21: 1293–1310, 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Knight:1976:DPA

[Kni76a] W. R. Knight. Draft proposed ANS FORTRAN. ACM SIGNUM Newsletter, 11(1):41–43, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knight:1976:LEC

[Kni76b] William R. Knight. A letter to the editor concerning draft proposed ANS FORTRAN. ACM SIGNUM Newsletter, 11(1):41–43, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knowlton:1970:EF

[Kno70] P. H. Knowlton. On the extensibility of Fortran. In J. T. Tou, editor, Software Engineering, Proc. Third Symposium on Computer and Information Sciences Held In Miami Beach, Florida, Dec., 1969, pages 225–251. Academic Press, New York, NY, USA, 1970.

Knowlton:1972:RUF

[Kno72] Ken Knowlton. A report on the use of FORTRAN-coded EXPLOR for the teaching of computer graphics and computer art. ACM SIG-PLAN Notices, 7(10):103–112, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knop:1973:AAR

[Kno73] Robert E. Knop. ACM Algorithm 441: Random deviates from the dipole distribution [G5]. Comm. ACM, 16(1):51, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knowlton:1975:MAF

[Kno75a] Ken Knowlton. MINI-EXPLOR — A Fortran-coded version of the EXPLOR language for mini (and large) computers. Computer Graphics, 9(3):31–42, Fall 1975. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Knowlton:1975:MEF

[Kno75b] Ken Knowlton. MINI-EXPLOR — a Fortran-coded version of the EXPLOR language for mini (and large) computers. Computer Graphics, 9(3):31–42, Fall 1975. CODEN CGRADI. ISSN 0097-8930.

Knuth:1962:HWCb

[Knu62] D. E. Knuth. History of writing compilers. In Digest of Technical Papers, ACM 62 National Conference, pages 43, 126. ACM Press, New York, NY 10036, USA, September 1962.

Knuth:1970:ESF

[Knu70] D. E. Knuth. An empirical study of FORTRAN programs. Technical Report CS-186, Stanford Univ., Dept. of Computer Science, Stanford, CA, USA, 1970.

Knuth:1971:ESF

[Knu71] Donald E. Knuth. An empirical study of FORTRAN programs. Software—Practice and Experience, 1(2):105–133, April/ June 1971. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Motivation for optimization.

Kolk:1974:FCP

[Kol74] Martin Leroy Kolk. FORTRAN calculation of proton induced ionization cross sections. Thesis (m.s.), East Texas State University, Commerce, TX, USA, 1974. vii + 38 pp.

Korhecz:1977:KSP

[Kor77] Imre Korhecz. Ketnyelvu szamitogepes programok: ALGOL 60 es FORTRAN programozasi nyelvu programok tervezese, irasa es elemzese. Egyetemi Szamitokozpont, Budapest, Hungary, 1977. ISBN 963-451-020-5. 368 pp. Two volumes.

Kornerup:1979:NRA

[Kor79] Peter Kornerup. A note on rational arithmetic. ACM SIG-MICRO Newsletter, 10(2):28, June 1979. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/
10.1145/1218003.1218004. See

[TG79b, TG79a].

Kotelly:1972:FPD

[Kot72] John Christopher Kotelly. A FOR-TRAN program that derives air temperature, density, and composition as a function of height and exospheric temperature. Environmental research papers 390, Air Force Cambridge Research Laboratories, Air Force Systems Command, USAF, Hanscom AFB, MA, USA, 1972. v + 37 pp.

Kazmier:1970:FEFa

[KP70a] Leonard J. Kazmier and Andreas S. Philippakis. Fundamentals of EDP and FORTRAN: a self-instructional manual. Mc-Graw-Hill, New York, NY, USA, 1970. ISBN 0-07-033416-1. xii + 179 pp.

Kazmier:1970:FEFb

[KP70b] Leonard J. Kazmier and Andreas S. Philippakis. Fundamentals of EDP and Fortran; a selfinstructional manual. McGraw-Hill, New York, NY, USA, 1970. xii + 160 pp.

Kernighan:1976:ST

[KP76] B. W. Kernighan and P. J.

Plauger. Software Tools. Addison-Wesley, Reading, MA, USA, 1976. ISBN 0-201-03669-X. 338 pp. LCCN QA76.6 .K42 1976.

Kernighan:1980:P

[KP80] B. W. Kernighan and P. J. Plauger. Programmierwerkzeuge. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-10419-4. ???? pp.

Kaesler:1963:FIP

[KPG63] Roger L. Kaesler, Floyd W. Preston, and Donald I. Good. Fortran II program for coefficient of association (match-coeff) using an IBM 1620 computer. Kansas Geological Survey Special distribution publication 4, University of Kansas, Lawrence, KS, USA, 1963. 7 pp.

Kieffer:1974:FIP

[KQS74] Susan W. Kieffer, William L. Quirin, and William Steinmetz. FORTRAN IV: programming for the CDC 3300. Open-file report 86-475, Adelphi University, Garden City, NY, USA, 1974. 159 pp.

Krishna:1969:FIP

[KR69] K. Krishna and S. Rajamani. A Fortran II program for analysing a two dimensional field by IBM 1620. Scientific report / India Meteorological Dept. 93, Institute of Tropical Meteorology, Poona, India, 1969. 5 + [4] pp.

Kral:1972:NAP

[Krá72a] J. Král. A new additive pseudorandom number generator for extremely short word-length. Information Processing Letters, 1(4): 164–167, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [?].

Krauss:1972:RCH

[Kra72b] Ray Herbert Krauss. Recognition of constrained hand printed Fortran symbols. Thesis (m.s.), Department of Computer Engineering, Case Western Reserve University, Cleveland, OH, USA, 1972. vii + 71 pp.

Kramer:1974:FPA

[Kra74] William P. Kramer. FORTRAN programs for averaging, filtering, plotting and spectral analysis of current meter observations on the I.B.M. 370/155. GSO technical report 74-1 NML/GSO, University of Rhode Island, Graduate School of Oceanography, Narragansett, RI, USA, 1974. 95 pp.

Kwon:1977:FPR

[KRB77] Byung Doo Kwon, Albert J. Rudman, and Robert F. Blakely. Fortran program for reduction of gravimeter observations to Bouguer anomaly. Geophysical computer program 5; geological survey occasional paper 23, Geological Survey, Bloomington, IN, USA, 1977. 32 pp.

Kwon:1978:FPC

[KRB78] Byung Doo Kwon, Albert J. Rudman, and Robert F. Blakely. FOR-TRAN program for correlation of stratigraphic time series: part 2, power spectral analysis. Geophysical computer program 6; geological survey occasional paper 26, Indiana Geological Survey, Bloomington, IN, USA, 1978. 50 pp.

Kremser:1966:BRW

[Kre66a] H. Kremser. Book report: W. Prager, Introduction to FOR-TRAN. Programming and Numerical Methods. Computing, 1(3): 286, 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kremser:1966:BRDa

[Kre66b] H. Kremser. Book review: D. D. McCracken und W. S. Dorn, Numerical Methods and Fortran Programming. Computing, 1(1): 90, 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Krilanovich:1971:RNF

[Kri71] M. Krilanovich. RFC 119: Network Fortran subprograms, April 21, 1971. URL ftp://

                                        ftp.internic.net/rfc/rfc119.
                                        txt;   ftp://ftp.math.utah.
                                        edu/pub/rfc/rfc119.txt. Sta-

tus: UNKNOWN. Not online.

Krohn:1975:PAC

[Kro75] Howard E. Krohn. A parallel approach to code generation for Fortran like compilers. ACM SIGPLAN Notices, 10(3):146–152, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kellner:1978:IAS

[KRS78] Richard G. Kellner, Theodore N. Reed, and Ann V. Solem. An implementation of the ACM/ SIGGRAPH Proposed Graphics Standard in a multisystem environment. Computer Graphics, 12 (3):308–312, August 1978. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Krumbein:1967:FIC

[Kru67] William Christian Krumbein. Fortran IV computer programs for Markov chain experiments in geology. Computer contribution 13, State Geological Survey, The University of Kansas, Lawrence, KS, USA, 1967. 38 pp.

Krumbein:1968:FIC

[Kru68] William Christian Krumbein. Fortran IV computer program for simulation of transgression and regression with continuous-time Markov models. Computer contribution 26, State Geological Survey, The University of Kansas, Lawrence, KS, USA, 1968. 38 pp.

Kruskal:1969:EPR

[Kru69] J. B. Kruskal. Extremely portable random number generator. Comm. ACM, 12(2):93–94, February 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kennedy:1968:TSF

[KS68] Michael Kennedy and Martin B.

Solomon. Ten statement FOR-TRAN plus FORTRAN IV for the IBM 360 using the WATFOR compiler. Technical report, Department of Computer Science, University of Kentucky, Lexington, KY, USA, 1968. 1 v pp.

Kennedy:1970:TSF

[KS70] Michael Kennedy and Martin B. Solomon. Ten statement Fortran plus Fortran IV for the IBM 360, featuring the WATFOR and WAT-FIV compilers. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1970. ISBN 0-13-903419-6. xvi + 381 pp. LCCN QA76.5 .K39. Second ed. published in 1974 under title: Ten statement Fortran plus Fortran IV.

Kreitzberg:1972:EF

[KS72a] Charles B. Kreitzberg and Ben Shneiderman. Elements of Fortran. Harcourt, Brace, Jovanovich, College and School Division, San Diego, CA, USA, June 1972. ISBN 0-15-522156-6. vi + 121 pp. LCCN QA76.73.F25K74. US$13.30. URL http://www.

                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=0155221566.

Kreitzberg:1972:EFS

[KS72b] Charles B. Kreitzberg and Ben Shneiderman. The elements of Fortran style: techniques for effective programming. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1972. ISBN 0-15-522156-6. vi + 121 pp. LCCN QA76.73.F25 K74.

Kreitzberg:1974:FPS

[KS74] Charles B. Kreitzberg and Ben Shneiderman. FORTRAN programming: a spiral approach. Harcourt, Brace, Jovanovich, San Diego, CA, USA, revision April 1974 edition, 1974. 443 pp.

Kennedy:1975:TSF

[KS75a] Michael Kennedy and Martin B.

Solomon. Ten statement Fortran plus Fortran IV; sensible, modular, and structured programming with WATFOR and WATFIV. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1975. ISBN 0-13-903385-8. xxvii + 579 pp. LCCN QA76.73.F25 K45 1975. First ed. published in 1970 under title: Ten statement Fortran plus Fortran IV for the IBM 360, featuring the WATFOR and WAT-FIV compilers.

Kreitzberg:1975:FPS

[KS75b] Charles B. Kreitzberg and Ben Shneiderman. FORTRAN programming: a spiral approach, with WATFOR/WATFIV and standard FORTRAN. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1975. ISBN 0-15-528012-0. x + 437 pp. LCCN QA76.73.F25 K75 Sci-Eng.

Kunzi:1967:NMM

[KTZ67] H. P. Künzi, H. G. Tzschach, and C. A. Zehnder. Numerische Methoden der mathematischen Optimierung, mit ALGOL- und FORTRAN-Programmen. (German), volume Band 8 of Leitfäden der angewandten Mathematik und Mechanik. Teubner, Stuttgart, Leipzig, 1967. 151 pp.

Kunzi:1968:NMM

[KTZ68] Hans Paul Kunzi, H. G. Tzschach, and C. A. Zehnder. Numerical methods of mathematical optimization with ALGOL and FOR-TRAN programs. Computer science and applied mathematics; a series of monographs and textbooks. Academic Press, New York, NY, USA, June 1968. ISBN 0-12-428850-2. x + 171 pp. LCCN ???? US$53.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124288502.

Kunzi:1971:NMM

[KTZ71] Hans Paul Kunzi, H. G. Tzschach, and C. A. Zehnder. Numerical methods of mathematical optimization with ALGOL and FORTRAN programs. Computer science and applied mathematics. Academic Press, New York, NY, USA, corrected and augmented ed. edition, 1971. ISBN 0-12-428850-2. viii + 219 pp. LCCN QA402.5 .K8131 1971. Translated by Werner C. Rheinboldt and Cornelie J. Rheinboldt.

Kubicek:1973:AAL

[Kub73] Milan Kubicek. ACM Algorithm 470: Linear systems with almost tridiagonal matrix [F4]. Comm. ACM, 16(12):760–761, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1966:CAE

[Kuk66] H. Kuki. Comments on the ANL evaluation of the OS/360 FOR-TRAN math function library. ???? SSD 169, C4773, SHARE Secretary Distribution, ????, 1966. 47– 53 pp.

Kuki:1967:CAE

[Kuk67] H. Kuki. Comments on the ANL evaluation of the OS/360 FORTRAN math function library. SHARE Secretary Distribution, SSD 169(C4773):47–53, 1967.

Kuki:1972:AAC

[Kuk72a] Hirondo Kuki. ACM Algorithm 421: Complex gamma function with error control [S14]. Comm. ACM, 15(4):271–272, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1972:CGF

[Kuk72b] Hirondo Kuki. Complex gamma function with error control [S14]. Comm. ACM, 15(4):262–267, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuo:1972:CAN

[Kuo72] Shan Sun Kuo. Computer Applications of Numerical Methods. Addison-Wesley, Reading, MA, USA, June 1972. ISBN 0-201-03956-7. ???? pp. LCCN ????

Kuo:1973:ALF

[Kuo73] Shan Sun Kuo. Assembler Language for Fortran, Cobol, and Pl/I Programmers: IBM 370/360. Addison-Wesley, Reading, MA, USA, December 1973. ISBN 0-201-03954-0. ???? pp. LCCN ???? US$38.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0201039540.

Kuo:1974:ALF

[Kuo74] Shan S. (Shan Sun) Kuo. Assembler language for FORTRAN, COBOL, and PL/I programmers: IBM 370/360. Addison-Wesley, Reading, MA, USA, 1974. ISBN 0-201-03954-0. xiv + 573 pp. LCCN QA76.73.A8 K86.

Kam:1971:DDS

[KW71] Alan C. H. Kam and Charles F. Wall. DYNA: dynamic storage allocation in FORTRAN for the IBM/ 360 operating system. Research report 53, Dept. of Political Science, University of Hawaii, Honolulu, Honolulu, HI, USA, 1971. i + 13 pp.

Kramer:1975:FGP

[KW75] William P. Kramer and Robert H. Weisberg. FORTRAN graphics programs for physical oceanographic and time series data. Marine technical report 46, University of Rhode Island, Kingston, RI, USA, 1975. 92 pp.

Lagerlof:1974:IRR

[Lag74] Rolf O. E. Lagerlöf. Interpolation with rounded ramp functions. Comm. ACM, 17(8):476–479, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lun:1972:SDH

[lAL72] Kei leung Albert Lun. System design of a hardware Fortran compiler. Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1972. [11] + 138 pp.

Lalande:1975:FPS

[Lal75] Marc E. Lalande. A FORTRAN program to solve the Bretschneider wind-wave relationships for deep water. Technical memoranda TEC 816 Canada, Environmental Canada, Atmospheric Environment Service, Downsview, Ontario, Canada, 1975. 12 pp.

Lambird:1971:SEA

[Lam71a] Robert J. Lambird. String enhancements to ANSI Standard Fortran (part 1). ACM SIGPLAN Notices, 6(1):5–12, January 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lamoitier:1971:LFIa

[Lam71b] Jean-Pierre Lamoitier. Le Langage Fortran IV. Bibliothèque technique Philips. Dunod, Paris, France, 1971. xii + 129 pp.

Lamoitier:1971:LFIb

[Lam71c] Jean-Pierre Lamoitier. Le Langage Fortran IV. Bibliothèque technique Philips. Dunod, Paris, France, 1971. xvi + 248 pp.

LaMotte:1972:MSR

[LaM72] L. R. LaMotte. Miscellanea: The SELECT routines: a program for identifying best subset regressions. Applied Statistics, 21(1):92– 93, 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lamoitier:1974:EPF

[Lam74] Jean-Pierre Lamoitier. Exercices de programmation en Fortran IV.

Dunod technique. Dunod, Paris, France, 1974. ISBN 2-04-008160-7. viii + 144 pp.

Lamoitier:1977:EPF

[Lam77] Jean-Pierre Lamoitier. Exercices de programmation en Fortran IV. Dunod université. Ouvrages fondamentaux (serie orange). Dunod, Paris, France, 1977. ISBN 2-04-009305-2. 148 pp.

Lamoitier:1978:UPP

[Lam78] Jean-Pierre Lamoitier. Uprazhneniia po programmirovaniiu na Fortrane IV. Mir, Moscow, Russia, 1978. 162 pp.

Lansford:1972:PFIa

[Lan72a] Lansford Publishing Company. Programming in Fortran IV, 1972.

Lansford:1972:PFIb

[Lan72b] Lansford Publishing Company. Programming in Fortran IV, 1972.

Lanam:1980:PGE

[Lan80] Douglas H. Lanam. A package for generating and executing Fortran programs with Macsyma. Master of science, plan ii, Dept. of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1980. various pp.

LaPlace:1972:PLP

[LaP72] André M. L. LaPlace. PL/I list processing used as interactive system support. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(22):10–24, March 1972.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lapscher:1978:TDL

[Lap78] Fernand Lapscher. Technique du langage FORTRAN: description et pratique. Collection Méthodes. Hermann, Paris, France, 1978. 328 pp.

Larner:1963:DIP

[Lar63a] R. Larner. Design of an integrated programming and operating system, part IV: The system’s FORTRAN compiler. IBM Systems Journal, 2:311–321, September–December 1963. CO-DEN IBMSA7. ISSN 0018-8670.

Larsson:1963:EAF

[Lar63b] Robert D. (Robert Dustin) Larsson. Equalities and approximations, with Fortran programming. John Wiley and Sons, New York, London, Sydney, 1963. x + 158 pp.

Larmouth:1967:TFRa

[Lar67a] J. (John) Larmouth. The T3 Fortran reference manual. Technical report, University Mathematical Laboratory, Cambridge, UK, 1967. 67 pp.

Larmouth:1967:TFRb

[Lar67b] J. (John) Larmouth. The T3 Fortran reference manual. Technical report, University Mathematical Laboratory, Cambridge, UK, 1967. 67 pp.

Larson:1967:FPM

[Lar67c] Glenn Arthur Larson. A Fortran program for the minimization of sequential logic networks. Thesis (m.s.), George Washington University, Washington, DC, USA, 1967. iv + 76 pp.

Larson:1969:FPA

[Lar69] S. Larson. Fortran programs for analysis and prediction of sea and swell in semiclosed seas and coastal waters. Technical Report ????, Fleet Numerical Weather Central, ????, 1969. 49 pp.

Larmouth:1973:SF

[Lar73a] J. Larmouth. Serious FOR-TRAN. Software—Practice and Experience, 3(2):87–107, April/ June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Larmouth:1973:SFP

[Lar73b] J. Larmouth. Serious FOR-TRAN — part 2. Software—Practice and Experience, 3(3):197–225, July/September 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lasker:1971:SFP

[Las71] Leslie Lasker. SATRAP: a Fortran program for coupled channel calculations using a surface delta interaction potential. Thesis (m.s.), University of Minnesota, Minneapolis, MN, USA, 1971. iii + 106 pp.

Lathrop:1979:UTF

[Lat79] Scott Lathrop. Using a terminal, file usage, Fortran program statements, Fortran control statements,

Lau:1980:FCM

[Lau80] H. T. (Hang Tong) Lau. A FOR-TRAN code for maximum matching in graphs. Technical report CS-80-13, Vanderbilt University, Dept. of Computer Science, Nashville, TN, USA, 1980. 11 pp.

Lawson:1977:BLA

[Law77] Chuck L. Lawson. Basic linear algebra subprograms for FORTRAN usage. Technical Report SAND-77-0898, Sandia Laboratories, Albuquerque, NM, USA, 1977. 38 pp. Available through the National Technical Information Service.

Lawson:1978:BLA

[Law78] C. Lawson. Basic Linear Algebra Subprograms for FORTRAN usage. ACM Transactions on Mathematical Software, 5(3):308–323, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lawson:1979:NT

[Law79] Charles L. Lawson. No title. In ACM [ACM79], pages 38–41.

Lehman:1968:DCF

[LB68] Richard S. Lehman and Daniel Edgar Bailey. Digital computing; Fortran IV and its applications in behavioral science. John Wiley and Sons, New York, London, Sydney, 1968. ISBN 0-471-52400-X. xxiv + 303 pp.

Little:1970:STS

[LB70] T. W. Little and W. A. Boutwell.

Survey tabulation and summarization program FORTRAN IV. Research report 5, Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1970. 41 pp.

LeCureux:1977:MFP

[LB77] Floyd LeCureux and James Burnett. Modularized Fortran programming. ????, ????, 1977. xii + 482 pp.

Lundstrom:1980:CMA

[LB80] Stephen F. Lundstrom and George H. Barnes. A controllable MIMD architecture. In Proceedings of the 1980 International Conference on Parallel Processing, pages 19– 27. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1980.

Lawson:1966:FIS

[LBG66] C. L. Lawson, N. Block, and R. D. Garrett. FORTRAN IV subroutines for contour plotting. Technical memo / Section 314 106, Jet Propulsion Laboratory, Pasadena, CA, USA, 1966. 53 pp.

Lee:1980:FPS

[LBM^{+}80] G. Lee, T. Boreham, B. Minns, F. Smith, and R. Soderstrom. FORTRAN programming standards. ACM SIGPLAN Notices, 15 (2):51–63, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lambird:1975:EIL

[LC75] Robert J. Lambird and Lawrence E. Cornish. The EXEC 8 implementation of List Processor-N. ACM SIGPLAN Notices, 10(10):25–36, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledgard:1978:FSP

[LC78] Henry F. Ledgard and Louis J.

Chmura. Fortran With Style: Programming Proverbs. Hayden computer programming series. Hayden Book Co., Rochelle Park, NJ, USA, March 1978. ISBN 0-8104-5682-6. 164 pp. LCCN QA76.73.F25 .L38. US$14.95.

URL http://www.cbooks.com/
sqlnut/SP/search/gtsumt?source=
&isbn=0810456826.

Leavenworth:1964:FIS

[Lea64] B. M. Leavenworth. FORTRAN IV as a syntax language. Comm. ACM, 7(2):72–80, February 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leach:1967:FPT

[Lea67] Lanse M. Leach. FORTRAN to PL/1 translator: Written in PL/1 on the IBM system/360 using recursive descent and PL/1 character strings as the major tools of the translation. Technical Report 33-78-2, Computation Center, Stanford University, Stanford, CA, USA, 1967. various pp.

LRI:1970:RRM

[Lea70] Leasco Response Incorporated.

Response/360 reference manual for the programming language: FORTRAN. Leasco Response Incorporated, Washington, DC. USA, 1970. various pp.

Leath:1975:FPT

[Lea75] Charles Leath. FORTRAN to PL/1 translator. Masters project in lieu of thesis, University of Missouri — Columbia, Columbia, MO, USA, 1975. 63 pp.

Leathers:1978:SAA

[Lea78] Burton L. Leathers. Statistical algorithms: Algorithm AS 131: Tabulating frequency distributions for variables with structured code sets. Applied Statistics, 27(3):359–362, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.
                                        cmu.edu/apstat/131. See remark

[?].

Leach:1980:FP

[Lea80] John Timothy Leach. A FOR-TRAN 77 preprocessor. Thesis (m.s.), Florida Atlantic University, Boca Raton, FL, USA, 1980. vii + 108 pp.

Lecht:1966:PFC

[Lec66a] Charles Philip Lecht. The programmer’s Fortran 11 and 1V: a complete reference. McGraw-Hill, New York, NY, USA, 1966. xx + 162 pp. With a foreword by Robert Bemer.

Lecht:1966:PFI

[Lec66b] Charles Philip Lecht. The programmer’s Fortran II and IV: a complete reference. McGraw-Hill, New York, NY, USA, 1966. xx + 162 pp.

Lecht:1968:FII

[Lec68] Charles Philip Lecht. FORTRAN II y IV del programador; una referencia completa. Compania Editorial Continental, Mexico. Mexico, 1968. 181 pp.

Ledgard:1975:PPF

[Led75] Henry F. Ledgard. Programming proverbs for FORTRAN programmers. Hayden computer programming series. Hayden Book Co., Rochelle Park, NJ, USA, 1975. ISBN 0-8104-5820-9. 130 pp. LCCN QA76.73.F25,L41.

Lee:1967:AC

[Lee67a] John A. N. Lee. The Anatomy of a Compiler. Reinhold computer science series. Reinhold Pub. Corp., New York, NY, US, 1967. xi + 275 pp. LCCN QA76.5 .L4135.

Lee:1967:SCF

[Lee67b] R. M. Lee. A short course in Fortran IV programming, based on IBM operating System/360, basic Fortran IV. McGraw-Hill, New York, NY, USA, 1967. viii + 235 pp.

Lee:1969:FIP

[Lee69] P. J. Lee. FORTRAN IV programs for canonical correlation and canonical trend-surface analysis. Computer contribution 32, University of Kansas, Lawrence, KS, USA, 1969. 46 pp.

Lee:1972:SCB

[Lee72] Robert M. Lee. A short course in basic Fortran IV programming based on the IBM System/360 and System/370. McGraw-Hill, New York, NY, USA, 1972. ISBN 0-07-036998-4. viii + 239 pp. LCCN QA76.73.F25 L431.

Lee:1974:FPS

[Lee74a] Hoi-Lam Lee. Fortran program for solving the algebraic matrix Riccati equation. Thesis (m.s.), Iowa State University, Ames, IA, USA, 1974. 99 pp.

Lee:1974:AC

[Lee74b] John A. N. Lee. The Anatomy of a Compiler. Computer science series. Van Nostrand Reinhold, New York, NY, USA, second edition, 1974. ISBN 0-442-24733-8. x + 470 pp. LCCN QA76.6 .L37 1974.

Lee:1974:CSP

[Lee74c] Joseph Arnold Lee. A comparative study of programming languages: APL, BASIC, COBOL, FORTRAN. Thesis (m.s.), Marquette University, Milwaukee, WI, USA, 1974. 37 + [31] pp.

Lee:1977:CFP

[Lee77] John A. N. Lee. Considerations for future programming language standards activities. Comm. ACM, 20(11):788–794, November 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leland:1974:FH

[Lel74] Laurence Arthur Leland. Fortran hydro. Thesis (m.s.c.e.), University of Massachusetts, (Amherst, Boston, Dartmouth: which ??) MA, USA, 1974. viii + 75 pp.

Lemke:1975:FFT

[Lem75] Kevin Lemke. FORTE: a FOR-TRAN text editor. Honors program senior thesis, ????, ????, 1975. iii + 75 pp.

Leonard:1974:FMH

[Leo74] Marcia White Leonard. A FOR-TRAN model of human acid-base balance regulatory mechanisms. Thesis (m.m.s.), Brown University, Providence, RI, USA, 1974. iii + 43 pp.

Lepley:1976:SAS

[Lep76] M. Lepley. A simplified algorithm to solve geometric programming problems using FORTRAN. Technical report USERDA/EY-76-S-02-2895*000/Tr/76/6, Carnegie-Mellon University, Pittsburgh, PA, USA, 1976. [4] + 16 [19] pp.

Lerew:1972:FPM

[Ler72] Lloyd Eugene Lerew. A Fortran psychometric model. Thesis (m.s.), Dept. of Agricultural Engineering, Michigan State University of Agriculture and Applied Science, East Lansing, MI, USA, 1972. [1] + vii + [46] pp.

Lesk:1972:GID

[Les72] Arthur M. Lesk. Generation of interactive displays from FORTRAN using the PDP-10/LSD-1 computer graphics system. Software— Practice and Experience, 2(3):259– 273, July/September 1972. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lesk:1973:FPS

[Les73] A. M. Lesk. A FORTRAN program for the solution of simultaneous linear Boolean inequalities by the algorithm of Hammer and Rudeanu. Journal of Computational Physics, 12(1):150–152, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999173901770.

DeLeuw:1979:UTCa

[Leu79a] Cather De Leuw. Urban traffic control system first generation FORTRAN IV overlay software (extended version): executive summary. Report — Federal Highway Administration FHWA-TS-79-222, Dept. of Transportation, Federal Highway Administration, [Office of Research and Development], Office of Development, Implementation Division, Washington, DC. USA, 1979. v + 45 pp.

DeLeuw:1979:UTCb

[Leu79b] Cather De Leuw. Urban traffic control system first generation FORTRAN IV overlay software (extended version): volume 6, program flow charts. Implementation package — Federal Highway Administration FHWA-IP-79-5, Federal Highway Administration, Office of Development Implementation Division, Washington, DC. USA, 1979. 42 pp.

Levenq:1971:FLS

[Lev71] Jean-François Levenq. Le Fortran un langage scientifique des ordinateurs. Science-poche 44. Dunod, Paris, France, 1971. XIV-113 pp.

Lewis:1963:CMF

[Lew63] Theodore S. Lewis. Character manipulation in FORTRAN. Comm. ACM, 6(2):65, February 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lewart:1973:AAA

[Lew73] C. R. Lewart. ACM Algorithm 463: Algorithms SCALE1, SCALE2, and SCALE3 for determination of scales on computer generated plots [J6]. Comm. ACM, 16(10):639–640, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lewis:1979:FPG

[Lew79a] Anthony Lewis. The FORTRAN programmer’s guide. Prime Computer, Inc., Framingham, MA, USA, 1979. 290 pp.

Lewis:1979:PMU

[Lew79b] V. Ellen Lewis, editor. Proceedings of the 1979 MACSYMA Users Conference: [held in] Washington, DC, USA, June 20–22, 1979. Massachusetts Institute of Technology, Laboratory for Computer Science, Cambridge, MA, USA, 1979.

Lewis:1980:FRG

[Lew80a] Anthony Lewis. The FORTRAN reference guide. Prime Computer, Inc., Framingham, MA, USA, 1980. 95 pp.

Lewis:1980:BFS

[Lew80b] Robert Jay Lewis. Business Fortran: a Structured Approach. Wadsworth, Pacific Grove, CA, USA, December 1980. ISBN 0-534-00778-3. ???? pp. LCCN ???? US$30.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534007783.

Lewis:1980:PPP

[Lew80c] William E. Lewis. Problem-solving principles for programmers: applied logic, psychology, and grit. Hayden Book Company, Rochelle Park, N.J., 1980. ISBN 0-8104-5138-7. 163 pp. LCCN QA76.6 .L49.

Lewis:1980:PSP

[Lew80d] William E. Lewis. Problem-Solving Principles for Programmers: Applied Logic, Psychology and Grit. Howard W. Sams, Indianapolis, IN 46268, USA, December 1980. ISBN 0-8104-5138-7. ???? pp. LCCN ????

Loveman:1975:POT

[LF75] David B. Loveman and Ross A. Faneuf. Program optimization — theory and practice. ACM SIGPLAN Notices, 10(3):97–102, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lawson:1978:SPR

[LF78] C. L. Lawson and J. A. Flynn. SFTRAN3 programmer’s reference manual. Technical Report 1846-98, Jet Propulsion Laboratory, Pasadena, CA, USA, December 1978.

Lamoitier:1973:LFI

[LG73] Jean-Pierre Lamoitier and G. Gabrielle.

Le langage Fortran IV. Dunod université. Dunod, Paris, France, 2e edition, 1973. xiii + 261 pp.

Locs:1974:FED

[LG74] G. Locs and J. M. Gary. A Fortran extension for data display. IEEE Transactions on Computers, C-23(12):1257–1263, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672437.

Lamoitier:1978:LFI

[LG78] Jean-Pierre Lamoitier and G. Gabrielle.

Le langage Fortran IV. Dunod université. Dunod, Paris, France, 3e, nouv. tirage edition, 1978. ISBN 2-04-010214-0. xiii + 261 pp.

Loew:1975:FIP

[LGF75] W. Loew, A. P. Gutierrez, and L. A. Falcon. A Fortran IV program to store and retrieve weather data for use in weather driven ecosystems. Technical report, International Center for Biological Control, Univ. of California Berkeley and Riverside, Albany, NY, USA, 1975. 5 + [4] + 24 pp.

Leitch:1965:FIP

[LH65] James L. Leitch and Nancy S. Hagen. A Fortran IV program for the evaluation of post-irradiation survival times. Office of Technical Services, Washington, DC. USA, 1965. various pp.

Hao:1980:FCH

[lH80] Po lin Hao. FORTRAN cheng hsu she chi. Jen min yu tien chu pan she: Hsin hua shu tien Pei-ching fa hsing so fa hsing, Pei-ching, ti 1 pan edition, 1980.

Lawson:1979:ABL

[LHKK79a] C. L. Lawson, R. J. Hanson, D. R. Kincaid, and F. T. Krogh. Algorithm 539: Basic Linear Algebra Subprograms for Fortran usage [F1]. ACM Transactions on Mathematical Software, 5(3):324–325, September 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lawson:1979:BLA

[LHKK79b] C. L. Lawson, R. J. Hanson, D. R. Kincaid, and F. T. Krogh. Basic Linear Algebra Subprograms for Fortran usage. ACM Transactions on Mathematical Software, 5 (3):308–323, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lyon:1980:UAFa

[LHLM80a] Gordon Lyon, Frances E. Holberton, J. (John) Larmouth, and M. D. McIlroy. Using ANS FOR-TRAN. NBS handbook 131, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1980. vi + 100 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Lyon:1980:UAFb

[LHLM80b] Gordon Lyon, Frances E. Holberton, J. (John) Larmouth, and M. D. McIlroy. Using ANS FOR-TRAN. NBS handbook 131, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, 1980. vi + 100 pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Lillyquist:1968:FCT

[Lil68] Michael Jerome Lillyquist. FOR-TRAN computation of a table for the SPAN decision-making method in dyads. Thesis (m.a.), University of Arizona, Tucson, AZ, USA, 1968. 88 pp.

Lill:1971:NAM

[Lil71] Shirley A. Lill. Note on Algorithm 46: a modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 14(1):106, February 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_01/tiff/106.tif. See

[?, ?].

Limited:1978:ID

[Lim78] International Computers Limited.

Introduction to DAP-Fortran. Document Number AP20, July 1978.

Lindsay:1971:RAF

[Lin71] D. S. Lindsay. A rounded arithmetic FORTRAN compiler for CDC 6000 machines. Report ????, Department of Computer Science, University of California, Berkeley, CA, USA, December 1971. ???? pp.

Linz:1972:AAF

[Lin72] Peter Linz. ACM Algorithm 427: Fourier cosine integral [D1]. Comm. ACM, 15(5):358–360, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Linder:1976:CTS

[Lin76] William H. Linder. COMPUTER-TUTOR: From a student project to a self-paced CAI/CMI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):57–60, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Lipschutz:1977:PF

[Lip77] Seymour Lipschutz. Programming With Fortran. McGraw-Hill, New York, NY, USA, September 1977. ISBN 0-07-037984-X. ???? pp. LCCN ???? US$12.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
007037984X.

Lipschutz:1978:PFS

[Lip78] Seymour Lipschutz. Programming With Fortran (Schaum’s Outlines). McGraw-Hill, New York, NY, USA, May 1, 1978. ISBN 0-07-037984-X. 314 pp. LCCN ???? US$12.95.

Little:1974:DIA

[Lit74] Cindy E. Little. The design and implementation of an additional access method for the FORTRAN programmer. Thesis (m.s.), University of Arkansas, Fayetteville, Fayetteville, AR, USA, 1974. 180. pp.

Luter:1971:MOCb

[LJ71a] James G. Luter and Marvin L.

Johnson. Mathematics oriented computer programming, with basic FORTRAN IV. Anaheim Pub. Co., Fullerton, CA, USA, 1971. various pp.

Luter:1971:MOCa

[LJ71b] James G. Luter and Marvin L.

Johnson. Mathematics oriented computer programming with basic FORTRAN IV: instructor’s guide and answer manual. Anaheim Pub. Co., Fullerton, CA, USA, 1971. 140 pp.

Ljungkvist:1980:PEF

[Lju80] Sten Ljungkvist. PASCAL and existing FORTRAN files. ACM SIGPLAN Notices, 15(5):54–55, May 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Law:1974:ICU

[LK74] Victor J. Law and Daniel B.

Killeen. Introduction to computing using the FORTRAN language. Technical report, Tulane University, New Orleans, LA, USA, 1974. viii + 276 pp.

Larsen:1965:MSD

[LL65] Robert P. Larsen and M. Morris Larson. Modeling and simulation of digital networks. Comm. ACM, 8(5):308–312, May 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lowry:1969:OCO

[LM69] Edward S. Lowry and C. W. Medlock. Object code optimization. Comm. ACM, 12(1):13–22, January 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liddiard:1970:MMF

[LM70] Lawrence Liddiard and E. James Mundstock. MNF (MiNnesota FORTRAN) reference manual: for CDC 6000/7000/Cyber series computers. Technical report, University of Minnesota, Minneapolis, MN, USA, 1970. various pp.

Liu:1976:SSD

[LM76] Philip S. Liu and Frederic J. Mowle. Selection schemes for dynamically microcoding Fortran programs. ACM SIGARCH Computer Architecture News, 4(4): 122.6, January 1976. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Laurila:1969:FPT

[LML69] Simo H. Laurila, Surendra Pratap Mathur, and L. Li. Fortran programs for transformation between rectangular and geographic map coordinates. HIG report 69-5, Hawaii Institute of Geophysics, University of Hawaii, Honolulu, HI, USA, 1969. v + 31 pp.

Lanzarone:1977:IAP

[LMP77] Gaetano A. Lanzarone, Marco Maiocchi, and Roberto Polillo. Introduzione alla programmazione strutturata: il caso del Fortran, Cobol e Assembler. Collana dei “Quaderni di informatica”; 4. F. Angeli, Milano, Italy, 1977. 221 pp.

Lutz:1977:FIL

[LO77] Peter H. Lutz and Daniel E. Oldman. FOIL, and intermediate language for FORTRAN version 1.0. Technical report 128, State University of New York at Buffalo, Dept. of Computer Science, Buffalo, NY, USA, 1977. 29 + 5 + 7 + 9 + 6 pp.

Loeckx:1974:ALP

[Loe74] Jacques Loeckx, editor. Automata, Languages and Programming, number 14 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., August 1974.

Lott:1971:FFC

[Lot71] Richard W. Lott. Fortran: a first course. Granite State Publishers, Nashua, NH, USA, 1971. 136 pp.

Loudon:1967:RPD

[Lou67] Thomas Victor Loudon. The Rokdoc package: description and listing of a library of routines in Fortran 4 for statistical analysis, summary and display of data concerning sedimentary rocks. Computation in sedimentology, no. 5; University of Reading geological reports, no. 1, Sedimentology Research Laboratory, University of Reading, Reading, UK, 1967. 152 pp.

Louis:1973:DNL

[Lou73] F. Louis. Calcul des dérivées successives d’une fonction d’une variable réelle. applications aux développements de Taylor, de Bürmann et de Thiele. cours donné dans le cadre du programme d’enseignement académique du CERN, 1969–1970, CERN 73-17. In Éléments d’analyse numérique et leur programmation en FORTRAN IV. Première partie. (French), pages vi + 61. CERN—Organisation Européenne pour la Recherche Nucléaire, Geneva, Switzerland, 1973.

Loudon:1974:AGD

[Lou74] T. V. Loudon. Analysis of geological data using ROKDOC, a Fortran IV package for the IBM 360/ 65 computer. Institute of Geological Sciences. Report no. 74/1 Report (Institute of Geological Sciences (Great Britain)); 74/1. H.M. Stationery Office, London, UK, 1974. iv + 131 pp.

Love:1968:FPD

[Lov68] Wesley Dwight Love. A FOR-TRAN program for the discrete Fourier analysis of aperiodic waveforms. Thesis (m.s.), Virginia Polytechnic Institute, Blacksburg, VA, USA, 1968. 40 pp.

Lovas:1975:ICP

[Lov75] Charles M. Lovas. An instructional computer program library. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1): 129–132, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Lowery:1976:FCE

[Low76] Janice R. Long Lowery. A Fortran compiler efficiency meter with application to the improvement of program speeds at the source level. Thesis (m.s.), Dept. of Computer Science, College of Natural Sciences and Mathematics, University of Houston, Houston, TX, USA, 1976. 73 pp.

Liddell:1971:IFP

[LP71] Heather Mary Liddell and Anthony J. Powell. Introduction to FORTRAN programming. Engineering science monographs. Harrap, London, UK, 1971. ISBN 0-245-50522-9 (hardcover), 0-245-50529-6 (paperback). vii + 2–154 pp. LCCN QA76.73.F25L53 1971.

Land:1973:FCM

[LP73] Ailsa H. (Ailsa H.) Land and S. (Susan) Powell. Fortran Codes for Mathematical Programming: Linear, Quadratic and Discrete. John Wiley and Sons, New York, London, Sydney, June 1973. ISBN 0-471-51270-2. 249 pp. LCCN QA402.5.L33. US$83.95. URL http://www.

                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=0471512702.

Lentini:1974:VOF

[LP74] M. Lentini and V. Pereyra. A variable order finite difference method for nonlinear multipoint boundary value problems. Mathematics of Computation, 28(128): 981–1003, October 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lipschutz:1978:SOT

[LP78] Seymour Lipschutz and Arthur Poe. Schaum’s outline of theory and problems of programming with Fortran. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 1978. ISBN 0-07-037984-X. 314 pp. LCCN QA 76.73 F25L57.

Lipschutz:1979:TPP

[LP79] Seymour Lipschutz and Arthur Poe. Teoria y problemas de programacion con Fortran; incluye Fortran estructurado. Serie de compendios Schaum. McGraw-Hill, New York, NY, USA, 1979. ISBN 968-451-017-9. 314 pp.

Lipschutz:1979:PFTa

[LPJ79a] Seymour Lipschutz, Arthur Poe, and Sylvie Jacoud. Programmation Fortran, théorie et applications. Serie Schaum. MacGraw-Hill, Ediscience, New York, Paris, 1979. xi + 314 pp.

Lipschutz:1979:PFTb

[LPJ79b] Seymour Lipschutz, Arthur Poe, and Sylvie Jacoud. Programmation FORTRAN: théorie et applications [375 problemes resolus]. Série Schaum; Schaum’s outline series. Ediscience-McGraw Hill, Auckland, New Zealand, 1979. ISBN 2-7042-1013-6. ix + 314 pp.

Lynch:1977:CTI

[LR77] Robert Emmett Lynch and John Rischard Rice. Computers, Their Impact and Use: Structured Programming in Fortran. Holt, Rinehart, and Winston, New York, NY, USA, January 1977. ISBN 0-03-088525-6. ix + 453 pp. LCCN QA76 .L893. US$12.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0030885256.

Larkin:1971:DFA

[LS71a] P. A. (Peter Anthony) Larkin and W. J. Douglas Stephen. Du formol au Fortran; la biologie au Canada. Conseil des sciences du Canada. Étude spéciale 18, Information Canada, Ottawa, Ontario, Canada, 1971. 87 pp.

Larkin:1971:FFS

[LS71b] P. A. (Peter Anthony) Larkin and W. J. Douglas Stephen. From formalin to Fortran; some facts and figures about basic biology in Canada. Background study for the Science Council of Canada. Science Council of Canada. Special study 18, Information Canada, Ottawa, Ontario, Canada, 1971. 79 pp.

Lyon:1975:STI

[LS75] Gordon Lyon and Rona B. Stillman. Simple transforms for instrumenting FORTRAN decks. Software—Practice and Experience, 5 (4):347–358, October/December 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luckmann:1976:NFA

[LS76] M. Luckmann and R. Steinbrueggen. Non-rational functions an additional SAC-1 subsystem. Technical report, Technische Universität München (??), München, Germany, 1976.

Liu:1973:AAE

[LT73] C. N. Liu and D. T. Tang. ACM Algorithm 452: Enumerating combinations of m out of n objects [G6]. Comm. ACM, 16(8):485, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Larson:1975:MMP

[LT75] Gale L. Larson and Fred J. Taylor. MPLIB = a mathematical program library for optimization techniques with FORTRAN. Reference manual 75-2, Dept. of Electrical Engineering for Computer Center, the University of Texas at El Paso, El Paso, TX, USA, 1975. 31 pp.

Lancaster:1976:RF

[LT76] Ronald L. Lancaster and Richard T. Thomas. Reading Fortran.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(4):6–7, December 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Larrabee:1980:FPC

[LTB80] R. D. Larrabee, W. Robert Thurber, and W. Murray Bullis. A FORTRAN program for calculating the electrical parameters of extrinsic silicon. Semiconductor measurement technology NBS special publication 400-63, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC, USA, October 1980. iv + 48 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Ludeman:1969:PUP

[Lud69] M. M. Ludeman. PL/I utility programs for Formac applications. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (12):53–58, July 1969. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Luengo:1966:IPL

[Lue66] R. Luengo. Introduction a la programmation en language Fortran IV. Technical report, Faculté des sciences, Orsay, France, 1966. 48 pp.

Lumsden:1977:FPI

[Lum77] Donald Fred Lumsden. A FOR-TRAN preprocessor for the implementation of modular organization. Technical report, Computer Science Dept., University of Missouri, Columbia, MO, USA, 1977. 50 pp.

Lurie:1973:SFI

[LV73] D. Lurié and C. Vandoni. Statistics for FORTRAN identifiers and scatter storage techniques. Software—Practice and Experience, 3 (2):171–177, April/June 1973. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Locs:1977:FPN

[LV77] Gyula Locs and J. Vigassy. A FORTRAN programozasi nyelv. Muszaki Konyvkiado, Budapest, Hungary, 4. kiad edition, 1977. ISBN 963-10-1808-3. 335 pp.

Ledley:1966:FIP

[LW66] Robert Steven Ledley and James Bruce Wilson. Fortran IV Programming. McGraw-Hill, New York, NY, USA, 1966. ISBN 0-07-036973-9. x + 229 pp. LCCN ???? US$6.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0070369739.

Lyczak:1980:EPS

[Lyc80] Richard A. Lyczak. Elementary programming for statistics: FOR-TRAN and BASIC programming for statistical analysis. Duxbury Press, North Scituate, MA, USA, July 1980. ISBN 0-87872-230-0. x + 197 pp. LCCN QA276.4.L93.

Lynah:1963:FLP

[Lyn63] Francis Pelzer Lynah. FORTRAN language program for the calculation of droplet trajectories. Thesis (m.s.), Drexel University, Philadelphia, PA, USA, 1963. 31 pp.

Lyon:1974:FA

[Lyo74] Gordon Lyon. A FORTRAN analyzer. NBS technical note 849, U. S. National Bureau of Standards, Washington, DC. USA, 1974. iii + 23 pp. For sale by the Supt. of Docs., U. S. Govt. Print. Off.

Lyon:1980:UAFc

[Lyo80] Gordon Lyon. Using ANS Fortran. NIST handbook 131, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1980. vi + 100 pp. For sale by National Technical Information Service.

Lytle:1975:PLL

[Lyt75] Paul William Lytle. A procedure for linking linear programming to Fortran for past-optimal budgeting. Staff paper 9, Dept. of Agricultural Economics, University of Nebraska, Lincoln, NE, USA, 1975. 7. pp.

Matthews:1978:FKD

[MA78] Richard H. Matthews and James O.

Alben. Fitting kinetic data for two independent saturable terms by Multifit II, a general purpose curve fitting program in FOR-TRAN IV. Journal of Theoretical Biology, 72(2):219–230, May 22, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0022519378900905.

MacGowan:1964:FST

[Mac64] Roger A. MacGowan. FOR-TRAN subroutines for time series data reduction. Comm. ACM, 7 (3):153–157, March 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

MacKinnon:1967:FPS

[Mac67] Mary Matilda Sonntag MacKinnon. A FORTRAN program for the SPAN technique of group decision making. Thesis (m.a.), University of Arizona, Tucson, AZ, USA, 1967. 98 pp.

MacNeilage:1968:DGTa

[Mac68a] D. C. MacNeilage. DISPLAY — a guide to the TRACK FORTRAN debugging system. Rand Corporation. Research memorandum RM-5618, Rand Corporation, Santa Monica, CA, USA, 1968. 34 pp.

MacNeilage:1968:DGTb

[Mac68b] D. C. MacNeilage. DISPLAY, a guide to the TRACK FOR-TRAN debugging system. Research memorandum RM-5618-PR, Rand Corp., Santa Monica, CA, USA, 1968. ix + 34 pp.

Macdonald:1969:FPS

[Mac69] P. D. M. Macdonald. Fortran programs for statistical estimation of distribution mixtures; some techniques for statistical analysis of length-frequency data. Fisheries Research Board of Canada. Technical report 129, Fisheries Research Board of Canada, Ottawa, Ontario, Canada, 1969. iii + 45 pp.

MacCracken:1970:FTA

[Mac70a] Daniel D. MacCracken. Fortran in der technischen Anwendung [Fortran with engineering applications]. Ein Lehrbuch in 29 Fallstudien aus der Praxis. Carl Hanser, München, Germany, 1970. 319 pp.

Macleod:1970:SFI

[Mac70b] I. A. Macleod. SP/1 — a FOR-TRAN integrated string processor. The Computer Journal, 13 (3):255–260, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_03/130255.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/255.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/256.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/
257.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_03/tiff/259.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
03/tiff/260.tif.

Macleod:1971:MFM

[Mac71] I. A. Macleod. MP/1 — a FORTRAN macroprocessor. The Computer Journal, 14(3): 229–231, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

                                        URL http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        14/Issue_03/140229.sgm.abs.
                                        html;   http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        14/Issue_03/tiff/229.tif;
                                        http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_14/Issue_
                                        03/tiff/230.tif;   http://www3.
                                        oup.co.uk/computer_journal/
                                        hdb/Volume_14/Issue_03/tiff/
                                        231.tif.

MacKimmie:1973:VWP

[Mac73] Christine K. MacKimmie. A van Wijngaarden property grammar for basic Fortran. Thesis (m.sc.), University of Calgary, Calgary, Alberta, Canada, 1973. iii + 93. pp.

Macnab:1974:F

[Mac74] Donald S. Macnab. Fortran. Blackie, Glasgow, Scotland, 1974. ISBN 0-216-89766-1, 0-216-89682-7 (educational ed.). ix + 214 pp. LCCN QA76.73.F25M33.

Maghsoodloo:1971:IHS

[Mag71] S. Maghsoodloo. An investigation by high speed sampling of the frequency distribution of rank correlation. Computing, 8(1–2): 1–12, 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Malcolm:1970:AFA

[Mal70] Michael Malcolm. An algorithm for floating-point accumulation of sums with small relative error. Technical Report STAN-CS-70-163, Computer Science Department, Stanford University, Stanford, CA, USA, 1970. 21 pp.

Malcolm:1972:ARP

[Mal72] Michael A. Malcolm. Algorithms to reveal properties of floating-point arithmetic. Comm. ACM, 15(11):949–951, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [?].

Malkosh:1977:IPP

[Mal77] Menachem Malkosh. Internal procedure parameters in structured FORTRAN precompilers. ACM SIGPLAN Notices, 12(9):45–51, September 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mansoori:1963:FPC

[Man63] Iraj D. Mansoori. A FORTRAN program for a class of dynamic programming problem: research project. Thesis (m.s. in electrical engineering), University of California, Berkeley, Berkeley, CA, USA, September 1963. various pp.

Mancino:1964:CFC

[Man64] O. G. Mancino. Characteristics of the FORTRAN CEP language. Comm. ACM, 7(7):423–424, July 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mann:1969:IFP

[Man69] Richard A. Mann. An IBM 1130 Fortran primer. International Texbook Co., Scranton, PA, USA, 1969. ISBN 0-7002-2237-5. vii + 216 pp. LCCN QA76.8.I125 M3.

Maniotes:1971:BFSa

[Man71] John. Maniotes. Beginning Fortran: Simplified, 12-Statement Programming. Howard W. Sams, Indianapolis, IN 46268, USA, June 1971. ISBN 0-8104-5870-5. LCCN ???? US$10.35. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0810458705.

Manifold:1972:CF

[Man72a] George O. Manifold. Calculating with Fortran. Charles E. Merrill Publishing Co., Columbus, OH, USA, 1972. ISBN 0-675-09181-0. x + 190 pp. LCCN QA 76.73 F25 M35 General Coll.

Mann:1972:FIP

[Man72b] Richard A. Mann. A FORTRAN IV primer. Intext Educational Publishers, New York, NY, USA, 1972. ISBN 0-7002-2412-2. vii + 207 pp. LCCN QA76.73.F25 M36.

Mann:1974:IFP

[Man74] Richard A. Mann. An IBM 1130 Fortran primer. Intext Educational Publishers, New York, NY, USA, second edition, 1974. ISBN 0-7002-2455-6. xii + 276 pp. LCCN QA76.8.I125M3 1974.

Marill:1966:CFI

[Mar66] Edgardo Julio Marill. A conversational FORTRAN interpreter for the IBM system/360 time-sharing system. Thesis (m.s. in engin.), University of Florida, Gainesville, FL, USA, 1966. iv + 64 pp.

Marlow:1971:FSS

[Mar71] S. Marlow. Fortran subroutines for the solution of linear equations, inversion of matrices and evaluation of determinants. AERE, R 6899 AERE-R (Series); 6899. Theoretical Physics Division, Atomic Energy Research Establishment, Harwell, Berkshire, UK, 1971. ISBN 0-7058-0131-4. 61 pp.

Marateck:1977:F

[Mar77a] Samuel L. Marateck. Fortran. Academic Press, New York, NY, USA, June 1977. ISBN 0-12-470460-3. xvi + 671 pp. LCCN QA76.73.F25 M37. US$19.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0124704603.

Marateck:1977:IMF

[Mar77b] Samuel L. Marateck. Instructor’s manual for FORTRAN. Academic Press, New York, NY, USA, 1977. 72 pp.

Mark:1978:FPS

[Mar78a] Robert K. Mark. FORTRAN program for Shapiro-Wilk W test for normality for Honeywell Multics System. Open-file series 78- 1069, U.S. Geological Survey, Menlo Park, CA, USA, 1978. 10 pp.

Martin:1978:ESF

[Mar78b] John Joseph Martin. Enhancement of SPARKS, a FORTRAN preprocessor. Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1978. ii + 133 pp.

Martin:1980:FBS

[Mar80] Edley Wainright Martin. Fortran for Business Students: a Programmed Instruction Approach. John Wiley and Sons, New York, London, Sydney, October 1980. ISBN 0-471-04622-1. ???? pp. LCCN ???? US$50.40. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0471046221.

MITCC:1960:AFS

[Mas60] Massachusetts Institute of Technology.Computation Center. Additions to the FORTRAN source language. Memorandum CC-153, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 12 pp.

MITCCL:1962:SPR

[Mas62] Massachusetts Institute of Technology.Cooperative Computing Laboratory. Symbol pattern recognition within Fortran: the SHADOW IV system. Technical report, MIT Cooperative Computing Laboratory, Cambridge, MA, USA, 1962. 52 pp.

MUCC:1971:FUU

[Mas71] Massachusetts University Computing Center. FORTRAN: UMASS user’s manual. Technical report, University Computing Center, Amherst, MA, USA, 1971. various pp.

Mather:1972:FIS

[Mat72a] Paul M. Mather. FORTRAN IV subprogram for the simultaneous varimax rotation in factor analysis. Computer applications in the natural and social sciences 13, section B, Computer Applications, Nottingham, UK, 1972. B 6 + II-4 pp.

Mathur:1972:BDC

[Mat72b] F. P. Mathur. A brief description and comparison of programming languages FORTRAN, ALGOL, COBOL, PL/I, and LISP 1.5 from a critical standpoint. Technical report, Jet Propulsion Laboratory, California Inst. of Technology, Pasadena, CA, USA, 1972. 13. pp.

Mathews:1989:SCA

[Mat89] J. Mathews. Symbolic computational algebra applied to Picard iteration. Mathematics and Computer Education, 23(2):117–122, Spring 1989. CODEN MCEDDA. ISSN 0730-8639.

Maurer:1972:PIL

[Mau72a] W. D. Maurer. The Programmer’s Introduction to Lisp. American Elsevier, New York, NY, USA, 1972. ISBN 0-444-19572-6. LCCN QA76.73.L23 M38.

Maurer:1972:SEB

[Mau72b] W. D. Maurer. A semantic extension of BNF. INTJC3, Sect. A, 3 (2):157–176, September 1972.

Maurer:1977:TPC

[Mau77] W. D. Maurer. The teaching of program correctness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):142–144, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Maynard:1972:CPM

[May72] Jeff Maynard. Computer programming made simple. W. H. Allen, London, 1972. ISBN 0-491-00882-1 (hardcover), 0-491-00872-4 (paperback). xix + 300 pp. LCCN QA76.6.M393.

May:1973:PBA

[May73a] Phillip T. May. Programming Business Applications in Fortran IV. Houghton Mifflin, Boston, MA, USA, June 1973. ISBN 0-395-14047-1. vii + 462 pp. LCCN HF5548.5.F2 M35. US$24.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0395140471.

Mayes:1973:BRB

[May73b] D. G. Mayes. Book review: More Fortran Programs for Economists, by Lucy Joan Slater. Journal of the Royal Statistical Society. Series A (General), 136(2):265–266, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2345121.

Mazur:1977:FPD

[Maz77] Jacob Mazur. A FORTRAN program for desmearing small-angle X-ray scattering curves. NBS technical note 936, Dept. of Commerce, National Bureau of Standards, Institute for Materials Research, Washington, DC. USA, 1977. iv + 77 + [1] pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Mazlack:1978:UIF

[Maz78] Lawrence J. Mazlack. The use of interactive Fortran. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):260–265, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Marsaglia:1968:OLRa

[MB68a] George Marsaglia and T. A. Bray.

One-line random number generators and their use in combinations. Report ??, Boeing Scientific Research Laboratories, Seattle, WA, USA, March 1968. 12

pp. URL http://www.dtic.mil/
docs/citations/AD0667956.

Marsaglia:1968:OLRb

[MB68b] George Marsaglia and T. A. Bray.

One-line random number generators and their use in combinations. Comm. ACM, 11(11): 757–759, November 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCrackin:1964:FPA

[MC64] Frank L. McCrackin and James P.

Colson. A Fortran program for analysis of ellipsometer measurements and calculation of reflection coefficients from thin films. NBS technical note 242, Dept. of Commerce, National Bureau of Standards: U.S.G.P.O., Washington, DC. USA, 1964. ii + 42 pp.

McNally:1970:FPC

[MC70] William D. McNally and James E. Crouse. Fortran program for computing coordinates of circular arc single and tandem turbomachinery blade sections on a plane. NASA technical note NASA TN D-6020, National Aeronautics and Space Administration, Washington, DC. USA, 1970. iii + 46 pp. For sale by the Clearinghouse for Federal Scientific and Technical Information, Springfield, VA, USA.

More:1980:ABA

[MC80a] J. J. Moré and M. Y. Cosnard. Algorithm 554: BRENTM, A FOR-TRAN subroutine for the numerical solution of nonlinear equations. ACM Transactions on Mathematical Software, 6(2):240–251, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1980:ABF

[MC80b] J. J. Moré and M. Y. Cosnard. Algorithm 554: BRENTM, A Fortran subroutine for the numerical solution of nonlinear equations [F5]. ACM Transactions on Mathematical Software, 6(2):240–251, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1980:BAF

[MC80c] J. J. Moré and M. Y. Cosnard. BRENTM: A Fortran subroutine for the numerical solution of systems of nonlinear equations. ACM Transactions on Mathematical Software, 6(2):240–251, June 1980.

McAdams:1977:SEFa

[McA77a] Sister Joseph Kieran McAdams. A study of the effectiveness of FRT-DAP, a CAI drill-and-practice system, used in conjunction with video-tapes in teaching Fortran. Computer science thesis (Ph.D.), University of Missouri — Rolla, Rolla, MO, USA, 1977. 174 pp.

McAdams:1977:SEFb

[McA77b] Sister Joseph Kieran McAdams.

A study of the effectiveness of FRTDAP, a CAI drill-and-practice system, used in conjunction with video-tapes in teaching FOR-TRAN. Technical report, University of Missouri, Rolla, MO, USA, 1977. viii + 163 pp.

McCowan:1962:VDP

[MCB^{+}62] D. W. (Douglas W.) McCowan, F. J. A. Corbató, Michael J. Bailey, Kalon L. Kelley, and T. H. Dupree. A visual display package for use within Fortran. Technical note 20, Cooperative Computing Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1962. 18 pp.

McCracken:1961:GFP

[McC61] Daniel D. McCracken. A guide to FORTRAN programming. John Wiley and Sons, New York, London, Sydney, 1961. viii + 88 pp. LCCN QA76.5 .M17 1961.

McCracken:1962:GFP

[McC62] Daniel D. McCracken. A Guide to FORTRAN programming. John Wiley and Sons, New York, London, Sydney, 1962. 88 pp.

McCracken:1963:PF

[McC63] Daniel D. McCracken. Programacion FORTRAN. Editorial Limusa, Mexico, DF, Mexico, 1963. ISBN 968-18-0184-9. 112 pp.

McCormack:1964:FPW

[McC64a] C. McCormack. Fortran programs written in 1964 in NSF Summer Institute. Adelphi College, Dept. of Graduate Mathematics, Garden City, NY, USA, 1964. various pp.

McCracken:1964:GFP

[McC64b] Daniel D. McCracken. A guide to FORTRAN programming. John Wiley and Sons, New York, London, Sydney, 1964. 88 pp.

McCracken:1965:GFI

[McC65a] Daniel D. McCracken. A guide to Fortran IV programming. John Wiley and Sons, New York, London, Sydney, 1965. viii + 151 pp.

McCracken:1965:GFP

[McC65b] Daniel D. (Daniel Delbert) Mc-Cracken. A guide to Fortran 1V programming. John Wiley and Sons, New York, London, Sydney, 1965. 151 pp.

McCalla:1967:INM

[McC67a] Thomas Richard McCalla. Introduction to Numerical Methods and Fortran Programming. John Wiley and Sons, New York, London, Sydney, January 1967. ISBN 0-471-58125-9. xiii + 359 pp. LCCN ???? US$27.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471581259.

McCracken:1967:FEA

[McC67b] Daniel D. McCracken. Fortran With Engineering Applications. John Wiley and Sons, New York, London, Sydney, June 1967. ISBN 0-471-58236-0. x + 237 pp. LCCN ???? US$26.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471582360.

McCracken:1967:GFI

[McC67c] Daniel D. McCracken. A guide to Fortran IV programming. John Wiley and Sons, New York, London, Sydney, 1967. viii + 151 pp.

McCameron:1968:FLP

[McC68a] Fritz A. McCameron. Fortran; logic and programming. R. D. Irwin, Homewood, IL, USA, 1968. x + 246 pp.

McCammon:1968:UF

[McC68b] Mary McCammon. Understanding FORTRAN. Crowell, New York, NY, USA, 1968. ix + 370 pp.

McCammon:1969:FIP

[McC69a] Richard B. McCammon. FOR-TRAN IV program for nonlinear estimation. Number 34 in Computer contribution. State Geological Survey, Lawrence, KS, USA, 1969. 20 pp.

McCrackin:1969:FPA

[McC69b] Frank L. McCrackin. A Fortran program for analysis of ellipsometer measurements. National Bureau of Standards. Technical note 4, U.S. Govt. Print. Office, Washington, DC. USA, 1969. iv + 79 pp.

McCameron:1970:FI

[McC70a] Fritz A. McCameron. FORTRAN IV. The Irwin-Dorsey series in information processing. R. D. Irwin, Homewood, IL, USA, 1970. x + 215 pp.

McCameron:1970:SMF

[McC70b] Fritz A. McCameron. Solutions manual for Fortran IV. R. D. Irwin, Homewood, IL, USA, 1970. v + 145 pp.

McConalogue:1971:AAI

[McC71] D. J. McConalogue. Algorithm 67: An axis invariant procedure to provide slopes at points by local three-point parametric curve fitting. The Computer Journal, 14(2):209, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_02/tiff/209.tif. See

note [?].

McCracken:1972:GFI

[McC72a] Daniel D. McCracken. A Guide to Fortran IV Programming. John Wiley and Sons, New York, London, Sydney, second edition, June 1972. ISBN 0-471-58281-6. xiii + 288 pp. LCCN QA76.5 .M168 1972. US$40.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471582816.

McCracken:1972:GFP

[McC72b] Daniel D. McCracken. A Guide to Fortran Programming. John Wiley and Sons, New York, London, Sydney, August 1972. ISBN 0-471-58212-3. ???? pp. LCCN ???? US$17.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471582123.

McCracken:1973:PFI

[McC73] Daniel D. McCracken. Programacion Fortran IV. Editorial Limusa, Mexico, DF, Mexico, 1973. 165 pp.

McCameron:1974:FI

[McC74a] Fritz A. McCameron. Fortran IV.

Irwin-Dorsey information processing series. R. D. Irwin, Homewood, IL, USA, revised edition, June 1974. ISBN 0-256-01582-1. xvi + 319 pp. LCCN QA 76.73 F25M12 1974. US$15.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0-256-01582-1.

McCoy:1974:TRC

[McC74b] Kevin C. McCoy. Testing the readability and composability of computer programs: methodology. ACM SIGPLAN Notices, 9(10):28, October 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCracken:1974:IMD

[McC74c] Daniel D. McCracken. Instructor’s manual for Daniel D. McCracken, A simplified guide to FORTRAN programming. John Wiley and Sons, New York, London, Sydney, 1974. ISBN 0-471-58293-X. 79 pp.

McCracken:1974:PF

[McC74d] Daniel D. McCracken. Programacion Fortran. Editorial Limusa, Mexico, DF, Mexico, 1974. 112 pp.

McCracken:1974:SGF

[McC74e] Daniel D. McCracken. A Simplified Guide to Fortran Programming. John Wiley and Sons, New York, London, Sydney, June 1974. ISBN 0-471-58292-1. 277 pp. LCCN QA76.73.F25 M32. US$38.85. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0-471-58292-1.

McCuen:1975:FPC

[McC75] Richard H. McCuen. Fortran Programming for Civil Engineers. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1975. ISBN 0-13-329417-X. xv + 448 pp. LCCN TA345.M32. US$34.00. URL http://www.

                                        cbooks.com/sqlnut/SP/search/
                                        gtsumt?source=&isbn=013329417X.

McCameron:1978:F

[McC78a] Fritz A. McCameron, editor. Fortran 4. R. D. Irwin, Homewood, IL, USA, third edition, October 1978. ISBN 0-256-02036-1. ???? pp. LCCN ???? US$16.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0256020361.

McCameron:1978:FI

[McC78b] Fritz A. McCameron. FORTRAN IV. The Irwin series in information and decision science. R. D. Irwin, Homewood, IL, USA, third edition, 1978. ISBN 0-256-02037-X. xvi + 375 pp.

McCarthy:1978:HL

[McC78c] John McCarthy. History of LISP.

ACM SIGPLAN Notices, 13(8): 217–223, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McConalogue:1978:ACI

[McC78d] D. J. McConalogue. Algorithm 102: Convolution integrals involving probability distribution functions. The Computer Journal, 21(3):270–272, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.
oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/
270.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/271.
tif;   http://www3.oup.co.uk/

                                        computer_journal/hdb/Volume_
                                        21/Issue_03/tiff/272.tif.

McCracken:1978:PFI

[McC78e] Daniel D. McCracken. Programacion Fortran IV. Editorial Limusa, Mexico, DF, Mexico, 2d, translated to Spanish edition, 1978. 272 pp.

McCracken:1978:PFS

[McC78f] Daniel D. McCracken. Programacion FORTRAN simplificada. Editorial Limusa, Mexico, DF, Mexico, 1978. 328 pp.

McCabe:1979:QQF

[McC79] C. Kevin McCabe. QWIKTRAN, quick FORTRAN for micros, minis, and mainframes. Dilithium Press, Forest Grove, OR, May 1979. ISBN 0-918398-24-X. xvi + 220 pp. LCCN QA76.73.F25 M26. US$12.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        091839824X.

McDaniel:1980:IF

[McD80] Kenneth L. McDaniel. Introduction to Fortran. Naval Education and Training Program Development Center, Pensacola, FL, USA, 1980 edition, 1980. 126 pp.

McGregor:1967:CNC

[McG67] B. McGregor. Conversion of the neutron code TDC to FORTRAN IV. Technical Report AAEC/TM 370, Research Establishment, Australian Atomic Energy Commission, Lucas Heights, NSW, Australia, 1967. 6 pp.

McGraw-Hill:1970:EF

[McG70] McGraw-Hill Book Company. Essentials of FORTRAN, 1970.

McGuire:1976:EEE

[McG76a] Robin K. McGuire. EQRISK, evaluation of earthquake risk to site: Fortran computer programs for seismic risk analysis. Open-file report 76-67, U.S. Geological Survey, Reston, VA, USA, 1976. 91 pp.

McGuire:1976:FCP

[McG76b] Robin K. McGuire. Fortran computer program for seismic risk analysis. Open-file report 76-67, U.S. Geological Survey, Reston, VA, USA, 1976. iii + 90 pp.

Mcgettrick:1980:DPL

[Mcg80] A. D. Mcgettrick. The Definition of Programming Languages. Cambridge University Press, Cambridge, UK, 1980. ISBN 0-521-29585-8.

McIlroy:1971:AFC

[McI71] M. D. McIlroy. ANS FORTRAN charts. Computing Science Technical Report 13, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1971. ?? pp.

McKinley:1980:BFS

[McK80] Joe W. McKinley. Beginning FOR-TRAN. Matrix Publishers, Portland, OR, USA, 1980. ISBN 0-916460-11-8. x + 242 pp.

McLeod:1973:FID

[McL73] J. M. McLeod. Fortran IV data summaries for the Swaine jack pine sawfly life system. Laurentian Forest Research Centre. Rapport d’information LAU-X 3, Centre de recherches forestières des Laurentides, Sainte Foy, PQ, Canada, 1973. 59 pp.

McLain:1978:AVA

[McL78] D. H. McLain. Algorithm 100: Vector approximation to curves. The Computer Journal, 21(2):178– 180, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/
178.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/179.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
21/Issue_02/tiff/180.tif.

McMains:1966:FFC

[McM66] Forrest McMains. A Fortran flowchart computer program. Information report 8, Picatinny Arsenal, Dover, NJ, USA, 1966. 61 pp.

McMains:1967:CFP

[McM67] Forrest McMains. Contour: a Fortran program to produce contour maps from a matrix. Technical memorandum 1796, Data Processing Systems Office, Picatinny Arsenal, Dover, NJ, USA, 1967. 33 pp.

McCracken:1964:NMF

[MD64] Daniel D. McCracken and William S. Dorn. Numerical methods and FORTRAN programming, with applications in engineering and science. John Wiley and Sons, New York, London, Sydney, 1964. xii + 457 pp.

McCracken:1966:MNP

[MD66a] Daniel D. McCracken and William S.

Dorn. Metodos numericos y programacion FORTRAN: con aplicaciones en ingenieria y ciencias. Editorial Limusa, Mexico, DF, Mexico, 1966. ISBN 968-18-0827-4. 476 pp.

McCracken:1966:NMF

[MD66b] Daniel D. McCracken and William S.

Dorn. Numerical methods and Fortran programming: with applications in engineering and science. John Wiley and Sons, New York, London, Sydney, 1966. 457 pp.

McCracken:1968:NMF

[MD68] Daniel D. McCracken and William S.

Dorn. Numerical methods and FORTRAN programming with applications in engineering and science. John Wiley and Sons, New York, London, Sydney, 1968. 457 pp.

McCracken:1973:MNP

[MD73] Daniel D. McCracken and William S.

Dorn. Metodos numericos y programacion FORTRAN; con aplicaciones en ingenieria y ciencias. Editorial Limusa-Wiley, Mexico, DF, Mexico, 1973. 476 pp.

Meek:1972:CSE

[Mee72] B. L. Meek. Correspondence: On “Suggested Extensions to FOR-TRAN IV”. The Computer Journal, 15(3):267, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
15/Issue_03/150267.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/267.tif. See

[Fin72b].

Meeks:1974:ASE

[Mee74] Arthur Lee Meeks. Analysis of the statement execution sequence of DO(s) with extended ranges in FORTRAN programs. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1974. 32 pp.

Meek:1978:FPAb

[Mee78a] B. L. (Brian Lawrence) Meek. Fortran, PL/I, and the Algols. Elsevier North-Holland, New York, NY, USA, 1978. ISBN 0-444-19464-9. viii + 291 pp. LCCN QA76.73.F25 .M43.

Meek:1978:FPAc

[Mee78b] B. L. (Brian Lawrence) Meek. Fortran, PL/I, and the Algols. Macmillan computer science series. MacMillan Publishing Company, New York, NY, USA, 1978. ISBN 0-333-24385-4. viii + 291 pp.

Meek:1978:FPI

[Mee78c] Brian Meek. Fortran, PL/I and the Algols. Unwin Brothers Limited, Old Woking, Surrey, UK, 1978. ISBN 0-333-24385-4.

Meek:1978:FPAa

[Mee78d] Brian L. (Brian Lawrence) Meek.

Fortran, PL/I and the Algols. Macmillan computer science series. MacMillan Publishing Company, New York, NY, USA, 1978. ISBN 0-333-24385-4. viii + 291 pp. LCCN QA76.73.F25 M43 1978b. UK£12.00.

Meek:1979:FPA

[Mee79] B. Meek. Fortran, PL-I and the Algols. American Elsevier, New York, NY, USA, September 1979. ISBN 0-444-19464-9. ???? pp. LCCN ???? US$48.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0444194649.

Mei:1968:APM

[Mei68] Harry T. Mei. Algorithmic process manual, Fortran IV/CDC3300. Lamar State College of Technology, Beaumont, TX, USA, 1968. various pp.

Mei:1969:FIC

[Mei69] Harry T. Mei. FORTRAN IV:

CDC3300 for engineers and scientists. Lamar State College of Technology, Beaumont, TX, USA, 1969. various pp.

Meissner:1971:RF

[Mei71] Loren P. Meissner. Rudiments of FORTRAN. Addison-Wesley, Reading, MA, USA, 1971. x + 109 pp.

Meissner:1974:CSE

[Mei74] Loren P. Meissner. A compatible “structured” extension to Fortran. ACM SIGPLAN Notices, 9 (10):29–36, October 1974. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1975:EFC

[Mei75a] L. P. Meissner. On extending FORTRAN control structures to facilitate structured programming. ACM SIGPLAN Notices, 10(9): 19–30, September 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1975:P

[Mei75b] L. P. Meissner. On preprocessors. ACM SIGPLAN Notices, 10(12): 39–??, December 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1976:PCS

[Mei76] L. P. Meissner. Proposed control structures for extended FOR-TRAN. ACM SIGPLAN Notices, 11(1):16–21, January 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1977:F

[Mei77] L. P. Meissner. FORTRAN 77. ACM SIGPLAN Notices, 12(1): 93–94, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meitner:1978:FPC

[Mei78] Peter L. Meitner. FORTRAN program for calculating coolant flow and metal temperatures of a fullcoverage-film-cooled vane or blade. NASA technical paper 1259, National Aeronautics and Space Administration, Scientific and Technical Information Office, Washington, DC. USA, 1978. iv + 82 pp. For sale by the National Technical Information Service.

Melkanoff:1962:FPE

[Mel62] Michel A. Melkanoff. A Fortran program for elastic scattering analyses with the nuclear optical model. University of California publications in automatic computation 1, University of California Press, Berkeley, CA, USA, 1962. 116 pp.

Merwin:1958:AFP

[Mer58a] Marjorie Merwin. Addenda to the FORTRAN programmer’s reference manual. Memorandum CC-93, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, May 1, 1958. 4 pp.

Merwin:1958:DOF

[Mer58b] Marjorie Merwin. Description of the output of the Fortran compiler. Memorandum CC-71, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, February 17, 1958. 5 pp.

Merwin:1960:DFA

[Mer60a] Marjorie Merwin. Description of the FAP assembly program used with the FORTRAN monitor system for the 709 computer. Memorandum CC-161, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 34 pp.

Merwin:1960:NFI

[Mer60b] Marjorie Merwin. Notes on the FORTRAN II compiled program. Memorandum CC-152, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 7 pp.

Merrill:1974:FCM

[Mer74] R. Gregg Merrill. FORTRAN concordance manual. NOAA technical memorandum ERL SEL-32, Space Environment Laboratory, Boulder, CO, USA, 1974. 19 pp.

GarciaMerayo:1977:LF

[Mer77] Felix Garcia Merayo. El lenguaje Fortran. Paraninfo, Madrid, Spain, 2a. corregida edition, 1977. 295 pp.

Merriam:1978:FCPa

[Mer78a] C. W. (Charles Wolcott) Merriam. FORTRAN computer programs: solutions to optimization problems arising in feedback control. D. C. Heath and Company, Lexington, MA, USA, 1978. ISBN 0-669-01995-X. xiii + 351 pp. Distributed by Teakfield Ltd.

Merriam:1978:FCPb

[Mer78b] C. W. (Charles Wolcott) Merriam. Fortran Computer Programs: Solutions to Optimization Problems Arising in Feedback Control. Lexington Books, Lexington, MA, USA, April 1978. ISBN 0-669-01995-X. xiii + 350 pp. LCCN TJ216.M41. US$26.95. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=066901995X.

Merriam:1978:FCPc

[Mer78c] C. W. (Charles Wolcott) Merriam. FORTRAN computer programs: solutions to optimization problems arising in feedback control. Lexington Books, Lexington, MA, USA, 1978. ISBN 0-669-01995-X. xiii + 350 pp.

Merchant:1979:AFP

[Mer79] Michael J. Merchant. The ABC’s of Fortran Programming. Wadsworth, Pacific Grove, CA, USA, March 1979. ISBN 0-534-00634-5. ix + 357 pp. LCCN ???? US$29.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0534006345.

Messier:1973:IAF

[Mes73] Raymond Messier. Introduction au Fortran. R. Messier, Trois-Rivi‘eres, Qu’ebec, Canada, 1973. 126 pp.

Messier:1974:WFI

[Mes74] Raymond Messier. WATFIV et FORTRAN IV. R. Messier, Trois-Rivi‘eres, Qu’ebec, Canada, 1974. 109 pp.

Metcalf:1980:FPO

[Met80] Michael Metcalf. FORTRAN program optimization. Technical Report ????, European Organization for Nuclear Research, Geneva, Switzerland, 1980. vi + 41 pp.

Malcom:1968:CBF

[MG68] Robert E. Malcom and Malcolm H. Gotterer. Computers in business; a Fortran introduction. International Textbook Co., Scranton, PA, USA, 1968. x + 267 pp.

Manning:1970:FIP

[MG70] William A. Manning and Robert S. Garnero. A Fortran IV Problem Solver. McGraw-Hill, New York, NY, USA, January 1970. ISBN 0-07-039918-2. viii + 167 pp. LCCN ???? US$8.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0070399182.

Mawson:1971:DFI

[MG71] Joseph C. Mawson and Paul J. Godfrey. DIVERSE: a FORTRAN IV program to calculate diversity indices of stream bottom organisms. Technical report, Water Resources Research Center, University of Massachusetts, Amherst, NY, USA, 1971. 11 + [12] pp.

Malcom:1973:CAF

[MGL73] Robert E. Malcom, Malcolm H. Gotterer, and Frank Luh. Computers in administration: a Fortran IV introduction. Intext Educational Publishers, New York, NY, USA, 1973. ISBN 0-7002-2426-2. ix + 374 pp. LCCN HF5548.5.F2 M3 1973. Freely adapted from an earlier book, Malcom and Gotterer: Computers in business … 1967.

Marxer:1972:ECP

[MH72] Ellen Marxer and Donald L. Hartford. Elements of computer programming: FORTRAN. Delmar Publishers, Albany, NY, USA, 1972. 223 pp.

Marxer:1973:ECP

[MH73] Ellen Marxer and Donald L. Hartford. Elements of computer programming: FORTRAN, teacher’s manual. Delmar Publishers, Albany, NY, USA, 1973. iii + 82 pp.

Martinson:1975:SAA

[MH75a] E. O. Martinson and M. A. Hamdan. Statistical algorithms: Algorithm AS 87: Calculation of the polychoric estimate of correlation in contingency tables. Applied Statistics, 24(2):272–278, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/87. See re-

mark [?].

Meissner:1975:BSM

[MH75b] Loren P. Meissner and Ruth L.

Hinkins. B4Tran: a structured mini-language approach to the teaching of Fortran. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):200–205, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Maly:1978:FCS

[MH78] K. Maly and A. R. Hanson. Fundamentals of the Computing Sciences. Supplementary Volume. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1978.

Maniotes:1971:BFSb

[MHH71a] John Maniotes, Harry B. Higley, and James N. Haag. Beginning FORTRAN: simplified, 12statement programming. Hayden computer programming series. Hayden Book Co., Rochelle Park, NJ, USA, 1971. 248 pp.

Maniotes:1971:BFSc

[MHH71b] John Maniotes, Harry B. Higley, and James N. Haag. Beginning FORTRAN: simplified, 12statement programming teacher’s guide and answer book. Hayden computer programming series. Hayden Book Co., Rochelle Park, NJ, USA, 1971. 30 pp.

Mendicino:1968:LC

[MHM^{+}68] Sam F. Mendicino, Robert A. Hughes, Jeanne T. Martin, Frank H. McMahon, John E. Ranelletti, and Richard G. Zwakenberg. The LRLTRAN compiler. Comm. ACM, 11(11): 747–755, November 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manson:1964:FPF

[MI64] Vincent Manson and John Imbrie.

Fortran program for factor and vector analysis of geologic data using an IBM 7090 or 7094/1401 computer system. Special distribution publication 13, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1964. 46 pp.

Maine:1975:FPD

[MI75a] Richard E. Maine and Kenneth W.

Iliff. A FORTRAN program for determining aircraft stability and control derivatives from flight data. NASA technical note NASA TN D-7831, National Aeronautics and Space Administration, Washington, DC, USA, 1975. vi + 171 pp. For sale by the National Technical Information Service.

Miller:1975:NMF

[MI75b] John F. Miller and Frank Ingenito. Normal mode FORTRAN programs for calculating sound propagation in the ocean. NRL memorandum report 3071, National Technical Information Service, Springfield, VA, USA, 1975. iii + 107 pp.

Maine:1980:UMM

[MI80] Richard E. Maine and Kenneth W.

Iliff. User’s manual for MMLE3, a general FORTRAN program for maximum likelihood parameter estimation. NASA technical paper 1563, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1980. iv + 85 pp. For sale by the National Technical Information Service.

Microsoft:1979:MFD

[Mic79a] Microsoft. Microsoft FORTRAN 80 documentation. Microsoft, Inc., Bellevue, WA, USA, 1979. various pp.

Microsoft:1979:MFU

[Mic79b] Microsoft. Microsoft FORTRAN 80 user’s manual, 1979.

Middleton:1974:COP

[Mid74] A. G. Middleton. Cost-oriented program optimisation. Information Processing Letters, 2(6):167– 170, April ??, 1974. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mikami:1973:FOP

[Mik73] E. Y. Mikami. A FORTRAN optimization program using the epsilon method. NWC technical publication 5573, Naval Weapons Center, China Lake, CA, USA, 1973. 26 pp.

Miller:1968:CNF

[Mil68] Victor Julian Miller. Complex number FORTRAN. Thesis (m.s.), Newark College of Engineering, Newark, NJ, USA, 1968. 48 pp.

Miller:1973:EFS

[Mil73a] Edward F. Miller. Extensions to FORTRAN to support structured programming (ITRAN). ACM SIGPLAN Notices, 8(6):63–64, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millstein:1973:CSI

[Mil73b] R. Millstein. Control structures in Illiac IV Fortran. Comm. ACM, 16(??):622–627, ?? 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Millstein:1975:CSI

[Mil75] R. E. Millstein. Control structures in Illiac IV Fortran. Comm. ACM, 6(10):157–164, October 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Misra:1978:ALC

[Mis78a] Alok Chandra Misra. Authoring languages for computer-assisted instruction: implementation of FORTRAN topics in ETL, an advanced CAI language, and experimentation with CS 1011-tutorials. Thesis (m.s.), Worcester Polytechnic Institute, Worcester, MA, USA, 1978. 133 pp.

MSUTETCC:1978:FPR

[Mis78b] Mississippi State University.Thomas E.Tramel Computing Center. FORTRAN programmer’s reference. Computing Center Reference Manual Series 1, Thomas E. Tramel Computing Center, Mississippi State University, Mississippi State, MS, USA, 1978. various pp.

Mitchell:1965:FIP

[Mit65] J. K. Mitchell. FORTRAN II program for precipitation data analysis. Agricultural engineering research report, University of Illinois, College of Agriculture, Agricultural Experiment Station, Agricultural Engineering Department, Urbana, IL, USA, 1965. 43 pp.

Mills:1968:FPC

[MK68] J. C. Mills and Colin Harold Leslie Kennard. Fortran programmes for crystallographers. Technical Report 1(1), Queensland University Computer Centre, Brisbane, Queensland, Australia, 1968. v pp.

McCullagh:1970:SFI

[MK70] M. J. McCullagh and Cuchlaine A. M. King. Spitsym, a FOR-TRAN IV, computer program for split simulation. Computer contribution 50, University of Kansas, Lawrence, KS, USA, 1970. 20 pp.

Morgan:1973:ALP

[MK73] Howard L. Morgan and James C. Kinard. ASAP: a language and philosophy for teaching file processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1): 21–23, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Malcolme-Lawes:1970:PF

[ML70a] D. J. Malcolme-Lawes. Programming — Fortran. Pergamon Press, Oxford, 1970. ISBN 0-08-015508-1. v + 122 pp. LCCN QA76.5.M1918 1970.

Malcolme-Lawes:1970:P

[ML70b] D. J. Malcolme-Lawes. Programming– Fortran. Pergamon Press, Oxford, NY, USA, ???? edition, 1970. ISBN 0-08-015508-1. v + 122 pp.

McCarthy:1958:RTF

[MM58] John McCarthy and Marjorie Merwin. Routines for turning Fortran program into Fortran functions. Memorandum CC-60, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, January 3, 1958. 5 pp.

Mancino:1965:ISF

[MM65] O. G. Mancino and M. Morandi Checchi. The internal structure of the FORTRAN CEP translator. Comm. ACM, 8(3):149–151, March 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:1969:FIP

[MM69] Charles O. Morgan and Jesse M.

McNellis. FORTRAN IV program, KANS, for the conversion of general land office locations to latitude and longitude coordinates. Special distribution publication 42, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1969. 24 pp.

Machura:1973:AAR

[MM73a] Marek Machura and Andrzej Mulawa. ACM Algorithm 450: Rosenbrock function minimization [E4]. Comm. ACM, 16(8):482–483, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [?, ?, ?, ?].

Machura:1973:ARF

[MM73b] Marek Machura and Andrzej Mulawa. Algorithm 450: Rosenbrock function minimization [E4]. Comm. ACM, 16(8):482–483, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [?, ?, ?, ?].

Millstein:1975:IIF

[MM75] Robert E. Millstein and Charles A. Muntz. The ILLIAC IV FOR-TRAN compiler. ACM SIGPLAN Notices, 10(3):1–8, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:1978:SFW

[MM78] John B. Moore and Leo J. Makela. Structured Fortran with WATFIV: text and reference. Reston Publishing Co., Inc., Reston, VA, USA, 1978. ISBN 0-8359-7101-5. xvi + 552 pp.

McKinley:1980:BF

[MM80] Joe W. McKinley and Merl K. Miller. Beginning Fortran. Matrix Publishers, Portland, OR, USA, July 1980. ISBN 0-916460-11-8 (paperback). x + 242 pp. LCCN QA76.73.F25 .M4. US$15.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0916460118.

Meissner:1980:FFS

[MO80] Loren P. Meissner and Elliott Irving Organick. Fortran 77: Featuring Structured Programming.

Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, third edition, January 1, 1980. ISBN 0-201-05499-X. xi + 500 pp. LCCN QA76.73 .F25 O73 1980. US$11.95; US$32.25. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

020105499X. Second ed. (1974) published under title: Fortran IV. Includes index.

Mochel:1969:FPP

[Moc69] Myron G. Mochel. FOR-TRAN programming, programs and schematic storage maps. Clarkson College of Technology, Potsdam, NY, USA, 1969. viii + 229 pp.

Mochel:1970:FPP

[Moc70] Myron G. Mochel. Fortran programming, programs, and schematic storage maps. McGraw-Hill, New York, NY, USA, 1970. ISBN 0-07-042635-X. ix + 192 pp. LCCN QA76.5 .M538.

Mochel:1971:FPP

[Moc71a] Myron G. Mochel. Fortran programming, programs, and schematic storage maps. McGraw-Hill, New York, NY, USA, 1971. ISBN 0-07-042635-X (paperback). xi + 192 pp.

Mochel:1971:SMA

[Moc71b] Myron G. Mochel. Solutions manual to accompany FOR-TRAN programming, programs, and schematic storage maps. Mc-Graw-Hill, New York, NY, USA, 1971. 125 pp.

Modi:1974:CRF

[Mod74] Mukund U. Modi. Coupling realtime Fortran to DDC for hierarchic process control. Thesis (m.e.), Widener College, Chester, PA, USA, 1974. 41 + x pp.

Mohilner:1977:UPF

[Moh77] Patricia R. Mohilner. Using Pascal in a Fortran environment. Software—Practice and Experience, 7 (3):357–362, May/June 1977. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Moler:1971:MCF

[Mol71] Cleve B. Moler. Matrix computations with Fortran and paging. Technical Report STAN-CS-71-196, Computer Science Department, Stanford University, Stanford, CA, USA, 1971. 13 pp.

Moler:1972:LES

[Mol72a] C. B. Moler. Linear equation solver. Comm. ACM, 15(4):274– ??, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). ACM Algorithm 423.

Moler:1972:MCF

[Mol72b] C. B. Moler. Matrix computations with Fortran and paging. Comm. ACM, 15(4):268–270 (or 268–274??), April 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1972:AAL

[Mol72c] Cleve B. Moler. ACM Algorithm 423: Linear equation solver [F4]. Comm. ACM, 15(4):274, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Monro:1975:SAA

[Mon75] Donald M. Monro. Statistical algorithms: Algorithm AS 83: Complex discrete fast Fourier transform. Applied Statistics, 24(1): 153–160, March 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/83.

Monro:1977:CFP

[Mon77a] Donald M. Monro. Computing With Fortran: a Practical Course. Edward Arnold, London, UK, June 1977. ISBN 0-7131-2546-2. vi + 242 pp. LCCN QA 76.73 F25 M757 1977. US$14.95; US$3.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0713125462.

Monro:1977:CFI

[Mon77b] Donald M. Monro. Computing with FORTRAN IV: a practical course. Bits Press, Cleveland, OH, USA, 1977. 242 pp.

Montgomery:1978:EBE

[Mon78] Peter L. Montgomery. Evaluation of Boolean expressions on one’s complement machines. ACM SIG-PLAN Notices, 13(12):60–72, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Monro:1979:CFI

[Mon79] Donald M. Monro. Computing with FORTRAN IV: a practical course. Edward Arnold, London, UK, reprinted 1979 with minor corrections edition, 1979. ISBN 0-7131-2546-2. vi + 242 pp.

Moore:1960:IDP

[Moo60] Donald P. Moore. IBM 709/ 7090 data processing system bulletin FORTRAN assembly program (FAP) for the IBM 709/7090. IBM Corporation, White Plains, NY, USA, 1960. 77 pp.

Mooers:1969:DDL

[Moo69] Calvin N. Mooers. Data descriptive languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(1):31–36, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Moon:1971:LPF

[Moo71] B. A. M. Moon. List-processing in plain FORTRAN. Australian Computer Journal, 3(3), August 1971. CODEN ACMJB2. ISSN 0004-8917.

Moore:1975:WFP

[Moo75] John B. Moore. WATFIV: Fortran programming with the WAT-FIV compiler. Reston Publishing Co., Inc., Reston, VA, USA, June 1975. ISBN 0-87909-876-7. xvi + 492 pp. LCCN ????

Moore:1976:FCP

[Moo76] William R. Moore. Fortran calculation of proton induced X-ray production cross sections. Thesis (b.s.), East Texas State University, Commerce, TX, USA, 1976. 22 pp.

Moore:1977:IFC

[Moo77] James Lum Moore. An interactive FORTRAN compiler for the TICK system. Thesis (m. s.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1977. vi + 86 pp.

Moran:1970:IF

[Mor70] James E. Moran, Jr. IBM 1130 FORTRAN. OTI, Klamath Falls, OR, USA, 1970. 67 pp.

Moruzzi:1971:AFT

[Mor71] V. L. Moruzzi. APL/FORTRAN translations. Technical report, IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1971. 9 pp.

Morrison:1973:UPP

[Mor73] J. E. Morrison. User program performance in virtual storage systems. IBM Systems Journal, 12(3): 216–237, 1973. CODEN IBMSA7. ISSN 0018-8670.

Morrison:1975:SCM

[Mor75] W. H. (William H.) Morrison.

String and character manipulation routines for UNIVAC 1108 FOR-TRAN users. Technical Report GAS/OP/MCL; TM-248, U. S. General Services Administration, Office of Preparedness, Washington, DC. USA, 1975. 33 pp.

Morawski:1979:EFP

[Mor79] Paul Morawski. EDIT: a FOR-TRAN program maintenance system for the Texas Instruments ASC-7. Technical report, Naval Research Laboratory, Washington, DC, USA, 1979. iii + 13 pp.

Moses:1964:AFP

[Mos64] Arden T. Moses. Applications of FORTRAN programming. Moronics Co., Las Cruces, NM, USA, 1964. various pp.

Mostek:1978:XFI

[Mos78] Mostek Corporation. XFOR-80 FORTRAN IV cross assembler. Mostek Corp., Carrollton, TX, USA, 1978. ii + 18 + [26] pp.

Mothershed:1966:OFC

[Mot66] C. T. Mothershed. ORSEF: a Fortran code for the calculation of multi-stage flash evaporation desalination plant designs. Ornl-tm-1542, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1966. v + 126 pp.

Mottl:1979:DPN

[Mot79] J. Mottl. Description of a program for nonlinear programming. The Computer Journal, 22(3): 256–261, August 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
22/3/256.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/
256.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/257.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
22/Issue_03/tiff/258.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_
03/tiff/259.tif;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/
260.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/261.
tif.

Moursund:1970:PAS

[Mou70] David G. Moursund. Problem analysis and solution using Fortran IV. Wadsworth, Pacific Grove, CA, USA, 1970. 275 pp.

Melbourne:1965:SCD

[MP65] A. J. Melbourne and J. M. Pugmire. A small computer for the direction processing of FORTRAN statements. The Computer Journal, 8(1):24–27, April 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
08/Issue_01/080024.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/24.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
25.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/26.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/
27.tif.

Marlow:1972:FSD

[MP72] S. Marlow and M. J. D. (Michael James David) Powell. A Fortran subroutine for drawing a curve through a given sequence of data points. U.K.A.E.A. Research Group. Report, R7092. A.E.R.E., Harwell, Berkshire, UK, 1972. ISBN 0-7058-0022-9. 51 pp.

Marlow:1973:FSP

[MP73] S. Marlow and M. J. D. Powell. A FORTRAN subroutine for plotting a cubic spline function. Technical Report R7470, AERE Harwell, Harwell, Berkshire, UK, 1973.

Meeson:1979:OFP

[MP79] Reginald Meeson and Arthur Pyster. Overhead in FORTRAN preprocessors. Software—Practice and Experience, 9(12):987– 999, December 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Meyer:1967:CLP

[MR67] Albert R. Meyer and Dennis M. Ritchie. The complexity of loop programs. In Solomon Rosenthal, editor, ACM ’67: Proceedings of the 1967 22nd national conference, pages 465–569. ACM Press, New York, NY 10036, USA, January 1967. ISBN 1-4503-7494-8.

                                        URL https://dl.acm.org/doi/
                                        abs/10.1145/800196.806014.

Mawson:1970:TCP

[MR70a] J. C. Mawson and R. J. Reed. Three computer programs: Backcalculation, condition factor, and stomach content, CDC 3600 Fortran/Format. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (1):156–157, January 1970. CO-DEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.
com/doi/full/10.1139/f70-015.

Muthukrishnan:1970:CDT

[MR70b] C. R. Muthukrishnan and V. Rajaraman. On the conversion of decision tables to computer programs. Comm. ACM, 13 (6):347–351, June 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercer:1973:AGP

[MR73] Andrew Mercer and Azriel Rosenfeld. An array grammar programming system. Comm. ACM, 16(5):299–305, May 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marlow:1978:FSC

[MR78] S. Marlow and John Ker Reid. A Fortran subroutine for comparing two files. AERE-R 8971, Computer Science and Systems Division, Atomic Energy Reseach Establishment, Harwell, Berkshire, UK, 1978. [1] + 24 pp. Distributed by HMSO.

Makarenko:1973:F

[MRS73] G. I. (Grigorii Ivanovich) Makarenko, A. V. (Aleksandr Vladimirovich) Rakitskii, and A. I. (Albert Ivanovich) Saltykov. FORTRAN. Novoe v zhizni, nauke, tekhnike. Seriia “Mathematika, kibernetika”, 11 Novoe v zhizni, nauke, tekhnike. Seriia Mathematika, kibernetika; 1973, no. 11. Znanie, Moskva, USSR, 1973. 64 pp.

McCormick:1964:NMF

[MS64] John Michael McCormick and Mario George Salvadori. Numerical methods in FORTRAN. Prentice-Hall applied mathematics series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1964. 324 pp.

Mitalas:1966:FIP

[MS66] G. P. Mitalas and D. G. Stephenson. Fortran IV programs to calculate radiant energy interchange factors. Computer program 25, National Research Council, Canada, Division of Building Research, Ottawa, Ontario, Canada, July 1966. 13 + A13 pp.

Murrill:1968:FIP

[MS68] Paul W. Murrill and Cecil L. Smith. Fortran IV programming, for engineers and scientists. International Textbook Co., Scranton, PA, USA, 1968. ix + 313 pp.

Martin:1969:TSM

[MS69] Donald C. Martin and Beverly Saylor. Telecourse students manual for introduction to FORTRAN programming. Technical report, Division of Continuing Education and Industrial Extension Service, North Carolina State University, Raleigh, NC, USA, 1969. various pp.

Mueller:1970:FP

[MS70a] K. H. Mueller and I. Streker. Fortran Programmierungsanleitung. Bibliographisches Institut, Mannheim, Germany, 1970. ISBN 3-411-00804-0.

Murrill:1970:IFIa

[MS70b] Paul W. Murrill and Cecil L.

Smith. An introduction to FOR-TRAN IV programming; a general approach. International Textbook Co., Scranton, PA, USA, 1970. ISBN 0-7002-2266-9. xi + 276 pp. LCCN QA76.5 .M85. Includes appropriate implementation information and error codes for WATFOR and WATFIV compilers.

Murrill:1970:IFIb

[MS70c] Paul W. Murrill and Cecil L.

Smith. An introduction to FOR-TRAN IV programming: a general approach. Thomas Y. Crowell, New York, NY, USA, second edition, 1970. ISBN 0-7002-2469-6. x + 389 pp.

Michelangeli:1971:FII

[MS71] P. Michelangeli and Jacques Subsol. FORTRAN IV (I.B.M. version H). Compagnie internationale de services en informatique, Saclay, France, 1971. 188 pp.

Mickunas:1973:PGS

[MS73a] M. Dennis Mickunas and Victor B. Schneider. A parsergenerating system for constructing compressed compilers. Comm. ACM, 16(11):669–676, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Murrill:1973:FIPa

[MS73b] Paul W. Murrill and Cecil L.

Smith. Fortran IV programming for engineers and scientists.

Harper & Row, New York, NY, USA, second edition, 1973. ISBN 0-7002-2419-X. xi + 322 pp. LCCN QA76.73.F25 M87 1973.

Murrill:1973:FIPb

[MS73c] Paul W. Murrill and Cecil L. Smith. Fortran IV programming for engineers and scientists. Intext Educational Publishers, New York, NY, USA, second edition, 1973. ISBN 0-7002-2419-X. xi + 322 pp. LCCN QA76.73.F25 M87 1973.

Murrill:1973:ICS

[MS73d] Paul W. Murrill and Cecil L. Smith. Introduction to Computer Science. Intext Educational Publishers, New York, NY, USA, 1973. ISBN 0-7002-0219-6.

Murrill:1973:SMF

[MS73e] Paul W. Murrill and Cecil L. Smith. Solutions manual for Fortran IV programming for engineers and scientists. Harper and Row, Publishers, New York, NY, USA, second edition, 1973. 66 pp.

MacCallum:1974:MLS

[MS74a] K. J. MacCallum and L. T. Shafe. A mixed language system POP-2 and FORTRAN. Software— Practice and Experience, 4(2):145– 154, April/June 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Manter:1974:IMS

[MS74b] M. Manter and J. Sarni. An introductory manual to STOFI, data management system. Technical report, University of California (Berkeley CA), Lawrence Berkeley Lab, UCB (Berkeley CA), Berkeley, CA, USA, January 1974.

Murray-Shelley:1975:CPE

[MS75a] Richard Murray-Shelley. Computer programming for electrical engineers. McGraw-Hill, New York, NY, USA, 1975. ISBN 0-07-084060-1. 151 pp. LCCN QA76.6.M88.

Murrill:1975:IFIb

[MS75b] Paul W. Murrill and Cecil L.

Smith. Introduction to Fortran IV Programming. John Wiley and Sons, New York, London, Sydney, second edition, June 1975. ISBN 0-471-60341-4. ???? pp. LCCN ???? US$35.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471603414.

Murrill:1975:IFIa

[MS75c] Paul W. Murrill and Cecil L.

Smith. An introduction to FOR-TRAN IV programming: a general approach. Intext Educational Publishers, New York, NY, USA, second edition, 1975. ISBN 0-7002-2469-6. x + 389 pp. LCCN QA76.73.F25M88 1975.

Murrill:1975:SMA

[MS75d] Paul W. Murrill and Cecil L.

Smith. Solutions manual to accompany An introduction to Fortran IV programming: a general approach, 2d edition. Thomas Y. Crowell, New York, NY, USA, 1975. ISBN 0-7002-2479-3. 85 pp.

Maron:1977:FIV

[MS77a] N. (Neil) Maron and G. G. Sutherland. Fortran interface to VEctor programming. Technical report, Lawrence Livermore Laboratory, Livermore, CA, USA, 1977. 9 pp.

Merchant:1977:AFP

[MS77b] Michael J. Merchant and John R. Sturgul. Applied FORTRAN programming: with standard FOR-TRAN, WATFOR, WATFIV, and structured WATFIV. Wadsworth, Pacific Grove, CA, USA, 1977. ISBN 0-534-00497-0. xiii + 544 pp. LCCN QA76.73.F25M46 1977.

Maron:1978:FIV

[MS78] Neil Maron and G. G. Sutherland. Fortran interface to vector programming. UCID 17477, University of California, Lawrence Livermore Laboratory Technical Information Dept., Livermore, CA, USA, 1978. 9 pp. For sale by the National Technical Information Service.

Miura:1979:NFP

[MS79] Hirokazu Miura and Lucien A. Schmit. NEWSUMT: a Fortran program for inequality constrained function minimization: users guide. Technical report, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA 23665, Hampton, VA, USA, 1979. ii + 44 pp.

Melkanoff:1961:FPE

[MSNC61] Michel A. Melkanoff, David S.

Saxon, John S. Nodvik, and David G. Cantor. A Fortran program for elastic scattering analyses with the nuclear optical model. University of California publications in automatic computation 1, University of California Press, Berkeley, CA, USA, 1961. iii + 116 pp.

Melkanoff:1966:SFP

[MSR66] Michel A. Melkanoff, Tatsuro Sawada, and Jacques Raynal. SEEK: a FORTRAN program for automatic searches in elastic scattering analyses with the nuclear optical model. Report 66-10, Dept. of Engineering, University of California, Los Angeles, CA, USA, 1966. iii + 105 pp.

Meyer:1978:S

[MSS78a] Meyer, Schaller, and B. Schmidt.

SIM-queue: . . .. Arbeitsberichte des IMMD 11, Universität Erlangen, Erlangen, Germany, 1978.

Musielak:1978:VSG

[MSS78b] H. Musielak, B. Schmidt, and M. Stößel. Vergleich von Simulationssprachen GPSS-Fortran — GPSS — GASP IV. Arbeitsberichte des IMMD und El. Rechenanlagen 21, 1979, 23-28, 11, 9:55, 1978.

Mizzi:1972:PPL

[MT72] F. Mizzi and P. Toussaint. Physictran: a programming language for physicists, economists and engineers extending the possibilities of Fortran V in automatic solving. In Freiman et al. [FGR72], pages 364–369. ISBN 0-7204-2063-6. LCCN QA75.5 .I532

                                        1971. URL https://dblp.org/
                                        db/conf/ifip/ifip71-1.html.

McCabe:1975:FIP

[MT75] A. McCabe and N. J. Topping. FORTRAN IV program for the mapping of spatial data using a grey intensity scale. Report 17, University of Rhodesia Institute of Mining Research, Salisbury, Rhodesia, 1975. 13 pp.

Martello:1978:AAS

[MT78] S. Martello and P. Toth. Algorithm 37. Algorithm for the solution of the 0-1 single knapsack problem. Computing, 21(1):81– 86, 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

MtLebanon:1979:CGF

[Mt.79] Mt.Lebanon School District (Pa.). The Curriculum guide for FOR-TRAN IV: Mt. Lebanon High School. Mt. Lebanon School District, Pittsburgh, PA, USA, 1979. 14 pp.

Moore:1980:FSA

[MT80] John B. Moore and Gilles Tellier. FORTRAN structure avec WAT-FIV. Edition de l’École polytechnique de Montréal, Montréal, PQ, Canada, 1980. xv + 601 pp.

Minor:1975:FPC

[MU75] Lewis G. Minor and John R. Underwood. A FORTRAN program for the calculation of the state transition matrix as a linear combination of real time functions (EAT). Technical report, U.S. Army Missile Command, Redstone Arsenal, AL, USA, 1975. 44 pp. Distributed by the National Technical Information Service.

Mueller:1966:FSH

[Mue66] Dennis J. Mueller. FORTRAN subroutines for Householder’s method in the complex case and eigensystems of Hermitian matrices. Technical Report ANL-7231, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1966. 31 pp.

Mueller:1969:PER

[Mue69] D. Mueller. Programmierung Elektronischer Rechenanlagen. Bibliographisches Institut, Mannheim, Germany, 3 edition, 1969.

Muecke:1975:FPW

[Mue75] Arnold Henry Muecke. A Fortran preprocessor which uses the Reinwald-Soland algorithm to translate decision tables into optimal source code. Thesis (m.s.), Dept. of Computer Science, College of Natural Sciences and Mathematics, University of Houston, Houston, TX, USA, 1975. i + 47 + 67 pp.

Mullish:1968:MPFa

[Mul68a] Henry Mullish. Modern programming: Fortran IV. Blaisdell Pub. Co., Waltham, MA, USA, 1968. ISBN 0-471-00388-3. xii + 132 pp.

Mullish:1968:MPFb

[Mul68b] Henry Mullish. Modern programming: Fortran IV. Xerox College Publishing, Lexington, MA, USA, 1968. xii + 132 pp.

Muller:1980:SFT

[Mul80a] Diane R. Muller. A self-paced FORTRAN tutorial. Thesis (m.s.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1980. ii + 91 + [2] pp.

Mullish:1980:MPF

[Mul80b] Henry Mullish. Modern programming: Fortran IV. John Wiley and Sons, New York, London, Sydney, 1980. xii + 132 pp.

Murphy:1966:TCC

[Mur66] Harry M. Murphy, Jr. TIDY, a computer code for renumbering and editing FORTRAN source programs. AS 642 099 ????, Clearinghouse for Federal Scientific and Technical Information, Springfield, VA, USA, 1966. 93 pp.

Murata:1970:FNY

[Mur70] Haruo Murata. FORTRAN ni yoru puroguramingu tekunikku. CS raiburari; 1. Sogo Tosho, Tokyo, Japan, 1970. 2 + 7 + 227 pp.

Murray:1971:IRL

[Mur71] John H. Murray. Inversion of rational Laplace transforms using FORTRAN IV. Thesis (m.e.), Brigham Young University, Provo, UT, USA, 1971. xi + 123 pp.

Murach:1977:BDPb

[Mur77a] Mike Murach. Business data processing: COBOL and BASIC/ FORTRAN versions also; overhead projector foils. Science Research Associates, Chicago, IL, USA, second edition, 1977. various pp.

Murach:1977:BDPc

[Mur77b] Mike Murach. Business data processing: COBOL and BASIC/ FORTRAN versions also; student workbook. Science Research Associates, Chicago, IL, USA, second edition, 1977. ISBN 0-574-21112-8. 126 pp.

Murach:1977:BDPa

[Mur77c] Mike Murach. Business data processing with BASIC and FOR-TRAN. Science Research Associates, Chicago, IL, USA, second edition, 1977. ISBN 0-574-21115-2. x + 575 pp. LCCN HF 5548.5 B3 M93 1977. First ed. published in 1973 under title: Business data processing and computer programming. Includes index.

Murach:1980:IFA

[Mur80] Mike Murach. An introduction to FORTRAN: to accompany Business data processing. Science Research Associates, Chicago, IL, USA, 1980. ISBN 0-574-21285-X. viii + 118 pp.

Malan:1966:FOR

[MV66] Willemien Malan and R. Van der Walt. Fire-2; a one-dimensional reactor computing program in Fortran IV for multi-region, multigroup reactors. PEL 88288, Atomic Energy Board, Pretoria, South Africa, 1966. 5 pp.

Maisel:1969:IED

[MW69] Herbert Maisel and Donald L. Wright. Introduction to electronic digital computers; with emphasis on the System/360, FORTRAN IV, and PL/1. McGraw-Hill, New York, NY, USA, 1969. xv + 395 pp.

Masri:1971:FPC

[MW71a] F. N. Masri and I. R. Williams. I. A Fortran program for calculating degenerate Raman bands of symmetric tops with an adaptation for infrared bands. Computer Physics Communications, 2 (5):298, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465571900051.

Masri:1971:IFP

[MW71b] F. N. Masri and I. R. Williams. II. A Fortran program for calculating degenerate Raman bands of spherical tops with an adaptation for infrared bands. Computer Physics Communications, 2 (5):299, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465571900063.

McBride:1975:TFI

[MW75] James R. McBride and David J.

Weiss. TETREST: a Fortran IV program for calculating tetrachoric correlations. Research report 75-2, Department of Psychology, University of Minnesota, Minneapolis, MN, USA, 1975. 39 pp. Distributed by National Technical Information Service.

Myers:1973:SFS

[Mye73] Wayne L. Myers. A system of FORTRAN subroutines for simplified input to computers. Research report 196, Michigan State University, Agricultural Experiment Station, East Lansing, MI, USA, 1973. 23 pp.

Mardia:1975:SAAc

[MZ75] K. V. Mardia and P. J. Zemroch. Statistical algorithms: Algorithm AS 84: Measures of multivariate skewness and kurtosis. Applied Statistics, 24(2):262–265, June 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/84.

Nag:1978:NFL

[Nag78] Nag. NAg FORTRAN library manual. NAg [Downers Groves, Ill], Oxford, UK, mark 7 edition, 1978. 5 v. pp. Distributed by NAg (USA).

Nag:1980:NFM

[Nag80a] Nag. NAg FORTRAN mini manual. NAg [Downers Groves, Ill], Oxford, UK, mark 8 edition, 1980.

337 pp. Distributed by NAg (USA).

Nagata:1980:FLM

[Nag80b] Hiroyasu Nagata. Formal: a language with a macro-oriented extension facility. Computer Languages, 5(2):65–76, ???? 1980. CODEN COLADA. ISSN 0096-0551.

Nake:1968:BRR

[Nak68] F. Nake. Book report: T. R. McCalla, Introduction to Numerical Methods and FORTRAN Programming. Computing, 3(1): 84, 1968. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Nakagawa:1977:RMF

[Nak77] Tomoyasu Nakagawa. Reference manual of FORTRAN program ILLOD-(AND-OR-B) for optimal AND-OR networks. Report UIUCDCS-R-77, 883, Dept. of Computer Science, University of Illinois, Urbana, IL, USA, 1977. 35 + [80] pp.

NCCL:1970:AUM

[Nat70a] National Computing Centre Limited, Manchester, UK. APSE user manual: FORTRAN version, 1970. various pp.

NCCL:1970:SFP

[Nat70b] National Computing Centre Limited. Standard Fortran programming manual. Computer standards series. National Computing Centre, Manchester, UK, 1970. ISBN 0-85012-021-7. [3] + 152 pp. LCCN QA76.73.F25 N37. Distributed in the U.S.A. by Science Associates/International, Inc., 23 East 26th Street, New York, N.Y. 10010. Errata sheet (3 p.) inserted.

NCCL:1972:SFP

[Nat72] National Computing Centre Limited. Standard Fortran programming manual. Computers and the professional. National Computing Centre, Manchester, UK, second edition, 1972. ISBN 0-85012-103-5, 0-85012-063-2 (paperback).

[5] + 152 + 3–44 pp. LCCN QA76.73.F25 N37 1972.

NCR:1973:NCF

[Nat73] National Cash Register Company.

NCR century Fortran: student text. National Cash Register Co., Dayton, OH, USA, second edition, 1973. 155 pp.

Naur:1975:PLN

[Nau75] Peter Naur. Programming languages, natural languages, and mathematics. Comm. ACM, 18 (12):676–683, December 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Papers from the Second ACM Symposium on Principles of Programming Languages (Palo Alto, Calif., 1975).

Navon:1978:TNA

[Nav78] I. M. Navon. TWISK 25, a nonlinear alternating direction implicit FORTRAN IV program for solving the shallow-water equations. Technical report, National Research Institute for Mathematical Sciences, Pretoria, South Africa, May 1978. 24 + 27 pp.

Norrod:1970:SCFa

[NBH70a] Valmer Norrod, Sheldon Blecher, and Martha Horton. A self-study course in Fortran program[m]ing. NASA contractor report NASA CR-1478, National Aeronautics and Space Administration, Washington, DC. USA, 1970. various pp. For sale by the Clearinghouse for Federal Scientific and Technical Information, Springfield, VA, USA.

Norrod:1970:SCFb

[NBH70b] Valmer Norrod, Sheldon Blecher, and Martha Horton. A self-study course in Fortran program[m]ing. NASA contractor report NASA CR-1478, National Aeronautics and Space Administration, Washington, DC. USA, 1970. various pp. For sale by the Clearinghouse for Federal Scientific and Technical Information, Springfield, VA, USA.

Nelson:1975:PPF

[NC75] James M. Nelson and Charles E. Cohn. Parallel processing in FOR-TRAN with floating-point hardware. Software—Practice and Experience, 5(1):65–68, January/ March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ng:1976:RTI

[NC76] T. S. Ng and A. Cantoni. Run time interaction with Fortran using mixed code. The Computer Journal, 19(1):91–92, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

NCR:1969:NCF

[NCR69] NCR Corporation. NCR Century FORTRAN reference manual. Century Software Distribution Center, NCR, Dayton, OH, USA, 1969. various pp.

NCR:1970:NCF

[NCR70] NCR Corporation. NCR Century Fortran. National Cash Register Co., Dayton, OH, USA, 1970. 147 pp.

Nebe:1971:DSW

[Neb71] H. H. Nebe. Datenverarbeitung für Studenten der Wirtschafts- und Sozialwissenschaften. IBM Corporation, New York, NY, USA, second edition, 1971.

Neely:1975:EAF

[Nee75] Peter M. Neely. Editorial: After FORTRAN, what? Software—Practice and Experience, 5 (1):1–4, January 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nehrkorn:1974:MRR

[Neh74] K. Nehrkorn. Methoden Der Realisierung Einer Reellen Intervallarithmetik In Einem Fortran-System. Wiss. Z. Techn. Hochsch. Karl-Marx-Stadt, 16: 469–474, 1974. CODEN WZTKAY. ISSN 0372-7610.

Newnham:1967:FPS

[New67] R. M. Newnham. A FORTRAN programme to simulate pulpwood harvesting machines. Forest Management Institute. Information report FMR-X-7, Dept. of Forestry and Rural Development, Ottawa, Ontario, Canada, 1967. 32 pp.

Newnham:1972:FPE

[New72] R. M. Newnham. A FORTRAN program for estimating the operating cost per available machine hour of a logging machine. Forest Management Institute. Information report FMR-X-42, Canadian Forestry Service, Dept. of the Environment, Ottawa, Ontario, Canada, 1972. v + 22 pp.

Newman:1973:RKD

[New73] M. Newman. Reading of keywordlabelled data in random order and in free format with a computer program in IBM Fortran IV: user’s guide to sub-routine INP3F. CSIR guide K34, Council for Scientific and Industrial Research, Pretoria, South Africa, 1973. ISBN 0-7988-0444-0. 37 pp. Distributed by National Technical Information Service.

Newsted:1975:GAP

[New75] Peter R. Newsted. Grade and ability predictions in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(2):87–91, June 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Newnham:1976:FPL

[New76] R. M. Newnham. The FOR-TRAN program for LOGPLAN: a model for planning logging operations. Forest Management Institute. Information report FMR-X-78, Canadian Forestry Service, Dept. of the Environment, Ottawa, Ontario, Canada, 1976. iii + 77 pp.

Nickerson:1974:FFP

[Nic74] Robert C. Nickerson. Fundamentals of FORTRAN programming. Winthrop Publishers, Cambridge, MA, USA, 1974. ISBN 0-87626-300-7. xi + 300 pp.

Nicholls:1975:SDP

[Nic75a] J. E. Nicholls. The Structure and Design of Programming Languages. Addison-Wesley, Reading, MA, USA, 1975. ISBN 0-201-14454-9.

Nickerson:1975:FFP

[Nic75b] Robert C. Nickerson. Fundamentals of FORTRAN programming. Winthrop Publishers, Cambridge, MA, USA, 1975. ISBN 0-87626-300-7. xi + 300 pp. LCCN ????

Nickerson:1975:IMF

[Nic75c] Robert C. Nickerson. Instructor’s manual for Fundamentals of FOR-TRAN programming. Winthrop Publishers, Cambridge, MA, USA, 1975. 63 + 45 pp.

NZC:1978:FRM

[Nic78] Nicolet Zeta Corp. FORTRAN reference manual: fundamental plotting subroutines. Publication 431-014, Nicolet Zeta Corp., Concord, CA, USA, 1978. various pp. Release 5.3 of March 1978.

Nickerson:1980:FFPa

[Nic80a] Robert C. Nickerson. Fundamentals of FORTRAN programming. Little, Brown and Co., Boston, MA, USA, second edition, 1980. ISBN 0-316-60644-8. xi + 450 pp. LCCN QA76.73.F25N49 1982. Reprinted, 1982.

Nickerson:1980:FFPb

[Nic80b] Robert C. Nickerson. Fundamentals of FORTRAN programming. Winthrop Publishers, Cambridge, MA, USA, second edition, 1980. ISBN 0-87626-301-5. xi + 450 pp. LCCN QA76.73.F25N49 1980.

Nickerson:1980:IMF

[Nic80c] Robert C. Nickerson. Instructor’s manual for Fundamentals of FOR-TRAN programming. Winthrop Publishers, Cambridge, MA, USA, second edition, 1980. ISBN 0-87626-302-3. iii + 130 pp.

Nielsen:1968:FPB

[Nie68] Gordon L. Nielsen. FORTRAN primer for business and economics. D. H. Mark Pub. Co., Braintree, MA, USA, 1968. x + 161 pp.

Nielsen:1971:FPB

[Nie71] Gordon L. Nielsen. FORTRAN primer for business and economics. d.h. mark publications. General Learning Press, Morristown, NJ, USA, 1971. x + 161 pp.

Niemeyer:1972:SSH

[Nie72a] G. Niemeyer. Die Simulation von Systemablaeufen mit Hilfe von Fortran IV — GPSS auf Fortran-Basis. Walter de Gruyter, New York, NY, USA, 1972. ISBN 3-11-004024-7. 239 + [2] pp.

Niemeyer:1972:SSM

[Nie72b] G. Niemeyer. Die Simulation von Systemablaeufen mit Hilfe von Fortran IV — GPSS auf Fortranbasis. Walter de Gruyter, New York, NY, USA, 1972. ISBN 3-11-004024-7.

Niessner:1972:RAE

[Nie72c] Herbert Niessner. Remark on “Algorithm 343: Eigenvalues and eigenvectors of a real general matrix”. Comm. ACM, 15(6):466, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [?].

Niemeyer:1975:PF

[Nie75] G. Niemeyer. Programmierkurs Fortran. Franz Vahlen, München, Germany, 1975. ISBN 3-8006-0529-5.

NDKG:1969:FPN

[Nih69] Nihon Denki Kabushiki Gaisha.

FORTRAN puroguramingu nyumon. Number 2 in Iidipiesu nyumon shirizu. Nihonnoritsukyokai, Tokyo, Japan, 1969. 249 pp.

Nikolai:1978:USF

[Nik78] P. J. Nikolai. USA Standard Fortran implementation of simultaneous iteration for real diagonizable matrices. SIAM Review, 20(3): 631–632, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nirschl:1969:IFP

[Nir69] N. Nirschl. Introduction to FOR-TRAN programming for the IBM 1130. ????, De Pere, WI, USA, 1969. v + 32 pp.

Nishimura:1978:ZKF

[Nis78]  Hirohiko Nishimura.  Zinbun

kagakuno FORTRAN sitizyusiti. Tokyodaigakushuppankai, Tokyo, Japan, 1978. ISBN 4-13-062060-6. 234 pp.

Nyhoff:1968:FES

[NL68] Larry R. Nyhoff and Sanford Leestma. FORTRAN 77 for engineers and scientists. Collier Macmillan, London, second edition, 1968. ISBN 0-02-388631-5 (paperback). xii + 590 pp. LCCN QA76.73.F25 N9 1988 Check local catalog for specific location.

Nakagawa:1971:RMF

[NL71] Tomoyasu Nakagawa and Hung-Chi Lai. Reference manual of Fortran program ILLOD-(NOR-B) for optimal NOR networks. Technical Report UIUCDCS-R-71-488, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, December 1971. ii + 33 + [69] pp. Replaced by 1129.

Nishisato:1975:OFI

[NL75] Shizuhiko Nishisato and Kuo-Sing Leong. OPSCAL: a Fortran IV program for analysis of qualitative data by optimal scaling. Measurement of evaluation of categorical data technical report 3, Dept. of Measurement and Evaluation, Ontario Institute for Studies in Education, Toronto, Ontario, Canada, 1975. 119 pp.

Navon:1970:FAF

[NM70] I. Michael Navon and U. Muller.

FESW — a finite-element Fortran IV program for solving the shallow water equations. Advances in engineering software, 1(??):77–84, 1970. CODEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Notto:1974:FC

[NM74] Ralph W. Notto and Robert C.

Marien. FORTRAN consultant. Data Tactics, Gaithersburg, MD, USA, 1974. 79 pp.

Navon:1978:TFF

[NM78] I. M. Navon and U. Muller.

TWISK 29: FESW-a finiteelement FORTRAN IV program for solving the shallow-water equations. Technical Report ????, National Research Institute for Mathematical Sciences, Pretoria, South Africa, May 1978. 33 + 26 pp.

Nance:1972:IFR

[NO72] Richard E. Nance and Claude Overstreet. Implementation of Fortran random number generators on the SRU 1108. Technical report CP-72022, Computer Science/Operations Research Center, Institute of Technology, Southern Methodist University, Dallas, TX, USA, 1972. 14 pp.

Nance:1975:IFR

[NO75] Richard E. Nance and Claude Overstreet, Jr. Implementation of Fortran random number generators on computers with one’s complement arithmetic. Journal of Statistical Computation and Simulation, 4(3):235–243, ???? 1975. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. URL http://

                                        www.tandfonline.com/doi/abs/
                                        10.1080/00949657508810126.

Nolan:1971:FIC

[Nol71] Richard L. Nolan. FORTRAN IV computing and applications. Addison-Wesley, Reading, MA, USA, 1971. xi + 365 pp.

Norment:1963:CFP

[Nor63] H. G. Norment. A collection of FORTRAN programs for crystal structure analysis. Nrl report 5885, U.S. Naval Research Laboratory, Washington, DC, USA, April 29, 1963. 71 pp.

Nordstrom:1966:PAD

[Nor66] Edwin Alan Nordstrom. Part I: Acid dissociation constants of phenol in light and heavy water. part II: A Fortran II program for plotting contour maps. Thesis (m.s.), Dept. of Chemistry, Massachusetts Institute of Technology, Cambridge, MA, USA, 1966. [6] + 41 pp.

NorthernVirginiaEducationalTelevision:1970:SMF

[Nor0 ] Northern Virginia Educational Television. Speak to me in FOR-TRAN, 1970 (??).

Newnham:1969:FPS

[NS69] R. M. Newnham and S. Sjunnesson. A FORTRAN program to simulate harvesting machines for mechanized thinning. Forest Management Institute. Information report FMR-X-23, Canadian Forestry Service, Dept. of Fisheries and Forestry, Ottawa, Ontario, Canada, 1969. x + 48 + [25] pp.

Newton:1976:TBP

[NS76] Glen E. Newton and J. Denbigh Starkey. Teaching both PL/I and Fortran to beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):106–107, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

Nordstroem:1971:LFF

[NSB71] M. Nordstroem, E. Sandewall, and D. Breslaw. LISP F1, a FOR-TRAN implementation of LISP 1.5. Report, Uppsala Universitet, Dept. of Computer Sciences, Datalogilaboratoriet, Inst. f. Informationsbehandling, Uppsala, Sweden, 1971.

Nuttall:1976:FPM

[Nut76] Albert H. Nuttall. FORTRAN program for multivariate linear predictive spectral analysis, employing forward and backward averaging. NUSC technical document 5419, Naval Underwater Systems Center, New London, CT, USA, 1976. 22 pp.

Nutt:1978:CPF

[Nut78] Gary J. Nutt. A comparison of PASCAL and FORTRAN as introductory programming languages. ACM SIGPLAN Notices, 13(2): 57–62, February 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ng:1978:FPM

[NY78] P. H. Ng and G. Young. A 1900 FORTRAN post mortem dump system. Software—Practice and Experience, 8(4):421–427, July/ August 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nydegger:1968:ICP

[Nyd68] Adolph C. Nydegger. An introduction to computer programming, with an emphasis on FOR-TRAN IV. Addison-Wesley, Reading, MA, USA, 1968. ix + 269 pp.

Obrien:1975:CBM

[O’b75] J. A. O’brien. Computers in Business Management: An Introduction. R. D. Irwin, Homewood, IL, USA, 1975. ISBN 0-256-01671-2.

Obregon:1970:CMT

[Obr70] Roberto Obregon. Computers: a management tool; data processing fundamentals, Cobol and Fortran programming languages. Matrix, Inc., Dallas, TX, USA, 1970. iv + 178 pp.

Obregon:1971:CMT

[Obr71] Roberto Obregon. Computers:

a management tool; introduction to computer science, Fortran IV and ANSI Cobol programming languages. Matrix, Inc., Dallas, TX, USA, 2d, rev. edition, 1971. xi + 209 pp.

ODonnell:1965:FCN

[O’D65] Hugh Wilson O’Donnell. Flexural constants for non-uniform members utilizing step functions and FORTRAN II computer language. Thesis (m.s. in civil engineering), Louisiana Polytechnic Institute, Ruston, LA, USA, August 1965. v + 59 pp. School renamed to Louisiana Technical University in 1970.

ODonohue:1974:SLA

[O’D74] Alice Ann O’Donohue. A series of learning activity packages to facilitate the learning of FOR-TRAN IV for competent high school students. Masters’ thesis, State University of New York, College of Arts and Science Professional Studies Division, Oswego, NY, USA, 1974. v + 40 pp.

Oertel:1971:FPU

[Oer71] Alfred Charles Oertel. Fortran programmes for use in fluorescent X-ray analyses of silicates, soils, and vegetable matter. Commonwealth Scientific and Industrial Research Organization. Division of Soils. Technical paper 4, CSIRO, Melbourne, Victoria, Australia, 1971. ISBN 0-643-00870-5. 24 pp.

Osterweil:1976:DVE

[OF76] Leon J. Osterweil and Lloyd D. Fosdick. DAVE: a validation error detection and documentation system for Fortran programs. Software—Practice and Experience, 6(4):473–486, October/December 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Organick:1978:PLS

[OFP78] Elliott I. Organick, Alexandra I. Forsythe, and Robert P. Plummer. Programming Language Structures. Academic Press, New York, NY, USA, 1978. ISBN 0-12-528260-5.

Ondrick:1969:FIC

[OG69] Charles W. Ondrick and John C. (John Cedric) Griffiths. Fortran IV computer program for fitting observed count data to discrete distribution models of binomial, Poisson and negative binomial. Computer contribution 35, Kansas Geological Survey, Lawrence, KS, USA, 1969. 20 pp.

Ojakangas:1970:FIP

[Oja70] Dennis Roger Ojakangas. FOR-TRAN IV program for simulating geologic development of sedimentary basins. Computer contribution 49, University of Kansas, Lawrence, KS, USA, 1970. 52 pp.

OKins:1964:DAT

[O’K64] Charles Eugene O’Kins. The development of analytic techniques and FORTRAN programs for the evaluation of plane motion mechanisms. Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1964. 99 pp.

Ohtsuki:1972:ATS

[OK72] Tatsuo Ohtsuki and Hiroshi Kato.

An accuracy test system for FOR-TRAN library functions. NEC Res. Development, 24:1–15, 1972.

SaezOlivito:1971:RFI

[Oli71] Angel Saez Olivito. Rutinas en FORTRAN IV para programacion lineal, de utilizacion en la empresa agropcuaria. Ganaderas del Ebro. Comunicaciones 4, Instituto de Economia y Producciones Ganaderas del Ebro, Zaragoza, Spain, 1971. 89 pp.

Ollman:1971:EFI

[Oll71] Mark Eugene Ollman. Extension of FORTRAN by the introduction of a pre-processor. Thesis (m. s.), Chico State College, Chico, CA, USA, 1971. Two volumes.

OLeary:1966:FIM

[OLS66] Mont O’Leary, R. H. Lippert, and Owen T. Spitz. FORTRAN IV and MAP program for computation and plotting of trend surfaces for degrees 1 through 6. Computer contribution 3, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1966. 48 pp.

Organick:1974:FI

[OM74] Elliott Irving Organick and Loren P. Meissner. Fortran IV. Addison-Wesley, Reading, MA, USA, second edition, January 1974. ISBN 0-201-05503-1. ix + 293 pp. LCCN QA76.5 .O72 1974. US$15.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0-201-05503-1. First ed. pub-

lished in 1966 under title A Fortran IV primer.

ONeill:1974:SPI

[O’N74] Dennis M. O’Neill. SFOR — a precompiler for the implementation of a FORTRAN-based structured language. ACM SIGPLAN Notices, 9(12):22–29, December 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ono:1979:BAF

[Ono79a] Kiyoshi Ono. BFORT — A Fortran system with arbitrary precision integer and real arithmetic. Technical report, Department of Physics, University of Tokyo, Tokyo, Japan, January 1979.

Ono:1979:BFS

[Ono79b] Kiyoshi Ono. BFORT — A Fortran system with arbitrary precision integer and real arithmetic. Technical report, Department of Physics, University of Tokyo, Tokyo, Japan, January 1979.

Olivieri:1975:CPA

[OR75] Peter Olivieri and Michael W. Rubin. Computers and Programming: a Neoclassical Approach. McGraw-Hill, New York, NY, USA, 1975. ISBN 0-07-047692-6.

Oldehoeft:1977:MTI

[OR77] R. R. Oldehoeft and R. V. Roman. Methodology for teaching introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):123–128, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Organick:1961:DEE

[Org61a] Elliott Irving Organick. Drill exercises and example problems for A primer for programming with the FORTRAN language. Technical report, Computing and Data Processing Center, University of Houston, Houston, TX, USA, 1961. v + 105 pp.

Organick:1961:PPF

[Org61b] Elliott Irving Organick. A primer for programming with the FOR-TRAN language. Technical report, Computing and Data processing Center, University of Houston, Houston, TX, USA, 1961. vi + 101 pp.

Organick:1963:FP

[Org63] Elliott Irving Organick. A FOR-TRAN primer. Addison-Wesley series in computer science and information processing. Addison-Wesley, Reading, MA, USA, 1963. vi + 186 pp.

Organick:1966:FIP

[Org66a] Elliott Irving Organick. A Fortran IV primer. Addison-Wesley series in computer science and information processing. Addison-Wes-ley, Reading, MA, USA, 1966. viii + 263 pp.

Organick:1966:FP

[Org66b] Elliott Irving Organick. A FOR-TRAN primer. Addison-Wesley series in computer science and information processing. Addison-Wesley, Reading, MA, USA, 1966. 186 pp.

Organick:1972:FI

[Org72] Elliott Irving Organick. Fortran IV. Fondo Educativo Interamericano, Mexico, DF, Mexico, 1972. vii + 301 pp.

Osin:1975:FI

[Osi75] Luis Osin. FORTRAN IV. Universidad de la Republica, Division Publicaciones y Ediciones, Montevideo, Uruguay, 1975. 89 pp.

Osin:1977:FI

[Osi77] Luis Osin. FORTRAN IV. Universidad de la Republica, Division Publicaciones y Ediciones, Montevideo, Uruguay, 1977. 89 pp.

Oster:1962:MTB

[Ost62] C. A. Oster. More on testing BCD words with FORTRAN. Comm. ACM, 5(11):545, November 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oster:1964:LBM

[Ost64] C. A. Oster. Limited bit manipulation using FORTRAN II.

Comm. ACM, 7(12):719–721, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Osyczka:1976:AOS

[Osy76] Andrzej Osyczka. An algorithm of optimization for a special class of networks. Computing, 16(1–2): 77–97, 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Olszewski:1980:PPV

[OT80] E. A. Olszewski and W. J. Thompson. A program for perspective views of three-dimensional surfaces. Computer Physics Communications, 21(2):185–193, December 2, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465580900399.

Ott:1978:FPC

[Ott78] Wayne Ott. A Fortran program for computing the pollutant standards index (PSI). Environmental monitoring series EPA-600/4-78-001, Environmental Protection Agency, Office of Research and Development, Office of Monitoring and Technical Support, Monitoring Technology Division, Washington, DC. USA, May 1978. vi + 65 pp.

Overstreet:1972:FVP

[Ove72] Claude Overstreet. A FORTRAN V package for testing and anaylsis of pseudorandom number generators. Technical report CP-72009, Computer Science/Operations Research Center, Institute of Technology, Southern Methodist University, Dallas, TX, USA, 1972. 7 + [24] pp.

Oviedo:1977:SDS

[Ovi77] Enrique I. Oviedo. The syntactic definition of structured FOR-TRAN. Technical report 120, State University of New York at Buffalo, Dept. of Computer Science, Buffalo, NY, USA, 1977. 60 pp.

IBM:1962:PPF

[Owe62] Robert H. Owens. Programmer’s primer for FORTRAN for IBM data processing systems. Technical report, University of New Hampshire, Durham, NH, USA, 1962. 89 pp.

Owens:1965:LEA

[Owe65a] David Owens. Letter to the Editor: Avoiding bit manipulation in FORTRAN. Comm. ACM, 8 (11):649, November 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Owens:1965:LEC

[Owe65b] David Owens. Letter to the Editor: comments on recent FOR-TRAN ideas. Comm. ACM, 8 (12):788, December 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Owens:1970:EXF

[Owe79] D. H. (David H.) Owens. Extended XDS FORTRAN IV: making FORTRAN easier to use. Xerox Data Systems, El Segundo, CA, USA, 1970 (or 1979??). ii + 16 pp.

Packard:1969:PGH

[Pac69] Hewlett Packard. A Pocket Guide to Hewlett–Packard Computers. Cupertino, CA, USA, 1969.

Page:1974:AAM

[Pag74a] R. L. Page. ACM Algorithm 479:

a minimal spanning tree clustering method [Z]. Comm. ACM, 17(6):321–323, June 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?, ?].

Pager:1974:EUP

[Pag74b] David Pager. On eliminating unit productions from LR(k) parsers. In Loeckx [Loe74], pages 242–254.

Palme:1968:CBS

[Pal68] Jacob Palme. A comparison between Simula and Fortran. BIT (Nordisk tidskrift for informationsbehandling), 8(3):203–209, September 1968. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.
com/openurl.asp?genre=article&
issn=0006-3835&volume=8&issue=
3&spage=203.

Pankhurst:1970:CPG

[Pan70] R. J. Pankhurst. A computer program for generating diagnostic keys. The Computer Journal, 13(2):145–151, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_02/130145.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/145.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/146.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/
147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
13/Issue_02/tiff/149.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_
02/tiff/150.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/
151.tif.

Parks:1970:FIP

[Par70] James Marshall Parks. FORTRAN IV program for Q-mode cluster analysis on distance function with printed dendrogram. Computer contribution 46, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1970. 32 pp.

Partridge:1975:DDW

[Par75] D. Partridge. Dynamic database which automatically removes unwanted generalisation for the efficient analysis of language features that exhibit a disparate frequency distribution. The Computer Journal, 18(1):43–48, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Park:1977:EME

[Par77] Soo Kyung Park. An evaluation of ML/I (EPS) macros for structured FORTRAN extensions. Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1977. 59 + 15 pp.

Parker:1978:MDM

[Par78] R. A. Parker. More on detecting misspelt identifiers in Fortran. Software—Practice and Experience, 8(1):109, January 1978. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Patterson:1967:CRF

[Pat67] Keith Patterson. The causes of revolution: a Fortran model of Orwell’s Animal Farm. Thesis (b.s.), Dept. of Political Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1967. 66 + 12 pp.

Patterson:1973:AAA

[Pat73a] T. N. L. Patterson. ACM Algorithm 468: Algorithm for automatic numerical integration over a finite interval [D1]. Comm. ACM, 16(11):694–699, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:1973:BRB

[Pat73b] T. N. L. Patterson. Book review: Fortran techniques: A. Colin Day, Cambridge University Press, 1972, £2.50. Computer Physics Communications, 5(5):397, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465573900672.

Patterson:1974:UMF

[Pat74] M. R. Patterson. A user’s manual for the FORTRAN IV version of the Wisconsin Hydrologic Transport Model. Technical Report ORNL-NSF-EATC-7, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1974. vii + 252 pp.

Patel:1977:MSD

[Pat77] Arvind D. Patel. A modular symbolic differentiation system written in Fortran. Thesis (m.s.), Dept. of Computer Science, College of Natural Sciences and Mathematic, University of Houston, Houston, TX, USA, 1977. ca. 200 pp.

Paulin:1971:FAL

[Pau71a] G. Paulin. Fortran-Training, Aufgaben mit Loesungen. VEB Verlag Technik, Berlin, Germany, 1971.

Paulin:1971:FAM

[Pau71b] G. Paulin. Fortran-training, Aufgaben mit Loesungen. Veb Verlag Technik, Berlin, Germany, 1971.

Pawlicki:1965:F

[Paw65a] Gerard S. Pawlicki. FORTRAN, 1965.

Pawlicki:1965:FPD

[Paw65b] Gerard S. Pawlicki. FORTRAN programming of digital computers,

Payne:1964:FP

[Pay64] Norman Gregson Payne. 1620 FORTRAN programming. Where did this appear?, 1964.

Payne:1970:FTP

[Pay70] W. H. Payne. Fortran Tausworthe pseudorandom number generator. Comm. ACM, 13(1):57, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pereyra:1967:TCS

[PB67] Walter T. Pereyra and Dina Barzel. Trawl catch summarization program, IBM 7090/7094, Fortran IV. Transactions of the American Fisheries Society, 96(3):363– 364, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1973:FPW

[PB73a] James Laurie Parker and Marilyn Bohl. FORTRAN programming and WATFIV. Science Research Associates, Chicago, IL, USA, 1973. xvi + 284 pp.

Peller:1973:WFF

[PB73b] Ildiko C. Peller and Anne K.

Baron. WASP: a flexible FOR-TRAN IV computer code for calculating water and steam properties. NASA technical note NASA TN D-739, National Aeronautics and Space Administration; National Technical Information Service, Washington, DC. USA, 1973. 116 pp.

Paige:1974:USP

[PB74a] M. R. Paige and J. P. Benson. The use of software probes in testing FORTRAN programs. Computer, 7(7):40–47, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Presser:1974:ECD

[PB74b] L. Presser and J. Benson. Evaluation of compiler diagnostics. The Computer Journal, 17(2):121–123, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Pollack:1964:GFI

[PC64] Seymour V. Pollack and Sadi Crawford. A guide to FORTRAN IV. Volunteer Braille Services, Washington, DC, USA, 1964. various pp.

Pore:1967:FPC

[PC67] N. A. Pore and R. A. Cummings. A Fortran program for the calculation of hourly values of astronomical tide and time and height of high and low water. Technical report, Weather Bureau, Systems Development Office, Techniques Development Laboratory, ????, 1967. 17 pp.

Pereira:1971:PNBa

[PC71a] R. Agonia Pereira and Duarte Costa Cabral. Programacão NEAT, BA-SIC, FORTRAN e ALGOL. Centro de Calculo Cenctifico, Oeiras, Portugal, 1971. 336 pp.

Pereira:1971:PNBb

[PC71b] R. Agonia Pereira and Duarte Costa Cabral. Programmacão NEAT, BASIC, FORTRAN e ALGOL. Centro de Calculo Cientifico, Oeiras, Portugal, 1971. 336 pp.

Parkin:1978:EEB

[PC78a] A. Parkin and R. B. Coats.

EDSIM-event based discrete simulation using general purpose languages such as Fortran. The Computer Journal, 21(2):122–127, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
21/2/122.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/
122.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
21/Issue_02/tiff/124.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
02/tiff/125.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/
126.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/127.
tif.

Pohl:1978:ICS

[PC78b] Jens G. Pohl and Jeff Conrad.

Introduction to computer systems and programming in architecture and construction: FORTRAN IV and BASIC. EDUCOL, San Luis Obispo, CA, USA, 1978. xiii + 432 pp.

Chiu:1979:FCH

[pC79] Yu pu Chiu. FORTRAN cheng hsu she chi. Ko hsueh chu pan she: Hsin hua shu tien Pei-ching fa hsing so fa hsing, Pei-ching, ti 1 pan edition, 1979. vi + 496 pp.

Pesamosca:1976:FEP

[PCR76] Giancarlo Pesamosca, M. L. Lo Cascio, and G. Ravaioli. FOR-TRAN: elementi di programmazione. La goliardica, Roma, Italy, 1976. 170 pp.

Pierce:1967:IMM

[PD67] Frances Pierce and C. J. Dicostanzo. Instantaneous mortality and maturation schedules IBM 1620, Fortran II. Transactions of the American Fisheries Society, 96 (3):366, July 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1976:WH

[PD76] Rex L. Page and Richard L. Didday. WATFIV for humans. West Publishing Company, St. Paul, MN, USA, 1976. ISBN 0-8299-0100-0. vii + 492 pp. LCCN QA76.73.F25P33. US$11.95.

Page:1980:FH

[PD80a] Rex L. Page and Richard L. Didday. Fortran 77 for humans. West Publishing Company, St. Paul, MN, USA, 1980. ISBN 0-8299-0271-6. xi + 476 pp. LCCN QA76.73.F25 P3. US$10.95. Two volumes.

Page:1980:IMA

[PD80b] Rex L. Page and Richard L. Didday. Instructor’s manual to accompany FORTRAN 77 for humans. West Publishing Company, St. Paul, MN, USA, 1980. 88 pp.

Peck:1977:USD

[Pec77] John C. Peck. The university — a systems development center for state government or how to solve the education vs. training problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):1– 5, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Pennington:1968:CIC

[Pen68] Jack D. Pennington. A computerassisted instruction (CAI) of FOR-TRAN IV. Thesis (m.s.), University of Southwestern Louisiana, Lafayette, LA, USA, 1968. 105 pp.

Pennington:1970:ICM

[Pen70] Ralph H. Pennington. Introductory Computer Mehtods and Numerical Analysis. MacMillan Publishing Company, New York, NY, USA, January 1970. ISBN 0-02-393830-7. ???? pp. LCCN ????

Perritt:1972:PFI

[Per72a] Roscoe D. Perritt. PL/1 and FORTRAN IV computer exercises: accounting problems and practice case to accompany Financial Accounting and Managerial Accounting. Macmillan Co.,; Collier-Macmillan, New York, London, 1972. 103 pp.

Perritt:1972:SMP

[Per72b] Roscoe D. Perritt. Solutions manual for PL/1 and FORTRAN IV computer exercises: accounting problems and practice case to accompany Financial Accounting and Managerial Accounting. Macmillan; Collier-Macmillan, New York, NY, USA, 1972. 72 pp.

Perlis:1975:ICS

[Per75] A. J. Perlis. Introduction to Computer Science. Harper & Row, New York, NY, USA, second edition, 1975. ISBN 0-06-045128-9.

Perde:1977:SRA

[Per77] Michael Lee Perde. A study of rational arithmetic modeled for the Fortran IV language. Thesis (m.s.), University of Arkansas, Fayetteville, Fayetteville, AR, USA, 1977. 79 pp.

Perlis:1978:ASD

[Per78] Alan J. Perlis. The American side of the development of Algol. ACM SIGPLAN Notices, 13(8):3– 14, August 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perry:1980:FCF

[Per80] William J. Perry. FABSTAT: a card-imaged FORTRAN program to be used for axially distributed planar-structural data. Open-file report 81-158, U.S. Dept. of the Interior, Geological Survey, Reston, VA, USA, 1980. ???? pp.

Peterson:1974:IPL

[Pet74] W. W. Peterson. Introduction to Programming Languages. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1974. ISBN 0-13-493486-5.

Petersen:1976:EF

[Pet76] Thomas Michael Henry Petersen.

Elementary Fortran. The M. and E. computer programming series. MacDonald and Evans, London, UK, 1976. ISBN 0-7121-0548-4. xiv + 162 pp.

Petrus:1980:PFS

[Pet80] Octavian Petrus. Programarea in FORTRAN: stil in programare. “Junimea”, Iasi, ??, 1980. 205 + [3] pp.

Paulik:1965:WLR

[PG65] G. J. Paulik and Lawrence E.

Gales. Weighted linear regression for two variables, IBM 709, Fortran II. Transactions of the American Fisheries Society, 94(2):196, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1966:FIP

[PG66] Jack Warren Pierce and Donald I.

Good. FORTRAN II program for standard-size analysis of unconsolidated sediments using an IBM 1620 computer. Special distribution publication — State Geological Survey 28, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1966. 19 pp.

Pearson:1967:FCP

[PG67] K. G. Pearson and A. R. Glasson. A FORTRAN 11 computer program for steady state aeroelastic evaluation of slender non-rolling rocket vehicles. Technical note, hsa 128, Weapons Research Establishment (Australia), Adelaide, South Australia, 1967. 54 pp.

Pepin:1963:IAF

[PH63] Gerard Roger Pepin and Frank Earl Herin. Improvement of AFIT 1620 FORTRAN. Thesis (m.s.), Air Force Institute of Technology, Wright-Patterson Air Force Base, OH, USA, 1963. 154 pp.

Peterson:1971:FII

[PH71] W. Wesley (William Wesley) Peterson and Jean L. Holz. Fortran IV and the IBM 360. McGraw-Hill, New York, NY, USA, 1971. ix + 205 pp.

Prudom:1977:SPC

[PH77] A. Prudom and M. A. Hennell. Some problems concerning the automatic translation of Fortran to Algol 68. ACM SIGPLAN Notices, 12(6):138–143, June 1977. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Phelizon:1976:FIC

[Phe76] Jean François Phelizon. FOR-TRAN IV: [communication avec l’ordinateur]. Collection Informatique. Economica, Paris, France, 1976. ISBN 2-7178-0066-2. 66 pp.

Phelizon:1979:PAI

[Phe79] Jean François Phelizon. Programmation en Assembleur: initiation a partir du Fortran. Masson, Masson, France, 1979. ISBN 2-225-63481-5. 182 pp. LCCN QA76.73.A8P5. FF75.00.

Phillips:1967:FPR

[Phi67] R. E. Phillips. A Fortran program for relativistic kinematic calculations in two-body nuclear reactions. Technical report, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1967. iii + 25 pp.

Phillips:1971:CEPa

[Phi71a] Thomas Gordon Phillips. A comparison of the effectiveness of a programmed text and a computer based display unit media in teaching Fortran IV. Thesis (ed. d.), University of Missouri, (Columbia, Rolla, Kansas City, St. Louis: which ??), MO, USA, 1971. 536 pp.

Phillips:1971:CEPb

[Phi71b] Thomas Gordon Phillips. A comparison of the effectiveness of a programmed text and a computer based display unit media in teaching FORTRAN IV. Technical report, University of Missouri at Columbia, Columbia, MO, USA, 1971. 554 pp.

Picker:1966:FAS

[Pic66] William A. Picker. A FORTRAN autoprogram for solving ordinary linear differential equations with constant coefficients using exact z-transforms of (l/s)n and trapezoidal convolution. Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1966. 65 pp.

Piessens:1973:AAG

[Pie73] Robert Piessens. ACM Algorithm 453: Gaussian quadrature formulas for Bromwich’s integral [D1]. Comm. ACM, 16(8):486–487, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piessens:1974:AAC

[Pie74] Robert Piessens. ACM Algorithm 473: Computation of Legendre series coefficients [C6]. Comm. ACM, 17(1):25, January 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pillsbury:1970:CAA

[Pil70] Wilbur F. Pillsbury. Computer augmented accounting. South-Western Pub. Co., Cincinnati, OH, USA, 1970. ISBN 0-538-01716-3 (v. 1), 0-538-01717-1 (v. 2), 0-538-01718-X (v. 3), 0-538-01307-9 (v. 1, 2 ed.). various pp.

Pinneo:1973:CST

[Pin73] Robert Orin Pinneo. A comparative study of time-sharing vs. batch processing in the teaching of introductory programming in FOR-TRAN. Thesis (Ph.D.), Oregon State University, Corvallis, OR, USA, 1973. [10] + 138 pp.

Pindor:1980:FPD

[Pin80] Andrzej Pindor. Fortran program for deconvolution of experimental data on electrical resistivity of metals and lattice specific heat. Prace Instytutu Fizyki, nr. 87 0137-6594. Zaklad Narodowy im. Ossolinskich, Wroclaw, Poland, 1980. 20 pp.

Pitman:1979:FLT

[Pit79] K. M. Pitman. A FORTRAN ⟶ LISP translator. In Lewis [Lew79b], page ??

Presberg:1975:PIP

[PJ75] David L. Presberg and Neil W.

Johnson. The Paralyzer: IV-TRAN’s parallelism analyzer and synthesizer. ACM SIGPLAN Notices, 10(3):9–16, March 1975. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plummer:1976:WFIb

[PJT76a] L. Niel Plummer, Blair F. Jones, and Alfred H. Truesdell. WATEQF — a FORTRAN IV version of WA-TEQ: a computer program for calculating chemical equilibrium of natural waters. Water resources investigations 76-13, Dept. of the Interior, Geological Survey, Water Resources Division, Reston, VA, USA, 1976. [1] + iii + 61 pp.

Plummer:1976:WFIa

[PJT76b] L. Niel Plummer, Blair F. Jones, and Alfred H. Truesdell. WATEQF — a FORTRAN IV version of WATEQ a computer program for calculating chemical equilibrium of natural waters, user’s guide. Water resources investigations; 76-13 Water resources investigations (Geological Survey (U.S.)); 76- 13. Dept. of the Interior, Geological Survey, Water Resources Division, Reston, VA, USA, 1976. [1] + iii + 61 pp.

Pritsker:1967:GIF

[PK67] A. Alan B. Pritsker and Philip J. Kiviat. GASP II; a Fortran based simulation language. Technical report, Dept. of Industrial Engineering, Arizona State University, Tempe, AZ, USA, 1967. 71 pp.

Pritsker:1969:SGI

[PK69] A. Alan B. Pritsker and Philip J. Kiviat. Simulation with GASP-II: a FORTRAN based simulation language. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1969. ISBN 0-13-810432-8. xii + 332 pp. LCCN QA76.5 .P73.

Plouff:1975:DFF

[Plo75] Donald Plouff. Derivation of formulas and FORTRAN programs to compute gravity anomalies of prisms. Technical report, U.S. Geological Survey, Reston, VA, USA, 1975. 90 pp. Distributed by National Technical Information Service.

Plouff:1977:PDF

[Plo77] Donald Plouff. Preliminary documentation for a FORTRAN program to compute gravity terrain corrections based on topography digitized on a geographic grid. Reports-Open file series — United States Geological Survey; 77-535, U.S. Geological Survey, Reston, VA, USA, 1977. 45 pp.

Plumb:1961:IFS

[Plu61] S. C. (Stephen C.) Plumb. IBM FORTRAN: self teaching. IBM Corporation, Poughkeepsie, NY, USA, second edition, 1961. various pp.

Plumb:1963:IFS

[Plu63] S. C. (Stephen C.) Plumb. IBM FORTRAN: self-teaching. IBM Corporation, Poughkeepsie, NY, USA, third edition, 1963. 345 pp.

Plumb:1964:IFP

[Plu64] Stephen C. (Stephen C.) Plumb.

Introduction to Fortran; a Program for Self-Instruction. Mc-Graw-Hill, New York, NY, USA, January 1964. ISBN 0-07-050350-8. vii + 203 pp. LCCN ???? US$13.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070503508.

Plumb:1965:FIS

[Plu65] S. C. (Stephen C.) Plumb.

FORTRAN for IBM system 360; a programmed instruction course. Science Research Associates, Chicago, IL, USA, 1965. 274 pp.

Pleasant:1980:RFI

[PMBK80] James C. Pleasant, L. M.

McDowell-Boyer, and G. G. Killough. RAGTIME, a FORTRAN IV implementation of a timedependent model for radionuclides in agricultural systems: first progress report. Technical report, U.S. Nuclear Regulatory Commission: Available from GPO Sales Program, Division of Technical Information and Document Control, U.S. Nuclear Regulatory Commission; National Technical Information Service, Washington, DC, USA, 1980. vii + 109 pp.

Plumb:1968:FISb

[PN68a] S. C. (Stephen C.) Plumb and David E. Napper. FORTRAN for IBM System/360. Science Research Associates, Chicago, IL, USA, 1968. 16 pp.

Plumb:1968:FISc

[PN68b] S. C. (Stephen C.) Plumb and David E. Napper. FORTRAN for IBM System/360: a programmed instruction course. Science Research Associates, Chicago, IL, USA, 1968. xiii + 274 pp.

Plumb:1968:FISa

[PN68c] S. C. (Stephen C.) Plumb and David E. Napper. FORTRAN/ 360 for IBM system. Science Research Associates, Chicago, IL, USA, 1968. 274 pp.

Percy:1965:SFP

[PNK65a] John H. Percy, Dhirendra Ramachandra Navaratna, and Stanley Klein. SABOR I: a FOR-TRAN program for the linear elastic analysis of thin shells of revolution under axisymmetric loading by using the matrix displacement method. ASRL TR 121-5, Aeroelastic and Research Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1965. v + 99 pp.

Percy:1965:SIF

[PNK65b] John H. Percy, Dhirendra Ramachandra Navaratna, and Stanley Klein. SABOR III: a FOR-TRAN program for the linear elastic analysis of thin shells of revolution under axisymmetric loading by using the matrix displacement method. ASRL TR 121-6, Aeroelastic and Research Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1965. iii + 99 pp.

Pollack:1965:GFIa

[Pol65a] Seymour V. Pollack. A guide to Fortran IV. Columbia University Press, New York, NY, USA, 1965. 260 pp.

Pollack:1965:GFIb

[Pol65b] Seymour V. Pollack. A guide to Fortran IV. Columbia University Press, New York, NY, USA, 1965. 290 pp.

Pollard:1978:FCP

[Pol78] David D. Pollard. FORTRAN computer program for calculation of stress-intensity factors, stresses, and displacements associated with a fluid-pressurized fracture near the earth’s surface. Open-file series 78-160, U.S. Geological Survey, Reston, VA, USA, 1978. 25 pp.

Pomeranz:1974:AAE

[Pom74] John Pomeranz. ACM Algorithm 487: Exact cumulative distribution of the Kolmogorov–Smirnov statistic for small samples. Comm. ACM, 17(12):703–704, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [?].

Poore:1962:PCM

[Poo62] Jesse H. Poore, Jr. Pracniques: Character manipulation in 1620 FORTRAN II. Comm. ACM, 5 (12):602, December 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Poole:1974:PAC

[Poo74] P. C. Poole. Portable and adaptable compilers. In Bauer and Eickel [BE74], pages 427–497. LCCN QA76.6 .C6281.

Potter:1966:CCF

[Pot66] Stephen Potter. Correction to a comparison of Fortran language implementation. Comm. ACM, 9(6):412, June 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Powell:1968:FSS

[Pow68] M. J. D. Powell. A Fortran subroutine for solving systems of nonlinear algebraic equations. Technical report, Theoretical Physics Division, Atomic Energy Research Establishment, Harwell, Berkshire, UK, 1968. 54 pp.

Powell:1969:FSI

[Pow69] M. J. D. Powell. Fortran subroutine to invert a rectangular matrix of full rank. Report AERE R 6072, Theoretical Physics Division, Atomic Energy Research Establishment, Harwell, Berkshire, UK, 1969. 13 pp.

Powell:1970:FSS

[Pow70a] M. J. D. Powell. A Fortran subroutine for solving systems of nonlinear algebraic equations. In Numerical Methods for Nonlinear Algebraic Equations (Proc. Conf., Univ. Essex, Colchester, 1969), pages 115–161. Gordon and Breach, New York, NY, USA, 1970.

Powell:1970:FSUa

[Pow70b] M. J. D. Powell. A FORTRAN subroutine for unconstrained minimization, requiring first derivatives of the objective function. Report AERE R6469, Her Majesty’s Stationary Office, London, UK, 1970. 46 pp.

Powell:1970:FSUb

[Pow70c] M. J. D. Powell. A FORTRAN subroutine for unconstrained minimization, requiring first derivatives of the objective function. Report AERE R6494, Her Majesty’s Stationary Office, London, UK, 1970. ISBN 0-11-981135-9. ???? pp.

Power:1974:FFL

[Pow74] Marian E. Power. Flex: a Fortran language extension. Thesis (m.sc.), University of Manitoba, Winnipeg, Manitoba, Canada, 1974. 119. pp.

Pavlak:1977:AFE

[PP77] Raymond Pavlak, Jr. and Adir Pridor. Assembler in a Fortran environment with a new debugging aid. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(4):63–69, December 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prager:1965:IBF

[Pra65] William Prager. Introduction to basic FORTRAN programming and numerical methods. A Blaisdell book in the pure and applied sciences. Blaisdell Pub. Co., Waltham, MA, USA, 1965. ix + 203 pp. LCCN QA76.5 .P7.

Pratt:1975:PLD

[Pra75] T. W. Pratt. Programming Languages: Design and Implementation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1975. ISBN 0-13-730432-3. xiv + 530 pp. LCCN QA76.7 .P71.

Preston:1970:FIP

[Pre70] D. A. Preston. FORTRAN IV program for sample normality tests. Computer contribution 41, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1970. 28 pp.

Presser:1979:FT

[Pre79] Leon Presser. Fortran tools. In ACM [ACM79], pages 61–62.

Price:1969:EBF

[Pri69] Wilson T. Price. Elements of basic Fortran IV programming, as implemented on the IBM 1130/1800 computers. Holt, Rinehart, and Winston, New York, NY, USA, 1969. ISBN 0-03-076560-9. xi + 387 pp. LCCN QA76.5 .P72. Published in 1975 under title: Elements of Fortran IV programming. “Selected references”: p. 2.

Price:1975:EFI

[Pri75] Wilson T. Price. Elements of Fortran IV programming. Rinehart Press, San Francisco, CA, USA, second edition, 1975. ISBN 0-03-089502-2. x + 301 pp. LCCN QA76.73.F25P74 1975.

Prime:1977:FPG

[Pri77a] Prime Computer, Inc. The FOR-TRAN programmer’s guide. Prime Computer, Inc, Framingham, MA, USA, 1977. 319 pp.

Prime:1977:PPD

[Pri77b] Prime Computer, Inc. Prime preliminary documentation release: PDR3057 FORTRAN programmer’s guide. Prime Computer, Inc., Framingham, MA, USA, 1977. ca. 350 pp.

Perrott:1980:SEU

[PRO80] R. H. Perrott, A. K. Raja, and P. C. O’Kane. Simulation experiment using two languages. The Computer Journal, 23(2):142–146, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Peterkin:1974:FBA

[PS74] Craig R. Peterkin and David H. Scuse. FORTRAN with business applications. WATFAC series in computer science and computer applications Waterloo Foundation for the Advancement of Computing. WATFAC series in computer science and computer applications. Waterloo Foundation for the Advancement of Computing, Waterloo, Ontario, Canada, 1974. ISBN 0-919884-01-6. xv + 332 pp. LCCN HF 5679 P646c.

Pomponiu:1978:FAS

[PS78] C. Pomponiu and M. Sararu. Fourier analysis with splines, a FORTRAN program. Computer Physics Communications, 16(1): 93–112, December 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465578901133.

Pienaar:1967:PMD

[PT67] L. V. Pienaar and John Arthur Collingwood Thomson. A program for multiple discriminant analysis (FORTRAN 1130). Technical report — Fisheries Research Board of Canada 40, Biological Station, Nanaimo, BC, Canada, 1967. 32 pp.

Pienaar:1968:EYP

[PT68] L. V. Pienaar and John Arthur Collingwood Thomson. Equilibrium yield per recruitment: a 1130 Fortran version of program no. FRG 708.

Technical report — Fisheries Research Board of Canada 92, Biological Station, Nanaimo, BC, Canada, 1968. 22 pp.

Pienaar:1969:PMD

[PT69] L. V. Pienaar and John Arthur Collingwood Thomson. A program for multiple discriminant analysis (FORTRAN 1130) extended version. Technical report — Fisheries Research Board of Canada 112, Biological Station, Nanaimo, BC, Canada, 1969. 26 pp.

Pienaar:1973:TPU

[PT73] L. V. Pienaar and John Arthur Collingwood Thomson. Three programs used in population dynamics: WVONB — ALOMA — BHYLD (FORTRAN 1130). Fisheries Research Board of Canada. Technical report 367, Pacific Biological Station, Nanaimo, BC, Canada, 1973. 33 pp.

Parker:1977:CDG

[PTM77] A. (Andrew) Parker, Roger Till, and Margaret Moss. The computation and display of geotechnical properties for sediment cores using BASIC and FOR-TRAN. Computer-based packages for teaching earth science. report no.4; Reading University geological reports no. 10, Geology Dept., University of Reading, Reading, UK, 1977. ISBN 0-7049-0280-X. 63 pp.

Pullin:1964:FAT

[Pul64] D. Pullin. A FORTRAN to ALGOL translator. The Computer Journal, 7(1):24–27, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/070024.sgm.
abs.html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/24.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
07/Issue_01/tiff/25.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/
26.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/27.tif.

Payne:1974:FPC

[PV74] G. L. Payne and P. L. Von Behren. Fortran program to calculate finite-range no-recoil DWBA transfer cross sections. Computer Physics Communications, 7 (1):13–37, January 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/001046557490054X.

Pyle:1962:CMF

[Pyl62] I. C. Pyle. Character manipulation in FORTRAN. Comm. ACM, 5(8):432–433, August 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pyle:1963:DF

[Pyl63] I. C. Pyle. Dialects of FORTRAN. Comm. ACM, 6(8):462–467, August 1963. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

QUBCC:1971:IF

[Que71] Queen’s University of Belfast.

Computer Centre, Belfast (University Rd), Northern Ireland. Introduction to FORTRAN, 1971. [3] + 38 pp.

QUBCC:1977:FE

[Que77] Queen’s University of Belfast.

Computer Centre, Belfast, Northern Ireland. FORTRAN by example, 1977. 112 pp.

Rabinowitz:1962:RAL

[Rab62] Irving N. Rabinowitz. Report on the algorithmic language FOR-TRAN II. Comm. ACM, 5 (6):327–337, June 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

RCA:1970:IIF

[Rad70] Radio Corporation of America, Camden, NJ, USA. IFOR (Interactive FORTRAN) terminal user aid, 1970. 30 pp.

Radford:1975:CPF

[Rad75] Arthur S. Radford. Computer programming, Fortran. Computer science studies Teach yourself books. Teach Yourself Books, London, UK, 1975. ISBN 0-340-19495-2 (paperback). 253 pp.

Radford:1976:CPF

[Rad76a] Arthur S. Radford. Computer programming/Fortran. Teach yourself books. Hodder and Stoughton, London, UK, 1976. ISBN 0-340-19495-2. 253 pp.

Radue:1976:TEF

[Rad76b] J. E. Radue. On the teaching and evaluation of a Fortran service course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(2):32– 35, June 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

RadioShack:1979:FTM

[Rad79] Radio Shack. FORTRAN [TRS-80 model I], 1979.

Radford:1980:CPF

[Rad80] Arthur S. Radford. Computer Programming FORTRAN. Fodors, seventh edition, July 1980. ISBN 0-679-10378-3. ???? pp. LCCN ???? US$8.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0679103783.

RadioShack:1980:FTM

[Rad88] Radio Shack. FORTRAN TRS-80 model III, 1980 (or 1988??).

Raffenetti:1979:CSF

[Raf79] Richard C. Raffenetti. A comparative study of the Fortran development environment provided by the VAX/VMS and VAX/UNIX operating systems. Technical memorandum 346, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1979. iii + 29 pp. Reproduced by the National Technical Information Service.

Rajaraman:1977:CPF

[Raj77] V. Rajaraman. Computer programming in Fortran IV. Prentice-Hall of India, New Delhi, India, 1977. 235 pp.

Ralston:1971:FFC

[Ral71a] Anthony Ralston. Fortran and the first course in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(4):24–29, December 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ralston:1971:FIP

[Ral71b] Anthony. Ralston. Fortran IV Programming: a Concise Exposition. McGraw-Hill, New York, NY, USA, January 1971. ISBN 0-07-051164-0. xi + 177 pp. LCCN ???? US$17.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0070511640.

Ranelletti:1965:DFS

[Ran65] John E. Ranelletti. Dynamic format specifications. Comm. ACM, 8(8):508–510, August 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rapp:1966:FPCa

[Rap66a] Richard H. Rapp. A FORTRAN program for the computation of the disturbance components of gravity. Scientific report no. 10; report no. 76, Ohio State University Research Foundation, Columbus, OH, USA, 1966. v + 61 pp.

Rapp:1966:FPCb

[Rap66b] Richard H. Rapp. A FORTRAN program for the computation of the normal gravity and gravitational field of the earth. Scientific report no. 3; report no. 52, Ohio State University Research Foundation, Columbus, OH, USA, 1966. 35 pp.

Rapp:1966:FPP

[Rap66c] Richard H. Rapp. A FORTRAN program for the prediction of mean anomalies. Scientific report no. 4; report no. 63, Ohio State University Research Foundation, Columbus, OH, USA, 1966. 49 pp.

Rapp:1966:FPU

[Rap66d] Richard H. Rapp. A FORTRAN program for the upward continuation of gravity anomalies. Scientific report no. 5; report no. 62, Ohio State University Research Foundation, Columbus, OH, USA, 1966. 68 pp.

Ratzer:1972:FC

[Rat72] Gerald F. G. Ratzer. A FOR-TRAN course. Holt, Rinehart and Winston of Canada; distributed in the U.S. by Winston Press [Minneapolis, MN], Toronto, Ontario, Canada, 1972. ISBN 0-03-928115-9. ix + 211 pp. LCCN QA76.73.F25 R37. Distributed in the U.S. by Winston Press, Minneapolis, MN, USA.

Raun:1968:IFC

[Rau68] Donald L. Raun. An introduction to Fortran computer programming for business analysis. Dickenson series in computer and information science. Dickenson Pub. Co., Encino, CA, USA, 1968. xvi + 318 pp.

Rauhauser:1978:LPG

[Rau78] Russ C. Rauhauser. An LALR(1)

parser grammar for FORTRAN. Technical Report Contract no. AD/A-056 839CU-CS-129-78, U.S. Army Research Office, Durham, NC, USA, 1978. iv + 55 pp. Reproduced by National Technical Information Service, U.S. Dept. of Commerce.

Rawlinson:1977:SPM

[Raw77] Jerry Dean Rawlinson. A sequential PASCAL manual for FOR-TRAN programmers. Thesis (Master’s), Kansas State University, Manhattan, KS, USA, 1977. 143 pp.

Raynal:1963:PFP

[Ray63] J. Raynal. Programme Fortran pour la diffusion elastique de deutons avec un modele optique contenant des termes tensoriels. Commissariat a l’énergie atomique. Rapport 2287, Centre d’études nucléaires de Saclay, Gif sur Yvette, France, 1963. 162 pp.

Rudman:1976:FPCa

[RB76a] Albert J. Rudman and Robert F.

Blakely. Fortran program for correlation of stratigraphic time series. Occasional paper — Indiana Geological Survey 14; geophysical computer programs 3, Dept. of Natural Resources, Bloomington, IN, USA, 1976. 31 pp.

Rudman:1976:FPGa

[RB76b] Albert J. Rudman and Robert F. Blakely. Fortran program for generation of synthetic seismograms. Occasional paper — Indiana Geological Survey 13; geophysical computer programs 2, Dept. of Natural Resources, Bloomington, IN, USA, 1976. 27 pp.

Rudman:1976:FPGb

[RB76c] Albert J. Rudman and Robert F. Blakely. FORTRAN program for generation of synthetic seismograms. Geophysical computer program 2; geological survey occasional paper 13, Printed for authority of the State of Indiana, Bloomington, IN, USA, 1976. 27 pp.

Rudman:1976:FPCb

[RBK76] Albert J. Rudman, Robert F. Blakely, and Byung Doo Kwon. FORTRAN program for correlation of stratigraphic time series. Geophysical computer program; 3, 6 geological survey occasional paper; 14, 26, State of Indiana, Dept. of Natural Resources, Geological Survey, Bloomington, IN, USA, 1976. ???? pp.

Rudman:1975:FPU

[RBp75] Albert J. Rudman, Robert F. Blakely, and Fortran program. Fortran program for the upward and downward continuation and derivatives of potential fields. Geophysical computer program 1; geological survey occasional paper 10, State of Indiana, Dept. of Natural Resources, Geological Survey, Bloomington, IN, USA, 1975. 23 pp.

Roberts:1975:ARM

[RCL75] K. V. Roberts, J. P. Christiansen, and J. W. Long. Adiabatic relaxation to 1D MHD pressure equilibrium. EQUIL: a Fortran module and test program. Computer Physics Communications, 10(5):264–281, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465575900971.

Ryan:1966:CSI

[RCM66] J. L. Ryan, R. L. Crandall, and M. C. Medwedeff. A conversational system for incremental compilation and execution & in a timesharing environment. fjcc, pages 1– 21, 1966.

Rohl:1975:CBS

[REC75] J. S. Rohl, H. D. Ellison, and R. J.

Collins. An in-core batching standard FORTRAN compiler for large ICL 1900 machine. Software— Practice and Experience, 5(1):19– 28, January/March 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reed:1968:MEF

[Ree68] Alan Charles Reed. A method for extending FORTRAN V for the interactive graphical solution of numerical problems. Technical report 4-10, Computer Science [Dept.], Information Processing Systems University of Utah, Salt Lake City, UT, USA, 1968. vi + [25] pp.

Reed:1971:MEF

[Ree71] Alan Charles Reed. A method for extending FORTRAN V for the interactive graphical solution of numerical problems. Thesis (m.s.), Dept. of Electrical Engineering, Computer Science, University of Utah, Salt Lake City, UT, USA, 1971. vi + 53[1] pp.

Reese:1972:NMF

[Ree72] Oliver W. Reese. Numerical method and FORTRAN program for the solution of an axisymmetric electrostatic collector design problem. NASA technical note D-6959, National Aeronautics and Space Administration, Washington, DC, USA, 1972. iii + 65 pp. For sale by the National Technical Information Service.

Reeves:1973:BRB

[Ree73] C. M. Reeves. Book review: Numerical methods with FORTRAN IV case studies, W. S. Dorn and D. D. McCracken, Wiley, London, 1972. No. of pages: 447. Price: £6.00. Software—Practice and Experience, 3(2):190–191, April/ June 1973. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reeves:1975:BRB

[Ree75] C. M. Reeves. Book review: Digital Computing and Numerical Methods (With Fortran IV, Watfor, and Watfiv Programming) Brice Carnahan and James O.

Wilkes, John Wiley, 1973. No. of pages: 477. Price: £7.00. Software—Practice and Experience, 5 (1):111, January 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reeves:1976:BRB

[Ree76] C. M. Reeves. Book review: Computer Science: Programming in FORTRAN with WATFOR WAT-FIV. The Computer Journal, 19 (2):172a–172, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/
content/19/2/172-a.full.pdf+
html.

Reeves:1979:APS

[Ree79] A. P. Reeves. An array processing system with a Fortran-based realization. Computer Graphics and Image Processing, 9(3):267– 281, 1979. CODEN CGIPBG. ISSN 0734-189X.

Reid:1972:TFS

[Rei72a] John Ker Reid. Two FORTRAN subroutines for direct solution of linear equations whose matrix is sparse, symmetric and positivedefinite. Theoretical Physics Division, Atomic Energy Research Establishment, Harwell, Berkshire, UK, June 1972. ISBN 0-7058-0092-X. 32. pp. AERE — R 7119, HL72/2370 (C13).

Reid:1972:GFP

[Rei72b] Thomas A. Reid. A guide to FOR-TRAN programming and UNI-WAFT. Computer Center, University of Western Australia, Nedlands, W.A. 6009, Australia, second edition, 1972. ISBN 0-9599452-0-2. 216 pp. LCCN QA76.73.F25R45 1972.

Reifer:1976:SFD

[Rei76] Donald J. Reifer. The structured Fortran dilemma. ACM SIGPLAN Notices, 11(2):30–32, February 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:1979:FMF

[Rei79] John Reid. Functions for manipulating floating-point numbers. ACM SIGNUM Newsletter, 14(4): 11–13, December 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Reid:1980:CDP

[Rei80] John Reid. Complex double precision in association with Fortran 77. ACM SIGNUM Newsletter, 15(1):16–17, March 1980. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rejchrt:1972:SFG

[Rej72] Vladimir J. Rejchrt. Signal flow graph and a Fortran program for Haar–Fourier transform. IEEE Transactions on Computers, C-21(9):1026–1027, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009086.

Rens:1965:FPC

[Ren65] Frank J. Rens. A FORTRAN program for coordinate mapping using IBM 7090 computer. United States. Office of Naval Research. Geography Branch. Technical report 10, Northwestern University, Evanston, IL, USA, 1965. 1 v pp.

Reynolds:1968:TWT

[Rey68] William K. Reynolds. Truckweight table W-3 for IBM 1620, model II. New Mexico State Highway Dept., Planning Research Section Special Studies Unit, Santa Fe, NM, USA, 1968. various pp.

Reyment:1969:FIP

[Rey69] R. A. Reyment. FORTRAN IV program for the generalized statistical distance and analysis of covariance matrices for the CDC 3600 computer. Computer contribution 39, Kansas Geological Survey, Lawrence, KS, USA, 1969. 42 pp.

Reynolds:1977:CSH

[Rey77] R. A. Reynolds. Character string handling in Fortran. The Computer Journal, 20(4):325–329, November 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Rule:1973:FIP

[RFP73] Wilfred P. Rule, Robert G. Finkenaur, and Farrell G. Patrick. Fortran IV programming. Prindle, Weber and Schmidt, Boston, MA, USA, 1973. vii + 504 pp.

Raubenheimer:1968:GFP

[RG68] L. J. Raubenheimer and G. Gilat. OR GHX a Fortran program for calculating frequency distribution functions of hexagonal closepacked crystals. PEL 178, Atomic Energy Board, Pelindaba, Pretoria, South Africa, 1968. 27 pp.

Rink:1976:ASC

[RG76] R. A. Rink and B. P. Guru. Analytical solutions for a class of nonlinear differential equations using Altran. BIT (Nordisk tidskrift for informationsbehandling), 16(2): 161–171, June 1976. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.
com/openurl.asp?genre=article&
issn=0006-3835&volume=16&issue=
2&spage=161.

Rigby:1977:DFI

[RG77] D. Rigby and A. M. Giblin. DIS-PLOT — a Fortran IV subroutine to plot three-dimensional data. Commonwealth Scientific and Industrial Research Organization. Minerals Research Laboratories. Technical communication 62, CSIRO Division of Mineralogy, North Ryde, NSW, Australia, 1977. ISBN 0-643-02068-3. 8 + [15] pp.

Rose:1976:PCP

[RH76] L. L. Rose and H. Hellerman. Portable character processing in Fortran and fixed integer environments. IEEE Transactions on Software Engineering, SE-2(3):176–185, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702363.

Robillard:1978:EPS

[RH78] Pierre N. Robillard and R. C.

(Richard C.) Holt. Exercices de programmation structurée. Technical Report EP-78-E-23, École polytechnique de Montréal, Montréal, PQ, Canada, 1978. v + 93 pp.

Riccat:1973:SFP

[Ric73] Alain Riccat. Un sous-programme FORTRAN pour la résolution des équations de Barraud Riccati discrètes. (French). Technical report, Ann. École Nat. Sup. Méc., 1973. 57–83 pp. Deuxième semestre.

Ridolfi:1967:ADF

[Rid67] Pierluigi Ridolfi. Applicazioni del Fortran: nel calcolo numerico, nella ricerca operativa e nella statistica. Number 5 in Collana di matematica e statistica applicate ai problemi aziendali. F. Angeli, Milano, Italy, 1967. 176 pp.

Ridolfi:1968:FTE

[Rid68] Pierluigi Ridolfi. Il Fortran: Teoria ed esercizi. Number 6 in Collana di matematica e statistica applicate ai problemi aziendali. F. Angeli, Milano, Italy, 1968. 167 pp.

Ridolfi:1969:FTA

[Rid69] Pierluigi Ridolfi. El Fortran: teoria y aplicaciones. La empresa moderna. Iberico Europea de Ediciones, S.A., Madrid, Spain, 1969. 197 pp.

Ridolfi:1978:ADF

[Rid78] Pierluigi Ridolfi. Applicazioni del Fortran nel calcolo numerico, nella ricerca operativa e nella statistica. Number 21 in Collana di informatica. F. Angeli, Milano, Italy, 6. edition, 1978. 176 pp.

Ridler:1979:FRM

[Rid79] Philip Ridler. A FORTRAN reference manual. Pitman Publishing Ltd., London, UK, 1979. 120 pp.

Rindfleisch:1977:DFO

[Rin77] Daniel H. Rindfleisch. DUMP-Interpretationen bei Fehlern in OS- und OS/VS-Programmen der Systeme IBM/360 und IBM/370: ein methodischer Leitfaden Zur Lösung von Datei- und Programmproblemen am Beispiel von COBOL-, FORTRAN-, PL/1und ASSEMBLER-Programmen. Hanser, München, Germany, 1977. ISBN 3-446-12372-5. xii + 273 pp.

Ring:1979:FPN

[Rin79] Martin D. Ring. FORTRAN programs for numerical and graphical analysis of elastic response, radiated sound pressure, and reflected sound pressure for plane plate structures with fluid on one side. NRL memorandum report 3885, Naval Research Laboratory, Washington, DC, USA, 1979. iv + 76 pp.

Ripley:1977:PPA

[Rip77] G. D. Ripley. Program perspectives: a relational representation of measurement data. IEEE Transactions on Software Engineering, SE-3(4):296–300, July 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Ritchie:1968:LSF

[Rit68] Earl Dean Ritchie. LPI: a system for Fortran list processing. Thesis (m.s.), Washington State University, Pullman, WA, USA, 1968. v + 37 pp.

Reuter:1978:SEU

[RJAS78] Eric K. Reuter, John P. Jeter, J. Wayne Anderson, and Bruce D. Shriver. Some experiments using interval arithmetic. In IEEE SCA ’78 [IEE78], pages 75–80. ISSN 1063-6889. URL http://www.

acsel-lab.com/arithmetic/arith4/
papers/ARITH4_Reuter.pdf. IEEE

catalog no. 78CH1412-6C.

Richardson:1973:AAC

[RK73] Joel A. Richardson and J. L.

Kuester. ACM Algorithm 454: The complex method for constrained optimization [E4]. Comm. ACM, 16(8):487–489, August 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?].

Ropes:1969:FIP

[RMM69] Leverett H. Ropes, Charles O.

Morgan, and Jesse M. McNellis.

FORTRAN IV program for synthesis and plotting of water-quality data. Special distribution publication 39, State Geological Survey; The University of Kansas, Lawrence, KS, USA, 1969. iii + 59 pp.

Robbins:1962:FBD

[Rob62] D. K. Robbins. FORTRAN for business data processing. Comm. ACM, 5(7):412–414, July 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robinson:1967:MTS

[Rob67a] Enders A. Robinson. Multichannel Time Series Analysis with Digital Computer Programs. Holden-Day, San Francisco, CA, USA, 1967. xxiii + 298 pp. LCCN QA276 .R58.

Robinson:1967:MFP

[Rob67b] George A. Robinson. MACRO-FROTRAN, a facility for programmerdefined macro-instructions in FORTRAN programs. Technical report, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1967. 39 pp. Available from Clearinghouse for Federal Scientific and Technical Information, Springfield, VA, USA.

Robinson:1968:FFC

[Rob68] G. S. Robinson. Forex; a Fortran compilation subroutine for the IBM 360. Technical Report AAEC/E190, Research Establishment, Australian Atomic Energy Commission, Lucas Heights, NSW, Australia, 1968. 16 + 10 pp.

Roberts:1969:PSF

[Rob69] K. V. Roberts. The publication of scientific Fortran programs. Computer Physics Communications, 1(1):1–9, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465569900113.

Roberts:1979:FPS

[Rob79] F. D. K. Roberts. A Fortran program solution of the constrained l [infinity] linear approximation problem. Internal report 132, University of Victoria, Dept. of Mathematics, Victoria, BC, Canada, 1979. [11]. pp.

Rochkind:1970:TFS

[Roc70] M. M. Rochkind. Transferable FORTRAN subroutine for rapid extended sorting. IEEE Transactions on Computers, C-19(3):270– 272, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

RodriguezL:1976:CRA

[Rod76] Oscar Rodriguez L. Comparison of the Remez algorithm written in the FORTRAN and the PL/I languages. Thesis (m.s.), Ball State University, Muncie, IN, USA, 1976. vi + 100 pp.

Rogers:1980:DMA

[Rog80] G. Rogers. Dynamic 3D modeling for architectural design. Computer Aided Design, 12, 1:13–20, 1980. CODEN CAIDA5. ISSN 0010-4485.

Rohl:1973:PFC

[Roh73] J. S. (Jeffrey Soden) Rohl. Programming in Fortran: a course of ten television lectures. Manchester University Press, Manchester, UK, 1973. ISBN 0-7190-0555-8. 127 pp. LCCN QA76.73.F25R63.

SalazarRomero:1975:FTP

[Rom75] Jose Humberto Salazar Romero. Fortran, teoria y practica. Editorial Trillas, Mexico, DF, Mexico, preedicion edition, 1975. 141 pp.

Rosen:1966:PSL

[Ros66] Saul Rosen, editor. Programming Systems and Languages. McGraw-Hill computer science series. Mc-Graw-Hill, New York, NY, USA, 1966. xv + 734 pp. LCCN QA76.5 .R53. Reissued in 1967. Contains a reprint of the first published paper on Fortran [BBB^{+}57] on pp. 29– 47.

Rosin:1971:FNC

[Ros71] Robert F. Rosin. Fortran and the new culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 3(1):10–11, March 1971. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosen:1972:PSL

[Ros72] Saul Rosen. Programming systems and languages (1965–1975). Comm. ACM, 15(7):591–600, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roscoe:1973:FPF

[Ros73] John T. Roscoe. The Funstat package in Fortran IV. Holt, Rinehart, and Winston, New York, NY, USA, 1973. ISBN 0-03-002016-6. vi + 135 pp. LCCN HA29 .R785.

Rosin:1978:ASH

[Ros78] Robert F. Rosin. ACM SIGPLAN history of programming languages conference PL/I language summary. ACM SIGPLAN Notices, 13(8):225–226, August 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rotta:1971:FIG

[Rot71] Julius C. Rotta. FORTRAN IV-Rechenprogramm fur grenzschichten bei kompressiblen ebenen und achsensymmetrischen stromungen. Deutsche Luft-und Raumfahrt. Forschungsbericht 71-51, Deutsche Forschungs-und Versuchsanstalt fur Luft-und Raumfahrt, Porz-Wahn, 1971. 82 pp. Summary in English. AVA-FB 7113.

Roussel:1975:PMR

[Rou75] P. Roussel. Prolog: Manuel de reference et utilisation. Tr, Groupe d’Intelligence Artificielle Universit’e d’Aix-Marseille II, Marseille, France, 1975.

Rowland:1976:SCS

[Row76] Stuart W. Rowland. Some comments on structured FORTRAN.

ACM SIGPLAN Notices, 11(10): 43–48, October 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Randers-Pehrson:1974:IGF

[RP74] Glenn Randers-Pehrson. Interactive graphics FORTRAN subroutines for editing and manipulating data files within an application program. Technical report, Feltman Research Laboratory, Picatinny Arsenal, Dover, NJ, USA, 1974. ii + 15 pp. Distributed by National Technical Information Service.

Restrepo-Posada:1979:FPI

[RPE79] Pedro Juan Restrepo-Posada and Peter S. Eagleson. Fortran programs for the interpretation and analysis of NOAA hourly precipitation data tapes. Technical note no. 22; technical report t79-2., Massachusetts Institute of Technology, Ralph M. Parsons Laboratory for Water Resources and Hydrodynamics, Cambridge, MA, USA, 1979. 44 + [1] pp.

Reyment:1970:FIP

[RR70] R. A. Reyment and Hans-Ake Ramden. FORTRAN IV program canonical variates analysis for CDC 3600 computer. Computer contribution 47, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1970. 40 pp.

Rice:1973:ICF

[RR73a] John K. Rice and John Rischard Rice. Introduction to computing with Fortran. Holt, Rinehart, and Winston, New York, NY, USA, 1973. ISBN 0-03-086217-5. ix + 384 pp. LCCN QA76.73.F25 R51.

Robers:1973:AAD

[RR73b] P. D. Robers and S. S. Robers.

ACM Algorithm 458: Discrete linear L_{1} approximation by interval linear programming [E2]. Comm. ACM, 16(10):629–631, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenblatt:1973:SFPa

[RR73c] Lisa Rosenblatt and Judah I. (Judah Isser) Rosenblatt. Simplified FORTRAN programming: with companion problems. Addison-Wesley, Reading, MA, USA, 1973. various pp.

Rosenblatt:1973:SFPb

[RR73d] Lisa Rosenblatt and Judah I. (Judah Isser) Rosenblatt. Simplified Fortran programming: with companion problems. Addison-Wesley, Reading, MA, USA, 1973. ISBN 0-201-06511-8 (paperback). 336 pp.

Reger:1969:SPF

[RS69] J. Reger and M. Schmöller. Simulationsprogramme in der Programmiersprache Fortran zur Bestimmung der Leistungsfähigkeit von Koppelanordnungen. Informationen Fernsprech-Vermittlungstechnik, 5, 1, 1969.

Robinson:1972:CFS

[RS72] Edward L. Robinson, Jr. and Henry A. Scarton. CONTOR: a FORTRAN subroutine to plot smooth contours of a singlevalued arbitrary three-dimensional surface. Journal of Computational Physics, 10(2):242–251, October 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999172900642.

Robinson:1980:IAS

[RS80] Sally S. Robinson and M. L. Soffa. An instructional aid for student programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 12(1): 118–129, February 1980. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 11th SIGCSE Symposium on Computer Science Education.

Ross:1969:DFC

[RSBR69] Tony Alan Ross, Garry D. Smith, Roscoe Allen Boyer, and Richard Douglas Ross. Development of a Fortran computer program to design school bus routes. Technical report, University of Mississippi, Oxford, MS, USA, 1969. various pp.

Rosen:1965:PPU

[RSD65] Saul Rosen, Robert A. Spurgeon, and Joel K. Donnelly. PUFFT

— the Purdue University fast FORTRAN translator. Comm. ACM, 8(11):661–666, November 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubin:1978:SAA

[RST78] Donald B. Rubin, Thomas W. F.

Stroud, and Dorothy Thayer. Statistical algorithms: Algorithm AS 120: a Fortran algorithm for the additive model in a two-way unbalanced manova. Applied Statistics, 27(1):92–97, March 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://lib.stat.cmu.edu/
apstat/120.

Robinson:1976:EAF

[RT76] S. K. Robinson and I. S. Torsun.

Empirical analysis of Fortran programs. The Computer Journal, 19 (1):56–62, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Robinson:1977:DAP

[RT77] S. K. Robinson and I. S. Torsun. Dynamic analysis of program performance (DAP) in a Fortran batch environment. Software —Practice and Experience, 7(3): 307–315, May/June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rubenstein:1969:FCPb

[Rub69a] Nanch Ordelheide Rubenstein. A Fortran computer program for transcribing Franconian rhythm. Thesis (Ph.D.), Washington University, St. Louis, MO, USA, 1969. 134 pp.

Rubenstein:1969:FCPa

[Rub69b] Nancy Ordelheide Rubenstein. A FORTRAN computer program for transcribing Franconian rhythm. Rubenstein, Saint Louis, MO, USA, 1969. v + 135 pp.

Rubenstein:1969:FCPc

[Rub69c] Nancy Ordelheide Rubenstein. A Fortran computer program for transcribing Franconian rhythm. Thesis (Ph.D.), Washington University, St. Louis, MO, USA, 1969. v + 134 pp.

Rule:1966:FIP

[Rul66a] Wilfred P. Rule. Fortran IV programming. Prindle, Weber and Schmidt, Boston, MA, USA, 1966. vii + 221 pp.

Rule:1966:IFP

[Rul66b] Wilfred P. Rule. Introduction to Fortran programming. Prindle, Weber and Schmidt, Boston, MA, USA, 1966. vii + 167 pp.

Rule:1968:FIP

[Rul68a] Wilfred P. Rule. Fortran IV programming. Prindle, Weber and Schmidt, Boston, MA, USA, 1968. vii + 221 pp. LCCN QA76.5 .R8.

Rully:1968:IGD

[Rul68b] A. D. Rully. Interactive graphics in data processing: a subroutine package for FORTRAN. IBM Systems Journal, 7(3/4):248–256, 1968. CODEN IBMSA7. ISSN 0018-8670.

Rule:1980:FPA

[Rul80] Wilfred P. Rule. Fortran, a Practical Approach With Style and Structure. Prindle, Weber and Schmidt, Boston, MA, USA, third edition, June 1980. ISBN 0-87150-290-9. ix + 534 pp. LCCN QA76.73.F25 R84 Sci-Eng. US$37.80. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0871502909.

Russell:1978:CCS

[Rus78] Richard M. Russell. The Cray-1 computer system. Comm. ACM, 21(1):63–72, January 1978. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Russ:1979:ESP

[Rus79] O. W. Russ. ECHO, a simple procedure to echo input data from any existing FORTRAN code. Technical Report K/CSD/TM-27, U.S. Dept. of Energy, Washington, DC, USA, January 1979. 23 pp.

Rao:1978:NFP

[RV78] K. Srinivasa Rao and K. Venkatesh.

New Fortran programs for angular momentum coefficients. Computer Physics Communications, 15(3–4):227–235, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465578900930.

Ralston:1969:MMD

[RW69] Anthony Ralston and H. S. Wilf.

Mathematische Methoden für Digitalrechner. R. Oldenbourg, München, Germany, 1969.

Ralston:1976:SFE

[RW76] Anthony Ralston and Jerrold L.

Wagener. Structured Fortran: an evolution of standard Fortran. Technical report 107, State University of New York at Buffalo, Dept. of Computer Science, Amherst, NY, USA, 1976. 86 pp.

Ripley:1977:SAM

[RW77] G. David Ripley and James Wm White. A survey and analysis of minicomputer Fortran dialects. Technical report, Dept. of Computer Science, University of Arizona, Tucson, AZ, USA, 1977. iii + 55 pp.

Ryder:1972:FVM

[Ryd72] B. G. Ryder. The FORTRAN Verifier: Motivation and implementation. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1972. ?? pp.

Ryder:1974:PV

[Ryd74] Barbara G. Ryder. The PFORT verifier. Software—Practice and Experience, 4(4):359–377, October–December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryder:1973:FVU

[Ryd3 ] B. G. Ryder. The FORTRAN Verifier: User’s guide. Computing Science Technical Report 12, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 1973 (??). ?? pp.

Rudman:1977:FPG

[RZB77] Albert J. Rudman, Robert Ziegler, and Robert F. Blakely. Fortran program for generation of earth tide gravity values. Geophysical computer program 4; geological survey occasional paper 22, State of Indiana, Dept. of Natural Resources, Geological Survey, Bloomington, IN, USA, 1977. 14 pp.

Schneck:1973:FFO

[SA73] P. B. Schneck and Ellinor Angel. A FORTRAN to FOR-TRAN optimising compiler. The Computer Journal, 16(4):322– 330, November 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
16/Issue_04/160322.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/322.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/323.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/
324.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_04/tiff/325.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
16/Issue_04/tiff/326.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
04/tiff/327.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/
328.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_04/tiff/329.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
16/Issue_04/tiff/330.tif.

Sandell:1974:MCT

[SA74] Nils Richard Sandell and Michael Athans. Modern control theory: a self-study subject: manual of Fortran computer subroutines for linear, quadratic, Gaussian designs. Center for Advanced Engineering Study, Massachusetts Institute of Technology, Cambridge, MA, USA, 1974. 193 pp.

Sabin:1976:PSE

[Sab76] M. A. Sabin. Portability-some experiences with FORTRAN. Software—Practice and Experience, 6(3):393–396, July/September 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sadowsky:1972:MOS

[Sad72] George Sadowsky. MASH, an online system for socio-economic microsimulation of the U.S. household sector. Technical report, Urban Institute, Washington, DC, USA, 1972.

Sakoda:1964:DMDa

[Sak64a] James Minoru Sakoda. DYSTAL manual: dynamic storage allocation language in FORTRAN. Technical report, Sociology Computer Laboratory, Dept. of Sociology and Anthropology, Brown University, Providence, RI, USA, 1964. ix + 160 + [7] + [64] pp.

Sakoda:1964:DMDb

[Sak64b] James Minoru Sakoda. DYS-TAL manual; dynamic storage allocation language in FORTRAN.

Technical report, The Laboratory, Providence, RI, USA, 1964. ix + 194 + 68 pp.

Sakoda:1965:DMD

[Sak65] James Minoru Sakoda. DYSTAL manual: dynamic storage allocation language in FORTRAN. Technical report, Brown University, Dept. of Sociology and Anthropology, Providence, RI, USA, 1965. ix + 308 pp.

Sakoda:1970:DMD

[Sak70] James Minoru Sakoda. DYS-TAL manual: dynamic storage allocation language in FORTRAN. Technical report, Sociology Computer Laboratory, Brown University, Providence, RI, USA, 1970. 223 pp.

Sakoda:1979:DGP

[Sak79] James M. Sakoda. DYSTAL 2: a general purpose extension of FOR-TRAN. ACM SIGPLAN Notices, 14(1):77–90, January 1979. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salina:1970:SFI

[Sal70] E. Salina. SQUIRREL; a FOR-TRAN IV one-dimensional fewgroup diffusion-depletion code which includes the effects of local power and water density. Technical report, Commission of the European Communities, Luxembourg, Luxembourg, 1970. 62 + 70 pp.

Sale:1971:AIC

[Sal71a] A. H. J. Sale. Algorithm 65: An improved clustering algorithm. The Computer Journal, 14(1): 104–106, February 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://www3.
oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
104.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/105.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_01/tiff/106.tif.

Sale:1971:CFS

[Sal71b] A. H. J. Sale. The classification of FORTRAN statements. The Computer Journal, 14(1): 10–12, February 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
14/Issue_01/140010.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/10.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/
11.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/12.tif.

Salem:1976:CFF

[Sal76] M. Salem. Converter of FOR-TRAN format and data statements to standard form. Computer Physics Communications, 11(2):199–209, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465576900527.

Salomon:1977:DFF

[Sal77a] David Salomon. A design for Fortran to facilitate structured programming. ACM SIGPLAN Notices, 12(1):95–100, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salomon:1977:FIP

[Sal77b] K. B. Salomon. A Fortran IV program which determines that region of a polygon within a polygonal boundary. Computers and Geosciences, 4:53–63, 1977. CO-DEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Salomon:1978:FIP

[Sal78] K. B. Salomon. A Fortran IV program which determines that region of a polygon within a polygonal boundary. Computers and Geosciences, 4(1):53–63, 1978. CO-DEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Sammet:1969:PLH

[Sam69] Jean E. Sammet. Programming Languages: History and Fundamentals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1969. xxx + 785 pp. LCCN QA76.5 .S213.

Sanderson:1970:CL

[San70] Peter C. Sanderson. Computer Languages: a Practical Guide to the Chief Programming Languages. Newnes-Butterworths, London, UK, 1970. ISBN 0-408-04303-2. vii + 200 pp. LCCN QA76.5.S216 1970.

Sanderson:1973:ICB

[San73] Peter C. Sanderson. Interactive Computing in Basic: an introduction to interactive computing and a practical course in the BASIC language. Butterworths, London, UK, 1973. ISBN 0-408-70527-2. vii + 161 pp. LCCN QA76.73.B3 S261 1973.

Sander:1974:FIP

[San74] D. M. Sander. FORTRAN IV program to calculate air infiltration in buildings. DBR computer program 37, National Research Council of Canada, Division of Building Research, Ottawa, Ontario, Canada, 1974. 7 + 4 + 12 + 25 + [5] pp.

Sandwick:1978:FCP

[San78] Jane P. Sandwick. A Fortran compiler for the PDP-11/ 34 minicomputer. Thesis (m.s.), Ball State University, Muncie, IN, USA, 1978. iii + 107 pp.

Sassman:1969:FFC

[Sas69] Paul Anthony Sassman. A fast FORTRAN compiler for the RCA Spectra 70. Thesis (m.s.), Newark College of Engineering, Newark, NJ, USA, 1969. 34 pp.

Sass:1974:FFP

[Sas74a] C. Joseph Sass. Fortran Four Programming and Applications. Holden-Day, San Francisco, CA, USA, June 1974. ISBN 0-8162-7473-8. ???? pp. LCCN ???? US$17.95. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0-8162-
7473-8.

Sass:1974:FIP

[Sas74b] C. Joseph Sass. Fortran IV programming and applications. Holden-Day computer and information sciences series. Holden-Day, San Francisco, CA, USA, 1974. ISBN 0-8162-7473-8. x + 350 pp. LCCN QA76.73.F25 S271.

Sawyer:1962:IFC

[Saw62] Craig D. (Craig Delany) Sawyer.

An IBM-709 Fortran code for counterflow heat exchanger transients. Thesis (m.s.), Dept. of Nuclear Engineering, Massachusetts Institute of Technology, Cambridge, MA, USA, 1962. various pp.

Schwar:1978:BPI

[SB78] James P. Schwar and Charles L.

Best. BASIC programming with an introduction to FORTRAN. Lafayette College, Easton, PA, USA, 1978. 306 pp.

Schallert:1979:PF

[SC79] William F. Schallert and Carol Reedy Clark. Programming in Fortran. Addison-Wesley, Reading, MA, USA, May 1979. ISBN 0-201-06716-1. v + 442 pp. LCCN QA76.73.F25S33. US$30.25. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0201067161.

Scanlon:1970:ICS

[Sca70] Raymond D. Scanlon. An interpolating cubic spline FORTRAN subroutine. Technical Report AD 702 910, Watervliet Arsenal, Watervliet, NY, USA, 1970. ii + 25 pp. Distributed by Clearinghouse for Federal Scientific and Technical Information.

Scarton:1971:DPF

[Sca71] Henry A. Scarton. Double precision FORTRAN subroutines to compute both ordinary and modified Bessel functions of the first kind and of integer order with arbitrary complex argument: J_{n}(x + jy) and I_{n}(x + jy). Journal of Computational Physics, 8(2):295– 299, October 1971. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Schroeder:1962:NSI

[Sch62] Hans Paul Schroeder. Numberwriting subroutine for IBM 1620 Fortran and the CCP incremental plotter. Thesis (b.s.), Dept. of Civil Engineering, Massachusetts Institute of Technology, Cambridge, MA, USA, 1962. 16 + [8] pp.

SMSG:1966:ACMd

[Sch66a] School Mathematics Study Group. Algorithms, computation, and mathematics. Algorithms, computation, and mathematics (Fortran supplement) Student text. School Mathematics Study Group, Stanford, CA, USA, revised edition, 1966. 132 pp.

SMSG:1966:ACMe

[Sch66b] School Mathematics Study Group.

Algorithms, computation, and mathematics. Algorithms, computation, and mathematics (Fortran supplement) Teacher’s commentary. School Mathematics Study Group, Stanford, CA, USA, revised edition, 1966. 109 pp.

SMSG:1966:ACMa

[Sch66c] School Mathematics Study Group.

Algorithms, computation and mathematics Fortran supplement. School Mathematics Study Group, Stanford, CA, USA, revised edition, 1966. 2 volumes.

SMSG:1966:ACMb

[Sch66d] School Mathematics Study Group.

Algorithms, computation and mathematics: (Fortran supplement) student text. School Mathematics Study Group, Stanford, CA, USA, revised edition, 1966. 132 pp.

SMSG:1966:ACMc

[Sch66e] School Mathematics Study Group.

Algorithms, computation and mathematics: (Fortran supplement) teacher’s commentary. School Mathematics Study Group, Stanford, CA, USA, revised edition, 1966. 102 pp.

Schofield:1967:MAF

[Sch67] C. F. Schofield. A manual of the Atlas Fortran V language. Technical report, University of London Atlas Computing Service, London, UK, 1967. 200 pp.

Schonbeck:1968:FIM

[Sch68] Rudolph G. Schonbeck. Fortran IV; for multi-programming systems, with emphasis on the GE-600 series computer. Addison-Wesley, Reading, MA, USA, 1968. 243 + 12 + 10 + 6 pp.

Schriber:1969:FCS

[Sch69] Thomas J. Schriber. FORTRAN case studies illustrating fundamental business applications. Technical report, University of Michigan, Ann Arbor, MI, USA, 1969. iii + [93] pp.

Schriber:1970:FCS

[Sch70] Thomas J. Schriber. Fortran case studies for business applications. John Wiley and Sons, New York, London, Sydney, 1970. various pp.

Schriber:1971:FCS

[Sch71] Thomas J. Schriber. Fortran case studies for business applications. John Wiley and Sons, New York, London, Sydney, 1971. various pp. Distributed by Ulrich’s Books, inc., Ann Arbor, MI, USA.

Schneck:1972:ARV

[Sch72a] Paul B. Schneck. Automatic recognition of vector and parallel operations in a higher level language. ACM SIGPLAN Notices, 7(11):

45–52, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Special Issue on Control Structures in Programming Languages.

Schumann:1972:MDB

[Sch72b] U. Schumann. MAPLIB: a data bank of FORTRAN functions describing material properties. Software—Practice and Experience, 2 (1):21–28, January/March 1972. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schwartz:1972:FPE

[Sch72c] Marvin F. Schwartz. A Fortran program for an exact test of independence in an M ×N contingency table. Thesis (m.s.), Air Force Institute of Technology, Wright-Patterson Air Force Base, OH, USA, 1972. 35 pp. Distributed by U.S. National Technical Information Service, AD 743-613.

Schur:1973:TCL

[Sch73] L. D. Schur. Time-Shared Computer Languages: an Introduction to Conversational Computing. Addison-Wesley, Reading, MA, USA, 1973. ISBN 0-201-06761-7. xiv + 561 pp. LCCN QA76.7 .S361.

Schrem:1974:EES

[Sch74] E. Schrem. Erlauterungen und erganzungen zu standard FOR-TRAN IV. Report 163, Institut fur Statik und Dynamik der Luft-und Raumfahrtkon-struktionen, Universität Stuttgart, Stuttgart, Germany, 1974. [iii] + 40 pp.

Schryer:1975:UGD

[Sch75] Norman L. Schryer. A user’s guide to DODES, a double precision ordinary differential equation solver. Computing Science Technical Report 33, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1975. ii + 18 pp.

Schryer:1976:NST

[Sch76] Norman L. Schryer. Numerical solution of time-varying partial differential equations in one space variable. Computing Science Technical Report 53, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 19, 1976. ii + 44 pp.

Schmidt:1977:GMN

[Sch77] B. Schmidt. GPSS-FORTRAN — die modellerstellung nach dem baukastenverfahren. Elektronische Rechenanlagen, 19, 1:21–24, 1977. CODEN ELRAA4. ISSN 0013-5720.

Schmidt:1978:RSR

[Sch78a] B. Schmidt. Rechnermodelle: Die Simulation von Rechneranlagen mit GPSS-Fortran. R. Oldenbourg, München, Germany, 1978. ISBN 3-486-22851-X. ???? pp.

Schmidt:1978:GVI

[Sch78b] Bernd Schmidt. GPSS-Fortran, Version II, volume 6. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1978. ISBN 3-540-09037-1. ???? pp.

Schrage:1979:MPF

[Sch79a] Linus Schrage. A more portable Fortran random number generator. ACM Transactions on Mathematical Software, 5(2):132–138, June 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Schulz:1979:IFC

[Sch79b] D. Schulz. The Intel FORTRAN-80 compiler for the 8080 and 8085 microprocessors. ACM SIGPLAN Notices, 14(4):64–76, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwartz:1979:CBU

[Sch79c] Ronald David Schwartz. A comparison between using a computerized statistical learning package and a Fortran programming language for teaching statistics to high school students. Thesis (Ph.D.), University of Akron, Akron, OH, USA, 1979. x + 122 pp.

Schiller:1980:SCP

[Sch80a] Donald C. Schiller. System capacity and performance evaluation. IBM Systems Journal, 19(1):46– 67, 1980. CODEN IBMSA7. ISSN 0018-8670.

Schmidt:1980:G

[Sch80b] B. Schmidt. GPSS-Fortran. John Wiley and Sons, New York, London, Sydney, 1980. ISBN 0-471-27881-5. xiii + 523 pp. LCCN QA76.73.G18 .S36.

Schmidt:1980:GFW

[Sch80c] B. Schmidt. Gpss Fortran (Wiley Series in Computing). John Wiley and Sons, New York, London, Sydney, June 1980. ISBN 0-471-27881-5. ???? pp. LCCN ????

Schmidt:1980:GF

[Sch80d] Bernd Schmidt. Gpss Fortran. Wiley series in computing. John Wiley and Sons, New York, London, Sydney, June 1980. ISBN 0-471-27881-5. xiii + 523 pp. LCCN QA76.73.G18 .S36. US$77.95; UK£18.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471278815.

Schmidt:1980:SSF

[Sch80e] K. Schmidt. SIGMUS — ein simulator-generator für multiprozessorsysteme. Angewandte Informatik, 10:402–408, 1980. ISSN 0013-5704.

Schofield:1980:HBS

[Sch80f] Charles G. Schofield. Homogenisation/blending systems design and control for minerals processing: (with FORTRAN programs), volume 2 of Series on bulk materials handling; v.2. Trans Tech Publications, Ltd., Clausthal, Germany; Rockport, MA, USA, December 1980. ISBN 0-87849-030-2. xiii + 315 (or xiii + 321??) pp. LCCN TS176 .S36 1980.

Schwartz:1980:CBU

[Sch80g] Ronald David Schwartz. A comparison between using a computerized statistical learning package and a Fortran programming language for teaching statistics to high school students. Thesis (Ph.D.), University of Akron, Akron, OH, USA, 1980. x + 122 pp.

SDS:1964:SSF

[Sci64] Scientific Data Systems. SDS 900 series: FORTRAN II reference manual. Scientific Data Systems, Santa Monica, CA, USA, 1964. vi + 78 pp.

SDS:1965:SSF

[Sci65] Scientific Data Systems, Santa Monica, CA, Santa Monica, CA, USA. SDS 900 series Fortran II operations manual, 1965. iv + 36 pp.

SRACDED:1969:CCM

[Sci69] Science Research Associates.Computer Related Education Dept. Computing concepts in mathematics. Science Research Associates, Chicago, IL, USA, 1969. 47 pp.

SCSP:1976:F

[Sco76a] Scottish Computers in Schools Project. Fortran. The computer: yours obediently; 3b. W. and R. Chambers Ltd., London, UK, 1976. ISBN 0-550-77114-X. 109 pp.

SCSP:1976:NUC

[Sco76b] Scottish Computers in Schools Project. Now use the computer: Fortran. The computer: yours obediently 3b. W. and R. Chambers Ltd., London, UK, 1976. ISBN 0-550-77106-9. 108 pp.

Scowen:1977:DFA

[Sco77a] R. S. Scowen. The diagnostic facilities in Algol and Fortran compilers. Technical report, Division of Numerical Analysis and Computing, National Physical Laboratory, Teddington, Middlesex, 1977. 43 pp.

Scowen:1977:DMI

[Sco77b] R. S. Scowen. On detecting misspelt identifiers in Fortran. Software—Practice and Experience, 7 (4):536, July 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Scott:1978:FAC

[Sco78] James H. Scott. A FORTRAN algorithm for correcting normal resistivity logs for borehole diameter and mud resistivity. Open-file series 78-669, U.S. Geological Survey, Denver, CO, USA, 1978. 12 pp.

Sampson:1966:FIT

[SD66] Robert J. Sampson and John C. Davis. FORTRAN II trendsurface program with unrestricted input for the IBM 1620 computer. Special distribution publication — Kansas. State Geological Survey 26, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1966. 12 pp.

Sampson:1967:TRS

[SD67] Robert J. Sampson and John C. Davis. Three-dimensional response surface program in Fortran II for the IBM 1620 computer.

Computer contribution 10, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1967. 20 pp.

Strickland:1972:IML

[SD72] Roger P. Strickland, Jr. and Jimmie D. Davis. Interfacing the MPS/ 360 linear programming routine with FORTRAN programs. Technical Report ????, U.S. Dept. of Agriculture, Economic Research Service, Washington, DC, USA, November 1972. iv + 30 pp.

Stevenson:1973:PFI

[SD73] Robert L. Stevenson and George R.

Dixon. Programming in FOR-TRAN IV. William Paterson College of New Jersey, Wayne, NJ, USA, 1973. vii + 156 pp.

Sheldon:1974:FIP

[SDH74] E. Sheldon, D. R. Donati, and H. R. Hiddleston. “MIA”, a Fortran-IV program for making spin and parity assignments to high-lying single and coherent twin nuclear levels from (alpha, nucleon) angular distributions in on-resonance, compoundnuclear, channel-spin-1/2 reactions. Computer Physics Communications, 8(3):199–219, October 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465574900976.

Shettel:1980:FFC

[SDZ80a] Don L. Shettel, Ralph F. D’Andrea, and Richard J. Zinkl. FACEDT, a FORTRAN computer program to process hydrogeochemical and stream-sediment reconnaissance data for multivariate statistical analysis. Technical Report GJBX 246(80), U.S. Dept. of Energy, Grand Junction, CO, USA, 1980. 41 pp.

Shettel:1980:LFC

[SDZ80b] Don L. Shettel, Ralph F. D’Andrea, and Richard J. Zinkl. LLLSRT a FORTRAN computer program for processing Lawrence Livermore Laboratory HSSR data. Technical Report GJBX 247(80), U.S. Dept. of Energy, Grand Junction, CO, USA, 1980. 54 pp.

Spiess:1974:PF

[SE74] W. E. Spiess and Gerd Ehinger. Programmieruebungen in Fortran. Walter de Gruyter, New York, NY, USA, 1974. ISBN 3-11-003777-7.

Sears:1979:OTFa

[Sea79a] Joel L. Sears. Optimization techniques in FORTRAN. Petrocelli Books, New York, NY, USA, 1979. ISBN 0-89433-034-9. xiii + 90 pp. LCCN QA76.73.F25 S38.

Sears:1979:OTFb

[Sea79b] Joel L. Sears. Optimization Techniques in Fortran. McGraw-Hill, New York, NY, USA, August 1979. ISBN 0-07-091042-1. ???? pp. LCCN ???? US$10.00. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0070910421.

Searle:1980:RFR

[Sea80] S. R. (Shayle R.) Searle. [Regdata:

FORTRAN regdata, PL/I regdata, S table]. Technical report, Biometrics Unit, Cornell University, Ithaca, NY, USA, 1980. various pp.

Sedgwick:1977:SCF

[Sed77] Arthur Sedgwick. Structuring control in Fortran. ACM SIGPLAN Notices, 12(11):55–60, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seeds:1975:FIB

[See75] Harice L. Seeds. Fortran IV for Business and General Applications. John Wiley and Sons, New York, London, Sydney, March 1975. ISBN 0-471-77109-0. viii + 422 pp. LCCN HF5548.5.F2 S344. US$27.45. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471771090.

Seidel:1972:FEC

[Sei72] Kenneth P. Seidel. Fortran, with emphasis on the CDC lower 3000 computer series. Goodyear Pub. Co., Pacific Palisades, CA, USA, 1972. ISBN 0-87620-321-7. xi + 144 pp. LCCN QA76.73.F25S4.

Seidel:1975:TEF

[Sei75]  Robert C. Seidel.  Transfer-

function-parameter estimation from frequency response data: a FORTRAN program. NASA Technical memorandum X-3286, National Aeronautics and Space Administration, Washington, DC. USA, 1975. 31 pp.

Selfridge:1972:PFI

[Sel72] Oliver G. Selfridge. A primer for Fortran IV on-line. MIT Press, Cambridge, MA, USA, 1972. ISBN 0-262-69035-7 (paperback). vi + 319 pp. LCCN QA76.5 .S4.

Self:1977:SCF

[Sel77] Linda Elizabeth Self. The structure and complexity of Fortran programs. Thesis (m.s.), University of Florida, Gainesville, FL, USA, 1977. v + 138 pp.

Seppanen:1975:FRR

[Sep75] Marvin S. Seppanen. A Fortran routine reorganizer. Information circular — Bureau of Mines 8696, U.S. Dept. of the Interior, Bureau of Mines, Pittsburgh, PA, USA, 1975. 62 pp.

SBC:1971:CFR

[Ser71] Service Bureau Company. Call/ 360: FORTRAN reference manual. Service Bureau Corporation, Harrison, NY, USA, 1971. iv + 130 pp.

Summers:1972:AAF

[SF72] Victor Summers and Dennis J. Frailey. The Algol and/or Fortran programmer’s guide to Jovial. Technical report CP-72017, Computer Science/Operations Research Center, Institute of Technology, Southern Methodist University, Dallas, TX, USA, 1972. iv + 20 pp.

Stucki:1975:NAC

[SF75] Leon G. Stucki and Gary L. Foshee. New assertion concepts for self-metric software validation. ACM SIGPLAN Notices, 10(6): 59–71, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schnupp:1976:SPP

[SF76] P. Schnupp and Christiane Floyd.

Software: Programmentwicklung und Projektorganisation. Walter de Gruyter, New York, NY, USA, 1976. ISBN 3-11-005953-3. ???? pp.

Shimasaki:1979:PPA

[SFIK79] M. Shimasaki, S. Fukaya, K. Ikeda, and T. Kiyono. A Pascal program analysis system and profile of Pascal compilers. In Proceedings of the 12th Hawaii International Conference on System Sciences, pages 85–90. ACM Press, New York, NY 10036, USA, 1979.

Sachs:1967:MFI

[SG67] Lester M. Sachs and Murrary Geller. MOSES, a FORTRAN IV system for polyatomic molecules. International Journal of Quantum Chemistry, 1(S1):445–455, January 16–21, 1967. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory.

Suarez:1969:FPA

[SG69] Carlos B. Suárez and Mario Gallardo. A Fortran program to analyze the rotational structure of diatomic molecules. Journal of Computational Physics, 4(3):424–426, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0021999169900102.

Steinberg:1978:FII

[SG78] Horst Steinberg and Jorg Grosholz. FORTRAN IV: e. problemorientierte Programmiersprache. PU Datenverarbeitung. Siemens-Aktiengesellschaft, [Abt. Verl.], Berlin, Germany, 2. Aufl. edition, 1978. 352 pp.

Shantz:1967:WUW

[SGM^{+}67] Peter W. Shantz, R. A. German, J. G. Mitchell, R. S. K. Shirley, and C. R. Zarnke. WATFOR— the University of Waterloo FOR-TRAN IV compiler. Comm. ACM, 10(1):41–44, January 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stark:1967:ECC

[SH67] A. E. Stark and Mary C. Halden. Estimation of catch composition from sample surveys by stepwise application of raising factors, CDC 3200, Fortran IV. Transactions of the American Fisheries Society, 96 (2):230, April 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1978:FIP

[SH78] Deane K. (Deane Kingsley) Smith and Mark Holomany. A FOR-TRAN IV program for calculating X-ray powder diffraction patterns– version 7. Technical report, College of Earth and Mineral Sciences, Dept. of Geosciences Pennsylvania State University, University Park, PA, USA, 1978. iii + 63 pp.

Shavell:1965:UFS

[Sha65] Zalman A. Shavell. The use of FORTRAN in subroutines with COBOL main programs. Comm. ACM, 8(4):221–223, April 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shaw:1971:FIBa

[Sha71a] Christopher J. Shaw. FOR-TRAN information bulletin. ACM SIGPLAN Notices, 6(6):6–22, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1971:FIBb

[Sha71b] Christopher J. Shaw. FOR-TRAN information bulletin. ACM SIGPLAN Notices, 6(10):50–65, November 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.
org/citation.cfm?id=1317452.

Sharma:1976:PFI

[Sha76] K. D. Sharma. Programming in FORTRAN IV. Affiliated East-West Press, New Delhi, India, 1976. vi + 240 pp.

Shapiro:1977:FIO

[Sha77] Michael D. Shapiro. Fortran 77 input-output seems out of touch. ACM SIGPLAN Notices, 12(10): 65–69, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheridan:1959:ATC

[She59] P. Sheridan. The arithmetic translator compiler of the IBM FOR-TRAN Automatic Coding System. Comm. ACM, 2(2):9–21, February 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//community.computerhistory.
org/scc/projects/FORTRAN/paper/
p9-sheridan.pdf.

Shepard:1970:OWA

[She70a] Donald Shepard. One way analysis of variance (with sub-group averages): description of a Fortran computer programme. Discussion paper — University of Nairobi, Institute for Development Studies, no. 101; technical paper — University of Nairobi, Institute for Development Studies no. 1, Institute for Development Studies, University of Nairobi, Nairobi, Kenya, 1970. 20 pp.

Sherman:1970:TCP

[She70b] Philip M. Sherman. Techniques in computer programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1970. ISBN 0-13-899039-5. xiii + 348 pp. LCCN QA76.5 .S47.

Shearin:1978:SFI

[She78a] K. Kay Shearin. SIMUDELT:

FORTRAN IV program for simulation of delta building. Thesis (m.s.), North Carolina State University, Raleigh, NC, USA, 1978. v + 133 pp.

Sherman:1978:ANF

[She78b] A. H. Sherman. Algorithm 533 NSPIV, a FORTRAN subroutine for sparse Gaussian elimination with partial pivoting (F4). ACM Transactions on Mathematical Software, 4(4):391–398, December 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sherman:1978:NFS

[She78c] A. H. Sherman. NSPIV, a FOR-TAN subroutine for sparse Gaussian eliminating with partial pivoting (F4). Collected Algorithms from ACM, Algorithm 533, pages 1–24, 1978.

Sherman:1978:NAF

[She78d] A. H. Sherman. NSPIV: A Fortran subroutine for sparse Gaussian elimination with partial pivoting. ACM Transactions on Mathematical Software, 4(4):391–398, December 1978.

Sherman:1978:ANA

[She78e] Andrew H. Sherman. Algorithm 533: NSPIV, A FORTRAN subroutine for sparse Gaussian elimination with partial pivoting. ACM Transactions on Mathematical Software, 4(4):391–398, December 1978. CODEN ACMSCU.

ISSN 0098-3500 (print), 1557-7295 (electronic).

Shneiderman:1976:EEP

[Shn76] B. Shneiderman. Exploratory experiments in programmer behavior. International Journal of Computer and Information Sciences, 5 (2):123–143, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Shneiderman:1977:EIP

[Shn77] Ben Shneiderman. Evaluating introductory programming textbooks: a guide for students, instructors, authors and publishers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(3):56–58, August 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue on the Eighth Technical Symposium on Computer Science Education.

Shogan:1976:UMP

[Sho76a] Andrew W. Shogan. A user’s manual for PERTNET: a FOR-TRAN code for analyzing stochastic PERT networks. Working papers in management science CP-390, Center for Research in Management Science, University of California, Berkeley, Berkeley, CA, USA, 1976. 10 pp.

Shogan:1976:UMR

[Sho76b] Andrew W. Shogan. A user’s manual for REBOUND: a FORTRAN code for bounding network reliability. Working papers in management science CP-391, Center for Research in Management Science, University of California, Berkeley, Berkeley, CA, USA, 1976. 14 pp.

Short:1980:TFA

[Sho80] Kent S. Short. Two FORTRAN applications of wind-driven Ekman water transport theory: upwelling index and storm tide. NOAA Western Region computer programs and problems NWS WRCP 15, National Oceanic and Atmospheric Administration, National Weather Service, Western Region, Salt Lake City, UT, USA, 1980. 8 pp.

Shukiar:1969:FPI

[Shu69] H. J. Shukiar. A FORTRAN programmer’s introduction to SIM-SCRIPT II. Rand Corporation. Memorandum RM-5937-PR Rand Corporation. Research memorandum RM-5937-PR. Rand Corp., Santa Monica, CA, USA, 1969. ix + 26 pp.

Shum:1975:ICF

[Shu75] Che-Kwan Shum. IBM 360/CDC 6600 FORTRAN IV program conversion software. Thesis (m.s. in engr.), University of Texas at Austin, Austin, TX, USA, 1975. vi + 27 pp.

Siciliano:1974:LFE

[Sic74] Antonio Siciliano. Il linguaggio Fortran: esposizione del Fortran ANSI e delle estensioni su vari sistemi con esercitazioni di analisi numerica. Zanichelli, Bologna, Italy, 1974. viii + 265 pp.

Sidik:1972:NFIa

[Sid72a] Steven M. Sidik. NAMER: a Fortran IV program for use in optimizing designs of two-level factorial experiments given partial prior information. Technical report, U.S. National Aeronautics and Space Administration, Washington, DC, USA, 1972. 66 pp.

Sidik:1972:NFIb

[Sid72b] Steven M. Sidik. NAMER: a Fortran IV program for use in optimizing designs of two-level factorial experiments given partial prior information. NASA technical note D-6545, U.S. National Aeronautics and Space Administration, Washington, DC, USA, 1972. 66 pp.

Siebert:1974:HFP

[Sie74] H. Siebert. Höhere Fortran Programmierung. Walter de Gruyter, New York, NY, USA, 1974. ISBN 3-11-00397-5 (invalid ISBN checksum).

SIGPLAN:1976:DPA

[SIG76] SIGPLAN: ACM Special Interest Group in Programming Languages. Draft proposed ANS FORTRAN: BSR X3.9, X3J3/76. ACM Press, New York, NY 10036, USA, 1976. 160 pp.

Signorino:1980:FPS

[Sig80] John Salvatore Signorino. A Fortran package for the simulation and spectral analysis of a channelized receiver and other systems. Electrical engineering thesis (m.s.), University of Missouri — Rolla, Rolla, MO, USA, 1980. x + 220 pp.

Sikes:1971:ACB

[Sik71] Glenn H. Sikes. The analysis of continuous beams for highway bridges IV computer program Fortran IV-F: a user’s manual. Technical Report ????, State Highway Department of Georgia, Bridge Division, Atlanta, GA, USA, 1971. various pp.

Sillay:1961:LGF

[Sil61] Frank I. Sillay. Load and go FORTRAN: a program development and teaching system for the IBM 1620. Publication 150, Massachusetts Institute of Technology, Dept. of Civil Engineering, Civil Engineering Systems Laboratory, Cambridge, MA, USA, 1961. 33 + [9] pp.

Silver:1971:SFI

[Sil71] Gerald A. Silver. Simplified FOR-TRAN IV programming. Harcourt, Brace, Jovanovich, San Diego, CA, USA, 1971. ISBN 0-15-581049-9. vi + 361 pp. LCCN QA76.5 .S552. Second ed. (c1976) published under title: Simplified ANSI FOR-TRAN IV programming.

SVS:1980:AF

[Sil80] Silicon Valley Software, Inc. Apple Fortran, 1980.

Simpson:1966:TSC

[Sim66] Stephen Milton Simpson. Timeseries computations in FORTRAN and FAP. Addison-Wesley, Reading, MA, USA, 1966. 1120 pp. LCCN QA276 .S52.

Simpson:1976:FTA

[Sim76a] J. C. Simpson. Fortran translation of algorithm 409 discrete Chebyshev curve fit. ACM Transactions on Mathematical Software, 2(1): 95–97, March 1976.

Simpson:1976:AFT

[Sim76b] Joseph C. Simpson. Algorithm 501: Fortran translation of Algorithm 409, discrete Chebychev curve fit [E2]. ACM Transactions on Mathematical Software, 2(1):95–97, March 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Fut78].

Sinclair:1973:DEF

[Sin73] Annette Sinclair. Determination of extremal functions in H^{p} by a Fortran program. SIAM Journal on Numerical Analysis, 10(1):137– 146, March 1973. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Singh:1978:FPM

[Sin78] Devindar Singh. A FORTRAN program for multicrop farms to determine operations schedule, field machinery requirements and costs. Agricultural economics report 331, Dept. of Agricultural Economics, Michigan State University, East Lansing, MI, USA, 1978. 110 pp.

Sites:1978:PTS

[Sit78] Richard L. Sites. Programming tools: statement counts and procedure timings. ACM SIGPLAN Notices, 13(12):98–101, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1962:FASa

[SJ62a] Robert E. (Robert Elijah) Smith and Dora E. Johnson. Fortran autotester: a self-training course designed to emancipate the scientist and engineer from the need for the professional programmer. John Wiley and Sons, New York, London, Sydney, 1962. ISBN 0-471-80337-5. 176 pp. LCCN QA 76.5 S65f 1962a.

Smith:1962:FASb

[SJ62b] Robert E. (Robert Elijah) Smith and Dora E. Johnson. Fortran autotester: a self-training course designed to emancipate the scientist and engineer from the need for the professional programmer. John Wiley and Sons, New York, London, Sydney, 1962. 176 pp.

Smith:1963:FAS

[SJ63] Robert E. (Robert Elijah) Smith and Dora E. Johnson. Fortran autotester: a self-training course designed to emancipate the scientist and engineer from the need for the professional programmer. John Wiley and Sons, New York, London, Sydney, 1963. 179 pp.

Smith:1972:FTP

[SJ72] Robert E. (Robert Elijah) Smith and Dora E. Johnson. FOR-TRAN: texto programado. Serie Limusa Wiley de libros programados. Limusa-Wiley, Mexico, DF, Mexico, 1972. 158 pp.

Schlaegel:1969:FIP

[SK69] Bryce E. Schlaegel and D. L. Kulow. Fortran IV program for stand table projection of forest growth. Bulletin 583T, West Virginia University Agricultural Experiment Station, Morgantown, WV, USA, 1969. 18 pp.

Scarborough:1980:IOF

[SK80] Randolph G. Scarborough and Harwood G. Kolsky. Improved optimization of Fortran object programs. IBM Journal of Research and Development, 24(6):660–676, November 1980. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Skeel:1979:ILB

[Ske79] R. D. Skeel. An incomplete list of better FORTRAN codes for IVPS in ODES. ACM SIGNUM Newsletter, 14(2):22–24, June 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Slater:1967:FPE

[Sla67] Lucy Joan Slater. Fortran programs for economists. Cambridge University. Dept. of Applied Economics. Occasional papers 13, Cambridge University Press, Cambridge, UK, 1967. 150 pp.

Slater:1971:FSB

[Sla71] Lucy Joan Slater. First steps in basic Fortran. Chapman and Hall, Ltd., London, UK, 1971. ISBN 0-412-10120-3. viii + 104 pp. LCCN QA76.5 .S556.

Slater:1972:MFP

[Sla72] Lucy Joan Slater. More Fortran programs for economists, volume 34 of Occasional papers / University of Cambridge. Department of Applied Economics. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-09722-3 (paperback). vii + 148 (or vii + 146??) pp. LCCN HG136.C3 O3 v.34.

Slezak:1975:MIB

[Sle75] Robert Slezak. A machine independent basic FORTRAN compiler for mini-computer implementation. Thesis (m.s.), Illinois Institute of Technology, Chicago, IL, USA, 1975. v + 53 pp.

Slimick:1971:CSI

[Sli71] John Slimick. Current systems implementation languages: One user’s view. ACM SIGPLAN Notices, 6(9):20–28, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sloan:1968:FCC

[Slo68] Karen F. Sloan. Fortran and Cobol: a comparison of computer languages. Honors college honors papers, Department of Mathematics, Kent State University, Kent, OH, USA, 1968. 29 pp.

Stein:1964:CP

[SM64] M. L. Stein and W. D. Munro.

Computer Programming. Academic Press, New York, NY, USA, 1964.

Stein:1966:FIPa

[SM66a] Marvin L. Stein and William D. Munro. A Fortran introduction to programming and computers, including Fortran IV. Academic Press, New York, NY, USA, 1966. vi + 122 pp.

Stein:1966:FIPb

[SM66b] Marvin Leonard Stein and William Delmar Munro. A Fortran introduction to programming and computers, including Fortran 1V. Academic Press, New York, NY, USA, 1966. 122 pp.

Stein:1968:CP

[SM68] M. L. Stein and W. D. Munro. Computer Programming. Academic Press, New York, NY, USA, 4 edition, 1968.

Schriber:1970:FAB

[SM70] Thomas J. Schriber and Laurence A. Madeo. Fortran applications in business administration. Technical report, University of Michigan. Graduate School of Business Administration, Ann Arbor, MI, USA, 1970.

Schick:1972:FE

[SM72a] William Schick and Charles J. Merz. Fortran for engineering. Mc-Graw-Hill, New York, NY, USA, 1972. ISBN 0-07-055276-2. xv + 367 pp. LCCN TA345 .S53.

Schick:1972:SMA

[SM72b] William Schick and Charles J. Merz. Solutions manual to accompany FORTRAN for engineering. McGraw-Hill, New York, NY, USA, 1972. 110 pp.

Sturgul:1972:AFI

[SM72c] John R. Sturgul and Michael J.

Merchant. Applied Fortran IV programming. Wadsworth, Pacific Grove, CA, USA, 1972. ISBN 0-534-00128-9. xiii + 399 pp. LCCN QA76.73.F25 S85.

Skop:1973:FIP

[SM73a] Richard A. Skop and James Mark.

A Fortran IV program for computing the static deflections of structural cable arrays. NRL report 7640, Naval Research Laboratory, Washington, DC, USA, 1973. iii + 90 pp.

Svehla:1973:FIC

[SM73b] Roger A. Svehla and Bonnie J.

McBride. FORTRAN IV computer program for calculation of thermodynamic and transport properties of complex chemical systems. NASA technical note NASA TN D-7056, National Aeronautics and Space Administration; National Technical Information Service, Washington, DC. USA, 1973. iv + 174 pp.

Smith:1975:FAD

[SM75] J. L. Smith and H. G. MacKenzie.

FORDATA, a data base management package under FORTRAN in the Cyber76. Technical report, CSIRO, Canberra, Australia,

Saltykov:1976:PIF

[SM76a] A. I. (Albert Ivanovich) Saltykov and G. I. (Grigorii Ivanovich) Makarenko. Programmirovanie na iazyke Fortran. Bibliotechka programmista. Nauka, Moscow, Russia, 1976. 254 + [2] pp.

Sturgul:1976:AFI

[SM76b] John R. Sturgul and Michael J. Merchant. Applied Fortran IV programming. Wadsworth, Pacific Grove, CA, USA, second edition, 1976. ISBN 0-534-00440-7. xvii + 391 pp. LCCN QA76.73.F25S85 1976.

Sheldon:1971:CTD

[SMD71] E. Sheldon, S. Mathur, and D. Donati. Computation of total, differential and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bσ) and (a, bσ−σ). (III) Fortran translations of the Algol programs ‘Mandy’ and ‘Barbara’. Computer Physics Communications, 2(5):272–287, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/article/
                                        pii/0010465571900038.

Smith:1963:CMF

[Smi63a] D. D. Smith. Character manipulation in 7090 FORTRAN. Comm. ACM, 6(8):440, August 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1963:FPC

[Smi63b] Deane Kingsley Smith. A FOR-TRAN program for calculating Xray powder diffraction patterns. Technical report, University of California, Lawrence Radiation Laboratory, Livermore, CA, USA, 1963. iii + 72 pp. Distributed by Clearinghouse, U.S. Department of Commerce.

Smith:1966:GDP

[Smi66] F. G. (Frederick Gordon) Smith.

Geological data processing using Fortran IV. Harper’s geoscience series. Harper & Row, New York, NY, USA, 1966. xv + 284 pp.

Smith:1967:BF

[Smi67a] Robert E. (Robert Elijah) Smith.

The bases of FORTRAN. Control Data Institute, Minneapolis, MN, USA, 1967. vii + 253 pp.

Smith:1967:SCU

[Smi67b] Robert E. (Robert Elijah) Smith.

School computer-use plan; orientation course. Control Data Corp., Minneapolis, MN, USA, 1967. vii + 281 pp.

Smith:1968:CC

[Smi68] Robert E. (Robert Elijah) Smith.

Competence course. Control Data Corp., Minneapolis, MN, USA, 1968. vi + 296 pp.

Smith:1970:BT

[Smi70a] Robert E. (Robert Elijah) Smith.

The biorhythm theory. Control Data Corp., Minneapolis, MN, USA, 1970. 58 pp.

Smith:1970:CES

[Smi70b] Robert E. (Robert Elijah) Smith. Computer explorer series, in Fortran and Basic. Control Data Corporation, Minneapolis, MN, USA, 1970. various pp.

Smith:1970:DHT

[Smi70c] Robert E. (Robert Elijah) Smith. Distances here to there. Control Data Institute, Minneapolis, MN, USA, 1970. vi + 74 pp.

Smith:1970:EA

[Smi70d] Robert E. (Robert Elijah) Smith. Excursion in astrology. Control Data Corp., Minneapolis, MN, USA, 1970. 150 pp.

Smith:1970:F

[Smi70e] Robert E. (Robert Elijah) Smith. Factorials. Control Data Corp., Minneapolis, MN, USA, 1970. 58 pp.

Smith:1970:IM

[Smi70f] Robert E. (Robert Elijah) Smith. Interest in money. Control Data Corp., Minneapolis, MN, USA, 1970. 50 pp.

Smith:1970:LAN

[Smi70g] Robert E. (Robert Elijah) Smith. A look at numerology. Control Data Corp., Minneapolis, MN, USA, 1970. 118 pp.

Smith:1970:PH

[Smi70h] Robert E. (Robert Elijah) Smith. Population holocaust. Control Data Corp., Minneapolis, MN, USA, 1970. 65 pp.

Smith:1970:PF

[Smi70i] Robert E. (Robert Elijah) Smith.

Primes and factors. Control Data Corp., Minneapolis, MN, USA, 1970. 69 pp.

Smith:1970:RN

[Smi70j] Robert E. (Robert Elijah) Smith.

Random numbers. Control Data Corp., Minneapolis, MN, USA, 1970. 82 pp.

Smith:1970:GT

[Smi70k] Robert Elijah Smith. Game theory. Control Data Corporation, Minneapolis, MN, USA, 1970. 81 pp.

Smith:1971:BF

[Smi71a] Robert E. (Robert Elijah) Smith.

The bases of FORTRAN. Control Data Institute, Minneapolis, MN, USA, 1971. vii + 253 pp.

Smith:1971:BA

[Smi71b] Robert E. (Robert Elijah) Smith.

Big arithmetic. Control Data Corporation, Minneapolis, MN, USA, 1971. vi + 82 pp.

Smith:1971:SD

[Smi71c] Robert E. (Robert Elijah) Smith.

Sorting data. Control Data Corporation, Minneapolis, MN, USA, 1971. vi + 120 pp.

Smith:1971:SP

[Smi71d] Robert E. (Robert Elijah) Smith.

Statistical procedures. Control Data Corporation, Minneapolis, MN, USA, 1971. v + 137 pp.

Smith:1971:YH

[Smi71e] Robert E. (Robert Elijah) Smith. Your handwriting. Control Data Corp., Minneapolis, MN, USA, 1971. vi + 105 pp.

Smith:1972:BT

[Smi72a] Robert E. (Robert Elijah) Smith. The biorhythm theory. Control Data Corp., Minneapolis, MN, USA, 1972. 88 pp.

Smith:1972:LP

[Smi72b] Robert E. (Robert Elijah) Smith. Linear programming. Control Data Institute, Minneapolis, MN, USA, 1972. vi + 160 pp.

Smith:1972:M

[Smi72c] Robert E. (Robert Elijah) Smith. Metrology. Control Data Institute, Minneapolis, MN, USA, 1972. vi + 114 pp.

Smith:1972:VE

[Smi72d] Robert E. (Robert Elijah) Smith. Visual elements. Control Data Corp., Minneapolis, MN, USA, 1972. vi + 182 pp.

Smith:1973:BFS

[Smi73a] Robert E. (Robert Elijah) Smith. The bases of FORTRAN: a selftraining approach to computer programming. Control Data Corporation, Minneapolis, MN, USA, fifth edition, 1973. vii + 282 + [113] pp.

Smith:1973:PR

[Smi73b] Robert E. (Robert Elijah) Smith. Palm reading. Control Data Institute, Minneapolis, MN, USA, 1973. vi + 106 pp.

Smith:1973:PH

[Smi73c] Robert E. (Robert Elijah) Smith.

Population holocaust. Control Data Corp., Minneapolis, MN, USA, second edition, 1973. 74 pp.

Smith:1973:P

[Smi73d] Robert E. (Robert Elijah) Smith.

Posters. Control Data Institute, Minneapolis, MN, USA, 1973. v + 112 pp.

Smith:1973:SG

[Smi73e] Robert E. (Robert Elijah) Smith.

Simulated gaming. Control Data Institute, Minneapolis, MN, USA, 1973. 128 pp.

Smith:1977:PFD

[Smi77] David N. Smith. Proposals for Fortran data structures. ACM SIGPLAN Notices, 12(9):85–102, September 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smillie:1978:IAA

[Smi78] K. W. Smillie. Introducing APL, ALGOL and FORTRAN. Technical report, Dept. of Computing Science, University of Alberta, Edmonton, Alberta, Canada, 1978. i + 61 pp.

Smith:1979:NFC

[Smi79] D. H. (David Huston) Smith. New FORTRAN computer programs to acquire and process isotopic mass spectrometric data. ORNL/TM 7002, Dept. of Energy, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1979. ix + 87 pp. For sale by the National Technical Information Service.

Smith:1980:EVT

[Smi80] Kathryn A. (Kathryn Anne) Smith. Evaluation of verification and testing tools for FOR-TRAN programs. NASA technical memorandum 80205, National Aeronautics and Space Administration, Scientific and Technical Information Branch, Washington, DC, USA, 1980. 21 pp. For sale by the National Technical Information Service.

Steel:1965:FPD

[SMM65] Mary Nan Steel, Frank L. Mc-Crackin, and John Mandel. A FORTRAN program for determining an empirical expression or a quantity measured at combinations of several levels of each of two variables. [U.S.] National Bureau of Standards. Technical note no. 259. United States Government Printing Office, Washington, DC, USA, 1965. 61 pp. For sale by the Superintendent of Documents, Govt. Print. Off.

SoftwareHouse:1980:SDB

[Sof80] Software House. System 1022 data base management system: user’s reference manual: host language interface, Fortran, COBOL. Software House, Cambridge, MA, USA, 1980. 129 pp.

Solberg:1964:VIS

[Sol64] Kjell Oddvar Solberg. “VOIFLO I;” a steady state FORTRAN code for the hydraulics of a boiling loop.

Report 85, Institutt for atomenergi, Kjeller, Norway, 1964. 59 pp.

Solvberg:1969:F

[Sol69] Arne Solvberg. Fortran. Technical report, Norges Tekniske Høgskole, Tapir, Norway, 1969. 154 pp.

Solntseff:1978:PLI

[Sol78] N. Solntseff. Programming languages for introductory computing courses: a position paper. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 10(1):119–124, February 1978. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Papers of the SIGCSE/ CSA Technical Symposium on Computer Science Education.

Somervaille:1971:CTF

[Som71] Ian James Somervaille. Computer translation from FORTRAN to PL1. UNICIV report R-73, School of Civil Engineering, University of New South Wales, Kensington, NSW, Australia, 1971. ISBN 0-85841-013-3. 33 pp.

Southwell:1967:FIL

[Sou67] R. L. Southwell. FORTRAN IV library subroutines. Technical report, West Virginia University Conputer Center, Morgantown, WV, USA, 1967. 35 pp.

Southwell:1968:FIL

[Sou68] R. L. Southwell. FORTRAN IV library subroutines. Technical report, West Virginia University Computer Center, Morgantown, WV, USA, 1968. 40 pp.

Soukup:1971:CPF

[Sou71] J. Soukup. A computer program for finding a general distribution function of several sets of random data (UNIVAC 1108 — Calcomp 663 — Fortran V.). Canada. Geological Survey. Paper 71-20, Dept. of Energy, Mines and Resources, Ottawa, Ontario, Canada, 1971. v + 72 pp.

Soylemez:1971:CGF

[Soy71] Mustafa Ismail Ekrem Sezer Soylemez. Computer generated FOR-TRAN programs to prepare material and energy balances for process units. Thesis (Ph.D.), University of Pennsylvania, Philadelphia, PA, USA, 1971. xxxii + 255 pp.

Sterling:1970:CCS

[SP70] Theodor D. Sterling and Seymour V. Pollack. Computing and Computer Science: a First Course With Fortran IV. MacMillan Publishing Company, New York, NY, USA, 1970. ISBN 0-02-417110-7. xvi + 398 pp. LCCN ???? US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0024171107.

Skordalakis:1978:CF

[SP78] E. Skordalakis and G. Papakonstantinou. Coroutines in FOR-TRAN. ACM SIGPLAN Notices, 13(9):76–84, September 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sparks:1973:SAAa

[Spa73] D. N. Sparks. Statistical algorithms: Algorithm AS 58: Euclidean cluster analysis. Applied Statistics, 22(1):126–130, March 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/58. See remark

[?].

Spaeth:1975:AOR

[Spa75a] Helmuth Spaeth. Ausgewaehlte Operations Researd-algorithmen in Fortran. R. Oldenbourg, München, Germany, 1975. ISBN 3-486-20121-2.

Sparks:1975:ISL

[Spa75b] Russell Owen Sparks. An investigation of source language recovery for FORTRAN IV. Thesis (m.s.), University of Arkansas, Fayetteville, AR, USA, 1975. iv + 215 pp.

Spath:1975:AOR

[Spa75c] Helmuth Spath. Ausgewahlte Operations Research-Algorithmen in FORTRAN. Verfahren der Datenverarbeitung. R. Oldenbourg, München, Germany, 1975. ISBN 3-486-20121-2. 115 pp. LCCN T57.6.S63.

Spath:1976:CA

[Spä76] H. Späth. L_{1} cluster analysis.

Computing, 16(4):379–387, 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Spath:1979:ACL

[Spä79a] H. Späth. Algorithm 39. Clusterwise linear regression. Computing, 22(4):367–373, 1979. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). See correction [?].

Spath:1979:AOF

[Spa79b] Helmuth Spath. Ausgewahlte Operations-research-Software in FORTRAN. R. Oldenbourg, München, Germany, 1979. ISBN 3-486-23911-2. 163 pp.

Spath:1980:CAA

[Spa80] Helmuth Spath. Cluster analysis algorithms for data reduction and classification of objects. Computers and their applications. Halsted Press, New York, USA, June 1980. ISBN 0-85312-141-9, 0-470-26946-4 (Halsted). 226 pp. LCCN QA278.S6813. UK£15.00. Translation of Cluster-Analyse-Algorithmen zur Objektklassifizierung und Datenreduktion by Ursula Bull.

Sperry:1966:TMS

[Spe66a] Sperry Rand Corporation. Univac Division, New York, NY, USA. 1107 thin-film memory system: Fortran IV programmers reference manual, 1966. various pp.

Sperry:1966:MSF

[Spe66b] Sperry Rand Corporation. Univac Division, New York, NY, USA. 1108 multi-processor system: Fortran V programmers reference manual, 1966. various pp.

Sperry:1967:F

[Spe67] Sperry Rand Corporation. Univac Division, New York, NY, USA. FORTRAN, 1967. various pp.

Spencer:1969:IMP

[Spe69a] Donald D. Spencer. Instructor’s manual for programming with USA standard FORTRAN and FOR-TRAN IV. Blaisdell Pub. Co., Waltham, MA, USA, 1969. 26 pp.

Spencer:1969:PUS

[Spe69b] Donald D. Spencer. Programming with USA standard FOR-TRAN and FORTRAN IV. A Blaisdell book in pure and applied mathematics. Blaisdell Pub. Co., Waltham, MA, USA, 1969. xi + 243 pp.

Sperry:1969:FVP

[Spe69c] Sperry Rand Corporation, Univac Division, New York, NY, USA. FORTRAN V programmers reference, 1969. various pp.

Spencer:1970:SSF

[Spe70a] Donald D. Spencer. SIMPLE-TRAN, a subset of FORTRAN. Abacus Computer Corp., Daytona Beach, FL, USA, 1970. 33 pp.

Sperry:1970:UFF

[Spe70b] Sperry Rand Corporation.Univac Division. UNIVAC fundamentals of FORTRAN. Sperry Rand Corporation, New York, NY, USA, 1970. various pp.

Sperry:1973:FVL

[Spe73a] Sperry Rand Corporation.Univac Division. FORTRAN V library: programmer reference. Sperry-UNIVAC Computer Systems, New York?, 1973. ???? pp.

Sperry:1973:FVP

[Spe73b] Sperry Rand Corporation.Univac Division. FORTRAN V: programmer reference. Sperry-UNIVAC Computer Systems, New York?, 1973. ???? pp.

Sperry:1974:FTP

[Spe74a] Sperry Rand Corporation.Univac Division. FORTRAN translator,: programmer reference. Sperry-UNIVAC Computer Systems, St. Paul, MN, USA, 1974. ???? pp.

Sperry:1974:UFF

[Spe74b] Sperry Rand Corporation.Univac Division. Univac fundamentals of FORTRAN: programmer reference. Sperry-UNIVAC Computer Systems, St. Paul, MN, USA, 1974. ???? pp.

Sperry:1976:FAP

[Spe76a] Sperry Rand Corporation.Univac Division. FORTRAN (ASCII): programmer reference. UNIVAC Computer Systems, New York?, 1976. ???? pp.

Sperry:1976:FVP

[Spe76b] Sperry Rand Corporation.Univac Division. FORTRAN V programmer reference. Sperry Univac Computer System, King of Prussia, PA, USA, 1976. various pp.

Spencer:1977:FP

[Spe77a] Donald D. Spencer. FORTRAN programming. Camelot Pub. Co., Ormond Beach, FL, USA, 1977. ISBN 0-89218-006-4, 0-89218-007-2. 128 pp. LCCN QA76.73 .F25S621.

Spencer:1977:FW

[Spe77b] Donald D. Spencer. FORTRAN workbook. Camelot Pub. Co., Ormond Beach, FL, USA, 1977. ISBN 0-89218-018-8. 46 pp.

Spencer:1977:IMF

[Spe77c] Donald D. Spencer. Instructor’s manual for FORTRAN programming. Camelot Pub. Co., Ormond Beach, FL, USA, 1977. ISBN 0-89218-011-0. 32 pp.

Spencer:1977:PSF

[Spe77d] Donald D. Spencer. Problem Solving With Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, April 1977. ISBN 0-13-720094-3. xiv + 320 pp. LCCN QA76.73 .F25S64. US$29.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0137200943.

Spencer:1978:VMT

[Spe78a] Donald D. Spencer. Visual masters for teaching FORTRAN programming. Camelot Pub. Co., Ormond Beach, FL, USA, 1978. ISBN 0-89218-035-8. 64 pp.

Sperry:1978:FAP

[Spe78b] Sperry Rand Corporation.Univac Division. FORTRAN (ASCII): programmer reference. UNIVAC Computer Systems, St. Paul, MN, USA, 1978. various pp.

Sperry:1979:FAL

[Spe79] Sperry Rand Corporation.Univac Division. FORTRAN (ASCII) level 9R1: programmer reference. Sperry-UNIVAC Computer Systems, St. Paul, MN, USA, 1979. ???? pp.

Speelpenning:1980:CFP

[Spe80a] B. Speelpenning. Compiling Fast Partial Derivatives of Functions Given by Algorithms. PhD thesis, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, January 1980.

Spencer:1980:FP

[Spe80b] Donald D. Spencer. Fortran Programming. Camelot Pub. Co., Ormond Beach, FL, USA, second edition, January 1980. ISBN 0-89218-042-0. 128 pp. LCCN QA76.73.F25 S62 1980 Sci-Eng. US$11.93. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0892180420.

Sperry:1980:FAL

[Spe80c] Sperry Rand Corporation.Univac Division. FORTRAN (ASCII) level 9R1: summary. Sperry-UNIVAC Computer Systems, St. Paul, MN, USA, 1980. l v. pp.

Sperry:1980:FIO

[Spe80d] Sperry Rand Corporation.Univac Division. Fortran IV OS/3: programmer reference. Sperry Univac, Blue Bell, PA, USA, 1980. various pp.

Sperry:1980:SFA

[Spe84] Sperry Corporation. Series 1100 FORTRAN (ASCII) level 11R1: release description. Sperry, New York, NY, USA, 1980 (or 1984??). various pp.

Sperry:1969:MSC

[Spe9 ] Sperry Rand Corporation. Univac Division, New York, NY, USA. 1108 multi-processor system: conversational FORTRAN V service language, provisional language specification, 1969 (??). various pp.

Spira:1965:CVZ

[Spi65] Robert Spira. Check values, zeros and Fortran programs for the Riemann zeta function and its first three derivatives. Technical report 1, University of Tennessee. University Computing Cente, Knoxville, TN, USA, 1965. 154 pp.

Spiess:1970:EPF

[Spi70] Wolfgang E. Spiess. Einfuhrung in das Programmieren in Fortran. De Gruyter Lehrbuch. W. de Gruyter and Co., Berlin, Germany, 1970. 196 pp.

Spicer:1972:SAA

[Spi72] C. C. Spicer. Statistical algorithms: Algorithm AS 52: Calculation of power sums of deviations about the mean. Applied Statistics, 21(2):226–227, June 1972. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://lib.stat.cmu.edu/
apstat/52.

Spinks:1980:PIF

[Spi80] Colegate Villaret Spinks. Programmed instruction: FORTRAN tutor. Thesis (m.s.), Computing Science, Texas A and M University, College Station, TX, USA, 1980. x + 249 pp.

Squires:1970:CPP

[Squ70] R. H. Squires. A computer program for the presentation of pollen data; manual of the pollen diagram plotting program (RHS 1) written in IBM FORTRAN IV for use with an IBM 1130 under disk monitor system (version 2) requiring a 1403 line printer and a 1627 (model 1) graph plotter. Occasional papers series 11, Dept. of Geography, University of Durham, Durham, England, 1970. v + 53 pp.

Spiess:1972:EPF

[SR72] Wolfgang Eckehard Spiess and Friedrich Georg Rheingans. Einfuhrung in das Programmieren in FORTRAN. De Gruyter Lehrbuch. Walter de Gruyter, New York, NY, USA, 1972. 216 pp.

Sheldon:1973:CTD

[SR73] E. Sheldon and V. C. Rogers. Computation of total and differential cross section for compound nuclear reactions of the type (a, a), (a, a′), (a, b), (a, γ), (a, γ − γ), (a, bγ) and (a, bγ − γ): (IV) FORTRAN program ‘CINDY’. Computer Physics Communications, 6(3):99–131, September 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465573900490.

Spiess:1974:EPF

[SR74] W. E. Spiess and F. G. Rheingans.

Einführung in das Programmieren in Fortran. Walter de Gruyter, New York, NY, USA, 4 edition, 1974. ISBN 3-11-005747-6.

Smith:1976:SLP

[SR76] Carol Smith and Jon Rickman. Selecting languages for pedagogical tools in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 8(3):39– 47, July 1976. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 6th SIGCSE Symposium on Computer Science Education.

SrinivasaRao:1969:LEF

[Sri69] K. Srinivasa Rao. Lectures on elements of Fortran programming. Matscience report 68, Institute of Mathematical Sciences, Madras, India, 1969. 108 pp.

Swingle:1966:PAP

[SS66] Wayne E. Swingle and H. S.

Swingle. Population analysis programs, IBM 1620, Fortran/ Format. Transactions of the American Fisheries Society, 95(2):232– 233, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sage:1968:IPT

[SS68a] F. H. Sage and D. V. Smith. The introduction of a paging technique into the symmetric list processor, SLIP. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ?? (9):34–46, April 1968. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Solinas:1968:MPF

[SS68b] G. C. Solinas and G. Silini. MO-TUS — programma Fortran per la elaborazione di dati sperimentali relativi all’accrescimento di tumori trapiantabili in animali di laboratorio [di] G. C. Solinas [e] G. Silini. Technical Report RT-BIO (68) 1, Comitato nazionale per l’energia nucleare, Roma, Italy, 1968. 39 pp.

Schoch:1970:PF

[SS70] Henning Schoch and Wolfgang Schatz. Programmierung in FOR-TRAN. BSB Teubner, Leipzig, Germany, 1970. 406 pp.

Schlechtendahl:1972:CRF

[SS72] E. G. Schlechtendahl and U. Schumann. Correspondence: Remarks on “Faults in functions, in ALGOL and FORTRAN”. The Computer Journal, 15(2):147, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
15/Issue_02/150147.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/147.tif. See

[Hil71].

Saylor:1973:AAC

[SS73] Paul E. Saylor and James D. Sebastian. ACM Algorithm 460: Calculation of optimum parameters for alternating direction implicit procedures [D3]. Comm. ACM, 16 (10):633–635, October 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stern:1974:FSA

[SS74] Robert A. Stern and Nancy B.

Stern. FORTRAN supplement to accompany Principles of data processing. John Wiley and Sons, New York, London, Sydney, 1974. ISBN 0-471-82325-2. v + 137 pp.

Swarztrauber:1975:EFSa

[SS75a] Paul Swarztrauber and Roland Sweet. Efficient FORTRAN subprograms for the solution of elliptic partial differential equations. National Center for Atmospheric Research (U.S.) Technical note TN-109, Atmospheric Technology Division, National Center for Atmospheric Research, Boulder, CO, USA, 1975. 139 pp.

Swarztrauber:1975:EFSb

[SS75b] Paul Swarztrauber and Roland Sweet. Efficient FORTRAN subprograms for the solution of elliptic partial differential equations (abstract). ACM SIGNUM Newsletter, 10(4):30, December 1975. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Silver:1976:SAF

[SS76] Gerald A. Silver and Joan B. Silver. Simplified ANSI FORTRAN IV programming. Harcourt, Brace, Jovanovich, San Diego, CA, USA, second edition, 1976. ISBN 0-15-581040-5. xi + 335 pp. LCCN QA76.73.F25 S56 1976 Sci-Eng. First ed. published in 1971 under title: Simplified FORTRAN IV programming. Includes index.

Schultheis:1978:ANS

[SS78a] H. Schultheis and R. Schultheis. An algorithm for non-smoothing contour representations of twodimensional arrays. Computing, 19(4):381–387, 1978. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stern:1978:PFC

[SS78b] Robert A. Stern and Nancy B. Stern. Programacion Fortran: complemento del texto, Principios de procesamiento de datos. Editorial Limusa, Mexico, DF, Mexico, 1978. 135 pp.

Swartzrauber:1979:AEF

[SS79a] P. N. Swartzrauber and R. A. Sweet. Algorithm 541: Efficient FORTRAN subprograms for the solution of separable elliptic partial differential equations. ACM Transactions on Mathematical Software, 5(3):352– 364, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Swartztrauber:1979:AEF

[SS79b] Paul N. Swartztrauber and Roland A. Sweet. Algorithm 541: Efficient Fortran subprograms for the solution of separable elliptic partial differential equations [D3]. ACM Transactions on Mathematical Software, 5(3):352–364, September 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Swarztrauber:1979:EFS

[SS79c] P. N. Swarztrauber and R. A.

Sweet. Efficient Fortran subprograms for the solution of separable elliptic partial differential equations. ACM Transactions on Mathematical Software, 5(3):352– 364, September 1979.

Saran:1977:CPN

[SSS77] Swami Saran, Prabhat K. Swamee, and Kiran K. Singh. Computer programming and numerical methods: [for engineers]. Sarita Prakashan, Meerut, India, 1977. 378 pp.

Steuerwalt:1978:CRF

[SSS78] Michael Steuerwalt, Paul Swartztrauber, and Roland Sweet. Certification report on “Efficient FOR-TRAN subprograms for the solution of elliptic partial differential equations”. Technical Report LA-7524-MS, U.S. Dept. of Energy, Washington, DC, USA, October 1978. 17 pp.

Setzer:1972:CFI

[SST72] Valdemar W. Setzer, Imre Simon, and Kowaltowski Tomasz. Curso de Fortran IV basico. Editora Edgard Blucher, São Paulo, Brazil, 1972. 102 pp.

Sander:1973:FIP

[ST73a] D. M. Sander and G. T. Tamura. A FORTRAN IV program to simulate air movement in multistorey buildings. Computer program 35, National Research Council of Canada, Division of Building Research, Ottawa, Ontario, Canada, 1973. 50 pp.

Sparks:1973:CFS

[ST73b] D. N. Sparks and A. D. Todd. A comparison of FORTRAN subroutines for calculating latent roots and vectors. Applied Statistics, 22 (2):220–225, 1973. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1973:SAAb

[ST73c] D. N. Sparks and A. D. Todd. Statistical algorithms: Algorithm AS 60: Latent roots and vectors of a symmetric matrix. Applied Statistics, 22(2):260–265, June 1973. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.
                                        cmu.edu/apstat/60. See corrigen-

dum [?].

Stanfield:1960:TPF

[Sta60] R. Stanfield. Three packages of FORTRAN subroutines, MI DET1 …, MI EGN1 …, MI QRT1. Memorandum CC-158, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 3 pp.

Stammler:1965:FVN

[Sta65] Rudi J. J. Stamm’ler. The FOR-TRAN version of the neutron thermalization code K-7 THER-MOS for the CDC-3600 computer. Kjeller report 101, Institutt for atomenergi, Kjeller, Norway, 1965. 14 pp.

SCC:1969:SIF

[Sta69] Standard Computer Corporation, ???? Standard IC-4000 FOR-TRAN IV reference manual, 1969. various pp.

Stacey:1974:FIC

[Sta74] G. M. Stacey. Fortran interface to the CODASYL Database Task Group specifications. The Computer Journal, 17(2):124–129, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Stark:1975:CPH

[Sta75] Peter A. Stark. Computer Programming Handbook. Tab Books, Blue Ridge Summit, PA, USA, September 1975. ISBN 0-8306-5752-5 (hardcover), 0-8306-4752-X (paperback). 506 pp. LCCN QA76.6.S7. US$12.95.

Steinberg:1960:MPF

[Ste60a] J. R. Steinberg. MI POIF: a FOR-TRAN version of MI POIS (CC-149). Memorandum CC-156, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 1 pp.

Steinberg:1960:MSC

[Ste60b] J. R. Steinberg. MI scope:

a conversion to FORTRAN of MI APLT. Memorandum CC-154, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 1 pp.

Sternlight:1960:SF

[Ste60c] David Sternlight. Self-programming with FORTRAN. Thesis (b.s.), Dept. of Economics and Social Sciences, Massachusetts Institute of Technology, Cambridge, MA, USA, 1960. 22 pp.

Steffan:1970:AFF

[Ste70] William Lee Steffan. An automatic FORTRAN flow chart generator for the GE-425. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1970. 195 pp.

Steffensen:1972:FIP

[Ste72a] R. J. Steffensen. A FORTRAN IV program for thermochemical calculations involving the elements Al, B, Be, C, F, H, Li, Mg, N, and O and their compounds. Thesis, Purdue University, Lafayette, IN, USA, 1972. xii + 143 pp.

Stewart:1972:XSC

[Ste72b] J. M. Stewart. The X-ray system of crystallographic programs for any computer having a PID-GIN FORTRAN compiler, version of June 1972. Technical report, University of Maryland, Computer Science Center, College Park, MD, USA, 1972. 283 pp.

Steinert:1973:FGC

[Ste73] Mark Steinert. FORTRAN G compatible Calcomp plot routines for the WATFIV compiler. Thesis (m.sc.), University of Manitoba, Winnipeg, Manitoba, Canada, 1973. [7] + 82 + [82] pp.

Steriadi:1974:MAL

[Ste74] Mariana Steriadi. Morphological aspects in the language FOR-TRAN IV. (Romanian). Studii şi cercetări Matematice, 26:755–761, 1974. ISSN 0039-4068, 0567-6401.

Steingraber:1975:FFS

[Ste75a] Jack Steingraber. Fortran Fundamentals: a Short Course. Hayden computer programming series. Howard W. Sams, Indianapolis, IN 46268, USA, October 1975. ISBN 0-8104-5860-8. 90 pp. LCCN QA76.73.F25S8.

US$7.25.  URL http://www.
cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0810458608.

Stevens:1975:CFL

[Ste75b] K. G. Stevens, Jr. CFD — A Fortran-like language for the Illiac IV. ACM SIGPLAN Notices, 10 (3):72–76, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steels:1976:FOP

[Ste76a] L. Steels. The FORLI. OLB package for list processing in FORTRAN 4: a user’s manual. Technical report, Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1976. 35 pp.

Stewart:1976:AHA

[Ste76b] G. W. Stewart. Algorithm 406 HQR3 and EXCHNG: Fortran subroutines for calculating and ordering and eigenvalues of a real upper Hessenberg matrix. ACM Transactions on Mathematical Software, 2:275–280, 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stewart:1976:AHE

[Ste76c] G. W. Stewart. Algorithm 506: HQR3 and EXCHNG: Fortran subroutines for calculating and ordering the eigenvalues of a real upper Hessenberg matrix [F2]. ACM Transactions on Mathematical Software, 2(3):275–280, September 1976. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Steele:1977:FAM

[Ste77]  Guy Lewis Steele, Jr.  Fast

arithmetic in MACLISP. Report A. I. MEMO 421, Massachusetts Institute of Technology, A. I. Lab., Cambridge, Massachusetts, September 1977.

Stewart:1978:SAF

[Ste78a] G. W. Stewart. SRRIT — A FOR-TRAN subroutine to calculate the dominant invariant subspaces of a real matrix. Technical Report TRR-514, University of Maryland, Department of Computer Science, College Park, MD, USA, 1978.

Stewart:1978:SFS

[Ste78b] G. W. Stewart. SRRIT — A FOR-TRAN subroutine to calculate the dominant invariant subspaces of a real matrix. Technical Report TRR-514, University of Maryland, Department of Computer Science, College Park, MD, USA, 1978.

Steuerwalt:1979:CEF

[Ste79] Michael Steuerwalt. Certification of “Algorithm 541: Efficient Fortran subprograms for the solution of separable elliptic partial differential equations [D3]”. ACM Transactions on Mathematical Software, 5(3):365– 371, September 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

SIT:1980:SFD

[Ste80] Stevens Institute of Technology.

SITGO-10/ 20: a FORTRAN ’77 debugging compiler: version 3A. Stevens Institute of Technology, Hoboken, NJ, USA, 1980. various pp.

Stiegler:1962:ACF

[Sti62] A. D. Stiegler. An algebraic compiler for the FORTRAN assembly program. Comm. ACM, 5 (11):545, November 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stipanuk:1972:GTS

[Sti72] Jeane Grayce Stipanuk. Generation of ten statement FOR-TRAN programs by using conditional probabilities. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1972. 88 pp.

Stock:1971:FBA

[Sto71] Karl F. Stock. FORTRAN in bibliothekarischer Anwendung. 3 Fallstudien mit ausfuhrlicher Programmdokumentation. [Wienerstr. 260: K. F. Stock], Graz, Austria, 1971. vi + 112 pp.

Stock:1976:IZZ

[Sto76] Karl F. Stock. Inventarisierung, Zugangsstatistik, Zuwachsverzeichnis im off-line-Verfahren: mit Dokumentation eines Programmsystems in FORTRAN. K. F. Stock, Graz, Austria, 1976. 94 pp.

Stodola:1980:PPL

[Sto80] Frank W. Stodola. The PLUS programming language. ACM SIGPLAN Notices, 15(1):146–155, January 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stroud:1978:TSP

[Str78] Richard Manson Stroud. Translation of the SPARKS preprocessor from FORTRAN to SPARKS. Thesis (m.s.), Kansas State University, Manhattan, KS, USA, 1978. 175 pp.

Stuart:1968:FP

[Stu68] Fredric Stuart. Fortran programming. John Wiley and Sons, New York, London, Sydney, 1968. ISBN 0-471-83477-7. xix + 353 pp. LCCN QA76.5 .S8.

Stuart:1970:FPa

[Stu70a] Fredric Stuart. Fortran programming. John Wiley and Sons, New York, London, Sydney, 1970. xix + 346 pp.

Stuart:1970:FPb

[Stu70b] Fredric Stuart. Fortran Programming. John Wiley and Sons, New York, London, Sydney, revised edition, June 1970. ISBN 0-471-83466-1. xix + 371 pp. LCCN QA76.5 .S8 1970. US$41.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0471834661.

Stuart:1971:WWF

[Stu71] Fredric Stuart. Watfor/Watfiv Fortran Programming. John Wiley and Sons, New York, London, Sydney, June 1971. ISBN 0-471-83471-8. xv + 239 pp. LCCN QA76.5 .S821 1971.

Sun:1973:ALF

[Sun73] Kuo Shan Sun. Assembler Language for Fortran, Cobol, and PL/I Programmers: IBM 370/ 360. Addison-Wesley, Reading, MA, USA, December 1973. ISBN 0-201-03954-0. ???? pp. LCCN ????

Shapiro:1974:IVS

[SW74] Stuart C. Shapiro and Douglas P.

Witmer. Interactive visual simulators for beginning programming students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 6(1): 11–14, February 1974. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 4th SIGCSE symposium on Computer science education.

Sundstrom:1975:EIP

[SW75] Arne Sundström and Lars Westerling. ENEP: an interactive program for solving one-dimensional continuum mechanical problems using a graphic terminal (abstract). ACM SIGNUM Newsletter, 10(4):38–39, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Szalajka:1979:ITP

[SW79] Walter S. Szalajka and Philip Walch. Integrated theory and practice — an approach to the first computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 11(1):45–48, February 1979. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 10th SIGCSE Symposium on Computer Science Education.

Soloway:1980:PPP

[SW80] Elliot M. Soloway and Beverly Woolf. Problems, plans, and programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 12(1): 16–24, February 1980. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 11th SIGCSE Symposium on Computer Science Education.

Swartz:1972:BFI

[Swa72] D. Swartz. B72-22 FORTRAN IV: a modem approach. IEEE Transactions on Computers, C-21(11):

1251, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672082.

Sweers:1967:SFI

[Swe67] H. E. Sweers. Some FORTRAN II programs for computer processing of oceanographic observations. Technical report 37, NATO Subcommittee on Oceanographic Research, Bergen, Norway, 1967. [3] + 105 pp.

Swingle:1964:ILP

[Swi64a] Wayne E. Swingle. Instructions for length-weight programs for IBM 1620 in FORTRAN: format (FOR-TRAN 1). Zoology-Entomology Department series. Fisheries 1, Agricultural Experiment Station, Auburn University, Auburn, AL, USA, 1964. 19 pp.

Swingle:1964:LWRa

[Swi64b] Wayne E. Swingle. Length–weight relationships I, IBM 1620, Fortran/Format. Transactions of the American Fisheries Society, 93(3): 318–319, July 1964. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1964:LWRb

[Swi64c] Wayne E. Swingle. Length–weight relationships II, IBM 1620, Fortran/Format. Transactions of the American Fisheries Society, 93(3): 319, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swinburne:1972:NFC

[Swi72] Swinburne College of Technology.Computer Studies Dept. Notes for FORTRAN courses. Swinburne College Press, Hawthorn, Victoria, Australia, second edition, 1972. 102 pp.

Seitz:1968:AAM

[SWL68] R. N. Seitz, L. H. Wood, and C. A. Ley. AMTRAN: automatic mathematical translation. Interactive Systems for Experimental Applied Mathematics, pages 44–65, 1968.

Simard:1977:AAT

[SYR77] Albert J. Simard, G. A. Young, and R. D. Redmond. AIRPRO: an air tanker productivity computer simulation model application the Fortran program (summary). Forest Fire Research Institute (Canada) Information report FF-X-6, Forest Fire Research Institute, Canadian Foresty Service, Dept. of Fisheries and the Environment, Ottawa, Ontario, Canada, 1977. 17 pp.

SEL:1973:SFIa

[Sys73a] Systems Engineering Laboratory. System 85/86 FORTRAN IV compiler reference manual. Technical report, Systems Engineering Laboratories, Ft. Lauderdale, FL, USA, July 1973. 152 pp.

SEL:1973:SFIb

[Sys73b] Systems Engineering Laboratory. System 85/86 FORTRAN IV compiler reference manual. Technical report, Systems Engineering Laboratories, Ft. Lauderdale, FL, USA, July 1973. 152 pp.

Starostenko:1980:MAF

[SZ80] Vitalii Ivanovich Starostenko and Antonina Nikolaevna Zavorotko. Metodika i Algol (Fortran)– programmy ustoichivogo resheniia obratnykh lineinykh i nelineinykh zadach gravimetrii. “Nauk. dumka,”, Kiev, USSR, 1980. 102 pp.

ICL:1966:FPT

[Tab66] International Computers and Ltd.

Programmed Learning Dept. Tabulators. 1900 Fortran: programmed text. International Computers and Tabulators, High Wycombe (Bucks.), 1966. various pp.

Tajiri:1965:UCS

[Taj65] Kazuo Tajiri. The use of COBOL subroutines in FORTRAN main programs. Comm. ACM, 8(4):233, April 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tamura:1966:FPC

[Tam66] T. (Taro) Tamura. Fortran programs for the Clebsch–Gordan and Racah coefficients and the ninej symbol. Technical report, Oak Ridge National Laboratory, Tennessee, Oak Ridge, TN, USA, 1966. 11 pp.

Tan:1978:EFP

[Tan78a] Chen-Bang Tan. The evolution of FORTRAN and the proposed ANSI standard. Thesis (m.s.), Department of Computer Science, Mississippi State University, Mississippi State, MS, USA, 1978. vii + 62 pp.

Tandem:1978:FPM

[Tan78b] Tandem Computers. FORTRAN programming manual. Tandem Computers Inc., Cupertino, CA, USA, 1978. various pp.

Tandem:1980:TFR

[Tan80a] Tandem Computers. Tandem FORTRAN 77 reference manual. Tandem Computers Inc., Cupertino, CA, USA, 1980. various pp.

Tanik:1980:SDM

[Tan80b] Murat M. Tanik. Software development monitoring graphs. ACM SIGSOFT Software Engineering Notes, 5(2):34–37, April 1980. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tanik:1980:TEP

[Tan80c] Murat M. Tanik. Two experiments on a program complexity perception by programmers. ACM SIGPLAN Notices, 15(9): 64–66, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1968:COF

[Tay68] R. Taylor. Correspondence: Orion FORTRAN compiler. The Computer Journal, 10(4):388, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://www3.oup.

co.uk/computer_journal/hdb/
Volume_10/Issue_04/100388.sgm.
abs.html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/388.

tif. See [TH64, ?].

Taylor:1976:FIO

[Tay76] R. Taylor. Fast input/output of variable-length arrays in FOR-TRAN IV. The Computer Journal, 19(3):273, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
19/Issue_02/tiff/273.tif.

Taylor:1977:TPB

[Tay77] Robert P. Taylor. Teaching programming to beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 9(1):88–92, February 1977. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Special issue for the Seventh Technical Symposium on Computer Science Education.

Taylor:1980:FIL

[Tay80] J. C. (John Charles) Taylor. A Fortran IV least squares computer program for the profile refinement of cubic powder diffraction patterns with cubic harmonic functions. Technical Report AAEC/E488, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1980. ISBN 0-642-59689-1. 19 pp.

Tryon:1965:UMB

[TB65] Robert Choate Tryon and Daniel Edgar Bailey. User’s manual of the BC TRY system of cluster and factor analysis; tape version for IBM 709, 7090, 7094 programs (FOR-TRAN II), July 31, 1965. ????, Berkeley, CA, USA, 1965. 157 pp.

Tremblay:1980:SFW

[TB80] Jean-Paul Tremblay and Richard B. Bunt. Structured FORTRAN WATFIV-S programming. Mc-Graw-Hill, New York, NY, USA, 1980. ISBN 0-07-065171-X. xiv + 419 pp. LCCN QA76.73.F25 .T74 1980.

Thomson:1970:IPM

[TC70] John Arthur Collingwood Thomson and J. R. Calaprice. IBM 1130 programs for multiple discrimination analysis of X-ray spectroscopy data (FORTRAN). Technical report 212, Fisheries Research Board of Canada, Biological Station, Nanaimo, BC, Canada, 1970. 41 pp.

Tweedale:1975:DUI

[TC75] A. Tweedale and C. Chan. On the design of a user interface to large FORTRAN subroutine packages under IBM 360/370 OS. ACM SIGNUM Newsletter, 10(1):15–17, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See note [Boy75].

Tryon:1978:SLF

[TD78] Peter V. Tryon and Janet R. Donaldson. STATLIB: a library of FORTRAN subroutines for statistical analysis of experimental data. Technical report, National Bureau of Standards, Center for Applied Mathematics Statistical Engineering Laboratory, Boulder, CO, USA, 1978. various pp.

Teague:1972:CPF

[Tea72] Robert Teague. Computing problems for Fortran solution. Canfield Press, San Francisco, CA, USA, 1972. ISBN 0-06-388672-3. ix + 245 pp. LCCN QA76.73.F25T4.

Teague:1974:FDA

[Tea74] Robert Teague. Fortran, a discovery approach. Canfield Press, San Francisco, CA, USA, 1974. ISBN 0-06-388670-2 (invalid ISBN checksum). xiv + 270 pp. LCCN QA76.73.F25 T42.

Berlin:1972:KAS

[Tec72] Technische Universität Berlin.

Kursmaterialien zur Automatisierung: Spezielle Themen der Fortran- und Algol-Programmierung. Technical report, Technische Universität Berlin, Berlin, Germany, 1972.

Tellier:1980:SCP

[Tel80] Gilles Tellier. Solutions a certains problèmes de FORTRAN structuré avec WATFIV. École polytechnique de Montréal, Montréal, PQ, Canada, 1980. 130 pp.

Thacker:1979:R

[TG79a] W. I. Thacker and G. W.

Gorsline. Response. ACM SIGMICRO Newsletter, 10(2):

29, June 1979. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/
10.1145/1218003.1218005. See

[TG79b, Kor79].

Thacker:1979:MPR

[TG79b] William I. Thacker and G. W. Gorsline. Micro programming rational arithmetic operations. ACM SIGMICRO Newsletter, 10(1):10– 13, March 1979. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/
10.1145/1217236.1217238. See

critical comments [Kor79] and response [TG79a].

Tennent:1962:FPH

[TH62] Richard C. Tennent and Lester Abram Herr. A FORTRAN program for the hydraulic design of circular culverts. Hydraulic engineering circular 7, U.S. Dept. of Commerce, Bureau of Public Roads, Washington, DC, USA, 1962. 38 pp.

Taylor:1964:FSO

[TH64] R. Taylor and D. A. Harragan. The FORTRAN system for Orion. The Computer Journal, 7(2):114–116, July 1964. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
07/Issue_02/070114.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/114.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
02/tiff/115.tif;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/

116.tif. See correspondence [?, Tay68].

Tharp:1977:CCF

[Tha77] A. L. Tharp. A comparison of COBOL, FORTRAN, PL-I and SPITBOL. Computer Languages, 2(4):171–178, ???? 1977. CODEN COLADA. ISSN 0096-0551.

Therrien:1968:PFG

[The68] J. J. Therrien. Plain FORTRAN; a guide to compatibility in computer programming. Report ????, Inland Waters Branch, Dept. of Energy, Mines and Resources, Ottawa, Ontario, Canada, 1968. iv + 19 pp.

Thompson:1965:FPC

[Tho65] D. Thompson. A Fortran program to calculate the flow field and performance of an axially symmetric deLaval nozzle. NASA technical note NASA TN D-2579, National Aeronautics and Space Administration, Scientific and Technical Information Division, Washington, DC, USA, 1965. viii + 150 pp.

Thompson:1966:TFT

[Tho66] Bruce (Bruce L.) Thompson. Fortrain: Fortran trainer. Control Data Corp., Minneapolis, MN, USA, teacher’s, with student workbook edition, 1966. iii + 177 pp.

Thongurai:1968:DNS

[Tho68] Kamthorn Thongurai. A digital network simulator using FOR-TRAN. Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1968. [7] + 52 pp.

Thomas:1971:PIG

[Tho71] Paul A. V. Thomas. A programmed instruction guide for FORTRAN IV with WATFOR and WATFIV. Technical report, Electrical Engineering Dept., University of Windsor, Windsor Ontario, Canada, 1971. 85 pp.

Thomas:1972:SAA

[Tho72a] Donald G. Thomas. Statistical algorithms: Algorithm AS 50: Tests of fit for a one-hit vs. two-hit curve. Applied Statistics, 21(1):103–112, March 1972. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/50.

Thomas:1972:IFU

[Tho72b] Paul A. V. Thomas. An introduction to FORTRAN using WAT-FOR or WATFIV: a programmed instruction text. Technical report, Electrical Engineering Dept., University of Windsor, Windsor Ontario, Canada, 1972. 169 pp.

Thorin:1972:ECP

[Tho72c] Marc Thorin. Exercises commentes de programmation en langage Fortran, a l’usage des debutants. Masson, Masson, France, 1972. vii + 62 pp.

Thorin:1978:ECF

[Tho78] Marc Thorin. Exercices commentés de Fortran: Syntaxe élementaire, approfondissement et style. Masson, Masson, France, 2e rev. et augm. edition, 1978. 109 pp.

Throssell:1979:PCP

[Thr79] Terrence Throssell. PARMS:

a computer program to modify FINSYS-2 TABLE-2 and other FORTRAN programs. USDA Forest Service research note INT 273, Intermountain Forest and Range Experiment Station, Ogden, UT, USA, 1979. 10 pp.

Taylor:1972:SIU

[TI72] Lawrence W. Taylor and Kenneth W. Iliff. Systems identification using a modified Newton– Raphson method: A Fortran program. NASA technical note D-6734, U.S. National Aeronautics and Space Administration, Washington, DC, USA, 1972. 72 pp.

Tippett:1976:ARB

[Tip76] Peter Van Rensselaer Tippett.

Analysis of the relationships between learning style profiles and the learning of FORTRAN basic skills. Thesis (m.b.a.), San Diego State University, San Diego, CA, USA, 1976. v + 96 pp.

Tjugen:1968:IET

[Tju68] Karsten Tjugen. Innforing i EDB Tekniske beregninger, FORTRAN programmering. Universitetsforlaget, Oslo, Norway, 1968. various pp.

Tobey:1965:BMF

[Tob65] George D. Tobey. Bit manipulation in Fortran language. Comm.

ACM, 8(6):378, June 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tokko:1968:SMP

[Tok68] Mok Tokko. The student manual for programming in FORTRAN IV. Thesis (m.a.), Kansas State Teachers College of Emporia, Emporia, KS, USA, 1968. 111 pp.

Tomasso:1971:PCF

[Tom71] Carlo Tomasso. A practical course in Fortran IV computer programming. Foulsham, London, UK, 1971. ISBN 0-572-00766-3 (paperback). 126 pp.

Torrey:1969:AFD

[Tor69] Howard Levi Torrey. An all FOR-TRAN digital simulation program with an analog structure. Thesis (Ph.D.), Iowa State University, Ames, IA, USA, 1969. 115 pp.

Tou:1970:SEP

[Tou70] Julius T. Tou, editor. Software engineering: proceedings of the Third Computer and Information Sciences Symposium, Miami Beach, FL, USA, 1969. Academic Press, New York, NY, USA, 1970. LCCN QA76.5 .C61255 1969.

Torsun:1977:NIF

[TR77] I. S. Torsun and S. K. Robinson. Non-‘Interpretive’ FOR-TRAN input/output. Software— Practice and Experience, 7(2):205– 213, March/April 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Trivisonno:1973:FIC

[Tri73] Roger J. Trivisonno. Fortran IV computer program for calculating critical speeds of rotating shafts. NASA technical note NASA TN D-7385, National Aeronautics and Space Administration, Washington, DC, USA, 1973. iii + 52 pp. For sale by the National Technical Information Service.

Tripathi:1979:RFI

[Tri79] V. S. Tripathi. RANTEST — a Fortran IV program for testing randomness of uniform pseudorandom numbers. Computers and Geosciences, 5(??):251– 268, ???? 1979. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Trotter:1964:FCP

[Tro64] Charles Earl Trotter. A Fortran computer program designed to identify the physical facilities for public secondary school instructional materials centers. Thesis, University of Tennessee, Knoxville, TN, USA, 1964. 162 pp.

Trower:1966:FPC

[Tro66a] W. Peter Trower. A FOR-TRAN program for calculating kinematic and dynamic quantities of particle interactions and decays. Journal of Computational Physics, 1(1):144–145, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999166900167.

Trower:1966:FSC

[Tro66b] W. Peter Trower. A FOR-TRAN subroutine for calculating the range-energy relation of charged particles in chemical elements. Journal of Computational Physics, 1(1):145–147, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0021999166900179.

TRW:1973:UTCa

[TRW73a] TRW Inc. Transportation and Environmental Operations, Houston, TX, USA. Urban traffic control system FORTRAN IV software documentation, 1973. [384] pp.

TRW:1973:UTCb

[TRW73b] TRW Inc. Transportation and Environmental Operations, Houston, TX, USA. Urban traffic control system FORTRAN IV software interface manual, 1973. [256] pp.

TRW:1973:UTCc

[TRW73c] TRW Inc.Transportation and Environmental Operations. Urban traffic control system: FOR-TRAN IV software documentation. Federal Highway Administration, Washington, DC. USA, 1973. ca. 400 pp. Available through National Technical Information Service.

TRW:1973:UTCd

[TRW73d] TRW Inc.Transportation and Environmental Operations. Urban traffic control system: FOR-TRAN IV software interface manual. Federal Highway Administration, Washington, DC. USA, 1973. ca. 250 pp. Available through National Technical Information Service.

Taylor:1973:FPG

[TS73] Judy Taylor and Paul Spencer. A Fortran program for grade crossing collision studies using computer graphics: final report. Report — Federal Railroad Administration FRA-ORD and D-74-6, U. S. Dept. of Transportation, Federal Railroad Administration Office of Research, Development and Demonstrations: Available through the National Technical Information Service, Washington, DC. USA, 1973. ca. 160 pp.

Taylor:1976:DSP

[TS76] Fred J. Taylor and Steve L. Smith.

Digital Signal Processing in Fortran. D. C. Heath and Company, Lexington, MA, USA, June 1976. ISBN 0-669-00330-1. xii + 403 (or xi + 402??) pp. LCCN TK5102.5 .T34. US$33.50. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0669003301.

Turner:1980:FFP

[TT80] Danny W. Turner and F. Eugene Tidmore. FACES — a FORTRAN program for generating Chernofftype faces on a line printer. The American Statistician, 34(3):187, August 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        jstor.org/stable/2683888.

Tucker:1977:PL

[Tuc77] Allen B. Tucker.Jr. Programming Languages. McGraw-Hill, New York, NY, USA, 1977. ISBN 0-07-065415-8. xv + 439 pp. LCCN QA76.7 T8 1977.

Tuggle:1975:HPC

[Tug75] Francis D. Tuggle. How to program a computer, using Fortran IV. Grid Inc., Columbus, OH, USA, 1975. ISBN 0-88244-082-9. 173 pp. LCCN QA76.73.F25T83.

Turner:1968:FIP

[Tur68] A. Keith Turner. Fortran IV programs to develop contour maps of 3-dimensional data. Project C-36-72A. File no. 1-6-1; no. 6 Joint Highway Research Project (Ind.) (series); JHRP-68-6., Engineering Experiment Station, Joint Highway Research Project, Purdue University, West Lafayette, IN, USA, 1968. 86 pp.

Turner:1969:GSFa

[Tur69a] A. Keith Turner. The GCARS system Fortran IV programmers manual. Project C-36-72A. File 1-6-1; no. 25-27 Joint Highway Research Project (Ind.) (Series); no. 69-25 — 69- 7., Engineering Experiment Station, Joint Highway Research Project, Purdue University, West Lafayette, IN, USA, 1969. ???? pp.

Turner:1969:GSFb

[Tur69b] A. Keith Turner. The GCARS system Fortran IV users manual.

Project C-36-72A. File 1-6-1; no. 24 Joint Highway Research Project (Ind.) (Series); no. 69-24., Engineering Experiment Station, Joint Highway Research Project, Purdue University, Lafayette, IN, USA, 1969. iii + 30 pp.

Turner:1973:IFI

[Tur73] Robert L. Turner. Introduction to FORTRAN IV. Idaho Research Foundation, Moscow, ID, USA, 1973. ix + 237 pp.

Tomasso:1971:CPP

[TW71] C. Tomasso and Ian D. Watt.

Computer programming: a practical course in Fortran IV. W. Foulsham and Co., London, New York, Toronto, etc., 1971. ISBN 0-572-00766-3. 126 pp.

Tymshare:1968:FI

[Tym68] Tymshare, Inc. FORTRAN IV.

Tymshare, Los Altos, CA, USA, revision 2 edition, 1968. 52 pp.

Tymshare:1970:SFS

[Tym70] Tymshare, Inc., Palo Alto, CA, USA. SUPER FORTRAN: a superset of H. Level FORTRAN IV, 1970. x + 180 pp.

Ubell:1976:FSC

[Ube76] Michael Ubell. FORTED, a syntax checking Fortran editor for the Princeton UNIX Fortran system: research project. Thesis (m.s. in electrical engineering), University of California, Berkeley, Berkeley, CA, USA, May 1976. various pp.

Ulery:1974:SLO

[UK74] Dana L. Ulery and H. M. Khalil. Survey of language-oriented systems for numerical linear algebra. The Computer Journal, 17(1):82– 88, February 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Unger:1969:BRD

[Ung69] C. Unger. Book review: D. D. McCracken FORTRAN with Engineering Applications. Computing, 4(1):88, 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

USOEP:1968:EII

[Uni68a] United States.Office of Emergency Planning. EXEC-8 input-output interface for FORTRAN V. NREC technical manual — Office of Emergency Planning 206, The Office, Washington, DC. USA, 1968. 33 pp.

UPESE:1968:EDL

[Uni68b] Université de Paris.Ecole supérieure d’électricité. Elements du langage FORTRAN IV 1130. Technical Report 2136, Ecole Supérieure d’Electricité, Malakoff, France, 1968. 80 pp.

UCS:1969:TFL

[Uni69a] United Computing Systems, Inc., Kansas City, MO, USA. Timesharing FORTRAN language: reference manual, 1969. viii + 128 pp.

UCBCC:1969:CFG

[Uni69b] University California, Berkeley.

Computer Center, Berkeley, CA, USA. CAL 6400 Fortran guide, preliminary edition, revision 2 edition, 1969. various pp.

UMECD:1969:WLR

[Uni69c] University of Missouri.Engineering Computer Dept. WATFOR library reference manual: FOR-TRAN IV supplied functions and WATFOR.WATLIB. Lucas Bros., Columbia, MO, USA, 1969. 230 pp.

Univac:1970:FF

[Uni70] Univac. Fundamentals of FOR-TRAN. Sperry Rand, New York, NY, USA, 1970. various pp.

USBC:1971:CCF

[Uni71] United States.Bureau of the Census. Cenfor: Census Fortran manual. Dept. of Commerce, Bureau of the Census, Washington, DC. USA, 1971. various pp.

UN:1972:FL

[Uni72] University of Natal. Fortran language. Technical report, University Computer Centre, Natal, South Africa, 1972. 71 pp.

UCBCC:1973:CRF

[Uni73] University California, Berkeley.

Computer Center, Berkeley, CA, USA. CAL RUN Fortran guide, 1973. various pp.

UCBCC:1974:CRF

[Uni74a] University California, Berkeley.Computer Center. CAL RUN Fortran guide. Technical report, University of California, Computer Center, Berkeley, CA, USA, 1974. various pp.

UTEPCC:1974:FCP

[Uni74b] University of Texas at El Paso.Computer Center. FORTRAN: callable plot subroutines. Reference manual 74-5, University of Texas at El Paso, El Paso, TX, USA, 1974. 20 pp.

USNDADPESO:1975:TFB

[Uni75a] United States.Navy Dept.Automatic Data Processing Equipment and Selection Office. Software Development Division. Transferability of FORTRAN benchmarks. Technical Report AD/A-039 741, Department of the Navy, Automatic Data Processing Equipment Selection Office, Software Development Division, Washington, DC. USA, 1975. 11 pp. Reproduced by National Technical Information Service.

UMCC:1975:EFF

[Uni75b] Université de Montréal.Centre de calcul. Elements de Fortran II fondamental: manuel de reférence: Ordinateur CDC 3100. Université de Montréal, Centre de calcul, Montréal, PQ, Canada, 1975. 51 + A-2 + 12 pp.

UTACC:1977:FDA

[Uni77] University of Texas at Austin. Computation Center, Austin, TX, USA. Fortran debugging aids, UT-CDC 6000, 1977. 398 pp.

USMA:1978:CFP

[Uni78] United States Military Academy.Academic Automation Division. Cobol and Fortran programmer course (CBT [i.e. CBL-FTN]): for use at the United States Military Academy. Academic Automation Division, Office of the Dean of the Academic Board, United States Military Academy, West Point, NY, USA, 1978. A14–1–A14–51 + A15–1–A15–6 pp.

UWDCS:1979:VMC

[Uni79] University of Waterloo.Dept.of Computing Services. VM/370 CMS Fortran user’s guide. Technical report, Dept. of Computing Services, University of Waterloo, Waterloo, Ontario, Canada, 1979. ii + 28 pp.

USNBS:1980:FCS

[Uni80a] United States.National Bureau of Standards. FORTRAN: category, software standard; subcategory, programming language. FIPS pub 69, National Bureau of Standards, Washington, DC, USA, 1980. 4 pp. For sale by the National Technical Information Service.

UWDCS:1980:VCF

[Uni80b] University of Waterloo. Dept. of Computing Services, Waterloo, Ontario, Canada. VM/ 370 CMS Fortran user’s guide, 1980. iii + 35 pp.

UWMACC:1980:DDS

[Uni80c] University of Wisconsin — Madison.Academic Computing Center.

DULPDX/DULPLX, SIMPDX/ SIMPLX: routines for ASCII Fortran: reference manual for 1100 series computers. Operations research series 1505, Academic Computing Center, University of Wisconsin — Madison, Madison, WI, USA, 1980. various pp.

UMCC:1970:MMT

[Uni0 ] University of Michigan Computing Center. MTS: the Michigan Terminal System. Technical report, University of Michigan Computing Center, Ann Arbor, MI, USA, 1970 (??). various pp.

Upchurch:1972:MMR

[Upc72] James Kimble Upchurch. MINI-FOR: minicomputer real-time FORTRAN programming through simulation. Thesis (m.s. — computer science), University of Arizona, Tucson, AZ, USA, 1972. 122 pp.

USAEC:1961:TRN

[U.S61] U.S.Atomic Energy Commission. [Technical reports on natural radiation, fallout shelters, blast effect, computer random number generators, radioactive tracers, nuclear reactors, irradiated foods, nuclear powered vehicles, stratospheric circulation, FORTRAN, nuclear detectors, switching circuits, measurement of airplane stability, and fire testing of radioisotope containers]. U. S. Atomic Energy Commission, Washington, DC, USA, 1961.

MIL-STD-1753:1978:MSF

[U.S78] U.S. Department of Defence. Military Standard 1753: FORTRAN, DoD Supplement To American National Standard X3.9-1978, November 9, 1978, 1978. Available on the World-Wide Web

at http://observer.gsfc.nasa.
gov/iteams/Standards/milspec.
html.

Villaveces-Atuesta:1962:PSF

[VAB62] Alvero Villaveces-Atuesta and N. W. (Norman William) Bryan. Plot subroutines for 1620 FOR-TRAN. Publication 156; Technical report T62-4, Massachusetts Institute of Technology, School of Engineering Dept. of Civil Engineering, Civil Engineering Systems Laboratory, Cambridge, MA, USA, 1962. 5 + [6] + [1] + [4] + [1] pp.

VandePol:1966:FPO

[Van66] David Wilfred Van de Pol. A FORTRAN program to optimize air-cycle cooling systems. Thesis (m.s.), Massachusetts Institute of Technology. Dept. of Mechanical Engineering, Cambridge, MA, USA, 1966. 136 + 5 pp.

VanKlink:1968:FIF

[Van68a] K. H. Van Klink. Fortran II–Fortran IV conversion of Methuselah. Technical Report AAEC/TM455, Australian Atomic Energy Commission, Research Establishment, Lucas Heights, NSW, Australia, 1968. 26 + 19 pp.

Vandome:1968:BRB

[Van68b] P. Vandome. Book review: Fortran Programs for Economists, by Lucy Joan Slater. Journal of the Royal Statistical Society. Series A (General), 131(3):441–442, ???? 1968. URL http://www.

jstor.org/stable/2343539.

VanHoa:1973:ACP

[Van73a] Tran Van Hoa. Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 16(3):273–276, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/
273.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/274.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
16/Issue_03/tiff/275.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_
03/tiff/276.tif. See note [?, ?].

Vanderplaats:1973:CFP

[Van73b] Garret N. Vanderplaats. CON-MIN, a FORTRAN program for constrained function minimization user’s manual. NASA technical memorandum X-62282, Ames Research Center and U.S. Army Air Mobility R and D Laboratory, Moffett Field, CA, USA, 1973. 60 pp. Reproduced by National Technical Information Service.

Varner:1977:FPD

[Var77] Ruth N. Varner. FORTRAN program to determine length of gage blocks using single wavelength interferometry. NBS technical note 956, Dept. of Commerce, National Bureau of Standards, Institute for Basic Standards, Washington, DC. USA, 1977. [2] + 51 + [1] pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Vasan:1972:MDC

[Vas72] Malore Sreeni Vasan. A method of developing computer programs in FORTRAN IV for simulation of feedback systems with applications. Thesis (m.s. in engineering management), University of Dayton, Dayton, OH, USA, 1972. 187 pp.

Vegelius:1971:SFP

[Veg71] Jan Vegelius. SIEGEL: a FOR-TRAN program for nonparametrical methods. Report 109, Dept. of Psychology, University of Uppsala, Uppsala, Sweden, 1971. 50. pp.

Vegelius:1974:CFP

[Veg74] Jan Vegelius. CORALL: a FOR-TRAN program for correlation measures. Report 147, Dept. of Psychology, University of Uppsala, Uppsala, Sweden, 1974. 42 + 8. pp.

Veinott:1966:PDT

[Vei66] Cyril G. Veinott. Programming decision tables in FORTRAN, COBOL or ALGOL. Comm. ACM, 9(1):31–35, January 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Veldman:1967:FPB

[Vel67] Donald J. Veldman. Fortran programming for the behavioral sciences. Holt, Rinehart, and Winston, New York, NY, USA, 1967. x + 406 pp.

Verzuh:1959:FMA

[Ver59] Frank M. Verzuh. FORTRAN modifications and additional format features. Memorandum CC-147, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1959. 3 pp.

Veronis:1965:NUD

[Ver65] George Veronis. A note on the use of a digital computer for doing tedious algebra and programming. Comm. ACM, 8(10):625–626, October 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vigue:1977:CPF

[VG77] F. Vigue and J. Gunther. Cours de programmation Fortran IV. IREM (Institute de recherche sur l’enseignement des mathématiques), Université Louis Pasteur, Strasbourg, France, 1977. 138 pp.

Vignes:1969:TPP

[VHP69] J. Vignes, B. Hallopeau, and M. La Porte. Theorie et pratique de la programmation Fortran. Publications de l’Institut français du petrole. Societé des Editions Technic, Paris, France, 1969. 220 pp.

Vickers:1964:RFC

[Vic64] Frank Dow Vickers. RW — 300 FORTRAN, a compiling system for the RW-300 digital computer. Thesis, University of Florida, Gainesville, FL, USA, 1964. v + 157 pp.

Vichova:1970:FPS

[Vic70a] Ellen Vichova. Fortran program for the solution of polynomial equations with real coefficients by the method of adaptive strategy. Technical Report PURE, no. 44, Ceskoslovenska akademie ved. Ustav radiotechnicky a elektroniki, Praha, Czechoslovakia, 1970. 22 pp.

Vickers:1970:FIM

[Vic70b] Frank D. Vickers. Fortran IV, a modern approach. Holt, Rinehart, and Winston, New York, NY, USA, 1970. ISBN 0-03-083060-5. xvi + 207 pp. LCCN QA76.5 .V49.

Vickers:1973:FIE

[Vic73] Frank D. Vickers. Fortran IV: un enfoque moderno. Editorial Diana, Mexico, DF, Mexico, 1973. 221 pp.

Vickers:1977:FIE

[Vic77] Frank D. Vickers. Fortran IV: un enfoque moderno. Editorial Diana, Mexico, DF, Mexico, 1977. 221 pp.

Vickers:1978:FIMa

[Vic78a] Frank D. Vickers. Fortran IV: a Modern Approach. Kendall/Hunt Pub., Dubuque, IA, USA, second edition, June 1978. ISBN 0-8403-1829-4. xii + 235 pp. LCCN QA76.73.F25V48 1978. US$10.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0840318294.

Vickers:1978:FIMb

[Vic78b] Frank D. Vickers. Fortran IV: a modern approach. Kendall/Hunt Pub., Dubuque, IA, USA, second edition, 1978. ISBN 0-8403-1829-4. xii + 235 pp. LCCN QA76.73.F25V48 1978.

Vinturella:1975:IF

[Vin75] John B. Vinturella. Introduction to FORTRAN. Petrocelli/Charter, New York, NY, USA, 1975. ISBN 0-88405-301-6. xii + 266 pp. LCCN QA76.73.F25V54 1975.

Vignes:1972:TPP

[VL72] Jean Vignes and Michel Laporte. Theorie et pratique de la programmation Fortran. Collection langages et algorithmes de l’informatique. Editions Technip, Paris, France, 3e edition, 1972. 220 pp.

vonMeerwall:1975:FCA

[vM75] E. D. von Meerwall. A Fortran code for automatic spectrum analysis on medium-scale computers. Computer Physics Communications, 9(6):351–359, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465575900156.

vonMeerwall:1976:SFP

[vM76] E. D. von Meerwall. A simple Fortran program to interpret cubic X-ray powder diffraction data. Computer Physics Communications, 11(3):331–337, June/ August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465576900205.

vonMeerwall:1977:FPS

[vM77] E. D. von Meerwall. A Fortran program to simulate quadrupoledistorted NMR powder patterns. Computer Physics Communications, 13(2):107–115, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465577900364.

vonMeerwall:1978:FPR

[vM78a] E. D. von Meerwall. A Fortran program for routine analysis of magnetic susceptibility data. Computer Physics Communications, 15(3–4):237–245, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0010465578900942.

vonMeerwall:1978:FPC

[vM78b] E. D. von Meerwall. A Fortran program to collect histograms of microscopic scalar interactions. Computer Physics Communications, 13 (5–6):421–427, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465578900395.

vonMeerwall:1979:FPP

[vM79] E. D. von Meerwall. A FOR-TRAN program to perform signal averaging, multichannel scaling, and pulse-height analysis. Computer Physics Communications, 18(3):417–426, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465579900110.

Schenck:1963:FMH

[vNS63] Hilbert van Nydeck Schenck. Fortran methods in heat flow. Ronald Press Co., New York, NY, USA, 1963. viii + 289 pp.

vanOosterom:1978:THT

[vO78] A. van Oosterom. Triangulating the human torso. The Computer Journal, 21(3):253–258, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/
253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
21/Issue_03/tiff/255.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_
03/tiff/256.tif;   http://www3.

oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/
257.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/258.
tif.

Vowels:1974:AFI

[Vow74] R. A. Vowels. ALGOL 60 and FORTRAN IV. J. Wiley and Sons Australasia, Sydney, Australia; New York, NY, USA, 1974. ISBN 0-471-91192-5. 173 pp. LCCN QA76.73.A24 V68.

Vowels:1977:AFI

[Vow77] R. A. Vowels. ALGOL 60 and FORTRAN IV. John Wiley and Sons, New York, London, Sydney, 1977. ISBN 0-471-91192-5 (paperback). x + 173 pp. LCCN QA76.73.A24 V68.

Vowels:1978:FIA

[Vow78] R. A. Vowels. FORTRAN IV y ALGOL 60. Editorial Limusa, Mexico, DF, Mexico, 1978. 202 pp.

Vujisic:1975:FPR

[VP75] B. R. Vujisić and D. S. Pesić. A FORTRAN program for rotational analysis of the spectra of diatomic molecules. Journal of Computational Physics, 19(4):432–434, December 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/002199917590073X.

Valiaho:1976:PSR

[VP76] Hannu Valiaho and Timo Pekkonen. A procedure for stepwise regression analysis: (with a program in FORTRAN V). Akademie-Verlag, Berlin, Germany, 1976. 90 + [1] pp.

Vallance:1980:BRB

[VP80a] D. M. Vallance and Peter H. Prowse. Book reviews: Programming in Standard Fortran 77, by A. Balfour and D. Marwick, 1979; Issues in Data Base Management, by Herbert Weber and Anthony I. Wasserman, 1979. The Computer Journal, 23(2):152, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
23/Issue_02/tiff/152.tif.

Vallance:1980:BRP

[VP80b] D. M. Vallance and Peter H. Prowse. Book reviews: Programming in Standard Fortran 77, by A. Balfour and D. Marwick, 1979; Issues in Data Base Management, by Herbert Weber and Anthony I. Wasserman, 1979. The Computer Journal, 23(2):152, May 1980. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
23/Issue_02/tiff/152.tif.

Voigtlander:1967:AGP

[VR67] Clyde W. Voigtlander and Alan C. Roochvarg. Age and growth program, Control Data 3600, Fortran 63–Fortran IV. Transactions of the American Fisheries Society, 96(3): 364–366, July 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Verma:1980:MPF

[VS80] S. B. Verma and Maithili Sharan.

Multiple precision floating-point computation in FORTRAN. Software—Practice and Experience, 10 (3):163–173, March 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanNiekerk:1966:FIC

[VV66] Helena Van Niekerk and J. W. Von Backstrom. A Fortran IV computer code for calculation of CIPW norms and niggli values. Technical Report PEL 126, Atomic Energy Board, Pelindaba, South Africa, 1966. 24 pp.

Wagner:1970:FAI

[Wag70] Robert A. Wagner. Finiteness assumptions and intellectual isolation of computer scientists. Comm. ACM, 13(12):759–760, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagener:1975:SFP

[Wag75] J. L. Wagener. Structured FOR-TRAN programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 7(1):206–211, February 1975. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 5th SIGCSE symposium on Computer science education.

Wagener:1980:FPP

[Wag80a] Jerrold L. Wagener. Fortran 77: Principles of Programming.

John Wiley and Sons, New York, London, Sydney, January 1980. ISBN 0-471-04474-1. xi + 370 pp. LCCN QA76.73 .F25 W34 1980. US$37.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0471044741.

Wagner:1980:FPL

[Wag80b] Neal R. Wagner. A FORTRAN preprocessor for the large program environment. ACM SIG-PLAN Notices, 15(12):92–103, December 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wahlstedt:1968:FIP

[Wah68] Warren C. Wahlstedt. Fortran IV program for computation and display of principal components. Computer contribution 21, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1968. 27 pp.

Waite:1974:O

[Wai74] W. M. Waite. Optimization. In Bauer and Eickel [BE74], pages 549–602. LCCN QA76.6 .C6281.

Walton:1963:DFA

[Wal63] Eric Lee Walton. Differentiation of Fortran arithmetic expressions on a digital computer. Thesis (m.a.), University of Florida, Gainesville, FL, USA, 1963. iv + 83 pp.

Waller:1968:TFI

[Wal68] Thomas R. Waller. Two FOR-TRAN II programs for the univariate and bivariate analysis of morphometric data. United States. National Museum Bulletin; 285 Bulletin (United States National Museum); 285. Smithsonian Institution Press, Washington, DC. USA, 1968. v + 55 pp. For sale by the Supt. of Docs. U.S. Govt. Print. Off.

Walton:1970:DEF

[Wal70] James Leroy Walton. The design of an educational FORTRAN processor. Thesis (m.s.), Mississippi State University, Mississippi State, MS, USA, 1970. ix + 165 pp.

Walker:1972:ICS

[Wal72] Terry M. Walker. Introduction to Computer Science: Fortran Language Programming. Allyn and Bacon, Needham Heights, MA, USA, June 1972. ISBN 0-205-03578-7. xi + 244 pp. LCCN ???? US$13.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0205035787.

Walker:1975:FFP

[Wal75] Terry M. Walker. Fundamentals of Fortran Programming: With Watfor/Watfiv. Allyn and Bacon, Needham Heights, MA, USA, August 1975. ISBN 0-205-04885-4. ix + 325 pp. LCCN QA76.73.F25W34. US$11.95. URL http://www.

cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0205048854.

Walker:1980:PCSa

[Wal80a] Henry M. Walker. Problems for computer solutions using FOR-TRAN. Winthrop computer systems series. Winthrop Publishers, Cambridge, MA, USA, 1980. ISBN 0-87626-654-5. x + 203 pp. LCCN QA43 .W34.

Walker:1980:PCSb

[Wal80b] Henry M. Walker. Problems for Computer Solutions Using Fortran. Little, Brown and Co., Boston, MA, USA, June 1980. ISBN 0-316-91834-2. ???? pp. LCCN ???? US$19.75. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0316918342.

Wang:1978:EPF

[Wan78] J. Y. Wang. The evaluation of periodic functions with large input arguments. ACM SIGNUM Newsletter, 13(4):7–8, December 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Warner:1969:FIP

[War69] Jeffrey Warner. Fortran IV program for construction of P1 diagrams with the Univac 1108 computer. Computer contribution 33, Kansas Geological Survey, Lawrence, KS, USA, 1969. 38 pp.

Warner:1975:PDG

[War75] D. D. Warner. A partial derivative generator. Computing Science Technical Report 28, Bell Telephone Laboratories, Murray Hill, NJ, USA, 1975.

Warner:1979:DDI

[War79] James R. Warner. DIGRAF: Device independent graphics from Fortran. Technical report, Graphics Development Group, University Computing Center University of Colorado, Boulder, CO, USA, 1979. various pp.

Watters:1968:FPC

[Wat68] John Watters. Fortran programming; a complete course in writing Fortran programs. Heinemann, London, UK, 1968. ISBN 0-435-77800-5. [5] + 354 pp. LCCN QA 76.5 W38 General Coll.

Watkins:1973:MTD

[Wat73a] S. L. Watkins. Masked threedimensional plot program with rotations. Comm. ACM, 17(??):520– 523, ?? 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watkins:1973:MTP

[Wat73b] S. L. Watkins. Masked threedimensional plot program with rotations. Comm. ACM, 17:520–523, 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watts:1975:FIP

[Wat75] Raymond D. Watts. A Fortran IV program for analytic continuation of VLF electromagnetic data. Open-file report 75-159, U.S. Dept. of the Interior, Geological Survey, Reston, VA, USA, 1975. 28 pp.

Waters:1976:SUM

[Wat76] Richard C. Waters. A system for understanding mathematical FORTRAN programs. A. I. Memo 368, Massachusetts Institute of Technology, A. I. Lab., Cambridge, MA, USA, August 1976. 78 pp.

Weldon:1965:ABN

[WB65] Roger J. Weldon and Robert L. Baker. Applications of binary numbers in computer routines. Comm. ACM, 8(5):315–318, May 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitehouse:1971:ICP

[WB71] Leonard George Whitehouse and C. J. Bellamy. An introduction to computer programming. Monash University Computer Centre, Melbourne, Victoria, Australia, 1971. ISBN 0-909549-00-1. various pp. LCCN QA76.73.F25W46.

Williams:1968:FPC

[WCT68] I. R. Williams, M. Craig, Jr., and C. L. Thompson. A Fortran program for calculating the solid angle subtended by one circular disk at another. Journal of Computational Physics, 2(3):332–333, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999168900612.

Whitten:1975:MCI

[WD75] D. E. Whitten and P. A. D. Demaine. A machine and configuration independent Fortran: Portable Fortran PFortran. IEEE Transactions on Software Engineering, SE-1(1): 111–124, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312825.

Wichmann:1979:PCG

[WD79] B. A. Wichmann and J. Du Croz. Program to calculate the GAMM measure. The Computer Journal, 22(4):317–322, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilcox:1976:DIT

[WDT76] Thomas R. Wilcox, Alan M. Davis, and Michael H. Tindall. Design and implementation of a table driven, interactive diagnostic programming system. Comm. ACM, 19(11):609–616, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Webb:1978:CP

[Web78] J. T. Webb. Coral 66 Programming. NCC Publications, Manchester, UK, 1978. ISBN 0-85012-193-0. ix + 110 pp. LCCN QA76.73.C29W4.

Wedel:1975:FTI

[Wed75] Dorothy Wedel. Fortran for the Texas Instruments ASC system. ACM SIGPLAN Notices, 10(3): 119–132, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1964:ISP

[Weg64] P. Wegner, editor. Introduction to System Programming: proceedings of a symposium held at the London School of Economics, July, 1962. Academic Press, New York, NY, USA, 1964. LCCN QA76 .W43 1964.

Wegner:1966:SWD

[Weg66] Louis H. Wegner. The “special weighted distribution problem” of linear programming an application and FORTRAN program. Memorandum RM-4867-PR, Rand Corporation, Santa Monica, CA, USA, 1966. vii + 49 pp.

Weingarten:1965:SCL

[Wei65] Fred W. Weingarten. Simulation of computer logic by Fortran arithmetic. Comm. ACM, 8 (8):516–517, August 1965. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weiss:1966:VPIa

[Wei66a] Ruth A. Weiss. BE VISION, a package of IBM 7090 FORTRAN programs to draw orthographic views of combinations of plane and quadric surfaces. Journal of the ACM, 13(2):194–204, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weiss:1966:VPIb

[Wei66b] Ruth A. Weiss. BE VISION, a package of IBM 7090 FORTRAN programs to draw orthographic views of combinations of plane and quadric surfaces. Journal of the ACM, 13(2):194–204, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weinert:1967:SFC

[Wei67] A. E. Weinert. A SIMSCRIPT-FORTRAN case study. Comm. ACM, 10(12):784–792, December 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weiss:1969:CUF

[Wei69] Eric A. Weiss. Computer usage; 360 Fortran programming. McGraw-Hill/computer usage series. McGraw-Hill, New York, NY, USA, 1969. xix + 232 pp.

Weiss:1973:PPU

[Wei73] Eric A. Weiss. PL/1 para usuarios de Fortran: introduccion a un lenguaje de aplicacion general. Limusa-Wiley, Mexico, DF, Mexico, 1973. 129 pp.

Weinberg:1975:XFI

[Wei75] Lynn A. Weinberg. XLFIT, a FORTRAN IV computer program for the calculation of optimal phase boundaries. Thesis (m.a.), Univ. of Cincinnati, Cincinnati, OH, USA, 1975. vi + 21 pp.

Wells:1970:CTF

[Wel70a] M. Wells. Correspondence: Towards Fortran VI? The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/120.

tif. See [HL70].

Welschinger:1970:IAL

[Wel70b] Ph. Welschinger. Initiation au langage Fortran. Centre de Calcul de l’Esplanade, Strasbourg, France, 1970. 56 + 39 pp.

Werkema:1965:FID

[Wer65] G. J. Werkema. Fortran II data reduction program for gas phase radiolysis studies. Technical Report RFP-624, Dow Chemical Company, Golden, CO, USA, 1965. iv + 22 pp.

Wertz:1972:SSP

[Wer72] H. J. Wertz. SUPER-CODEX: Supervisor plus compiler of differentiable expressions. Technical report, Mathematics Research Center, University of Wisconsin — Madison, Madison, WI, USA, 1972.

West:1969:TFI

[Wes69] Marie West. 940 time-sharing FORTRAN II TAP subroutine manual. ESSA technical memorandum ERLTM-SDL 14, Space Disturbances Laboratory, Boulder, CO, USA, 1969. 41 pp.

Wetherell:1979:APF

[Wet79] Charles Wetherell. Array processing for Fortran. Technical report, Lawrence Livermore Laboratory, Livermore, CA, USA, 1979. various pp.

Wetherell:1980:APF

[Wet80] Charles Wetherell. Array processing for FORTRAN. Technical Report UCID-30175, National Technical Information Service, Springfield, VA, USA, 1980. vi + 64. pp.

Wasserbauer:1975:FPP

[WG75] Charles A. Wasserbauer and Arthur J. Glassman. FORTRAN program for predicting off-design performance of radial-inflow turbines. NASA technical note NASA TN D-8063, National Aeronautics and Space Administration, Washington, DC, USA, 1975. 54 pp. For sale by the National Technical Information Service.

Walker:1973:PPO

[WH73] Justin C. Walker and Charles E.

Hughes. POPSS — a parametric operating system simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):166–169, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Whitten:1968:FIC

[Whi68] E. H. Timothy (Eric Harold Timothy) Whitten. FORTRAN IV CDC 6400 computer program to analyze subsurface fold geometry. Computer contribution 25, State Geological Survey, University of Kansas, Lawrence, KS, USA, 1968. 46 pp.

White:1971:VFC

[Whi71a] E. G. White. A versatile Fortran computer program for the capture–recapture stochastic model of G. M. Jolly. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):443–445, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

                                        cdnsciencepub.com/doi/full/
                                        10.1139/f71-056.

White:1971:CPC

[Whi71b] Ellen Gould Harmon White. A computer program for capturerecapture studies of animal populations: a Fortran listing for the stochastic model of G. M. Jolly. Special publication — Tussock Grasslands and Mountain Lands Institute 8, Lincoln College, Tussock Grasslands and Mountain Lands Institute, Christchurch, New Zealand, 1971. xiii + 33 pp.

Whitney:1972:AAM

[Whi72] V. Kevin M. Whitney. ACM Algorithm 422: Minimal spanning tree [H]. Comm. ACM, 15(4):273–274, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?].

White:1976:LSS

[Whi76] J. W. White. An off-line simulation system for development of real-time FORTRAN programs. International Journal of Computer and Information Sciences, 5(1): 59–79, March 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Widner:1979:CIS

[Wid79] Stephen John Widner. A computer-assisted instruction system for teaching elementary FOR-TRAN programming concepts.

Thesis (m.s.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1979. v + 79 pp.

Wiener:1975:FPR

[Wie75] Howard L. Wiener. A FOR-TRAN program for rapid computations involving the non-central chi-square distribution. NRL memorandum report 3106, Naval Research Laboratory; Springfield, VA, USA, Washington, DC, USA, 1975. 9 + 10 + 37 pp. Distributed by National Technical Information Service.

Willmes:1965:FPR

[Wil65] Henry Willmes. A Fortran program for the reduction of differential cross section data. Technical report P-72, prepared under contract nonr-233 (44) for the Office of Naval Research, Univ. of California, Los Angeles, CA, USA, 1965. iii + 20 pp.

Wilner:1967:AFT

[Wil67] Wayne T. Wilner. ALGOL to FORTRAN translator. Report CGTM 15 (CS239), Stanford University, Stanford, CA, USA, Spring 1967. URL https://

www.slac.stanford.edu/vault/
collvault/greylit/cgtm/CGTM15.
pdf.

Wilf:1969:PDC

[Wil69] Herbert S. Wilf. Programming for a digital computer in the FOR-TRAN language. Addison-Wesley, Reading, MA, USA, 1969. vii + 86 pp.

Wilkinson:1972:FIP

[Wil72a] Robert Hayden Wilkinson. A FORTRAN IV program for the analysis of steady-state time series from gyros and other instruments. Report E-2700, M.I.T. Charles Stark Draper Laboratory, Cambridge, MA, USA, 1972. vii + 81 pp.

Williamson:1972:HLP

[Wil72b] H. Williamson. ACM algorithm 420: Hidden-line plotting program. Comm. ACM, 15(2): 100–103, February 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williamson:1972:HP

[Wil72c] H. Williamson. Hidden-line plotting. Comm. ACM, 15:100–102, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williamson:1972:AAH

[Wil72d] Hugh Williamson. ACM Algorithm 420: Hidden-line plotting program [J6]. Comm. ACM, 15 (2):100–103, February 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [?, ?, ?, ?, ?].

Williamson:1972:AHL

[Wil72e] Hugh Williamson. Algorithm 420: Hidden-line plotting program [J6]. Comm. ACM, 15(2): 100–103, February 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [?, ?, ?, ?, ?].

Willoughby:1973:SAT

[Wil73] Theodore C. Willoughby. Student attitudes toward computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 5(1):145–148, February 1973. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 3rd SIGCSE symposium on Computer science education.

Williams:1974:BRBb

[Wil74] H. P. Williams. Book review:

FORTRAN Programs for Mathematical Programming, by A. Land; S. Powell. Journal of the Royal Statistical Society. Series A (General), 137(4):640–641, ???? 1974. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.
org/stable/2344740.

Wilson:1975:BRB

[Wil75] Eveline Wilson. Book review:

FORTRAN Techniques with Special Reference to Non-Numerical Applications, by A. Colin Day. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):149–150, June 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL

http://www.jstor.org/stable/
2987673.

Wiley:1976:FEa

[Wil76a] G. R. Wiley. Fortran extended.

Techsearch Inc., Adelaide, South Australia, Australia, 1976. ISBN 0-909386-05-6. x + 142 pp.

Wiley:1976:FEb

[Wil76b] G. R. Wiley. Fortran extended. Techsearch, Adelaide, South Australia, Australia, second edition, 1976. ISBN 0-909386-07-2. x + 144 pp.

Williams:1976:SPA

[Wil76c] Paul Lyle Williams. Structured programming as an aid in learning FORTRAN programming useful factors for predicting success in a FORTRAN programming course, and the interaction involved. Thesis (Ph.D.), Iowa State University, Ames, IA, USA, 1976. 165 pp.

Williams:1977:SFP

[Wil77a] Neil Kent Williams. Structured FORTRAN preprocessor: a project. Thesis (m.s.), California State University, Chico, Chico, CA, USA, 1977. vi + 82 pp.

Willmott:1977:WFIa

[Wil77b] Cort J. Willmott. WATBUG: a FORTRAN IV algorithm for calculating the climatic water budget. Publications in climatology 30(2), Laboratory of Climatology, Elmer, NJ, USA, 1977. 55 pp.

Willmott:1977:WFIb

[Wil77c] Cort J. Willmott. WATBUG: a FORTRAN IV algorithm for calculating the climatic water budget. Contribution 23, Water Resources Center, University of Delaware, Newark, DE, USA, 1977. 55 pp.

Willmott:1977:WFIc

[Wil77d] Cort J. Willmott. WATBUG: A FORTRAN IV algorithm for calculating the climatic water budget. C. w. thorntwhaite associates. laboratory of climatology publications in climatology, v.30, no.2 laboratory of climatology (c.w. thornthwaite associates). publications in climatology; v.30, no.2., Laboratory of Climatology, Elmer, NJ, USA, 1977. 55 pp.

Williams:1980:RAF

[Wil80a] J. D. Williams. ROSENET: A Fortran IV program for production of Rose diagrams compatible with Gould or Calcomp plotting facilities. Computers and Geosciences, pages 95–103, 1980. CO-DEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Williams:1980:RFI

[Wil80b] J. D. Williams. Rosenet: a Fortran IV program for production of Rose diagrams compatible with Gould or Calcomp plotting facilities. Computers and Geosciences, pages 95–103, 1980. CO-DEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Winterbon:1974:CFP

[Win74] K. B. Winterbon. CHEBINQ: a FORTRAN program to calculate Chebyshev inequalities and Gaussian integration parameters. Technical report, Chalk River Nuclear Laboratories, Chalk River, Ontario, Canada, 1974. 41 pp.

Winheim:1975:DHU

[Win75] I. A. Winheim. A description of how UNIX Fortran works. Report 1063 (TM 75-8234.3), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 10, 1975. ?? pp.

Winkler:1979:PBP

[Win79] J. F. H. Winkler. Das Prozesskonzept in Betriebssystemen und Programmiersprachen. Interner Bericht 3/79, Institut für Informatik 3, Universität Karlsruhe, Karlsruhe, Germany, 1979. 1–52 pp.

Wishart:1969:FIP

[Wis69] D. (David) Wishart. FORTRAN II programs for 8 methods of cluster analysis (CLUSTAN I). Computer contribution 38, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1969. 112 pp.

Wittman:1974:HUE

[Wit74] Edward Wittman. How to use the extended FORTRAN on the I.T.S. time-sharing computer, 1974.

Witherell:1979:AEMb

[Wit79a] Robert Joseph Witherell. An analysis of the effectiveness of methods and instruction used in the teaching of Fortran IV computer programming. Thesis (ed. d.), Boston University, Boston, MA, USA, 1979. xii + 152 pp.

Witherell:1979:AEMa

[Wit79b] Robert Joseph Witherell. An analysis of the effectiveness of methods of instruction used in the teaching of Fortran IV computer programming. Thesis, Boston University, Boston, MA, USA, 1979. xii + 152 pp.

Witherell:1979:AEMc

[Wit79c] Robert Joseph Witherell. An analysis of the effectiveness of methods of instruction used in the teaching of Fortran IV computer programming. Thesis (ed. d.), Boston University, Boston, MA, USA, 1979. xii + 152 pp.

Witherell:1979:AEMd

[Wit79d] Robert Joseph Witherell. An analysis of the effectiveness of methods of instruction used in the teaching of Fortran IV computer programming. Thesis (ed. d.), Boston University, Boston, MA, USA, 1979. xii + 152 pp.

Wendel:1977:FED

[WK77] Irv. K. Wendel and Richard L.

Kleir. FORTRAN error detection through static analysis. ACM SIGSOFT Software Engineering Notes, 2(3):22–28, April 1977. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wyatt:1976:PEP

[WLO76] W. T. Wyatt Jr., D. W. Lozier, and D. J. Orser. A portable extended precision arithmetic package and library with Fortran precompiler. ACM Transactions on Mathematical Software, 2(3): 209–231, September 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wegner:1960:MFL

[WM60] Peter Wegner and G. Miedel. MI-FLL — FORTRAN to linking loader system post-editor: SAP coded for the 704. Memorandum CC-163, Massachusetts Institute of Technology, Computation Center, Cambridge, MA, USA, 1960. 15 pp.

Wolcott:1977:FIP

[WM77] N. M. Wolcott and F. L. Mc-Crackin. A Fortran IV program to draw enhanced graphic characters. Computer Graphics, 11 (2):121–127, July 1977. CODEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wood:1969:FIC

[WM9 ] Craig A. Wood and E. P. Miles. Fortran IV for the CDC 6400/ 6500/6600 computers. Technical report, Florida State University, Tallahassee, FL, USA, 1969 (??). iv + 192 pp.

Walsh:1972:FIP

[WMM72] John E. (John Edward) Walsh, Robert J. Mandel, and Hans J. Matthias. Fortran IV: programming and problem solving. I. Pitman, Toronto, Ontario, Canada, 1972. [vi] + 246 pp.

Wolf:1968:SFS

[Wol68a] Kurt Bernardo Wolf. A set of Fortran subroutines for handling bases of group representations. Journal of Computational Physics, 2(3):334–335, February 1968. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0021999168900624.

Wolfe:1968:BRB

[Wol68b] Shirley Wolfe. Book review: The Bases of Fortran: A Self-Training Approach to Computer Programming by R. E. Smith. Technometrics, 10(2):412–413, May 1968. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.
org/stable/1267064.

Wolfe:1973:LPF

[Wol73] Carvel S. Wolfe. Linear programming with Fortran. Scott, Foresman and Company, Glenview, IL, USA, 1973. ISBN 0-673-07797-7. vii + 231 pp. LCCN T57.74 .W65 Sci-Eng.

Wolcott:1978:FCP

[Wol78a] Merriam Charles Wolcott. Fortran Computer Programs: Solutions to Optimization Problems Arising in Feedback Control. Lexington Books, Lexington, MA, USA, April 1978. ISBN 0-669-01995-X. ???? pp. LCCN ????

Wolcott:1978:FIE

[Wol78b] Norman M. Wolcott. FORTRAN IV enhanced character graphics. NBS special publication; Computer science and technology 500-32, U.S. Dept. of Commerce, National Bureau of Standards, Washington, DC. USA, 1978. 8 + 24 + 20 pp. For sale by the Supt. of Docs., U.S. Govt. Print. Off.

Wong:1967:FPD

[Won67] P. Wong. A FORTRAN program to determine absorbed organ dose due to inhalation of radionuclides. Technical Report USNRDL-TR-67-7, AD649288, U.S. Naval Radiological Defense Laboratory, San Francisco, CA, USA, January 11, 1967. iv + 106 pp.

Woolley:1977:FAC

[Woo77a] J. D. Woolley. FORTRAN: a comparison of the new proposed language (1976) to the old Standard (1966). ACM SIGPLAN Notices, 12(7):112–125, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woolley:1977:FCN

[Woo77b] John D. Woolley. FORTRAN: a comparison of the new proposed language (1976) to the old standard (1966). ACM SIGPLAN Notices, 12(7):112–125, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wortman:1969:NSF

[Wor69] John D. Wortman. NLPROG: a set of Fortran programs to find the minimum of a constrained function. U.S. Army Ballistic Research Laboratory Memorandum report 1958, Clearinghouse for Federal Scientific and Technical Information, Springfield, VA, USA, 1969. 39 pp.

Worth:1976:NF

[Wor76a] Thomas Worth. Non-technical FORTRAN. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1976. ISBN 0-13-623678-2. 356 pp. LCCN QA76.73.F25W67.

Worth:1976:NTF

[Wor76b] Thomas Worth. Non-Technical Fortran. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1976. ISBN 0-13-623678-2. ???? pp. LCCN ???? US$34.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0136236782.

Warner:1978:DFI

[WPK78] James R. Warner, Margaret A.

Polisher, and Robert N. Kopolow. DIGRAF — a FORTRAN implementation of the proposed GSPC standard. Computer Graphics, 12 (3):301–307, August 1978. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930.

Wray:1970:FIC

[Wra70] William B. Wray. Fortran IV CDC 6400 computer program for constructing isometric diagrams. Computer contribution 44, Kansas Geological Survey, University of Kansas, Lawrence, KS, USA, 1970. 58 pp.

Wright:1966:CFL

[Wri66] Donald L. Wright. A comparison of the FORTRAN language implementation for several computers. Comm. ACM, 9(2):77–79, February 1966. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:1972:VSP

[Wri72] T. Wright. Visible surface plotting. Comm. ACM, 17(??):152–157, ?? 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wrigley:1977:PSMa

[Wri77a] Neil Wrigley. Probability surface mapping: an introduction with examples and Fortran programs, volume 16 of Concepts and techniques in modern geography. Geo Abstracts, University of East Anglia, Norwich NR4 7TJ, UK, 1977. ISBN 0-902246-88-7. ISSN 0306-6142. 72 + [6] pp. LCCN G70.23 .C65 no.16. UK£1.50.

Wrigley:1977:PSMd

[Wri77b] Neil Wrigley. Probability surface mapping: an introduction with examples and Fortran programs. Concepts and techniques in modern geography 16, Geo Abstracts, University of East Anglia, Norwich, UK, 1977. ISBN 0-902246-88-7. 72 + [6] pp.

Wrigley:1977:PSMe

[Wri77c] Neil Wrigley. Probability surface mapping: an introduction with examples and Fortran programs. Concepts and techniques in modern geography 16, Geo Abstracts, University of East Anglia, Norwich, UK, 1977.

Wrigley:1977:PSMf

[Wri77d] Neil Wrigley. Probability surface mapping: an introduction with examples and Fortran programs. Concepts and techniques in modern geography 16, Geo Abstracts, University of East Anglia, Norwich, UK, 1977. ISBN 0-902246-88-7. 72 + [6] pp.

Wenzel:1971:MMC

[WS71] A. Wenzel and Th. Scheidemandel. Multivariate Modelle in der chemischen Industrie. (German) [Multivariate models in chemical industry]. Computing, 8(1–2):40– 61, 1971. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Witmer:1973:IVF

[WS73] Douglas P. Witmer and Stuart Charles Shapiro. The interactive visual FORTRAN interpreter. Technical report 6, Computer Science Department, Indiana University, Bloomington, IN, USA, 1973. 6 pp.

Wu:1973:CPI

[Wu73a] Margaret Schlosser Wu. Computers and programming, an introduction. Appleton-Century-Crofts, New York, NY, USA, 1973. ISBN 0-390-96193-0. viii + 280 pp. LCCN QA76.5.W8.

Wu:1973:BPF

[Wu73b] Nesa L’abbe Wu. Business programming in FORTRAN IV. Wm. C. Brown Publishers, Dubuque, IA, USA, 1973. xii + 305 pp.

Wu:1973:SMA

[Wu73c] Nesa L’abbe Wu. Solutions manual to accompany Business programming in FORTRAN IV. Wm. C. Brown Publishers, Dubuque, IA, USA, 1973. 43 pp.

Wu:1977:BPFa

[Wu77a] Nesa L’abbe Wu. Business programming in FORTRAN IV. Wm. C. Brown Publishers, Dubuque, IA, USA, second edition, 1977. ISBN 0-697-08123-0. xviii + 336 pp. LCCN HF5548.2 .W81 1977. Rev. ed. published as: ANSI FOR-TRAN IV and FORTRAN 77. 3rd ed. c1982. Includes index.

Wu:1977:BPFb

[Wu77b] Nesa L’abbe Wu. Business programming in FORTRAN IV: solutions manual. Wm. C. Brown Publishers, Dubuque, IA, USA, second edition, 1977. 66 pp.

XDS:1970:EFIa

[Xer70a] Xerox Data Systems. Extended FORTRAN IV-H: reference manual for XDS Sigma 5/7 computers. Xerox Data Systems, El Segundo, CA, USA, 1970. v + 97 pp.

XDS:1970:EFIb

[Xer70b] Xerox Data Systems. Extended FORTRAN IV reference manual for XDS Sigma 5/7 computers. Xerox Data Systems, El Segundo, CA, USA, 1970. v + 151 pp.

XDS:1970:XBF

[Xer70c] Xerox Data Systems. Xerox basic FORTRAN and basic FORTRAN IV: language and operations reference manual. Xerox Data Systems, El Segundo, CA, USA, 1970. iv + 61 pp.

Xerox:1971:XEF

[Xer71a] Xerox Corporation. Xerox extended FORTRAN IV: operations reference manual: Sigma 5-9 computers. Xerox Data Systems, El Segundo, CA, USA, 1971. iv + 109 pp.

XDS:1971:XEFa

[Xer71b] Xerox Data Systems. Xerox Extended FORTRAN IV-H: operations reference manual. Xerox Data Systems, El Segundo, CA, USA, 1971. iii + 32 pp.

XDS:1971:XEFb

[Xer71c] Xerox Data Systems. Xerox Extended FORTRAN Library: technical manual. Xerox Data Systems, El Segundo, CA, USA, 1971. iii + 171 pp.

Xerox:1973:XEF

[Xer73] Xerox Corporation. Xerox Extended FORTRAN Library: technical manual: Sigma 5-9 computers. Xerox Corp., El Segundo, CA, USA, 1973. iii + 183 pp.

Xerox:1974:XAF

[Xer74a] Xerox Corporation. Xerox ANS FORTRAN IV: operations reference manual. Xerox, El Segundo, CA, USA, 1974. iii + 42 pp.

Xerox:1974:XEF

[Xer74b] Xerox Corporation. Xerox extended FORTRAN IV: technical manual. Xerox, El Segundo, CA, USA, 1974. 698 pp.

Xerox:1975:XEFa

[Xer75a] Xerox Corporation. Xerox Extended FORTRAN IV: language reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. vi + 178 pp.

Xerox:1975:XEFb

[Xer75b] Xerox Corporation. Xerox Extended FORTRAN IV: operations reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. vi + 133 pp.

Xerox:1975:XFD

[Xer75c] Xerox Corporation. Xerox FOR-TRAN debug package (FDP): reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1975. iv + 65 pp.

Xerox:1976:CAF

[Xer76a] Xerox Corporation. CP-V ANS FORTRAN: language reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1976. vii + 204 pp.

Xerox:1976:XAF

[Xer76b] Xerox Corporation. Xerox ANS FORTRAN: operations reference manual. Honeywell Information Systems, Inc., Waltham, MA, USA, 1976. viii + 135 pp.

Yarbrough:1962:IDO

[Yar62] Lynn D. Yarbrough. Input data organization in FORTRAN. Comm. ACM, 5(10):508–509, October 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yates:1971:SAA

[Yat71a] F. Yates. Statistical algorithms:

Algorithm AS 43: Variable format in Fortran. Applied Statistics, 20 (2):213–216, June 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:

//lib.stat.cmu.edu/apstat/43.

See corrigenda [Yat71b].

Yates:1971:SAC

[Yat71b] F. Yates. Statistical algorithms: Corrigenda: Algorithm AS 43: Variable format in Fortran. Applied Statistics, 20(3): 346, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/43. See

[Yat71a].

Yates:1971:VFF

[Yat71c] F. Yates. Variable format in Fortran. Applied Statistics, 20(2):213– 216, June 1971. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yzerdraat:1969:FPF

[YHE69] W. Yzerdraat, K. Hooper, and B. D. Erdtmann. Fortran programs for faunal analysis. Geological paper 69-3, Carleton University, Dept. of Geology, Ottawa, Ontario, Canada, 1969. 106 pp.

Yohe:1972:AAH

[Yoh72] J. M. Yohe. ACM Algorithm 428: Hu-Tucker minimum redundancy alphabetic coding method [Z]. Comm. ACM, 15(5):360–362, May 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [?].

Yohe:1978:SIA

[Yoh78] J. M. Yohe. A semi-portable interval arithmetic package for Fortran. Freiburger Intervall-Ber. 78/3, Universität Freiburg, Freiburg, Germany, 1978.

Yohe:1979:IAP

[Yoh79a] J. M. Yohe. The interval arithmetic package — multiple precision version. MRC Technical Summary 1908, University of Wisconsin, Madison, January 1979.

Yohe:1979:NAP

[Yoh79b] J. M. Yohe. Nonstandard arithmetic packages for Fortran. MRC Technical Summary 1907, University of Wisconsin, Madison, Madison, WI, USA, 1979.

Yohe:1980:PSI

[Yoh80] J. M. Yohe. Portable software for interval arithmetic. In Alefeld and Grigorieff [AG80], pages 211–229. CODEN COSPDM. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84. In cooperation with R. Albrecht, U. Kulisch, and F. Stummel.

York:1964:AFM

[Yor64] E. J. York. Atlas FORTRAN manual. Technical Report AERE-R. 4599 (pt. 1), Theoretical Physics Division, United Kingdom A.E.A. Research Group, Harwell, Berkshire, UK, 1964. various pp.

Youngman:1976:EF

[You76] Michael Brendon (Michael Brendon) Youngman. Elements of Fortran. School of Education, University of Nottingham, Nottingham, England, 1976. ISBN 0-85359-038-9. [4] + 61 pp. LCCN QA76.73.F25Y68.

Yucel:1980:BPF

[YP80] M. Nadir Yücel and Boğos Pinar.

Bit processing with FORTRAN. ACM SIGPLAN Notices, 15(9): 58–60, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yong:1975:CPA

[YS75] Marian Y. Y. Yong and Robert A.

Skillman. A computer program for analysis of polymodal frequency distributions (ENORM-SEP), FORTRAN IV. Fishery Bulletin, 73(3):681, 1975. CO-DEN FSYBAY. ISSN 0090-0656 (print), 1937-4518 (electronic).

URL https://spo.nmfs.noaa.
gov/content/computer-program-
analysis-polymodal-frequency-
distributions-enormsep-fortran-
iv;   https://spo.nmfs.noaa.
gov/sites/default/files/pdf-
content/fish-bull/yong.pdf.

Yurukoglu:1976:IVS

[Yur76] K. Tanju (Kadir Tanju) Yurukoglu. Iktisat ve sosyal bilimler icin FORTRAN programlama dili. Bogazici Universitesi Yayinlari, Bebek, Istanbul, 1ci baski edition, 1976. 135 pp.

Zaalouk:1969:BFC

[Zaa69] M. Gamal Zaalouk. BURNAPAN — a FORTRAN code for depletion of burnable poisons inside a cyclindrical pin with anisotropic neutron incidence. Kjeller report KR-138 Institutt for atomenergi. Kjeller report KR-138. Institutt for atomenergi, Kjeller Research Establishment, Kjeller, Norway, 1969. 39 +

[14] pp.

Zakrajsek:1977:AGE

[Zak77] E. Zakrajsek. Algorithm 93: Generalised eigenvalue problem. The Computer Journal, 20(1): 86–91, February 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/86.
tif; http://www3.oup.co.uk/
computer_journal/hdb/Volume_
20/Issue_01/tiff/87.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
88.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/89.tif;   http:
//www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/
90.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/91.tif.

Zalewski:1973:SFU

[Zal73] Steve J. Zalewski. Semi-structured FORTRAN using control path expansion: research project. Master of science, plan ii, Dept. of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1973. 26 pp.

Zavisca:1973:SFP

[Zav73] Ernest G. Zavisca. Simplified Fortran programming; an inductive approach. Kendall/Hunt Pub., Dubuque, IA, USA, 1973. xi + 203 pp.

Zaider:1978:RFP

[ZD78] M. Zaider and J. F. Dicello.

RBEOER: a Fortran program for the computation of RBEs, OERs, survival ratios, and the effects of fractionation using the theory of dual radiation action. Informal report LA 7196-MS, Dept. of Energy, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, May 1978. 19 pp. For sale by the National Technical Information Service.

Zilinskas:1978:SAA

[Žil78] A. Žilinskas. Statistical algorithms: Algorithm AS 133: Optimization of one-dimensional multimodal functions. Applied Statistics, 27(3):367–375, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.
cmu.edu/apstat/133.

Zimmerman:1969:EFA

[Zim69] Carl Duane Zimmerman. Extended FORTRAN algebraic manipulator with applications to linear problems of physics. Thesis (Ph.D.), University of Minnesota, Minneapolis, MN, USA, 1969. various pp.

Zinsmeister:1979:FTS

[Zin79] G. E. Zinsmeister. Fortran on Time-Sharing. Engineering Press, San Jose, CA, USA, June 1979. ISBN 0-910554-27-7. ???? pp. LCCN ???? US$9.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0910554277.

Zlatev:1979:LAF

[ZN79a] Z. Zlatev and H. Nielsen. LLSS01 — A Fortran subroutine for solving least squares problems (User’s guide). Technical Report 79-07, Institute of Numerical Analysis, Technical University of Denmark, Lyngby, Denmark, 1979. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from math.liu.se

in pub/references.

Zlatev:1979:LFS

[ZN79b] Z. Zlatev and H. Nielsen. LLSS01 — A Fortran subroutine for solving least squares problems (User’s guide). Technical Report 79-07, Institute of Numerical Analysis, Technical University of Denmark, Lyngby, Denmark, 1979. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from math.liu.se

in pub/references.

Zohni:1972:FPC

[Zoh72] O. Zohni. A FORTRAN program for the computation of the generalized Talmi coefficients. Computer Physics Communications, 3(1): 61–68, January/February 1972.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/article/
pii/0010465572900550.

Zohar:1980:CSS

[Zoh80] S. Zohar. Correction to “Fortran subroutines for the solution of Toeplitz sets of linear equations”. IEEE Trans. Acoustics, Speech, and Signal Processing, 28 (5):601, 1980. CODEN IETABA. ISSN 0096-3518.

Zorn:1968:SFI

[Zor68] William Ernest Zorn. STREDO:

a FORTRAN IV package for student cumulative guidance records for kindergarten through junior college. Thesis (m.s.), University of Southwestern Louisiana, Lafayette, LA, USA, 1968. 104 pp.

Zinkl:1980:FCP

[ZSD80] Richard J. Zinkl, Don L. Shettel, and Ralph F. D’Andrea. FOR-TRAN computer programs to process Savannah River Laboratory hydrogeochemical and streamsediment reconnaissance data. Technical Report GJBX 248(80), U.S. Dept of Energy, Grand Junction, CO, USA, 1980. 96 pp.

Zalotai:1978:FPF

[ZSF78] Lajos Zalotai, Laszlo Seres, and Pal Fejes. FORTRAN programok fizikai kemiai feladatok megoldasara. Muszaki Kiado, Budapest, Hungary, 1978. ISBN 963-10-2161-0. 440 pp. LCCN QD455.3.E4Z34.

Zinsmeister:1976:FTC

[ZSW76] G. E. Zinsmeister, Fred D. Stockton, and Lee A. Webster. FOR-TRAN on time-sharing: (CDC Cyber edition). ????, Northampton, MA, USA (??), 1976. 137 pp.

Zinsmeister:1977:FTC

[ZSW77] G. E. Zinsmeister, Fred D. Stockton, and Lee A. Webster. FOR-TRAN on time-sharing: CDC CY-BER edition. ????, Northampton, MA, USA, 1977. 136 pp.

Zinsmeister:1979:FT

[ZSW79] G. E. Zinsmeister, Fred D. Stockton, and Lee A. Webster. FOR-TRAN on time-sharing. Engineering Press, San Jose, CA, USA, 1979. ISBN 0-910554-27-7. iii + 103 pp. LCCN QA76.73.F25 Z56. US$4.95.

Zlatev:1976:SFI

[ZT76] Zahari Zlatev and Per Grove Thomsen. ST — a Fortran IV subroutine: for the solution of large systems of linear algebraic equations with real coefficients by use of sparse technique. Technical Report 76-05, Numerisk Institut, Danmarks Tekniske Højskole, Copenhagen, Denmark, 1976. 65 pp.

Zwakenberg:1975:VEL

[Zwa75] R. G. Zwakenberg. Vector extensions to LRLTRAN. ACM SIGPLAN Notices, 10(3):77–86, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zwass:1980:PFS

[Zwa80] Vladimir Zwass. Programming in Fortran: Structured Programming With Fortran IV and Fortran 77. Barnes & Noble, December 1980. ISBN 0-06-460194-3. ???? pp. LCCN ???? US$8.95. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0064601943.